A Complete Bibliography of Publications in Concurrency and Computation: Practice and

Experience: 2000–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

27 September 2023 Version 2.01

Title word cross-reference

(α, k) [LXW17]. (P U)^{2}M^{2} [CSL^{+}18]. 0 [SSIH19]. 1 [SSIH19, VDL^{+}15]. 16 × 16 [TPGC15]. 2 [CCW06, PDC16]. 2pq [CL14]. 3 [ACIC^{+}13, Boc19, CSMK19, CPE^{+}19, DCG11, EMEY14, KSM^{+}08a, MBP16, MCY^{+}07, MJL01, OLG^{+}15, PSLC11, PSCK^{+}15, QSZL18, RWK17, TTR^{+}10, WSZ^{+}18, YBC^{+}07, ZLKK17]. 5 [AVS^{+}19]. ^{2} [YNX^{+}16]. ^{T M} [YL01, ZJKL10]. n [LSP15]. _{T P} [LTK17]. c [HW16]. ℓ [DHV03]. G(d) [WCA08]. K [LFZ07, SS19c, DHV03, GR13, KH12, PGL^{+}17, TLX^{+}17, XLL^{+}18]. ↔ [GRW^{+}19]. LU [DFLL14]. N [BDH15, CGK14, GGV14, SSB^{+}14, TL14, AS15, LLRS19, PGL^{+}17, PDCA17]. s [PGL^{+}17]. t [HJM^{+}11]. t/k [XYL18]. x [IR11].

-anonymity [LXW17]. -ary [PGL^{+}17]. -body [CGK14, GGV14, SSB^{+}14, TL14]. -Cube [AS15]. -D [ZLKK17].

-dimensional [AVS^{+}19]. -direct [PGL^{+}17]. -gram [PDCA17]. -indirect [PGL^{+}17]. -level [ZLKK17]. -mean [HW16]. -means [TLX^{+}17, SS19c]. -mer [GR13]. -nearest [KH12]. -of- [LLRS19]. -out-of- [DHV03]. -private [HJM^{+}11]. -thread [BDH15]. -wait-freedom [IR11].

.NET [BHW05, HLB10, LHB^{+}19].

/OpenMP [VDL^{+}15].

1 [RMP^{+}13a]. 1.1 [OA02]. 1.2 [CG01]. 1.3 [MP04]. 10th [Kni06, WT15]. 128-processor [LL01]. 12th [Fox17a]. 1394 [HON04]. 14th [GJ17]. 1516 [MP04]. 15th [GZX17]. 1605 [Ano06]. 17 [Ano06]. 18th [PCC17]. 1940 [DKMM14].

2 [BS04, BB13, BÇG14, JLT06, LXW^{+}16, LSK04]. 2.0 [CBHTE11, DWC09, DH15, FP09, LVN^{+}12, MWL^{+}15, PFC^{+}09, ZL09, Zic12]. 2.0-based [MWL^{+}15]. 2.2 [HRR^{+}11]. 2000 [LL01, LSK04, PIH04, Wis02]. 2001 [ZGRSC10]. 2003 [ZMJZ10]. 2005 [WC08]. 2007 [BL09a]. 2008 [BL09b, NCD^{+}08]. 2009 [BL11a, SHT11]. 2010 [BL11b, Bou13, HTBR12]. 2011 [BL13b]. 2012 [BL13a, HTW14, Hou12, QFG14]. 2013 [AF14, LBW14, PDD14, WDGK15, WT15]. 2014 [CČJ^{+}16, FB16, HTBR16, LBS15]. 2015 [LBT16]. 2016 [Che19]. 2017 [Du18c, OKG18, SG19]. 2018 [Wu18]. 21st [BHJ^{+}16]. 21st-century [BHJ^{+}16]. 2D [ZZZ^{+}15]. 2D-DWT [ZZZ^{+}15]. 2nd [FZ08].

3.0 [DBB^{+}16]. 30.7 [SLM^{+}10]. 369 [GKS09]. 3D [SL14, Che18, LXW^{+}16]. 3G [KCS07]. 3rd [CC09].

4.0 [JCP15]. 40 [DAC^{+}18].

5G [GLL16].

6 [OCC^{+}05]. 6.1 [ZMJZ10]. 6.2 [ZMJZ10]. 600 [LSK04]. 6LoWPAN [BNNH19]. 6th [Run10].

77 [AL04].

8 [KS19c, SAdB^{+}16]. 802.11s [BOB13]. 802.21 [WCLH12]. 8th [LB19].

90 [FSPC^{+}02]. 90/HPF [FSPC^{+}02]. 95 [vWAH^{+}02]. ’99 [TM01]. 9th [WCWB19].

A-LDA [GLD17]. A2 [FNBS16]. AAA [BT18, MML^{+}17]. AAA-based [MML^{+}17]. AAAA [WBB^{+}07]. AAH [GQR16]. ABC [BPL12]. ABC-GA [BPL12]. abnormal [GBXL17, SRR19]. abnormity [LFW19]. abort [ASP19]. ABS [SAdB^{+}16]. absorption [WWY19]. abstract [AHM06, CTY15, DWC09]. abstraction [IAH^{+}15, JMF09, LFG05, OKW18, SGM18, WP12]. abstractions [VS02, VFAD18]. Accelerate [YXLZ16, FBV^{+}13, MTHK14, TSL^{+}19]. accelerated [ANPR16, BDW14, CGK^{+}16, CMMB13, CP14, DCD^{+}14, IOOH12, JLH^{+}16, LZL^{+}17b, LS15, MCB14, MPSGD14, NSN^{+}17, PGdCJ^{+}18, QSX^{+}17, RK15, SBC15, TDM^{+}15, TPT^{+}18, ZWL^{+}17, ADK^{+}16]. Accelerating [AdCPdSD17, BKLH09, DFTHD18, DMC^{+}18, DCK12, EDB^{+}14, GC19, KW18, KHF^{+}17, LL16a, MKO^{+}17b, NNH^{+}14, PRCV16, PSV19, RCA^{+}11, SJISVR17, SNK^{+}15, TB12, WKB^{+}19, ZCD^{+}12, BP17, CCO15a, ISO^{+}14, PPdSTB17, SAD13, SIOS02]. Acceleration [ZO14, ABG^{+}13, BYDC19, KC13, KPNS18, PZ11, ZAB^{+}19]. Accelerator [PSS^{+}19]. accelerators [ADF^{+}13, BKSM^{+}15, BHKW12, CGST17, HJB12, HCKF15, SRF13, YOBS16]. accelerometers [ZZL^{+}18]. acceptance [ALL^{+}15, HLA^{+}18]. Access [AK01, RCB^{+}04, SW11, AFGL09, ATKH^{+}17, AC02, AV07, AAF^{+}07, BDI^{+}07, BHA^{+}15b, CSL08, CJC^{+}18, CLH^{+}16, DFC12, DKMV07, GvHKK11, GBSHA01, JSR^{+}19a, JSG17, KKDS19, KFS^{+}06, LZW13, LLLS18, LCMY13, MLL^{+}11, MTGZ17, MCB14, MD02, OTG^{+}07, RR01, Sch04, SKNH09, SS07, SW12, SCLK15, TYHL12, sTzNL16, WLW11, WYBS16, WFKS18, XHH12, YBO10, ZYN^{+}07, ZXW16b, ZZL^{+}18]. Access-controlled [RCB^{+}04]. accesses [LPC^{+}14, NPL19]. accessible [OK15]. Accessing [GKP^{+}09, Wit10]. account [RSPV17, VS19]. Accounting [GEJ^{+}08, HGT14, MAS^{+}14, SAC^{+}07]. accounts [WBB^{+}07]. accumulator [LZY^{+}16]. accumulator-based [LZY^{+}16]. accuracy [DFLL14, EMEY14, QLZX19, TLM17]. Accurate [BCK^{+}09, GW15, SM19a, WLF19, Xu19, AAF^{+}07, FOTW04, GQR16, HTWW19, KD19, MCC16, TCP^{+}05, VB16]. accurately [VŠC17]. ACES [Run10]. achieve [CAG^{+}13, PQP13, YLLZ09]. Achieving [CLW^{+}19, CBPP02, DPP03, DFLL14, SSZ13, WLW11, ZYN^{+}07, XTLG08]. ACID [CEMR19]. ACM [Fox01, Fox05]. acoustic [MS07, OFR^{+}17]. ACPI [XRD^{+}17]. ACPI-compliant [XRD^{+}17]. Acquisition [SZD19, CMCAA17, CL19]. across [AAE^{+}09, BPdM06, CC15, XLQL18]. action [KS19a, WGQ^{+}18]. Active [PLL14, RM11, ZHT08, ZTGW17, BDP18, DT17, Pun01, SM04, XM02, ZL06]. ActiveSpaces [DZJ^{+}15]. activities [SR19c, WWL^{+}17b]. activity [BFH17, BDMM^{+}05, GQR16, Yu18]. actor [BAT13, CWM18]. actuator [ZCXH17]. acyclic [AS17]. ad [CNPP09, Den07, DK19a, DA15, EB10, HKA^{+}15, IHB15, KOO12, KKK10, KABD07, MTM19, MSMA19, MLRR09, PY19, QWW^{+}16, Sha15, SK17, WXSH19, YWM^{+}10]. ad-hoc [Den07, DK19a, IHB15, KABD07, MTM19, PY19]. ADAGE [YR15]. adaptability [DT15a, SPSNvS07, ZBZH11]. Adaptable [CZL^{+}17, dRL10, ML19a, PGO^{+}04]. Adaptation [LLH^{+}09, RCR^{+}15, AAHRW04, GFBR10, LW05, MvNK^{+}06, RWK17, WO02, WRC09, WFHT17, XLT^{+}17, Zhu18]. Adapting [LBTE14, ZYH09, KL02]. Adaptive [AWR17, ADF^{+}19, dOCPFJ13, CZG16, EME19, FNI17, GVC10, GLS^{+}19, KR06, LPS^{+}09, LGQ^{+}17, LCW^{+}17, LFW19, LCMY13, PGK11, PCD15, RHRB13, SZ18, TCBR^{+}10, WML^{+}19, AR16, BJ01, BCD^{+}02, BB12, BVIB19, BM08, BFVRC15, CEMR19, CRB09, CLL^{+}19, CCW^{+}15, DDP^{+}06, GvDHS12, HKRR08, HHKA14, JN03, JyLdZ^{+}18, KW19b, KFD18, LLMM19, LSJ16, LC18, LB11, MB18, MV16, Nak02, NC05, RHBK11, SGD^{+}18, SYMA17, SWD^{+}17, SG18, VB16, WCR^{+}14, WLW14, WLL03a, XYSW18, YCL11, YLD13, YYLC19, YESG^{+}17b, YESG^{+}17a, YESG^{+}19, ZH15, ZCXH17, dOOO^{+}12]. adaptively [LPSF11, PWMX16]. adaptivity [VD05]. ADD [PSS^{+}19]. adder [NK19, SS19a]. Adding [SRN^{+}15, vRS05]. address [ADK^{+}16, CXW17, HKS^{+}12, WHX19, ZDB^{+}14]. Addressing [CBBCD08]. Adelson [BBB16]. Adelson-Velskii [BBB16]. ADIOS [LLT^{+}14]. Adjusting [YYCH10, JKZ03, YYC10]. adjustment [SWS^{+}18, ZWL^{+}19]. adjustment-based [SWS^{+}18]. administrative [Kri19]. admission [DMA13, XCL09, ZCC^{+}06]. Adolescent [CS09]. adoption [HLX^{+}16, HSV^{+}19]. advance [ET09]. Advanced [FR02, Fox12, LBFS17, MCAB^{+}02, NC05, SRdS09, SF10, AP06, DN19, FSPC^{+}02, Fer15, LAC^{+}08, LL13, QLL10, SSU18, SE01, Xha18]. Advancements [MRJ^{+}14]. Advances [Ano15a, BCX15, BLXE19, DSH18, MLY10, NPM19, PIGK16, SFN12, XCHK14, XHZ12, XCHY13, Zha08, BBM19, DM15, DL17, LMH^{+}14, LL15, LNBL17, MPSGD14, Not16a, PZO19, QD17, RS13, SPSB19, SRTG^{+}07, WS17, WS19b, YMLR16]. Advancing [KMJ14]. advection [JPS17]. Adversaries [LDZ^{+}19]. advertisement [XY17]. aerial [ABK^{+}18, LZH^{+}15]. AES [CLH^{+}11, FLYL16, PZ11]. AEZ [ASB^{+}19]. AF [KT19a]. affected [KM19]. affecting [HMM^{+}09]. affinity [KB13, LXK^{+}19]. affinity-aware [LXK^{+}19]. against [AYSZ14, BSB^{+}03, CLW^{+}19, HYLG15, KKS12, LWY^{+}17, LLRS19]. age [ABDO09, HT15, ZS19]. Agent [KKS12, RK01, CCCW13, CAC15, CKC09, CGN15, CN16, DMRS15, DPST06, EFA^{+}17, GPVCdBRO12, HM16, JLHH14, KZY^{+}18, KFD18, KN19, LCT16, MT08, OM06a, PGK11, PRS01, RHL^{+}18, TL19]. Agent-based [KKS12, CAC15, CKC09, KZY^{+}18, KFD18, LCT16, MT08]. agent-oriented [DMRS15]. agents [ADSV16, AKW04, CJ12, DK19b, EJF^{+}16, FRdOR^{+}19, NPTT06, QKSJ07, SZMF19, SNEP14, YCW08, YG19, ZMZD11]. agents-based [ADSV16]. aggregate [FGC06]. aggregation [DXWD16, DK19a, DZC16, LWSZ19, NJ15, TLWZ14, WJXZ18, ZGX11]. aggressive [VJHB05]. agnostic [VKM^{+}09]. agreement [KDW^{+}17, PY19, XXCY19, ZZC15]. agreements [DPGA11, YS07]. agricultural [AAV^{+}15, CHX^{+}19, HAJL16]. ahead [FBS16]. AHM [WAS07, WC08]. AI [ZS19]. aided [BAZ09, CL19, KBDA19, LGdVH13, SM19c, WLFX17]. Aimed [CZ15a]. air [ZL19, ZDL19]. Airavata [BAC^{+}15, PMG^{+}15]. aircraft [ZQMC19]. airport [BDZ19]. AL [CZQ17]. AL-DDCNN [CZQ17]. Alamos [WJLD09]. alarm [WML^{+}19]. Alchemist [GRW^{+}19]. alert [RCM12, WZXZ12]. alerting [VZB19]. algebra [ADI^{+}14, BHL^{+}09, BHJ^{+}16, BLL18, CGGH17, HLYD12, KLDB10, PHCR09, SD15, SLB08, SB17]. Algebraic [BFK^{+}17, BDE^{+}19, ODS^{+}13]. algebras [CMD11]. algebric [MQOQOH01]. algorithm [AA16, AS17, ACGG06, ACCM17, BYDC19, Bao19, BKH08, BY12, BKND16, BDF15, BT04, BDH16, CHP17, CSMS^{+}19, CMMB13, CCW^{+}15, CEM^{+}17, CDF^{+}17, DCJ12, DHV03, DS19, DCCZ18, DPS07, DLM13, Du18b, Du18a, EN19, EB18, EN16, FH13, FTT15, GZG^{+}16, GCWE15, GRQ19a, GLZ19, GM04, Guo19, GAS19, Has17, HZHP09, HSL19, HBD18, HdV13, HW16, JC07, JLH^{+}16, JLL18, JSR19b, JKV^{+}15, KBDA19, KA16, KH12, KHHC13, KKW^{+}14, KYBV17, KT19c, LLMM19, LZZ^{+}17, Li04, LC09, LDZ14b, LH17, LXYC17, LHHJ18, LWZ^{+}19, LWSZ19, LLT19b, LLH^{+}15, LYC16, LZBF17, LC18, LMX^{+}18, LHZX19, LWP19, LZZ^{+}15, LLQL14, MS17c, MT19b, MKKB04, MCCD18, MLRR09, MLYL17, MDL^{+}10, MNL15, NIIU17, OAS^{+}15, PLY13, PN19, PWH18, PCF^{+}17, PCT04, PSCK^{+}15, PV15, PZ17, QSZL18, QLZX19, QW17, RP19, RNJM17, Ree01, RSM01, RVVPD^{+}17, RIWS17, SR19a, SAB15]. algorithm [SJVR15, SAD13, SBC15, SS15a, SMM18, SIS19, SZ18, SR17, SdSL18, SG18, SZD19, SM19c, TZK16, TNIB17, TJ17b, TJ17a, TYL^{+}15, TLX^{+}17, TJD^{+}17, TRW07, Viv03, WBD^{+}19, WLLL15, WBZ10, WJ12, WZJD13, WCR^{+}14, WZS^{+}15, WJYH16, Wan18a, WF18, WZLL18, Wan18b, WLF19, cWsThY19, WW19, WZLQ16, WLL03a, WLL03b, WRDZ13, WZYG19, XYL18, XLG19, XMJ17, YGW17, YWL^{+}17a, YWLQ18, Yan19a, YYZH19, YYC^{+}19, YT19, YZXW17, YZ19, YL19, ZW09, ZY12, ZQK15, ZYW^{+}16, ZXW19, ZGZ19, ZL19, ZXC^{+}19, ZWXS19, ZLW^{+}19, ZZZX19, ZYLT06, ZFJ16, ZQW^{+}17, ZCH^{+}18, ZYLY18, ZFXJ19, ZWL^{+}17, ZLC17b, ZDX12, dARP17, dCRS11, sCR19, KAA19]. Algorithmic [SKK01, WS17, WS19b, BGV^{+}01, Cho01]. Algorithms [BJ18, CGBNM17, Fox10, GW19, SNM15, AJY^{+}15, ABDO09, ALMT19, AMVOSGAC17, ATİ17, AMS17, ABDR13, AT18, BYN^{+}17, BDL^{+}15, BEDK18, BB12, BCM15, BDTdS13, BDH18, CDA09, CMVRRVGI17, CCTW11, CCP^{+}15, DS04, DLT^{+}16, EOD^{+}19, EAGVBVDS11, FLYL16, FLMRC02, FRKS12, GYM14, GLM^{+}16, HLYD12, HT15, HR06, IQOvdG13, JdM12, JM07, JKZ03, KRW17, KHZN06, KR04, KR11, LLMK18, LF17, LWW06, LLH^{+}17, LXK^{+}19, LB11, LLY19c, MB17, MSD^{+}18, MHLC^{+}05, MTK16, MJL01, MB14, OO18, PSRR14, PP17, RTMZ13, RB17, SRS16, SGM18, SER15, SFH13, SFT15, SSU18, TMZ07, VS02, XTB17, YCW08, ZS17, ZL12, ZLZ^{+}17]. algorithms-by-blocks [IQOvdG13]. aligning [SL14]. alignment [AMHC11, BS04, LLB04, LS15, SRF13, TDL^{+}18, ZCL^{+}18]. alkali [WYY^{+}19]. all-pairs [SSB^{+}14]. all-to-all [HR18, JKD19, ZJKL10]. allies [LDZ^{+}19]. allocating [ME08, PBK19]. Allocation [HJTX17, BHD13, CA06, CCSS10, CLL^{+}19, CFTT17, DFPT06, EdPG^{+}10, FXX16, GEJ^{+}08, GS04a, ITK09, JL10, Jon09, KZY15, KZY^{+}18, LBV16, LC09, LDPZ14, LYF^{+}17, MS17a, PLC^{+}19, RPK08, Sha15, SS18, SN18, SKJ17, SdSL18, TAMC19, TAB^{+}06, TXZ^{+}17, TSBR10, VDB09, VGN^{+}16, WRLS12, WHXB19, XLZD13, YYC10, YCL11, YLC11, YYC^{+}19, YPLJ11, YL01, Zen19a, ZJL13, ZLH^{+}15, ZWMT12, vdKEL10]. allocations [SB17]. almost [BK05]. almost-whole [BK05]. alternating [ZF18]. alternative [BFU07, ELM^{+}16, Kuł14]. alternatives [AM01]. Alting [WBM^{+}10]. altitude [ABK^{+}18]. Alto [DC19a]. ALU [NK19]. Amadeus [BPB08]. Amazon [MSL^{+}14, MDH^{+}16, PRNM19]. AMBA [MS10]. Ambient [dMd^{+}17]. AMC [CCW^{+}15]. American [GG07, DCJ12, HLCW15, PW12, TZKH12]. among [BFU07, MÖO17]. amplifier [SM19c]. AMSBA [PIGK16]. AMULET1 [The01]. AMUSE [LDS^{+}08]. analyses [BAD^{+}11, DZ13, SMBT07, TCH^{+}13]. Analysing [LLT09, Sch04]. Analysis [AM07, CLZX10, CLW^{+}15, DXWC16, GHMX13, GG07, GGR^{+}10, HLYD12, KNT^{+}01, LMGZ19, MFG^{+}13, PQP13, PB19b, ULS03, YYXL19, Yan19a, AA16, ABF^{+}10, AHB^{+}10, Aia15, AAQAR^{+}17, AAF^{+}07, ANCA19, AC06, AHH14, BRK^{+}17, Bai17, BGGS14, BMV03, BBCG02, BBdS^{+}17, BKH08, BRWB06, BDP^{+}14, BLSP11, BWEB14, BDMM^{+}05, CM05, CC13, CSBL12, CGIP16, CXPL15, CSMK19, CM18, CMD17, DCG11, DIM18, DDX^{+}06, DXHL17, cDrLyC^{+}19, DL17, DLW19, DPP^{+}19, EMS15, FBH^{+}01, FMM08, FBC10, FBV^{+}17, FGG^{+}18, GHLS19, GYB^{+}11, Ger05, GMT07, GO10, GLZ19, GPW03, GSWJ19, GYP^{+}16, GC18, Guo19, HJB12, HGT14, HJH19, HFR^{+}17, HPS12, ISS^{+}02, IAH^{+}15, JLHH14, JZMD19, KvGS^{+}14, KR15, KHM^{+}11b, KYBV17, KCZ^{+}05, KSB^{+}19, KSM^{+}19, KAMB19, LLRS03, LAC^{+}08, LWG^{+}15, LZW^{+}17a, LH17, LGJ17, LLLyL16, LL16c, LQL^{+}09, LJML10, LFHT15, LSW07, LPC^{+}14, LGD15, MRMC15, MSL^{+}14, ASB^{+}19]. analysis [MTT15, MO02a, MS10, MSPDP19, MÍM19, MP17, MJD15, MJD17, MRP^{+}18, MWHW16, MDV07, MWW10, MWLS11, NLYZ12, OAS^{+}15, OFR^{+}17, PWC^{+}14, PFC14, PVR^{+}09, PLL17, PSV19, PSS^{+}18, PPP10, PRNM19, Puf13, QZH16, QZYZ16, RVD^{+}12, RVRD10, dRRdCRR16, RVVPD^{+}17, RS07, RGB^{+}15, SSE19, SGJ^{+}17, SAOKM04, SLV12, SER15, SGCA^{+}16, SLGL16, SIS19, SWD^{+}15, SM09, SWB12, SWL^{+}01, SGR19, SB19b, SWD^{+}17, TNH15, TNI16, TQL^{+}14, THQ19, TWN07, TWB13, TF03, WCA08, WZC16, WBC^{+}02, WWG^{+}11, WMDM07, WKL^{+}11, WCH^{+}07, XWFH08, XTZ10, XYER16, XZZ^{+}16a, XBB13, XZZ16b, YWL^{+}17b, YC19b, YHH13, YL19, Zen19b, ZCC^{+}06, ZPG10, ZH16, Zhe16, Zhu19, CKOG10, MCSML07]. analysis-driven [HPS12]. analytic [TPV17]. analytical [CS13, ID18, JAA08, LC17, PRS01, RGAK15, TYHL12, WTN07, ZHM^{+}17]. analytics [BM16, BWHS18, DSH18, DWG19, GQH17, KMJ14, Li18, LL19c, MYS19, PRCV16, SMS^{+}19, TTPJ16, WYZ^{+}17, YCQF19, YLH^{+}19]. analyze [FCY17, HLZD18, HWZX08]. Analyzer [CV07]. Analyzing [DT15a, HKG08, IÁE11, MÍM19, PB19a, SdOVM16, UGM18, LZL17a, RR15, XZHW09]. and/or [ECP18]. Android [AA16, CJC^{+}18, CL16, KGP^{+}19, KGK17, MKO^{+}17b, PCL17, QXJS17, YWL^{+}17b]. animal [CMT13]. anisotropic [Du18a]. annealing [DS19, HXY^{+}12, MK15b, WYZ12, WZLL18]. annotate [FHO^{+}15]. annotation [CHH18, MHK^{+}18, SRL^{+}14, WOH^{+}13]. annotations [IS10, MGS19, vRGNP09]. anomalies [SLV12]. anomalous [ZLZ^{+}17]. anomaly [AWR17, BMPP17, CRB^{+}17, HPD^{+}15, KMBR19, LZL17a, MW18, PFC14, PVCS18, RS16, WW19]. anonymity [LXW17, WWS^{+}12, YZXW17]. anonymization [SWZ12]. anonymized [DOJ^{+}19]. anonymous [BT18, MML16, SM19b, XXCY19, YZW^{+}15, YYK^{+}19]. Answering [GR13, TGS14, CZWH07, HHWZ08]. ant [MS18, TV14, YZ19]. Anti [LHZS19]. Anti-noise [LHZS19]. Anticipative [YYCH10]. ants [PCS^{+}12]. anycast [SPSNvS07]. AODV [KKK10]. AODV-RIP [KKK10]. AOI [RGCC15]. AOI-cast [RGCC15]. Apache [GC19, GRW^{+}19, PMG^{+}15, WZLL18]. APART [GG07]. APCIE [CČJ^{+}16]. APEC [Ano02]. APEX [SS07]. APEX-Map [SS07]. APF [Li19]. API [APHB16, YWL^{+}17b]. APIs [CS15]. app [Fer15]. Appearance [TNH15, TNI16]. appliance [JK10]. appliances [LLL15]. Application [BBA18, Che18, DK19b, GLZ19, KJ19a, MT19a, OKO18, PHGK10, RVD^{+}12, RNAD19, RO12a, RO12b, SWD^{+}15, TDM^{+}02, AV07, ARPPM17, BHD13, BvIF10, BT18, BAZ18, CRC^{+}15b, CCO15a, dOCPFJ13, CNAQ18, CZ15b, CM07a, CKBB14, DZL^{+}19, DZL^{+}17a, DL17, ESG11, FCY17, FGG^{+}18, FJG^{+}13, GA08, GSV03, GAE^{+}06, GWVP^{+}14, HLX^{+}16, HVM^{+}15, HKAC14, HYLG15, HK02, IHB15, JZJW15, JN03, KOK14, KSM^{+}08a, KS19c, KA11, LLD19, Li19, LHBW15, LXK^{+}19, MKB01, MAK18, MvWL^{+}10, NZKK11, OTO18, PWC^{+}14, PSG03, PLW^{+}18, QZH16, RNJM17, RMCHMG15, SBBE07, SLD^{+}12, SM03, SBDP15, SİM^{+}07, SVN12, TSL^{+}19, TKA^{+}02, TY15, TK10, TBK^{+}15, VSR^{+}09, VSB^{+}15, WXY10, WSZ^{+}18, XHH12, XS19, XM02, YDS^{+}14, YBC^{+}07, YZ19, ZS01, ZDA^{+}07, ZYL10, ZSS18, ZKJ^{+}07, ZZZ^{+}15, ZLC17b, dRL10, vAVS12, RTPPH12]. application-aware [DZL^{+}17a]. Application-driven [RVD^{+}12]. Application-level [BBA18]. application-network [LLD19]. application-runtime [GA08]. Application-specific [RO12a, RO12b, ZS01, ZLC17b]. Applications [AI17, BL17, CL08, CC09, EN09, EH18, Fed13, LWL17, LB19, PPST09, PC17b, STS19, SNM15, TM01, Tur04, Wu18, YWT^{+}12, ACJ10, ALKD16, ABtGT^{+}12, AA19, AMSS15, AK01, ASS^{+}05, AMGCC17, ALMT19, ACFT15, AC06, Ang08, ACG15, AFG16, Ano06, AAV^{+}15, AAE^{+}09, BH16, BMV03, BFR05, BCD^{+}02, BEQOR13, BBdS^{+}17, BBK11, BSP11, BR04, BFM^{+}06, BFVRC15, BAG17, BPD06, BSB^{+}03, CML^{+}10, CEH^{+}06, CGK^{+}16, CGST17, CV07, CGBNM17, CDMS15, CSBL12, CGIP16, CSL12, CWMZ06, CA06, CKC09, CGI17, CN02, CSPM13, CM18, CSWB11, DJM12, DFPT06, DHH^{+}13, DKMV07, DvNM^{+}11a, ET15, EPB14, EMS11, EDSV09, EFY17, EDBS08, EABVGV14, EMS15, EJF^{+}16, FBH^{+}01, FE17, FT06, FNBS16, FO18, FS18, GFBR10, GLS^{+}19, GTA10, GWC^{+}11, HJV^{+}19, HKS19, HFDJ10, HKS^{+}12, HLHC12, ISO^{+}14, JOC^{+}15, JCK^{+}13, JKL^{+}17]. applications [JZZL06, JK06, KMBR19, KTR11, KQR^{+}17, KKM^{+}06, LBOE18, LBTE14, Lan17, LHL10, LL05, LPH09, LLWS09, LDPZ14, LHHJ18, LMX^{+}18, LLL16, MWL^{+}13, MHJH16, MMMP01, ML19a, MKIO04, MLC04, MBC^{+}14, MSCS18, MMS07, MCG18, MSS16, MDH^{+}16, MKO^{+}17b, MCSML07, MK15b, MT09, NSBR07, NDT^{+}16, Not16a, OSK^{+}01, OK18, Par02, PWWR05, PS05, PTL^{+}16, PFC^{+}09, PRV11, PSS^{+}18, PK17, PH12, PB16, PIGK16, PSIP16, QCB17, RRBB11, RBP12, RMCN^{+}07, dRRdCRR16, RTMZ13, RMG^{+}10, RM11, RO12a, RO12b, SRS16, SM02, SV09, SAB15, SFLS04, SRM^{+}15, SG16, SIOS02, SAdB^{+}16, SHG^{+}07, SCBH09, SF16, SGV12, SM09, SD11a, SFH13, SFT15, SS15b, SE01, SK18, SCLK15, SZXG19, SVN12, SM19c, TKB09, TCDMR^{+}17, UR04, VDB09, VBW06, VCP16, VFAD17, WL11a, WAD12, WMC17, WLF19, WT10, WK07, WJKS18, WMDM07]. applications [WCL^{+}10, WSWL12, XXLL17, XYS17, YT15, YYC^{+}19, YK10, YJZZ19, YL01, ZBP06, ZBE17, ZBEM18, ZP06, ZP19, ZYLT06, ZHGX16, ZZ17, ZWL^{+}19, dCGKG06, vHMB08, GTGT11, HF17, LTL^{+}17]. Applied [WT15, DBH^{+}17, DAB09b, JZMD19, JKD19, MCB14]. apply [FMS11]. Applying [AMVOSGAC17, BHD13, CAG^{+}13, CBP^{+}04, CKBB14, GRQ19b, KW11, LWW06, MCY^{+}07, SGSC08, VSKK09, dCHMJ12, ASG^{+}08, sCR19]. approach [AOK19, Air17, AAHRW04, AR16, AMSS15, AML^{+}15, AAHA18, AMP^{+}18, ACMM06, ATİ17, AMS17, ADD^{+}05, AUHWJ19, BA18, BBG17, BTCB16, BPL12, BKM^{+}07a, BAZ09, BPS19, BCC^{+}05, BJC17, CWZL13, CLW^{+}19, CAC15, CLMM12, CG10, CKOG10, CLYC16, CCLP19, CH19, CWMW15, CLZ^{+}17, CLS14, CHH18, CL07, CBP^{+}04, DST11, DMRS15, DGA^{+}10, DED07, DAB09b, DS15, ELM^{+}16, ESZ09, EAGVBVDS11, EFA^{+}17, FHH15, FMS15, zGWXT09, GKP^{+}19, GH08, dAGC11, GVC10, GIL17, HKA^{+}15, HSJ^{+}18, HM16, HAA^{+}17, KL19, KR15, KT19b, KF18, KHZ^{+}15, KHVK17, KTM^{+}09, LBV16, LWF^{+}15, LZH^{+}15, LWT^{+}16, LGQ^{+}17, LZW17b, LWY15, LLL^{+}19, LPA^{+}08, LSL^{+}17, MJZ17, MY17, MZJ^{+}19, MTA^{+}07, MSS16, MJD17, MCXP15, MPVT17, MK15b, MLVBW12, MSG10, ORdSL13, OZI19, PFC14, PZHS18, PV04, PDCA17, PGK11, PCD15, PSIP16, PSW11, PBK19, PS19b, PME^{+}08]. approach [RSSM06, RHuR^{+}19, RGCC15, RMCHMG15, SJVR15, SB18, SD15, SS17b, SA19, TAMC19, THF15, TYTY15, TJZ^{+}19, TTR^{+}10, VGL06, VLW19, VH12, VFG11, VRDTB^{+}16, VO15, WBHW08, XWFH08, XZHW09, XDL^{+}11, XWD^{+}12, XDE^{+}04, XBZ10, XHCL15, YA04, YT15, YZ10, YHH13, Zen19a, ZYFZ19, ZLY^{+}13, ZH16, ZMYA18, ZLT^{+}16, ZFT08, ZQMC19, ZDR^{+}18]. approaches [AAF17, ABS16, BDT01, BCM^{+}05, COC18, MPSGD14, MJ19, OK18, PGP^{+}10, RBDI17, SİM^{+}07, Sod05, VLJ17, YDB^{+}13]. Approaching [IAH^{+}15]. approximate [GG09, GE08, IPGCMW18, WGY^{+}19]. approximation [BEDK18, CCTW11, CS13, JVMN19, WJ12]. approximations [CNP^{+}15]. apps [SSC^{+}16, YWL^{+}17b]. Araport [HVM^{+}15]. arbitrary [HP11, KMA04]. arbitration [LGL^{+}17]. architecting [AMABS18, Mit17a]. Architectural [BCC^{+}05, MCCG11]. Architecture [CLH^{+}11, IC19, MP04, Nel05, AP10, Akt18a, ADSV16, CT12, CLL14, CCL^{+}17, CHX^{+}19, CS17, CM07a, CJ12, CSB^{+}16, CKL19, CBIGL19, CMT13, CKNW06, DDF16, DBGA16, DXZ^{+}16, FAM^{+}18, Fer15, FNBS16, FTT15, GWW^{+}10, GAM17, GW15, Has17, HCK^{+}08, JVPI18, JLCA07, Kar14b, KHZN06, KJ19b, KPS14, KG19, KAMB19, LHC14, LSH^{+}16, LGQ^{+}17, MLS^{+}15, May10, NJM19, OCC^{+}05, PRS16, PSLC11, PS19a, PSJM13, ROA^{+}07, RW10, RCR^{+}15, RGL^{+}15, RHS17, SDB02, SPLO06, SB19a, SPW09, WCZX16, WL02, WLL14, XL17, ZW17, ZFT08, ZWW14, BBCG02, KKJH03, Zho06]. architecture-level [WCZX16]. Architectures [FK19, MN10, AHP^{+}13, ABC^{+}16, ABC^{+}15, ACS10, ATNW11, BOF15, BG14, BSP11, BIK^{+}11, BKZ^{+}13, BS10, BRCV16, BLKD08, CRC15a, CL18, CACC11, Cha03, CKL17, CNG13, CZG16, DFTHD18, DDM16, DCK12, EOD^{+}19, FBH^{+}01, FN13, FS18, GGV14, GVC10, HMM^{+}09, HLYD12, HBKM06, HdV13, JKD19, JKV^{+}15, JPS17, KB06, KMG^{+}18, LGP19, LL16a, LF17, MCP^{+}12, MLC04, MO02b, NO02, OAS^{+}15, PZ11, Par02, PLM^{+}19, PA18, PHCR09, PH12, RLMG16, RWK17, RHBK11, RGB^{+}15, SCRV11, SHT^{+}17, SRM13a, SF16, SFH13, SFT15, SHC^{+}16, SJPB17, STL^{+}15, SEF^{+}14, TYL^{+}15, VDL^{+}15, WS17, WS19b, YR15, vdABST10]. archives [ZKR^{+}07]. Archiving [Wit10]. Area [CS09, BMA03, DLW19, GHMX13, IC19, NLFA19, NS19, PS19a, RMP13b, XPBS11, XY17, ZMJZ10]. area-efficient [NS19, PS19a]. areas [RRR15]. Argus [FGC06]. arithmetic [KPNS18]. ark [BDG08]. ARM [BYDC19, BBdS^{+}17, CJX^{+}19, MOF15, MSPDP19]. ARM-FPGA [BYDC19]. arrangement [DBR13]. Array [CGK14, CM05, GvDHS12, Kuł14, LGFM05, WBO16, MDV07, RVVPD^{+}17]. arrays [AR19, CAD^{+}18, Kes04, LK03, MMG03, NNH^{+}14, TBK06]. arrhythmia [BTCB16]. arrive [LGCJ^{+}13]. arrow [GE08]. arrow-type [GE08]. art [BKZ^{+}13, Boc19, DHC13, MRS08]. artifact [ZZZ^{+}15]. artifacts [ZWW14]. Artificial [KC15, Bri16, CH19, DMM^{+}07, JZMD19, PB12, Pac16, PN19]. ary [PGL^{+}17]. ASCI [KHW05, RBBH02, YBC^{+}07]. Asian [LL16a]. ASKALON [FJP^{+}05]. ASL [GF07]. ASP [OISS07]. aspect [BAVM11, MSS16]. aspect-oriented [BAVM11, MSS16]. aspects [ACS10, GMPT15]. assemblies [TMR^{+}07]. assembling [BPD06]. Assembly [PC17b, CPS^{+}14]. assess [CHP17]. Assessing [HAvI13, LHBW15, MMSG17, MAS^{+}14, ZBZH11, BBB^{+}14, MDX14, PRS16]. assessment [CLMM12, FRU12, GGS^{+}16, HLY18, OTT19, YJL12]. asset [DCJ12]. assign [PB12]. assignment [LLYL09, MS19, SPJ14, Yos06]. assignments [TKZQ17]. assimilation [vHvdSvL03]. assist [KSK19, KKK^{+}19]. assistance [VZB19]. Assisted [dMd^{+}17, CJX^{+}19, Gog11, KM19, MWPX17, MPVT17, OTT19, RMG^{+}10, VEJD17]. ASSL [VH12]. associated [CLZX10, ZWLY16]. Association [KT19c, DBH^{+}17, HZH^{+}19, WWL^{+}15, WW19, XLMH14, YBX^{+}17, Zhe16, LLX15b]. associative [Kri13]. assumption [LZC14]. assurance [AL04]. astigmatism [DYY^{+}19]. astrophysics [BSC^{+}15, SWB12, vLRF^{+}02]. Astroturfing [PDCA17]. Asymmetric [ZLA^{+}15, GA09]. asynchronization [FSG19]. Asynchronous [GvDHS12, WK12, dCPD13, DPK10, IPGCMW18, KW11, LL19c, PH12, QLC04, The01]. ATF [RG19]. athlete [PdCMdS^{+}12, XS19]. ATIS [LB19]. Atmospheric [HGT14, RZL^{+}19, SPZ^{+}10]. atomicity [WKZL19]. atoms [GSB^{+}12]. atria [KSM^{+}08a]. attack [ANH16, AYSZ14, CYD^{+}15, FCY17, GYS^{+}17, GKM19, HYLG15, VRDTB^{+}16, Yan19b]. attack-defense [FCY17]. Attacks [LH14, BGdCCA11, CXW17, DCG15, Eng15, EA12, ECP18, FIO15, HAK19, TJZ^{+}19, VT15, XTZ10]. attempts [WLZ11]. attenuation [LLY^{+}19a]. Attitude [CH04]. attitudes [ZQR^{+}19]. Attribute [BXLJ16, JSG17, HWL18, LFZ^{+}17, LFWS15, LJW^{+}17, Nam19, PS19b, WZC16, YSQM19, ZWY^{+}19, SACRGL18]. Attribute-based [JSG17, LFZ^{+}17, LFWS15, LJW^{+}17, Nam19, WZC16, YSQM19, ZWY^{+}19, SACRGL18]. Auction [SN18, HK01, HJM^{+}11, SO16, WRDZ13, ZLH^{+}15]. Auction-based [SN18, SO16]. Auctioning [GD06]. auctions [Kar14b]. audit [YXL17]. auditing [FJG^{+}13, HZY^{+}19, LLSL15, LBY^{+}16, QXJS17, WZL^{+}17b, YCSY19]. augmentation [LFPP17]. augmented [BDP18, TLPS18]. Augmenting [AL04, KCS07]. Australia [Run10, SBB^{+}15]. authenticated [FIO15, XWXC14, ZZC15]. Authentication [BNNH19, AaBT16, Aia15, AZF^{+}12, AL04, BT18, EA12, FLL^{+}14, GLC^{+}04, HZC^{+}14, JNUH17, KDW^{+}17, LCM^{+}17, MMS17, NR17, NLYZ12, SVY19, SA19, Tan15, WDLL10, WXSH19, XHH12, XXCY19, YHHS16, YYK^{+}19, ZQD16, ZZL^{+}18]. author [HPK^{+}18]. authority [BFG14, ZWY^{+}19]. authorization [CZO^{+}08, GRSB09, KWL^{+}04, MTGZ17, SW11, TSL15]. authorizations [Rua15]. authorship [LCM^{+}17]. Auto [Sør13, FE17, PNJP19, RG19, TV14, WGQ^{+}18]. auto-combination [WGQ^{+}18]. auto-parallelization [PNJP19]. Auto-tuning [Sør13, FE17, RG19]. AutoDockCloud [EB14]. autoencoder [WTY^{+}19, YTL19]. autoencoder-based [WTY^{+}19, YTL19]. automata [Mos19]. Automated [AAF^{+}07, JAU19, LLY^{+}19b, BTCB16, CMW02, CK13, CVK15, CHH18, DS15, GWHJL19, KAM11, MAK18, RCXS09, RM03, SGSC08, SMBT07, SE01, SM19c, ZYFZ19]. Automatic [BD08, BKND16, CZ19, FMS08, Ger05, GG07, GO10, KKTHL13, LXP18, MFGE19, RPRG17, SD03, SPBL06, TLM17, WMDM07, BR04, BFK^{+}17, HL19, JCVU15, KGGT12, LM07, LHHJ18, LLLyL16, LLZ^{+}17a, NFG19, QEB^{+}10, SL18, TJF14, TDM^{+}19, WXY10, WFHT17, YSWZ17]. Automatically [SSC^{+}16, ZZ11, AAP13, ML19a, YOBS16]. automating [BHD13]. Automation [HBG^{+}06]. automobile [YZ19]. Autonomic [MBP^{+}05, TCBR11, CRCC09, CR12, CSL12, DED07, HM16, LDS^{+}08, RRBB11, RBP12, ZDR^{+}18]. autonomic-computing [ZDR^{+}18]. Autonomics [Pat08]. Autonomous [FZ07, Zhu07, BT18, JDB16, KOO12, Kri13, SJ19, SAC^{+}07]. autoscaling [VŠC17]. AutoSLAM [CVK15]. Autotuning [BAG17, NDL17, AHK^{+}15, Lan17, PAC^{+}17]. auxiliary [LWZ^{+}17].

Availability [XPWF15, ANTZ09, DGM18, FBC10, KKTHL13, MLG15, MS18, PV15]. available [KM03]. avalanche [PIH04]. avatars [TNH15, TNI16]. AVC [RSMFE^{+}12]. AVO [WZYG19]. avoid [CM06]. Avoidance [ZKWK17, SWS^{+}18, VKM^{+}09, YESG^{+}19]. avoiding [WS09]. AVPredictor [WKZL19]. Award [Bou13]. aware [AA19, ABC19, And13, ACIC^{+}13, AMAB17, ACCM17, AMS17, BDE^{+}19, BKS18, BFM^{+}06, BPB08, BHKW12, CEH^{+}06, CGST17, CRCC09, CSL^{+}18, CLQ^{+}17, CC15, CZG16, CLT^{+}16, CLH13, CFTT17, CDN15, DFG^{+}18b, DMRS15, DCP^{+}17, DPGA11, DHC11, DZL^{+}17a, DXM^{+}17, DYW16, DA15, EN19, EK19, EOD^{+}19, EQORS19, FA18, GYM14, GMVRGS15, GAW09, GBJ19, HKS19, HAAWA^{+}16, IHA^{+}15, IRB19, ID18, JZL14, JZL15, Jon09, KC15, KV12, KL12a, KBB11, Lan17, LBdM^{+}16, LWYM16, LDXC13, LFH08a, LYC16, LSJ16, LXK^{+}19, MTM19, MS13, MRL16, MSP^{+}13, MDB^{+}17, PCF^{+}17, PGC^{+}19, PPBB14, PRV11, QLD^{+}11, RHZ^{+}17, RIWS17, RAFD14, RGB^{+}15, STO17, SJW18, SBDP15, SGV12, TKZQ17, TDM^{+}19, VP19, WSL15, WQS^{+}16, WJXZ18, WDT18, WDW^{+}15, WSW^{+}12, XDJL18, YBO10, YSC^{+}17, ZLY^{+}13, ZWH^{+}17, ZFXJ19, XWPM19, YCWH07]. awareness [CAC15, LBOE18, RH07, TWZ^{+}19, YCQF19]. AXC [CBIGL19]. AxML [SLM04]. Azure [CTAB16, KBT^{+}14, LRS15, XBB13]. Azure-based [KBT^{+}14].

B [CJX^{+}19, IS10, MS10, PLL17]. B-mode [PLL17]. B.E. [BHH09, VSR^{+}09]. Babylon [vHMB08]. back [LMS18]. backbone [NSSAK13, NSSAK16]. Backfilling [LGCJ^{+}13, WGZL06]. background [CZJY19, KMA04]. backoff [WYQ^{+}13]. backtracking [FH13, PGdCJ^{+}18]. backup [BDP18, GÖ18, LYS18, LWYZ19, ZXW16a]. bacterium [ALVY05]. bag [PCGE18, PRV11, ZS17, NB12]. bag-of-task [PCGE18]. bag-of-tasks [PRV11, ZS17, NB12]. Baidu [ZHX^{+}19]. balance [HWL18, SBDP15]. Balanced [MS17a, LZZ^{+}17, MS17b, WQS^{+}16, ZLLL11]. Balancing [DT17, FSG19, WTL^{+}16, AS15, AR16, APHB16, BGV^{+}01, DBR13, DCCZ18, EB18, EOD^{+}19, FJ05, FE18, FT06, GRQ19a, GCL08, KKTHL13, KYM17, KR04, LJL^{+}17, MKIO04, PBK19, QCB17, SVB19, WJYH16, WLL03b, XBZ10, XTB17, ZEB10]. band [DK19a]. banded [BHL^{+}09]. bands [HCKF15]. bandwidth [GDD^{+}04, Hic18, LWF^{+}15, PIH04, YLR^{+}13, ZWMT12]. bar [AMRT14]. Baraglia [Ano06]. Bargaining [HJTX17, YCZ^{+}13, SPJ14]. Bargaining-based [YCZ^{+}13]. barrier [TZKH12]. barriers [KCBO17, WBM^{+}10]. BASARIM [UA18]. BASARIM-2017 [UA18]. base [cDrLyC^{+}19, XZH^{+}16]. Based [HJTX17, MN10, ALKD16, AaBT16, AFGL09, ABC^{+}16, AQRA^{+}18, ASWR12, AR16, AM15, AK01, AMRW06, ABC^{+}08a, AKG13, AC02, AAQAR^{+}17, AMABS18, ADSV16, AMP^{+}18, ALL^{+}15, ASS19, AUHWJ19, ABG^{+}13, Bai17, BTCGL17, BCI^{+}18, Bao19, BDZ19, BM10, BBG17, BEDK18, BTCB16, BOB13, BVIB19, BLL18, BKCP09, BXLJ16, BKH08, BZD16, Boc19, BAZ09, BBB16, Bou13, BAZ18, BCC^{+}05, BLDW16, BWEB14, BJC17, BHPS14, CMW02, CRB^{+}17, Can06, CYD^{+}15, CSL^{+}18, CAC15, CC10, CRC^{+}15b, CGMJ^{+}19, CR12, CJC^{+}18, CWL03, CA06, CY07, CWYX17, CLW^{+}18, CCLP19, CLH19, CH19, CLL^{+}19, CWMW15, CXT^{+}18, CL19, CVK15, CM06, CKC09, CW07, CL07, CSB^{+}16, CM02, CGB^{+}06, CNPP09, CRGR^{+}12, CMT13, CLX^{+}12, CDF^{+}17, CMD17, DD17, DVD^{+}12, DCJ12, DHV03, DBR13, DBGA16, DXG13, DRS^{+}13, DLX^{+}16, DK19a, DS19, DCY^{+}08, DG11, DGR^{+}07, DAC12]. based [DPM17, DHM14, DHH^{+}13, DPS07, DNB19, DBH^{+}17, DAC^{+}18, DH13, DJ19, DWG19, DC19b, Dra15, Du18b, Du18a, DYY^{+}19, DRF07, DT15b, EB18, EPB14, EMEY14, ET09, EFY17, ECP18, EAGVBVDS11, EFA^{+}17, FE17, FHH15, FPHZ20, FXX16, FIO15, FSM^{+}19, FJZ^{+}14, FPC15, FAPC16, FVRM15, FK19, FH13, FN13, GS08, GYM14, GHLS19, GDJ16, GED^{+}18, GMMT17, God12, GSR^{+}19, GIVRC^{+}10, GS04a, GE08, GIL17, GBG^{+}14, GYS^{+}17, GSWJ19, GHB^{+}06, GAB19, GLD17, GGC19, GPZ04, GKP^{+}09, HFDJ10, HZC^{+}14, HZH^{+}19, HZHP09, HXY^{+}12, HAJL16, HLF^{+}17, HR18, Hoh06, HCS18, HSHT14, HM16, HLL^{+}15, HLCW15, HLZD18, HWR03, HFTQ13, HLC19, HYGF19, HML19, HZY^{+}19, HGB^{+}08, HW16, HCK^{+}08, HY12, JC07, JKM^{+}17, JNUH17, JBL15, JTD^{+}19, JQSP08, JJGL13, JLHH14, JML^{+}16, JZMD19, JR19, JZZL06, JWW17, JLL18, JSG17, JPWH02, JSS07, KL19, KKDS19, KT19a]. based [KC15, KBDA19, KM19, KHZN06, KGGT12, KR15, KAA19, KB17, KHHC13, KHL17b, KQR^{+}17, KBT^{+}14, KT19b, KJŚ^{+}15, KKWZ15, KZY15, KZY^{+}18, KKS12, KABD07, Kri05, Kri13, KPS14, KFD18, KT19c, KR11, KBH^{+}15b, KAMB19, KSC12, LVN^{+}12, LSXL17, Lan17, LGL16a, LLN^{+}14, LHL10, LM08, Li04, LLH^{+}09, LWC12, LMKT13, LDZ^{+}14a, LLL15, LDZ^{+}15, LLG^{+}15, LLX^{+}15a, LZY^{+}16, LWYM16, LWY^{+}16, LCT16, LGQ^{+}17, LH17, LZW17b, LXYC17, LWZ^{+}17, LFZ^{+}17, LWQS19, LWZ^{+}19, LLY^{+}19a, LCW^{+}19, LSZ19, LZG^{+}19, LLT19b, LL19a, LWW^{+}19, LFWS15, LGL^{+}17, LWB13, LLH^{+}15, LLY^{+}19b, LMCL19, LHB^{+}19, LHT^{+}09, LWLZ11, LZC14, LLX15b, LGG16, LJW^{+}17, LJL^{+}17, LML^{+}18, LLH^{+}18, LC18, LMX^{+}18, LLLS18, LFW19, LL19b, LLL^{+}19, LHZX19, LWP19, LAL02, LSW07, LMDP19, LPC^{+}14, LZZ^{+}15, LCJ14, LHXY08, LXL^{+}09, LL19c, LLY19c, LSL^{+}17, MLL^{+}11, MLS^{+}15, MWPL15, MRY^{+}16, MY17, MYY18, MZG19, MB17, MRMC15, MS17a, MS18, MHLC^{+}05, MZ06, MMO^{+}16]. based [MB12, MMMP01, MSST15, MZW^{+}16, MBB19, MD19, MK15a, MTT15, MT19b, MCY^{+}10, MKAKG14, MCCD18, MRJ^{+}14, MML^{+}17, MB14, MWL^{+}15, MW18, MJS19, MT08, Mos19, MLYL17, MWL18, MS19, MNL15, MSG10, NIIU17, Nam19, NNK^{+}07, NNvVdA09, NC05, NMKB03, NJ05, OKO18, OLG^{+}15, OZI19, PWC^{+}14, PSRR14, PFC14, PYKL16, PCsHL18, PRCV16, PGL^{+}17, PDCA17, PWH18, PLW^{+}18, PCF^{+}17, PAdS^{+}17, PSICU18, PCT04, PDC16, PPC^{+}15, PSHL11, PC17a, PSS^{+}18, PK17, PSC^{+}17, PB16, PCD15, PSW11, PGW^{+}08, PME^{+}08, PJW^{+}14, QLLS15, QXXZ16, QSZL18, Qia19, QZDJ16, QLZX19, QML^{+}17, RM19, RGAK15, RBO^{+}02, RSKK19, RG19, RR15, RLZ15, RZL^{+}19, RMCN^{+}07, RGCC15, RSMFE^{+}12, RIWS17, RWK17, Roj19, RCLSK16, RG17, RHS17, RCT03, RRWS08, SJB14, SACRGL18, SS17a, SAB15, SR19b, SBBE07, SRM^{+}15, SPR^{+}07, SGD15, SAMS19, SS15a]. based [SIS19, SARL13, SWS^{+}18, SCLL19, SS18, SWLJ17, SN18, SZ18, SSIH19, SACJ04, SR17, SB19a, SPBL06, SHC^{+}16, SM19b, SO16, SWW^{+}16, Soo16, SLM04, STL^{+}15, SWD^{+}17, SWZ^{+}18, SR19c, SC07b, SS17b, SC19, SW12, SG18, SZD19, TWZ^{+}19, TAMC19, TJ17b, TZYL13, TQL^{+}14, TZLC15, TYL^{+}15, TYTY15, TLX^{+}17, THQ19, TCP^{+}05, TFG^{+}12, TDL^{+}18, TV14, TSBR10, TBTZ18, TBK^{+}15, TBH^{+}18, VYKM19, VS02, VDPC03, VDB09, VRSJ15, VP19, VRDTB^{+}16, VFAD18, VO15, VvSI07, WBD^{+}19, WKB^{+}19, WYZ12, WZ04, WKT08, WLDL08, WRC09, WDLL10, WRLS12, WJ12, WZJD13, WZZL13, WCR^{+}14, WZS^{+}15, WZC16, WFHT17, WLFX17, WWL^{+}17a, Wan18a, WDQ^{+}18, WF18, WZLL18, Wan18b, WFKS18, WML^{+}19, WTY^{+}19, WWX^{+}19, cWsThY19, WW19, WXSH19, WWC^{+}19, WHX19, WGY^{+}19, WFS^{+}19, WJP14, WBC^{+}17, WZXZ12, WK07, WJKS18, WCLH12, WRDZ13, WWL^{+}17b, WT18, WYY^{+}19, WXLM19, XHH12, XWFH08, XDL^{+}11, XWD^{+}12]. based [XBZ10, XZZ^{+}16a, XLG19, XZJ11, XZH^{+}16, XDP18, XYSW18, Xu19, XXY^{+}16, YCZ^{+}13, YWL^{+}17a, YWLQ18, YHYY19, YTF^{+}01, YHK09, YP10, YWC11, YLWZ18, YYXL19, YSQM19, YKA^{+}19, YYZH19, YTL19, Yan19b, YC19a, YT15, YBZ^{+}15, YKD^{+}15, YLZ18, YYLC19, YZXW17, YYWQ19, YLEB14, YZ10, YHH13, Yu18, YZ19, YCSY19, YL19, YYL^{+}12, YG19, ZK08, ZF18, ZW09, Zen19a, ZAB^{+}19, ZP06, ZCC^{+}06, ZEB10, ZLLL11, ZTM12, ZJL13, ZlC15, ZLH^{+}15, ZYW^{+}16, ZQD16, ZZ18, ZMYA18, ZBZ^{+}18, ZHC^{+}18, ZLH^{+}18, ZCZ^{+}19, ZJT^{+}19, ZL19, ZWY^{+}19, ZXC^{+}19, ZHJ19, ZLL19, ZZY^{+}19, ZM13, ZLT^{+}16, ZCLL19, ZLG^{+}19, ZFT08, ZACG16, ZWW^{+}18, ZSS18, ZZZX19, ZQMC19, ZBZH11, ZFJ16, ZCXH17, ZQW^{+}17, ZCL^{+}19, ZFXJ19, ZXXN06, ZZL^{+}17a, ZCS06, ZWMT12, dOOO^{+}12, dMd^{+}17, vHKT^{+}11, vNMW^{+}05]. bases [NZKK11]. Basic [ZX11, CGGH17, SKNH09]. basin [DLM13]. basing [JB19]. basis [CQXW14, JLQ^{+}17, KF01, SPZ^{+}10, ZXX17]. bat [AMS17, KT19a]. batch [LWLZ11, MHRI14, SRS16, SR19a, SAB15, SVN12, VLMPS^{+}18, WCLC13]. batch-of-stochastic-tasks [SR19a]. batch-of-tasks [SRS16]. batchsize [Hey19]. Batchsubmit [MHRI14]. bathymetry [MMG^{+}18]. battery [CLH13]. Bayesian [NJM19, WZT11, WZJD13]. BC [LBY^{+}16]. BDS [HYQ17]. BDS-IDC [HYQ17]. be [PCL17, EMS11]. beacon [AAF17, WT18]. beacon-enabled [AAF17]. beaconless [JWZ13]. beam [ZGZ19]. beamformer [PL15]. Bear [ON01, ON02]. Bee [MS17b, KC15, PN19]. bees [DBH^{+}17]. before [JW10, LSS15, PWJ10]. behavior [AAF^{+}07, CZJY19, CCS14, GGR^{+}10, GYS^{+}17, KL02, LF15, LSZ19, MSV^{+}10, RS16, SS18, YWL^{+}17b]. behavioral [IÁE11]. behaviors [DJ19, GBXL17, HL19, KCD19, ZQR^{+}19, XLL^{+}18]. behaviour [ADK^{+}17, MDX14]. beings [GQR16]. belief [KHL17b]. Bench [SPQ^{+}17]. benchmark [BCD^{+}10, BG04, CLL14, DS02, EHSU07, GPW03, GPW05, KB18, MvWL^{+}10, SCS17b, SPQ^{+}17, DLP03]. Benchmarking [BSB^{+}03, GFG^{+}09, MP05, BCM^{+}05, DMR^{+}07, Dik07, ZS01, ZCL14]. benchmarks [KHM^{+}11a, NNON02, SZT18, SCC^{+}10]. benefit [Zen19b]. benefits [SIRP17]. benevolent [XLT^{+}17]. Best [CS09, LME^{+}19, PB07b, PK08, GRGP12, MS17c]. better [LWW06, VAC^{+}07]. between [Hun15, IÁBE11, JPS17, KHW05, Kri13, RG18, SZT18, WDQ^{+}18, XZZ^{+}16a, XLYX11a, ZYL10]. BFG2 [AFR09]. bi [KSPM12, LOKW^{+}10]. bi-criteria [KSPM12]. bi-material [LOKW^{+}10]. Bias [GC18]. Bias-Sentiment-Topic [GC18]. bicycle [LHZX19]. bidirectional [LWG^{+}15]. Big [BTCB16, BM16, LWL15, LLLyL16, LLL16, RLC16, SG19, APHB16, BA18, BLXE19, CY15, CZJY19, DS19, DM15, DS17, DL17, DSH18, DWG19, DXM^{+}17, ESG17, GLS^{+}19, HWQ^{+}16, HL19, JLQ^{+}17, KT19c, LTL^{+}17, LYF^{+}17, Li18, LSZ19, LGL^{+}17, LBY^{+}16, LMX^{+}18, LDZ^{+}19, LMDP19, MYS19, PCsHL18, PIGK16, QWW^{+}16, QZYZ16, RM19, TDC18, WJYH16, WQL^{+}18, WS19a, XAK16, XSMZ16, XYER16, XGXH15, YZXW17, ZBE17, ZBEM18, ZLN^{+}13, ZSS18, Zhu18, ZS19, HYQ17, Mar19, NJM19, PSIP16, SG16]. Big-Data [SG19, ESG17, PIGK16]. BigData [ZH16]. BIGhybrid [AFG16]. bilevel [LZZ^{+}15]. Bilinear [LZY^{+}16]. Bilinear-map [LZY^{+}16]. binaries [GWHJL19]. binary [CL14, CGR19, HTWW19, HAA^{+}17, LCM12, LCW^{+}17, LLH^{+}18, LLL^{+}19, MPS11, PDCA17, SR17, ZZ14]. binding [DXZ^{+}16, SSU18]. BindMe [SSU18]. BinLPT [PGC^{+}19]. Bio [OK18, WSL15, ABG^{+}13, CBHTE11, CSL12, CP14, GPVCdBRO12]. Bio-inspired [OK18, WSL15, ABG^{+}13, CSL12, CP14, GPVCdBRO12]. bio-science [CBHTE11]. biochemical [KOK14, LTM^{+}14]. Biocompute [CBHTE11]. biodiversity [ABB^{+}15, WSP17]. bioextract [LGD15]. biofuel [HLL^{+}15]. bioinformatic [GvHKK11]. bioinformatics [BAD^{+}11, GFG^{+}09, HSRN11, LBTE14, PRC^{+}14, SFLS04, VRSJ15]. bioinspired [HdV13, LGdVH13]. biological [AHP^{+}13, GR13, KKW^{+}14, SKA^{+}14, YGG14]. Biologically [PCS^{+}12, HAE09]. Biologically-inspired [PCS^{+}12]. Biology [BA04, Mar05, CCP^{+}15, GPP^{+}18, LTM^{+}14, MPSGD14, THM^{+}11, WOH^{+}13]. Biomashups [HSRN11]. biomass [HLL^{+}15]. biomass-to-biofuel [HLL^{+}15]. biomedical [GWC^{+}11, WJKS18, YOBS16]. biomedicine [VRSJ15]. biometric [SWZ^{+}18, SC19, ZQD16]. biometrics [LH14]. Biomimetic [ZWZ^{+}18]. bionic [DCCZ18]. bioscience [HCG07]. BioScope [HLL^{+}15]. biosensor [SR19b]. bipartite [PPR19, XWX^{+}17]. bipartitioning [MUKY18]. Birds [PCS^{+}12]. bit [HLO^{+}16]. bit-parallel [HLO^{+}16]. bitmap [YWBS19]. bitonic [PSHL11]. bitstream [WYZ12, Zhe16]. BitTorrent [JJGL13, LNKZ08, TWN07]. BitTorrent-like [TWN07]. black [HW14, KC13, BHPS14]. Blacklight [CPS^{+}14]. BladeCenter [SLM^{+}10]. blame [BWEB14]. BLAS [Sør13]. BLAST [Kri05, SL14, YHK09]. blind [CLS14, LLQL14]. Blinn [DG11]. Block [Has17, ABV05, ADF^{+}19, KYBV17, LGP19, LB11, PZZ08, PZZ10, SAD13, SGR19, TPGC15, TQL^{+}14, XGC19, MDL^{+}10]. block-Jacobi [ADF^{+}19]. block-structured [LB11]. block-Toeplitz [ABV05]. Blockchain [HLC19, HZY^{+}19]. Blockchain-based [HLC19]. blockcipher [CMMS17]. Blocking [LGY17, Cho01, ESGQ^{+}11, KYBV17, YESG^{+}19]. blocks [IQOvdG13, Tan12]. blog [LWY^{+}17, LWP19]. Bloom [ATKH^{+}17]. BLOR [LWF^{+}15]. Blue [EMS11, KB18, RGL^{+}15]. Bluetooth [CNPP09, WCCL05]. Bluetooth-based [CNPP09]. BM [WSZ^{+}18]. BM-MOPSO [WSZ^{+}18]. board [ABDO09, ZJS11]. Boas [Kuł14]. body [CGK14, GGV14, SSB^{+}14, TL14]. boldly [LSS15]. Boltzmann [BFM^{+}10, CGK^{+}16, CNAQ18, MWLS11, QWZ^{+}19, RMW19, VLJ17, ZAB^{+}19]. bone [BCA^{+}10, THM^{+}11]. bones [CSC^{+}17]. bookmarking [God12]. Boosted [ADK^{+}17]. Boosting [ACIC^{+}13]. Border [DT15b]. BoT [SRS16]. both [HJH19]. botnets [KKS12]. bottleneck [PCsHL18]. bound [CMMB13, CT11b, Cuz11, FOTW04, GLM^{+}16, GMMT17, MCB14, PSIP16, SSIH19, SBDP15]. bounded [DZ13, LC09, PAdS^{+}17, WYZ^{+}17]. bounding [MCB14, YZW^{+}15]. bounds [FMP10, LGFM05, vRGNP09]. box [XHCL15]. BPEL [HLZD18, Ley06, Slo06, TMF^{+}10]. BPEL4WS [CKNW06]. brain [BDMM^{+}05, EMEY14, PVR^{+}09, WCZ^{+}18]. BRAMS [JPS17]. branch [CMMB13, GLM^{+}16, GMMT17, JKL19, MCB14, Mit19, PSJM13, SSIH19, SBDP15]. branch-and-bound [CMMB13, GLM^{+}16, GMMT17, MCB14, SSIH19, SBDP15]. Brand [DJ19]. Brazil [PS13]. Brazilian [GBMM15]. breaches [Kin04]. Breaking [WWS^{+}12, YZ19]. breast [VJ19]. BRGP [LZZ^{+}17]. bridge [MMSN^{+}01, VDL^{+}15]. Bridging [BKWM19, RSSM06, HMFK15, Hun15, MTHK14]. bring [ADM06]. Bringing [MD19]. Broadband [DAC12, RDP10]. broadcast [DÇK^{+}18, KHZN06, LL10, MTK16, XZH^{+}17]. broadcast-based [KHZN06]. broadcasting [AKMZ13, KMA04, LLKC08, ZQK15]. broadcasts [KGK17, KCS07]. broker [BKM^{+}07b, GCZ^{+}17, VBW06, AC02, ACC^{+}07, CEM^{+}08]. Brokering [DPGA11, ET09, KD15, PGW06, TSBR10, YLC11]. Browsing [CBQ^{+}11, mLGP03, LXL^{+}09, MM19, YYK^{+}19]. Broyden [PV04]. BSCHEF [MRY^{+}16]. BSNet [HFTQ13]. BSP [SGCA^{+}16, dCGKG06]. Bucket [WDLL10]. Bucket-based [WDLL10]. Budget [TLF17, KS19b, KD15, ZFWJ19]. budget-constrained [ZFWJ19]. Budget-constraint [TLF17]. buffer [CDN15, LWW06]. bufferless [GGLD11]. bugs [DS02]. Building [ASG^{+}08, CZ11, CJZZ10, HZL^{+}16, HYL^{+}19, KKL06, LS19a, RCXS09, SS19c, Tan12, WNN^{+}15, XZH^{+}16, YR15, ART14, ACS10, BAS07, Bri16, CWMZ06, DH15, HKG08, MSL^{+}14, MST15, NRW04, OTG^{+}07, PWWR05, SNEP14, SLD^{+}12, TMF^{+}10, VRMB13, ZWL^{+}13]. built [WWL^{+}15]. Bulk [YIN19, BDT01, Kes04, LLY^{+}19a, MWPX17, TNIB17, YB12, GDD^{+}04]. bulk-synchronous [Kes04]. bulletin [ABDO09]. bursting [ACC^{+}15]. Bursty [LWT^{+}16, GHMX13, KMA04, VO15]. bus [LWSZ19, LLC^{+}15b, LYC16, MS10, ZTZ^{+}18, ZWMT12]. bus-subway [ZTZ^{+}18]. business [CHH18, HFTQ13, IÁE11, IÁBE11, LFH08a, MWJ^{+}10, SvDO15, XLZD13, XWH^{+}17, AK01, DKMV07]. busy [TY15]. buying [ZLH^{+}15]. buying-based [ZLH^{+}15]. Bypassing [RG17]. bytecode [Cog03, Cog04, KN01, SD03].

C [Tan12, VDL^{+}15, Bou06, BSB^{+}03, GDMT^{+}12, IS10, KS04, KW01, KS05, MRY^{+}16, NTK08, PS07, SCBH09, SHST13, SSB^{+}14, TNIB17]. C# [BHW05, WLR05]. C-DBLP [MRY^{+}16]. C-RAN [SHST13]. C2C [LL19b, XZJ11]. C2C-PAKA [XZJ11]. C2CU [TNIB17]. C4.5 [MLYL17]. caBIG [HLB10]. Cache [HT15, AAF^{+}07, BBC16, CPB07, CFPJ^{+}17, CCSS10, CCW^{+}15, CS17, DP14, DC19b, GSG06, Gog11, GLS^{+}19, HPS12, ID18, KKGO04, KSC12, MY17, PRU14, SJW18, SC07b, SS17b, WFJ^{+}17, WFKS18, WZXZ12, ZZD^{+}17, dRL10]. cache-based [WZXZ12]. cache-efficient [Gog11]. Cache-oblivious [HT15]. Caches [AS19, BYN^{+}17, RAFD14]. Caching [Hic18, CWL03, LR05, SCLL19, SNB^{+}01, dRL10]. Cactus [DvdS06]. CAFS [WZXZ12]. caGrid [TMF^{+}10]. Cairns [Run10]. calculation [MFGE19, Str11, WYW^{+}17]. calculations [BGGL07, BDTdS13, PIH04, RGL^{+}15]. calculus [KCW09, QLF^{+}06]. calendar [LZC08]. Calibrating [SNEP14]. calibration [BLDW16, DLZ16, KD19, LCJ14, YSWZ17]. California [GGR^{+}10]. call [BM07, EPB14, GSR^{+}19, GLD17, HTWW19, WJKS18]. call-graph [EPB14]. call-path [BM07]. callgraph [CMW02]. callgraph-based [CMW02]. calls [KGP^{+}19, RG17, YWL^{+}17b]. camera [LCJ14]. CamFlow [KKK^{+}19]. campaign [HFDJ10]. campus [DDX^{+}06, MTHK14, CRB09]. Can [MTHK14, PCL17]. cancelable [LH14]. cancellation [CW09]. cancer [DMM^{+}07, VJ19]. candidate [QZYZ16]. CANFIS [MJ19]. canvas [SvDO15]. CAP [SPQ^{+}17]. capabilities [AL04, BCI^{+}09, MSP^{+}19, OK15, SPG08, VZB19]. capability [ABG^{+}05, RVD^{+}12, RBB^{+}09, XLMH14, YGL05]. capable [PRU14]. capacitor [LCW^{+}19]. capacity [CLZ^{+}17, GEJ^{+}08, LWYZ19]. capacity-constrained [LWYZ19]. capitals [LGJ17]. capping [HJV^{+}19]. CAPTCHA [HCBRM16, OKO18, OTO18]. capture [BD08, FMS08, MCY^{+}07, SGSC08, XS19]. Capturing [OORVB14, SP16]. carbon [AHB^{+}10, BBSW17, HM12]. carbon-flux [AHB^{+}10]. card [XZJ11, XDP18]. cardiac [RCA^{+}11]. cardinality [LLG^{+}15]. cardiovascular [KM19]. cards [AVS^{+}19]. caring [KF15]. Carlo [ATVLM14, CCO15a, GQH17, KDC17, NDT^{+}16, RDP10, SS15c, WZJD13]. CARMEN [WHW10]. carry [SS19a]. CartaBlanca [PCVZ^{+}04, VDPC03]. cascade [YHASZ19]. case [BPT^{+}16, BDMM^{+}05, CMCAA17, DT01, EDB^{+}14, EMB11, GFG^{+}09, GPP^{+}18, GRS^{+}17, GIL17, HKS^{+}12, HPS12, HCK^{+}08, KOK14, LBTE14, LLN^{+}14, LLH^{+}09, MCP^{+}12, NNK^{+}07, NNvVdA09, PRC^{+}14, RCC17, RNAD19, RGL^{+}15, SGD15, SvDO15, SIRP17, SE01, SLC^{+}18, SLC^{+}19, TMF^{+}10, TDC18, The01, WST^{+}17, ZL19, dABV08, dARP17]. case-based [LLH^{+}09, NNK^{+}07, NNvVdA09]. case-study [GPP^{+}18]. cast [RGCC15, WYQ^{+}13]. catalogues [SK08]. Catalyzer [HCG07]. catastrophe [BRCV16]. catchment [DLM13]. categorization [KGGT12]. causal [BMA03, MGM^{+}08]. Causality [CW09]. CBIR [PPP10, VYKM19]. CBMR [LH17]. CC [Cha03]. CC-NUMA [Cha03]. CCA [AAW^{+}02, AKM^{+}06, GCN09, GLC07]. CCD [Bao19]. CCGrid [OKG18]. CCGrid-Life [OKG18]. CCGrid’2007 [CS09]. CCJ [NMKB03]. CCLRC [ACMA07]. CCN [SQS^{+}19]. ccNUMA [CBPP02]. CDL [XDL^{+}11]. CDMA [MS07]. cell [KRW17, MLVB05, MNL15, QH10, RDP10, SR19b, VDL^{+}15, BHH09, DAC12, EMS11, KD07, SSK11, VSR^{+}09, ZDC^{+}09]. cell-based [MNL15, SR19b]. Cell/B.E. [BHH09, VSR^{+}09]. cellular [Mos19]. CENSOR [CKL19]. Census [DKMM14]. Center [HGT14, BKZ^{+}13, CYD^{+}15, DGW16, DMW^{+}10, FAM^{+}18, GSR^{+}19, HSM14, IRB19, LLC^{+}15a, LGY17, PLW^{+}18, ZWH^{+}17]. centered [AHB^{+}10]. centers [AMAB17, BB12, CRB^{+}17, DGL^{+}12, GLD17, HSL19, JZL14, JZL15, KTB17, LWYZ19, ROQL18, SZG^{+}19, SPSB19, XLQL18]. Central [WHW10]. Centrality [AMP^{+}18, BOF15, CEB^{+}18]. centralized [CRC15a, DKMV07, WGZL06]. centric [CN16, KSM^{+}08b, Kri13, PBF15, SBJ^{+}15, SCLL19, SSIH19, ZJS^{+}17]. centroid [FRKS12]. century [BHJ^{+}16]. CERE [PAC^{+}17]. certificate [LDZ^{+}14a]. certificate-based [LDZ^{+}14a]. certificateless [DXWD16, WWC^{+}19]. certificates [BAD^{+}11]. certification [BFG14, HY12]. certified [XWXC14]. CFD [FBV^{+}13, GHLS19]. CG [ABF^{+}17]. CGC2011 [CL13]. Chain [LXP^{+}12, CHX^{+}19, HAJL16, HLL^{+}15, KSR14, LWC12, LS19b, ZACG16]. Chain-to-chain [LXP^{+}12]. chaining [TSA^{+}19]. challenge [CBBCD08, GH08, HSBMR08, LS14, PBD^{+}15, MLA^{+}08, SKS^{+}08]. challenged [FP09]. Challenges [YWT^{+}12, ZQH12, BCA^{+}10, CSAC19, Dik07, DHC13, FBV^{+}13, LLT^{+}14, PHY^{+}18, PCJ17, PCB^{+}18, PT12, SN18, WJJM17, LF15]. CHAMELEON [DZZL19]. Chan [YHJ^{+}14]. change [BB19, JLQ^{+}17]. changes [PWJ10]. changing [SWU08, ZCL^{+}19]. Channel [VP19, Du18a, DXZ^{+}16, HKB07, KT19a, LWG^{+}15, LWW06, MTM19, MS07, SCLK15, ZKWK17]. chaos [MSV^{+}10]. chaotic [LWW^{+}19]. Character [TJD^{+}17, ZSS18]. characteristic [KHW05]. Characteristics [LZW^{+}17a, DAC^{+}18, PIH04, WLZ11, ZDHJ18]. Characterization [JVMN19, dOCPFJ13, HKS^{+}12, RGL^{+}15, SCC^{+}10, SMS^{+}19, dP06, vAVS12]. Characterizing [HKAC14, MSN^{+}19]. Charm [BBK11]. CHARMM [NCWD^{+}04]. Chasm [RSSM06]. chatbots [ML19b]. Chay [KCD19]. Chebyshev [LWW^{+}19]. check [LDZ^{+}15, vRGNP09, LCC^{+}03]. Checking [PNB04, BCCM16, CAC^{+}08, Guo19, HFF07, LCC^{+}03, MK12, PAdS^{+}17, SZR16, YGL05]. checkpoint [AG17a, Jon09, PGB03, BDB^{+}13].

Checkpoint-on-Failure [BDB^{+}13]. Checkpointing [LX08, dCGKG06, ALYD17, BBB^{+}14, GÖ18, KAL07, MJ11, MB18, RMG^{+}10, SBS19, SGV12, SK18, YCW08, ZWL^{+}19]. checkpointing-enabled [SGV12]. checkpointing/recovery [MB18]. checks [LGFM05]. chemical [HPHB^{+}15, YYZH19]. cheminformatics [CBQ^{+}11]. Cherenkov [RVVPD^{+}17]. China [ZGRSC10, JW10, LMGZ19, MZS^{+}10, SLC^{+}18, SLC^{+}19, YQL^{+}15, ZZYW10, ZL19]. Chinese [HLX^{+}16, JHCH19, LGJ17, LMCL19, THQ19, ZQD^{+}17, ZCLL19]. Chip [PS19a, GGFPGB14, GA09, LLN^{+}14, MCP^{+}12, MST13, Puf13, RS12, SPS17, XLL^{+}15]. chip-multiprocessors [RS12]. chips [HTHW16, SSM04]. Chiron [ODS^{+}13]. choice [CHZ10, CHZ12, SSMB15, WBM^{+}10]. Cholesky [ZDG^{+}14]. choose [PLY13]. Choosing [BFU07]. chord [BKH08, CCG^{+}08]. chord-based [BKH08]. Chord-like [CCG^{+}08]. Choreography [Ley06, ZDC15]. chosen [LZC14]. chosen-ciphertext [LZC14]. chunking [STO17]. churn [WTN07]. CILogon [BFG14]. CIM [DLX^{+}16]. cipher [WYL14]. ciphers [LGP19, TQL^{+}14]. ciphertext [LFWS15, LZC14, WZC16, WLFX17]. ciphertext-policy [LFWS15, WZC16]. ciphertexts [ZWY^{+}19]. circuit [AMSR14, CKRO13, IKP19, MOK04]. circuit-switched [MOK04]. circuits [AMSR14, GLC^{+}04, Sin10]. circulation [RZL^{+}19]. citations [CZ19, DK19b]. citizen [HAvI13]. city [BKLH09, MCG18, XYS17, ZXW19, MZG19, WKL^{+}11, YYWQ19, ZP19]. City-based [MZG19]. civil [ZQMC19, HCBRM16]. claimed [WFS^{+}19]. clairvoyant [BCM15, dSGD14]. class [DP19, God12, HWR03, JOK^{+}18, KHL17b, LL18, LLT^{+}14, SGM18, SRF13]. CLASSe [MML^{+}17]. classes [Bac03, GG09, WMA07]. Classification [KBE07, DLJ15, DS19, DP19, God12, HYLG15, HHKA14, KTHA18, KT19b, LZL^{+}17b, LLJ18a, LLJ18b, LLY^{+}19b, LGQS12, LC18, LWP19, MPS11, MSM^{+}14, MJ19, PLZ14, Pla08, QXXZ16, RS16, SN16, SRR19, VJ19]. classified [CZL^{+}17]. classifiers [HZL^{+}16, LCM^{+}17]. Classifying [HPK^{+}18, ZCLL19]. classroom [GRGP12]. ClearSpeed [GSB^{+}12]. Client [Hic18, BYN^{+}17, FHH15, GGC19, PB07a, PRS01]. client-server [GGC19, PRS01]. Client-side [Hic18, FHH15]. clients [MWJ^{+}10]. climate [WJYH16, Zho06, ZBC^{+}07, ZDC^{+}09, ZCD^{+}12]. climbing [YT19]. clinical [CCLP19, KSM^{+}08a, KSM15]. cloaking [KHHC13]. Clock [BHH09, JK13, DCA17]. clocks [TAI^{+}11]. clone [LKKL16, ZWL^{+}17]. CLORIFI [LKKL16]. closed [BLDW16, LXYC17]. closed-form [BLDW16]. closer [MZK16]. Cloud [AOK19, CR13, CKL19, CPSP17, DWG19, EBMD13, GWC^{+}11, HSHT14, JM19, JRHJ16, LFPP17, LV12, PPC^{+}15, RCC17, SRAG16, TZLC15, VRMB13, WLFX17, Yu18, ZBE17, ZBEM18, ZZL^{+}18, Zhu18, AaBT16, AaBT17, ACC^{+}15, AMBT17a, AWR17, Air17, Akt18a, AG17a, AG17b, AJY^{+}15, ACG15, AMAB17, ANH^{+}19, BYN^{+}17, BTCGL17, BV16, BCX15, BHD13, BLXE19, BZD16, BWHS18, BXQ17, CSMB15, CMCAA17, CRB^{+}17, CCC^{+}16, CYD^{+}15, CEMR19, CLQ^{+}17, CSL12, CL13, CJZ^{+}15, CXPL15, CLH^{+}16, CLL^{+}19, CJZZ10, CZ15b, CVK15, CRV15, DIM18, DD16, DRS^{+}13, DM15, DS17, DZL^{+}17b, DSH18, DZZL19, DC19b, DXZ^{+}16, DCG15, EN19, EB18, ETR^{+}13, FHO^{+}15, FHH15, FCY17, FLYL16, FPC15, FGZ^{+}18, FTR15, GQH17, GQJL18, GRQ19b, GRQ19a, GCSB19, GMPT15, GSR^{+}19, GPP^{+}18, GCZ^{+}17, GWVP^{+}14, GÖ18, GAS19, HKS19, HKA19a, HAAWA^{+}16, HSJ^{+}18, HPD^{+}15, HZY^{+}19, HHPL16]. cloud [IHB15, IRB19, KKDS19, KC15, KSK19, KB17, KM13, KYM17, KTB17, KKK^{+}19, KCKC15, KBT^{+}14, KKT13, KZY^{+}18, KMRT18, KSK17, KG19, LZZ^{+}17, LWC12, LLLJ14, LLL15, LWL15, LWG^{+}15, LZY^{+}16, LZW^{+}17a, LHHJ18, LLT^{+}19a, LWW^{+}19, LS19b, LDXC13, LW13, LQL^{+}15, LBY^{+}16, LHLH16, LZBF17, LRS15, LSMVML15, MWPX17, MTGZ17, MYY18, MS13, MBMB18, MB12, MSST15, Mar19, MCP^{+}12, MK15a, MTT15, MB18, MW18, MPVT17, MMG^{+}18, MAK18, NR17, Nam19, PLY13, PYKL16, PRCV16, PRD^{+}13, PLW^{+}18, PWS19, PT12, PPR19, PRNM19, PBK19, Qia19, QGZL18, QCB17, RHRB13, RBP12, RBNG15, RSKK19, RBT18, RHuR^{+}19, dRRdCRR16, RB17, RLDZ13, RHS17, SVB19, SIST18, SR19b, STO17, SBC15, SN18, SZG^{+}19, SPJ14, SPSB19, SWP17, SKB^{+}17, SWW^{+}16, Soo16, SMFM18, SGL^{+}17, SKA^{+}14, SB19b, SA19, SCLK15, TZ16, THF15, TSL15, TLF17, TXZ^{+}17, TY15, TBTZ18]. cloud [TPV17, VŠC17, WBD^{+}19, WLW11, WL12, WYBS16, WLZ17, WSZ^{+}18, WLZ^{+}18, WZLQ16, WZL^{+}17b, WWG^{+}11, WS19a, WSWL12, WNN^{+}15, XRD^{+}17, Xha18, XS19, XBB13, XXX15, XWH^{+}17, XTB17, YSL^{+}15, YLD13, YXL17, YSC^{+}17, YZCT17, YT15, YBZ^{+}15, YCQF19, YCSY19, YYL^{+}12, ZYZ^{+}12, ZLN^{+}13, ZDC15, ZNT^{+}16, ZZ18, ZQD^{+}17, ZYZC17, ZRB19, ZFJ16, ZWH^{+}17, ZCH^{+}18, ZYLY18, ZZL^{+}19, dOOO^{+}12, dMd^{+}17, BB12, CR12, CMS17, EH18, ESG11, KBB11, KMG^{+}18, LCW^{+}17, MDH^{+}16, OKO18]. Cloud-aided [WLFX17]. cloud-assisted [MWPX17]. Cloud-based [HSHT14, CLL^{+}19, FPC15, GSR^{+}19, KKDS19, RSKK19, RHS17, TBTZ18, dOOO^{+}12]. Cloud-enabled [CKL19]. cloud-HPC [KMRT18]. cloud-integration [GMPT15]. cloud-of-things [CMCAA17]. cloud/fog [LWW^{+}19]. cloudlet [YBZ^{+}15, YBX^{+}17]. cloudlet-based [YBZ^{+}15]. CloudMe [TDC18]. CloudMon [LLL15]. clouds [BB15, CTAB16, CMS17, DXM^{+}17, EQW^{+}18, GVK12, GYP^{+}16, HM16, JMF09, KS19b, KMBR19, KOK14, KSPM12, LBdM^{+}16, LZW^{+}16, LFWS15, LGL16b, MDB^{+}17, MK15b, OKP16, PCGE18, PC17b, PB16, PRP^{+}15, SM11, SFCAV16, SYMA17, SK18, TPT^{+}18, sTzNL16, VGN^{+}16, WP12, XLT^{+}17, YLR^{+}13, YNX^{+}16, ZZ15, ZLH^{+}15, ZS17, ZBZ^{+}18, ZH15, WNN^{+}15, EMS15, HYQ17, SBP12]. CloudSim [VŠC17]. Club [SLT^{+}06]. Cluster [DMR^{+}07, ACC^{+}12, BYDC19, BBdS^{+}17, CWL03, CGGH17, DSO^{+}01, DT17, ELM^{+}16, FHO^{+}15, FJP^{+}05, FS18, HWZ^{+}15, HON04, JAA08, JCJ17, Jon09, KF01, KSC12, LGL16a, LXRJ13, MG09a, MFG^{+}13, MSS16, NO02, PDY14, RPK08, SF16, SLM^{+}10, WJYH16, WDT18, XLQL18, YCL11, YK10, ZP07, ZYZ06, EEK^{+}04]. cluster-based [CWL03]. cluster-computing [ELM^{+}16]. clustered [DMA13, GSG06, LR05, LYL07, ZACG16]. Clusterfile [IT03]. Clustering [BDY03, KT19b, KT19c, ASE^{+}17, DRS^{+}13, DLX^{+}16, HKA19a, HW16, Jun16, KOO12, KPS14, LWSZ19, LWY15, LHZX19, MT19b, NJZZ19, OZI19, PB19b, PCS^{+}12, SS15a, SIS19, SS17b, TLX^{+}17, YSW19, YZXW17, ZMYA18]. Clustering-based [KT19b, OZI19]. clusters [AD02, BDL^{+}15, BB04, BdL06, BHBD13, CZL^{+}17, CFP^{+}03, DWC^{+}15, GKSR14, HR06, ITK09, IQOvdG13, KB13, LJL^{+}17, LSP15, LKYS04, ML19a, MOF15, MJD17, MVML11, MWLS11, OLG^{+}15, PSRR14, PKB03, PSS^{+}18, PSJM13, RMP13b, RCLSK16, RHL^{+}18, SV09, SdSL18, SS15c, TKZQ17, TFG^{+}12, TMAG03, WLW14, XDJL18, YWC11, ZSZ15, ZCL^{+}18]. CMOS [GZHF19, NK19]. CMS [FMS15]. CNN [He19, HCS18]. CNN-based [HCS18]. Co [KT19c, CCSS10, CCLP19, Jon09, ME08, VDB09, YYC10, ZCZ^{+}19, LMS18, MDV07]. co-allocating [ME08]. co-allocation [CCSS10, Jon09, VDB09, YYC10]. Co-Array [MDV07]. Co-Clustering [KT19c]. co-scheduling [ZCZ^{+}19]. co-training [CCLP19]. coalesce [NRR15]. Coalition [CEM^{+}17, BCdlCT06, HAAWA^{+}16]. coallocation [ET09]. COAP [KG19]. coarse [CDA09, ID18, QML^{+}17, Yos06]. coarse-grain [Yos06]. coarse-grained [CDA09, ID18]. coarse-to-fine [QML^{+}17]. coastal [ABC^{+}08a, ZDA^{+}07]. Code [vWAH^{+}02, AAP13, BFM^{+}10, CSC^{+}17, CC15, CXW17, DVD^{+}12, DZJ^{+}15, FED03, FNBS16, FBV^{+}13, GE06, GAB19, GGC19, JCVU15, JFI^{+}08, KB06, LBFS17, LKKL16, LSH^{+}16, MLVB05, MK12, Ogi02, PSW11, RLRG15, SPH13, TLM17, VDL^{+}15]. code-based [GAB19]. code-bones [CSC^{+}17]. code-copying [GE06]. codec [ZJT^{+}19]. coded [XWPM19]. CoderLabs [TBTZ18]. Codes [IC19, AAF^{+}07, GBFP09, HR18, HTWW19, KBG^{+}09, LDZ^{+}15, LLH^{+}18, Nev17, PWMX16, PWMX17, PHCR09, RCB03, RMCHMG15, SCLK15, Tan15, ZDB^{+}14]. CoDIMS [PGO^{+}04]. coding [AZF^{+}12, CSWB11, KCS07, PAM^{+}15, Tan15, XGC19, ZSL^{+}15]. coefficient [KKW^{+}14, WWY19]. coefficients [NDT^{+}16, XYER16, ZZ14]. coevolution [BPL12, WJT^{+}14]. coevolutionary [LZBF17]. CoG [PTCN07]. Cognitive [LSY^{+}12, DLJ15, DK19a, FXX16, LCMY13, OKO18, OK18, TZYL13, WRLS12, WRDZ13, YCZ^{+}13, ZZY^{+}19, ZLQ^{+}18]. Cogset [VJK13]. coherence [SC07b]. coherent [CS17, PRU14]. COIL [OOTK01]. Coimbatore [PB19b]. COISA [AMB^{+}17]. Cold [LWC12, TWW^{+}19]. cold-start [TWW^{+}19]. Collaborating [LWQS19, XMJ17]. Collaboration [SNM15, AHB^{+}10, ACF^{+}07, CWZL13, DCEK15, DT15a, FWU^{+}04, GRSB09, JSG17, Kri13, LLD19, LZW^{+}17a, PB19a, TBTZ18, VAC^{+}07, YLEB14, ZM13]. collaborations [NRW04, TMS^{+}12]. Collaborative [LF15, AMSS15, ACFT15, BHA^{+}15b, CBHTE11, CH04, CPE^{+}19, DFG^{+}18b, Dra15, DR15, EMB11, LH17, LGD15, MST15, NAP^{+}07, PBD^{+}15, PML^{+}05, SHG^{+}07, SAM^{+}17, TZLC15, TSL15, VLW19, WBHW08, WTY^{+}19, WW08, Xha18, YAG19, ZDA^{+}07, SPR^{+}07]. collaboratory [MP02, ZSL^{+}10, vLRF^{+}02]. collapse [CCS14]. collapsible [ZCW^{+}18]. collecting [BDL^{+}15]. collection [ANPR16, BCK^{+}09, BDZ19, BML08, HM03, KSN16, TC12]. collection-oriented [BML08]. Collections [Tan12]. Collective [CHPvdG07, LCFkL05, FA18, HKRR08, NMKB03, SSZ13, WZL^{+}17a]. collectives [RTET15]. collector [JM19, Kal11, Puf13]. collectors [AP06]. collision [TZYL13]. collisional [RMCHMG15]. collusion [ZZL^{+}19]. collusion-resisting [ZZL^{+}19]. colocation [BPT^{+}16]. colony [KC15, MS18, PN19, TV14, YZ19]. color [KDC17, LYL^{+}19, VYKM19]. colored [PP17]. coloring [AMZ19, CLF^{+}17]. colour [KL19]. column [Has17]. combination [CDF^{+}17, ECP18, GFL04, LS19b, WGQ^{+}18, ZHW^{+}16]. Combinatorial [MP17, SCS17a]. combined [GHLS19, JyLdZ^{+}18, KKGO04, LWY15, ZGL19]. Combining [MGI17, PAM^{+}15, AMS17, THF15, THQ19, VJHB05]. command [WKL14]. command-line [WKL14]. comments [LWP19, THQ19]. commerce [DJ19, PRS16]. commercial [YQL^{+}15]. commitment [BF07]. Commodity [MTD^{+}02, PvLV^{+}02, vLGL^{+}02, RVRD10, SPMP11, vLFGL01]. Common [WD07, LF17, PBF15, WLR05, Zho06]. Communicating [May10]. Communication [ABF^{+}17, CDN15, DWC^{+}15, GPVCdBRO12, NRR15, SRM^{+}15, ALYD17, AD02, BBPV05, BBCG02, BBK11, CHPvdG07, CS17, DBR13, DPST06, DBB^{+}16, DXHL17, ETR^{+}15, GKPT13, GRS06, HKRR08, HLHC12, JKD19, JKZ03, LCA^{+}19, LZW13, LCFkL05, MG09b, MPT07, MMSN^{+}01, NMKB03, NS19, PH12, PCL17, QXJS17, RIFR10, SJ19, TTD^{+}11, TNH15, TNI16, TKA^{+}02, TV14, TW07, TAI^{+}11, UGM18, VP19, YJZZ19, ZJKL10, ZLA^{+}15, ZWMT12]. Communication-aware [CDN15]. Communication-hiding [DWC^{+}15]. communication-induced [ALYD17]. Communications [Ano15a, MLY10, MN10, NPM19, BC16, BBW19, CLH13, DZC16, FA18, FK19, FN13, Fox12, MRL16, MFGE19, RTET15, SCGZ19, SSZ13, ZHM^{+}17]. communities [EMB11, JJGL13, JWY^{+}17, MRY^{+}16, MZA19, PFC^{+}09, RWK^{+}02, WYY^{+}19, ZX09]. Community [WD07, vLRF^{+}02, BSC^{+}15, BDI^{+}07, CLL^{+}18, CWYX17, HLF^{+}17, HYL^{+}19, LZWD^{+}15, SFCAV16, SCLL19, WDW^{+}15, WBB^{+}07, Yu18, ZSWS18, ZXX17]. community-based [HLF^{+}17]. commuting [AKG13]. compact [AMB^{+}17, IHA^{+}15, PPR19, PZ17]. company [YWLQ18]. Comparative [SSE19, ZXW19, EQW^{+}18, HPVRPF14, LLN^{+}14, LL18, NJ15, PBF15, SAOKM04, SBC15, SCBH09, VJ19]. compare [DHM14, KW01]. compare-and-swap [DHM14]. Comparing [GSB^{+}12, KF01, KSG11, LKJ03, SHT^{+}17, NTK08]. Comparison [GPS^{+}07, LF17, PH12, BB13, CKD^{+}19, CEB^{+}18, Dra15, Fer13, dFdOSR^{+}19, GPW05, KIM^{+}03, KHW05, KR04, KC06, LCM^{+}17, MKIO04, PGB03, PSHL11, PRS01, RBDI17, SM02, SZT18, SCRV11, SSK11, SSV19, TMF^{+}10, TZ16, ZKA07]. comparison-based [PSHL11]. comparisons [BKZ^{+}13, CLYC16]. compass [RGCC15]. Compatibility [SM02, HLZD18, IÁBE11]. Compensation [SE01, DYY^{+}19, TLM17]. competition [MME13, XTZ10]. competitive [KZY^{+}18, MSD^{+}18, TBK06]. compilation [AC09, KKGO04, MGI17]. Compiler [CCC12b, KL12b, RMCA12, TJF14, TXY^{+}16, WMvP^{+}09, BAG17, FE18, LHC^{+}07, LYL07, MCAB^{+}02, MLP04, RMG^{+}10]. compiler-assisted [RMG^{+}10]. Compilers [Kni06, TFDA07, GE06, MSP^{+}19, SM02, YTF^{+}01, KB12]. Compiling [BB04, BK05, CCO^{+}15b, NNON02]. completeness [ZX11]. completion [NNK^{+}07]. Complex [BCD^{+}10, BFM^{+}10, BDY02, CLL^{+}18, Cog04, CMD17, DJM12, DXHL17, JHCH19, Kar14a, KSS^{+}17, LHL10, RW10, RHBK11, SGR19, Wan18b, ZLAa^{+}17]. complexities [MvWvM^{+}17]. complexity [CL14, FAM^{+}18, HZHP09, HK02, NFG19, OKM10]. compliance [FGG^{+}18]. compliant [dFMSPSW06, XRD^{+}17]. Component [Ber07, JPWH02, JSS07, SBBE07, Zho06, ALKD16, AKM^{+}06, CGB^{+}06, DGR^{+}07, EJF^{+}16, HML20, JLCA07, LLH19, MST^{+}05, MB12, PFC14, PPC^{+}15, PB16, SARL13, SVN12, TMR^{+}07, VDPC03, WK07, XWFH08]. Component-based [JPWH02, JSS07, CGB^{+}06, DGR^{+}07, MB12, PPC^{+}15, PB16, SARL13, VDPC03, WK07, XWFH08]. Component-oriented [SBBE07, JLCA07]. Components [JZZL06, AFR09, BN19, CGGH17, GBSHA01, KL12a, LSW07, PXY^{+}07, QH10, RE03, RBO^{+}02, SPLO06, SGD^{+}18, SLB08, YBC^{+}07]. composable [YL16]. Composing [EABVGV14]. composite [CEMR19, GYM14, RCKV12, ZLY^{+}13]. composition [CHH18, DCP^{+}17, EAGVBVDS11, KL12a, NPTT06, Pre01, RG18, XDL^{+}11, XWD^{+}12, YLD13]. Compositional [DAB09b, MSS16, RG18]. compositions [LLX^{+}15a, PPBB14, SGD15]. compound [SAP16]. Comprehensive [BDP^{+}14, MRP^{+}18, WKZL19, XLHT17, JAA08, PGC^{+}19, VSK17, dP06, dMd^{+}17]. Compressible [QWZ^{+}19, WJLD09]. compression [CMMS17, CXC^{+}18, CS13, FNI17, LSE^{+}13, UMD^{+}13]. Compressive [JyLdZ^{+}18, CL19, PHY^{+}18]. Compromised [VS19]. Computacionais [DC19a]. Computation [Du18c, FH01, TH10, ABDP15, BP17, CP14, DFTHD18, ETR^{+}13, EJD15, FLMRC02, GSV03, HZHP09, JVPI18, LRLY17, LG08, LLRS19, LSP15, LPA^{+}08, LL19c, MB16, NRR15, PSCK^{+}15, PXY^{+}07, Riz04, TWB13, WLWX14, WLWX16, WSRM12, YSQM19, ZP07, ZZ14, ZKJ^{+}07, ZZL^{+}17a, ZZL^{+}17b, vRKS03, Ano06]. computation-offloading [JVPI18]. Computational [BA04, DDE^{+}12, HBH02, Mar05, Qiu11, QFG14, QFT14, RBBH02, TCDMR^{+}17, vdS06b, AUHWJ19, BFM^{+}06, BPD06, CAD^{+}18, CNKJ18, CKC09, CCP^{+}15, CDP17, DBR13, DS07, DMD16, FAM^{+}18, FP02, FMS08, KV12, KBG^{+}09, KKWZ15, LMH^{+}14, MP02, MAdS^{+}10, MPSGD14, MTVF14, MD02, NAP^{+}07, PW12, PSG03, PB07a, PYF02, PCC17, PV15, RCB03, SNH15, Sha15, SR17, TP14, TRH^{+}02, TV14, VDL^{+}15, WGZL06, YHK09, YLC11, YTL19, ZSL^{+}10, vHKT^{+}11, vdS06a, GTGT11]. computationally [GPV09, RMCN^{+}07]. computations [BCI^{+}09, DIK14, DKJ13, EFY17, GGV14, GDMT^{+}12, GEBA17, GS04b, KFD18, LSXL17, LWW^{+}19, MCP^{+}12, MRS08, NNH^{+}14, NPL19, NDL17, OCC^{+}05, RMCA12, Rav16, RLMG16, RCA^{+}12, RPRG17, SAP16]. Compute [MDH^{+}16, BAC^{+}15, MÖO17, SKNH09, VLMPS^{+}18, ZWW14]. Computer [BM04, GAM17, Nel05, SNM15, Wu18, ZCW^{+}18, AKW04, BHJ^{+}16, CPG^{+}16, Che18, CPXA06, DMW^{+}10, DCA17, FBV^{+}17, FJG^{+}13, GQ04, He19, LGdVH13, LHC14, LWL15, MCP^{+}12, MO02b, NSSAK13, NSSAK16, PSJM13, RGAK15, SRM13a, SS18, WAD12, WCZ^{+}18, ZDC^{+}09, HF17]. computer-aided [LGdVH13]. computer/digital [LHC14]. Computers [Kni06, TFDA07, BCM^{+}07, BCC^{+}05, DSO^{+}01, ON01, ON02, PCVZ^{+}04, RVRD10, RMCN^{+}07, RSTV05, RLRG15, SSK11, SS15c, VdSK^{+}05]. Computing [ACF^{+}07, Ano15a, ACD02, BJ18, Ber07, BRCV16, CR13, CM07b, FZ07, GM10, GPPR17, IBvA^{+}02, JX06, KB12, LV12, MLY10, NPM19, PHGK10, PW05, RR11, SN06, SCNH07, STS19, SFN12, Tho07, VC16, ZYH09, ZQH12, AOK19, AaBT16, Akt18a, AML^{+}15, AdSCdR^{+}19, AMGCC17, AJY^{+}15, ADF^{+}13, Ano06, ATİ17, AKM^{+}06, ABG^{+}05, ANH^{+}19, BYN^{+}17, BGGS14, BTCGL17, BFR05, BHM^{+}12, BCX15, BCD^{+}10, BLXE19, BHQOS15, BZD16, BBM19, BKM^{+}07b, BDG^{+}10, BPT^{+}16, BWW^{+}08, BXQ17, BHKW12, BPD06, BAGS02, BM02, CLW^{+}19, CGBNM17, CLQ^{+}17, CL13, CJZZ10, CZ15b, CLS14, CAG^{+}13, CL07, CMB06, CT16, CN02, CPSP17, CBP^{+}04, CGB^{+}06, COdO^{+}11, CMD17, Dab09a, DN19, Dam11, DD16, DC19a, DRS^{+}13, DED07, DWC09, DM15, DS17, DSH18, DWG19, DC19b, DKJ13, DCG15, EB18, ELM^{+}16, EDBS08, Erw02, ETR^{+}13]. computing [EQORS19, FJP^{+}05, FJ05, FMS11, FMT16, Fox12, FB16, GFBR10, GQH17, GRQ19b, GKSR14, GKG^{+}04, GAM17, GBMM15, GS04a, GWC^{+}11, GLD17, GVP^{+}14, HKS19, HSM14, HSJ^{+}18, HQoS11, HWR03, HLC19, IHB15, JRHJ16, JCK^{+}13, JLL18, JPWH02, JK13, KDC17, KC15, KMBR19, KBB17, KB17, KM13, KMJ^{+}17, KYM17, KSM^{+}08a, KKT13, Kri05, LLMK18, LGLA15, LBV16, LWC12, LLLJ14, LL15, LSS15, LDXC13, LLH^{+}15, LQL^{+}15, LBY^{+}16, LZBF17, LAL02, LMOT10, MTGZ17, MZG19, MHJH16, MAS16, MS17a, MS17b, MS18, MB12, MK15a, MDH^{+}16, MM17, MSB17, MB18, MJD15, MJS19, MMG^{+}18, MM10, MZK16, MGR02, NR17, Nam19, NNK^{+}07, NC05, NJ05, OISS07, PW12, PYKL16, PLC^{+}19, PRD^{+}13, PIAH12, PSS^{+}19, PC14, PRC^{+}14, PK17, PT12, PBK19, QZDJ16, QGZL18, QLL10, RRBB11, RHRB13, RVD^{+}12, RBP12, RBNG15, RSSM06, RHZ^{+}17]. computing [RCA^{+}12, RBB^{+}09, RNAD19, RB17, RLC16, RCLSK16, SRS16, SM04, SL10, SG16, SBC15, lSsCY17, SQS^{+}19, SRAG16, SFH13, SFT15, Soo16, SRL^{+}14, SWHL16, SS07, SNGR18, SAM^{+}17, TTD^{+}11, TSA^{+}19, TZ16, TKZQ17, TWSM05, TTL05, TTPJ16, TY15, TSKM18, TSS18, TBH^{+}18, UA18, VD05, VFAD18, WKB^{+}19, WLLL16, WZ04, WCA08, WYBS16, WYZ^{+}17, WFJ^{+}17, WLZ17, WML^{+}19, WWY19, WWC^{+}19, WZLQ16, WSW^{+}12, WS19b, XCL09, XPWF15, XADLC15, XLYL17, XS19, XBB13, XXX15, XBM14, XYS17, XWH^{+}17, XTB17, YCL11, YDB^{+}13, YBX^{+}17, YLEB14, ZBE17, ZBEM18, ZH08, ZZ16, ZZC^{+}17, ZWT^{+}18, ZP19, ZYY^{+}19, ZQD^{+}17, ZYZC17, ZRB19, ZKJ^{+}07, ZCD^{+}12, ZYH12, ZZ17, ZDR^{+}18, ZCH^{+}18, ZYLY18, ZFWJ19, ZXXN06, ZWW14, ZJS11, dAAVS12, dARP17, dMd^{+}17, BM12, GJ17, GZX17, SANB08, WLL03a, WLL03b]. computing-based [DWG19, MS17a, MS18]. conception [PBD^{+}15]. Concepts [DMW^{+}10, SP16, Sch04]. Concurrency [Ano06, FH01, TH10, BVGVEA11, BMS^{+}09, BT04, CAC^{+}08, CM02, FR02, HL06, Hoa10, LSW07, LWL^{+}19, TRW07, TDM^{+}19, WJH06, dCHMJ12]. Concurrent [AFGL09, BHM^{+}12, BH05, KFD18, SW09, Tan12, AKG13, ACGG06, ABS16, ADK^{+}17, BL04, CL10, CGIP16, CBB^{+}19, DZM^{+}15, GM04, IR11, JK10, Kar14a, Kar14b, KIM^{+}03, Kuł14, LPSF11, LDPZ14, LSW07, MKIO04, MISV13, MS05, MCC16, NRR15, RCKV12, SSZ14, WZZL13, WCC04, WO14, WL11b, YKA^{+}19]. Condensed [BIK^{+}11]. condition [IR11, SWLJ17]. conditional [FBS16, SKB^{+}17]. conditions [LBDS15, LFG05, MFGE19, SWU08]. Condor [TTL05, LTM^{+}14]. cone [ZGZ19]. cone-beam [ZGZ19]. Conference [AF14, Che19, Du18c, FZ08, LB19, WDM14, Wu18, ZC19, CL13, DR15, GWD15, PDD14, PCC17, UA18, WDGK15, WT15, Fox01, Fox05, HF17]. confidentiality [XBW^{+}15]. Configurable [SRF13, CGB^{+}06, GKPT13, WZ04, YDL09]. configuration [AMVOSGAC17, BDP18, CKRO13, GBSHA01, KKTHL13, KAM11, YYXL19]. configurations [PTL^{+}16]. configured [STWSP12]. Configuring [ERZ^{+}11, GCSB19]. confinement [PNB04]. conflict [BAS07]. conflict-free [BAS07]. conflicts [HDX^{+}17]. conformal [QSZL18]. Conformance [SKR17]. Confuga [DT17]. Congestion [LLC^{+}15a, WDW^{+}15, ALL^{+}15, BVIB19, LWSZ19, WMA07, YESG^{+}17b, YESG^{+}17a]. Congestion-aware [WDW^{+}15]. Congestion-free [LLC^{+}15a]. Conjugate [JWW17, MDL^{+}10, SK09, SSK11]. connected [BF07, MRY^{+}16, MvWvM^{+}17]. Connecting [MH07, BSP11]. connectionist [YTL19]. connectivity [CNPP09]. conquer [CCW06, NDL17, YA04, ZLT^{+}16]. consciousness [LLYL09]. consensus [BFG01]. conservative [BGdCCA11, DVB14]. consideration [TSA^{+}19, XBW^{+}15]. considerations [FA18, KBH^{+}15b]. considering [LWS19, MS17b, TYHL12, XWX^{+}19]. Consistency [OCS01, ADM06, ANTZ09, CY07, GKPT13, HWY^{+}17, VSK17, WNT02]. consistent [PQP13]. consolidation [ACG15, ACG17, AMAB17, BB12, BB15, IRB19, LBdM^{+}16, ZBZ^{+}18]. constant [ZWY^{+}19]. constant-size [ZWY^{+}19]. Constrained [XZT^{+}11, EQW^{+}18, JAU19, KSR14, LLT09, LWYZ19, LZBF17, MHLC^{+}05, PCGE18, QW17, ZWXS19, ZH15, ZFWJ19, ZLA^{+}15]. Constraint [YSW19, DAC12, GAE^{+}06, LWFL14, LGL16b, LNCY11, MSB17, RC09, SMH^{+}19, SKK02, TLF17]. constraint-based [DAC12]. constraints [ACG15, ACG17, AAE^{+}09, CY07, Cuz11, Hun15, KS19b, KZY15, LLG^{+}15, MS05, RIWS17, SAMS19, TKK^{+}11, TCDMR^{+}17, XXLL17]. construct [CCCC06, zGWXT09]. Constructing [AVS^{+}19, WKL^{+}11, ZlC15, CLL14, KRW17, RRR15, WCR^{+}14, XZJ11, ZM13, ZZ11]. construction [GCO^{+}14, IKP19, KJ19a, LFZ^{+}17, LCW^{+}17, SBBE07, SN16, WXY10, WBO16, YWL^{+}17a, YLR^{+}13, YYWQ19]. consumer [CZJY19, LSZ19]. consuming [ZQD^{+}17]. consumption [ADI^{+}14, ADMQO14, CLH19, FMT16, GYP^{+}16, HLB10, NSSAK13, NSSAK16, RR15, dRRdCRR16, SB18, XXLL17]. contact [XM02]. container [BTCGL17, BPdM06, GKP^{+}19]. container-based [BTCGL17]. containers [EK19, MAK18, WJKS18, YYZH19]. Contaminant [YGW17]. contaminants [VLF^{+}13]. Contemporary [SNM15]. Content [Zic12, BM10, CCK^{+}17, CHZ10, CHZ12, CDF^{+}17, JQSP08, LNKZ08, MWPL15, MWPX17, PF12, PZZ08, PZZ10, RSPV17, SGSC08, TSBR10, YQL^{+}15, ZW09]. content-based [JQSP08, MWPL15, TSBR10]. contention [BBK11, BDH18, DHM14, WYQ^{+}13, XCL09]. Context [And13, BKS18, CAC15, IHA^{+}15, CMT13, DHC11, DCFC08, HPS05, KR15, LMCL19, LS05, LCMY13, PAdS^{+}17, Sod05, ZLY^{+}13, ZJT^{+}19, ZDC^{+}09]. context- [DCFC08]. Context-aware [And13, BKS18, IHA^{+}15, DHC11, ZLY^{+}13]. Context-awareness [CAC15]. context-bounded [PAdS^{+}17]. context-free [LS05]. contexts [DPST06]. Contextual [GAE^{+}06, KM13, TWW^{+}19, PPdSTB17]. contextualization [PLJ18]. contiguous [PMAL14]. continuity [CN16]. Continuous [DDM16, LWYM16, LCT16, TMP16, TB12, TSKM18, ZSZ^{+}14]. contourlet [PJW^{+}14]. contours [PLL14]. contract [HLC19]. contracts [BWEB14, Dam11]. Control [CN16, IÁBE11, NSSAK13, NSSAK16, AFGL09, ALZR11, ALL^{+}15, ACG15, AFG^{+}05, BCD^{+}02, BEQOR13, BMS^{+}09, BT04, BHA^{+}15b, BJC17, CSL08, CJC^{+}18, CLH^{+}16, DMA13, DZ13, FJ05, FR02, GBSHA01, IS10, JSG17, KKDS19, Kar16, KKK^{+}19, KN19, LM08, LXP^{+}12, LZL^{+}17b, LCW^{+}19, Li19, LLLS18, LWL^{+}19, MLL^{+}11, MLG15, MTGZ17, MABP13, QLZX19, SARL13, Sin10, SW11, SW12, TL19, TJF14, TDM^{+}19, sTzNL16, TBK^{+}15, WMA07, WLW11, WYW^{+}17, XCL09, XHH12, YBO10, ZCC^{+}06, ZYN^{+}07, ZZ18, ZJT^{+}19, ZXW16b, ZCXH17, ZQR^{+}19, ZZL^{+}18]. control-based [LM08]. controlled [KBB17, RCB^{+}04, TV14, ZMZD11]. controller [ACG17, EME19, LWW06, NIIU17]. controllers [ZW17, ZTGW17]. Controlling [CHP17, dSGD14, dRC10]. convection [JN03, MO02a]. convection-dominated [JN03]. converge [WYQ^{+}13]. converge-cast [WYQ^{+}13]. convergence [CGBNM17, KMRT18, LZW^{+}16, WLP^{+}17, XWX^{+}19]. conversational [ML19b]. conversations [PCL17]. converter [LCM12, lSsCY17]. convertible [XWXC14]. conveying [MG09b]. convolution [LCW^{+}19, PDC16]. convolutional [CGMJ^{+}19, LLT19b, LLZ^{+}17b, QSX^{+}17, Wan18a, YLZ18, ZLW^{+}18]. Cooperation [Ano02, PRT09, KOOB15, MZW^{+}16, OKBO19]. Cooperative [GdMK^{+}18, GCL08, HK07, HJTX17, IOOH12, JX06, JLL18, QLS13, RIP18, SK17, Bou06, BDP18, CPB07, CWL03, DA15, EZJ^{+}18, FXX16, KT19a, KIM^{+}03, KKS12, MKIO04, SE01, WLLL16, WBZ10, XZH^{+}17, ZZY^{+}19]. coopetition [CDH^{+}15]. Coordinated [NB12, YZR14, Sod05]. Coordinating [CSL08, RE03, Pun01, UR04, YLLZ09]. Coordination [CCT15, OM06b, BHBD13, CW11b, LLLyL16, MZ06, OM06a, SNB^{+}01, TCH^{+}13, YLL^{+}18]. coprocessor [DWC^{+}15]. coprocessors [SSE19]. copying [GE06, HM03]. CORBA [BMV03, DPP03, MMSN^{+}01, NDP^{+}05, OSK^{+}01, PvLV^{+}02, SNB^{+}01]. Core [ZQH12, AYN^{+}14, ART14, AMTM17, ABC^{+}15, AAW^{+}02, ACCM17, AMZ19, BGGL07, BHBD13, BUVS10, BRCV16, CPEA18, CLH^{+}11, CZG16, CZL^{+}17, CLF^{+}19, CS17, CZL12, CLRB15, DLZ16, DLK^{+}18, ELM^{+}16, GLM^{+}16, GdMK^{+}18, HTHW16, HKAC14, HvNJB15, HFR^{+}17, IZXM09, JSR19b, JPS17, KSG11, KW19b, LGLA15, LL16a, LQL^{+}09, MGBC16, MCP^{+}12, MM17, MSB17, OAS^{+}15, PZ11, RMP^{+}13a, RLMG16, RHBK11, SPW09, SPQ^{+}17, STL^{+}15, SEF^{+}14, TYL^{+}15, VDL^{+}15, WJYH16, XZ09, ZYH09, ZYH12, ZZL^{+}17a]. cores [BKSM^{+}15, DXZ^{+}16, GPPR17, HCD^{+}18, HT15, ZQZ^{+}16, ZDR^{+}18]. Cori [DAC^{+}18, HCD^{+}18]. corpus [ZLQ^{+}18]. corpuses [SMM18]. Correct [CT16]. correction [LLZ^{+}17b, ZYLT06]. correctness [ASS08]. Correlated [BHBD13, MOK04, PHY^{+}18]. correlation [CHP17, cDrLyC^{+}19, KKW^{+}14, LWQS19, LWS19, PWJ10, RCM12, XHH12, XYER16, ZLY^{+}13]. corrosion [cWsThY19]. cortex [CP14]. CoS [KSPM12]. coscheduling [DRS^{+}13, Sod05]. cosine [ZSWS18, ZHC^{+}18]. cosmological [WDG^{+}14]. Cost [EQW^{+}18, ESGQ^{+}11, HKS19, HSL19, LWYZ19, LSMVML15, PC14, SL10, Akt18b, AMS17, BCF12, BXLJ16, CHP17, GWVP^{+}14, HLHC12, LGY17, LLT^{+}19a, MS13, QWZ^{+}19, SGCA^{+}16, WSL15, YDL09, YBX^{+}17, Zen19b, ZLN^{+}13, ZLZ15, ZFJ16, ZQW^{+}17]. cost-aware [AMS17, WSL15]. cost-benefit [Zen19b]. cost-driven [LGY17]. Cost-effective [ESGQ^{+}11, HKS19, ZLN^{+}13]. Cost-efficient [LWYZ19, ZFJ16]. Cost-minimizing [HSL19]. count [KVGH11]. counter [LPC^{+}14]. counter-based [LPC^{+}14]. countermeasure [ZQD^{+}17]. countermeasures [AAI12]. counting [AP06, WLF19]. country [KSB^{+}19]. country-scale [KSB^{+}19]. County [WYY^{+}19]. coupled [AFG^{+}05, HC07, JK06, LC17, SV09, VDPC03, ZJS^{+}17]. Coupling [AFR09, Zho06, AvdADtH09, Boe12, FRB^{+}06, ISS^{+}02, ZDB^{+}14]. course [ZL06]. Courses [GBB^{+}15, LMH^{+}14]. Cover [CSL^{+}19]. coverage [XY17]. covert [QXJS17]. CPC [Kni06, KB12]. CPPC [RMG^{+}10]. CPSocio [ZX11]. CPSocio-SLN [ZX11]. CPSs [SZMF19]. CPU [ACG17, BEQOR13, CSMS^{+}19, DXZ^{+}16, EOD^{+}19, FTT15, GGV14, Has17, HLCW15, Kar16, LLMM19, LDZ14b, MJD17, Mit17b, PDY14, PRG15, SD15, VLJ17, VLW19, WLLL16, WDG^{+}14, WJYH16, WWLD18, XMJ17, ZDX12]. CPU-cores [DXZ^{+}16]. CPU-MIC [MJD17]. CPU/graphics [GGV14]. CPU/multi [SAP16]. CPUs [JdM12, LC17, RCLSK16, SEF^{+}14]. crawler [DH13, GDJ16, CMS17]. crawling [PZZ08, PZZ10]. CRAY [PSG03, BS04, BB13, BÇG14, BWHS18, BBW19, Cla18, CKD^{+}19, DAC^{+}18, HCD^{+}18, Hic18, JKL19, KSM^{+}19, KB18, KAMB19, LKJ03, LSK04, MWRK18, MH18, TH19]. CrayPat [KAMB19]. Creating [CDH^{+}15, CS15, DEF08, OGA^{+}06, RBO^{+}02]. creation [PLY13]. credibility [AAQAR^{+}17, ZW09]. credit [SR19c, XDP18]. criteria [KS19b, KSPM12, SVS^{+}08, WJ12]. criterion [MLWA19, TJ17b]. Critical [HL13, WK12, FAPC16, HAN19, LL10, MWPX17, QML^{+}17, RS12, SDH^{+}17, ZQW^{+}17, LWC17]. crop [PB19b]. Cross [GRSB09, HM16, WRLS12, YLL^{+}18, ZBC^{+}07, ASE^{+}17, CCLP19, DCJ14, ET09, GW15, HKA^{+}15, LPG^{+}14, MYS19, MD02, XZJ11, YLD13, ZDC15]. cross-architecture [GW15]. cross-cloud [YLD13, ZDC15]. Cross-clouds [HM16]. cross-currency [DCJ14]. cross-disease [CCLP19]. Cross-domain [GRSB09, ASE^{+}17, LPG^{+}14]. cross-Grid [ET09]. Cross-layer [WRLS12, YLL^{+}18, HKA^{+}15]. cross-media [MYS19]. Cross-organization [ZBC^{+}07]. cross-platform [MD02]. cross-realm [XZJ11]. crossbars [LLN^{+}14]. crossed [WLQL16]. crossing [CZQ17]. crowd [LLL^{+}19]. crowdsensing [SWLJ17]. crowdsourcing [LYF^{+}17, WJJM17, XZH^{+}16, ZXRZ19]. crowdturfing [LNBL17]. Crunching [GTL06]. cryptanalysis [WYL14]. crypto [CLH^{+}11]. crypto-core [CLH^{+}11]. cryptographic [ABDP15, BYDC19, OO18, QZDJ16]. cryptographically [HJM^{+}11]. cryptography [BOB13, BBB16, GAB19, NLYZ12, OK18, OTO18]. cryptography-based [BOB13]. cryptosystem [sCR19]. crystalline [XBB13]. CSC [LXP^{+}12]. CSE [DZZL19]. CSE2015 [PCC17]. CSFS [HYX05]. CSI300 [ZHX^{+}19]. CSP [MS10]. CSS [WCWB19]. CTL [BCCM16]. Cube [EJD15, WLQL16, AS15]. cubic [PMAL14]. cuckoo [sCR19, LS19b]. CUDA [BY12, BAG17, CLYC16, DCD^{+}14, ER12, FJZ^{+}14, GWVP^{+}14, HP11, HLO^{+}16, KVGH11, KPS14, MMO^{+}16, PAdS^{+}17, PGdCJ^{+}18, PSHL11, PSV19, TNIB17, TVCB18, VLMPS^{+}18, ZZZ^{+}15]. CUDA-enabled [PSV19]. CUDA-quicksort [MMO^{+}16]. CUDA-streams [TVCB18]. CUG [MH18, TH19]. cultivating [HAN19]. Cultural [PCJ17, GIL17, PC17a, YGW17]. cumulative [CH04]. curbing [LNBL17]. currency [DCJ14]. Current [TFDA07, Dik07, EDB^{+}14, GKSR14, HFR^{+}17, MG09a]. curve [BBB16, LBH07]. custom [BA18, PZ11]. Customer [JZL15]. customisable [GBJ19]. customized [CSMB15]. customizing [FRKS12]. cut [RNJM17, SS15a]. cuThomasBatch [VLMPS^{+}18]. cuThomasVBatch [VLMPS^{+}18]. Cyber [SZ11, ANCA19, DZW^{+}11, GQH17, GOLL17, KM19, LCC^{+}18, SZMF19, WWL^{+}15, ZX11]. cyber-infrastructure [WWL^{+}15]. cyber-physical [ANCA19, GOLL17, LCC^{+}18, SZMF19]. CyberGIS [HLL^{+}15, HYL^{+}19, LPW15, YLH^{+}19, PWC^{+}14]. cyberinfrastructure [BFG14, CW07, HLL^{+}15, IUCH^{+}17, KHM^{+}11b, LGD15, MvWvM^{+}17, PRC^{+}14]. cyberinfrastructure-based [HLL^{+}15]. cyberinfrastructures [MRJ^{+}14, PSC^{+}17]. Cybersecurity [FGG^{+}18, GQH17, PZO19]. Cyberspace [WCWB19, LNG^{+}16, LNBL17]. cycle [KD10, NQL^{+}17]. cycle-scavenging [KD10]. cyclic [RS12]. cycling [CGKW13]. cyclotomic [CL14]. Cyclotron [KD10]. cylinder [LLY^{+}19a].

D [CCW06, OLG^{+}15, RWK17, VDL^{+}15, ACIC^{+}13, Boc19, CSMK19, CPE^{+}19, DCG11, EMEY14, KSM^{+}08a, MBP16, MCY^{+}07, MJL01, PSLC11, PDC16, PSCK^{+}15, QSZL18, TTR^{+}10, WSZ^{+}18, YBC^{+}07, ZLKK17]. D3 [JKL^{+}17]. DAC [HPD^{+}15, ABFL17, ABMdAMF19]. DAC-Hmm [HPD^{+}15]. DAG [RRR15]. DAG-schedules [RRR15]. DAI [AKK^{+}07, AAB^{+}05]. daily [MAVG16, SS18]. DALP [LWYM16]. damage [Wan18a, ZYL10]. damping [JAU19]. dark [Du18a]. DARPA [SCC^{+}10]. DART [DPK10]. DartGrid [CWMZ06]. Data [ABB^{+}15, CLT^{+}16, CXT^{+}18, EPB14, GS04b, GPZ04, HYQ17, KPS14, Lan17, LY14, MLS^{+}15, Mar19, MP04, PJ18, PB07b, PK08, PS13, SG19, AaBT17, AKK^{+}07, AHB^{+}10, AMGCC17, ANPR16, AC08, APHB16, ADM06, AMAB17, dRADFG17, BA18, BA19, BC16, BDG08, BTCB16, BCF12, BPL^{+}19, BM16, BLXE19, BKWM19, BB12, BB04, BV11, BLL^{+}19, BKZ^{+}13, BPS19, BZdR^{+}10, BSZ09, BHA^{+}15b, BWHS18, BMPP17, BDMM^{+}05, CMCAA17, CEH^{+}06, CRB^{+}17, CV07, CYD^{+}15, CLNR18, CBHTE11, CGBNM17, CY15, CT12, CCSS10, CLH^{+}16, CLF^{+}19, CZJY19, CBQ^{+}11, CFV^{+}08, CT11b, CCP^{+}15, CTAB16, CM18, Cuz11, CS13, DCG11, DIM18, DFLNP07, DLX^{+}16, DGW16, DS19, DIK14, DCY^{+}08, DGL^{+}12, DPK10, DZJ^{+}15, DDF^{+}17, DM15, DZC16, DS17, DL17, DSH18, DWG19, DC19b, DXM^{+}17, DS15, DYW16, DA15, EN19, EK19, ESG17]. data [EJD15, EJD17, FVLS15, FAM^{+}18, FAB^{+}07, FNI17, GM17, GLM^{+}16, GSR^{+}19, GD07, GvDHS12, GTL06, GLS^{+}19, GSWJ19, GKP^{+}09, HMFK15, HKA19b, HVM^{+}15, HKA^{+}15, HAAWA^{+}16, HZHP09, HWQ^{+}16, HAJL16, HSL19, HL19, HR18, HCG07, HYL^{+}19, HWZ^{+}15, HZY^{+}19, HLB10, HAvI13, IÁE11, IÁBE11, IRB19, JCP15, JLQ^{+}17, JFI^{+}08, JKL^{+}17, JLL18, JZL14, JZL15, KBDA19, KTB17, KMJ14, KOOB15, KT19b, KKL06, KTZ^{+}18, KC18, KCZ^{+}05, KB13, KB18, LBOE18, LSE^{+}13, LSS05, LL10, LTL^{+}17, LPH09, LLLJ14, LWF^{+}15, LLC^{+}15a, LWL15, LZY^{+}16, LGY17, LRLY17, LHHJ18, Li18, LWYZ19, LSZ19, LFWS15, LGL^{+}17, LLLyL16, LL16c, LWLZ11, LPW15, LBY^{+}16, LHLH16, LXW17, LHH^{+}17, LMX^{+}18, LDZ^{+}19, LLRS19, LMDP19, LMOT10, LPG^{+}14, LLL16, LGD15, MWL^{+}13, MMW16, MY17, MYY18, MZJ^{+}19, Mal05, MHK^{+}18, ML19a, MTT15, MDB^{+}17, MRS03, MWRK18, MISV13, MYS19, MÍM19, MCB14, MRP^{+}18, MCXP15, MGM^{+}08]. data [MWHW16, MSM^{+}14, MFGE19, MLVBW12, Nam19, NCD^{+}08, NDT^{+}16, OOTK01, OO18, OHJ13, OBD^{+}18, PCsHL18, Pat08, PDY14, PLW^{+}18, PMG19, PWS19, PHCR09, PC17b, PQP13, PIGK16, PdCMdS^{+}12, PS07, PXY^{+}07, PRU14, QXXZ16, QWW^{+}16, QGZL18, QZYZ16, RM19, ROQL18, RSKK19, RBT18, RKS02, RLZ15, RJ01, RLC16, RCCH19, SACRGL18, SIST18, SK04, SGD15, SAdB^{+}16, SBJ^{+}15, SSIH19, SZG^{+}19, SM19a, SPSB19, SdOVM16, SRAG16, SMBT07, SB19a, SB18, SM19b, SKA^{+}14, SCV^{+}08, SWZ12, SLC^{+}18, SLC^{+}19, TZK16, TTV08, TJ17b, TJ17a, TZLC15, TDC18, TCDMR^{+}17, TC12, UMD^{+}13, VSK17, VŠ11, VBW06, WWS^{+}12, WLW11, WSL15, WZS^{+}15, WJYH16, WQS^{+}16, WYBS16, WYZ^{+}17, WMC17, WQL^{+}18, WFKS18, WW19, WBC^{+}02, WBC^{+}17, WZL^{+}17b, WS19a, XCHK14, XPS^{+}15, Xha18, XAK16, XSMZ16, XYER16, XZH^{+}17, XBB13, XLHT17, XXX15, XDJL18, XGXH15, XLQL18, YYCH10, YWY^{+}10, YDB^{+}13]. data [YXL17, YR15, YCQF19, YWBS19, YZXW17, YNX^{+}16, YCSY19, YL19, YYL^{+}12, ZBE17, ZBEM18, ZLKK17, ZLN^{+}13, ZZ14, ZZ15, ZWL^{+}15, ZZ16, ZNT^{+}16, ZGS17, ZJS^{+}17, ZMYA18, ZXRZ19, ZCZ^{+}19, ZHJ19, ZWF^{+}06, ZSS18, ZHGX16, ZWH^{+}17, ZFW^{+}17, ZGX11, ZZL^{+}18, Zhu18, ZZL^{+}19, ZS19, vHvdSvL03, vRS05, vSB06, BD04, CM18, FVRM15, HYQ17, NJM19, PJ18, PSIP16, SG16, YYC10, ZLW^{+}19]. data* [JSR^{+}19a]. Data-aware [CLT^{+}16, Lan17, MDB^{+}17]. data-based [BTCB16, FVRM15]. data-centered [AHB^{+}10]. data-centric [SBJ^{+}15, SSIH19, ZJS^{+}17]. data-dependent [JFI^{+}08]. Data-driven [CXT^{+}18, LY14, BPL^{+}19, KCZ^{+}05, SRAG16, SB18, TCDMR^{+}17]. data-intensive [AMGCC17, CGBNM17, CTAB16, HAAWA^{+}16, JKL^{+}17, MWL^{+}13, WSL15, WQS^{+}16, YR15, ZWL^{+}15, ZWF^{+}06]. data-operation [LWLZ11]. data-oriented [QZYZ16]. data-parallel [BB04, GvDHS12, LPH09, MFGE19]. data-related [LLL16]. data-rich [LPW15]. data-sharing [ZZ15]. database [ABMdAMF19, AAB^{+}05, CWMZ06, CNG13, HmLGP03, LLB04, LLWS09, LW13, PN19, RPK08, dFMSPSW06, WLDL08, ZCL^{+}18]. Databases [GMF01, BGM03, GR13, MIGA18, OCS01, Rav16, SMFM18, SC07b, SW12, WDLL10]. datacenter [ZWL^{+}13, ZLZ15, ZWL^{+}15]. datacenters [ZCZ^{+}19]. Dataflow [WL11a, JDH^{+}18, PSS^{+}19, RSN^{+}19]. dataflows [KFD18, SdOVM16]. DataGrid [GTL06]. dataset [SL18, XYER16]. datasets [BMPS07, DP19, FCT^{+}02, HYL^{+}19, LTKF11, SMS^{+}19, TMP16, WCH^{+}07]. dataspaces [LHLH16]. Date [BCD^{+}02]. Davidson [RR11]. day [CAKH17]. Dayao [ZMJZ10]. DBLP [MRY^{+}16]. DCUDP [ZGS17]. DDCNN [CZQ17]. DDS [CRGR^{+}12]. DDS-based [CRGR^{+}12]. De-centralized [WGZL06]. de-disturbing [Du18a]. dead [CH04]. Deadline [LZBF17, EQW^{+}18, HKS19, KS19b, LWFL14, LGL16b, PCGE18, TCDMR^{+}17, ZH15]. deadline-aware [HKS19]. Deadline-constrained [LZBF17, PCGE18]. deadlines [VDB09]. Deadlock [LZC^{+}02, ASS19, EN16, YESG^{+}17b, YESG^{+}17a, YESG^{+}19]. deadlock-free [ASS19, YESG^{+}19]. deadlocks [BHA15a]. deal [CCO15a, SAMS19]. debate [YLLC18]. deblocking [JML^{+}16]. debris [WYW^{+}17]. debug [BLA^{+}14]. debugger [FSPC^{+}02]. Debugging [Ur07, WK07, HM04]. decentralization [HJH19]. Decentralized [YCQF19, ZXW16b, ACJ10, CZ11, CW11b, CL07, CCT15, DCEK15, DGM18, Del08, DS07, JDB16, PWMX17, SGD15, WZC16]. deception [BCdlCT06]. Decision [ART14, ABMMR19, ADK^{+}17, BXLJ16, BJC17, DCCZ18, GQJL18, GPS^{+}07, HLL^{+}15, LZZ^{+}15, MLYL17, MWL18, MLWA19, NNvVdA09, SS17a, WLZ17, GED^{+}18]. decision-making [ABMMR19, DCCZ18, GQJL18, GPS^{+}07, HLL^{+}15, SS17a]. decision-theoretic [BXLJ16]. Decisional [LZC14]. declarative [CMS17, ET15, HL06]. decoded [PCL17]. decoder [PMG19]. decoders [LDZ^{+}15]. decoding [LDZ14b, LLH^{+}18, Ree01, SJPB17]. decomposition [BJ01, BDV02, FH13, JN03, JKD19, LYL^{+}19, MGBC16, QML^{+}17, TSL^{+}19, ZLT^{+}16, dARP17]. decompression [FNI17, SHC^{+}16, SJPB17]. decoupled [DD16, JQSP08]. Decoy [MIGA18]. decryption [LJW^{+}17]. dedicated [CFPJ^{+}17]. Deep [NJM19, RM03, ZLW^{+}18, BLCC19, GLZ19, Hey19, HLG17, KTHA18, KHL17b, KSB^{+}19, Li18, LLJ18a, LLJ18b, LSC^{+}19, LMCL19, LDZ^{+}19, NJZZ19, QSX^{+}17, SRR19, VLW19, WF18, WCZ^{+}18, WGQ^{+}18, WWL^{+}17b, ZBZ^{+}18, ELM^{+}16]. deep-learning [WWL^{+}17b]. deep-learnt [BLCC19]. default [XDP18]. defeating [TJZ^{+}19]. defects [EPB14, JWW17, LZJ^{+}18]. defence [KKS12]. Defense [LWY^{+}17, AQRA^{+}18, FCY17]. defensive [GSK19]. Deferring [AM01]. Deficit [LJML10]. defined [FO18, GKM19, HLG17, LLC^{+}15a, TSA^{+}19, TKK^{+}11, ZTGW17]. definition [BGM03, EMS15, MCG^{+}08, OOTK01]. definitions [XWXC14]. deformation [LMGZ19, PNL10, ZGRSC10]. degree [GSWJ19, ZWLY16]. DEISA [Sod07]. Dekker [BDH16]. Delaunay [CCW04, CCW06, LCW^{+}17, PZ17]. Delay [WYZ^{+}17, ALL^{+}15, DZ13, LWFL14, OKBO19, RGCC15, WRLS12, WDW^{+}15, XW13, ZLG^{+}19]. Delay-bounded [WYZ^{+}17]. delay-tolerant [ALL^{+}15, OKBO19]. delays [DBR13]. delegatable [Rua15]. Delegation [GLL16]. delimitation [DLM13]. Delivering [YSL^{+}15, AG17b, MCC^{+}15, PSLC11, WL02]. delivery [CCK^{+}17, HKA^{+}15, VO15, YQL^{+}15]. Delphoi [MvNK^{+}06]. deluge [BDG08]. demand [ASWR12, BKWM19, CCSS10, CZG16, FSM^{+}19, KBDA19, LL10, LWYM16, MS17a, MS17b, MS18, VGN^{+}16, VO15]. demand-aware [CZG16, LWYM16]. demand-based [VO15]. demanded [ZS17]. demands [TGS14]. Democratization [MAK18]. Dempster [ECP18, JLQ^{+}17]. Demystifying [TPGC15, YQL^{+}15]. DENEB [FÁBE11]. Denial [ORdSL13, CYD^{+}15, VT15]. denial-of-service [CYD^{+}15, VT15]. denoising [ZYY^{+}19]. dense [ADI^{+}14, ADMQO14, BGGL07, BDR^{+}17, BHL^{+}09, BCI^{+}09, BLL18, CKL17, HLYD12, HGW18, KB13, KLDB10, MCP^{+}12, SD15, YDS^{+}14]. Density [ZH16, FGC06, KPS14, LW05, LDZ^{+}15]. density-based [KPS14]. density-driven [LW05]. Dependability [ANCA19, XDL^{+}11]. dependence [OKM10]. dependencies [KR11]. Dependency [Dra15, ALYD17, YYL^{+}12, ZXX17]. Dependency-based [Dra15]. dependent [BM08, CN16, JFI^{+}08, JM07, PB12]. dependent-failure [JM07]. Deploy [PSS^{+}19, CSMB15, YT15]. deploying [ABtGT^{+}12, PKB03, WCL^{+}10]. deployment [CLNR18, DZJ^{+}15, ESG11, KSB^{+}19, PLY13, RBBH02, WSWL12, YBX^{+}17, ZFWJ19]. deployments [WXLM19]. deregulated [CAC15]. derivatives [ATVLM14, DCJ14, WSRM12]. deriving [ATİ17]. descent [JLH^{+}16]. descripting [Bri16]. description [BH09, BFK^{+}17, CHH18, KS19a]. Desempenho [DC19a]. Design [BYN^{+}17, BKH08, BAZ09, CHX^{+}19, DIM18, EEK^{+}04, EFM17, FA18, FO18, GG09, GCN09, GLC07, HKA^{+}15, HG11, KL19, KKJH03, KVGH11, KN19, LCM12, LCC^{+}18, ML19a, MPT07, Nel05, PSS^{+}19, PCF^{+}17, PL15, Puf13, QW17, RGL^{+}15, SIS19, Sin10, Wan18a, YCL11, AM01, AAB^{+}05, BBGA03, CFPJ^{+}17, CWC10, CLH^{+}11, CHM15, DP14, DZL^{+}17b, Dra15, EME19, EB10, GTFA13, GRGP12, HZHP09, HXY^{+}12, HH19, HPS12, ISKvW02, JCVU15, JM07, KWL^{+}04, KAP13, KDW^{+}17, LGdVH13, LZW^{+}17a, LRS15, MKKB04, Mos19, PRC^{+}14, PPBB14, PMG^{+}15, PS19a, RCC17, SGM18, SAdB^{+}16, SBJ^{+}15, SCGZ19, SACJ04, SSM04, SWL^{+}01, SM19c, TKHA13, XBXS13, XL17, YYC10, YBB^{+}07]. designed [JOC^{+}15]. Designing [ABR^{+}06, PZH^{+}15, RBB^{+}09, Rua15, RLVRGÁ14, AMSR14, CWMW15, Mit17b, VGL16, XBZ10, ZCL14]. designs [MAS^{+}14]. desk [HFF07]. desktop [GKG^{+}04, LJHL10, TAB^{+}06, THF15, WJ09, ACJ10, LWY^{+}16]. destructive [EZJ^{+}18]. Detailed [SLGL16]. Detect [MRY^{+}16, GGC19]. Detecting [BA19, MAdS^{+}10, WWL^{+}17b, ZT09, GWHJL19, HPD^{+}15, HLZD18, LCH^{+}06]. Detection [AMSS15, GKM19, SLV12, SRR19, YWL^{+}17b, YZ19, AWR17, ABK^{+}18, AMP^{+}18, BA18, BLCC19, BTCB16, BMPP17, CSL^{+}19, CRB^{+}17, CCCC06, CYD^{+}15, CGI17, DLJ15, DDF^{+}17, Du18b, ECP18, FRKS12, GF07, GRTX18, HDDG09, HSJ^{+}18, HCS18, HYGF19, JLQ^{+}17, JKV^{+}15, KL19, KMBR19, LLL15, LWT^{+}16, LGQ^{+}17, LZL17a, LL18, LZJ^{+}18, Li19, LMCL19, LGQS12, LML^{+}18, LMDP19, LZC^{+}02, MW18, MPVT17, MMB^{+}17, MZA19, PFC14, PVCS18, PDCA17, PS07, RP19, RS16, RLDZ13, RG17, SZMF19, SL18, SM19b, VS19, VJ19, VRDTB^{+}16, WL11a, WZJD13, WDQ^{+}18, WF18, WKZL19, WML^{+}19, WW19, WBC^{+}17, Yan19b, ZBZ^{+}15, ZSWS18, ZXX17, ZJT^{+}19, ZL19, ZLZ^{+}17, ZLW^{+}18, ZLQ^{+}18, ZWW^{+}18]. detections [Qi17]. detector [EZJ^{+}18, JKV^{+}15, SS19b, YDL09]. determination [MJL01]. determining [FOTW04, Riz04]. Deterministic [ANH^{+}19, CDA09, BB12, SSMB15]. develop [SdSL18]. Developing [RHZ^{+}17, SRTG^{+}07, YAA07, LLT^{+}14, SMY^{+}15]. Development [FBH^{+}01, KSN16, KHZ^{+}15, MKKB04, MTVF14, Nov02, SFLS04, YLEB14, BR10, DCY^{+}08, DCFC08, FÁBE11, FPC15, HCC^{+}15, KA11, LGD15, MO02a, MGR02, PGP^{+}10, PSW11, SKK01, SKR17, VFAD17, VSB^{+}15, WYY^{+}19, ZGL19, vLRF^{+}02]. developments [DDF^{+}15, SFT15]. deviation [WDQ^{+}18]. deviations [RVRD10]. Device [TTD^{+}11, BR10, CEB^{+}18, GIVRC^{+}10, KKK^{+}19, MPR04, WWL^{+}17a, XHH12, ZHM^{+}17, ZWW14]. device-oriented [KKK^{+}19]. device-to-device [WWL^{+}17a, ZHM^{+}17]. devices [AA16, Aia15, CMCAA17, CL16, CFTT17, DGM18, ETR^{+}15, KSK19, LLLyL16, MBP16, MKO^{+}17b, PSLC11, PGF19, SWZ^{+}18, VLW19]. DevOps [Air17]. DEVS [BAZ09, IPGCMW18]. DF [RSN^{+}19]. DFA [TJD^{+}17]. dHPF [MCAB^{+}02]. DHT [ZFJ16]. DHT-based [ZFJ16]. DHTrust [XLL^{+}12]. DHTs [CCG^{+}08, ZLLL11]. diabetic [ZBZ^{+}15]. diagnosis [CMW02, DFH10, LM07, XYL18, XJZ13, YY19]. diagnostic [MKKB04]. diagnostics [ROA^{+}07]. diagonal [XLL^{+}15]. diagrams [WCR^{+}14]. dialect [Bac03]. Dictionary [KW19a]. difference [OFR^{+}17, PSCK^{+}15, ZCL^{+}19]. differencing [MWHW16, PH12]. different [Boe12, CSMK19, GMVRGS15, GCZ^{+}17, KCD19, MLS^{+}12, RVVPD^{+}17, VJ19, WMA07, ZWZ^{+}18]. Differential [MJS19, BJ01, CLW^{+}19, PS10, PIAH12, RLVRGÁ14, SZG^{+}19, TKB09, TCSBMG17, Wan18b, WYL14]. differentially [ZXRZ19]. differentiated [YESG^{+}17b, YESG^{+}17a]. differentiation [WYQ^{+}13, vdKEL10]. Diffie [LZC14]. diffusion [ČSMK17, CSMK19, ZLAa^{+}17]. Digging [DPS16]. digit [PPR19]. Digital [Zhu19, ASS08, CL01, CNAQ18, DLM13, He19, KHF^{+}17, LHC14, PPR19, lSsCY17, TGS14, VYK^{+}10, Xu19]. digitized [DKMM14]. dilatational [HTR10]. dimension [CBQ^{+}11, LML^{+}18]. dimensional [AVS^{+}19, AR19, CWYX17, DMC^{+}18, DCCZ18, DP19, GSB^{+}12, HLCW15, JQL^{+}15, JN03, JdM12, KOOB15, LLMM19, MMW16, MJZ17, MZJ^{+}19, MABP13, Ogi02, RCCH19, SWZ12, TBK06, Wan18b, WCH^{+}07, ZHJ19, ZM13, ZHZ^{+}13]. Dimensions [AvdADtH09, HP11]. Direct [AV07, BdL06, PGL^{+}17, WJ09, ZLF19]. directed [AS17, AR19]. directions [PMG^{+}15]. directive [NO02, RG19]. directive-based [RG19]. directive/MPI [NO02]. directory [JCP15]. disaster [LWYZ19, ZCLL19]. disasters [GRTX18]. DISCOVER [MMMP01]. Discovering [GD07, GBXL17, SKA^{+}14]. Discovery [KKW^{+}14, LHXY08, AMRW06, BM16, CLTT13, GZG^{+}16, GFG^{+}09, GWVP^{+}14, HVM^{+}15, LKKL16, LDXC13, LAM^{+}09, LLX15b, MLS^{+}15, MTHK14, ORDG15, OBD^{+}18, RCB^{+}04, RCXS09, RSTV07, SGG07, TLPS18, WQL^{+}18, WGG^{+}07, ZSZ15, SGG07]. discrete [FBS16, KW19b, MQOQOH01, SP16, SHP14]. discrete-time [MQOQOH01]. discrimination [GPVCdBRO12, HZH^{+}19, XLMH14]. DisCSP [PP17]. disease [CCLP19, Riz04]. diseases [KM19]. disjoint [ZWXS19]. disjunctively [QW17]. disk [DYY^{+}19, WCH^{+}07, WTL^{+}16, YYS15, ZBZ^{+}15]. disk-resident [WCH^{+}07]. disks [DXZ^{+}16]. Dispatching [CKSC10]. Dissecting [AUHWJ19]. dissemination [BLSP11, LWF^{+}15, MWPX17, MLRR09, PF12, RSPV17, WZS^{+}15]. disseminative [SW11]. Distance [YZW^{+}15, Zha19, BOF15, CMD17, HYGF19, ZGS17, ZWXS19]. distance-based [CMD17, HYGF19]. Distance-bounding [YZW^{+}15]. Distinguish [He19]. distinguished [EMB11]. DistMe [RTPPH12]. distribute [RHL^{+}18]. DIstributed [FK19, AOK19, ADSV16, AC09, Ano15a, BM12, BCCM16, CL10, CRB09, CT12, CPXA06, CM07b, CMD17, DFG18a, DSMM^{+}15, DBR13, DFH10, DC19b, EN09, FH13, FB16, GJ17, GZX17, GBD16, HFF07, JBL16, JCK^{+}13, Jos05, KMG^{+}18, MWL^{+}13, MN10, NSBR07, PHGK10, PDD14, RJ01, STS19, SCLK15, TWSM05, TTL05, TW07, TMZ07, Tur04, Ur07, VC16, XLWZ11, ZF18, ZWMT12, ACJ10, AAW^{+}02, AdSCdR^{+}19, ABDO09, AFG16, APHB16, AFT01, AMZ19, AT18, BGGL07, BFL^{+}10, BBCG02, BDF15, BAS07, BZdR^{+}10, BDV02, BYT^{+}12, BM02, BBGA03, BDMM^{+}05, CLMM12, CLNR18, CLR18, CACC11, CLTT13, CBPP02, CNG13, CZQ17, CKC09, CLH^{+}08, CGH^{+}06, CGN15, CN02, CPSP17, DCG11, DD17, DST11, DBGA16, DVB14, DLH01, DWC09, DKJ16, DvNM^{+}11a, DvNM^{+}11b, DL07, DZM^{+}15, EDBS08, EABVGV14, EBGS01, EJF^{+}16, EFA^{+}17, FBH^{+}01]. distributed [FJ05, FT06, FN13, FBS16, GGHR16, GDJ16, GKP^{+}19, GVC10, GLC07, GCL08, GLS^{+}19, GLD17, HWL18, HWR03, HKG08, IBvA^{+}02, JKL^{+}17, JSPE15, JZL15, KSN16, KAL07, Kes04, KTB17, KHM^{+}11b, KMJ14, KO06, KHZ^{+}15, KSM^{+}19, LL15, LWT^{+}16, LWL17, LRLY17, LCW^{+}19, Lia16, LMS18, LZC09, LLdA08, LBDS15, LMOT10, MvWvM^{+}17, MST^{+}05, MZ06, MMBP12, MLC04, MJ11, MFF04, MPSGD14, MRH14, MA15, MCSML07, MVML11, MP03, MDL^{+}10, Not16a, OSK^{+}01, OKW18, OHJ13, OAS^{+}15, OM06a, PCVZ^{+}04, PFC14, PRS16, PN19, PVR^{+}09, PHY^{+}18, PLC^{+}19, PWMX16, PWMX17, PAM^{+}15, PSC^{+}17, PQP13, RBO^{+}02, Rav16, RS11, RGCC15, RHD^{+}16, RM11, RO12a, RSTV05, RMCHMG15, SJB14, SK08, SFLS04, SLV12, SRM13a, SFCAV16, SG16, SMM18, SARL13, SCS17b, SFT15, SLM05, SAM^{+}17, SS19c, SHP14, SS15c, TTV08, TTL06, TCH^{+}13, TBK^{+}15]. distributed [TBH^{+}18, VGL16, VT15, VFAD18, WGZL06, WJYH16, WML^{+}19, WZLQ16, WW08, WTN07, XCHY13, XPWF15, XLT^{+}17, XW13, XXLL17, XLYL17, XLQL18, XWPM19, XLL^{+}12, YDL09, YYC^{+}19, YLJZ13, ZLKK17, ZCZ^{+}19, ZQZ^{+}16, ZW17, ZKR^{+}07, ZZ17, ZCL^{+}18, dSGD14, vHMB08, vLDW11, KAA19, TM01]. distributed-shared [BDV02]. Distributing [MT08]. Distribution [BD04, HMPPT13, MP04, QKSJ07, BJD^{+}19, BDZ19, CCC^{+}16, GM17, LLLyL16, LNKZ08, LFX^{+}08, MLG15, MZG19, ML19a, MAS^{+}14, MSG10, NPTT06, NTK08, PF12, QWW^{+}16, RKS02, RTPPH12, SGCG09, XGXH15, YGW17, YWL^{+}17a, YHHS16, YF13, ZLG^{+}19]. distributions [SRM^{+}15]. disturbing [Du18a]. divergence [CMMB13, DBH^{+}17]. divergences [CSPM13]. diverse [HMM^{+}09, VRSJ15]. Divide [ZLT^{+}16, CCW06, NDL17, YA04]. Divide-and-conquer [ZLT^{+}16, CCW06, NDL17]. divider [LCM12]. diving [SLC^{+}18, SLC^{+}19]. divisible [DL07, LYC16]. division [LZW13]. DMG [PB07b, PK08]. DNA [Bri16, CT16, dFdOSR^{+}19, HSHT14, LS15, MKKB04, MP17, SCRV11, SRF13, SER15, SSM04]. do [CHZ12]. Docker [Spa19]. docking [EDB^{+}14, EB14, PSICU18, TCP^{+}05]. Document [HKA19a, PLZ14, SIS19, YHASZ19]. documentation [vLDW11]. documents [CL01, CCLP19, LFH^{+}08b, Ros06, SS15a, ZSZ15, ZZ11]. DoD [MP04]. DOF [WRDZ13]. DOF-based [WRDZ13]. Domain [KCG^{+}19, ASE^{+}17, AHH14, BJ01, GRSB09, JN03, LFX^{+}08, LPG^{+}14, MGBC16, MHH16, PC17a, QH10, SJW18, SCGZ19, ZLQ^{+}18, ZTGW17]. domain-specific [MHH16]. domains [PGW06]. dominated [JN03, SJVR15, WBD^{+}19]. dot [Mos19]. double [JVMN19, WLLZ18]. double-fetch [WLLZ18]. doubling [WBO16]. down [DPS16, ML19b, lSsCY17]. downlink [LZW13]. dragonfly [YESG^{+}17b, YESG^{+}17a]. DRGs [LHB^{+}19]. drill [DPS16]. drill-down [DPS16]. DRIVE [CKSC10]. driven [BPL^{+}19, CLR18, CXT^{+}18, CM02, DMA13, DCFC08, EdPG^{+}10, FSG19, FMT16, GLMT15, HPS12, KHZ^{+}15, KCZ^{+}05, LFPP17, LW05, LTL^{+}17, LGY17, LY14, MT19a, PS10, PLL14, RVD^{+}12, SRAG16, SB18, TCDMR^{+}17, XDL^{+}11, ZlC15]. driver [BR10, ZZD^{+}17]. DroidAuditor [QXJS17]. drop [ALL^{+}15]. dropping [GGLD11]. drowsy [ZLW^{+}18]. drug [BBGA03, GWVP^{+}14, JvAB^{+}15]. DS [MS07]. DS/CDMA [MS07]. DSM [KBVP07, VHBB03]. DSMs [BKCP09]. DSP [KL12b, LYL07, LLYL09, SG18, WLL14]. DSS [GED^{+}18]. DTM [DFLNP07, MSN^{+}19, RSN^{+}19]. Dual [NLYZ12, BUVS10, JVMN19, YYZH19]. dual-core [BUVS10]. DUNE [EFY17]. duplex [ZZY^{+}19]. duplicate [RLDZ13]. Duplication [KBB17]. Duplication-controlled [KBB17]. during [HLW^{+}19, SSZ13]. duty [CGKW13, NQL^{+}17]. DVB [WWLD18]. DVFS [CGST17, PTL^{+}16, RR19, SR19a]. DVFS-enabled [SR19a]. DVMS [QLS13]. DVS [Hic18]. DWT [ZZZ^{+}15]. Dynamic [ALVY05, AMAB17, DLJ15, DK19b, FT06, GBSHA01, KTR11, KKL09, KSC12, MRS03, NPTT06, NCD^{+}08, NQL^{+}17, PYKL16, PGdCJ^{+}18, PSJM13, RPK08, RBT18, RSN^{+}19, SPJ14, SHC^{+}16, VCP16, VGN^{+}16, WZL^{+}17b, XWX^{+}19, YWY^{+}10, ZFWJ19, ASE^{+}17, BB12, BB15, BDF15, BB19, CK13, CY07, CWYX17, CJZZ10, DSO^{+}01, DMRS15, DZJ^{+}15, DKJ13, DMMA17, FÁBE11, GYB^{+}11, GLMT15, GD07, GD08, GPW03, IÁBE11, IRB19, JOC^{+}15, JKL^{+}17, KJŚ^{+}15, KR04, KYBV17, Kuł14, LOKW^{+}10, LK03, LCW^{+}19, LCW^{+}17, LJL^{+}17, LWS19, LSMVML15, MZ06, MSCS18, Mit19, NSBR07, NSN^{+}17, RHRB13, RCA^{+}12, Roj19, SKK02, Sod05, STL^{+}15, TMS^{+}12, TSKM18, WRDZ13, XCL09, XBXS13, XBZ10, XLYL17, XWH^{+}17, YIN19, YCSY19, YL01, Zen19a, ZP06, ZEB10, ZZY^{+}15, ZZZX19, ZZL^{+}18]. dynamic-memory [GYB^{+}11]. dynamical [GQ04, JAU19]. Dynamically [KL02, And13, GGFPGB14, HLYD12, KW19b, Li04, SWU08, WCL^{+}10]. dynamicity [ZX11]. dynamics [AaBT16, AHP^{+}13, BDW14, CAG^{+}13, CDP17, DLK^{+}18, GKS09, JB19, KCD19, KF11, LGL16a, LZG^{+}19, RCB03, TCP^{+}05, VCW13, WJLD09, XTZ10, ZNT^{+}16]. dynamics-based [AaBT16].

e-Business [DKMV07]. e-commerce [DJ19, PRS16]. e-course [ZL06]. e-Health [LRS15, LDS^{+}08]. e-infrastructure [ANK^{+}17, CPSP17, CPE^{+}19]. e-Learning [GBB^{+}15, ABR^{+}06, KKDS19, SW11]. e-Malaria [GKM^{+}08]. E-polling [KSK19]. e-Research [SHG^{+}07]. e-Science [DMM^{+}07, GBB^{+}15, HF05, WAS07, WC08, Xu08, KA11, LFH^{+}08b, BD08, CCK^{+}17, FGP^{+}11, PME^{+}08, RLS^{+}09, SM11, SBP12, SGV12, VBW06, WHW10, YDB^{+}13]. e-services [Hus15]. E2 [WYL14]. eager [NC05]. EAP [HZC^{+}14]. EAP-based [HZC^{+}14]. Early [HSV^{+}19, MFC18, SMY^{+}15, OKW15, SC07b, YWLQ18]. earth [Nak02, TKA^{+}02, TMAG03, AFR09, KHW05, Ogi02, PXY^{+}07, VGL06, Zho06, ZBC^{+}07]. Earthquake [Ano02, ZSL^{+}10, FKP^{+}02, HTR10, ISS^{+}02, JW10, MHRI14, OKM10, PWJ10, SNK^{+}15, TRH^{+}02, TTR^{+}10, YLEB14, YZ10, ZGRSC10, ZMJZ10]. earthquakes [MZS^{+}10]. EASI [WNN^{+}15]. EASI-CLOUDS [WNN^{+}15]. easily [PKB03]. easy [MTA^{+}07]. EasyGrid [BR04, BFM^{+}06]. eavesdrop [PCL17]. EC2 [PRNM19]. Ecce [SDB02]. EcForest [YHASZ19]. Ecology [SZ11]. eCommerce [AHH14]. Economic [ACS10, BAGS02, Fox10, SS15b, cDrLyC^{+}19, HJH19, KD15, MS13]. economically [MRS^{+}10]. ecosystem [HFTQ13, SAM^{+}17]. ecosystems [LFHT15]. ECperf [BG04]. EDA [LWK15]. edge [DED07, GSK19, JLL18, KSK19, MT19b, MJS19, SZG^{+}19, VYKM19, WML^{+}19, ZFWJ19]. edge-cloud [KSK19]. editing [VYK^{+}10]. Edition [FK19]. Editor [ZQH12]. Editorial [AI17, AF14, BL17, BG14, BL13a, BXQ17, CGBNM17, CR13, CL13, Din09, DS17, DKJ13, DKJ16, EN09, ESG17, FN13, FH01, Fox10, Fox17a, GAM17, GZX17, HYQ17, HF17, Lee09, LBT17, LBFS17, LWL17, Li17, LNBL17, MHJH16, MPSGD14, MH18, NPM19, OK18, PZO19, PCJ17, Pie08, PDD14, PC17b, QD17, RHJ13, RS13, SRM13a, SG16, SFH13, SNM15, TP14, WR17, WDM14, WDGK15, XCHK14, Xha18, XCHY13, XPWF15, XADLC15, XBM14, XYS17, ZBE17, ZZ17, LS14, McE10]. Editorials [ADF^{+}13, Ano15a, BH16, BCX15, BM16, BHQOS15, BL13b, CČJ^{+}16, DMD16, DP14, DM15, DL17, DR15, Fed13, Fox17b, FB16, GWD15, GJ17, HKVW16, HdV13, HTW14, HTBR16, Hus15, JRHJ16, KKT13, LBW14, LBS15, LBT16, LL13, LWL15, LL15, LNG^{+}16, LSS15, LS14, LLL16, MKO^{+}17a, MM17, MB15, Not16a, Not16b, OEP^{+}15, PCC17, PIGK16, PS13, QZH16, QFG14, QFT14, RBNG15, RHT13, RLC16, SF16, SRAG16, SFT15, VRMB13, VRSJ15, VC16, WT15, WS17, XAK16, XSMZ16, XBK17, YMLR16, ZZ16, Zhu15]. editors [HdV13]. Education [Che19, ABMMR19, Air17, AMRT14, LMH^{+}14, LPW15]. educational [LGJ17]. EEG [KOOB15, LML^{+}18]. Effect [CEB^{+}18, SC07a, WWY19, BGGS14, CAC^{+}08, CH19, JVMN19, KNT^{+}01, KKGO04, TV14, ZDHJ18]. Effective [EBGS01, WO02, BLL^{+}19, CM05, CCC12b, CS13, ESGQ^{+}11, HKS19, JK10, KQR^{+}17, LSE^{+}13, LCC^{+}18, MSP^{+}13, MCXP15, MA15, SS17a, WSZ^{+}18, YBX^{+}17, ZLN^{+}13]. effectiveness [CRB^{+}17, CTY15, Eng15, KAL07, LLdA08]. effects [BDW14, JAU19, PLC^{+}19, YG19, ZZYW10]. efficacy [LWW06]. Efficiency [PDD14, BBdS^{+}17, BN19, BDM18, CLL^{+}18, dAGC11, GA09, GCPS^{+}14, GVP^{+}14, HJH19, JPS17, KT19a, KS19c, LL19c, MM19, PZHS18, PLM^{+}19, QSX^{+}17, SSV19, SSZ13, TY15, Tru15, WR17, WCLH12, WTL^{+}16, XDJL18, XLL^{+}15, XL17, XHW^{+}19, YG19, ZYLY18]. Efficient [AD02, ANPR16, BB02, CCW04, CLF^{+}17, CGN15, DVL13, DOJ^{+}19, DZC16, GKS^{+}07, GP07, GÖ18, HZC^{+}14, HWQ^{+}16, HGW18, HC07, HLO^{+}16, LLKC08, LZT12, LDZ14b, LDZ^{+}15, LRLY17, LAM^{+}09, LLH^{+}18, MCG18, PZ11, RLDZ13, STO17, SZR16, TZG^{+}19, WYZ12, WLLL16, WHXzL15, XBXS13, YCW08, ZLLL11, ZSL^{+}15, ZYW^{+}16, AAHA18, ACGG06, AZF^{+}12, AMAB17, BD08, BF07, BG14, BB12, BB15, BAVM11, BT04, CLH^{+}11, CXC^{+}18, CHH18, CLH13, CGKW13, CS16, CS13, DCJ12, DRS^{+}13, DPP03, DHM14, DCA17, EME19, EA12, EFM17, FLL^{+}14, FIO15, zGWXT09, GTFA13, Gog11, GSK19, HJV^{+}19, HKA19a, HKA^{+}15, HTWW19, HCKF15, IC19, JLT06, JZL14, JWZ13, KS19a, Kar16, KBB17, KMG^{+}18, KKWZ15, KVGH11, KKL06, KHVK17, KTZ^{+}18, LLRS03, LLLJ14, LDPZ14, LWF^{+}15, LLL15, LCT16, LZL^{+}17b, LYF^{+}17, LWYZ19, LMX^{+}18, MTM19, MZJ^{+}19, MST13, MMB^{+}17, NK19, NS19, OGA^{+}01]. efficient [PA18, PPP10, PS19a, PBK19, PS07, ROQL18, RR19, RHuR^{+}19, RMP13b, SRS16, SR19a, SVB19, SRM^{+}15, SER15, SK04, Sha15, SHST13, SQS^{+}19, SGV12, STTW18, SKJ17, SYMA17, TJD^{+}17, VYKM19, WBZ10, WWLD18, WTY^{+}19, WZ16, WSWL12, WS19b, XRD^{+}17, XJZ13, XGXH15, XY17, XLQL18, YBO10, YLLZ09, YYC10, YYC^{+}19, YF13, ZXW16a, ZYZ06, ZY12, ZJL13, ZSZ^{+}14, ZZ15, ZS17, ZWT^{+}18, ZQZ^{+}16, ZWXS19, ZH15, ZZC15, ZHZ^{+}13, ZFJ16, ZHGX16, ZGX11, sCR19, vNMW^{+}05]. efficiently [RHL^{+}18, ZYH12]. eigenproblem [PV04]. Eigensolver [CKD^{+}19, AYN^{+}14, BWD15, RR11]. eigenvalue [BWD15, BIK^{+}11, GSV03, GKK09, YDS^{+}14]. Elastic [MDH^{+}16, MVML11, PB16, CLNR18, GYP^{+}16, LDXC13, MWPL15, SMFM18, WZYG19]. elasticity [dRRdCRR16]. elastohydrodynamic [GB07]. election [HLW^{+}19]. electric [CAC15, PWH18, WW19]. electromagnetic [AML^{+}15, XMJ17]. electromagnetics [PSG03]. Electron [CRC^{+}15b, GSB^{+}12]. electronic [CKRO13, GGFPGB14, JDH^{+}18, RGL^{+}15, SGL^{+}17]. electrophysiology [KSM^{+}08a]. electrostatic [VDL^{+}15]. element [BJ01, BCA^{+}10, CC13, CSTV06, GGR^{+}10, HKB07, JN03, KW19b, LLY^{+}19a, LHBW15, MO02a, OA02, PSV19, QH10, XM02, YYWQ19, ZGL19]. element-by-element [OA02]. elements [BHPS14, TGB^{+}10]. elevation [DLM13]. elevator [BDH18]. elicitation [RBDI17]. Elimination [LGFM05, AM01, DDF^{+}15, FED03, KS19c, LWW06, Tan12, TLX^{+}17]. elliptic [BJ01, BBB16, DVD^{+}12]. email [CAKH17]. Embedded [Fox17b, HTW14, MHJH16, VK12, Bri16, Fox17a, HXY^{+}12, KHZ^{+}15, LTK17, MSP^{+}13, MÖO17, RHT13, SDH^{+}17, STWSP12, SG18, VH12, WST^{+}17, XCHY13, YWY^{+}10]. embedding [Li04, NJZZ19, TJ17b, WLP^{+}17, YHASZ19, YLZ18, ZZZX19]. embeddings [YFL18]. embodiment [TNH15, TNI16]. Emde [Kuł14]. emergencies [AG17b]. emergency [GSR^{+}19, MSST15, NFG19, RHS17, SJ19, XZH^{+}16]. emergent [GGR^{+}10]. Emerging [Ang07, CY15, CS06, ZYH09, GLL16, WAD12, Qiu11, QFG14, QFT14]. EMF [Dra17]. Emmerald [AB01]. eMOLST [vLDW11]. Emotion [LML^{+}18, FLG19, RP19]. emotional [GLZ19]. empathy [HCBRM16]. Empirical [AHH14, Bok12, CHZ10, GM17, GGV14, LH17]. employing [HON04, TLX^{+}17]. emulated [VRDTB^{+}16]. emulation [NR08]. Enable [Air17, ADM06, BM10, BBGA03, KKL09, LS19a, TMS^{+}12]. enabled [AAF17, CKL19, DFLNP07, ETR^{+}15, FMM08, GHB^{+}06, HYX05, LAC^{+}08, ORdSL13, PSV19, PXY^{+}07, RSTV07, SR19a, SGV12, TWZ^{+}19, TZG^{+}19, WBD^{+}03, YJZZ19, ZYLT06]. enables [CDH^{+}15]. Enabling [ACC^{+}15, BDI^{+}07, CPS^{+}14, DPK10, DDF^{+}17, DZL^{+}17a, KJ19b, LCA^{+}19, LLLJ14, PF12, PML^{+}05, Spa19, SPSNvS07, BR04, DR15, FPR05, LPW15, RMCA12, dRRdCRR16]. enactment [OKP16]. eNB [BN19]. encoder [PMG19]. encoding [BSZ09, DXG13, SM19a]. encrypted [HKA19b, LZY^{+}16, LW13, PWS19, WS19a, ZZL^{+}19]. encryption [ATKH^{+}17, BZD16, CLH^{+}16, CMMS17, CZ15b, HKA19b, HL19, LGP19, LFZ^{+}17, LWW^{+}19, LFWS15, LZC14, LJW^{+}17, MMS17, MML16, Nam19, SKB^{+}17, WZC16, WLFX17, XWXC14, XXX15, YZCT17, YSQM19, ZWY^{+}19]. end [CGBNM17, CK13, GHLS19, GM10, JK13, LGL^{+}17, TMZ07, WL02, ZKJ^{+}07]. end-host [TMZ07]. end-to-end [CK13]. endurance [ABK^{+}18]. Energy [ABC19, AZF^{+}12, ANK^{+}17, BFH17, CGMJ^{+}19, CFTT17, EOD^{+}19, IRB19, JWZ13, Kar16, KHM^{+}11b, KKWZ15, LDPZ14, LZL^{+}17b, LWL^{+}19, MTM19, MABP13, MMB^{+}17, NK19, PLM^{+}19, PSICU18, PRV11, RIWS17, SRS16, SR19a, Sha15, SHST13, SKJ17, SYMA17, WWX^{+}19, XGXH15, YBO10, ZJL13, ZWT^{+}18, ALZR11, AAC^{+}15, ADI^{+}14, ADMQO14, ADSV16, AHK^{+}15, AMAB17, ABDR13, BG14, BB12, BB15, BBdS^{+}17, BN19, CGST17, CAC15, CLS14, CLH13, CGKW13, CSB^{+}16, DXM^{+}17, DMMA17, EN19, EZJ^{+}18, FMT16, GTFA13, GMVRGS15, GA09, GCPS^{+}14, GVP^{+}14, HJV^{+}19, HKA^{+}15, HSL19, JZL14, KT19a, KC15, KBB17, KD19, KTB17, KS19c, KN19, KCG^{+}19, LCW^{+}19, LYC16, MS13, MBMB18, ML19a, MAVG16, MWRK18, MST13, MCC16, NSSAK13, NSSAK16, OKBO19, PLL14, PTL^{+}16, PLL17, QSX^{+}17, ROQL18, RR15, RR19, RHuR^{+}19, RMCN^{+}07, dRRdCRR16, Roj19, RGB^{+}15]. energy [SNEP14, SQS^{+}19, SPQ^{+}17, SSZ13, TY15, Tru15, WBZ10, WCZX16, WDT18, WS19b, XRD^{+}17, XXLL17, XHW^{+}19, ZLF19, ZY12, ZQD^{+}17, ZH15, ZYLY18, ZGL07, PDD14]. Energy-aware [ABC19, CFTT17, EOD^{+}19, PRV11, RIWS17, CGST17, DXM^{+}17, EN19, JZL14, KC15, LYC16, MS13, RGB^{+}15, WDT18]. Energy-based [CGMJ^{+}19, PSICU18]. Energy-efficient [JWZ13, Kar16, LDPZ14, LZL^{+}17b, MMB^{+}17, SRS16, SR19a, SHST13, SKJ17, SYMA17, XGXH15, ZJL13, ZWT^{+}18, BG14, BB15, CLH13, CGKW13, GTFA13, HJV^{+}19, HKA^{+}15, KBB17, MST13, ROQL18, RR19, WBZ10, WS19b, XRD^{+}17, ZY12, ZH15]. energy-harvesting [OKBO19]. energy-proportional [ZLF19]. ENES [VGL06]. enforcement [Dam11, EBMD13]. enforcing [MLL^{+}11]. engagement [XLY^{+}16]. engine [CEM^{+}08, Kar14a, KSS^{+}17, MGI17, MMSN^{+}01, ODS^{+}13, PPBB14, RDP10, SGJ^{+}17, VJK13, WJP14, DAC12]. Engineering [AF14, MP02, BHJ^{+}16, HPS05, HY12, JCK^{+}13, LSS05, LBFS17, LM07, LHBW15, MHRI14, PCC17, SRAG16, TQL^{+}14, WYW^{+}17, XLY^{+}16, YLEB14]. engineering-level [XLY^{+}16]. engines [CLW^{+}18, IPGCMW18, LBOE18, WZ04, WZ16, XLYX11b]. enhance [dAGC11, Jon09, LWW06, MJD17]. Enhanced [AMBT17a, BDTdS13, DK19a, DNB19, DXWD16, FE17, FTT15, GSK19, HFR^{+}17, KT19a, KABD07, KT19c, LHHJ18, LLJ18a, LLJ18b, LLC^{+}15b, MRS^{+}10, NJZZ19, PZZ08, PZZ10, PLZ14, WO14, XL17, YHASZ19, YHJ^{+}14]. enhancement [DXM^{+}17, PBD^{+}15, PJW^{+}14, SVY19]. Enhancements [AM07, AKK^{+}07]. Enhancing [ADI^{+}14, JCJ17, mLGP03, DWC09, GCSB19, HYL^{+}19, MJS19]. enough [PLR^{+}14]. ensemble [BY12, QXXZ16, YSW19, ZKJ^{+}07]. ensuring [ALYD17, SGL^{+}17]. ENT [IUCH^{+}17]. enterprise [DZL^{+}17b, KD10, KM13, SLD^{+}12, ZDL19, LR05, YAA07]. enterprise-oriented [DZL^{+}17b]. enterprises [DR15, HJH19, HH19, LS19b]. ENTICE [GKP^{+}19]. entities [Kri13, LZW^{+}17a]. Entity [FLG19, CCLP19]. Entropy [CYD^{+}15, EPB14, LFW19, MLG15, PSIP16, SJPB17, SG18, YYWQ19]. Entropy-based [CYD^{+}15, EPB14]. entropy-weighted [YYWQ19]. entry [MMS07]. enumeration [CSMS^{+}19]. Environment [SMBT07, ABMMR19, ABR^{+}06, BAZ09, BPB08, CLX07, CJZ^{+}15, CGS15, CLH^{+}08, CH04, CGN15, CEG^{+}05, CN02, CLX^{+}12, CMS17, DKKL06, Den07, DL10, DCP^{+}17, EB18, Erw02, FTRA15, FLG19, GZG^{+}16, GRSB09, GD07, GAW09, GVP^{+}14, HL19, HR06, HAE09, HLL^{+}15, HH19, IZXM09, JZZL06, JLL18, Kes04, KB17, KMG^{+}18, LAC^{+}08, LWL17, LLT^{+}19a, LSZ19, LJHL10, MZS^{+}10, MLS^{+}15, MMMP01, MYDM06, MW18, MGR02, NR17, OGA^{+}06, PLY13, PVR^{+}09, PBK19, RHRB13, RSKK19, RBT18, RGV09, RRWS08, SVB19, SR19b, Sch02, SDB02, SM19b, SKB^{+}17, SAM^{+}17, VDdN^{+}07, VP19, WL12, WZZL13, WCCL05, WLL03a, XXX15, YDB^{+}13, YCWH07, ZYZ06, ZYZ^{+}12, ZWL^{+}15, ZYZC17, ZL06, ZDL07, vNMW^{+}05, CKSC10, HBG^{+}06, IBvA^{+}02, JM19, LV12, MCSML07, SPR^{+}07, vSB06]. Environmental [ZP19, BAT13, LTL^{+}17, LL19a, MWRK18, ZKR^{+}07]. Environments [CR08, CL08, CC09, EN09, RIP18, SRdS09, SF10, ABtGT^{+}12, ANCA19, AFG^{+}05, ARPPM17, BCK^{+}09, BMA03, BHA^{+}15b, BWW^{+}08, CLR18, CLQ^{+}17, CVK15, CPE^{+}19, CS13, DCG11, DFLNP07, DHC11, EN19, EAGVBVDS11, EABVGV14, FP02, FRU12, JRHJ16, JQSP08, JPWH02, JK13, KKK^{+}19, KJŚ^{+}15, KZY15, KZY^{+}18, LLMK18, LR05, LZC08, LZC09, LW13, LQL^{+}15, MT19b, MST15, MRS^{+}09, NNK^{+}07, PGK11, RRBB11, RWK^{+}02, RVRD10, RHuR^{+}19, RGCC15, RB17, SS17a, SN18, SM03, SGV12, SAP16, SA19, TZG^{+}19, TB12, UR04, VGL16, WBD^{+}19, WTL^{+}16, XRD^{+}17, YK10, ACF^{+}07]. EPIC [Nev17]. epidemiology [SMS^{+}19]. equal [DLW19]. equal-area [DLW19]. equation [BdL06, LTZ^{+}19, OLG^{+}15, PS10, SSK11, Wan18b]. equations [BJ01, BEQOR17, BDE^{+}19, CSTV06, FYKW15, GSV03, HKB07, KD07, MQOQOH01, PIAH12, TYL^{+}15, WLL14]. equilibrium [SPH13]. equipment [ZQD^{+}17]. equivalence [RG18]. Era [SG19, ELM^{+}16, ESG17, LDZ^{+}19]. erasure [PWMX16, XGC19, XWPM19]. erasure-coded [XWPM19]. ERD [Kri19]. ergonomics [SZT19]. ERLANG [NTK08]. ERP [HH19]. Errata [DKJ16, TNI16, WLWX16]. Erratum [Ano06, Ano15d, Ano15b, Ano15c, NSSAK16, YESG^{+}17b]. error [BZD16, KCS07, LWS19, SWS^{+}18, TLM17]. error-resilient [SWS^{+}18]. error-tolerant [BZD16]. errors [BDW14, BL04, LCH^{+}06]. eScience [Bou13, TH10, BDMM^{+}05, MKO^{+}17a, PS13]. Establishing [RCKV12, SJ19]. establishment [CVK15]. Esterel [Sin10]. Estimating [BOF15, JSPE15, DS15, HSHT14, ZWLY16]. estimation [AMHC11, ABG^{+}13, DCK12, LLC^{+}15b, LWS19, SWLJ17, VB16, WDT18, YWR^{+}19, YGG14, vEGW06]. Estimator [ZCXH17, HML19, HML20, LLH19, FPHZ20]. Estimator-based [ZCXH17]. ETAS [IRB19]. Ethernet [Gog11, ZLA^{+}15]. ETL [DYW16, TALT16]. ETNGRID [Ang07, CS06]. EUBrazilOpenBio [ABB^{+}15]. EULAG [RCR^{+}15]. Euro [BL09a, BL09b, BL11a, BL11b, BL13b, BL13a, LBW14, LBS15, LBT16, LBT17, Wis02]. Euro-Par [BL09a, BL09b, BL11a, BL11b, BL13b, BL13a, LBW14, LBS15, LBT16, LBT17, Wis02]. Europa [OTG^{+}07]. EuroPar [CM07b]. European [CRC^{+}15b, GG07, GTL06, KOK14, VGL06]. evacuation [LLL^{+}19]. Evaluating [AJY^{+}15, BDM18, CTY15, DAC^{+}18, FVLS15, KTHA18, MOF15, MSCS18, NPL19, OSK^{+}01, TKHA13, VSR^{+}09, VdSK^{+}05, GMVRGS15, KKV13, MSN^{+}19, MRS^{+}09, Rua15]. Evaluation [CGST17, dCPD13, FK19, HML20, JDH^{+}18, LLH19, MN10, SGJ^{+}17, XPS^{+}15, ACMM06, BLCC19, BNNH19, BD04, BUVS10, CRCC09, CMVRRVGI17, CKOG10, CNG13, CWYX17, CDdW17, CLW^{+}15, CMS17, DLPV07, DBB^{+}16, DCCZ18, DMA13, EEK^{+}04, EQW^{+}18, ESG11, FLYL16, FG16, FGZ^{+}18, FN13, FMP10, GS08, GGV14, GS04a, HG11, JOK^{+}18, KJŚ^{+}15, LBOE18, LGP19, LH05, Li04, LL16b, LHB^{+}19, LFH08a, LBDS15, LC17, MBP16, MDH^{+}16, MPT07, MAH^{+}02, Not16b, OCC^{+}05, PB12, PBD^{+}15, PGC^{+}19, PRS01, QLZX19, QW17, RGAK15, SM02, SFCAV16, SIS19, SCGZ19, SPQ^{+}17, TKA^{+}02, WMA07, WKT08, WSZ^{+}18, WJKS18, XWD^{+}12, YZ10, ZF14, ZCC^{+}06, ZDC15, ZL12, ZQMC19, ZDX12]. evaluations [TMP16]. evaluative [YYWQ19]. evasion [MPVT17]. Event [XXY^{+}16, CWZL13, DA19, FSG19, FP02, FBS16, GZHF19, GCN09, GRTX18, Kar14a, KW11, LWT^{+}16, LLX15b, LLZ^{+}17a, MLZ19, MT19a, PVCS18, SWD^{+}17, SHP14, TL19, VEJD17, WDQ^{+}18, WK12, WCLC13, YP10, ZlC15, ZFT08]. event-based [SWD^{+}17, YP10, ZFT08]. event-driven [FSG19, MT19a, ZlC15]. event-triggered [TL19]. events [SGJ^{+}17, XZH^{+}16]. Every [HLW^{+}19]. Everywhere [AJM12]. evidence [TLWZ14, YY19, ZW09]. evidence-based [ZW09]. evolution [Arz17, CAG^{+}13, RLVRGÁ14, SDB02, TKB09, TCSBMG17]. Evolutionary [Du18c, ZQLZ12, ADD^{+}05, CMVRRVGI17, CQXW14, CLW^{+}15, DST11, EOD^{+}19, JC07, KFD18, LC09, LZG^{+}19, LF17, Yan19a]. Evolvable [HXY^{+}12]. evolving [DJ19, ER12, FNBS16, Li04, QXXZ16]. exact [RLDZ13]. example [KF11]. examples [EFY17, HZL^{+}16]. Exascale [AS19, SG19, BDL^{+}15, DPP^{+}19, ESG17, PLC^{+}19]. ExaStencils [KHVK17]. exception [QLD^{+}11, TCBR^{+}10, WFHT17]. exchange [AR16, AKG13, CH19, DCJ14, FIO15, FBS16, GVK12, LVN^{+}12, LBY^{+}16, PSC^{+}17, QMK12]. exchanged [QLLS15]. Exchanges [AS15]. excitation [RCA^{+}11]. exclusion [BDH15, BDH16, BDH18, DHV03, HBD18]. executable [FED03]. executing [LPSF11]. Execution [SAP16, AHM06, ARPPM17, AAE^{+}09, BPB08, CEMR19, CZY^{+}18, CMB06, CCP^{+}15, DRS^{+}13, EJD15, FÁBE11, FOTW04, FM08, GdMK^{+}18, HPS12, KWL^{+}04, LM08, LPS^{+}09, LCT16, LF17, LY14, MDB^{+}17, MCG18, MYDM06, PPBB14, QLD^{+}11, RR19, RC09, RMCHMG15, TZK16, TNIB17, TBK^{+}15, XLZD13, YIN19, dOOO^{+}12]. executions [NB12, YCQF19, dSGD14]. executives [RS12]. exemplars [KB13]. exercise [GPS^{+}07]. exhaustive [DYY^{+}19, KHF^{+}17]. exhibits [WST^{+}17]. existing [SGD^{+}18, BDT01]. ExNa [WZ16]. exotic [DCJ14]. expedition [WSP17]. expensive [GPV09, ZYZ^{+}12]. Experience [Ano06, BHW05, SNB^{+}01, TH10, BCC^{+}05, CHPvdG07, DN19, GTL06, KQR^{+}17, KBH^{+}15b, LRS15, MKB01, MFC18, RSC^{+}15, TTL05, TDM^{+}02, WWG^{+}11, WJLD09, FH01]. Experiences [AHK^{+}15, BGV^{+}01, GBMM15, MSL^{+}14, SLD^{+}12, BCM^{+}05, NCWD^{+}04, SBJ^{+}15, SMY^{+}15, SRTG^{+}07]. experiment [BD08, HGB^{+}08]. Experimental [BRWB06, GBB^{+}15, LFX^{+}08, ZDHJ18, AFG16, DYY^{+}19, FPHZ19, GRS06, HWY^{+}17, Pan20, RGAK15, WOH^{+}13, ZPG10]. experimentation [ACG18]. Experiments [BCdlCT06, CT12, FLMRC02, KBT^{+}14, PPST09, ZBC^{+}07]. experts [MTHK14]. Explicit [KGK17, BVIB19, FIO15, JMF09, KW18, ZZC15]. exploit [Nev17]. exploitation [ADF^{+}13, FTT15, OGA^{+}01]. Exploiting [BP17, BCI^{+}09, BG14, CS13, DMMA17, EdPG^{+}10, Jun16, LPH09, ZQZ^{+}16, ABFL17, ACG15, ACCM17, JLL18, KTM^{+}09, Nob08, SPW09, TZLC15, VSB^{+}15, YZR14]. Exploration [Bao19, CLH19, FTT15, ABDP15, BYN^{+}17, KAP13, PDY14, PL15, WBC^{+}17]. Exploration/exploitation [FTT15]. explorations [JCVU15]. exploratory [HSRN11, XYER16]. Exploring [BS10, CSL12, CS17, DZJ^{+}15, HTHW16, LNKZ08, SZA08, ZYH09, ZYH12, XLL^{+}18, YBZ^{+}15]. exponentiations [TKS18]. export [MWRK18]. exposure [Dra17]. ExPregel [SNH15]. express [ZDL19]. expression [KHL^{+}17a]. expressions [MFGE19]. expressiveness [YS07]. expressivity [KS04]. expressways [XHD^{+}19]. Extended [DPS16, LXW17, SSZ14, AS15]. Extending [BVGVEA11, BEQOR17, BDB^{+}13, EK19, CJC^{+}18, FTR15, FMMD13, OKW18, PY19]. extensibility [Slo06]. extensible [CS15, DLZ16, RHD^{+}16]. extension [DIK14, HKRR08, SCV^{+}08, cWsThY19, ZG04]. extensions [ANTZ09, BDT01, ISKvW02, Pac16, SIOS02, ZZ14, vRKS03]. external [LSL^{+}17, ZYFZ19]. extract [CZL^{+}17, LWZ^{+}17]. Extracting [CSBL12, MGM^{+}08, ZCLL19]. extraction [BTCB16, CSL^{+}19, DCG11, DPK10, HLO^{+}16, LFX^{+}08, RVVPD^{+}17, TJ17a, WYZ12, YTL19]. Extractive [YHASZ19]. extrapolation [CH04, KRS11]. Extreme [BCA^{+}10, BBB^{+}14, DZJ^{+}15, EN16, RLRG15, TLX^{+}17, WQS^{+}16, YLL^{+}18]. extreme-scale [BBB^{+}14, EN16, YLL^{+}18]. extremely [MWW10]. Exudates [ZBZ^{+}15]. Eye [MRS^{+}09]. EZRP [KABD07].

F [KS04]. Face [WZJD13, Du18b, HDDG09, QSZL18, XHH12, MKX^{+}15]. FACE-IT [MKX^{+}15]. Facilitating [Qi17, VFAD17, LHHJ18]. facilities [ACMA07, BVGVEAFG11, MH07, WLDL08]. facility [WHXB19]. factor [JyLdZ^{+}18, ZSWS18, AS15]. Factor-Optical-Factor [AS15]. factorization [BLKD08, Cho01, DFLL14, JLH^{+}16, WZ04, WZL^{+}17a, YTD17, ZF18, ZDG^{+}14]. factorizations [ADMQO14]. Factors [HMM^{+}09]. fading [MS07]. Failure [EPA15, LFHT15, BN19, GMS09, JM07, MST13, SBS19, SZMF19, SC07a, SK18, TSA^{+}19, YDL09, ZWH^{+}17, BDB^{+}13]. failure-aware [ZWH^{+}17]. Failure-resilient [EPA15]. failures [GSR^{+}19, LSW07, MAS^{+}14, PWMX16]. fair [DXWC16, GGS^{+}16, HLC19, LLKC08, TZYL13]. Fairness [RCT03, NvV09, dSGD14]. FairThreads [Bou06]. Faithful [STL^{+}15]. Fake [JBL15]. Falcon [VRMB13]. false [ECP18, CACC11]. false-praise [ECP18]. family [Kac11, KHM^{+}11a, SLM04]. FarGo [GBSHA01]. farming [CKBB14]. farms [GVC10]. Fast [ACC^{+}12, BEDK18, DFC12, DÇK^{+}18, FYKW15, GCO^{+}14, HBD18, LLH^{+}17, NN07, PSHL11, WBO16, ZGZ19, AB01, CL19, DLZ16, DZL^{+}17b, FOTW04, HAA^{+}17, KW19a, KCB09, Kuł14, LZH^{+}15, LZJ^{+}18, LY14, MKKB04, MRH14, MWL18, NA15, PMAL14, SJVR15, TCP^{+}05, WZLQ16, ZSS18]. faster [LS15]. FastFlow [ART14]. fat [ESGQ^{+}11, LGY17, ZJKL10]. fat-tree [LGY17, ZJKL10]. fat-trees [ESGQ^{+}11]. Fault [AAE^{+}09, BV11, FD01, LHT^{+}09, ASB^{+}19, NDP^{+}05, ZJS11, ACJ10, ADM06, BF07, BHBD13, CCCC06, CJZ^{+}15, ET15, Fec12, GWHJL19, GGR^{+}10, GG19, HTR10, ISS^{+}02, KAL07, OKM10, PYKL16, PGL^{+}17, PGK11, ROA^{+}07, VYK^{+}10, XPWF15, XTLG08, YY19]. fault-tolerance [CJZ^{+}15]. Fault-Tolerant [NDP^{+}05, AAE^{+}09, FD01, ZJS11, ACJ10, BF07, Fec12, KAL07, PGL^{+}17]. faulted [PNL10]. faults [KF15, XM02, ZCXH17]. FC2Q [ACG15]. FCMS [ACG17]. FDK [ZGZ19]. FDTD [XMJ17]. feasibility [BdL06, HKG08, SS15b, SZA08]. feasible [ATİ17]. FEAST [TGB^{+}10]. Feature [Pre01, CSL^{+}19, DZL^{+}19, HAA^{+}17, KOOB15, MSM^{+}14, PWH18, SS19c, TJ17a, WGQ^{+}18, ZWLY16, ZLW^{+}19, ZWL^{+}17]. Feature-oriented [Pre01]. Features [KS05, vLGL^{+}02, BLCC19, BDY03, KBH^{+}15b, LSH^{+}16, LCM^{+}17, VYKM19, ZYZ16, ZWW^{+}18, ZHW^{+}16]. Feautrier [Viv03]. federated [BFG14, GRSB09, GVK12, KMG^{+}18, LHL10, LBdM^{+}16, MSST15, HF17]. federation [BTCGL17, MML^{+}17]. federations [DMRS15, HAAWA^{+}16]. feed [HH19]. FEM [OA02]. femtocell [WRDZ13]. Ferrini [Ano06]. fetch [MNL15, WLLZ18]. FFBAT [AMS17]. FFT [BHM^{+}12, DT01, JKM^{+}17, NS19]. FFTs [JKD19]. Fi [MB17]. fiber [KJ19b]. field [JVMN19, KMJ^{+}17, MZ06, NNH^{+}14, XXY^{+}16]. field-based [MZ06]. field-programmable [NNH^{+}14]. fields [GBXL17, HCKF15]. Fifteenth [BJ18]. File [PPR19, AC06, BGdCCA11, BKND16, BDT01, BAS07, DL10, DT17, DZM^{+}15, HYX05, HCK^{+}08, ID18, IT03, KKL09, Lia16, LHH^{+}17, LLYL09, Mit17b, PGF19, SACRGL18, SNB^{+}01, TWN07, Tru15, YYCH10, ZH08, AC08]. file-transfer [AC06]. files [CCC12a, LYL07, LLYL09, SdOVM16, ZZC^{+}17]. filling [LBH07]. filter [AA16, ATKH^{+}17, BY12, CDP17, Jos05, LWQS19, PWH18, PA18, SS19a, SS19b]. filtering [BHA^{+}15b, CAKH17, DFG^{+}18b, Du18a, IZXM09, JML^{+}16, LH17, SWD^{+}17, TZLC15, VŠ11, WTY^{+}19, YAG19]. filters [GPV09]. finance [DMD16, JHCH19, PW12, TP14, DDE^{+}12]. Financial [YWLQ18, GCO^{+}14, GQH17, GGS^{+}16, HJH19, MB16, QGZL18, RDP10, TTPJ16, YL19]. find [STTW18]. Finding [ATI14, BL04, CT11a, DS02, JCVU15, KB13, MLZ19, MSV^{+}10, KHM^{+}11b, LSL^{+}17, TZK16]. findings [GCPS^{+}14]. Fine [BVGVEAFG11, BHA^{+}15b, Hoa10, JCP15, KWL^{+}04, CLH^{+}16, CLX^{+}12, JSR^{+}19a, NNvVdA09, QML^{+}17, RAFD14, RLVRGÁ14, SK18, TNH15, TNI16, sTzNL16, WLW11, WZL^{+}17a, ZYN^{+}07]. Fine-grain [Hoa10, JCP15, NNvVdA09]. Fine-grained [BHA^{+}15b, KWL^{+}04, CLH^{+}16, CLX^{+}12, JSR^{+}19a, RAFD14, RLVRGÁ14, SK18, TNH15, TNI16, sTzNL16, WLW11, WZL^{+}17a, ZYN^{+}07]. finger [CCG^{+}08]. fingerprint [ZHW^{+}16]. Finite [XM02, BJ01, BCA^{+}10, BHPS14, CC13, CSTV06, JN03, LLY^{+}19a, LHBW15, MO02a, NNH^{+}14, OFR^{+}17, PSG03, PSV19, PH12, PSCK^{+}15, QH10, TGB^{+}10, ZGL19]. finite-differencing [PH12]. Finite-element [XM02, JN03, MO02a]. finite-volume [PSG03]. FIR [SS19a]. fire [ACCM17]. firefly [AMS17, DS19]. firewall [CWMW15]. FireWorks [JOC^{+}15]. First [MLA^{+}08, WJLD09, MKO^{+}17a, MSPDP19, PMAL14, CR08, CS06, DT15b]. fish [LKPM09]. fit [PMAL14]. fitness [BSP11, EOD^{+}19]. Fitting [GM17, Ley06, PLL14]. fixed [CY07, KW11, PSV19]. fixed-grid [PSV19]. fixed-priority [KW11]. fixed-time [CY07]. FLAPW [DFTHD18]. flash [LWF^{+}15, DVD^{+}12, AMABS18]. Flexibility [BKM^{+}07a]. Flexible [BAVM11, CGKW13, CJ12, HR18, ACG18, BM10, BFM^{+}10, CCL^{+}17, CEG^{+}05, DZL^{+}17a, IT03, LGL^{+}17, WNT02, dRC10, vNMW^{+}05]. flight [ZQMC19]. FLND [SS19b]. floating [BTG06, LCM12, TLM17]. floating-point [BTG06, TLM17]. flood [HGB^{+}08]. flooding [GS08, GKM19, LLH^{+}18]. flooding-based [GS08]. floorplanning [ACIC^{+}13]. flow [AMTM17, BZB17, DdB01, EFM17, GCWE15, GPS^{+}07, GKM19, HKB07, IÁBE11, KKK^{+}19, KHVK17, LW05, LXW^{+}16, LHHJ18, LL16c, LFW19, LHXY08, LXL^{+}09, MWLS11, RNJM17, SARL13, WYW^{+}17, Wan18b, ZYW^{+}16, ZDHJ18, ZGL07, vLDW11, GHB^{+}06]. flow-shop [AMTM17]. flows [BFM^{+}10, DGW16, SPZ^{+}10, XWX^{+}19]. fluid [BFM^{+}10, BDY02, BDY03, EFM17, GHLS19, HC07, LLY^{+}19a, MB14, MWLS11, RCB03, RCR^{+}15, ZLL19]. fluid-particle [BDY03]. fluids [BDY02]. FluMapper [PWC^{+}14]. flux [AHB^{+}10]. fly [PS07, YESG^{+}19]. flying [SK17]. FMIPv6 [WCLH12]. FMM [ABC^{+}16, MRH14]. focus [AHH14]. focused [DH13, PZZ08, PZZ10]. Fog [AOK19, HAN19, KM19, BJD^{+}19, CLW^{+}19, GSK19, HLC19, KSK19, LWW^{+}19, SPSB19, TSA^{+}19, WWC^{+}19, YBX^{+}17, YJZZ19, ZYY^{+}19, SWHL16]. Fog-to-Cloud [AOK19]. folded [QLLS15]. Folders [Ros06]. folding [NCWD^{+}04, TTD^{+}05]. folksonomy [FBYO12]. follow [PBD^{+}15, PdCMdS^{+}12]. follow-up [PBD^{+}15, PdCMdS^{+}12]. Followee [XLL^{+}18]. follower [ZCXH17]. following [LJPP16, XLL^{+}18]. food [MKX^{+}15]. footprint [DS15, SZR16]. foraging [DK19b]. forecast [ABC^{+}08a, VCW13, ZZYW10, ZMYA18]. forecasting [HHKA14, TTR^{+}10]. foreign [DCJ14]. Forensic [CMCAA17, RCC17, SB19b, TDC18]. forensic-by-design [RCC17]. forensics [TDC18]. foresight [WKL^{+}11]. forest [ACCM17, YHASZ19]. Foreword [AS19, BL17, DC19a, FK19, HF17, LB19, MH18, PCC17, STS19, WR17, XBCW19, ZRB19, ZZ17, ZC19, KKT13]. fork [RR15]. fork-join-based [RR15]. form [BLDW16]. Formal [KAP13, Aia15, BAZ09, EL01, GWHJL19, MSCS18, QLF^{+}06, XWXC14]. formalisms [WGP^{+}15]. formalized [XBZ10]. format [CXC^{+}18, CBIGL19]. Formation [CG10, BCdlCT06, CNAQ18, HAAWA^{+}16]. forms [BIK^{+}11, GKK09, ZXW16b]. formula [WXLM19]. formula-based [WXLM19]. formulae [vEGW06]. formulations [SSB^{+}14]. Fortran [BB02, BSB^{+}03, CCW04, FSPC^{+}02, ISKvW02, KS02, LCC^{+}03, MDV07, SIOS02, SD03, vWAH^{+}02]. Forum [GPW03, Lee09, SKNH09]. Forward [WCWB19, ATKH^{+}17, BDB^{+}13, OKW15]. forwarding [WDW^{+}15]. Fostering [VAC^{+}07]. Foundations [Nar05]. four [Li19, WCH^{+}07]. four-dimensional [WCH^{+}07]. four-wire [Li19]. Fourier [SP16, WCZ^{+}18]. Fourth [CW11a]. FPGA [BYDC19, GSB^{+}12, LDZ14b, LGQ^{+}17, PSS^{+}19, QSX^{+}17, TPT^{+}18, WZ04, YOBS16, ZDX12]. FPGA-accelerated [QSX^{+}17, TPT^{+}18]. FPGA-based [LGQ^{+}17, WZ04]. FPGAs [GC19]. fractional [ČSMK17, CSMK19]. fragmentation [LCMY13]. frame [MPHL03, TKHA13]. FRAMESELF [AM15]. Framework [Ber07, EFG^{+}03, ABMMR19, ATKH^{+}17, AM15, ALZR11, AAW^{+}02, ABC^{+}08b, ADK^{+}16, ARPPM17, BPdM06, BB15, BKCP09, BKS18, BR04, BAZ18, BAC^{+}15, BSZ09, BAG17, BMPP17, BBA18, BPD06, CCCW13, CKL17, CA06, CLW^{+}18, CVK15, CS15, CN16, CM02, CGB^{+}06, Cuz11, DZW^{+}11, EBMD13, EDBS08, EHSU07, FAPC16, FRB^{+}06, FMT16, FJG^{+}13, GQH17, GWW17, GD08, GLC07, GGC19, GSK19, HK02, HAE09, HLHC12, HFTQ13, IAH^{+}15, JZJW15, JM07, KDC17, Kri05, KTB04, KSK17, Ley06, LZL17a, LS19a, LGL^{+}17, LGG16, LMOT10, MB14, MDV07, MZA19, MV16, NMM^{+}10, Nev17, NRW04, OISS07, OTG^{+}07, PSRR14, PWWR05, PTL^{+}16, PRG15, PMG^{+}15, PSS^{+}18, QXJS17, QSMK04, RBO^{+}02, RG19, RSC^{+}15, RHD^{+}16, RSMFE^{+}12, RCLSK16, SR19b, SGM18, SPG08, SA19, SZL09, TTV08, TMP16, TKB16, TPV17, TBK^{+}15, UAW09]. framework [VS02, WFHT17, WJJM17, WQL^{+}18, WNT02, WLW14, WSWL12, XZJ11, YGL05, YR15, ZXW16a, ZS01, ZP06, ZLN^{+}13, ZSZ^{+}14, ZCL14, ZDB^{+}14, ZL12, ZHM^{+}17, AK01, Zho06, ZBC^{+}07]. Frameworks [HmLGP03, vdS06b, CNKJ18, GPS^{+}07, GGV14, LLT^{+}14, PPMH15, PBF15, PNJP19, SZT18, SK17, YWA07, vdS06a]. framing [CHZ10, CHZ12]. FRAPP [ZLN^{+}13]. free [ASS19, AR19, BT04, BAS07, CLL^{+}18, DDF^{+}17, JWY^{+}17, LLC^{+}15a, LS05, LLH^{+}17, WWC^{+}19, YESG^{+}19]. freedom [IR11, YESG^{+}17b, YESG^{+}17a]. frequencies [SCS17a]. Frequency [KS19c, AD15, CSBL12, DK19a, EA12, Kar16, LZW13, NLYZ12, PLZ14, XBXS13, YZW^{+}15, AMAB17]. frequency-band [DK19a]. frequency-division [LZW13]. frequency-improved [PLZ14]. frequency-inverse [PLZ14]. frequent [CT11a, KT19c, LXYC17, ZSZ^{+}14]. freshness [ASS08]. FRFB [XLL^{+}18]. friction [JAU19]. friend [HLF^{+}17]. friendliness [BKM^{+}07a]. friendly [CGR19]. Frontier [XADLC15, LWL15, TZK16]. Frontiers [ZYH09, ZYH12]. FRP [WZLQ16]. FSO [BKA19]. FSVM [XLG19]. FT [XTLG08]. FT-Grid [XTLG08]. fuel [SR19b]. Fulfilling [HHPL16]. Full [CSC^{+}17, RSTV05, DMMA17, MB16, NK19, Pla08, ZNT^{+}16, ZZY^{+}19]. full-adder [NK19]. full-duplex [ZZY^{+}19]. full-pixel [Pla08]. full-system [DMMA17]. fully [NFG19, OHJ13, SK04, ZABP18]. fun [vdS06b]. Function [Soo16, Boe12, CMMS17, HTWW19, KS19c, LGQS12, LMX^{+}18, LMDP19, PWJ10, PSM^{+}11, SJISVR17, TSA^{+}19, TQL^{+}14]. function-call [HTWW19]. functional [CC15, KS05, LDPZ14, MSP^{+}19, SRL^{+}14, TBH^{+}18]. functionally [PSV19]. functions [CMVRRVGI17, EOD^{+}19, FGZ^{+}18, LPSF11, QZDJ16, SPZ^{+}10, vWAH^{+}02]. fused [FTT15]. fusion [AT01, And13, GKS^{+}07, GIVRC^{+}10, GLZ19, KW18, KC18, LC18, PJ18, RMCHMG15, SZ18, SR19c, WLW14, WZXZ12]. Future [SCNH07, Arz17, DLP03, KAMB19, QD17, SN18, WKL^{+}11]. futures [BGGS14]. Fuzzy [CEMR19, ACG15, ACG17, BTCB16, CWYX17, ECP18, HW16, KT19c, LZL^{+}17b, LWW06, MMBP12, MLWA19, QLZX19, RSR06, SS19b, SC19, TVCB18, WBD^{+}19].

G [LCYJ08, MCWL06, RMP^{+}13a, RSTV07, YHK09]. G# [LGL16a]. G-1 [RMP^{+}13a]. G-BLAST [YHK09]. G-PASS [MCWL06]. GA [BPL12, FTT15]. GAF [PWWR05]. Galaxies [MCC^{+}15]. Galaxy [ACC^{+}15, MSL^{+}14]. Game [BJC17, HJTX17, PRC^{+}14, AOK19, CG10, CQXW14, CLW^{+}15, FCY17, FXX16, HAAWA^{+}16, LC09, LZG^{+}19, MZW^{+}16, SS17a, TXZ^{+}17, WWS^{+}12, WLZ17, XZH^{+}17, Yan19a, Zen19a, ZZY^{+}19]. game-based [MZW^{+}16, Zen19a]. game-theoretic [CG10]. games [BDP18, CRC15a, Ios11, JVPI18]. gap [CZQ17, Hun15, RSSM06, BHJ^{+}16]. garbage [AP06, BCK^{+}09, HM03, JM19, Kal11, Puf13]. garbled [IKP19]. gas [WJLD09]. Gaspar [MSS16]. gate [JVMN19, NNH^{+}14]. gates [IKP19]. Gateway [DT15b, WDGK15, BSC^{+}15, CM07a, CGK^{+}07, CDH^{+}15, GWD15, JvAB^{+}15, MRJ^{+}14, MWL^{+}15, MKX^{+}15, PGP^{+}10, PMG^{+}15, SBJ^{+}15, SvDO15, SBB^{+}15, SMY^{+}15, Sod07, VSB^{+}15, CGK^{+}07, GBB^{+}15, LPW15, PYF02]. Gateways [WD07, ACF^{+}07, GBMM15, HMFK15, LZWD^{+}15, MCC^{+}15, MTA^{+}07, MCD^{+}15, OTG^{+}07, SvDO15, Sod07, WBB^{+}07]. gather [MTK16]. GAUGE [HBG^{+}06]. Gauss [BEQOR17, Has17, KS19c, LL19c, Tan12]. Gaussian [DDF^{+}15, Du18a, HCKF15, VŠ11, YWL^{+}17a, ZLH^{+}18, Zhu19]. gaze [MRS^{+}09]. GbA [LZW17b]. GCC2004 [JX06]. GCE [Tho07]. GCF [FRB^{+}06]. GCViR [TSBR10]. ge [BTCGL17]. GEANT4 [CRC^{+}15b]. gear [GHLS19]. GeForce [QWZ^{+}19]. GenApp [BAC^{+}15]. gene [KHL^{+}17a, EMS11, RGL^{+}15]. Gene/L [EMS11]. Gene/P [RGL^{+}15]. General [ETR^{+}13, ABDP15, AdSCdR^{+}19, BSZ09, CSL^{+}19, FRB^{+}06, FBV^{+}13, LKPM09, MWPL15, PSRR14, PVCS18, RMP^{+}13a, RZL^{+}19, SNK^{+}15, TXY^{+}16, WQL^{+}18, WLW14]. General-purpose [ETR^{+}13, ABDP15, AdSCdR^{+}19, LKPM09, RMP^{+}13a, SNK^{+}15]. Generalized [PHY^{+}18, BCM^{+}07, BMS^{+}09, CL14, DFC12, KSM15]. Generate [DIK14]. Generate-map-reduce [DIK14]. generated [He19, YOBS16]. Generating [ER12, vHKT^{+}11, AAP13, Ios11, KHVK17]. Generation [LXL^{+}09, Aia15, Ang07, BFK^{+}17, CSC^{+}17, CC13, Can06, CLR18, CZ19, CDL08, CS06, CPS^{+}14, CEM^{+}17, DCD^{+}14, GPS^{+}07, HCKF15, ISS^{+}02, JAU19, KKTHL13, KB06, KBE07, KMJ14, KJ19b, LLZ^{+}17a, LMO15, LBH07, MSL^{+}14, MSPDP19, MK12, PPMH15, PWMX17, QEB^{+}10, UAW09, XW13, XBM14, Zhu19]. generations [AP06, RVD^{+}12]. Generative [HBG^{+}06]. generator [DYW16, TNIB17, vWAH^{+}02]. Generic [ALMT19, LL05, APHB16, dRADFG17, GvDHS12, MHK^{+}18, RG19, SO16, TKS18, XZJ11]. genes [COdO^{+}11]. Genetic [SAB15, TZK16, AS17, ACCM17, BYN^{+}17, BDTdS13, GYM14, GRQ19a, HW16, KKWZ15, KPS14, LHHJ18, LWW06, LZBF17, LLY19c, MHLC^{+}05, PN19, PCF^{+}17, PV15, RP19, Riz04, SJVR15, SZD19, TRW07, WLL03a, YWLQ18, ZL19]. genetic-based [KKWZ15]. genome [MKAKG14, WWL^{+}15]. genome-based [MKAKG14]. genome-wide [WWL^{+}15]. genomes [ALVY05, COdO^{+}11]. genomics [TGS14, MSL^{+}14]. genre [RP19]. geo [JZL15, MHK^{+}18, PAM^{+}15, SWS^{+}18, XLQL18, ZCZ^{+}19]. geo-distributed [JZL15, XLQL18, ZCZ^{+}19]. geo-opportunistic [SWS^{+}18]. geo-referencing [PAM^{+}15]. geo-related [MHK^{+}18]. geodynamics [ZKJ^{+}07]. geodynamo [DGJ11]. GeoFEM [FCT^{+}02, MO02a, MO02b, NO02]. GeoFEST [PNL10]. geographic [JWZ13, ZCLL19, WYW^{+}17]. geographical [ASG^{+}08, QCB17, ZSZ15]. geographically [KTB17]. geolocated [RHS17]. Geolocation [ZLG^{+}19]. Geometric [LC18, CLZ^{+}17, SJW18, ZF18, ZYLT06]. geometrical [FMS15, ZDHJ18]. geometries [BFM^{+}10]. geometry [ZP06]. geometry-based [ZP06]. GEONGrid [YBB^{+}07]. geophysics [TSL^{+}19]. Geosciences [PW05, MCY^{+}10]. geoscientific [BvIF10]. geospatial [BMPP17, DCY^{+}08, LPW15, YLH^{+}19, Pie08]. geotagged [Jun16]. GEP [LMX^{+}18]. gestures [JTD^{+}19]. Getting [Nob08]. GF [SAD13]. Gibraltar [CSWB11]. gigabyte [FCT^{+}02]. GIS [ABC^{+}08a]. GIS-based [ABC^{+}08a]. git [JKL19]. glass [JWW17, LZJ^{+}18]. gLite [KSM^{+}08b, KKV13]. Global [BFL^{+}10, FWU^{+}04, NDT^{+}16, AHB^{+}10, ADK^{+}16, BDMM^{+}05, HKS^{+}12, HBKM06, LF17, LLYL09, MT19a, NPL19, Ogi02, PRD^{+}13, TKB09, TBK^{+}15, VBW06, XRD^{+}17, YSL^{+}15, YCW08, ZDB^{+}14, ZHW^{+}16]. Global-scale [BFL^{+}10]. Global-view [NDT^{+}16]. Globus [ACFT15, DCY^{+}08, Jac02, Kri05, MSL^{+}14, MCC^{+}15]. Globus-based [DCY^{+}08, Kri05]. GMarte [AHM06]. GMP [SFLS04]. GNSS [CL19, LWZ^{+}17]. go [LSS15]. goal [JBL16]. goal-oriented [JBL16]. goals [TALT16]. goats [ZWZ^{+}18]. GOLD [PCH^{+}08]. gone [LSS15]. Google [MGI17]. Gossip [OHJ13, ABDO09, BDL^{+}15, NJ15, VvSI07, ZK08]. gossip-based [VvSI07, ZK08]. GP [LSP15]. GPAW [RGL^{+}15]. GPFlow [RRWS08]. GPGC [ZYLT06]. GPGPU [CCO15a, MKO^{+}17b, MMG^{+}18, PIAH12, WFJ^{+}17, ZWW14]. GPGPUs [BCI^{+}18, FO18, NPL19, PW12, SSB^{+}14]. GPS [LWZ^{+}19]. GPS/WLAN [LWZ^{+}19]. GPU [CSMS^{+}19, SPZ^{+}10, ADK^{+}16, ABG^{+}13, BP17, BG17, BEQOR13, BDE^{+}19, BFM^{+}10, BKSM^{+}15, CMVRRVGI17, CMMB13, CNAQ18, CSPM13, CBB^{+}19, CS16, DRZ13, DÇK^{+}18, DMC^{+}18, DBH^{+}17, ER12, EOD^{+}19, FA18, Fer13, FTT15, FNI17, GSB^{+}12, GMMT17, Has17, HGW18, HQoS11, HW16, IOOH12, ISO^{+}14, JML^{+}16, JLH^{+}16, KW19a, KH12, KMJ^{+}17, KHF^{+}17, LLMM19, LGP19, LOSJ17, LDZ14b, LSH^{+}16, LCT16, LZL^{+}17b, LXYC17, LLH^{+}15, LS15, LLH^{+}18, LSP15, MMO^{+}16, MD19, MSN^{+}19, MNL15, MÖO17, MWLS11, NRR15, NSN^{+}17, OFR^{+}17, PDY14, PRG15, PDC16, PGdCJ^{+}18, PL15, PLL17, PK17, PH12, QWZ^{+}19, RSC^{+}15, RSMFE^{+}12, RWK17, Roj19, SSIH19, SIRP17, SAP16, SD15, SN16, SS15c, TPGC15, TDM^{+}15, VLJ17, VLW19, VLF^{+}13, WLLL15, WLLL16, WDG^{+}14, WBO16, WWLD18, XMJ17, YIN19, YTD17, YT19, ZCL^{+}18, ZDX12, dCRS11]. GPU-accelerated [ADK^{+}16, CMMB13, IOOH12, JLH^{+}16, LZL^{+}17b, LS15, PGdCJ^{+}18]. GPU-aware [BDE^{+}19, FA18]. GPU-based [ABG^{+}13, DBH^{+}17, HW16, LLH^{+}18, MMO^{+}16, PDC16, PK17, RSMFE^{+}12, Roj19, SSIH19]. GPUs [ALKD16, ASS19, AHK^{+}15, BDR^{+}17, BCI^{+}09, BY12, CGMJ^{+}19, CLF^{+}17, CZL12, ETR^{+}13, GW15, HP11, HCKF15, JKM^{+}17, KVGH11, KW18, KB13, LC17, MLS^{+}12, PSICU18, PSV19, RS11, RPRG17, RCLSK16, Sør13, TXY^{+}16, VL17, VLMPS^{+}18, VFG11, YDS^{+}14, ZZZ^{+}15]. GPUSGD [JLH^{+}16]. GRADE [Kac11]. graded [PSV19]. gradient [JLH^{+}16, JWW17, KL19, SK09, SSK11, MDL^{+}10]. gradual [RC09]. graduate [MTVF14]. grain [Hoa10, JCP15, NNvVdA09, Yos06]. grained [BHA^{+}15b, CDA09, CLH^{+}16, CLX^{+}12, ID18, JSR^{+}19a, KWL^{+}04, MDL^{+}10, RAFD14, RLVRGÁ14, SK18, TNH15, TNI16, sTzNL16, WLW11, WZL^{+}17a, ZYN^{+}07]. gram [PDCA17]. grammar [PS10]. grammar-driven [PS10]. grammars [LS05]. GRAND [VDdN^{+}07]. Grande [Fox01, Fox05, GPW03, GPW05]. granularity [DKJ13, RCA^{+}12, TJF14, YKA^{+}19, dSGD14]. GRAPES [LXRJ13, LTZ^{+}19]. Graph [PS10, ZCZ^{+}19, AQRA^{+}18, AMZ19, BOF15, CLF^{+}17, CMD17, DZL^{+}17a, EPB14, Hoh06, HTWW19, JLL18, KRW17, LZZ^{+}17, LZW17b, LL19c, PZH^{+}15, SNH15, SKK02, SS15a, WGY^{+}19, ZBZH11, ZHGX16]. graph-based [AQRA^{+}18, Hoh06, JLL18, LZW17b]. graph-cut [SS15a]. Graph500 [FBV^{+}17]. GraphBAD [PVCS18]. graphic [MPSGD14]. Graphical [DT15b, Eng15, LPH09, PSRR14, RMP^{+}13a, VDL^{+}15]. Graphics [ADF^{+}13, CP14, DCJ14, DG11, KC13, MCB14, VCW13, ATVLM14, ACC^{+}12, ABDP15, BDW14, BHQOS15, CGIP16, CSWB11, DCJ12, GWW17, GGV14, JdM12, LKPM09, LDZ^{+}15, LLH^{+}15, MAS16, OLG^{+}15, PSCK^{+}15, RCA^{+}11, RCR^{+}15, RK15, SPMP11, SPZ^{+}10, SAD13, SNK^{+}15, lSsCY17, Str11, SEF^{+}14, TZKH12, WJT^{+}14, WCZX16, ZO14, ZDG^{+}14, LSXL17]. graphs [AS17, DOJ^{+}19, FLMRC02, FBYO12, GG19, HWL18, LSL^{+}17, MGM^{+}08, RG18, RHL^{+}18, TSKM18, ZQK15]. GRAPLEr [SAM^{+}17]. GRASP [AMTM17]. gravitational [SR17]. gravity [HTR10]. Gray [Bou13, cDrLyC^{+}19]. grease [ZCW^{+}18]. greedy [SQS^{+}19]. Green [MAS16, CL13, DZ13, HSL19, KSK17, PTL^{+}16, SQS^{+}19, ZRB19]. Gregory [vEGW06]. Grey [KB17]. GRFA [LLYL09]. GRID [Ang07, CS06, ACF^{+}07, ACD02, CL08, CC09, FKP^{+}02, GIVRC^{+}10, GHPR05, Lee09, MTD^{+}02, NNTH^{+}02, PvLV^{+}02, PC17b, QXXZ16, Tho07, vLGL^{+}02, AC08, BC16, BAD^{+}11, BKM^{+}07b, BFVRC15, BZdR^{+}10, BWW^{+}08, CLW^{+}19, CPB07, CHL15, CRC^{+}15b, CSL08, CY08, CFV^{+}08, CLX^{+}12, CS13, Dab09a, DIM18, DMRS15, DZC16, FHO^{+}15, FMS11, FTRA15, HR18, HGB^{+}08, IOOH12, Ios11, JQSP08, Kac11, KD10, KV12, KKT13, KZY15, KN19, KBH^{+}15b, KA11, LC09, LZC08, LLSL15, LLL^{+}19, MLS^{+}15, MB12, MAS^{+}14, ME08, MSV^{+}10, NNvVdA09, PVR^{+}09, PSV19, PV15, RRBB11, RHRB13, RSTV07, RGV09, SGM18, Sha15, SGV12, SR17, SKNH09, THF15, TV14, TSBR10, VDB09, VSK17, VSKK09, WCL^{+}10, WSW^{+}12, dRL10, dAAVS12, vdABST10, vLFGL01, ACJ10, AKK^{+}07, AC02, ACC^{+}07, AHM06, ABR^{+}06, AV07, ACMM06, AC06, AAB^{+}05, ADM06, AFG^{+}05]. Grid [BR04, BKM^{+}07a, BDG^{+}10, BPB08, BLSP11, BAGS02, BM02, BBGA03, CEM^{+}08, CV07, CLX07, CRB09, CWMZ06, CA06, CY07, CR08, CW07, CLH^{+}08, CL07, CMB06, CDL08, CBP^{+}04, CGB^{+}06, Cyb06, DDP^{+}06, DDX^{+}06, DCY^{+}08, DFPT06, Dik07, DPS07, DKMV07, ET09, Erw02, FJP^{+}05, FP02, FG06, FAB^{+}07, FZ07, FS07, FZ08, Fox10, GEJ^{+}08, Ger05, GKG^{+}04, GS04a, GD07, GAE^{+}06, GTL06, GHB^{+}06, GKP^{+}09, HK07, HBG^{+}06, HPS05, Hoh06, JZZL06, JX06, KA09, KWL^{+}04, KR06, KFS^{+}06, Kri05, LW05, LAC^{+}08, Ley06, LWL^{+}06, LX08, LZC09, LFH^{+}08b, MCWL06, MRS^{+}10, MCY^{+}07, MWJ^{+}10, MP02, MBP^{+}05, MCCG11, MPT07, MGR02, NAP^{+}07, NZKK11, NSBR07, NNK^{+}07, NCWD^{+}04, NPTT06, Nov02, NJ05, OISS07, PFU^{+}05, PML^{+}05, PWWR05, PB07a, PHGK10, PXY^{+}07, QLF^{+}06, QLC04, RWK^{+}02, ROA^{+}07, RBBH02, SWU08, SBBE07, SDB02, SM04]. Grid [SN06, SCNH07, SANB08, SRdS09, SF10, SL10, SLT^{+}06, SB19a, SPBL06, Slo06, SRTG^{+}07, TWSM05, TMS^{+}12, TMZ07, UR04, VD05, VDdN^{+}07, VBW06, WCA08, WKT08, WBC^{+}02, WL02, WD07, WBD^{+}03, XPBS11, XTLG08, YAA07, YHK09, YDB^{+}13, ZBP06, ZCC^{+}06, ZKA07, ZH08, ZWF^{+}06, ZBP07, ZYLT06, ZKJ^{+}07, ZXXN06, ZL06, Zhu07, ZDL07, ZL09, dCGKG06, vNMW^{+}05, vLDA07]. Grid-based [GIVRC^{+}10, QXXZ16, HR18, HGB^{+}08, CA06, DPS07, YHK09]. grid-enabled [RSTV07, GHB^{+}06, LAC^{+}08, PXY^{+}07, WBD^{+}03, ZYLT06]. Grid-Flow [GHB^{+}06]. Grid-wide [GEJ^{+}08]. grid/cloud [MB12]. Grid5000 [AFG16]. GridASP [OISS07]. GridBLAST [Kri05]. gridification [AAV^{+}15]. GridLab [KKM^{+}06]. GridPortlets [ZKA07]. GridRPC [ABC^{+}08b]. Grids [FP09, PB07b, PK08, ASS08, ADSV16, BM08, BKM^{+}07b, BHPS14, CPB07, CC10, CG10, CCSS10, CW11b, Cuz11, DST11, DPGA11, FBC10, GKSR14, dAGC11, GRS^{+}17, JMF09, KOK14, KBG^{+}09, KSPM12, KKWZ15, KLP^{+}08, KW19b, LHL10, MLG15, MAdS^{+}10, MLVBW12, MSG10, NO02, Nak02, SM11, SNGR18, SVN12, TZK16, VDPC03, WP12, XTLG08, YYCH10, YLC11, ZP07, Zhu15, ZS19, AAHRW04, AKW04, ADD^{+}05, ASG^{+}08, BFM^{+}06, BDMM^{+}05, CRCC09, Can06, DS07, DRF07, FPR05, GQ04, GD06, HF05, HCK^{+}08, KTB04, KTM^{+}09, LSS05, LHT^{+}09, Pie08, PGO^{+}04, RMCN^{+}07, SAC^{+}07, SCV^{+}08, SD11b, TTV08, TKB09, TAB^{+}06, VBW06, WGZL06, YHK09, YYC10, ZKR^{+}07, dABV08]. GridSAT [CW07]. GridSim [BM02, SCV^{+}08]. GRIDSpace [DWC09]. GridSphere [NRW04]. GridWay [CHL15]. Gromacs [KF11]. Group [GG07, MH18, RRR04, TH19, GPS^{+}07, GKPT13, HWQ^{+}16, KKDS19, LWSZ19, MMSN^{+}01, PLY13, PYKL16, PY19, RSKK19, RIFR10, TAI^{+}11, WLZ^{+}18, XPS^{+}15, ZLH^{+}15, ZLC17a, ZXW16b, ZZL^{+}17a]. group-based [PYKL16, ZZL^{+}17a]. group-choose [PLY13]. group-oriented [WLZ^{+}18]. Group-SPMD [RRR04]. grouping [CLW^{+}18, LLL^{+}19, WCR^{+}14, YL16]. grouping-based [CLW^{+}18]. grouping-proof [YL16]. groups [HPK^{+}18, MZG19, MCXP15, RRR04, SMM18, ZHC^{+}18]. groupware [XPS^{+}15]. growth [cDrLyC^{+}19]. gSched [CLQ^{+}17]. gSET [MWJ^{+}10]. GSM [MB17]. GSWABE [LS15]. guaranteed [ABDO09]. guarantees [ASS08, KD15, LGCJ^{+}13, PSM03, WRLS12]. guessing [FIO15]. Guest [Din09, EN09, McE10, RS13, HdV13, XHCL15, ZQH12]. GUI [QEB^{+}10]. Guided [CGGH17, KHL^{+}17a, SHT^{+}17, WJXZ18]. GUIs [MH07]. GWAP [CWC10]. GYSELA [RLRG15].

H [GCWE15]. H-SOFT [GCWE15]. H.264 [RSMFE^{+}12]. H.264/AVC [RSMFE^{+}12]. H.265 [JML^{+}16]. H5 [GSK19]. HA-PSLS [KM03]. HaDaap [XDJL18]. Hadoop [CLQ^{+}17, CLW^{+}18, JCJ17, KHL^{+}17a, LZJ^{+}18, LL16b, LJL^{+}17, PSC^{+}17, SZT18, SB19a, XDJL18, ZZC^{+}17]. Hadoop-based [PSC^{+}17]. halftoning [KHF^{+}17]. halo [BBW19]. halo-swapping [BBW19]. hand [JTD^{+}19, LLH^{+}17]. Handel [IS10]. Handel-C [IS10]. handlers [YF13]. Handling [AMB^{+}17, AAI12, KW11, PME^{+}08, QLD^{+}11, RCC17, TCBR^{+}10, WFHT17, WK12]. handoff [HZC^{+}14, MJ19, PS19b, WCLH12]. handover [YHHS16]. Hands [WAS07, WC08, Xu08]. haptic [TWZ^{+}19]. haptic-enabled [TWZ^{+}19]. Haralick [WCH^{+}07]. hard [BDW14, BLA^{+}14, LTK17, Puf13, VDB09]. Hardware [SPS17, VEJD17, ABMdAMF19, ADF^{+}13, BHKW12, CJX^{+}19, ER12, FRKS12, GTFA13, GPV09, HJB12, HXY^{+}12, HYGF19, JCVU15, LGdVH13, LGQ^{+}17, LPC^{+}14, MPR04, SHT^{+}17, SRF13, The01, TRW07, TGB^{+}10]. Hardware-assisted [VEJD17, CJX^{+}19]. hardware-aware [BHKW12]. hardware-oriented [TGB^{+}10]. harmonic [Li19, SEF^{+}14]. Hartley [LZJ^{+}18]. harvested [JWZ13]. harvester [GHLS19]. harvesting [CSB^{+}16, OKBO19]. hash [MA15, WTN07]. hashing [CZL12, KSC12, MIGA18, Zha19]. Haskell [TL14]. HBench [ZS01]. HCM [SdSL18]. HCW [BJ18]. HD [DZM^{+}15]. HDD [LHH^{+}17]. HDD/SSD [LHH^{+}17]. HDF5 [EFM17]. HDKV [ZHZ^{+}13]. Head [YESG^{+}19, ESGQ^{+}11]. Head-of-line [YESG^{+}19, ESGQ^{+}11]. header [GBXL17]. heading [LMS18]. healing [AA19, FMS11, GCSB19, MO15]. health [EPA15, RM19, SGL^{+}17, vLDW11, LRS15, LDS^{+}08]. heart [BTCB16, OKP16]. heavy [EOD^{+}19, GZHF19, RVRD10]. heavy-ion [GZHF19]. heavy-tails [RVRD10]. held [HF17]. Hellman [LZC14]. Hello [LLT^{+}14]. Helmholtz [BdL06, LXRJ13, LTZ^{+}19, OLG^{+}15]. help [AG17b]. helpful [GFL04]. helpfulness [ZTM12]. Heritage [PCJ17, GIL17, PC17a]. Heston [BLDW16]. Heterogeneity [BJ18, AMB^{+}17]. Heterogeneous [BJ18, FS18, SF16, VLJ17, VFAD18, YBX^{+}17, AHP^{+}13, ABC^{+}16, AGMR05, ALMT19, AMP^{+}18, Ano06, ATNW11, BFR05, BG14, BCM15, BHQOS15, BAG17, BHKW12, CHP17, CL18, CLQ^{+}17, CW11b, CLT^{+}16, CLF^{+}19, CPXA06, DLPV07, DFTHD18, DLK^{+}18, DKJ13, DL07, ELM^{+}16, EOD^{+}19, EAGVBVDS11, EJF^{+}16, FE18, FGZ^{+}18, FNBS16, FM08, GL19, GVC10, GCPS^{+}14, HCG07, ITK09, KMBR19, KSM15, LLMK18, LLMM19, LBTE14, LWL17, LJL^{+}17, MP17, MJD17, MRS08, MJ19, NZKK11, PSLC11, PPdSTB17, PSC^{+}17, PPP10, PS19b, RBO^{+}02, RMCA12, RCA^{+}12, SRS16, SR19a, SJISVR17, SdSL18, SHC^{+}16, SJPB17, SO16, SSMB15, STL^{+}15, SEF^{+}14, TLF17, WKB^{+}19, XXLL17, XLYL17, XDJL18, XLY^{+}16, YCL11, ZABP18, ZLKK17, ZYZ06, ZWT^{+}18, ZQW^{+}17, VFAD17]. HeteroPar [BJ18, CČJ^{+}16]. heuristic [AMTM17, GCWE15, LBV16, PPST09, SRM13b, YLR^{+}13, ZH15]. heuristics [Ano06, BFR05, BB12, EQW^{+}18, XXLL17, YPLJ11]. HEV [WWX^{+}19]. HEVC [JML^{+}16]. HEVC/H.265 [JML^{+}16]. hexahedral [WO02]. HF [SS19b]. HiBench [SZT18]. HiCOMB [Mar05]. HIDCC [HSJ^{+}18]. hidden [EMEY14, HPD^{+}15]. Hides [DZZL19]. hiding [DWC^{+}15, SM19a]. Hierarchical [LPG^{+}14, RSKK19, TCSBMG17, BJD^{+}19, BDV02, CXT^{+}18, EMEY14, GKSR14, GMMT17, LFZ^{+}17, LBY^{+}16, MRL16, PF12, SS15a, SMM18, SIS19, TW07, VS02, XJZ13, Yos06, ZF18, ZXW16b, ZLA^{+}15]. Hierarchically [GBD16]. hierarchies [DP14]. hierarchy [BP17]. High [AAP13, AP10, BA04, Ber07, BDT01, BXQ17, BDH15, BDH18, DRZ13, DA19, DDE^{+}12, EMEY14, EB14, GL19, GM10, Li18, LSS15, Mar05, MLY10, MB02, NTK08, PHGK10, PW05, PPBB14, PK17, RCB03, RMW19, SG19, SFN12, SFH13, ZKJ^{+}07, AC06, AC08, AKM^{+}06, BCD^{+}10, BHJ^{+}16, BFM^{+}10, BBM19, BDY03, BBD10, BDG^{+}10, BPT^{+}16, BDV02, BPD06, CGBNM17, CAD^{+}18, CLH^{+}11, CLF^{+}17, CLS14, CEG^{+}05, CFP^{+}03, CRGR^{+}12, Dam11, DC19a, DL10, DMD16, DP19, DPK10, DFLL14, DZM^{+}15, DA15, EDB^{+}14, ESG17, EMS15, ETR^{+}13, FPHZ19, FGC06, FMT16, Fox12, FJG^{+}13, GFBR10, GKSR14, GAM17, GBMM15, GCN09, GA08, GDD^{+}04, GVP^{+}14, HDDG09, HLHC12, HLCW15, HYL^{+}19, HY12, IC19, JOC^{+}15, JK13, KDC17, KKTHL13, Kar14a, KSM^{+}08a, KTR11, KOOB15, Kri05, KF11, KWK05, LLMM19, LL05, LCM12, LZG^{+}19]. high [LGL^{+}17, LLH^{+}15, LAL02, MMW16, MJZ17, MZJ^{+}19, MMMP01, MHH16, MDH^{+}16, MPT07, MO02b, MHRI14, MA15, MMG^{+}18, MCC16, Pan20, PSS^{+}19, PGF19, PS19a, QXXZ16, QSX^{+}17, QWZ^{+}19, RVRD10, RCCH19, RCLSK16, SRF13, lSsCY17, SFT15, SKA^{+}14, SRL^{+}14, SS07, SAM^{+}17, SWZ12, TTD^{+}11, TKZQ17, TFG^{+}12, TTPJ16, TSS18, UA18, UGM18, VS02, VJK13, VdSK^{+}05, VFAD18, WFJ^{+}17, WLF19, WHX19, WL02, WK07, XLL^{+}15, XL17, YC19b, ZZ16, ZGS17, ZQD^{+}17, ZLZ^{+}19, ZHZ^{+}13, ZCXH17, ZCL^{+}19, BB02, CCW04, KS02, MP04, RK01, SIOS02]. High-accuracy [EMEY14]. high-bandwidth [GDD^{+}04]. High-contention [BDH18]. high-density [FGC06]. high-dimensional [DP19, HLCW15, KOOB15, LLMM19, MMW16, MJZ17, MZJ^{+}19, RCCH19, SWZ12, ZHZ^{+}13]. high-efficient [CLH^{+}11]. High-end [GM10, CGBNM17, JK13, LGL^{+}17, WL02, ZKJ^{+}07]. high-integrity [KWK05]. High-level [AAP13, DA19, GL19, NTK08, BDV02, CAD^{+}18, MHH16, MPT07, MCC16, VFAD18]. high-order [ZCXH17]. High-Performance [Ber07, MLY10, PW05, SG19, AP10, BDT01, BDH15, DRZ13, LSS15, MB02, PPBB14, RCB03, RMW19, AC06, AC08, AKM^{+}06, BHJ^{+}16, BFM^{+}10, BBM19, BPD06, CEG^{+}05, CFP^{+}03, CRGR^{+}12, Dam11, DMD16, DZM^{+}15, ESG17, FJG^{+}13, GFBR10, GBMM15, GCN09, GA08, GVP^{+}14, HDDG09, HLHC12, HY12, KDC17, KSM^{+}08a, KTR11, LL05, LLH^{+}15, LAL02, MMMP01, MDH^{+}16, PSS^{+}19, QWZ^{+}19, RVRD10, SFT15, SS07, TTD^{+}11, TTPJ16, VS02, VdSK^{+}05, WFJ^{+}17, WK07, ZLZ^{+}19]. high-precision [ZCL^{+}19]. high-productivity [TFG^{+}12]. high-quality [CLF^{+}17]. high-resolution [BDY03, HYL^{+}19]. high-rise [YC19b]. High-speed [ZKJ^{+}07, DPK10, DA15, LZG^{+}19, UGM18, WLF19, ZGS17]. High-throughput [EB14, EDB^{+}14, FMT16, JOC^{+}15, Kri05, PGF19, SKA^{+}14, SAM^{+}17]. high-volume [MHRI14]. higher [Air17, BBSW17, JMF09]. Highly [MKAKG14, SC19, BWHS18, DCK12, HKVW16, KS19a, KM03, KSS^{+}17, KHL17b, LLMK18, TCP^{+}05, VCP16]. highways [SLC^{+}18, SLC^{+}19]. Hilbert [KHHC13, SB19a]. Hilbert-order [KHHC13]. hill [YT19]. HiPINEB [SG19]. HIRLAM [VCW13]. Hirschberg [JSR19b]. histogram [VYKM19]. hits [SSV19]. hitting [CSMS^{+}19]. HKE [LBY^{+}16]. HKE-BC [LBY^{+}16]. HLA [DBR13, FAPC16, MT08, ZG04]. HLA-based [DBR13, FAPC16]. HLog [LSP15]. Hmm [HPD^{+}15]. Hoare [vO01]. hoc [CNPP09, Den07, DK19a, DA15, EB10, HKA^{+}15, IHB15, KOO12, KKK10, KABD07, MTM19, MSMA19, MLRR09, PY19, QWW^{+}16, Sha15, SK17, WXSH19, YWM^{+}10]. HOG [KS19a]. HOL [Sch04, vO01]. hole [SWS^{+}18, WWY19]. holistic [GVP^{+}14, PAC^{+}17]. holonic [FD01]. home [PBD^{+}15, WT18, LMOT10]. home-therapy [PBD^{+}15]. Homomorphic [Tan15, CZL12]. Honey [MS17b]. hop [BAT13, DZ13, JKZ03, MS07, MA15]. hormone [PB12, Pac16]. Horovod [KSM^{+}19]. hospitality [SVB19]. hospitals [LHB^{+}19]. host [LLRS03, TMZ07]. host-parasite [LLRS03]. Hot [Man08, LLX15b, VKM^{+}09]. hot-spot [VKM^{+}09]. hotness [XDJL18]. hotness-aware [XDJL18]. hotspot [IRB19]. hours [LWSZ19]. HPC [CS15, BRK^{+}17, BHJ^{+}16, BBdS^{+}17, BDP^{+}14, CGST17, GMVRGS15, KMRT18, KAM11, LBJ^{+}19, LME^{+}19, LMS18, MRL16, MOF15, MSPDP19, MAK18, OTG^{+}07, PCB^{+}18, dRRdCRR16, RHL^{+}18, SWD^{+}15, SKNH09, Spa19, TZG^{+}19, TGB^{+}10]. HPC-Europa [OTG^{+}07]. HPC-GAP [BHJ^{+}16]. HPC-optimized [MSPDP19]. HPCG [KB18]. HPCS [SCC^{+}10]. HPCT [ABF^{+}10]. HPCx [ABG^{+}05]. HPF [DS02, FSPC^{+}02, ISKvW02, MAH^{+}02, NNON02, Ogi02, OA02, PSG03, SM02, SIOS02, vWAH^{+}02]. HPF/JA [ISKvW02, Ogi02, SIOS02]. HPF/SX [MAH^{+}02]. HPGMG [KB18]. HPJava [LCFkL05]. HPL [BCD^{+}10]. HSA [MD19]. HSA-based [MD19]. HTC [FMT16, LHL10]. HTC-Sim [FMT16]. HTCaaS [KQR^{+}17]. HTML5 [GSK19]. HTTP [GBXL17]. Huard [BEQOR17, Has17]. Hub [GBB^{+}15, LWY^{+}17, ZDL19]. hub-and-spoke [ZDL19]. HUBzero [MCD^{+}15]. huge [LTKF11]. human [Bao19, BCA^{+}10, CPG^{+}16, Che18, CGI17, GQR16, KS19a, KSM^{+}08a, LGQ^{+}17, LSS15, RK15, ZLQ^{+}18]. human-computer [CPG^{+}16]. Humanics [Nis18]. humans [GOLL17]. HVM [LTK17]. HWMP [BOB13]. Hybrid [Akt18a, ABB^{+}15, ZJT^{+}19, AP10, AR16, AML^{+}15, AMABS18, AFG16, BYDC19, BLDW16, CSMS^{+}19, CLT^{+}16, CCW^{+}15, CAKH17, CKRO13, DFG18a, Den07, DS19, EAGVBVDS11, FMS15, FTT15, GGFPGB14, GKS09, GKSR14, GdMK^{+}18, Has17, HSJ^{+}18, HR06, HAA^{+}17, KB18, LM08, LLB04, LG08, LGL16b, LHH^{+}17, MLS^{+}15, MB17, MB12, MB14, MJD15, NO02, PCGE18, PSS^{+}18, RM03, SJVR15, SS19b, SB18, SD15, TAMC19, THF15, TYL^{+}15, TAI^{+}11, WZJD13, WDG^{+}14, WWLD18, XDE^{+}04, XDP18, YWC11, ZK08, ZS17, ZLW^{+}19, ZCD^{+}12, ZCL^{+}18, BOB13, PGW06]. hybrid-enhanced [FTT15]. hybridism [BPL12]. HybridMR [THF15]. Hydra [COdO^{+}11, PA08]. hydraulic [MO02a]. HypE [SR19a]. Hyper [ZQK15, AMTM17]. hyper-heuristic [AMTM17]. Hyper-star [ZQK15]. hypercube [BDF15, QLLS15, XYL18]. hypercube-based [QLLS15]. hypercube-like [XYL18]. Hypercubic [HP11, Li04]. hypergraphs [AMP^{+}18]. hyperspectral [Pla08, PPP10, SPMP11]. hypersurfaces [RMCN^{+}07]. hypervisors [GVP^{+}14]. hypocenter [ZMJZ10]. hypre [EFY17]. i-Computing [WLL03a, WLL03b]. I/O [EFM17, BFL^{+}10, BKND16, BDT01, Hic18, JCJ17, LGL^{+}17, LQL^{+}09, LLT^{+}14, PLM^{+}19, PLW^{+}18, WTL^{+}16, YLL^{+}18]. I/O-intensive [PLW^{+}18]. IaaS [GCZ^{+}17, LLL15, RB17]. IBC [BOB13]. IBC-HWMP [BOB13]. IBD [Riz04]. Ibis [vNMW^{+}05]. IBM [SLM^{+}10, ZDC^{+}09]. ICA [LLQL14]. ice [ZGL19]. ICEC2017 [Du18c]. ICICTA [Hou12]. ICWM2009 [CW11a]. ID [YCSY19]. ID-based [YCSY19]. IDC [HYQ17]. Identification [GLC^{+}04, KGP^{+}19, ANH16, AD15, CZQ17, CXT^{+}18, COdO^{+}11, DZL^{+}19, DA19, EA12, HL19, HPK^{+}18, LLT19b, LGQS12, LLLS18, LFW19, LHZS19, LL19b, LLQL14, MHLC^{+}05, NLYZ12, Wan18a, XBXS13, YGW17, YHYY19, YZW^{+}15, ZLG^{+}19, Zhe16]. Identifying [QZYZ16, SSC^{+}16]. Identity [SWW^{+}16, ATKH^{+}17, BOB13, GPVCdBRO12, HZC^{+}14, SWZ12, VRDTB^{+}16, WLFX17, YWL^{+}17a, YNX^{+}16]. Identity-based [SWW^{+}16, BOB13, HZC^{+}14, WLFX17, YWL^{+}17a]. idle [GKG^{+}04]. IDP [CEM^{+}17]. IEEE [Mar05, PCC17, BOB13, HON04, MP04, WCLH12]. IEEE802.16e [CLH13]. IGSIM [SB19a]. II [SLM05]. illuminating [HTR10]. ILU [ABF^{+}17]. ILU-preconditioned [ABF^{+}17]. IMA [XHCL15]. Image [EMS11, SMBT07, XYSW18, AT01, ASWR12, BÇG14, CRB09, Che18, Du18a, EMEY14, HAK19, HLY18, JZMD19, JZZL06, KSN16, KMG^{+}18, KSM15, LLJ18a, LLJ18b, LYL^{+}19, LC18, LHZS19, MWL^{+}15, PLL14, PPP10, PJW^{+}14, SK04, TVCB19, WJ12, WLW14, WBD^{+}03, WW08, WCH^{+}07, Xu19, YWL^{+}17a, YHJ^{+}14, ZXW16a, ZWL^{+}13, ZGZ19]. image-based [ASWR12]. image-processing [JZZL06]. ImageCL [FE18]. imagery [Pla08, WCZ^{+}18]. images [Boc19, CZL^{+}17, GKP^{+}19, He19, SPMP11, SS19b, ZBZ^{+}15, ZWW14]. imaging [KSG11, PVR^{+}09, PLL17, WKB^{+}19]. imbalance [KHL17b, LL18, UGM18]. immunodominance [ZWL^{+}17]. Immutability [PS05]. immutable [NN07]. IMP [GBB^{+}15]. Impact [DS04, MB17, GSR^{+}19, LGJ17, PB19a, QCB17, ROQL18, SM09, SK18, UGM18, XHW^{+}19, ZDC^{+}09]. Impacts [LLY^{+}19a, HSV^{+}19, LLT19b]. Impala [CZY^{+}18]. imperative [SPBL06]. imperfect [KT19a, LWG^{+}15]. imperialist [MSD^{+}18]. implement [MÍM19, SNB^{+}01, Slo06]. Implementation [CL18, CLF^{+}19, KSM15, KD07, MLVB05, MAH^{+}02, PB12, SER15, SS19a, SLM^{+}10, TKA^{+}02, TMAG03, ACGG06, AAB^{+}05, AFT01, BPdM06, BDY02, BDV02, CC13, CHX^{+}19, CKNW06, DRZ13, DPST06, DL10, DBB^{+}16, FO18, GG09, GSB^{+}12, GCN09, GLC07, GPV09, GAB19, HG11, KL19, KWL^{+}04, KKJH03, KBVP07, KVGH11, Kri05, KN19, LDZ14b, LLH^{+}18, LRS15, MMO^{+}16, MKKB04, MWL18, MDL^{+}10, MRS^{+}09, PSG03, PMAL14, PDC16, RMW19, RGB^{+}15, SBJ^{+}15, lSsCY17, SACJ04, TALT16, TKB16, VDL^{+}15, VHBB03, WLLL15, WWLD18, YP10, YYC10, YCL11, YGG14, ZYW^{+}16]. implementations [AA16, BDM18, CACC11, CCW04, DS04, DDF^{+}15, ER12, HPVRPF14, LLdA08, TL14, WCC04, YBB^{+}07]. Implementing [CJX^{+}19, CKNW06, DLK^{+}18, JM19, KKV13, YL01, NNON02, PCT04, PGF19, RBB^{+}09, MRH14]. implications [WLZ11]. implicit [RSM01]. importance [BMPS07, SCLL19]. Impossible [WYL14]. Improve [LS19b, DMRS15, DDF^{+}17, GIVRC^{+}10, PMG19, SAB15, TRW07, ZSWS18, ZCW^{+}18, ZZD^{+}17, ZJL15]. Improved [RF15, WLZ^{+}18, YHHS16, YLWZ18, ZZL^{+}19, BV16, CBHTE11, EB18, JZMD19, KKK10, LXYC17, LWK15, LLY19c, MMO^{+}16, MS19, Nam19, PLZ14, SB19a, cWsThY19, WBC^{+}17, WZYG19, XHH12, XLG19, YSQM19, YZXW17, ZWW14]. improvement [CLL^{+}18, CL16, DXWC16, DLT^{+}16, MÍM19, MJ19, TWN07, XLZD13]. improvements [EA12, SVN12]. improves [ORDG15]. Improving [AYN^{+}14, BN19, CLH13, Cog03, ET15, FBS16, GA09, Hic18, HTI05, KCB09, LLD19, LSH^{+}16, LL19c, MSMA19, MB16, RSC^{+}15, WCL^{+}10, WCLH12, XJAJ18, YFL18, YYS15, ABMdAMF19, AAHA18, CPG^{+}16, CLS14, HKRR08, MM19, RMCHMG15, TCSBMG17, Tru15, XDJL18]. Improvised [RP19]. in-core [BGGL07]. In-memory [SGD^{+}18, ZJS^{+}17, MY17, SGJ^{+}17]. In-network [XGC19]. In-place [LTL^{+}17, DVL13, GW19, PSHL11]. In-VIGO [MTA^{+}07]. in-VM [ZYFZ19]. inAspect [ASS^{+}05]. Incentive [ZXXN06, CLW^{+}15, MZW^{+}16, MME13, WLP^{+}17]. Incentive-based [ZXXN06]. incentives [LPY^{+}08]. Incentivising [PRP^{+}15]. incident [GQH17, RCC17]. inclusive [DWC09]. incomplete [YYXL19, ZWXS19]. incompressible [HKB07, ZYW^{+}16]. Incorporating [WFJ^{+}17, XLZD13, HmLGP03, LMH^{+}14, vdKEL10]. increased [YS07]. Increasing [CLZ^{+}17, PHCR09, Hey19]. Incremental [BM07, Rav16, TSKM18, LWT^{+}16, TJ17a, WHX19]. indefinite [BDR^{+}17, YTD17]. independent [BKSM^{+}15, CDMS15, GPW03, LLLS18, PFC14]. index [DKMM14, HCC^{+}15, LW13, Qia19, SER15, TPV17, XHD^{+}19]. indexes [YWBS19]. indexing [ATSAK15, DXG13, MIGA18, PPR19, RM19, ZHW^{+}16]. indicator [PRD^{+}13]. indicators [DPS16, ZP19]. indices [WXLM19]. indirect [PGL^{+}17]. indirection [LGFM05]. Indirectly [CKSC10]. indiscriminate [YSC^{+}17]. individually [LF15]. indoor [KBH15a, MB17, YYLC19]. induced [ALYD17, GZHF19]. induction [WKB^{+}19]. inductive [FMS11]. industrial [JKZ03]. industry [Air17, QGZL18, ZQD^{+}17]. inefficiency [WMDM07]. inertia [LWL^{+}19]. inference [MKAKG14, SJVR15, SLM04, SLM05, TVCB18]. inferring [SMM18]. InfiniBand [VKM^{+}09, ZJKL10]. influence [CHZ12, GRS^{+}17, HLW^{+}19, SMS^{+}19]. influential [MLZ19]. informatics [TTR^{+}10, vLDW11]. Information [Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano19a]. Information [Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Che19, ECP18, LB19, Pie08, SARL13, TKS18, WYW^{+}17, ZC19, AI17, AP10, AR16, And13, ASG^{+}08, BDL^{+}15, CZ15a, CW09, DK19b, GD08, HAN19, HKA19a, HSM14, JB19, KT19a, KKK^{+}19, KKW^{+}14, KSC12, KTM^{+}09, LLKC08, LWG^{+}15, LWZ^{+}17, LSZ19, LFW19, LLY19c, MYS19, MLRR09, MWL18, Nis18, PHY^{+}18, PVCS18, PLR^{+}14, PAM^{+}15, PME^{+}08, QMK12, SCLL19, SWZ^{+}18, SW11, TWW^{+}19, TMS^{+}12, WAD12, Yan19a, Zen19b, ZCLL19, Boe12, HF17, LWL^{+}06]. information-based [KSC12, SWZ^{+}18]. Informatization [HH19]. Infostation [TW07]. infrastructure [ACMA07, AJY^{+}15, ANK^{+}17, CRC^{+}15b, CZO^{+}08, CWMZ06, CPSP17, CPE^{+}19, CMS17, Cyb06, DMA13, JvAB^{+}15, JKL^{+}17, JQSP08, KMJ14, KA11, MCWL06, MPT07, MPVT17, MP03, PFU^{+}05, PCH^{+}08, SACJ04, WWL^{+}15, WSP17, YDB^{+}13, ZWL^{+}13, ABB^{+}15, DR15, WLR05]. infrastructure-as-a-service [CMS17]. infrastructure-less [DMA13]. infrastructures [AWR17, ACG18, AFG16, CSMB15, CHL15, CXPL15, GWVP^{+}14, GKP^{+}09, Ios11, Kac11, LBV16, LSMVML15, MVML11, RLS^{+}09, THF15]. ingestion [SIST18]. inhalable [MZG19]. inherently [KA16]. inheritance [Lyo02]. inhibiting [BGGS14]. Initial [VDL^{+}15, MRS^{+}09, RBBH02]. initiated [AR16]. injection [GWHJL19]. inlining [LH05]. innovation [YLLC18]. Innovations [ACD02]. innovative [DS17, HF17]. inpainting [HLY18]. inpatients [LHB^{+}19]. Input [TJD^{+}17, AAI12]. input/output [AAI12]. inrush [YHYY19]. insider [DCG15]. Insights [HLX^{+}16, WLZ17]. inspection [HLG17, LSC^{+}19, YYZH19]. inspired [ABG^{+}13, CSL12, CP14, CT16, GPVCdBRO12, HAE09, OK18, PCS^{+}12, SR19a, SR17, TTPJ16, WSL15, ZLQ^{+}18]. installation [CGGH17, HLA^{+}18]. installations [LME^{+}19]. installment [DL07]. instance [KCKC15, MCWL06, TKB16, XWH^{+}17]. instance-intensive [XWH^{+}17]. instance-oriented [MCWL06]. instances [Ios11, VRDTB^{+}16, LMH^{+}14]. instantiation [CSC^{+}17]. instantiations [KCB09]. institutions [LGJ17]. Instruction [GSG06, LHC14]. instructions [AB01, PBSB04]. instrument [MH07]. instrumentation [BDMM^{+}05, RS07]. instruments [MH07]. insulated [LDZ^{+}14a]. insurance [GQH17, YWLQ18]. integer [GLM^{+}16, KVGH11]. InteGrade [CML^{+}10, CC10, GKG^{+}04, dCGKG06]. integrals [LLMM19]. Integrated [LWZ^{+}19, ABC^{+}08a, AMSR14, AFR09, BAC^{+}15, Fec12, GKS^{+}07, GLC^{+}04, HCD^{+}18, JZZL06, KB06, LZC09, PXY^{+}07, ROA^{+}07, Sch02, SB19a, VDL^{+}15, YGL05, YP10, ZBZ^{+}18]. integrates [SAM^{+}17]. Integrating [AP06, CRC15a, MML^{+}17, ZKR^{+}07, BGV^{+}01, BHW05, CLX07, DCY^{+}08, HCG07, MCD^{+}15, SKA^{+}14]. Integration [DvdS06, FHO^{+}15, SM03, TPV17, Boc19, BDV02, CLH^{+}08, GMPT15, GD08, GSK19, Qia19, RJ01, SGD^{+}18, SZR16, SS15c, WQL^{+}18, XLY^{+}16]. Integrity [AaBT17, AL04, BC16, CJZZ10, KWK05, SWW^{+}16, WZL^{+}17b, XHCL15, YNX^{+}16]. Intel [AB01, CLRB15, CBIGL19, DAC^{+}18, FNBS16, HCD^{+}18, MCP^{+}12, RGB^{+}15, SWB12, Tan12, VDL^{+}15]. Intelligence [Du18c, CH19, DCCZ18, JZMD19, PCS^{+}12]. Intelligent [BM12, BFVRC15, COC18, DDF16, ESZ09, KSK19, VC16, VZB19, YSWZ17, ZL19, DBGA16, DWG19, Hus15, KSN16, KKT13, KZY^{+}18, KN19, LXL^{+}09, SZG^{+}19, ULS03, WZT11, XCHK14, Zhu18, Bai17, HYQ17]. Intel(R) [GdMK^{+}18]. intensity [LLZ^{+}17b]. intensive [AMGCC17, CBHTE11, CGBNM17, CTAB16, GGHR16, HAAWA^{+}16, HZHP09, JKL^{+}17, LHHJ18, LCYJ08, MWL^{+}13, PLW^{+}18, RMCN^{+}07, SAdB^{+}16, TKA^{+}02, VJHB05, WSL15, WQS^{+}16, XWH^{+}17, YR15, ZWL^{+}15, ZZ16, ZWF^{+}06]. intent [KGK17]. intention [LL19a]. inter [HJB12]. inter-node [HJB12]. interacting [NAK^{+}15]. interaction [BPdM06, CPG^{+}16, HC07, IÁBE11, JSPE15, LL19b, MMMP01, MB14, MP03, YZR14, ZP06]. interactions [JQSP08, RCT03]. Interactive [VYK^{+}10, WJ09, WST^{+}17, CEH^{+}06, CZWH07, CLW^{+}18, GRGP12, HHWZ08, HLZD18, IBvA^{+}02, KTB04, MCY^{+}10, PWC^{+}14, PML^{+}05, VŠC17, ZZ18, vSB06]. interceptor [Ang08]. Interceptors [NMMS01, BMV03]. interchange [BBG17, ZQZ^{+}16]. interconnect [FAM^{+}18, GDD^{+}04]. Interconnection [AS15, NZKK11, SG19, ESG17, JAA08, KMA04, QLLS15, ZLF19]. Interconnects [AS19, CKRO13, PLC^{+}19]. interdependent [Sha15]. interdisciplinary [CN02]. Interest [ZACG16, CRC15a, DCJ14, SZT19, ZJT^{+}19]. interested [DK19b, XY17]. interesting [LJPP16]. Interface [KKJH03, AJMJS05, dRADFG17, DBB^{+}16, GRW^{+}19, GHB^{+}06, HRR^{+}11, Jac02, KOB01, OORVB14, ULS03, WCZ^{+}18, WKL14, AMHC11, SWL^{+}01]. Interfaces [WD07, LOKW^{+}10, vHKT^{+}11]. interfacing [ASS^{+}05]. interference [FPHZ19, LXK^{+}19, Pan20, VP19, WLL14, XHW^{+}19, YYZ^{+}17, ZKWK17]. interference-aware [VP19]. InterGrid [dABV08]. interleaved [GSG06]. interleaver [IC19]. intermediate [PGL^{+}17, YXLZ16, YYL^{+}12]. internal [ABFL17, KJ19b]. International [Ang07, BL17, BJ18, CR08, CL08, CC09, CW11a, CR13, Che19, CS06, DR15, Du18c, FZ08, GJ17, GZX17, IUCH^{+}17, Kni06, LB19, Mar05, PC17b, WCWB19, Wu18, CL13, Fox17a, PCC17, WT15, AF14, HF17]. Internet [ABMMR19, AA19, AD15, BKS18, CHX^{+}19, CLH19, CM18, DFG18a, Den07, GED^{+}18, GTA10, HSL19, IAH^{+}15, KJ19a, KC18, Li17, LWW06, LTKF11, LLY19c, MK15b, MB15, PCJ17, PJ18, RS13, RMCN^{+}07, RO12b, SMH^{+}19, SS15b, SRN^{+}15, XPWF15, Yu18, ZlC15, ZZY^{+}15, ZLC17a, dMd^{+}17].

Internet-based [RMCN^{+}07]. Internet-of-Things [KC18, MK15b]. internetworking [dABV08]. interoperability [ET09, GLC07, ZBC^{+}07]. interoperable [FÁBE11, MP02]. Interoperating [CHL15]. Interoperation [RLS^{+}09, HAA^{+}07]. interplay [SD11a]. interpolation [DMC^{+}18, MAVG16, YLWZ18]. interprediction [RSMFE^{+}12]. interpretable [XDP18]. interprocess [TV14]. intersection [Eng15, LZY^{+}16, YLWZ18]. interval [FLMRC02, LRLY17, SK18]. intra [HJB12, XPS^{+}15, CRC^{+}15b]. intra-group [XPS^{+}15]. intra-node [HJB12]. Intra-Operative [CRC^{+}15b]. intraclass [KS19a]. intrinsics [KL12b]. Introducing [JKM^{+}17, JKL^{+}17]. Introduction [HTBR12, HTW14, Pie08, PDD14, RHT13, Run10, SHT11, VK12, ZQH12]. intrusion [HSJ^{+}18, LLL15, RG17, SPW09, SM19b, WLZ11, WML^{+}19, WZXZ12]. intrusive [CLW^{+}19]. intuitive [GvHKK11, RRWS08]. invariants [CMD17]. inventory [LXP^{+}12]. inverse [GG09, PV04, PLZ14]. inverses [GE08]. inversion [BEQOR13, BEQOR17, RSTV05, SGR19, WZYG19]. inverter [GZHF19]. investigate [WJT^{+}14]. investigated [WWY19]. Investigating [HJV^{+}19]. Investigation [YWA07, BDW14, HK01, KKK10, TDC18, ZDHJ18]. invocation [MKB01, SS19c, BVGVEAFG11, NMMS01]. invoking [OK15]. IO [DL10, LGG16]. Ion [KF11, GZHF19]. IoT [IAH^{+}15, PCJ17, AMB^{+}17, BJC17, CKL19, CEB^{+}18, CDP17, DZZL19, GRTX18, GIL17, KM19, PC17a, SR19b, SCS17b, WLZ^{+}18, YHYY19, ZKWK17]. IoT-based [BJC17, GIL17]. IoTs [KKK^{+}19]. IP [PCsHL18, WHX19, YJL12, ZLG^{+}19]. ipcmd [WKL14]. iPlant [LGD15, WWL^{+}15]. IPMI [KD19]. iPortal [KBH^{+}15b]. IPv6 [DEF08, HLX^{+}16, ORdSL13]. IPv6-enabled [ORdSL13]. IQ [CEH^{+}06]. IQ-Services [CEH^{+}06]. irregular [AAF^{+}07, GdMK^{+}18, GPZ04, HR06, KR04, LYL07, Nev17, YWL^{+}17a]. IS-FMIPv6 [WCLH12]. ISABELA [LSE^{+}13]. Isabelle [Sch04, vO01]. Isabelle/HOL [Sch04, vO01]. ISCOPE [Fox05]. ISENGARD [KA11]. iShare [WTL^{+}16]. island [LF17]. islands [dABV08]. isolated [KD10, ZZD^{+}17]. isolation [CRB^{+}17, KMBR19, WTL^{+}16]. isolation-based [CRB^{+}17]. isolation-trees [KMBR19]. isosurface [DCG11]. Israeli [Guo19]. isssue [OKG18]. Issue [AS19, AHP^{+}13, Ang07, Ano02, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g]. Issue [Ano18h, Ano18i, Ano18j, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, AM07, BA04, BM12, BL17, BHD13, BM04, Ber07, BKZ^{+}13, BDB^{+}13, BL09a, BL09b, BL11a, BL11b, BL13b, BL13a, CWZL13, CCCW13, CČJ^{+}16, dOCPFJ13, CLTT13, CR08, CC09, CW11a, CKRO13, CAG^{+}13, CS09, CS06, CMT13, CM07b, CS13, DRZ13, DRS^{+}13, DVL13, DDE^{+}12, DLM13, DH13, EBMD13, EH18, ETR^{+}13, Fed13, FK19, FN13, Fox01, Fox05, FG06, FZ07, FS07, FZ08, GG07, GM10, GvHKK11, GZX17, GMF01, GHPR05, HL13, HQoS11, HF05, HTBR12, HTW14, HMPPT13, HFTQ13, JJGL13, JX06, KS02, KM13, KR06, Kni06, KB12, Lee09]. Issue [LBS15, LXRJ13, LMKT13, LB19, LV12, LDXC13, LW13, MWL^{+}13, MS13, Man08, MSP^{+}13, Mar05, MFG^{+}13, MISV13, MLY10, MN10, MLA^{+}08, NPM19, Nar05, Nel05, NSSAK13, ODS^{+}13, OM06b, PLY13, Par02, PRD^{+}13, PHGK10, PW05, Pie08, PB07b, PK08, Puf13, Qiu11, QFT14, QLS13, RMP^{+}13a, RHRB13, RK01, RTMZ13, Run10, SHT11, SN06, SCNH07, SANB08, SRdS09, SF10, SRF13, SFN12, SD11b, TM01, Tho07, TH19, TH10, TWB13, TFDA07, Tur04, Ur07, VK12, VCW13, WAS07, WZZL13, WC08, WCLC13, WD07, WDM14, Wis02, XZ09, XLWZ11, XBXS13, XW13, Xu08, XJZ13, YLD13, YLR^{+}13, YLJZ13, ZWL^{+}13, ZLY^{+}13, ZLN^{+}13, Zha08, ZYH09, ZYH12, ZHZ^{+}13, ZL09, vdS06b, AF14, CL08, CR13, CL13, Che19, DN19, DC19a, DKJ16, Du18c, EL01, ESG17, Fox17a, GTGT11, GWD15, HLX^{+}16, HYQ17, HdV13]. issue [HF17, Hus15, KWXY18, LBW14, LBT16, LBT17, LBFS17, LL13, MYS19, MH18, OEP^{+}15, PDD14, PCC17, QLL10, RHT13, TP14, TSS18, TBH^{+}18, UA18, WAD12, WR17, WCWB19, WDGK15, Xha18, XBCW19, XYS17, ZRB19, ZZ17, ZC19]. Issues [Nel05, vdS06a, AAI12, DP14, GB07, GLC07, MCCG11, SWHL16, YJZZ19]. Itanium [JLT06]. item [LH17, ZSZ^{+}14]. item-based [LH17]. items [CT11a]. itemset [LXYC17]. itemsets [HMM^{+}09]. iteration [TYL^{+}15]. iteration-based [TYL^{+}15]. Iterative [SAD13, AYN^{+}14, AAC^{+}15, ADF^{+}19, CSTV06, EDSV09, GSV03, HC07, JSS07, KKGO04, LLB04, NO02, Nak02, PSRR14, RPRG17, XWX^{+}19, YGG14, YT19, ZW09]. ITME2016 [Che19]. IVM [GMMT17]. IVM-based [GMMT17].

J2EE [BG04]. JA [ISKvW02, Ogi02, SIOS02]. JAC [HL06]. Jaccard [YWR^{+}19]. Jacobi [ADF^{+}19, KYBV17, RR11]. JADE [FRdOR^{+}19]. Jalfon [Guo19]. JaMP [KBVP07]. Japanese [SM02]. JASAG [AAV^{+}15]. Java [Fox01, Fox05, Fox17a, HTW14, VK12, KvGS^{+}14, SAdB^{+}16, AJMJS05, AK01, ASS^{+}05, AdSCdR^{+}19, AFT01, Bac03, BVGVEA11, BVGVEAFG11, BHW05, BDT01, BP03, BK05, BSB^{+}03, CM05, CG01, Cog03, Cog04, DVL13, EFG^{+}03, EL01, EABVGV14, ETR^{+}15, FR02, FT06, Fox17b, GYB^{+}11, GE08, GPW03, GPW05, GS04b, HL13, HL06, HYX05, KHM^{+}11a, KOB01, KBVP07, KSR14, KW01, KWK05, LH05, LXP18, LAL02, LLdA08, LSW07, LWC17, LTK17, LGFM05, Lyo02, MLVB05, MCY^{+}10, MGS19, MMG03, NMMS01, NC05, NMKB03, OGA^{+}01, PSM03, PPMH15, PSW11, Puf13, RTET15, RS12, RHT13, RCB03, RR01, Sch04, SDH^{+}17, SCBH09, SM03, SGV12, SKR17, SPS17, SS19c, TTD^{+}11, VDPC03, VHBB03, WCCL05, WJH06, WBM^{+}10, WK12, WCC04, XHH12, YP10, YKA^{+}19, ZS01, ZYZ06, vHMB08, vNMW^{+}05, vRKS03, vRS05]. Java [vLFGL01, vLGL^{+}02, vO01]. Java-based [AK01, MCY^{+}10, NC05, vNMW^{+}05]. JavaBeans [LR05, YAA07]. JavaNws [KW01]. JavaScript [GGC19, MGI17, VCP16]. JavaSymphony [FJ05]. Jaya [GAS19]. Jcluster [ZYZ06]. JCSP [WBM^{+}10]. Jeeg [MS05]. JEL [DvNM^{+}11a]. Jenkins [BBA18]. Jetstream [HSV^{+}19]. JI.FI [BLA^{+}14]. Jigsaw [CWL03]. Jim [Bou13]. JIT [GE06]. JML [MPHL03]. Job [BWW^{+}08, KSM^{+}08b, NNK^{+}07, BLSP11, EGGA^{+}04, GQ04, Jon09, KWL^{+}04, LL16b, MWRK18, NvV09, RMCHMG15, SR17, Sod05, TZK16, WGZL06, YCL11, ZF14]. job-centric [KSM^{+}08b]. job-scheduling [SR17]. jobs [CNP^{+}15, LGCJ^{+}13]. Join [PN19, ABMdAMF19, LFZ07, MJZ17, MZJ^{+}19, RR15, ABMdAMF19]. joins [BG17, MMW16, RCCH19, WJXZ18]. Joint [dRRdCRR16, ASE^{+}17]. Joint-analysis [dRRdCRR16]. JOPI [AJMJS05]. Jordan [KS19c, Tan12]. JParEnt [SJPB17]. JPEG [SHC^{+}16, SJPB17]. JR [OKW18]. JS4Cloud [MTT15]. JTRES [HTW14, SHT11]. JTRES2011 [RHT13]. JTRES2013 [Fox17b]. JTRES2014 [Fox17a]. Jump [WSWL12]. Jump-start [WSWL12]. Jupyter [YLH^{+}19]. just [MGI17]. just-in-time [MGI17]. JVM [CG01, SD03]. JVM98 [GPW05]. JVMs [STWSP12].

K-means [DLX^{+}16]. K-model [HY12]. Kahan [HFR^{+}17]. Kahan-enhanced [HFR^{+}17]. Kalman [BY12, PA18]. Kano [LL19b]. Karma [SPG08]. Kautz [ZLLL11]. Kava [Bac03]. KDDML [RSTV07]. KDDML-G [RSTV07]. Kepler [LGP19, LSH^{+}16, LAB^{+}06, NSN^{+}17]. kernel [DGR^{+}07, KNT^{+}01, KW18, NRR15, PZH^{+}15, SAP16, EFM17]. KernelHive [RCLSK16]. kernels [dOCPFJ13, CKL17, CBB^{+}19, FVLS15, FJG^{+}13, KCB09, NRR15, RBB^{+}09, SWD^{+}15, VS02]. Key [GSWJ19, WLZ17, ATKH^{+}17, AKG13, AYSZ14, BJD^{+}19, BZD16, BBB16, CLZ^{+}17, FIO15, HWQ^{+}16, KKDS19, KTZ^{+}18, KDW^{+}17, LYS18, LDZ^{+}14a, LZW^{+}16, LZC14, LBY^{+}16, PY19, QWW^{+}16, RSKK19, SCS17a, SGCG09, XW13, XXCY19, YHHS16, ZZC15, ZLC17a, ZHZ^{+}13, ZXW16b, ZFXJ19]. key-insulated [LDZ^{+}14a]. key-value [LZW^{+}16, ZHZ^{+}13]. Keystroke [AaBT16]. KeyValueServe [DIM18]. keyword [BZD16, CLH^{+}16, DXG13, YZCT17]. keyword-based [DXG13]. keywords [LFX^{+}08]. KFCM [WWX^{+}19]. kidney [CLH^{+}08]. kind [GG19]. Kinematic [ZWZ^{+}18]. Kirchhoff [AdCPdSD17]. kit [vLFGL01, Nov02, PvLV^{+}02, vLGL^{+}02]. KLONOS [DHH^{+}13]. knapsack [QW17, SSIH19]. Knights [DAC^{+}18]. Knijnenburg [OS09]. KNL [HLA^{+}18]. Knowledge [FZ08, LM07, SD11b, ZLAa^{+}17, ZL09, BM16, Can06, CLTT13, Cuz11, FGP^{+}11, KBDA19, KCG^{+}19, LHXY08, LXL^{+}09, NZKK11, OKO18, RSTV07, SKA^{+}14, SZL09, WLDL08, WQL^{+}18, WGY^{+}19, XZH^{+}16, YTF^{+}01, ZGL07, Zhu15, ZS19, FS07, LFH^{+}08b, NZKK11, SPR^{+}07, TMZ07, ZL06, Zhu07, ZDL07]. knowledge-based [Can06, KBDA19, OKO18, YTF^{+}01]. known [PCsHL18]. KOALA [ME08]. Krylov [MKSS16]. Kunlun [ZGRSC10]. KWATT [QEB^{+}10].

L [LFG05, EMS11]. label [AMZ19, DZL^{+}19]. labels [HZL19]. Laboratory [BBGA03]. Labs [GBB^{+}15, TBTZ18]. labyrinth [ZDHJ18]. LAF [DXG13]. lake [SAM^{+}17]. Lambda [KCW09]. land [WYY^{+}19]. Landing [DAC^{+}18]. Landing-based [DAC^{+}18]. Landis [BBB16]. landscape [KHM^{+}11b]. Language [FE18, WLR05, BGM03, CJ12, CT16, GLZ19, HDX^{+}17, HAA^{+}07, KS04, KMJ^{+}17, LCFkL05, MTGZ17, MRH14, Nob08, OOTK01, PTCN07, RSSM06, SKR17, TMAG03, Wit10, ZYL^{+}08]. languages [BDV02, CGK14, Hoh06, LWB13, MSP^{+}19, SPBL06]. LAR [XWPM19]. Large [AML^{+}15, HTR10, KBT^{+}14, LW05, PDD14, RIP18, SL18, SVN12, AAE^{+}09, BH09, BGGL07, BCM^{+}07, BZdR^{+}10, BMPS07, CEH^{+}06, CHM15, CBQ^{+}11, CGN15, CWM18, CPS^{+}14, CDH^{+}15, DVD^{+}12, DLX^{+}16, DOJ^{+}19, DLM13, DP19, DZM^{+}15, EBGS01, ERZ^{+}11, EJD17, FHO^{+}15, FAPC16, FBV^{+}13, HFDJ10, HWQ^{+}16, HLF^{+}17, HWL18, JAA08, JCK^{+}13, JPWH02, KRS11, KCZ^{+}05, LME^{+}19, LBV16, LXRJ13, LXW^{+}16, LFZ^{+}17, MZJ^{+}19, MvWvM^{+}17, MSCS18, MFG^{+}13, MCY^{+}10, MKAKG14, MB14, MJD15, MJD17, Not16a, PTL^{+}16, PAM^{+}15, QZYZ16, QLS13, RVRD10, RKS02, RRR15, SNH15, SK09, SLV12, SMM18, SCBH09, SGCG09, SSIH19, SLM05, TJ17b, TJ17a, TTL06, TBK06, TRH^{+}02, TB12, VL17, WYZ^{+}17, WZXZ12, WSWL12, XBXS13, YWBS19, YLEB14, YMLR16, Zha19, ZYZ06, ZQZ^{+}16, ZHGX16, dCRS11]. large-data [CEH^{+}06]. Large-Scale [PDD14, HTR10, KBT^{+}14, LW05, AML^{+}15, BH09, BCM^{+}07, CHM15, CBQ^{+}11, CGN15, CWM18, CPS^{+}14, CDH^{+}15, DLX^{+}16, DP19, DZM^{+}15, ERZ^{+}11, FAPC16, HFDJ10, HWQ^{+}16, HLF^{+}17, HWL18, JAA08, JCK^{+}13, JPWH02, KCZ^{+}05, LBV16, LXRJ13, LXW^{+}16, MZJ^{+}19, MvWvM^{+}17, MSCS18, MCY^{+}10, MB14, MJD15, MJD17, Not16a, PAM^{+}15, QZYZ16, QLS13, SNH15, SK09, SMM18, SCBH09, SGCG09, TJ17b, TJ17a, TRH^{+}02, WYZ^{+}17, WZXZ12, WSWL12, XBXS13, YLEB14, YMLR16, Zha19, ZYZ06, ZHGX16, dCRS11]. large-sized [FHO^{+}15]. laser [DYY^{+}19]. latency [DGW16, ETR^{+}15, LWF^{+}15, MVWJ14, PRD^{+}13]. latency-sensitive [DGW16]. latent [SIS19, ZSWS18, DP19]. Latent-lSVM [DP19]. Latest [MPSGD14, SRM13a]. Lattice [LMKT13, CGK^{+}16, CNAQ18, MWLS11, QWZ^{+}19, RMW19, ZAB^{+}19, BFM^{+}10, VLJ17]. Lattice-based [LMKT13]. lattice-Boltzmann [MWLS11, RMW19, VLJ17]. LAUP [BNNH19]. law [XL17]. Layer [HJTX17, EB10, HKA^{+}15, KKJH03, KR11, SKS^{+}08, Tru15, WRLS12, YLL^{+}18]. layer-based [KR11]. layered [DWC09, LDZ^{+}15, LGQS12, LLH^{+}18, OKP16]. layers [LZL17a]. layout [HP11, IT03, LHH^{+}17, TBK06, WFKS18]. LBM [VL17]. LC [LLYL09]. LC-GRFA [LLYL09]. LCG [NCD^{+}08]. LDA [GLD17]. LDPC [LLH^{+}18, SCLK15]. LEACH [MSMA19]. LEAD [CM07a]. leader [ZCXH17]. leader-follower [ZCXH17]. leadership [JOK^{+}18, LLT^{+}14]. leading [DWC09]. Lean [BPS19]. leap [JTD^{+}19]. Learning [GBB^{+}15, Mar19, ORDG15, ZWW^{+}18, ZL06, ABR^{+}06, BHD13, CCCW13, CLL^{+}19, CZQ17, DFC12, DÇK^{+}18, DH13, ERZ^{+}11, FHO^{+}15, FE17, FSM^{+}19, GLZ19, HAN19, HZL19, HCBRM16, Hey19, JyLdZ^{+}18, KKDS19, KTHA18, KSB^{+}19, LLM19, LLH^{+}09, Li18, LLJ18a, LLJ18b, LS19a, LDZ^{+}19, MAVG16, MYS19, MSB17, MCXP15, OHJ13, PZO19, PRC^{+}14, Roj19, SNEP14, SGJ^{+}17, SİM^{+}07, TWZ^{+}19, THQ19, VLW19, VRDTB^{+}16, WF18, WML^{+}19, WWL^{+}17b, XJAJ18, YAG19, YZR14, YG19, ZBZ^{+}18, ZLH^{+}18, SW11]. Learning-based [ZWW^{+}18, FE17, TWZ^{+}19, WML^{+}19, ZLH^{+}18]. learnt [BLCC19]. least [ABV05, DMC^{+}18, MLL^{+}11, ZF18]. least-squares [ABV05, DMC^{+}18]. Lecture [Bou13]. legacy [BR04, MMS07]. legends [BH05]. Legion [NNTH^{+}02, NCWD^{+}04]. length [CL14, MNL15, XXLL17]. less [DMA13, FNI17]. Lesser [ON01, ON02]. lessons [LLT^{+}14, OGA^{+}06]. Level [MP04, AAP13, BPL12, BDV02, BBA18, CAD^{+}18, CK13, CCSS10, CCC12a, CCW^{+}15, DA19, DNB19, DPGA11, GCO^{+}14, GL19, GPW05, HJB12, KM03, KKJH03, KAP13, KJŚ^{+}15, KMRT18, LGLA15, LLKC08, LPY^{+}08, LHH^{+}17, LWC17, MHH16, MGI17, MPT07, MJD15, MCC16, NTK08, OGA^{+}01, OTT19, Pac16, Sør13, TTD^{+}11, VS02, VFAD18, WBZ10, WCZX16, WFKS18, XLY^{+}16, YS07, ZLKK17, ZZZ^{+}15, dCHMJ12, dRC10]. levels [CSB^{+}16, GKPT13, JMF09, KG19, SLB08]. levelset [FYKW15]. Leveraging [BBW19, KOOB15, PGF19, TWW^{+}19, GKG^{+}04, LGD15, Mit17c]. LFTM [MMBP12]. LFU [BBC16]. LHCb [SRTG^{+}07]. libraries [ASS08, BHL^{+}09, CL01, MD02, TTD^{+}11]. library [AMHC11, CAD^{+}18, CSWB11, GDMT^{+}12, HKRR08, JKM^{+}17, KS05, ON01, ON02, SSU18, YB12, VFAD17, vWAH^{+}02]. Life [LGD15, Qiu11, QFG14, QFT14, ACC^{+}15, Bou13, GvHKK11, OGA^{+}06, RTPPH12, SR19c, YWLQ18, OKG18]. Lifemapper [WSP17]. lifetime [CLH13, DMA13, KCBO17]. lifetime-aware [CLH13]. lifetime-driven [DMA13]. ligand [EDB^{+}14, TCP^{+}05]. Light [JSR^{+}19a, BJC17, LLT19b, ON01]. light-weight [ON01]. lightpath [MvWvM^{+}17]. lightpath-connected [MvWvM^{+}17]. Lightweight [BNNH19, FLB^{+}05, NR08, SWP17, Bac03, BC16, BBB16, CJZ^{+}15, CCL^{+}17, FLL^{+}14, JZJW15, KN01, ON02, QWW^{+}16, RBB^{+}09, WZXZ12, XXCY19, vRS05]. like [CCG^{+}08, KOB01, TWN07, XYL18]. likelihood [SLM04]. likelihood-based [SLM04]. limitation [RSPV17]. limited [KTZ^{+}18]. limits [BGGS14]. Linda [BDM18, Men03, WCC04]. line [CRC15a, DMR^{+}07, ESGQ^{+}11, zGWXT09, HK01, NA15, VB16, WFKS18, WKL14, YYZH19, YESG^{+}19]. line/off [zGWXT09]. linear [AAC^{+}15, ABC19, ADI^{+}14, ADF^{+}19, BHL^{+}09, BLL18, CC13, CL18, CL14, CNP^{+}15, CGGH17, DK09, DLH01, HLYD12, HAA^{+}17, JSS07, KD07, KLDB10, MSK19, Nak02, OHJ13, PZH^{+}15, SD15, SLB08, YSWZ17, ZWT^{+}18]. linear-time [DLH01, PZH^{+}15]. linearizability [Low17]. linearly [PHY^{+}18]. lines [FPHZ19, Pan20]. Linguistic [OO18, ECP18, MCG^{+}08, MMBP12, OTO18]. Link [LLX15b, IHB15, LXL^{+}09, PZH^{+}15, WRLS12, ZLF19, ZM13, Zhu07, ZYL^{+}08, ZZ11, ZX11]. link-based [ZM13]. linked [LBOE18, FVRM15]. links [LFZ07]. LINPACK [BCD^{+}10, DLP03]. Linux [EEK^{+}04, BdL06, Kar16, KF01, MKO^{+}17b, PKB03, WJKS18]. Linyphi [DEF08]. list [DFG17, RCXS09, WLLL15, ZQW^{+}17]. Lists [PPdSTB17]. literature [FVRM15, SN18, SKA^{+}14]. Liu [FPHZ20, HML19]. live [EJD15, MY17, MYY18, RMP13b]. liveness [IR11]. Living [dMd^{+}17]. LMI [EME19]. Load [FED03, MS17b, WQS^{+}16, AS15, AR16, APHB16, BGV^{+}01, CLW^{+}19, CW11b, DBR13, DL07, EB18, EOD^{+}19, FJ05, FE18, FT06, FGC06, GRQ19a, GCL08, KKTHL13, KL02, KYM17, KR04, LM08, LJL^{+}17, ML19a, MKIO04, PBK19, QCB17, SVB19, SBDP15, VP19, WJYH16, WLL03b, XBZ10, XTB17, YSL^{+}15, YZ10, ZYL10, ZEB10, ZMYA18]. Load-balanced [WQS^{+}16]. load-balancing [EB18, FT06]. load/unload [YZ10, ZYL10]. loading [LOKW^{+}10]. loads [LYC16]. local [AMHC11, BY12, DAC12, KHF^{+}17, LSXL17, LW05, LLYL09, PLL14, PWH18, SL18, TJ17b, WW08, ZBZ^{+}18, ZLG^{+}19, ZHW^{+}16]. locality [BMPS07, EK19, FJ05, KRS11, Mit17c, PWMX17, PLR^{+}14, WQS^{+}16, WGY^{+}19, XWPM19]. locality-aware [EK19, WQS^{+}16, XWPM19]. locality-based [WGY^{+}19]. localization [BAT13, CGI17, KBH15a, ZY12, ZCL^{+}19]. Locally [YWR^{+}19]. LoCATe [BCI^{+}18]. LoCATe-based [BCI^{+}18]. location [BAZ18, GMS09, GBJ19, JBL15, LWYM16, LLT19b, LWY15, MJS19, NLFA19, PWC^{+}14, PC17a, SWS^{+}18, TZLC15, TC17, WHXB19, XZZ16b, YKD^{+}15, ZSZ15, ZDL19]. location-allocation [WHXB19]. location-based [BAZ18, LWYM16, PWC^{+}14, PC17a, TZLC15]. lock [ASS19, AR19, DDF^{+}17, LWB13, ON02]. lock-based [ASS19, LWB13]. lock-free [AR19, DDF^{+}17]. Lockahead [MFC18]. locking [MFC18, YKA^{+}19]. locks [KNT^{+}01, SPS17]. loess [ZCW^{+}18]. log [FBYO12]. logarithmic [LCM12]. logging [BMA03, BBD10, BHBD13, RM11, SB19b, YLLZ09]. logic [BH09, DLH01, FMS11, HWZX08, LWW06, SS19b, WBD^{+}19, vO01]. logic-based [WBD^{+}19]. logical [HYGF19, MTGZ17]. logistic [HAA^{+}17]. logistics [cDrLyC^{+}19, GRQ19a, LWC12, Qia19, Yan19a]. logon [BFG14]. logs [WDQ^{+}18, WCLC13]. LoM2HiS [EBMD13]. Long [ZHX^{+}19, ABK^{+}18, dFdOSR^{+}19, GM17, QZYZ16, SCRV11, SVN12, ZGS17, ZXX17]. long-distance [ZGS17]. long-endurance [ABK^{+}18]. long-running [SVN12]. long-short [ZXX17]. long-tailed [GM17]. look [FBS16]. look-ahead [FBS16]. lookaside [CDN15]. lookup [KW19a, MA15, WHX19, WTN07]. loop [CSC^{+}17, CPXA06, DAB09b, GOLL17, HBKM06, OGA^{+}01, PCF^{+}17, PGC^{+}19, WDQ^{+}18, YWC11]. loop-level [OGA^{+}01]. loop-scheduling [CPXA06]. loops [JLT06, TXY^{+}16, ZLKK17]. Loosely [Sod05]. loss [BRCV16, CSAC19, FNI17, XBW^{+}15]. loss-less [FNI17]. loss-tolerant [XBW^{+}15]. losses [LLN^{+}14]. lossless [TZG^{+}19]. Low [ALYD17, BCF12, ETR^{+}15, LGLA15, Akt18b, BDE^{+}19, FAM^{+}18, GHMX13, LDZ^{+}15, MBMB18, MA15, MKO^{+}17b, NK19, PLM^{+}19, PA18, PS19a, QWZ^{+}19, SPQ^{+}17, SM19c, VS02]. Low-cost [BCF12, QWZ^{+}19]. low-density [LDZ^{+}15]. low-energy [MBMB18]. Low-latency [ETR^{+}15]. Low-level [LGLA15, VS02]. low-noise [SM19c]. low-power [MKO^{+}17b, NK19, PLM^{+}19, PS19a, SPQ^{+}17]. low-rank [BDE^{+}19, PA18]. lower [AVS^{+}19, FOTW04]. lower-dimensional [AVS^{+}19]. LR [ALVY05]. LR-PCR [ALVY05]. LRFU [BBC16]. LRU [BBC16]. LSB [DZZL19]. lSVM [DP19]. LTE [BN19]. LTE-A [BN19]. LTTng [MD19]. LU [WZ04]. lubrication [GB07]. Lucene [DKMM14]. lunar [LLT19b]. lurkers [AMP^{+}18]. LURR [YZZ^{+}10]. Lustre [DL10, MFC18]. Lyapunov [BEQOR17]. LZ [AL04]. LZ-77 [AL04].

M [ZMJZ10, ZMJZ10]. MAC [EB10, GHMX13]. Machine [FE17, HCBRM16, MAVG16, Roj19, TWZ^{+}19, AM15, AFT01, DÇK^{+}18, DXM^{+}17, ERZ^{+}11, GPW03, HTHW16, IRB19, KTB17, KCKC15, KMG^{+}18, LF15, LS19a, LLY^{+}19b, MS17c, MW18, PZO19, PCB^{+}18, SNEP14, SGJ^{+}17, STO17, SİM^{+}07, SKJ17, SYMA17, TVCB18, VRDTB^{+}16, WKT08, WJ12, WZJD13, WML^{+}19, WJLD09, XJAJ18, YBZ^{+}15, YYC^{+}19, ZXW16a, ZLZ15, ZLH^{+}15, ZBP07, ZWH^{+}17, ZYLY18, BP03, GKP^{+}19, LTK17, SGV12]. machine-based [WKT08]. machine-to-machine [AM15, STO17, SKJ17]. machines [AMAB17, BB12, BB15, BWHS18, BBW19, CCL^{+}17, EMS15, God12, GKG^{+}04, GdMK^{+}18, GE06, HKS19, KBB11, SJB14, VRDTB^{+}16, WDT18, XHCL15, XTB17, ZS01]. macromolecular [ABC^{+}15]. MacZ [CGKW13]. made [BDH16, LGCJ^{+}13, MTA^{+}07, STWSP12]. magnetic [EMEY14, KSM15, WKB^{+}19]. magneto [SCS17a]. magnetohydrodynamics [SPH13]. magnetosphere [Ogi02]. magnetostatic [KMJ^{+}17]. Magnum [SVB19]. Mahalanobis [HYGF19]. maintained [MZ06]. maintaining [ABDO09, BDF15]. maintenance [MST15, YDL09, ZLZ^{+}19, Kin04]. Make [DGW16, dRC10]. Make-way [DGW16]. makes [DKKL06]. Making [MTHK14, ABMMR19, BJC17, DCCZ18, GQJL18, GPS^{+}07, HLL^{+}15, SS17a]. Malaria [GKM^{+}08]. malicious [DLJ15, HCS18, LGQS12, LWY^{+}17, YWL^{+}17b]. Malleable [EDSV09]. malware [HZH^{+}19, HAA^{+}17, KGP^{+}19, LGQS12, NJZZ19]. malwares [HPK^{+}18]. mammograms [SRR19]. manage [VAC^{+}07, XCL09]. manageable [PKB03]. managed [BAC^{+}15, CEG^{+}05]. Management [Boe12, CPB07, CL08, CC09, CW11a, CR13, FAB^{+}07, LV12, PB07b, PK08, YK10, AM15, AG17a, APHB16, AKW04, BDL^{+}15, BBB16, BAGS02, BM02, CSL^{+}18, CRC15a, CR12, CLL14, CJX^{+}19, CGN15, Den07, DMW^{+}10, DS07, DHM14, DMM^{+}07, DXZ^{+}16, EGGA^{+}04, ECP18, GCSB19, GD08, GBG^{+}14, GTA10, GKM19, HMFK15, HWQ^{+}16, HTI05, HAE09, HM16, HFTQ13, JCP15, JBL16, JM19, JJGL13, JSG17, KSS^{+}17, KCKC15, KMG^{+}18, KO06, KF18, KTB04, KTZ^{+}18, KN19, LFPP17, LME^{+}19, LHHJ18, LZC09, LNCY11, LAB^{+}06, LHXY08, LDS^{+}08, MABP13, MBP16, MMW^{+}12, Men03, MB18, MVML11, MS19, NDT^{+}16, OCS01, PYKL16, PVCS18, Pat08, PAM^{+}15, PS13, PXY^{+}07, RRBB11, RSKK19, RAFD14, SMH^{+}19, SIST18, SGSC08, SACJ04, SPJ14, SWW^{+}16, TCDMR^{+}17, TC12, VCP16, VvSI07, WYBS16, WWX^{+}19, WNT02, YWY^{+}10, YLR^{+}13, YESG^{+}17b]. management [YESG^{+}17a, YYS15, YLJZ13, Yu18, ZABP18, ZYFZ19, ZLC17a, ZXW16b, BD04, MP04]. manager [MRS^{+}10]. Managing [BZdR^{+}10, GGFPGB14, HK02, JKL19, Kes04, Zic12, DFLNP07, HCG07, Mit17b, SMY^{+}15, VH12]. MANET [AG17a, BKA19, KBDA19, MBB19, SM19b]. MANETs [ASE^{+}17]. Manhattan [ZWXS19]. Manhattan-distance-constrained [ZWXS19]. manner [ZCL^{+}18]. Manual [AdCPdSD17]. Manufacturing [Bai17, FD01, GRQ19a, LZW^{+}17a, LS19b, LLY19c]. Many [COdO^{+}11, ZYH09, ZQH12, BHBD13, BRCV16, CPEA18, CLF^{+}19, CS17, CZL12, CLRB15, DLZ16, DLK^{+}18, ELM^{+}16, GPPR17, HCD^{+}18, HT15, HvNJB15, HFR^{+}17, JSR19b, JPS17, KQR^{+}17, LGLA15, LL16a, MCP^{+}12, MM17, RLMG16, SPQ^{+}17, VDL^{+}15, YLY04, ZXC^{+}19, ZYH12]. Many-Core [ZQH12, ZYH09, BHBD13, BRCV16, CPEA18, CLF^{+}19, CS17, CZL12, CLRB15, DLZ16, DLK^{+}18, ELM^{+}16, HvNJB15, HFR^{+}17, JSR19b, JPS17, LGLA15, LL16a, MCP^{+}12, MM17, RLMG16, SPQ^{+}17, ZYH12]. many-objective [ZXC^{+}19]. many-task [KQR^{+}17]. many-to-many [YLY04]. Manycore [LOSJ17, CKL17, CGGH17, SSMB15, RGB^{+}15]. Manycores [BL17, BH16]. map [AJY^{+}15, Ano06, BFR05, DIK14, LZY^{+}16, YYLC19, MWL18, SS07]. map-reduce [DIK14, MWL18]. mapping [Ano06, BFR05, CT11b, CDN15, DST11, GFL04, KSB^{+}19, KTM^{+}09, LLT^{+}19a, MPS11, PZ11, QSZL18, RMCN^{+}07, SW09, SSU18, VSR^{+}09, ZDR^{+}18]. mappings [ATİ17]. MapReduce [AFG16, BCCM16, DLX^{+}16, DGL^{+}12, DAL15, Fed13, IHA^{+}15, KF15, KAA19, LWFL14, LH17, LWZ^{+}17, LMX^{+}18, LPG^{+}14, MMW16, MLYL17, RM19, RTMZ13, SGCA^{+}16, SB19a, THF15, TLX^{+}17, UMD^{+}13, VJK13, YXLZ16, YWT^{+}12, ZF18, ZCL14, ZLT^{+}16]. MapReduce-based [DLX^{+}16, LMX^{+}18, RM19]. MapReduce-supported [DGL^{+}12]. maps [Del08, LWW^{+}19, LHXY08, Riz04, XDE^{+}04]. maps-based [LWW^{+}19]. marching [FRKS12]. margin [TJ17b]. marginal [WFS^{+}19]. Marine [MMG^{+}18, LLRS03]. maritime [WTEG17]. Market [VDB09, CAC15, GS04a, JHCH19, KD15]. Market-based [VDB09, GS04a]. marketing [DFG^{+}18b]. markets [GCO^{+}14, GVK12, MRS^{+}10]. MARKOV [MBB19, EMEY14, HPD^{+}15, WYY^{+}19, ZACG16]. Markov-based [WYY^{+}19]. Markovian [DPS07, XWD^{+}12, ZHM^{+}17]. MASA [dFdOSR^{+}19]. MASA-OpenCL [dFdOSR^{+}19]. mashup [WZT11]. masking [PJW^{+}14]. mass [BRWB06, HKG08, WJP14, YG19]. massive [CZL^{+}17, EFM17, FLYL16, MMW16, MWL^{+}15, MCXP15, PWC^{+}14, PW12, SZL09, ZWL^{+}13]. Massively [BÇG14, BS10, BCC^{+}05, CRC15a, CZL12, FBV^{+}13, JKV^{+}15, RGL^{+}15, SAB15, SRM13b, WT10]. master [ACIC^{+}13, CAG^{+}13, PRV11]. master-worker [ACIC^{+}13]. matching [EN16, HTWW19, MWPL15, PQP13, RTMZ13, TJD^{+}17, THQ19, YSWZ17, ZZY^{+}15]. matchmakers [DHC13]. matchmaking [WHXzL15]. MATE [MCSML07]. Matera [GIL17]. material [HH19, JVMN19, LOKW^{+}10, LLY^{+}19a, LHLH16, NAP^{+}07, Sod07]. materials [PSV19, XBB13]. Mathematical [HJH19]. Mathematics [WT15]. MATLAB [PIAH12]. matrices [AKG13, CHP17, MUKY18, SGR19, WZ04]. Matrix [ALKD16, BEQOR13, AB01, Akt18b, ADMQO14, AHK^{+}15, BCI^{+}09, BEQOR17, CKL17, CXC^{+}18, CWMW15, DS04, ER12, FJZ^{+}14, GWW17, GLM^{+}16, GDMT^{+}12, GS04b, GW15, GR14, GW19, HGW18, HT15, JHCH19, JLH^{+}16, KHZN06, KS19c, LL19c, MRL16, MCP^{+}12, NA15, OAS^{+}15, PIAH12, PLR^{+}14, SAD13, TDM^{+}15, VS02, VFG11, WZL^{+}17a, YDS^{+}14, ZF18]. matrix-matrix [AB01]. matrix-sign-function [KS19c]. Matrix-Vector [ALKD16, GW15]. matter [YYWQ19]. matter-element [YYWQ19]. matters [MZG19]. Max [RNJM17, BTG06]. Max-flow [RNJM17]. maxflow [BÇG14]. maximization [JZL15, KTB17, LCYJ08, ZS17]. Maximizing [CBB^{+}19, KCBO17, MRS^{+}10, MS18, PV15]. Maximum [YSWZ17, BRCV16, SLM04, TJ17b, WFS^{+}19]. Maxwell [LGP19]. May [Run10, JW10]. MBSA [CCL^{+}17]. MCDM [SAMS19]. MCSA2018 [Wu18]. MDPC [GAB19]. ME [XHH12]. MEAD [NDP^{+}05]. mean [HW16, SC07a, CKOG10]. Means [SS19c, DLX^{+}16, GMPT15, TLX^{+}17]. measure [AMBT17a, DNB19, NFG19, TTL06]. measurement [BCC^{+}05, BSZ09, BDP^{+}14, CJZZ10, FPHZ20, GBXL17, HFDJ10, HML19, JJGL13, KNT^{+}01, LWS19, MWW10, TPV17, WWL^{+}17a, XHCL15, Xu19]. measurement-based [BCC^{+}05, JJGL13]. measurements [JZMD19]. measures [TALT16]. Measuring [GSR^{+}19, dFMSPSW06, Tan12, XLYX11a, XLMH14, HCC^{+}15]. mechanical [Wan18a, YZ19]. mechanism [AS17, BKM^{+}07b, CLL^{+}18, CSL^{+}18, CJX^{+}19, CLH13, DDX^{+}06, DZL^{+}17a, FPHZ19, FT06, GHLS19, HKA^{+}15, JSR^{+}19a, KKDS19, KGGT12, KYM17, LWY^{+}16, LLF08, LLSL15, MMBP12, MML^{+}17, ON02, Pan20, RIWS17, SGCG09, SIRP17, SYMA17, TXZ^{+}17, WTEG17, WLP^{+}17, ZLH^{+}15, dAAVS12, YYCH10]. mechanisms [ASP19, CW09, CCT15, CLW^{+}15, GP07, GÖ18, LMGZ19, MME13, OSK^{+}01, OKW15, PGK11, RHZ^{+}17, SN18, ZWZ^{+}18, ZYZC17]. media [DA19, DFG^{+}18b, FLG19, GEBA17, HLW^{+}19, KT19c, MLZ19, MYS19, MZA19, PWC^{+}14, PLJ18, PDCA17, XZH^{+}16, YG19]. median [ZDL19]. mediation [SGD15, Kin04]. mediator [OOTK01, RJ01]. medical [AG17b, DXWC16, GSR^{+}19, KM19, KSG11, KT19b, LCC^{+}18, WNN^{+}15]. Medicine [Che19]. medium [ABK^{+}18, YBO10]. medium-altitude [ABK^{+}18]. Meeting [TKK^{+}11, WAS07, WC08, Xu08]. megabyte [HSHT14]. megabyte-scale [HSHT14]. Melling [WWY19]. meltdown [LBJ^{+}19]. members [LZWD^{+}15]. membership [YWBS19]. Membrane [QLF^{+}06]. Memoization [RSN^{+}19, MB16]. MEMoMR [YXLZ16]. Memory [SBDP15, AAW^{+}02, AMABS18, ACG17, BP17, BB02, BDV02, CACC11, CJX^{+}19, CBPP02, CLH^{+}11, CLT^{+}16, DFC12, DVL13, DC19b, DS15, DLT^{+}16, EME19, GTFA13, GYB^{+}11, HTI05, JLT06, KBDA19, KO06, KC06, LCA^{+}19, LLdA08, LPC^{+}14, LSL^{+}17, MY17, MVWJ14, MLC04, MLP04, NPL19, PCVZ^{+}04, RCM12, RLRG15, SGJ^{+}17, SGD^{+}18, SSV19, SS07, SS15c, VL17, WS09, WFKS18, WMvP^{+}09, XJAJ18, YGL05, YWY^{+}10, YYS15, YHH13, ZJS^{+}17, ZXX17, ZLQ^{+}18, ZDR^{+}18, ZHX^{+}19]. memory-aided [KBDA19]. memory-based [AMABS18]. memory-supported [RCM12]. mer [GR13]. Mesh [BOB13, OKM10, CC13, DEF08, Fer13, Fer15, LB11, RLMG16, SGD^{+}18, VLJ17, WO02, XJZ13, XWX^{+}19, YHHS16, ZWXS19]. meshes [FYKW15]. mesoscale [BDY02]. message [AD02, BCM^{+}07, BMA03, BBD10, BHBD13, CMMS17, EN16, Gog11, HdV13, MP05, NMKB03, OKW15, PFU^{+}05, RMG^{+}10, RM11, SVS^{+}08, SSZ14, WKL14, WDW^{+}15, AMHC11, SWL^{+}01]. message-oriented [MP05]. message-passing [BCM^{+}07, RMG^{+}10, RM11, SVS^{+}08, SSZ14]. messages [LCM^{+}17, ZQZ^{+}16]. meta [BKCP09, HPHB^{+}15, XZZ^{+}16a]. meta-analysis [XZZ^{+}16a]. meta-predictor [BKCP09]. meta-workflows [HPHB^{+}15]. Metadata [AFPO08, DVL13, DXZ^{+}16, GD08, GBG^{+}14, KKL09, SK08, dCHMJ12]. metagenomics [WWG^{+}11]. metaheuristics [GIVRC^{+}10, MM17, PSICU18]. metaphoric [PdCMdS^{+}12]. metascheduler [CRCC09, CHL15]. Meteor [JQSP08]. metering [ŠŽH17]. Method [BVGVEAFG11, NMMS01, Akt18b, AS15, BJ01, BDZ19, BV16, BEQOR17, BGM03, CACC11, CJX^{+}19, CW11b, CNAQ18, CL19, CNP^{+}15, cDrLyC^{+}19, DXM^{+}17, DLW19, DYY^{+}19, EN19, FCY17, FNI17, FOTW04, GPW05, HLF^{+}17, HLZD18, HWZX08, HYGF19, JHCH19, JVMN19, KS19a, KO06, KJŚ^{+}15, KZY15, KZY^{+}18, KC13, KW19b, LWZ^{+}17, Li19, LDXC13, LHBW15, LFH08a, LSW07, LCJ14, MKB01, MO02a, MRH14, NJZZ19, PPR19, PCD15, PBK19, QLF^{+}06, QLD^{+}11, RMW19, Roj19, SSIH19, SGR19, SWZ^{+}18, SS19c, SLC^{+}18, SLC^{+}19, TCP^{+}05, WBD^{+}19, WZJD13, WTY^{+}19, WFS^{+}19, YYC10, YLD13, YGG14, YSWZ17, YYXL19, YY19, YYWQ19, YZ10, ZAB^{+}19, ZZYW10, ZDC15, ZBZ^{+}18, ZHC^{+}18, ZXRZ19, ZGL19, ZLL19, ZWW14, AS15]. method-level [GPW05]. methodological [GVC10, MCCG11]. methodologies [PPST09]. Methodology [LG08, ZRB19, ANK^{+}17, FTRA15, HvNJB15, KOO12, MHK^{+}18, MDX14, MSCS18, MJ19, PCF^{+}17, RLC16, SC07b, TWB13, Zen19b]. Methods [BBSW17, GGS^{+}16, MÖO17, Qiu11, QFG14, QFT14, AM01, BFK^{+}17, DFTHD18, DGJ11, DS17, Dra15, GMVRGS15, GSV03, GRS^{+}17, GCPS^{+}14, HR18, JSS07, KJ19b, KRS11, KW18, LW05, LY14, MGBC16, MMB^{+}17, MKSS16, MB02, QH10, SE01, YDB^{+}13]. metric [CT16, NvV09, WLW14]. metrics [FJG^{+}13, GGS^{+}16, OORVB14, vAVS12]. metro [NFG19]. MFCC [LLLS18]. MFCC-based [LLLS18]. MFIX [GPS^{+}07]. MG [WWG^{+}11].

MG-RAST [WWG^{+}11]. MGRIT [MCCD18]. MHD [Ogi02]. MIC [CSMS^{+}19, HLCW15, MJD17]. Micro [MISV13, ADSV16, BCA^{+}10, KN19, LWY^{+}17, LWP19, WWY19, ZABP18]. micro-blog [LWY^{+}17, LWP19]. micro-finite [BCA^{+}10]. micro-grids [ADSV16]. micro-perforated [WWY19]. micro-server [ZABP18]. Micro-transactions [MISV13]. microalgae [GED^{+}18]. microarray [MSM^{+}14, ZLW^{+}19]. microarrays [MKKB04]. Microbial [SR19b, WYY^{+}19]. microblog [GC18, LWT^{+}16, LMCL19, ZCLL19]. micromachining [QLZX19]. microprocessors [MST13]. microscopic [XTZ10, ZCW^{+}18]. microsecond [AHP^{+}13]. Microservices [YJZZ19]. Microservices-enabled [YJZZ19]. Microsoft [TH10]. microtext [PLJ18]. Middleware [AJM12, ANTZ09, BCM^{+}05, KR06, MvNK^{+}06, MFF04, Nar05, PC17b, SN06, SCNH07, SM11, SBP12, AvdADtH09, AdSCdR^{+}19, AHM06, Ang08, CEH^{+}06, CC10, CMB06, CM02, CBP^{+}04, CGB^{+}06, CRGR^{+}12, DDP^{+}06, DvNM^{+}11b, ERZ^{+}11, FRdOR^{+}19, FGP^{+}11, GKG^{+}04, HGB^{+}08, JQSP08, JZJW15, KKV13, MP05, MB12, NJ05, PGO^{+}04, QLC04, RE03, RS11, RDP10, VSKK09, XPBS11, ZWF^{+}06, dCGKG06, dMd^{+}17, vHMB08, SANB08]. Migrate [YBZ^{+}15]. migrating [KBG^{+}09]. Migration [AdCPdSD17, ACC^{+}12, DC19b, GMS09, Jon09, KM13, KTB17, MYY18, MSP^{+}13, MRS08, MP04, PCB^{+}18, RMP13b, SYMA17, WDT18, YBZ^{+}15, ZLZ15, ZYLY18]. military [VP19]. Millennium [VRMB13]. mimetic [OFR^{+}17]. MIMO [JKV^{+}15, KT19a]. min [RNJM17, BTG06]. min-cut [RNJM17]. MIN/MAX [BTG06]. mine [LMGZ19, YHYY19]. mini [Fer15, LHBW15]. mini-app [Fer15]. mini-application [LHBW15]. minimal [BVIB19, CSMS^{+}19, HMM^{+}09, PZHS18, YESG^{+}19]. minimization [AHK^{+}15, HLHC12, PC14, XXLL17, ZFW^{+}17]. Minimizing [TY15, DBR13, HSL19, JK13, YDL09]. Minimum [ZLZ15, BXLJ16, DLM13, LLT^{+}19a]. Minimum-cost [ZLZ15]. Mining [FBYO12, KT19c, WCLC13, XLYX11b, ZGST08, ADK^{+}17, BPS19, CV07, CT12, CCP^{+}15, CM18, DBH^{+}17, EPB14, FLG19, GSWJ19, GLD17, HMM^{+}09, HAJL16, LJPP16, LLG^{+}15, LXYC17, LMX^{+}18, LMOT10, Mal05, NJM19, SLC^{+}18, SLC^{+}19, TTV08, THM^{+}11, WWS^{+}12, WW19, YL19, ZX09, ZHJ19, ZSS18, ZKR^{+}07, LMOT10]. minority [AOK19]. mirroring [JVPI18]. misbehaving [MAdS^{+}10]. misinformation [BA19]. missions [ZJS11]. Mississippi [HBH02]. misuses [DDF^{+}17]. Mitigating [QCB17]. mitigation [GKM19, IRB19, ORdSL13]. mix [HLX^{+}16]. mixed [BDE^{+}19, CSTV06, DS04, KD07, Pla08, Roj19]. mixed-parallelism [DS04]. mixed-pixel [Pla08]. mixed-precision [BDE^{+}19]. mixing [Bou06]. mixture [LLM19, PPP10]. ML [KSM^{+}19]. MLC [AMABS18]. MLSL [KSM^{+}19]. MOALO [SM19c]. Mobile [CKC09, Du18b, MWJ^{+}10, OKBO19, VSB^{+}15, WJJM17, XHH12, AKMZ13, AA16, Aia15, AMSS15, AG17b, BYN^{+}17, BCI^{+}18, BAS07, CWXW16, CWYX17, CL16, CJ12, DD16, DCP^{+}17, DA15, EJF^{+}16, FRdOR^{+}19, GBSHA01, HKA^{+}15, JRHJ16, JLHH14, JVPI18, KOO12, KKK10, LYF^{+}17, LHT^{+}09, MTM19, MSMA19, MABP13, MBP16, MDX14, MM19, Not16b, PYKL16, PGK11, PCD15, PRS01, QKSJ07, QMK12, RSPV17, RHS17, Sha15, SR17, SKB^{+}17, SS15b, SSC^{+}16, TC17, VLW19, VT15, WHXzL15, WZS^{+}15, WWL^{+}17a, XY17, YCW08, YWM^{+}10, YCWH07, YNX^{+}16, Zen19a, ZMZD11, ZY12, ZZ18, ZFWJ19, vHMB08, DD16, JM19, MWJ^{+}10]. Mobile-Grid [MWJ^{+}10]. Mobility [Den07, CLR18, MBP16, MJ11]. modal [YC19b]. mode [AAF17, AR19, KB18, LHB^{+}19, PLL17, XS19]. mode-directed [AR19]. Model [CLR18, LGG16, MK12, YC19b, ABtGT^{+}12, ASWR12, AMGCC17, AKM^{+}06, ABG^{+}13, Bac03, BV16, BVGVEA11, BCCM16, BCdlCT06, BLL^{+}19, BXLJ16, BDY02, BBB^{+}14, BAZ09, BBD10, BDG^{+}10, BBSW17, CL01, CAC^{+}08, CTY15, CZWH07, CXPL15, CWXW16, CLF^{+}19, CN16, DD17, DCJ14, DWC09, DLZ16, DNB19, DHC11, EMEY14, EJD15, FCY17, Fec12, FBV^{+}17, GQ04, GD06, GWVP^{+}14, GC18, Guo19, GKM19, GVP^{+}14, HZHP09, HW16, HY12, ID18, JAA08, KS19a, KCD19, KA09, KV12, KCW09, KKK^{+}19, KHL17b, KHZ^{+}15, LVN^{+}12, Lan17, LLWS09, LKPM09, LXP^{+}12, LZT12, LLX^{+}15a, LZL^{+}17b, LLY^{+}19a, LSZ19, LZG^{+}19, LF17, LCW^{+}17, LHB^{+}19, LFH08a, LZC08, LZC09, LWLZ11, LXW17, LL19b, LLRS19, LSP15, MLS^{+}15, MLG15, MTGZ17, MS13, MRMC15, MHH16, MZW^{+}16, MBC^{+}14, ML19b, MÍM19, MGM^{+}08, MSV^{+}10, MCC16, MKSS16, NO02, PP17, PAdS^{+}17, PSW11, RZL^{+}19, RCR^{+}15, RHL^{+}18, SMH^{+}19]. model [SNH15, SS17a, SJW18, SKK01, SK04, SvDO15, SS18, SR17, SSZ14, SE01, SR19c, SS19c, SZR16, TYHL12, TCSBMG17, TFG^{+}12, TZKH12, TLPS18, TW07, TMAG03, TSKM18, Tru15, VCW13, WCZX16, WLZ17, WZL^{+}17a, WDQ^{+}18, WBB^{+}07, WYY^{+}19, XDL^{+}11, XTZ10, XLHT17, XDP18, XHD^{+}19, XXY^{+}16, YGL05, YXL17, Yan19a, YC19a, YLZ18, YHJ^{+}14, YZXW17, YYWQ19, YHH13, YLJZ13, ZQLZ12, ZCL14, ZSL^{+}15, ZH16, ZWLY16, ZXX17, ZZY^{+}15, ZLQ^{+}18, ZQMC19, ZCL^{+}18, Zhu18, ZWL^{+}19, ZCS06, Zhu07, dP06, vHvdSvL03, vdABST10, AFR09, PGW06, PXY^{+}07]. Model-based [LGG16, BAZ09, EMEY14, Lan17, PSW11, YHH13, ZQMC19]. model-driven [KHZ^{+}15, XDL^{+}11]. Modeling [ADMQO14, Bai17, CGIP16, DD16, DLH01, DAL15, DPP^{+}19, FPC15, JB19, MBC^{+}14, RR15, SPZ^{+}10, SZT19, WMA07, XRD^{+}17, XWX^{+}17, Zho06, ZBC^{+}07, ZYL^{+}08, ACC^{+}12, AHP^{+}13, BM02, CCC^{+}16, CLZX10, CSB^{+}16, Cuz11, Dra15, FRU12, GRQ19b, GAE^{+}06, GW15, JVMN19, LLX^{+}15a, LBDS15, PSIP16, RGAK15, SB18, SKR17, SB17, SAM^{+}17, TSL^{+}19, TMR^{+}07, TSS18, XWFH08, XM02, ZDA^{+}07, ZJT^{+}19, ZGL19, ZACG16, ZABP18]. Modelling [MS10, XHW^{+}19, BBPV05, BWHS18, BBGA03, Eng15, GPP^{+}18, IAH^{+}15, LG08, LJML10, PIGK16, RW10, dFMSPSW06, SCV^{+}08, VGL06, vSB06]. Models [BL17, BJ18, Fox10, LLMK18, OM06b, SRdS09, AGMR05, AFG^{+}05, ABDR13, BH16, BDY03, BAGS02, CLH^{+}08, CLRB15, DvdS06, DLM13, GRS^{+}17, HTHW16, HPD^{+}15, HWZX08, JDH^{+}18, JAU19, KTHA18, KD19, KSG11, KKGO04, KSB^{+}19, LLM19, LPA^{+}08, MAVG16, MMSG17, MLP04, MSG10, OKM10, OHJ13, PRNM19, Qia19, SNEP14, SK17, TSL15, TLWZ14, VYK^{+}10, WDT18, WCLC13, YTL19, YOBS16, ZlC15, Zho06, ZBC^{+}07, ZDC^{+}09, ZLA^{+}15, vdS06b]. Modern [Wu18, ABMdAMF19, BCI^{+}09, CGST17, HTHW16, MMSG17, PA18]. modes [JMF09, RR11]. Modified [IC19, SS19a, ZCW^{+}18, ZCH^{+}18]. Modifying [VŠC17]. Modular [MPHL03, CZO^{+}08, DBGA16, YF13]. modulation [LLQL14]. modules [FGC06, ISS^{+}02]. MOEA [ACIC^{+}13]. moldable [Hun15, SO16]. molecular [AHP^{+}13, BDW14, BBGA03, DCD^{+}14, DG11, DLK^{+}18, EB14, GKS09, GBG^{+}14, KF11, LGL16a, PSICU18, RMCN^{+}07, TCP^{+}05, WJT^{+}14]. moment [JW10]. MONC [BBW19]. monitor [BKH08, CCCC06]. Monitoring [CPG^{+}16, Akt18a, BFH17, BAT13, CLW^{+}19, CHX^{+}19, CLH19, FLB^{+}05, GIL17, HFDJ10, HGB^{+}08, JBL16, Kri19, LTL^{+}17, LLL15, LMDP19, LMGZ19, MT19a, MMB^{+}17, NMM^{+}10, QLC04, SR19b, SWD^{+}15, TBK^{+}15, XBXS13, YHYY19, ZSZ^{+}14, ZYZC17, MCSML07]. monitors [CMPT08]. Monte [CCO15a, ATVLM14, GQH17, KDC17, NDT^{+}16, RDP10, SS15c, WZJD13]. MOPSO [WSZ^{+}18]. morphological [KL19]. Morton [TBK06]. MoSGrid [HPHB^{+}15]. most [EZJ^{+}18]. motif [DRZ13, FMS15]. motion [ABG^{+}13, JTD^{+}19, Qi17, TNH15, TNI16]. motor [WCZ^{+}18]. move [Ros06]. movement [BCD^{+}02, XS19]. Mover [AC08]. Moving [LTKF11, ATSAK15, DMC^{+}18, LOSJ17]. MpCCI [JK06]. MPDATA [RIWS17, RWK17]. MPI [ABF^{+}17, BDB^{+}13, BR04, BBW19, CC10, CDMS15, DL10, DBB^{+}16, EDSV09, FA18, FMS15, FLB^{+}05, GRW^{+}19, HRR^{+}11, KC06, LGG16, LL01, LZC^{+}02, LKJ03, LCC^{+}03, LKYS04, LSK04, MTK16, MvWL^{+}10, NO02, NSBR07, PDY14, PTL^{+}16, QB12, WLR05, YWC11]. MPI-2 [LSK04]. MPI-CHECK [LCC^{+}03]. MPI-IO [DL10, LGG16]. MPI/RT [SKD^{+}04]. MPI/RT-1.1 [SKD^{+}04]. MPI2007 [MvWL^{+}10]. mpiBLAST [YHK09]. MPICH [LKJ03]. MPLS [TAMC19]. MPMD [KB18]. MPSO [FTT15]. MPSoCs [LCA^{+}19]. MPTCP [CSL^{+}18]. MPTCP-based [CSL^{+}18]. MQTT [KG19]. MR [SRM13b]. MR-search [SRM13b]. MRMOGA [JC07]. MS [CV07]. MS-Analyzer [CV07]. Ms8.1 [ZGRSC10]. MSBNs [AC09]. MTA [BS04]. MTA-2 [BS04]. MTD [YC19a]. Multi [AR19, BAT13, CCC12a, CWYX17, CNAQ18, CCTW11, DZL^{+}19, DL07, EJD17, KS19b, KH12, LLT^{+}19a, MM17, OKP16, RHuR^{+}19, SG18, TSL15, WJ12, WBD^{+}03, XZ09, ZWY^{+}19, ZYH09, ALKD16, AT01, AFGL09, AYN^{+}14, ART14, AMTM17, ACCM17, BPL12, BIK^{+}11, BKSM^{+}15, BDY03, BRCV16, CPEA18, dCPD13, CGMJ^{+}19, CKOG10, CZG16, CZL^{+}17, CLW^{+}18, CCW^{+}15, CZ15b, CGN15, CN16, DCJ12, DIM18, DS19, DMC^{+}18, DCCZ18, DLZ16, DP19, DWC^{+}15, DYY^{+}19, DXZ^{+}16, DA15, EFG^{+}03, EHSU07, EJF^{+}16, EFA^{+}17, GWW17, GKP^{+}19, GLM^{+}16, GMMT17, GPVCdBRO12, HJB12, HTHW16, HKAC14, HSL19, HFR^{+}17, HM16, HAA^{+}07, HAA^{+}17, IZXM09, JvAB^{+}15, JCVU15, JC07, JQL^{+}15, JL10, JSR19b, Jon09, JK10, JPS17, KSG11, KOOB15, KW19b, KN19, LDPZ14, LXW^{+}16, LZL17a, LPY^{+}08, LQL^{+}09, LSMVML15, LXK^{+}19, MGBC16, MHLC^{+}05, MS07, MFG^{+}13, MH07, MSB17, MML16, MLVBW12, MDL^{+}10]. multi [OLG^{+}15, OAS^{+}15, OM06a, PRS16, PZ11, PRT09, PTCN07, PSICU18, PS19b, Puf13, QCB17, RHBK11, RHL^{+}18, SKK02, SAD13, SLV12, SCGZ19, SAP16, SPW09, SWW^{+}16, SMFM18, STL^{+}15, SVN12, TYL^{+}15, TL19, TKS18, TMAG03, VGN^{+}16, VLF^{+}13, WLWX14, WJYH16, WLWX16, WFKS18, XLHT17, YCL11, YLC11, Yan19a, YKA^{+}19, YYZH19, ZSWS18, ZWL^{+}13, ZLL19, ZM13, ZQZ^{+}16, ZDR^{+}18, ZCL^{+}18, ZZZ^{+}15, ZJL15, ZZL^{+}17a, ZTGW17, dCRS11, vdKEL10, SAP16]. Multi- [ZYH09, CPEA18]. multi-agent [CGN15, CN16, EFA^{+}17, GPVCdBRO12, HM16, KN19, OM06a, RHL^{+}18, TL19]. multi-asset [DCJ12]. multi-attribute [PS19b]. Multi-authority [ZWY^{+}19]. multi-channel [DXZ^{+}16]. multi-class [DP19]. multi-cloud [LSMVML15, QCB17, SWW^{+}16]. multi-cluster [Jon09, YCL11]. multi-component [ALKD16, EJF^{+}16, SVN12]. multi-constraint [SKK02]. multi-coprocessor [DWC^{+}15]. Multi-core [XZ09, AYN^{+}14, ART14, AMTM17, ACCM17, BRCV16, CZG16, CZL^{+}17, GLM^{+}16, HTHW16, HKAC14, HFR^{+}17, IZXM09, JSR19b, JPS17, KSG11, KW19b, LQL^{+}09, MGBC16, MSB17, OAS^{+}15, PZ11, RHBK11, SPW09, STL^{+}15, SEF^{+}14, TYL^{+}15, WJYH16, ZZL^{+}17a]. multi-cores [BKSM^{+}15, ZQZ^{+}16, ZDR^{+}18]. multi-CPU [SAP16]. multi-CPU/multi-GPU [SAP16]. Multi-criteria [KS19b]. multi-datacenter [ZWL^{+}13]. Multi-dimensional [AR19, CWYX17, DMC^{+}18, DCCZ18, JQL^{+}15, ZM13]. multi-disk [DYY^{+}19]. multi-domain [SCGZ19, ZTGW17]. multi-exponentiations [TKS18]. multi-functional [LDPZ14]. Multi-GPU [CNAQ18, KH12, GMMT17, VLF^{+}13, ZCL^{+}18, dCRS11, SAP16]. multi-GPUs [CGMJ^{+}19, PSICU18]. multi-grained [MDL^{+}10]. multi-granularity [YKA^{+}19]. multi-graphics [GWW17, OLG^{+}15, SAD13]. Multi-hop [BAT13, MS07]. multi-information [Yan19a]. multi-infrastructure [JvAB^{+}15]. Multi-installment [DL07]. multi-kernel [SAP16]. Multi-label [DZL^{+}19]. multi-language [HAA^{+}07, PTCN07]. Multi-layered [OKP16]. multi-layers [LZL17a]. Multi-level [CCC12a, BPL12, CCW^{+}15, HJB12, LPY^{+}08, WFKS18, ZZZ^{+}15]. multi-linear [HAA^{+}17]. Multi-objective [RHuR^{+}19, SG18, DS19, GKP^{+}19, JCVU15, JC07, KOOB15, MHLC^{+}05, XLHT17, vdKEL10]. Multi-organization [CCTW11, PRT09]. multi-party [WLWX14, WLWX16]. multi-phase [LXW^{+}16]. multi-powermode [JL10]. multi-processor [AFGL09, MGBC16, Puf13]. multi-programmed [CZG16]. multi-query [CLW^{+}18]. multi-rate [DA15]. multi-relational [ZSWS18]. multi-resolution [BDY03, ZLL19]. Multi-resource [LLT^{+}19a]. multi-rotor [YYZH19]. Multi-scale [EJD17, WJ12, SLV12]. multi-server [CKOG10]. multi-service [MLVBW12]. multi-session [JK10]. multi-site [YLC11]. multi-source [HSL19]. multi-spectral [AT01]. multi-swarm [dCPD13]. Multi-tenancy [TSL15]. multi-tenant [DIM18, SMFM18, VGN^{+}16]. multi-threaded [BIK^{+}11, EFG^{+}03, EHSU07, TMAG03, ZJL15]. multi-tier [LXK^{+}19]. multi-tiered [PRS16]. multi-use [CZ15b, MFG^{+}13]. multi-user [AFGL09, MH07, MML16]. Multi-wavelength [WBD^{+}03]. multiagent [CCCW13, YZR14, ZCXH17]. multibody [XM02]. multicast [CQXW14, EBGS01, LGY17, LWYZ19, MMSN^{+}01, TMZ07, YLY04, ZLC17a]. multicasts [RGX^{+}17]. multicloud [JSG17]. multicomponent [MWLS11]. multicomputer [SAOKM04]. multicomputers [CDA09]. multicontroller [GKM19]. Multicore [DP14, ZQH12, ADMQO14, ATNW11, BHM^{+}12, BHKW12, BLKD08, CGIP16, CLYC16, CEM^{+}17, DRZ13, DJM12, DDM16, DFG17, EPB14, FP09, GGV14, HLYD12, JdM12, JKD19, KLDB10, LXRJ13, LS14, MHJH16, ML19a, MSP^{+}13, Nob08, PDY14, PPBB14, QB12, RVD^{+}12, SCRV11, SLD^{+}12, SSK11, SM09, SHC^{+}16, SJPB17, SW09, TYTY15, TKS18, WJ09, XLL^{+}15, XL17, YWC11, YB12, ZYH12, RGB^{+}15]. multicore-aware [PPBB14]. Multicores [BL17, BH16, CCS14, HT15, May10, PRU14]. multidesignated [AYSZ14]. multidimensional [CAD^{+}18, CWMW15, DÇK^{+}18, GBD16, LGL16a, LLQL14, MMG03, PDY14, SS15c, TMP16, ZZY^{+}15]. multidisciplinary [RWK^{+}02]. multigrid [BFK^{+}17, DVD^{+}12, GKSR14, GRS^{+}17, LW05, SJW18]. multihop [MBP16]. Multilayer [CAKH17, ZP07]. multilayered [CJC^{+}18]. multilayers [SR19c]. Multilevel [MS19, THM^{+}11, CSTV06, EJD15, Fec12, GB07, LGJ17, Nak02, RCLSK16, WXLM19]. multilinear [SZ18, ZXW16b]. Multimedia [BTG06, YWT^{+}12, ABR^{+}06, EB10, FWU^{+}04, HAE09, JL10, JK10, KBDA19, LLWS09, LZW13, MBMB18, PBSB04, RS11, ZBP06, ZCC^{+}06, LV12]. Multimethod [GPP^{+}18]. multimodal [MYS19]. Multiobjective [TKB09]. multiparty [PCT04, RCT03]. Multipath [ZLC17b, MTM19, NIIU17, SS15b, YWM^{+}10]. multipathing [CSL^{+}18]. multiperiodic [PQP13]. multiperson [ZLH^{+}18]. multiphase [GPS^{+}07, MWLS11, XWX^{+}19, YHJ^{+}14]. multiplayer [CRC15a]. Multiple [CY07, FBV^{+}13, WZS^{+}15, ZLC17a, AAE^{+}09, FRKS12, JC07, KL19, KB18, LL10, LZW13, LWYZ19, LLH^{+}17, Lyo02, MAS^{+}14, MLP04, NB12, PWMX16, SJ19, SK09, SCLK15, WLZ17, XLYL17, YDS^{+}14, ZSL^{+}15, ZWL^{+}15, ZNT^{+}16, ZWW^{+}18]. multiple-access [SCLK15]. multiplexer [Mos19]. multiplexing [BVGVEAFG11, GCZ^{+}17]. Multiplication [ALKD16, Akt18b, AHK^{+}15, CXC^{+}18, DS04, FJZ^{+}14, GWW17, GW15, GR14, HGW18, MRL16, NA15, OAS^{+}15, SAD13, TDM^{+}15, VS02]. multiplicity [LH14]. multiplier [SZ18]. multiply [AB01]. multipole [LY14, MRH14]. multiprocessor [CLT^{+}16, KBB17, KL02, LWB13, MT19b, SPS17, The01]. multiprocessors [AD02, CFPJ^{+}17, GA09, KC06, RF15, RS12, SWB12]. multiprogrammed [KL02, YL01]. MultiRace [PS07]. multirail [CFP^{+}03]. multirings [YKD^{+}15]. Multiscale [GBB^{+}15, CNAQ18, KW19b]. Multisensor [JLQ^{+}17, KC18]. multisite [DST11]. multitasking [IOOH12, MÖO17]. multithread [CGIP16]. multithreaded [ABC^{+}15, AAC^{+}15, AR19, BHA15a, BS10, BÇG14, GRS06, GA09, PS07, RS07, TKA^{+}02, WT10]. multithreading [BCM^{+}07, CCC12a, GE08, KIM^{+}03, LZW17b, MIGA18, MKIO04, PHCR09]. multithreading-based [GE08]. Multiuser [LZW13, ZJL13]. multivariate [DLJ15, ID18]. Multiversion [BMS^{+}09, BT04]. multiversioning [TJF14]. multiview [RK15]. multiwatermarking [WL12]. multiway [YKA^{+}19, vSB06]. museum [KJ19a, RBDI17, WST^{+}17]. music [LLD19]. mutual [BDH15, BDH16, BDH18, HBD18, HHPL16, JNUH17, MWL18, XLT^{+}17, XXCY19]. MVTC [BT04]. MyCoG.NET [PTCN07]. myExperiment [DGA^{+}10]. MyExperimentalScience [FMMD13]. MyPYTHIA [HCD^{+}02]. myVocs [GRSB09].

Nadu [PB19b]. Naming [GMS09]. nano [Mos19]. nano-scale [Mos19]. nanosecond [GCO^{+}14]. nanotubes [BBSW17]. NAS [NNON02]. NASA [MDH^{+}16]. Nash [HJTX17]. national [CW07, GBMM15, HSM14, HYL^{+}19, PCsHL18, HGT14]. native [CCP^{+}15, SW12]. natural [GLZ19, GRTX18]. Nature [TTPJ16]. Nature-inspired [TTPJ16]. Navier [DdB01, FBV^{+}13, GSV03, HKB07]. navigation [EJD17, FPHZ19, Pan20, SHT^{+}17, TWZ^{+}19, VZB19]. NCF [PWJ10]. nDT [AMVOSGAC17]. Near [XWH^{+}17, CCW04, JCVU15]. Near-optimal [XWH^{+}17]. near-perfect [JCVU15]. Nearest [CACC11, ATI14, KH12, ZZL^{+}19]. nearly [LZWD^{+}15]. Neat [BB15]. necessary [LFG05]. need [MRS^{+}09]. negative [HLW^{+}19, MKO^{+}17a]. negotiation [ADSV16, CK13, CDL08, KZY15, RCKV12]. negotiation-based [KZY15]. neighbor [KH12, QW17, ZZL^{+}19]. Neighborhood [WXSH19]. neighbors [ATI14, CACC11]. Neighbourhood [Del08, ANH16, ZQK15]. neighbourhood-pair [ANH16]. NERSC [DAC^{+}18, HCD^{+}18]. Nested [RHL^{+}18, TXY^{+}16, ZLKK17]. net [GHB^{+}06, MS19, PP17, SS17a, WDQ^{+}18, XWD^{+}12, SGG07]. net-based [MS19, SS17a, WDQ^{+}18, XWD^{+}12]. NetBuild [MD02]. netCDF [LGL^{+}17]. netCDF-based [LGL^{+}17]. nets [DPS07, MRMC15, TDL^{+}18, EB10]. nets-based [MRMC15]. NetSolve [ACD02]. Network [DCP^{+}17, HF17, Jon09, MRL16, XZ09, ZWH^{+}17, AI17, AAQAR^{+}17, ACGG06, AVS^{+}19, AZF^{+}12, AKW04, BDZ19, BBK11, BDF15, BSZ09, CEH^{+}06, CLL^{+}18, CRCC09, CCK^{+}17, CLZX10, CZ11, CWXW16, CDdW17, CKRO13, CSB^{+}16, CT11b, CDF^{+}17, CM18, CS13, DFLNP07, DGL^{+}12, DK19b, EFA^{+}17, GZG^{+}16, GBMM15, GGC19, GSK19, GKM19, HM12, HDX^{+}17, HL19, HYX05, HFTQ13, IZXM09, JCJ17, JHCH19, JWW17, JK10, Jun16, KHHC13, KHL17b, KJ19b, KKT13, KC18, KWXY18, LLD19, LJPP16, LDPZ14, LL15, LZL17a, LCW^{+}19, LLT19b, LL19a, LDXC13, LLLyL16, LYC16, LWY^{+}17, LAL02, LMDP19, MTM19, MTGZ17, MS07, MAS^{+}14, MRP^{+}18, NSSAK13, NSSAK16, NQL^{+}17, NJM19, Not16b, OORVB14, PFC14, PCsHL18, PZH^{+}15, PAM^{+}15, QD17, QZH16, SJ19, SFCAV16, ŠŽH17, SMS^{+}19, SK17, SZG^{+}19, SPW09, SR19c, SG18, SRR19, Tan15, TYTY15, TC17, TPV17, ULS03, WLZ11]. network [WMA07, WLP^{+}17, Wan18a, WGQ^{+}18, WWX^{+}19, cWsThY19, WL02, XWX^{+}17, XGC19, XTZ10, XHZ12, XADLC15, XBW^{+}15, XZZ^{+}16a, XLL^{+}15, XL17, YCZ^{+}13, YWLQ18, YLLC18, YY19, YLZ18, ZPG10, ZY12, ZSL^{+}15, ZWLY16, ZHC^{+}18, ZL19, ZGL19, ZLW^{+}18, ZWXS19, ZZZX19, ZKJ^{+}07, ZHGX16, ZCXH17, ZYL^{+}08, ZZ11, ZX11, LLX15b, PS19a]. Network-aware [DCP^{+}17, Jon09, MRL16, CEH^{+}06, CRCC09]. network-based [EFA^{+}17, HFTQ13, JWW17, LAL02, ZZZX19]. network-bound [CT11b]. network-enabled [DFLNP07]. Network-on [PS19a]. network-on-chip [XLL^{+}15]. networked [CRGR^{+}12, LLD19, LLL15, WR17]. Networking [SCGZ19, ZDL07, DAC^{+}18, JZJW15, LCM^{+}17, RS13, RLVRGÁ14, SCLL19, WHXzL15, Zhu07]. Networks [AM07, HJTX17, SG19, XLWZ11, AKMZ13, Aia15, AQRA^{+}18, AS15, AAF17, AAHA18, AMP^{+}18, ALL^{+}15, BA19, BJD^{+}19, BPdM06, BNNH19, BN19, BFH17, BBB16, BAT13, CLL^{+}18, CLX07, CGMJ^{+}19, CQXW14, CWYX17, CPD^{+}17, CLH13, CGKW13, CFP^{+}03, CFTT17, CCM^{+}17, CNPP09, CSAC19, CLW^{+}15, CEB^{+}18, CMD17, DLJ15, DLPV07, DGM18, DFC12, Del08, DGW16, Den07, DFH10, DK19a, DEF08, DXHL17, DMA13, DMM^{+}07, DA15, DPP^{+}19, EB10, ESG17, ETR^{+}15, FXX16, FAM^{+}18, FH13, FBV^{+}17, GS08, GHMX13, GGFPGB14, GLL16, GBJ19, HZC^{+}14, HKA^{+}15, HWQ^{+}16, HLF^{+}17, HCS18, HLG17, IHB15, JNUH17, JAA08, JBL15, JSPE15, JWY^{+}17, JWZ13, JKZ03, KOO12, KKK10, KA16, KCBO17, KKW^{+}14, KMA04, KABD07, KDW^{+}17, LCA^{+}19, Li04, LXP^{+}12, LL13, LLC^{+}15a, LWG^{+}15, LGY17, LFZ07, LAM^{+}09, LMO15, LCMY13, LLZ^{+}17b, LXL^{+}09, MSMA19, MVWJ14, MZ06, MDX14, MMBP12]. networks [MOK04, MLRR09, MO15, NLFA19, ORdSL13, OEP^{+}15, OKBO19, PB19a, PF12, PMB15, PY19, PCD15, PS19b, QLLS15, QSX^{+}17, QWW^{+}16, QKSJ07, QMK12, RCB^{+}04, RNJM17, RSPV17, RMP13b, RH07, SAOKM04, SCS17a, SWS^{+}18, SK17, SGCG09, SC07a, SAM^{+}17, TKHA13, TAMC19, TZYL13, TZG^{+}19, TLWZ14, Tru15, UGM18, VS19, VRDTB^{+}16, VvSI07, WTEG17, WBZ10, WYQ^{+}13, WZS^{+}15, WWL^{+}17a, WMC17, WCZ^{+}18, WXSH19, WDW^{+}15, XBK17, XBCW19, XW13, XBW^{+}15, XYL18, XJZ13, XGXH15, XZT^{+}11, XLL^{+}12, XLL^{+}18, YBO10, YGW17, YHHS16, YKD^{+}15, YYZ^{+}17, YESG^{+}17b, YESG^{+}17a, YESG^{+}19, YWM^{+}10, YQL^{+}15, YLJZ13, YMLR16, ZK08, ZLF19, Zen19a, ZSWS18, ZQLZ12, ZJL13, ZGS17, ZLAa^{+}17, ZDL19, ZZY^{+}19, ZACG16, ZFW^{+}17, ZHX^{+}19, ZGX11, ZLA^{+}15, ZCS06, ZKWK17, ZTGW17, dCRS11, dCHMJ12]. networks-on-chip [GGFPGB14]. Neural [EFA^{+}17, ACGG06, CGMJ^{+}19, DFC12, DMM^{+}07, JWW17, LCW^{+}19, LLT19b, LLZ^{+}17b, ML19b, QSX^{+}17, SRR19, TLPS18, Wan18a, WCZ^{+}18, WWX^{+}19, cWsThY19, YWLQ18, YY19, YYZ^{+}17, YLZ18, ZL19, ZGL19, ZLW^{+}18, ZZZX19, ZCXH17]. neuron [KCD19]. neuronal [dCRS11].

Neuroscience [BDMM^{+}05, SBJ^{+}15, SvDO15, SMY^{+}15]. neutrality [YLLC18]. Neutralizer [YDL09]. Neutron [CGK^{+}07, CGK^{+}07]. news [LLZ^{+}17a, PB19a, ZW09]. NEWT [CS15]. Newton [vEGW06]. Next [Ang07, Can06, CS06, Aia15, CDL08, CPS^{+}14, GPS^{+}07, KMJ14, KJ19b, MSL^{+}14, UAW09]. Next-generation [Ang07, Can06, CS06, CPS^{+}14, GPS^{+}07, KMJ14, MSL^{+}14]. ng [RHD^{+}16]. NIC [Gog11]. NIC-assisted [Gog11]. nickel [LMGZ19]. nm [NK19]. NMOS [GZHF19]. no [LSS15]. Noah [BDG08]. NoC [PS19a, GGLD11]. NoCs [ZLC17b]. Node [SBS19, SCLL19, CEB^{+}18, DLJ15, DDX^{+}06, HJB12, MLWA19, SS18, XGXH15, ZQLZ12, ZQW^{+}17]. nodes [AMVOSGAC17, DWC^{+}15, GSK19, LWY^{+}17, MMB^{+}17, PGL^{+}17, VT15, ZWLY16]. noise [GA09, LLY^{+}19a, LHZS19, PWJ10, SS19b, SM19c, XLYX11a]. noises [TL19]. Non [BVIB19, BCM15, YTD17, ANH^{+}19, CLW^{+}19, CL18, CLH^{+}11, CS17, DPS07, LLH^{+}18, LLL^{+}19, LLYL09, SJVR15, SSMB15, TVCB18, WBD^{+}19, XWD^{+}12, YWLQ18, YESG^{+}19, Zhu19, dSGD14]. non-binary [LLH^{+}18]. non-cache-coherent [CS17]. Non-clairvoyant [BCM15, dSGD14]. non-deterministic [ANH^{+}19, SSMB15]. non-dominated [SJVR15, WBD^{+}19]. non-Gaussian [Zhu19]. Non-GPU-resident [YTD17]. non-intrusive [CLW^{+}19]. non-life [YWLQ18]. non-linear [CL18]. non-Markovian [DPS07, XWD^{+}12]. non-memory [CLH^{+}11]. Non-minimal [BVIB19, YESG^{+}19]. non-singleton [TVCB18]. non-uniform [LLL^{+}19, LLYL09]. Nonblocking [RTET15]. nonblockingly [DGW16]. noncooperative [Zen19a]. Nonintrusive [TC12]. Nonlinear [LW13, CHM15, DS15, VDPC03, ZCXH17]. nonnegative [ZF18]. nonrepudiation [HHPL16]. nonseparable [CMVRRVGI17]. nonsubsampled [PJW^{+}14]. nonuniform [XGXH15]. NORIA [MRMC15]. normal [SRR19]. normalized [PLL14]. North [ZZYW10]. NoSQL [HWZ^{+}15, HWY^{+}17]. Note [Ano13, Ano15d, Ano15b, Ano15c, LS19b, RBNG15]. note-enterprises [LS19b]. nothing [BOF15]. notifications [BVIB19]. Novel [IC19, ZS17, ABFL17, BOB13, GWW17, GYS^{+}17, HZHP09, HXY^{+}12, HYLG15, HCG07, JML^{+}16, JN03, KT19a, LSH^{+}16, LWZ^{+}17, MJZ17, MTGZ17, Mos19, PZH^{+}15, PSIP16, RMCHMG15, SQS^{+}19, SLC^{+}18, SLC^{+}19, SM19c, VJ19, VRDTB^{+}16, WZL^{+}17a, WW19, WZXZ12, XZHW09, YLZ18, YYLC19]. Novo [LGL16a]. Novo-G# [LGL16a]. NPACI [NCWD^{+}04, PKB03]. Nsp [CLL14]. NT [KF01]. Nu [BTCGL17]. nuclear [DGR^{+}07, GIVRC^{+}10]. nucleation [HTR10]. null [MKO^{+}17a]. NUMA [Cha03, ZJL15]. number [KMRT18, LMO15, PPMH15]. numbers [RTPPH12]. Numerical [LOKW^{+}10, DFLL14, GSB^{+}12, HTR10, MAVG16, MB02, Nob08]. numerics [TGB^{+}10]. NVIDIA [LSH^{+}16, VLMPS^{+}18, VFG11]. NWChem [HKS^{+}12].

O [EFM17, LFG05, BFL^{+}10, BKND16, BDT01, Hic18, JCJ17, LGL^{+}17, LQL^{+}09, LLT^{+}14, LFG05, PLM^{+}19, WTL^{+}16, YLL^{+}18]. O-intensive [PLW^{+}18]. O2O [Yu18]. OaaS [FTR15]. OB [XHCL15]. Obfuscation [CZ15b]. Obituary [OS09]. Object [EB05, HWR03, SWL^{+}01, VJHB05, AM01, AJMJS05, Bac03, BGM03, BKCP09, BP03, CL01, CMPT08, CGS15, CM02, DLH01, GKG^{+}04, HmLGP03, HK02, HPS12, HCK^{+}08, JM19, JL10, KL19, KS04, LVN^{+}12, LK03, LH05, MP03, NMKB03, OCS01, ORdSL13, Pre01, QSMK04, RJ01, dFMSPSW06, SKK01, YB12, GMF01]. object-based [BKCP09, CM02, HCK^{+}08, NMKB03]. Object-oriented [EB05, HWR03, SWL^{+}01, AM01, CL01, CGS15, DLH01, GKG^{+}04, HmLGP03, HK02, KS04, OCS01, QSMK04, RJ01, SKK01, YB12, GMF01]. object-passing [AJMJS05]. objective [DS19, GKP^{+}19, JCVU15, JC07, KOOB15, KMRT18, MHLC^{+}05, RHuR^{+}19, SG18, XLHT17, ZXC^{+}19, vdKEL10]. Objects [TM01, ATSAK15, DPP03, DS15, FHO^{+}15, FRKS12, IR11, LL10, LOSJ17, MS05, NN07, Pun01]. oblivious [HT15]. Observation [SM09, YYXL19]. observations [LOSJ17, vHvdSvL03]. Observing [TV14]. obtain [AMVOSGAC17]. Ocean [JWY^{+}05, DvdS06, KDC17]. oceanographic [vHvdSvL03]. OCI [ZWL^{+}19]. OCR [SS17b]. octree [CC13]. ODE [KW18]. ODMG [dFMSPSW06]. OFDM [CLZ^{+}17, HJTX17, KJ19b]. OFDMA [NS19, ZJL13]. off [AAC^{+}15, DD17, DMR^{+}07, HBKM06, KAL07, XLYL17, YLLZ09, GHMX13]. off-line [DMR^{+}07, zGWXT09]. offer [BKWM19]. official [Cog03]. Offline [Dam11, LJW^{+}17, RLZ15, TY15]. Offline/online [LJW^{+}17]. offloading [JVPI18, MKO^{+}17b, XZH^{+}17]. offset [HM12]. OGC [DCY^{+}08]. OGCE [ZKA07]. OGSA [AKK^{+}07, AAB^{+}05, LHT^{+}09]. OGSA-based [LHT^{+}09]. OGSA-DAI [AKK^{+}07, AAB^{+}05]. OGSI [Slo06]. oil [KCZ^{+}05, MBP^{+}05, PML^{+}05]. OLAP [KLP^{+}08]. oligonucleotide [MKKB04]. Olympic [PdCMdS^{+}12]. OmpSs [ABF^{+}17]. on-board [ZJS11]. On-demand [ASWR12, CCSS10, FSM^{+}19, LL10, MS17a, MS17b, MS18]. on-line [CRC15a, zGWXT09, HK01, VB16]. on-line/off-line [zGWXT09]. on-the-fly [PS07]. One [God12, Hun15, KMA04, SZXG19, CS17, DBB^{+}16, KYBV17, LSK04, SKS^{+}08]. One-class [God12]. One-round [SZXG19]. one-sided [CS17, DBB^{+}16, KYBV17, LSK04]. One-to-all [KMA04]. Online [HZL19, HL19, KTB17, LGL16b, LLX15b, PCGE18, RS16, AQRA^{+}18, AAHA18, BDL^{+}15, BB12, CZJY19, CS13, DGM18, GGC19, HLF^{+}17, HSL19, Ios11, JSPE15, JWW17, Kar14b, LJW^{+}17, MCXP15, RS07, SZR16, SZT19, TC17, THQ19, TJZ^{+}19, TBTZ18, VS19, WYZ^{+}17, WJXZ18, ZW09, ZXW16a, ZLH^{+}18, ZFW^{+}17, ZWL^{+}19, dSGD14]. ontologies [FTR15]. Ontology [FTR15, MPS11, MST15, AM15, AHH14, DHC11, DH13, FTRA15, KGGT12, PME^{+}08, UAW09, XWD^{+}12]. ontology-based [AM15, KGGT12, PME^{+}08, XWD^{+}12]. ontology-learning-based [DH13]. OODB [mLGP03]. OOLKIT [ABF^{+}10]. Open [BFG01, BAZ18, BZB17, BDP^{+}14, DGA^{+}10, KMJ^{+}17, KZY15, MRJ^{+}14, Men03, MGM^{+}08, Nob08, PSLC11, PPC^{+}15, PPR19, TTL06, YWA07, ACF^{+}07, CEG^{+}05, DT15b, Lee09, MM10, SKNH09]. open-source [BDP^{+}14, Nob08, PPC^{+}15, TTL06, YWA07]. OpenACC [CGK^{+}16, JCP15].

OpenCL [ABDP15, CBIGL19, FE17, FVLS15, dFdOSR^{+}19, JKM^{+}17, LL16c, SAP16, WJP14, ZWL^{+}17]. OpenCL-accelerated [ZWL^{+}17]. OpenCL-based [JKM^{+}17, WJP14]. OpenFlow [GCWE15, NIIU17]. OpenFOAM [XWX^{+}19]. opening [LZC14]. OpenISA [AMB^{+}17]. OpenMP [CLYC16, CBPP02, FSG19, GG09, HDDG09, JCP15, KOB01, KBVP07, KBG^{+}09, KC06, LHC^{+}07, LL01, MLC04, Nob08, VDL^{+}15, YWC11]. OpenMP-like [KOB01]. OpenMP-oriented [MLC04]. OpenPOWER [KS19c]. OpenStack [BB15, MML^{+}17, TKB16]. OpenTuner [BAG17]. OpenUH [LHC^{+}07]. operating [Cha03, LBDS15, MW18, PT12, SZR16, YL01]. operation [LME^{+}19, LZG^{+}19, LWLZ11, LWL^{+}19, ON02, OKW18, PCVZ^{+}04, SRM^{+}15, SSMB15, YYS15]. operational [YGL05]. operations [AAI12, DHM14, DÇK^{+}18, HSV^{+}19, HKRR08, JLT06, KPNS18, KLDB10, LZY^{+}16, OK15, SGCA^{+}16, ZX11]. Operative [CRC^{+}15b]. operator [ABFL17, ABMdAMF19, CLNR18, DPS16]. operators [PN19]. opinion [CDdW17]. opinions [ZTM12]. Opportunistic [EB10, CC10, CPD^{+}17, CCM^{+}17, CEB^{+}18, DKKL06, FBC10, dAGC11, HM12, MBB19, NQL^{+}17, PGK11, SMH^{+}19, SWS^{+}18, TYHL12, ZQLZ12]. Opportunities [YWT^{+}12, LH05]. optic [SCS17a, ZBZ^{+}15]. Optical [AS15, LLN^{+}14, BKA19, CKRO13, FAM^{+}18, GDD^{+}04, KJ19b, OORVB14, RLVRGÁ14]. OPTICS [LHZX19]. Optimal [BB12, BDP18, CCCW13, CLNR18, KB06, KB17, Zen19a, AMVOSGAC17, CSBL12, CW11b, DKJ13, ER12, JR19, JL10, JKV^{+}15, KA16, LCW^{+}19, LS15, LQL^{+}15, PTL^{+}16, RCA^{+}12, SJ19, SJW18, SVY19, WXLM19, XWH^{+}17, YIN19, YYXL19, ZQK15, ZWL^{+}19]. optimality [Mal05, Viv03]. Optimally [MUKY18]. optimisation [EFM17, GCWE15, GKP^{+}19, GvDHS12, YOBS16]. optimism [LLT09]. Optimistic [SSMB15, RM11, XPS^{+}15]. Optimization [DVD^{+}12, KAA19, MYY18, MB18, MO02b, OA02, PSM^{+}11, PXY^{+}07, ZDX12, ALKD16, ANK^{+}17, ANH^{+}19, BLL18, BKS18, BSP11, dCPD13, CSL12, CLW^{+}18, CLF^{+}19, CL07, CEM^{+}17, DPM17, DBH^{+}17, EN19, EQW^{+}18, GWW17, GRQ19a, GPP^{+}18, HLL^{+}15, HAA^{+}07, JLT06, KT19a, KHL17b, KKT13, KYBV17, KSB^{+}19, LWY^{+}16, LZW^{+}17a, LLLyL16, LGL16b, LSMVML15, MB17, MS17a, MS18, MHLC^{+}05, MRL16, MS17c, MT19a, MBP^{+}05, MCB14, MCAB^{+}02, MP17, NRR15, PN19, PWH18, PLR^{+}14, QSMK04, RK15, SWU08, SL18, SD11a, SG18, TLX^{+}17, TPT^{+}18, TV14, VJHB05, WSL15, Wan18b, WCCL05, WMvP^{+}09, WZYG19, XDE^{+}04, YYZH19, YYZ^{+}17, YPLJ11, ZHT08, ZT09, ZS17, ZZC^{+}17, ZDL19, ZXC^{+}19, ZCH^{+}18, ZYLY18, sCR19, MS17b]. optimization-based [TV14]. optimizations [JCVU15, KKL09, LL16c, NPL19, SAdB^{+}16, VHBB03, VCW13, dARP17]. optimize [BBW19, LS19b, TLM17]. Optimized [GAB19, JKD19, KL12a, ZJKL10, ABF^{+}10, BWD15, FNI17, IHB15, JK10, KCB09, KMJ^{+}17, KKW^{+}14, KG19, LH17, ML19a, MSPDP19, MFC18, PDC16, PS19a, PS19b, VS02, YWLQ18, ZWL^{+}15]. Optimizer [KB17]. Optimizing [BH09, BYN^{+}17, BBK11, Cha03, CQXW14, CCG^{+}08, GE06, HM12, HWZ^{+}15, ITK09, KHL^{+}17a, KR11, LXK^{+}19, PSCK^{+}15, RKS02, RC09, RSMFE^{+}12, SK09, SRL^{+}14, TK10, VŠ11, XY17, ZYZ^{+}12, BDZ19, CSC^{+}17, DAL15, EDBS08, LF15, LXW^{+}16, LHC^{+}07, MSB17, TZK16, WTN07]. optimum [SS17b]. option [CCO15a, HLCW15, LL16a, TTPJ16, ZO14]. options [DCJ12, PW12, TZKH12]. Opto [PWH18]. Opto-electric [PWH18]. opus [SVB19]. orchestration [JDB16, LM08, MK15a, PPC^{+}15, RBNG15, SHP14]. order [BBSW17, CCM^{+}17, KHHC13, LW13, MSV^{+}10, PCT04, WFKS18, ZCXH17, RC09]. order-based [PCT04]. ordering [KYBV17, RMCHMG15]. organization [CCTW11, DDX^{+}06, MS19, PLY13, PRT09, ZBC^{+}07]. organizational [XZZ^{+}16a]. organizations [CG10, GRSB09, PCH^{+}08, ZYN^{+}07]. organized [KOO12, KYM17, LAM^{+}09]. organizing [HM16, PB12, RIFR10, XDE^{+}04, ZWMT12]. oriented [AM01, AAHRW04, ACS10, BR10, BGM03, BAVM11, BM08, BML08, CL01, CLTT13, CGS15, CLH^{+}08, CBIGL19, DMRS15, DLH01, DZL^{+}17b, EABVGV14, EB05, GYM14, GKG^{+}04, GMF01, HmLGP03, HK02, HWR03, HFTQ13, JBL16, JLCA07, KS04, KKK^{+}19, KJŚ^{+}15, KC18, LFH08a, MCWL06, MP05, MLC04, MSS16, OCS01, Pre01, QZYZ16, QSMK04, ROA^{+}07, RJ01, RW10, RHS17, RDP10, SBBE07, SKK01, SWL^{+}01, TTV08, TGB^{+}10, WBHW08, WZZL13, WLZ^{+}18, YLJZ13, YB12, ZFT08, Zhe16]. Origin [LL01, LSK04, PIH04]. origins [Arz17]. orthogonal [LZW13, LCJ14, RRR04]. orthologous [COdO^{+}11]. OSD [PGF19]. other [KHW05, Sod05]. out-degree [ZWLY16]. Out-of-core [ABC^{+}15, GdMK^{+}18]. out-of-the-box [XHCL15]. outbreak [SMS^{+}19]. outdated [HZL^{+}16]. outlier [SL18]. output [AAI12]. outreach [AMRT14]. outsourced [HKA19b, JSR^{+}19a, KTZ^{+}18, LJW^{+}17, QZDJ16, WDLL10, WYBS16, YLWZ18, YSQM19]. outsourcing [LSC^{+}19, LWW^{+}19, SKB^{+}17, SWW^{+}16, WZL^{+}17b]. over-decomposition [TSL^{+}19]. overcome [SMH^{+}19]. overflow [LWW06]. overhead [ALYD17, MA15, Tan12, YLLZ09]. overheads [LLdA08]. overlapped [GBFP09]. overlapping [PGW06, Yos06]. overlay [KA16, LWF^{+}15, RIFR10, RH07, SAM^{+}17, VvSI07]. overlays [BDF15]. overload [QCB17]. oversubscription [DGW16]. overview [DCG15, SWHL16]. owned [HJH19]. ownership [PNB04].

P [Ano06, ZDL19, Kac11, RGL^{+}15, CWL03]. P-GRADE [Kac11]. p-hub [ZDL19]. p-Jigsaw [CWL03]. P2P [Ang08, BGdCCA11, BKH08, CLX07, CZ11, CT11b, CLW^{+}15, Del08, JLHH14, KA16, RH07, XPBS11, XPS^{+}15, XBZ10, ZEB10, dAAVS12]. P2P-based [XBZ10, ZEB10]. P2Pedia [DCEK15]. P2PGrid [CLX07].

P2PPerf [EDBS08]. P2PScheMe [dAAVS12]. PaaS [DPM17, PB16]. PAC [WLL14]. Pacific [IUCH^{+}17, PC17b]. package [PSM^{+}11, Sch04, WO14, Zhu19]. package/access [Sch04]. packed [BGGL07]. PackedObjects [YKA^{+}19]. packet [ALL^{+}15, CKRO13, GGLD11, HLG17, LZL^{+}17b, LSC^{+}19, STO17, ZL12]. packing [RGX^{+}17]. pad [YWY^{+}10]. PADP [YXL17]. PADTAD [Ur07]. PAGCM [RZL^{+}19]. page [ID18, PZZ08, PZZ10, ZZD^{+}17]. PageRank [LSXL17, PCD15]. pair [ANH16]. pairing [WWC^{+}19]. pairing-free [WWC^{+}19]. pairs [QZYZ16, SSB^{+}14]. pairwise [AMHC11, SGR19, YAG19]. PAKA [XZJ11]. PALF [LYL07]. Palirria [VB16]. Palladio [BWEB14]. PALM [BPD06]. palmprint [SC19]. panels [WWY19]. Papers [AHP^{+}13, BHD13, BKZ^{+}13, BDB^{+}13, CWZL13, CCCW13, dOCPFJ13, CLTT13, CKRO13, CAG^{+}13, CMT13, CS13, DRZ13, DRS^{+}13, DVL13, DLM13, DH13, EBMD13, ETR^{+}13, HL13, HMPPT13, HFTQ13, JJGL13, KM13, LXRJ13, LMKT13, LDXC13, LW13, MWL^{+}13, MS13, MSP^{+}13, MFG^{+}13, MISV13, NSSAK13, ODS^{+}13, PLY13, PRD^{+}13, PB07b, PK08, Puf13, QLS13, RMP^{+}13a, RHRB13, RTMZ13, SRF13, TWB13, VCW13, WAS07, WZZL13, WC08, WCLC13, XBXS13, XW13, Xu08, XJZ13, YLD13, YLR^{+}13, YLJZ13, ZWL^{+}13, ZLY^{+}13, ZLN^{+}13, ZHZ^{+}13, GZX17, PDD14]. PAR-3D-BLAST [SL14]. paradigm [CKL19, CKBB14, FJ05, GED^{+}18, KSK19, PRS01, TBH^{+}18, WW19, ZBP06, ZF18, ZDC^{+}09]. paradigms [CS13, GWC^{+}11, LMS18, MLS^{+}12, PRS01]. Parallel [AMHC11, AMTM17, Ano15a, AMZ19, AT18, BGGS14, BJ18, BHQOS15, Bok12, BDY02, BLKD08, CPEA18, CC13, CMVRRVGI17, CMPT08, CACC11, CCW06, CSTV06, ČSMK17, DCG11, DSO^{+}01, FSM^{+}19, FLMRC02, dFdOSR^{+}19, FCT^{+}02, GKSR14, GA08, GSV03, GKK09, GJ17, GZX17, HLCW15, HM04, ISS^{+}02, JN03, JKV^{+}15, KRW17, Kni06, KLP^{+}08, KB12, LSXL17, LJPP16, LKPM09, LS05, LBH07, MMW16, MKB01, MSD^{+}18, MQOQOH01, MSK19, MSM^{+}14, NO02, Nak02, OLG^{+}15, PCVZ^{+}04, PIH04, Pla08, PPP10, PA08, QSMK04, RSM01, RTPPH12, SPMP11, SKK02, SG16, STS19, Str11, SN16, SEF^{+}14, TTD^{+}05, TFDA07, WZ04, WCH^{+}07, WT15, YAG19, YA04, YT19, YWBS19, ZP07, ZLZ^{+}17, ZZZ^{+}15, vAVS12, AAP13, AA16, ABF^{+}10, AML^{+}15, ABF^{+}17, ABC19, ABV05, AC09, Ano06, ADK^{+}16, ATİ17, ACIC^{+}13, dRADFG17, BGGL07, BJ01]. parallel [BFR05, BCD^{+}02, BG14, BBCG02, BB02, BB04, BCM^{+}07, BdL06, BV11, BKND16, BZB17, BCC^{+}05, BDV02, CML^{+}10, CDA09, CT11a, CLL14, CLYC16, CNKJ18, CCW04, CZL^{+}17, Cho01, CGS15, CNP^{+}15, CZL12, CSMK19, CLRB15, CCT15, CRV15, DDP^{+}06, DCJ12, DVB14, DLX^{+}16, DPP03, DS04, DdB01, DCK12, DLM13, DDF^{+}15, DZL^{+}17a, DLW19, DvNM^{+}11a, DT01, EMEY14, EFM17, EOD^{+}19, ESG11, FBH^{+}01, FJ05, Fec12, FLYL16, FMS15, FBV^{+}13, FYKW15, GWW17, GL19, GMT07, GG09, GMMT17, GQ04, GDMT^{+}12, GM04, GE08, GvDHS12, GWC^{+}11, GPZ04, HKVW16, HMM^{+}09, HPVRPF14, HdV13, HSHT14, HLO^{+}16, HW16, ITK09, ID18, ISO^{+}14, IPGCMW18, IT03, JC07, JML^{+}16, JLCA07, KGK17, KOB01, KHZN06, KM03, Kes04, KL12a, KTR11, KOOB15, KRS11, KPS14, KYBV17, KR11, LW05, LLRS03, LK03, LPH09, LM07]. parallel [LDZ14b, LL15, LGQ^{+}17, LWL17, LTZ^{+}19, LG08, LF17, LCW^{+}17, LLH^{+}17, LHH^{+}17, LMX^{+}18, LSP15, LB11, LL19c, MRL16, MST^{+}05, ML19a, MJL01, MRS03, MMSG17, MBC^{+}14, MSCS18, MT19a, MCG18, MO02a, MGS19, MKKB04, MM17, MCCD18, MPSGD14, MJD15, MCSML07, MCC16, MFGE19, MLYL17, MLWA19, MAK18, MvWL^{+}10, MDL^{+}10, NSBR07, NMM^{+}10, NC05, NvV09, NNON02, NDT^{+}16, ODS^{+}13, PW12, PSG03, PPMH15, PS10, PV04, PRC^{+}14, PSM^{+}11, PPST09, PGF19, PSS^{+}18, PT12, PAC^{+}17, PSCK^{+}15, PZ17, QW17, QH10, RR15, Ree01, RZL^{+}19, RR11, RGL^{+}15, RLVRGÁ14, SL14, dFMSPSW06, SV09, SAB15, SJW18, SHT^{+}17, SRM13a, SRM13b, SER15, SK04, SMM18, SIS19, SCBH09, lSsCY17, SM03, SL18, SBDP15, STTW18, SİM^{+}07, SGD^{+}18, SVS^{+}08, SLM05, SK18, SNGR18, SS15c, TYL^{+}15, TLX^{+}17, TCSBMG17, TY15, TKS18, TCH^{+}13]. parallel [TSKM18, TF03, TBH^{+}18, VCP16, WKB^{+}19, WLLL15, WDG^{+}14, WCR^{+}14, WBO16, WWLD18, WZLL18, WLF19, WHX19, WLW14, WMDM07, WLL03a, WS17, WS19b, XPBS11, XCHY13, XXLL17, XLYL17, YYCH10, YWC11, YGG14, YXL17, YYC^{+}19, Yos06, YL01, YB12, ZF14, ZYZ06, ZP06, ZSZ^{+}14, ZYW^{+}16, ZZC^{+}17, ZGZ19, ZYLT06, ZZ17, ZWL^{+}17, Zhu18, dCGKG06, dOOO^{+}12, vHvdSvL03, vHMB08, CM07b, JWY^{+}05, KT19c, PNL10, SMBT07, SJPB17, TL14, Ur07]. parallel-in-time [MCCD18]. parallel/distributed [MCSML07]. parallelisation [RVVPD^{+}17]. Parallelism [BPL12, VRSJ15, ABFL17, DS04, FJ05, GVC10, HJB12, MB14, NSN^{+}17, Nev17, OGA^{+}01, VB16, GDD^{+}04, MMS07, PGdCJ^{+}18]. Parallelism-based [VRSJ15]. Parallelization [GB07, HKB07, Riz04, SS15a, SSK11, TRH^{+}02, ZZL^{+}17a, ZZL^{+}17b, AUHWJ19, CEM^{+}17, DT15b, FSG19, HC07, JSR19b, LF17, MUKY18, PNJP19, PSJM13, SMBT07, TRW07]. parallelize [SJVR15, SPW09]. Parallelized [GPV09, MKAKG14]. Parallelizing [BHL^{+}09, BHPS14, CCP^{+}15, DK09, KSS^{+}17, LXW^{+}16, LZJ^{+}18, LLRS19, BY12, Dut17, YTF^{+}01]. ParaMEDIC [BFL^{+}10]. parameter [AAE^{+}09, GG19, HML20, ISO^{+}14, KHL^{+}17a, LLH19, RMCN^{+}07, SL18, TSL^{+}19, WZYG19, YGG14, YK10, Zhu18]. Parameterizable [ZCL14]. parameterized [CHM15, SS07]. parameters [BAG17, JCVU15, KCD19, OORVB14, WLZ17, YC19b, ZWZ^{+}18, ZDHJ18]. Parametric [vEGW06, IÁE11, KS04]. parasite [LLRS03]. Pareto [KB17, MHLC^{+}05, Mal05, RLVRGÁ14, TZK16]. Pareto-based [KB17, MHLC^{+}05]. parity [LDZ^{+}15, PK17]. parity-check [LDZ^{+}15]. Park [Mar19]. parsimony [SJISVR17]. Partial [ASP19, ZHW^{+}16, BJ01, DFLL14, KKW^{+}14, LWLZ11, MCG^{+}08, PS10, SGR19]. Partially [XLL^{+}15]. participatory [BvIF10, CGOF15]. Particle [KHL17b, AA16, BDY02, BDY03, dCPD13, CDP17, EN19, HR18, LWQS19, MLVB05, PWH18, QH10, RK15, VDL^{+}15, WZYG19, XDE^{+}04, ZHT08, ZT09, ZS17]. particle-in-cell [MLVB05, QH10, VDL^{+}15]. particulate [MZG19]. partition [BTCB16, DZL^{+}17a, DLW19, HWZ^{+}15, LZW17b, PZZ08, PZZ10, ZCZ^{+}19]. partition-based [ZCZ^{+}19]. partitioned [ZDB^{+}14]. partitioning [ARPPM17, GPZ04, LZZ^{+}17, LLL^{+}19, PMAL14, SJW18, SKK02, SHC^{+}16, XWX^{+}19, ZLKK17]. partitioning-based [GPZ04, SHC^{+}16]. Partner [HAJL16, HJTX17, MABP13, MBP16]. party [FIO15, HLC19, WLWX14, WLWX16, ZZC15]. Pascal [CCO^{+}15b, LGP19]. pass [CH19, MCWL06]. pass-through [CH19]. Passing [AMHC11, SWL^{+}01, AJMJS05, BCM^{+}07, Gog11, NMKB03, OKW15, RMG^{+}10, RM11, SVS^{+}08, SSZ14, HSBMR08]. passive [KJ19b, ZCL^{+}19]. password [Eng15, FIO15, SVY19]. password-based [FIO15]. past [DLP03]. pasts [BFU07]. patches [CSL^{+}19]. patent [QZYZ16]. Path [DT15b, BM07, CSL^{+}18, HLZD18, KC18, LZH^{+}15, MSMA19, RC09, SJ19, VO15, YYZH19, ZXW19]. paths [GP07, ZWXS19]. pathways [LTM^{+}14]. patients [KM19, ZBZ^{+}15]. Pattern [TTR^{+}10, ATKH^{+}17, dRADFG17, BBG17, BLL18, CGS15, DCFC08, FBC10, GdMK^{+}18, LWSZ19, SAdB^{+}16, THQ19, WZ16, ZHC^{+}18]. pattern-based [BBG17, BLL18]. patterns [AAF^{+}07, CT12, DT15a, GYS^{+}17, HHWZ08, LJPP16, SRR19, WMDM07, ZJKL10]. PAWN [JNUH17]. payload [JNUH17]. payload-based [JNUH17]. Payment [CDL08, Du18b, HH19, LHB^{+}19]. PBlaman [BWEB14]. PBS [Cla18]. PC [HON04, LKYS04, SNGR18]. PCA [HYGF19]. PCR [ALVY05]. PDA [PB07a]. PDE [ALKD16, EFY17]. PDE-based [ALKD16, EFY17]. PDEs [CNKJ18, JN03]. PDNOC [XLL^{+}15]. PE [KSM^{+}19]. peak [LM08, YZZ^{+}10]. PEARC17 [DN19]. pedestrian [CXT^{+}18]. PEDISWESA [FK19]. Peer [Man08, Zha08, BM10, BAZ18, CRC15a, DCEK15, DS07, DvNM^{+}11b, EDBS08, EB05, FG16, FPR05, GS08, LDXC13, LNKZ08, LFZ07, LAM^{+}09, MABP13, MME13, NR08, PGW^{+}08, QMK12, RGV09, SAC^{+}07, TLWZ14, Tru15, XLL^{+}12, ZK08, ZCS06, dP06]. Peer-to-Peer [Man08, Zha08, BAZ18, CRC15a, DCEK15, DS07, DvNM^{+}11b, EB05, FG16, FPR05, GS08, LDXC13, LNKZ08, LFZ07, LAM^{+}09, MABP13, MME13, NR08, PGW^{+}08, QMK12, RGV09, SAC^{+}07, TLWZ14, Tru15, XLL^{+}12, ZK08, ZCS06, dP06]. peer-to-peer-based [BM10]. peer-to-peer-distributed [EDBS08]. PeerfactSim.KOM [FG16]. PEGASUS [TBK^{+}15, KDG^{+}08, LPS^{+}09, MCD^{+}15]. PEKS [ZQD16]. PEN [XL17]. penalties [KF15, LFG05]. pencil [JKD19]. PENNANT [Fer15]. People [Li17, ZLC17a]. pepper [SS19b]. Peptide [MHLC^{+}05, WJP14]. peptides [MIGA18]. perception [MW18]. perfect [JCVU15]. perfectly [ZLKK17]. perforated [WWY19]. perform [CBIGL19]. performability [MS17b]. Performance [ALKD16, AHP^{+}13, AF14, AC06, AFG^{+}05, AM07, BA04, BB02, Ber07, BBdS^{+}17, BSP11, BY12, BD04, BUVS10, BLSP11, CML^{+}10, CGK^{+}16, CRCC09, CCW04, DDE^{+}12, DMA13, ESG11, FMM08, FGZ^{+}18, FK19, FN13, FJG^{+}13, GG07, GLMT15, GMVRGS15, GS04a, GRS^{+}17, Guo19, GHPR05, HJB12, HKVW16, HKS^{+}12, HK01, HFR^{+}17, IHB15, JFI^{+}08, JLHH14, KMBR19, KAL07, KS02, KC06, KYBV17, KSB^{+}19, KSM^{+}19, LBOE18, LLRS03, LSS05, LHL10, LGP19, Li04, LWG^{+}15, LL16b, LJML10, LKYS04, MST^{+}05, Mar05, MDH^{+}16, MLY10, MJ19, MWW10, MN10, MNL15, MWLS11, Nel05, NJ05, OCC^{+}05, OAS^{+}15, PFU^{+}05, PGB03, PHGK10, PW05, QB12, RGAK15, RK01, RMCN^{+}07, RVVPD^{+}17, SG19, SZT18, SFCAV16, SIOS02, SWB12, SFN12, SGR19, TSL^{+}19, TWN07, TMR^{+}07, WKT08, WXLM19, YWC11, YOBS16, ZPG10, AA16, AKK^{+}07, ABF^{+}10]. performance [ABDP15, AP10, ABMdAMF19, AAC^{+}15, ADI^{+}14, AC08, AKM^{+}06, BLCC19, BCD^{+}10, BHJ^{+}16, BB12, BFM^{+}10, BM08, BBM19, BS10, BDT01, BBD10, BDG^{+}10, BPT^{+}16, BWEB14, BXQ17, BDH15, BPD06, CMW02, CC13, CHP17, CPG^{+}16, dOCPFJ13, CKOG10, Cha03, CLYC16, CNKJ18, CBPP02, CNG13, CXPL15, CLS14, CL16, CEG^{+}05, CFP^{+}03, CKD^{+}19, CRGR^{+}12, CMS17, DD17, DLPV07, Dam11, DIM18, DRZ13, DC19a, DDX^{+}06, DS02, DMR^{+}07, DPS16, DL10, DMD16, DFLL14, DZM^{+}15, DLT^{+}16, DMMA17, DPP^{+}19, EGGA^{+}04, EMS11, ESG17, EMS15, ETR^{+}13, FBH^{+}01, FE17, FE18, FLYL16, FMP10, Fox12, FBS16, GFBR10, GWW^{+}10, Ger05, GF07, GMT07, GO10, GKSR14, GAM17, GGV14, GBMM15, GCN09, GA08, GWVP^{+}14, GTA10, GW15, GYP^{+}16, GVP^{+}14, HM12, HDDG09, HTHW16, HMM^{+}09, HSV^{+}19, HPS05, HTI05, HvNJB15, HLHC12, HML20, HY12]. performance [ID18, JCJ17, JWY^{+}05, KF15, KDC17, KA09, Kar14a, KHZN06, KHW05, KL12a, KCB09, KSM^{+}08a, KTR11, KW01, KF11, KAMB19, LLD19, LL05, LM07, LSH^{+}16, Li18, LLH19, LLY^{+}19a, LSS15, LHBW15, LLH^{+}15, LHB^{+}19, LFH08a, LQL^{+}09, LAL02, LXK^{+}19, LL01, LKJ03, LSK04, MSMA19, MBP16, Mal05, MMMP01, ML19a, MLVB05, MMSG17, MBC^{+}14, MSPDP19, MSB17, MJD17, MOK04, MO02b, MDV07, MA15, MMG^{+}18, MFC18, MKSS16, MB02, MM10, NMM^{+}10, Not16b, OFR^{+}17, PSRR14, PLM^{+}19, PSS^{+}19, PGC^{+}19, PPBB14, PLL17, PSS^{+}18, PK17, PBF15, PS19a, QXXZ16, QWZ^{+}19, RVRD10, dRRdCRR16, RCB03, RMW19, RGL^{+}15, RCLSK16, RM03, RGB^{+}15, SM02, dFMSPSW06, SAB15, SRF13, SER15, SCC^{+}10, SLGL16, SCBH09, lSsCY17, SCGZ19, SZG^{+}19, SSK11, SWD^{+}15, SM09, SİM^{+}07, SSB^{+}14, SFH13, SFT15, SPQ^{+}17, SB17, SRL^{+}14, STL^{+}15, SK18]. performance [SLM^{+}10, SWD^{+}17, SS07, TTD^{+}11, TKZQ17, TYHL12, TCSBMG17, TPT^{+}18, TTPJ16, TRW07, TSS18, TF03, UA18, VS02, VJK13, VDL^{+}15, VdSK^{+}05, WFJ^{+}17, WHX19, WK07, WTN07, WCL^{+}10, WTL^{+}16, XWD^{+}12, XJAJ18, XZZ^{+}16a, YYS15, YBC^{+}07, ZF14, ZCC^{+}06, ZCL14, ZZ16, ZL12, ZLZ^{+}19, ZZD^{+}17, ZJL15, ZDX12]. performance-aware [KL12a, LFH08a]. Performance-based [LHL10, NJ05, YWC11]. Performance-driven [GLMT15]. performance-energy [AAC^{+}15]. Performance-influence [GRS^{+}17]. performance-oriented [BM08]. performance/cost [GWVP^{+}14]. performances [CGIP16]. period [HLW^{+}19]. periodic [MFGE19, RF15]. peripheral [Sin10]. periscope [GO10, LGG16]. Perl [MTD^{+}02]. permeation [KF11]. PERMIS [CZO^{+}08]. permission [CJC^{+}18, YWL^{+}17b]. permission-based [CJC^{+}18]. permutation [AMTM17, MCB14, PSM^{+}11]. persistency [DFLNP07]. person [CZQ17]. personal [CGOF15, GHMX13, SVY19]. personalization [FHH15, WLDL08]. Personalized [LXW17, AMBT17a, AMBT17b, CLMM12, God12, Yu18, ZQR^{+}19]. perspective [CLW^{+}15, WDW^{+}15, YLLC18, ZSL^{+}10]. perspectives [LZWD^{+}15, VRSJ15]. perturbation [PZHS18]. pervasive [GZG^{+}16, JQSP08, RCB^{+}04]. pessimistic [ZQW^{+}17]. petaflop [GKS09]. petascale [EDB^{+}14]. Peter [OS09]. Petri [DPS07, EB10, GHB^{+}06, MRMC15, MS19, PP17, TDL^{+}18, WDQ^{+}18, XWD^{+}12]. Petri-net-based [GHB^{+}06]. petroleum [ZQD^{+}17]. PETSc [MCCD18]. PGAS [LGLA15, MRH14]. PGAS-FMM [MRH14]. Phantom [GBJ19]. phase [LXW^{+}16, Li19, MWLS11, ZCL^{+}19]. Phi [CBIGL19, KKW^{+}14, MTK16, RMW19]. phishing [CAKH17, HAK19]. Phi^{T M} [GdMK^{+}18]. photodiodes [PIH04]. photographic [Boc19]. photography [Boc19]. photonic [AMSR14, CSAC19, DPP^{+}19, GGFPGB14, MVWJ14]. photonic-electronic [GGFPGB14]. photonics [BG14]. phylogenetic [BAD^{+}11, SJVR15, SJISVR17, SLM04, SLM05]. phylogenies [MKAKG14]. Physical [HJTX17, SZ11, ANCA19, DZW^{+}11, GOLL17, GWVP^{+}14, IT03, LCC^{+}18, SZMF19, ZX11]. physics [BBdS^{+}17, Fer13, Sod07, TB12, VDPC03]. PI [ZZYW10, DCA17]. PID [LWW06]. Piecewise [PAC^{+}17]. pilot [RMCHMG15]. PIPE [SMBT07, ZYW^{+}16]. Pipeline [CGS15, TCBR11, WWG^{+}11]. pipelined [DKJ13, GPV09, MKSS16, RCA^{+}12]. Pipelines [AGMR05, GVC10, KKL06]. pipelining [TDM^{+}19, YWY^{+}10]. pivoting [DFLL14, Has17]. pixel [Pla08]. pixels [FRKS12]. place [BCI^{+}18, DVL13, GW19, LTL^{+}17, PSHL11]. placement [EMS15, JLL18, LXK^{+}19, MS17c, SHST13, WZLQ16, WSW^{+}12, XDJL18, XTB17, ZJS^{+}17, ZW17, ZWH^{+}17, ZFW^{+}17]. Planning [MLVBW12, BPB08, DHH^{+}13, JR19, LZH^{+}15, PPST09, XLZD13, YYZH19]. plant [DGR^{+}07, GED^{+}18]. planted [DRZ13]. Planting [CRV15]. plasma [RR11, Sod07]. plasmas [RMCHMG15]. plate [GHLS19]. Platform [GPW03, MZK16, ACFT15, AFG16, ATNW11, AMB^{+}17, BRK^{+}17, CSMB15, CJZZ10, CS15, DJM12, DCA17, FÁBE11, HVM^{+}15, LTL^{+}17, LHB^{+}19, MCC^{+}15, MW18, MD02, NO02, PPC^{+}15, PC17a, RCM12, TBTZ18, WWL^{+}15, WZLQ16, XBB13, YP10, CEG^{+}05]. platform-as-a-service [ACFT15]. Platforms [BJ18, AYN^{+}14, ALMT19, ATİ17, ACCM17, BEQOR13, BCM15, BHQOS15, CHP17, CJX^{+}19, KSR14, LQL^{+}09, MRL16, MB12, MTT15, OFR^{+}17, PRV11, QLS13, SER15, The01]. play [WYAB07]. pleasingly [GWC^{+}11]. PLPP [MMS07]. PLS [ZXW19]. Plug [WYAB07, BKM^{+}07b]. Plug-and-play [WYAB07]. plug-in [BKM^{+}07b]. Plugging [BKM^{+}07b]. plugin [CS15, MWRK18]. plus [Yu18]. PMaC [PTL^{+}16]. PMAM [BL17]. PMU [LWS19, YYXL19]. POGGI [Ios11]. point [BTG06, LCM12, LDZ14b, MMS07, OTG^{+}07, TLM17, WCR^{+}14, YZZ^{+}10]. point-set [WCR^{+}14]. points [LCJ14, QML^{+}17, Soo16, TLX^{+}17]. points-based [Soo16]. PoLAPACK [Cho01]. Polder [IBvA^{+}02]. policies [BBC16, KQR^{+}17, KKV13, NNvVdA09, OSK^{+}01]. policy [BA18, BBC16, CVK15, GLS^{+}19, HDX^{+}17, LFWS15, RAFD14, WZC16, YL01]. policy-based [CVK15]. polling [KSK19]. pollution [BGdCCA11]. polygon [YIN19]. polygonal [ZKWK17]. polygons [CZL^{+}17]. polyhedral [CSC^{+}17]. polyimide [ZCW^{+}18]. polymorphism [KS04]. polynomial [CH04, YLWZ18]. PON [KJ19b]. Pool [HR06, VCP16]. pools [KR04, TK10]. POP [JWY^{+}05]. popular [PBF15, PNJP19]. population [XBXS13]. porous [GEBA17, LLY^{+}19a]. port [cDrLyC^{+}19]. Portability [JPS17, ABDP15, CGK^{+}16, FE17, FE18, GFBR10, JWY^{+}05, MMSG17]. portable [BMV03, DPP03, DLZ16, DT01, LHC^{+}07, LTK17, RMG^{+}10]. Portal [GBB^{+}15, Nov02, SPR^{+}07, AHB^{+}10, AC02, ACC^{+}07, BAD^{+}11, BFM^{+}06, CW07, HCD^{+}02, HAA^{+}07, Kac11, KBH^{+}15b, MCY^{+}10, NRW04, PYF02, PGP^{+}10, YWA07, YBB^{+}07, YLEB14, ZDA^{+}07, ZKA07, vLDA07, ACMA07, CM07a, HBH02, NNTH^{+}02]. Portal-based [SPR^{+}07]. Portals [EMB11, GvHKK11, Tho07, BKM^{+}07a, BKM^{+}07b, MH07, NAP^{+}07, TDM^{+}02, YAA07]. portals/portlets [YAA07]. Portfolio [MSB17]. portfolios [BRCV16, WSRM12]. porting [DHH^{+}13, KOK14, WWG^{+}11]. portlet [WYAB07]. portlets [ACF^{+}07, YAA07]. pose [RK15]. Position [SWS^{+}18, MBB19, XZZ^{+}16a]. position-based [MBB19]. positioning [LWZ^{+}19, MB17, YYLC19]. positive [HZL^{+}16, LZWD^{+}15]. possession [YXL17, ZNT^{+}16]. possibilistic [MZA19]. possibilities [HGT14]. Possible [SCNH07, PFU^{+}05]. post [LTL^{+}17]. postquantum [YZCT17]. posts [LHB^{+}19]. postseismic [ZGRSC10]. potential [BHA15a, RMCN^{+}07, YZ10, ZZY^{+}19, ZGL07]. potentially [CSL^{+}18]. Power [EQORS19, KBB11, LBdM^{+}16, LWS19, MSP^{+}13, PSL^{+}16, TQL^{+}14, TPT^{+}18, TDM^{+}19, ADMQO14, AMSR14, CLH19, ČSMK17, CSMK19, CSAC19, DMW^{+}10, DGR^{+}07, DMMA17, GKG^{+}04, GYP^{+}16, HKVW16, HTHW16, HJV^{+}19, MWRK18, MFG^{+}13, MKO^{+}17b, NK19, PLM^{+}19, PLL17, PS19a, RHZ^{+}17, SLGL16, ŠŽH17, SB18, SPQ^{+}17, WRLS12, XL17, ZMYA18]. Power-aware [EQORS19, KBB11, LBdM^{+}16, MSP^{+}13, TDM^{+}19, RHZ^{+}17]. power-saving [MFG^{+}13]. powered [ADSV16, SR19b]. powermode [JL10]. PPAM [WT15]. PPEM [LLM19]. PPLTCAM [WHX19]. Practical [EA12, FLYL16, JWY^{+}05, LSC^{+}19, XW13, CSB^{+}16, HWZX08, LFZ^{+}17, ZYFZ19]. Practice [Ano06, DN19, FH01, KQR^{+}17, TH10, ASP19, BCCM16, CHPvdG07, Fox12, GTL06, Hun15, JCK^{+}13, LWC17, RKS02, RLC16, TTL05, TDM^{+}02, YDB^{+}13, ZRB19]. practices [GRGP12, LME^{+}19]. practitioners [HMPPT13]. PRAGMA [PC17b, Arz17, IUCH^{+}17, SWP17]. PRAGMA-ENT [IUCH^{+}17]. praise [ECP18]. Pre [AdCPdSD17, PWJ10, YWL^{+}17a, BJD^{+}19, SGCG09, WLP^{+}17, YHHS16]. pre-distribution [BJD^{+}19, SGCG09, YHHS16]. Pre-image [YWL^{+}17a]. Pre-seismic [PWJ10]. Pre-stack [AdCPdSD17]. pre-transformation [WLP^{+}17]. precedence [Hun15]. Precision [BLDW16, ADF^{+}19, BDE^{+}19, KD07, LCM12, Roj19, ZCL^{+}19]. Precision-tuning [BLDW16]. preconditioned [ABF^{+}17]. preconditioner [JN03]. preconditioning [ADF^{+}19]. predicates [ZYZ^{+}12]. predication [ZFXJ19]. predication-based [ZFXJ19]. predict [CDP17]. predictability [WLZ11, ZSL^{+}10]. predictable [HWQ^{+}16, LTK17, MDX14]. Predicting [BHA15a, DGM18, SİM^{+}07, DFC12, FBC10, WYY^{+}19, XDL^{+}11]. prediction [AD02, ACCM17, BPL12, BDTdS13, CDdW17, DMR^{+}07, DKMV07, DJ19, FE18, GPV09, HWL18, JFI^{+}08, KA09, KHL17b, LLX^{+}15a, LS05, MAVG16, Mit17c, Mit19, MV16, NNK^{+}07, PSRR14, SL10, Soo16, STL^{+}15, TZLC15, VGN^{+}16, WZL^{+}17a, WKZL19, cWsThY19, XDP18, ZTM12, ZYZ16, ZACG16, ZHX^{+}19]. Predictive [SMFM18, GED^{+}18, HML20, LLH19, ZXX17]. predictor [BKCP09, ZBZ^{+}18]. predistortion [SZ18]. predistribution [SCS17a]. preempting [SJB14]. preemption [KW11]. preemptive [Bou06, FRdOR^{+}19, KW11]. Preface [BM12, LL13, Nag10, NM10, LS14]. preference [RBDI17, SZG^{+}19]. preferences [ZZL^{+}17b]. prefetching [BKCP09, CM05, Lia16]. prefix [PPR19, WBO16]. prefix-doubling [WBO16]. Pregel [JR19]. pregnancy [LCC^{+}18]. Premia [CLL14]. Preparing [HCD^{+}18]. preprocessing [Akt18b, CV07, LQL^{+}15]. preprocessor [PBSB04]. presence [EZJ^{+}18, JAU19, LGFM05, MOK04]. present [DLP03]. preservation [BLL^{+}19, ZLN^{+}13]. preserving [AD15, BC16, DZC16, HKA19b, LLM19, LSC^{+}19, LW13, LXW17, QGZL18, TJ17b, TC17, WHXzL15, WZC16, WMC17, XZZ16b, YYK^{+}19, YNX^{+}16, Zha19]. Preventing [HAK19, Kin04]. prevention [LYS18, SZMF19, SPW09, WYW^{+}17]. Price [PGW06, BGGS14]. Price-sensitive [PGW06]. pricer [BLDW16]. prices [BGGS14]. Pricing [ATVLM14, PGW06, TZKH12, CCO15a, CL07, DCJ12, DCJ14, HLCW15, LL16a, MB02, PRNM19, TTPJ16, TKB16, ZO14]. primary [CP14, GÖ18]. primary-backup [GÖ18]. primer [SSM04]. primitives [ABDP15, BBCG02]. principal [HML20, LLH19]. principle [MLL^{+}11]. printing [WSZ^{+}18]. prior [Du18a, ZSWS18]. prioritization [KGK17]. prioritized [LHT^{+}09, YYS15]. priority [KW11, MT19b, XWH^{+}17, ZWMT12]. priority-based [MT19b, ZWMT12]. PRISM [VGL06]. Privacy [LLM19, NPM19, QGZL18, SGL^{+}17, WMC17, ZXC^{+}19, AAHA18, AD15, BA18, BC16, BLL^{+}19, BB19, CLW^{+}19, DZC16, DZZL19, GBJ19, HKA19b, JBL15, KWXY18, LWYM16, LSC^{+}19, LSZ19, LWY15, LXW17, MJS19, QWW^{+}16, TC17, WAD12, WHXzL15, WZC16, XAK16, XBK17, XBCW19, XZZ16b, YYK^{+}19, YKD^{+}15, YCQF19, YMLR16, YNX^{+}16, ZLN^{+}13, ZYY^{+}19, ZQR^{+}19, LDZ^{+}19]. privacy-based [MJS19]. Privacy-ensuring [SGL^{+}17]. Privacy-preserving [LLM19, QGZL18, WMC17, BC16, DZC16, HKA19b, LSC^{+}19, LXW17, TC17, WZC16, YYK^{+}19]. private [CFPJ^{+}17, DXWC16, ESG11, GLM^{+}16, HJH19, HJM^{+}11, JJGL13, PCsHL18, SYMA17, TZ16, YWR^{+}19, YLWZ18, ZXRZ19, TKS18]. private-shared [CFPJ^{+}17]. privilege [MLL^{+}11]. Pro [Cla18]. Proactive [VvSI07, CW09, HHKA14, IRB19, SZA08]. PROB [YP10]. probabilistic [ALL^{+}15, CXPL15, Guo19, LNCY11, YZ10, ZCS06]. probabilistically [LLT09]. probabilities [SK18]. probability [ZZL^{+}17b]. probable [BRCV16]. probe [MKKB04, SS07]. problem [AMTM17, ABV05, ACIC^{+}13, BPL12, BIK^{+}11, Bok12, CKRO13, CGK14, CS16, DRZ13, DdB01, FMS11, GP07, HC07, JPWH02, KH12, KHL17b, LSXL17, LAC^{+}08, LL18, LWK15, MPS11, MCB14, MME13, QW17, RGX^{+}17, RM03, RLVRGÁ14, SDB02, SSB^{+}14, TL14, WBD^{+}19, WLLL15, WLLL16, WHXB19, WZYG19, XLHT17, YIN19, YA04, vSB06]. problem-solving [JPWH02, LAC^{+}08, SDB02]. problems [BA18, BWD15, CW07, ČSMK17, CSMK19, CG01, CEM^{+}17, GF07, LZZ^{+}15, MSB17, PCsHL18, SSIH19, SD15, YDS^{+}14, ZS17]. procedure [KKK10]. procedures [NFG19]. Proceedings [Run10]. Process [BR10, ABMMR19, CWZL13, CMB06, CMD11, GWHJL19, HAN19, HRR^{+}11, HY12, ITK09, KSPM12, MBB19, May10, MÍM19, ON01, ON02, RR19, RW10, SB17, TPV17, WFHT17, WDQ^{+}18, XLZD13, ZLH^{+}18, ZLQ^{+}18]. Process-oriented [BR10, RW10]. processes [FÁBE11, HLZD18, IÁE11, IÁBE11, Jos05, SGG07, TALT16, XZHW09]. Processing [LSXL17, SMBT07, WT15, ATVLM14, ACC^{+}12, ADF^{+}13, dRADFG17, BG17, BDW14, BHQOS15, CLNR18, CY15, CRB09, CGIP16, CP14, CPSP17, CTAB16, CS13, DDP^{+}06, DCJ12, DCJ14, DG11, DZJ^{+}15, DZL^{+}17a, DWG19, DL07, DT15b, EMS11, EPA15, GWW17, GGV14, GLZ19, HWL18, HH19, JQL^{+}15, JZMD19, JdM12, JZZL06, Kar14a, KC13, KKL06, KLP^{+}08, LBOE18, LPS^{+}09, LTL^{+}17, LPH09, LOSJ17, LKPM09, LDZ^{+}15, LGL^{+}17, LLH^{+}15, LWLZ11, LSJ16, LMDP19, LPG^{+}14, MAS16, MS17a, MS18, MBMB18, MCB14, MK12, MPSGD14, MWL^{+}15, MCXP15, MMG^{+}18, OLG^{+}15, PSRR14, PPST09, Pla08, PSCK^{+}15, RMP^{+}13a, RP19, RLZ15, RS11, RCA^{+}11, RHD^{+}16, RCR^{+}15, RK15, SNH15, SPMP11, SPZ^{+}10, SAD13, SK04, lSsCY17, SL18, SCS17b, Str11, SEF^{+}14, TVCB19, TZKH12, VDL^{+}15, VCW13, WJT^{+}14, WCZX16, WJYH16, WCLC13, XPBS11, XCHK14, XLHT17, Xu19, XZT^{+}11]. processing [Yos06, ZLLL11, ZWL^{+}13, ZO14, ZHGX16, ZWL^{+}19, ZDG^{+}14, dRC10]. processor [ABDP15, AFGL09, CLRB15, GSG06, KD07, LHC14, LL01, MGBC16, MCP^{+}12, NS19, Puf13, RRR04, RMW19, YL01]. Processors [AS19, SSV19, ZYH09, AAC^{+}15, ADMQO14, BHM^{+}12, BBdS^{+}17, BHKW12, CPEA18, CGST17, CLF^{+}19, CSWB11, DLZ16, DLK^{+}18, GCPS^{+}14, HFR^{+}17, JLT06, KBE07, KKW^{+}14, KL12b, KLDB10, LGLA15, LYL07, LLYL09, MSPDP19, RVD^{+}12, RR19, SR19a, SNK^{+}15, SPW09, SPQ^{+}17, TYTY15, TKS18, WJ09, ZYH12, ZZL^{+}17a]. product [ER12, HFR^{+}17, PLR^{+}14, SZT19, VFG11]. Production [BPS19, GED^{+}18, NTK08, PSL^{+}16, RLS^{+}09]. productive [GBFP09]. Productivity [MLS^{+}12, YBC^{+}07, KSM^{+}19, LLY19c, TFG^{+}12]. products [CHX^{+}19, HAJL16, Yan19a]. profile [KWK05, MSG10, OZI19, SL10, SKNH09]. profile-based [MSG10]. Profiles [MG09b]. Profiling [CSPM13, BM07, BAVM11, TYTY15]. profit [DPM17, LFPP17, Yan19a, ZS17]. Program [JWY^{+}05, BPdM06, CLZ^{+}17, CRV15, HM04, KL02, KB18, OKW18, SLM04, SLM05, TNIB17, TRH^{+}02, TBK^{+}15, YYS15, ZJL15]. program-to-program [BPdM06]. programmability [DP14]. programmable [CSWB11, FRKS12, NNH^{+}14]. programme [TWB13]. programmed [CZG16]. programmes [ADK^{+}17]. Programming [BH16, BL17, CLTT13, CGH^{+}06, MCP^{+}12, PA08, RWK^{+}02, SRdS09, SF10, UR04, VFAD17, ALVY05, BB02, BAVM11, CAD^{+}18, CLYC16, CNP^{+}15, CLRB15, DK09, DWC^{+}15, EK19, EBGS01, EB05, FJ05, FMS11, GL19, GA08, GvDHS12, HDX^{+}17, HvNJB15, HR06, JDH^{+}18, JZZL06, JLCA07, KOB01, KIM^{+}03, KSG11, Kes04, KHL^{+}17a, KS05, LL05, LCFkL05, LWB13, MLS^{+}12, MHH16, MKIO04, MTT15, MMSG17, MGS19, MSB17, MRH14, NO02, PRG15, PBF15, Pre01, RRR04, RGV09, SK04, SPBL06, TFG^{+}12, TMAG03, TBH^{+}18, WO14, YIN19, YWC11, YB12, ZDB^{+}14, ZWT^{+}18, ZDC^{+}09, vNMW^{+}05]. programs [ABF^{+}10, ADK^{+}16, ABS16, AUHWJ19, BHA15a, BB04, BV11, BK05, BL04, CL10, DAL15, Dut17, EFG^{+}03, EL01, EHSU07, FSPC^{+}02, FSG19, FLB^{+}05, GRS06, GM04, HL13, ITK09, KO06, LL16c, LZC^{+}02, LCC^{+}03, MTVF14, NA15, PAdS^{+}17, PAC^{+}17, PS07, RR15, RS07, SVS^{+}08, SSZ14, TLM17, TF03, VJHB05]. Progress [FS07, BKM^{+}07a, BKM^{+}07b, KKM^{+}06]. project [GKM^{+}08, GSWJ19, WNN^{+}15, ELM^{+}16]. projection [MZJ^{+}19]. projections [KAMB19, YSW19]. projects [KKM^{+}06]. Prolog [AR19]. Prometheus [ACG18]. promoting [CNP^{+}15]. proof [SZXG19, YL16, ZLH^{+}15]. proofs [WLZ^{+}18]. propagation [ACCM17, AMZ19, CWXW16, GdMK^{+}18, KB13, OFR^{+}17, SR19c, TLWZ14]. properties [ANH16, ABDO09, CEMR19, CSL12, HTHW16, IÁE11, KM13, LLY^{+}19a, MPHL03, ZQK15]. property [ALYD17]. proportional [ZLF19]. proposal [FMS11]. proposed [CG01]. Prospects [PCB^{+}18]. protect [BGdCCA11, ZBP07]. Protecting [LWY15, WYBS16, YKD^{+}15, ATKH^{+}17, CLF^{+}19, SW11]. protection [ALZR11, CJZZ10, LWYM16, LSZ19, LL19a, LLLyL16, RR01, ZXC^{+}19]. protein [BPL12, BDTdS13, MPR04, NCWD^{+}04, SL14, SRL^{+}14, TCP^{+}05, TTD^{+}05, YA04, ZCL^{+}18, SHH^{+}14]. proteins [FMS15]. proteomics [CV07, KBH^{+}15b]. Protocol [BNNH19, AKG13, AD15, BKA19, BF07, BDB^{+}13, DXWC16, DK19a, EN16, EA12, FIO15, GHMX13, GBJ19, HLC19, IHB15, KABD07, KG19, MTM19, MABP13, MRMC15, MMSN^{+}01, NLYZ12, SQS^{+}19, SWW^{+}16, TZ16, TAI^{+}11, WMC17, XBXS13, XJZ13, YZW^{+}15, YLWZ18, YYK^{+}19, YWM^{+}10, YL16, ZZC15, Zhe16, ZBZH11, BOB13, DT15b]. protocols [ALYD17, Aia15, BBB^{+}14, BHBD13, DVB14, DT15b, GD06, JLHH14, LWB13, NJ15, PGB03, SS15b, SC07b, XZJ11]. prototype [FGC06, JLHH14, ULS03]. prototyping [GBMM15]. protozoan [COdO^{+}11]. Provable [ZNT^{+}16]. Provably [LDZ^{+}14a, FIO15, KTZ^{+}18, SZXG19, XWXC14]. Provenance [AA19, BML08, KDG^{+}08, MWHW16, MLA^{+}08, OBD^{+}18, BA18, BA19, BD08, CFV^{+}08, CBBCD08, FMS08, FM08, GH08, HSBMR08, LPA^{+}08, MMW^{+}12, MGM^{+}08, RNAD19, SGSC08, SPG08, TC12, ZGST08, KSM^{+}08b, SKS^{+}08]. provide [BHA^{+}15b, JMF09]. provider [TPV17]. providers [ANK^{+}17, EdPG^{+}10, NB12]. Providing [GvHKK11, GMPT15, HSM14, PSM03, YESG^{+}17a, BHA^{+}15b, JKV^{+}15, SGD15, YESG^{+}17b]. provision [WSL15]. provisioning [EBMD13, HHKA14, KBB11, LDPZ14, SJB14, SD11a, SKJ17, ZFJ16]. ProvManager [MMW^{+}12]. proxy [CLH^{+}16, LFWS15, LHBW15, SKB^{+}17, XWXC14, YZCT17]. pruned [dFdOSR^{+}19, LCW^{+}17]. pruning [GKS^{+}07]. PSEFC [KT19c]. pseudo [RTPPH12]. pseudo-random [RTPPH12]. pseudonym [BB19]. pseudorandom [PPMH15]. pseudospectral [DGJ11]. PSkel [PRG15]. PSLS [KM03]. PSO [EB18, KHL^{+}17a, ZCH^{+}18]. PSO-based [EB18]. psychology [CZJY19]. PT [MGS19]. PU [PLZ14]. Public [LZC14, ATKH^{+}17, BZD16, CDdW17, GWVP^{+}14, HZY^{+}19, LL19a, LFWS15, LHB^{+}19, LMOT10, ZHM^{+}17]. Public-key [LZC14, BZD16]. public-resource [LMOT10]. public-safety [ZHM^{+}17]. Publication [HLB10]. publications [GWD15, PB19a, WDGK15, WDM14]. publish [BBPV05, MWPL15, MWPX17, TKK^{+}11]. publish/subscribe [BBPV05, MWPL15, MWPX17, TKK^{+}11]. publishing [HCG07, LXW17, WYAB07]. PUEA [EZJ^{+}18]. pull [DT17]. pump [GHLS19]. Purchase [ZYZ16, DJ19, HH19]. Pure [GVK12, ASP19, VDPC03]. pure-Java [VDPC03]. purpose [ABDP15, AdSCdR^{+}19, ETR^{+}13, LKPM09, PSRR14, RMP^{+}13a, SNK^{+}15, SW12, TXY^{+}16]. purpose-based [SW12]. push [DT17]. puzzle [CWC10, Ios11]. PVFS [KKL09]. pyGlobus [Jac02]. PySy [WO14]. Python [Jac02, WO14].

Q [SR19a]. Q-HypE [SR19a]. QC [GAB19]. Qespera [MV16]. QNX [KF01]. QoC [DD17]. QoC-based [DD17]. QoE [DD17]. QoS [BPB08, CL07, CLX^{+}12, DMRS15, DXM^{+}17, GYM14, GMPT15, HAAWA^{+}16, LLX^{+}15a, LDXC13, PRD^{+}13, QLD^{+}11, RC09, RCKV12, SAMS19, TZLC15, TKK^{+}11, WRLS12, WSW^{+}12, XWFH08, XZHW09, YBO10, YLR^{+}13, YSC^{+}17, YCWH07, ZS17]. QoS-aware [YCWH07, BPB08, GYM14, LDXC13, QLD^{+}11, WSW^{+}12, YSC^{+}17]. QoS-based [CL07, CLX^{+}12]. QoS-demanded [ZS17]. QR [BLKD08]. QSQL [RCXS09]. quadrature [GSB^{+}12]. Quakesim [PGP^{+}10]. qualitative [LLRS03]. qualities [ZQMC19]. Quality [MCCG11, TALT16, AAHRW04, CLF^{+}17, CM02, CRGR^{+}12, GCSB19, HLY18, HAvI13, KTM^{+}09, LDPZ14, LGJ17, LL19b, OORVB14, PSM03, PME^{+}08, RBO^{+}02, STO17, WSZ^{+}18, YJL12, ZL19]. quality-of-service [CRGR^{+}12, KTM^{+}09]. quality-of-service-based [RBO^{+}02]. quantification [BCF12]. Quantifying [ROQL18]. Quantitative [BKZ^{+}13, GYB^{+}11, ACMM06, HCC^{+}15, JZMD19, vAVS12]. quantity [CZ15a]. Quantum [HPHB^{+}15, TZ16, dARP17, Mos19, NDT^{+}16, SR19a, SR17]. quantum-dot [Mos19]. quantum-inspired [SR19a, SR17]. quasi [LOKW^{+}10, NN07, SJW18]. quasi-cache-aware [SJW18]. quasi-immutable [NN07]. quasi-static [LOKW^{+}10]. quasicyclic [LDZ^{+}15]. quasigroup [KPNS18]. quaternions [CH04]. queries [BLA^{+}14, DDM16, GBD16, LOSJ17, LFZ07, LC17, TSKM18, YWBS19, ZYZ^{+}12]. Query [CLW^{+}18, SPG08, ANH^{+}19, CZY^{+}18, DDP^{+}06, FBYO12, JQL^{+}15, KLP^{+}08, LTL^{+}17, LWY15, LW13, MJZ17, MRS03, PN19, PLJ18, TMP16, WGY^{+}19, XLYX11b, XZT^{+}11, ZLLL11, ZZL^{+}19, RCXS09]. querying [GR13]. question [CZWH07, HHWZ08]. questions [GR13]. queue [ESGQ^{+}11, GRQ19b, MV16, PTL^{+}16]. queueing [MLVBW12]. queues [WKL14]. Queuing [DZ13, YHH13]. Quick [RCXS09]. quicksort [MMO^{+}16, MMO^{+}16]. quiescence [MCG^{+}08].

R [Ano06, SB19a, KW19a, PRCV16, PSM^{+}11]. R&E [PCsHL18]. R-based [PRCV16]. R-Tree-based [SB19a]. R-Trie [KW19a]. RACAM [YYC10]. race [PS07]. races [DDF^{+}17]. RADAR [GCSB19]. radial [SPZ^{+}10]. radiation [CSB^{+}16, ZWW14]. radio [AD15, DK19a, EA12, FXX16, FO18, JKZ03, KJ19b, LCMY13, NLYZ12, SHST13, TZYL13, XBXS13, YCZ^{+}13, YZW^{+}15, YYLC19, LSY^{+}12]. radio-frequency [AD15, YZW^{+}15]. radio-over-fiber [KJ19b]. radiological [WBC^{+}17]. Radiotherapy [CRC^{+}15b]. railway [LZG^{+}19, ZCW^{+}18, Zhu19]. raising [AMRT14]. RAN [SHST13]. random [AR16, ANPR16, DFC12, HMPPT13, HCKF15, JHCH19, Li04, LMO15, MZJ^{+}19, RTPPH12, YYZH19, Zhu19]. random-walk-based [Li04]. Randomforest [LWP19]. Randomized [AKMZ13, ABDO09]. range [GBD16, LOSJ17, XZT^{+}11, ZLLL11]. rank [BDE^{+}19, FSM^{+}19, MWL18, PA18, YAG19]. Ranked [PPdSTB17, BV16]. ranking [Del08, WFS^{+}19, PPdSTB17]. Ransomware [LYS18]. Rapid [HLA^{+}18, KD19, WSRM12, GBMM15, LLT19b, NTK08, WWG^{+}11, KT19c]. RAPL [KD19]. rare [KHL17b]. RARM [KT19c]. Raspberry [DCA17]. RAST [WWG^{+}11]. rate [BN19, CH19, DCJ14, DA15, GHMX13, MST13, XWX^{+}19, ZXW19, ZJT^{+}19, ZCL^{+}19]. rating [SWZ12]. ratio [YZ10, ZYL10]. rational [WLWX14, WLWX16]. rationality [LC09]. RAVE [GAW09]. Ravenscar [KWK05]. raw [HH19, SdOVM16]. RAxML [SLM05]. RAxML-II [SLM05]. Ray [MGBC16, SBC15]. Ray-tracing [MGBC16]. Rayleigh [MS07]. RBAC [PZHS18]. RBF [SPZ^{+}10, YWLQ18, ZZZX19]. RCS2 [WWLD18]. rCUDA [RSC^{+}15, SIRP17]. RDF [GKP^{+}09, LZZ^{+}17, UMD^{+}13]. RDMA [ETR^{+}15]. RDMA-enabled [ETR^{+}15]. re [ANH16, CLH^{+}16, CZ15b, CZQ17, CXT^{+}18, DBR13, LFWS15, SKB^{+}17, XXX15, XWX^{+}19, YZCT17, PPdSTB17]. re-arrangement [DBR13]. re-encryption [CLH^{+}16, CZ15b, LFWS15, SKB^{+}17, XXX15, YZCT17]. re-identification [ANH16, CZQ17, CXT^{+}18]. re-partitioning [XWX^{+}19]. Re-Ranking [PPdSTB17]. Re-sort [PPdSTB17]. reachability [CL10, HWL18]. reaction [YYZH19]. reactive [QLS13]. readings [CS13]. ready [KW19b]. Real [AT01, ABK^{+}18, EN09, Fox17b, HTW14, JTD^{+}19, RK15, SSM04, Tur04, VK12, YJL12, ZTM12, AWR17, BVGVEA11, BLA^{+}14, Bri16, BMPP17, CHX^{+}19, CSB^{+}16, Cuz11, DvNM^{+}11b, EPA15, EAGVBVDS11, EABVGV14, FBH^{+}01, FRKS12, FRdOR^{+}19, FO18, FLB^{+}05, FAB^{+}07, Fox17a, GGS^{+}16, GKK09, GTL06, KOO12, KHM^{+}11a, Kal11, KGK17, KvGS^{+}14, KBB11, KSR14, KWK05, LS19a, LWB13, LTK17, LSL^{+}17, MGBC16, MSP^{+}13, MFF04, MSP^{+}19, MÖO17, Not16a, OSK^{+}01, PSM03, PSW11, Puf13, PRU14, RS16, RF15, RHT13, RVVPD^{+}17, SIOS02, SK18, SPS17, SZR16, TBTZ18, WYZ^{+}17, XLY^{+}16, XHD^{+}19, XXY^{+}16, ZTZ^{+}18, ZG04, BJC17, NDP^{+}05, SKD^{+}04]. Real-Time [Fox17b, HTW14, Tur04, VK12, AT01, ABK^{+}18, EN09, JTD^{+}19, RK15, SSM04, YJL12, ZTM12, AWR17, BVGVEA11, BLA^{+}14, Bri16, BMPP17, CHX^{+}19, Cuz11, EPA15, EAGVBVDS11, EABVGV14, FRKS12, FRdOR^{+}19, FO18, FAB^{+}07, Fox17a, GGS^{+}16, KOO12, KHM^{+}11a, Kal11, KGK17, KvGS^{+}14, KBB11, KSR14, KWK05, LS19a, LWB13, LTK17, MGBC16, MSP^{+}13, MFF04, MSP^{+}19, MÖO17, Not16a, OSK^{+}01, PSM03, PSW11, Puf13, PRU14, RS16, RF15, RHT13, SK18, SPS17, SZR16, TBTZ18, WYZ^{+}17, XLY^{+}16, XHD^{+}19, ZTZ^{+}18, ZG04, NDP^{+}05, SKD^{+}04, BJC17]. real-time-analysis [RVVPD^{+}17]. real-world [DvNM^{+}11b, FBH^{+}01, LSL^{+}17, SIOS02]. realistic [SAOKM04]. reality [Boc19, BDP18]. realization [TGB^{+}10]. Realizing [FRKS12]. reallocation [RBT18]. realm [XZJ11]. reaming [Bou13]. reasoned [YY19]. Reasoning [PMB15, BH09, LLH^{+}09, LC18, NNK^{+}07, NNvVdA09]. rebroadcast [KBDA19]. receive [Gog11]. receiver [AR16, WWLD18]. receivers [LWZ^{+}17]. recipe [WLDL08]. Recipes [DH15]. reckoning [CH04]. Reclaiming [ABDR13]. reclustering [HM12]. recognition [BCI^{+}18, CCLP19, Du18b, GQR16, JWW17, KS19a, WCZ^{+}18, WGQ^{+}18, ZSS18]. recommendation [AMBT17a, CDF^{+}17, HLF^{+}17, HCD^{+}02, KT19c, LH17, OZI19, RLZ15, TWW^{+}19, WZL^{+}17a, ZX09, ZTZ^{+}18, XLL^{+}18]. recommendations [PRS16, RBDI17]. recommended [Yan19b]. recommender [AMBT17b, FVRM15, SZT19, XWX^{+}17]. Reconfigurable [BKA19, CGB^{+}06, KHZ^{+}15, LGL16a, LGQ^{+}17]. reconfiguration [PZHS18, TZG^{+}19]. reconstructed [ZHW^{+}16]. Reconstructing [CPE^{+}19]. reconstruction [FMS08, KSM15, MJL01, SBC15, XYSW18, ZGZ19, XWPM19]. record [LH14]. recordings [CMT13, LML^{+}18]. records [DXWC16, SGL^{+}17]. recovery [BDB^{+}13, KCS07, MG09a, MB18, PGB03, XZH^{+}17, YLLZ09, ZXW16a]. Recurrence [CM05, JZMD19]. recurring [SP16]. recursive [DIK14, DFLL14, NSN^{+}17]. Recursively [YYCH10, YYC10]. Recycling [WGG^{+}07]. red [HW14, KC13, BUVS10]. red/black [KC13]. Redesigning [BBD10]. Redistribution [DT01, CW11b, HR18, RCKV12]. reduce [AJY^{+}15, CCC12a, DIK14, GGLD11, LLdA08, NPL19, MWL18]. Reduced [TDL^{+}18, GA09, LHC14, TJD^{+}17]. Reducing [CMMB13, DBH^{+}17, Dra17, KMRT18, OKW18, VL17, BBK11, ESGQ^{+}11, NSSAK13, NSSAK16]. reduction [BCM15, BN19, BXLJ16, CBQ^{+}11, DÇK^{+}18, LLX^{+}15a, LML^{+}18, RMCA12, Roj19, WML^{+}19, XZHW09, ZBZH11, dARP17]. reduction-based [LLX^{+}15a]. reductions [BTG06, GPZ04]. redundance [TLX^{+}17]. redundancy [ASE^{+}17, FED03, MSN^{+}19, PWMX17, SC07a, XLYX11a]. redundant [ZWH^{+}17]. Reed [CSWB11, KCS07]. Reengineering [MMS07]. Refactoring [CM06, WZZL13]. reference [ABtGT^{+}12, AP06, HWR03, JAU19, vdABST10]. referencing [PAM^{+}15]. refined [KS19a, KGP^{+}19]. refinement [HvNJB15, LB11, VLJ17]. refining [IS10]. reflection [Arz17, SN18]. Reflections [SvDO15]. reflective [CBP^{+}04, HGB^{+}08]. regarding [KS19c, SWU08]. region [ATI14, GR14, MZS^{+}10, PB19b, ZJT^{+}19]. regional [cDrLyC^{+}19, PNL10]. Register [WLL14, CCC12a, LHC14, LYL07, LLYL09, Mit17b]. registration [QSZL18]. registry [FMM08]. regression [BPS19, BBA18, HAA^{+}17, PSRR14, ZLH^{+}18]. regression-based [PSRR14]. regridding [LB11]. regulating [SCS17a]. regulation [YLLC18]. regulatory [KHM^{+}11b]. reinforced [KCBO17]. reinforcement [BHD13, CCCW13]. reinsurance [BRCV16]. related [CZ19, LLL16, MHK^{+}18, MZA19, SdOVM16]. relational [cDrLyC^{+}19, SC07b, ZSWS18]. relations [XLMH14]. Relationship [ZYL10, XZZ^{+}16a]. relationships [AFGL09, HmLGP03]. Relative [SAC^{+}07, BLL^{+}19]. Relativistic [HW14]. relay [DZ13, SCLK15, TKHA13, WRLS12, ZPG10, ZJL13, ZZY^{+}19]. relay-based [ZJL13]. release [JW10]. Relevance [GOLL17, WFS^{+}19]. relevance-based [WFS^{+}19]. Reliability [Dab09a, SZMF19, BDZ19, CAG^{+}13, FPHZ20, HML19, MST13, XSMZ16]. reliability-based [BDZ19]. reliable [DA15, HKA^{+}15, KKL06, MSMA19, PYKL16, SS18, SM19a, VO15, XZH^{+}17, YLY04]. Relieve [LWSZ19]. Remote [HMFK15, NMMS01, ASWR12, AHM06, DXWC16, HWR03, MWL^{+}13, MWL^{+}15, MKO^{+}17b, PJW^{+}14, PRS01, RSC^{+}15, SHST13, SIRP17, SS19c, TAB^{+}06, WYAB07, XLHT17, YHJ^{+}14, ZWL^{+}13, ZYLT06, BVGVEAFG11]. remote-sensing [ZYLT06]. remotely [Pla08, SPMP11]. Removal [SS19b, ECP18, RC09, ZWW14, ZZZ^{+}15]. removing [LFG05, XLYX11a]. Rendering [SZG^{+}19, ASWR12, WJ09, ZZC^{+}17]. Rendezvous [Kri13]. Rendezvous-based [Kri13]. renewable [KTB17]. rental [GCZ^{+}17]. reordering [CBB^{+}19, GKK09, MUKY18, SAB15]. repackaged [AMSS15].

Repairing [PWMX16, AS17, XGC19]. Reparallelization [KBG^{+}09]. RePast [MT08]. repeated [LOSJ17]. replacement [BBC16, DLT^{+}16, SS17b]. replica [PSC^{+}17, VSK17, WSW^{+}12, YYC10, ZFXJ19]. replicas [BF07]. replicated [AaBT17, MKB01, TMS^{+}12, YCSY19, ZH08]. Replicating [Kal11]. replication [ASS08, BPdM06, CLNR18, DFLNP07, EN19, ESZ09, GMS09, GÖ18, MY17, NCD^{+}08, RBT18, SMFM18, XPS^{+}15, ZNT^{+}16]. reply [OKW15, OK15]. Report [MKO^{+}17a, WWG^{+}11]. Reporting [LRS15]. repositories [BH09]. repository [BM10, BPL^{+}19, FHO^{+}15, GKP^{+}19]. representation [CXT^{+}18, mLGP03, LFH^{+}08b, SLC^{+}18, SLC^{+}19, XYSW18]. representations [KSN16]. representing [BSZ09]. reproducibility [MWHW16]. reproducible [YLH^{+}19]. reproduction [ZHJ19]. reprogramming [XBW^{+}15]. Reputation [AMRW06, AAQAR^{+}17, CHZ10, MK15a, CZWH07, CLX^{+}12, XLL^{+}12, ZQLZ12]. Reputation-based [AMRW06, AAQAR^{+}17, MK15a]. reputations [SZA08]. request [BMV03]. Requests [CKSC10, LL10, RSR06]. require [KO06]. Requirements [FGG^{+}18, KBH^{+}15b, Can06, FPC15, MÍM19, MG09b, Sod07, SE01, VL17]. rerandomization [CXW17]. resampling [ZF14]. ResAna [KvGS^{+}14]. rescheduling [NB12]. Research [Boc19, CDdW17, CH19, CZJY19, cDrLyC^{+}19, Du18a, GZG^{+}16, HDX^{+}17, LCW^{+}19, LSZ19, LTZ^{+}19, LHZX19, LLY19c, Qia19, lSsCY17, WYW^{+}17, WF18, XS19, YL19, YG19, ZSS18, ZCL^{+}19, ACMA07, BPL^{+}19, DN19, DM15, EMB11, Fer15, LZWD^{+}15, LPW15, MKX^{+}15, PB19a, QLZX19, SBB^{+}15, WNN^{+}15, YTF^{+}01, HGT14, SHG^{+}07]. researchers [MTHK14]. Reservation [GCZ^{+}17, DFPT06, VDB09, VO15]. reservations [ET09, RSR06]. reservoir [KCZ^{+}05, LAC^{+}08, MBP^{+}05, PML^{+}05]. resident [WCH^{+}07, YTD17]. residential [MZG19]. residue [KPNS18]. resilience [XPWF15]. resiliency [SBS19]. Resilient [BDL^{+}15, SNGR18, ASE^{+}17, EPA15, SWS^{+}18]. resistance [FIO15, SZD19, ZQLZ12]. resisting [CXW17, ZZL^{+}19]. resolution [BDY03, EN16, HYL^{+}19, OLG^{+}15, WYQ^{+}13, XYSW18, ZABP18, ZLL19]. resolutions [JC07]. resolve [KS19a]. resonance [EMEY14, KSM15]. resonator [DYY^{+}19]. Resource [AC02, ACC^{+}07, CEM^{+}08, FBC10, LLF08, LQL^{+}15, Men03, NNvVdA09, RSR06, SJB14, TAB^{+}06, TCH^{+}13, YLC11, BHD13, BKM^{+}07b, BAC^{+}15, BDP^{+}14, BAGS02, BM02, CLQ^{+}17, CA06, CZ11, CLL^{+}19, CBB^{+}19, DFPT06, DS07, DvNM^{+}11a, EdPG^{+}10, ET09, EBMD13, FCY17, FXX16, GCSB19, God12, GVK12, GMVRGS15, GS04a, GAW09, HSM14, HHKA14, KC15, KvGS^{+}14, KF18, KSR14, KTB04, LFPP17, LVN^{+}12, Ley06, LC09, LLL15, LLT^{+}19a, LWB13, LAM^{+}09, LMOT10, MLS^{+}15, MRS^{+}10, NB12, PYKL16, PPC^{+}15, PGW06, PRP^{+}15, QLC04, RCB^{+}04, RBNG15, RSPV17, SMH^{+}19, SLV12, SN18, SPJ14, SGV12, SD11a, SdSL18, Soo16, SB17, TAMC19, TXZ^{+}17, TCDMR^{+}17, TK10, VDB09, VGN^{+}16, WZLQ16, WP12, WL11b, XLZD13, YPLJ11, Zen19a, ZJL13, ZM13, ZZY^{+}15, ZFJ16, ZLA^{+}15, dRC10, vdKEL10]. resource-aware [GAW09, SGV12]. resource-constrained [ZLA^{+}15]. resource-constraint [SMH^{+}19]. resource-efficient [LLL15]. Resources [WD07, BDI^{+}07, BFVRC15, CR12, CLH^{+}08, FHO^{+}15, GGFPGB14, GD06, GKP^{+}09, HKG08, Jun16, KBT^{+}14, KFS^{+}06, LBV16, MBMB18, NCWD^{+}04, OBD^{+}18, PBK19, SWU08, SWD^{+}15, SO16, VAC^{+}07, Wit10, XCL09, ZMZD11, ZBP07, ZDL07]. response [KMRT18, LWW06, MSST15, YZ10, ZYL10, Zhu19]. RESTful [ET15, CS15]. restricted [CLH^{+}16]. restriction [TXZ^{+}17]. restructuring [CLL^{+}18]. results [BG04, BCM^{+}05, CML^{+}10, GRS06, LLRS03, MKO^{+}17a, SLM^{+}10, VDL^{+}15, YXLZ16]. retargetable [PBSB04]. retinal [ZBZ^{+}15]. Retracted [FPHZ19, FPHZ20, HML20, Pan20, ZLW^{+}19]. Retraction [Ano12]. retransmission [KCS07]. retrieval [CHH18, CMT13, DXWC16, DNB19, HKA19a, LS15, MLRR09, PLJ18, PPP10, PPR19, SIST18, TSBR10, UAW09, WXLM19, XGXH15, YFL18, Zha19, TKS18]. returns [DFC12]. reuse [CXW17, LVN^{+}12, PHCR09, RSN^{+}19, WGG^{+}07, YXLZ16]. revealed [BDY03]. revenue [MRS^{+}10]. reverse [ACC^{+}12, ATI14, RGCC15, TQL^{+}14, RC09]. reversed [HZH^{+}19]. reversible [SS19a, SM19a]. review [ABS16, FVRM15, GRQ19b, IHA^{+}15, LGdVH13, MG09a, SN18, ZHC^{+}18]. revisited [BCK^{+}09]. Revisiting [DVB14]. Revocable [MML16, FLL^{+}14, WLFX17]. Revocation [WJH06, sTzNL16, ZWY^{+}19]. rewriting [GGC19]. rewriting-based [GGC19]. RF [BT04, SM19c]. RF-MVTC [BT04]. RFID [FLL^{+}14, SZXG19, YL16]. Riccati [BDE^{+}19, MQOQOH01, PIAH12]. rice [SZD19, MCAB^{+}02]. rich [LPW15]. RICS [TJD^{+}17]. RICS-DFA [TJD^{+}17]. right [LLH^{+}17]. right-hand [LLH^{+}17]. Rights [HCBRM16, GLL16]. rigorous [RTPPH12]. Rim [PC17b, IUCH^{+}17]. RIN [BB19]. ring [ZWW14, ZZZ^{+}15]. RIoTBench [SCS17b]. RIP [KKK10]. RIPPER [XDP18]. rise [YC19b]. Risk [FRU12, SR19c, BT04, CLL14, DPGA11, WSRM12, DCK12]. risk-aware [DPGA11]. risk-free [BT04]. Ritrovato [Ano06]. RMA [BBW19]. RMI [SS19c, WCCL05]. RNA [KHM^{+}11b, LS05]. RNN [JQL^{+}15]. road [KHHC13, RNJM17, SWLJ17, XZT^{+}11]. Roadrunner [WJLD09, GKS09]. roadside [YBZ^{+}15]. roadway [LMGZ19]. robin [GDJ16, LJML10]. robot [BAD^{+}11, DCCZ18, SHT^{+}17, ZZ18]. robotic [JAU19, ZJS11]. robotics [BCI^{+}18, CRGR^{+}12]. robots [SJ19]. Robust [AG17a, CHM15, CGI17, KBH15a, LPY^{+}08, SB17, CCCC06, ECP18, HZC^{+}14, LXP^{+}12, MSK19, NTK08, Sha15, SGV12, WRC09, XLL^{+}12, Yan19b, vdKEL10]. robustness [BCdlCT06, LNKZ08]. rock [CNAQ18, LLY^{+}19b]. Rocks [PKB03]. rod [YZ19]. rogue [AYSZ14]. Role [KM13, LLG^{+}15, DCFC08, HTR10, HPS05, MLL^{+}11, Mal05, SACJ04]. role-based [MLL^{+}11, SACJ04]. role-driven [DCFC08]. roles [PZHS18]. rollback [ALYD17, ASP19, MG09a, YLLZ09]. rollback-dependency [ALYD17]. rollback-recovery [MG09a]. Roofline [KAMB19]. roofline-based [KAMB19]. ROP [TJZ^{+}19]. rotating [DdB01, MO02a]. rotor [YYZH19]. rough [BXLJ16, YY19]. Round [LJML10, GDJ16, SZXG19, TQL^{+}14]. route [KKK10, KT19c]. router [LWL^{+}19]. routers [ZLG^{+}19]. routines [Cho01, CGGH17, LSK04, VLMPS^{+}18]. routing [BKA19, BVIB19, DK19a, DT15b, DA15, FXX16, GBJ19, IHB15, JWZ13, KA16, KABD07, LWF^{+}15, LLC^{+}15a, MTM19, MSMA19, MBB19, NQL^{+}17, PGL^{+}17, PS19a, RGCC15, SWS^{+}18, SQS^{+}19, TZG^{+}19, WZS^{+}15, YESG^{+}17a, YESG^{+}19, YWM^{+}10, ZJKL10, ZLC17b, YESG^{+}17b]. rover [SJ19]. RPC [BDG^{+}10]. RPS [MSMA19]. RPS-LEACH [MSMA19]. RSA [PY19, sCR19]. RSEL [FLL^{+}14]. RSS [YYLC19]. RSS-based [YYLC19]. RT [FRdOR^{+}19]. RT-1.1 [SKD^{+}04]. RT-JADE [FRdOR^{+}19]. RTS [SWLJ17]. RTSJ [HTI05, KW11]. Rule [KT19c, YTL19, DBH^{+}17, FSM^{+}19, LLH^{+}09, LZZ^{+}15]. rule-based [FSM^{+}19]. rule-sets [LZZ^{+}15]. rules [BTCB16]. rumor [LMCL19]. Run [LH05, VHBB03, AA19, LCH^{+}06]. Run-time [LH05, VHBB03, AA19, LCH^{+}06]. Running [ZQD^{+}17, Bao19, CML^{+}10, MÖO17, PSS^{+}18, RMP^{+}13a, SVN12]. Runtime [CEG^{+}05, LSJ16, MST13, ADI^{+}14, BEDK18, BKH08, CSC^{+}17, CCCC06, DT15a, GFBR10, GA08, JK13, LLdA08, MFGE19, RMCA12, STL^{+}15, SWD^{+}17, TMAG03, WRC09, WMvP^{+}09]. Runtime-aware [LSJ16]. runtimes [JFI^{+}08, MD19]. rupture [LOKW^{+}10]. rush [LWSZ19]. RV [MP05]. RW [BDH16]. RW-safe [BDH16]. RWS [LPA^{+}08].

S [GKP^{+}09, GKS09]. S3BD [WS19a]. SaaS [AMBT17b, FHH15, SK18]. SaaS-based [FHH15]. SABR [TZKH12]. SaC [SSB^{+}14, ZLN^{+}13]. SaC-FRAPP [ZLN^{+}13]. SaC/C [SSB^{+}14]. Safe [vRGNP09, BDH16, WL11b]. Safety [RS12, SDH^{+}17, WCWB19, CG01, LZG^{+}19, ZHM^{+}17, HL13, LWC17, WK12]. Safety-critical [RS12, SDH^{+}17, LWC17]. SAGA [SKNH09]. SAHARA [May18]. Sakai [SHG^{+}07]. sale [WFS^{+}19]. sale-ranking [WFS^{+}19]. saline [WYY^{+}19]. saline-alkali [WYY^{+}19]. salt [SS19b, SZD19]. salt-tolerant [SZD19]. sample [JyLdZ^{+}18, YWL^{+}17a]. sampled [PLR^{+}14]. sampling [CSBL12, dOCPFJ13, FSM^{+}19, WJXZ18]. SAMRAI [HK02]. Sandboxing [WJKS18]. sandwich [FP09]. Sandy [VDL^{+}15]. SAP [AK01]. Sapphire [HM03]. satellite [SMBT07, Zen19a]. satisfaction [GSWJ19, JZL15]. satisfaction-aware [JZL15]. satisfiability [CW07]. saving [CPEA18, MFG^{+}13, ZQD^{+}17]. SAVM [ZYFZ19]. SC-OCR [SS17b]. SCADA [ALZR11, LWS19]. Scalability [CSMK19, DGJ11, KRS11, LL01, AYN^{+}14, BHD13, BCA^{+}10, BZB17, CCS14, FE18, GB07, HKB07, HKAC14, LKYS04, LSK04, RVRD10, RLRG15, SGCA^{+}16, VRMB13, VDdN^{+}07, YLY04]. Scalable [AD15, BMA03, GEJ^{+}08, GM10, GKPT13, JQL^{+}15, JR19, KKT13, LB11, MB14, PY19, SCRV11, SMS^{+}19, UMD^{+}13, ZW17, ATKH^{+}17, BB02, BLL18, BWD15, BKH08, BDF15, BKLH09, BWHS18, BMPP17, CFPJ^{+}17, CTAB16, DLM13, FBV^{+}17, GWHJL19, HRR^{+}11, Kar14b, KHM^{+}11b, KMJ14, KSB^{+}19, KSC12, LXRJ13, MWPL15, MTT15, MCG18, MWW10, NMM^{+}10, PSG03, PSLC11, QLL10, RTET15, RZL^{+}19, SRF13, SGCG09, SD15, SS07, TMP16, TAI^{+}11, WYZ12, WLFX17, YLH^{+}19, ZLN^{+}13, ZlC15, ZGS17, dMd^{+}17, QH10]. scalar [CSTV06, FNBS16, HFR^{+}17]. Scalarm [BRK^{+}17]. Scalasca [GWW^{+}10]. Scale [KSM^{+}19, PDD14, RIP18, AHP^{+}13, AML^{+}15, BH09, BFL^{+}10, BCM^{+}07, BBB^{+}14, CLL^{+}18, CHM15, CBQ^{+}11, CGN15, CWM18, CPS^{+}14, CDH^{+}15, DVD^{+}12, DLX^{+}16, DP19, DZJ^{+}15, DZM^{+}15, EN16, EBGS01, ERZ^{+}11, EJD17, FAPC16, HFDJ10, HTR10, HWQ^{+}16, HLF^{+}17, HWL18, HSHT14, JAA08, JCK^{+}13, JWY^{+}17, JPWH02, KBT^{+}14, KCZ^{+}05, KSB^{+}19, LW05, LBV16, LXRJ13, LXW^{+}16, MZJ^{+}19, MvWvM^{+}17, MSCS18, MCY^{+}10, MB14, MJD15, MJD17, Mos19, Not16a, PTL^{+}16, PAM^{+}15, QZYZ16, QLS13, RLRG15, SNH15, SK09, SLV12, SMM18, SCBH09, SGCG09, TJ17b, TJ17a, TRH^{+}02, WJ12, WYZ^{+}17, WZXZ12, WSWL12, XBXS13, YLEB14, YMLR16, YLL^{+}18, Zha19, ZYZ06, ZHGX16, dCRS11]. scale-free [CLL^{+}18, JWY^{+}17]. SCALEA [TF03]. scaled [ML19b]. scaled-down [ML19b]. scales [WQS^{+}16]. Scaling [Hey19, SPH13, DMMA17, HWZ^{+}15, KS19c, PDY14, RPK08, SLD^{+}12, XLQL18, ZLF19, AMAB17]. Scaling-aware [AMAB17]. scan [ABFL17]. scatter [MTK16]. scattering [GSB^{+}12]. scavenging [KD10]. SCC [CLRB15]. SCCs [LSL^{+}17]. scenario [WST^{+}17]. scenarios [FGG^{+}18, LWW^{+}19, LSMVML15, VSR^{+}09]. scene [Wan18b]. Schedule [XXLL17, ABDR13, CNP^{+}15]. scheduled [HLYD12, SAB15]. scheduler [BM08, CLQ^{+}17, EME19, KMRT18, KAM11, ME08, PGC^{+}19, PK17, SO16, ZJL15]. schedulers [ADI^{+}14, KKWZ15, LL16b, NvV09, RO12a, RO12b, ZF14]. schedules [KBE07, LLH^{+}18, RRR15]. Scheduling [AAF17, AS17, BKSM^{+}15, DJM12, EJF^{+}16, GRS06, IQOvdG13, KLDB10, LL10, SRdS09, SF10, XLYL17, AJY^{+}15, ABC^{+}08b, ALMT19, Ang08, AMS17, ATNW11, BFM^{+}06, BKND16, BAGS02, BM02, CSC^{+}17, CHP17, CCC12a, CLT^{+}16, CZY^{+}18, CPXA06, CL07, CCTW11, DSO^{+}01, DXM^{+}17, Dra17, DKJ13, DRF07, EB18, ESZ09, EQW^{+}18, EABVGV14, EFA^{+}17, FRdOR^{+}19, GDJ16, GRQ19a, GSG06, GQ04, GMVRGS15, GA09, GLS^{+}19, GAS19, HKS19, HZHP09, HLG17, Hun15, IHA^{+}15, JZL14, JZL15, KS19b, KV12, KBB17, KB17, KW11, KQR^{+}17, KSPM12, KO06, KKV13, KR11, LF15, LHL10, LLKC08, LHC14, LWFL14, LGY17, LGL16b, LHT^{+}09, LCYJ08, LQL^{+}09, LJML10, LQL^{+}15, LYC16, LBY^{+}16, LSJ16, LZBF17, MS17b, MS18, MSP^{+}13, MDB^{+}17, MRS03, MK15b, NSBR07, NC05, ON02, PRT09, PCGE18, PCF^{+}17, PRV11, PV15, QLS13, RHRB13, RF15, RR19, RHuR^{+}19, RHZ^{+}17]. scheduling [RCA^{+}12, RB17, SRS16, SV09, SR17, SWP17, Sod05, SK18, TKB09, TZYL13, TYTY15, TLF17, TY15, TV14, VBW06, VP19, Viv03, WBD^{+}19, WGZL06, WRC09, WL11a, WZZL13, WQS^{+}16, XLT^{+}17, XLHT17, XWH^{+}17, YWC11, ZEB10, ZWL^{+}15, ZS17, ZWT^{+}18, ZCZ^{+}19, ZL12, ZH15, ZQW^{+}17, ZCH^{+}18, ZXXN06, dAAVS12]. schema [CT11b, SE01]. schema-mapping [CT11b]. scheme [AQRA^{+}18, AR16, ALL^{+}15, BJD^{+}19, BC16, BOB13, BZD16, BBB16, CC13, CCW06, CZY^{+}18, CDP17, DBR13, DXHL17, DA15, FLL^{+}14, ISO^{+}14, JNUH17, KS19c, KMA04, KTZ^{+}18, KDW^{+}17, LDZ^{+}14a, LWYM16, LFWS15, LJML10, LZC14, MMS17, NR17, OFR^{+}17, PWMX17, PY19, SCS17a, STO17, SWLJ17, SB19b, TZYL13, TC17, TKS18, sTzNL16, WYQ^{+}13, WZC16, WLFX17, WXSH19, WWC^{+}19, WZXZ12, XCL09, XHH12, XBZ10, XWXC14, XXX15, XXCY19, YSQM19, YKA^{+}19, Yos06, ZEB10, ZTZ^{+}18, ZWY^{+}19, ZGX11, vdKEL10]. Schemes [WS09, CPXA06, ESGQ^{+}11, zGWXT09, GCZ^{+}17, HKA19b, LX08, LWW^{+}19, LLLyL16, LWL^{+}19, ZDX12]. scheming [NQL^{+}17]. Scholarship [LVN^{+}12]. Scholes [BHPS14]. schooling [LKPM09]. schools [GKM^{+}08]. Schur [GKK09]. Science [BSC^{+}15, CGK^{+}07, DMM^{+}07, GWD15, GBB^{+}15, HF05, LZWD^{+}15, MTA^{+}07, Sod07, WAS07, WC08, WD07, Wu18, Xu08, ZYH09, ACF^{+}07, ACFT15, BKWM19, CBHTE11, CAD^{+}18, Che18, CDH^{+}15, DGA^{+}10, FKP^{+}02, GBMM15, HMFK15, HAvI13, JCK^{+}13, JKL^{+}17, KA11, LMH^{+}14, LWL15, LFH^{+}08b, LGD15, MCC^{+}15, MCD^{+}15, MWL^{+}15, MTVF14, MKX^{+}15, NAP^{+}07, Nak02, OTG^{+}07, PGP^{+}10, PMG^{+}15, PC17b, PCC17, RTPPH12, SvDO15, SRAG16, SLM^{+}10, VSB^{+}15, WHW10, WBB^{+}07, WBD^{+}03, YR15, YLEB14, ZSL^{+}10, ZWF^{+}06, ZYH12, vHKT^{+}11, BD08, CCK^{+}17, FGP^{+}11, GTGT11, HF17, PME^{+}08, RLS^{+}09, SPR^{+}07, SM11, SBP12, SGV12, VBW06, WHW10, WDGK15, YDB^{+}13]. Sciences [Qiu11, QFG14, QFT14, ACC^{+}15, GvHKK11, OGA^{+}06, Sod07]. Scientific [Ber07, LAB^{+}06, AFG^{+}05, AKM^{+}06, BBG17, BML08, BYT^{+}12, BSB^{+}03, CSMB15, CGH^{+}06, DRS^{+}13, DHH^{+}13, DT17, DCFC08, DK19b, GHB^{+}06, HJV^{+}19, HZHP09, HCD^{+}02, JPWH02, LSE^{+}13, LMH^{+}14, Lan17, LL05, LPH09, LS19a, LGL16b, LTKF11, LNCY11, LHLH16, LZBF17, LFH^{+}08b, MMMP01, MOF15, MMW^{+}12, MYDM06, MCD^{+}15, MRJ^{+}14, MM10, NAK^{+}15, ODS^{+}13, OCC^{+}05, PB19a, Par02, PFC^{+}09, PGO^{+}04, QLD^{+}11, RSSM06, RCXS09, RC09, RB17, RRWS08, SM02, SAB15, SM09, SD11a, SKA^{+}14, SGG07, TMF^{+}10, TMP16, TCBR^{+}10, TCBR11, TC12, WRC09, YWBS19, YK10, YYL^{+}12, ZP06, ZWL^{+}15, ZJS^{+}17, ZCZ^{+}19, ZDL07, dOOO^{+}12, vRKS03]. SciScope [BvIF10]. SCM [PS19a]. scope [BDB^{+}13]. scratch [YWY^{+}10]. scratch-pad [YWY^{+}10]. screening [GCPS^{+}14, JvAB^{+}15, KBT^{+}14]. script [MTT15]. script-based [MTT15]. Scripting [BYT^{+}12, Nob08]. SCRRM [DA15]. SCTP [DLPV07]. ScyFlow [MYDM06]. SDIVIP [YNX^{+}16]. SDK [CG01]. SDN [Bai17, CKL19, HDX^{+}17, IUCH^{+}17, PCsHL18, TAMC19, YC19a, ZW17]. SDN-based [TAMC19, YC19a]. SDN-IP-based [PCsHL18]. seals [ZDHJ18]. seamless [cWsThY19]. search [And13, AMVOSGAC17, BKS18, BMS^{+}09, BZD16, CMW02, CLH^{+}16, CGR19, DXG13, DAC12, DKMM14, DOJ^{+}19, EN19, FMS15, GS08, GKS^{+}07, HKA19b, IPGCMW18, KHF^{+}17, LS19b, MPR04, PWS19, PPST09, RIFR10, SRM13b, SER15, SPLO06, SR17, WMC17, WJP14, WZ16, WS19a, XLYX11b, XZT^{+}11, YPLJ11, ZK08, ZBZH11, ZHZ^{+}13, ZCL^{+}18, ZHX^{+}19, ZCS06, sCR19]. search-space [GKS^{+}07]. Searchable [HKA19b, MML16, YZCT17]. searches [LLB04, RM03]. searching [SL14]. SEC [GSK19]. SEC-H5 [GSK19]. SecNRCC [XBW^{+}15]. second [LGP19, Ang07, CL08, CR13]. secondary [LS05, ZGL19]. secrecy [ATKH^{+}17]. secret [CLZ^{+}17, TQL^{+}14, XW13]. Section [STS19, ZQH12, RBP12]. sector [PCsHL18]. Secure [ALZR11, DXHL17, FXX16, GQH17, GSK19, HZY^{+}19, KKDS19, LMO15, QZDJ16, SCS17a, SB19b, WS19a, ZMZD11, ZZ14, ZZ15, ASE^{+}17, AYSZ14, BOB13, BZD16, CSL^{+}18, CLMM12, CKL19, DK09, DZL^{+}17b, FLL^{+}14, FIO15, GKM19, HKA19a, HWQ^{+}16, JSR^{+}19a, KD10, KTZ^{+}18, KDW^{+}17, LLLJ14, LDZ^{+}14a, LCC^{+}18, LBY^{+}16, MWJ^{+}10, MG09b, NR17, PWS19, PY19, RSKK19, SWW^{+}16, SWZ^{+}18, SC19, SZXG19, WLW11, WLWX14, WLWX16, WZL^{+}17b, XCHK14, XWXC14, XBW^{+}15, XXX15, YZCT17, YWM^{+}10, YL16, ZYFZ19, ZNT^{+}16, ZQD16, ZGX11, ZZL^{+}19, KKK^{+}19]. Secure-CamFlow [KKK^{+}19]. Secured [SM19b, CK13, MBB19]. Securing [LNG^{+}16, VT15, XLWZ11, LNBL17]. Security [AKK^{+}07, BM04, Boe12, HJTX17, KV12, Kin04, LB19, NPM19, SK08, WZC16, WCWB19, XZ09, XAK16, XSMZ16, XBK17, XBM14, YWT^{+}12, Zha08, AI17, AAHA18, AMS17, BXQ17, CGOF15, CJC^{+}18, COC18, CGI17, DXWD16, FPC15, IZXM09, KKK10, KKK^{+}19, KJŚ^{+}15, KWXY18, LSZ19, LZC14, LLSL15, MCWL06, MKX^{+}15, Nam19, NLYZ12, Nis18, OKO18, OK18, OEP^{+}15, PHY^{+}18, PVCS18, PMB15, PMG19, SGJ^{+}17, SVY19, SW09, SWHL16, TWZ^{+}19, WTEG17, WAD12, WR17, WT18, XHZ12, XADLC15, XBCW19, YZXW17, YMLR16, YJZZ19, Zen19b, ZYY^{+}19, HYQ17]. Security-aware [KV12]. security-level [KJŚ^{+}15]. SEED [JZL14]. seeking [HAN19]. segment [FJZ^{+}14]. segment-based [FJZ^{+}14]. segmentation [ALVY05, BÇG14, EMEY14, LYL^{+}19, LLZ^{+}17b, WJ12, YHJ^{+}14]. Segregation [Ang08]. Seidel [LL19c]. Seidel-based [LL19c]. Seine [ZP06]. Seismic [JW10, ACC^{+}12, PWJ10, RSTV05]. seismogenic [MZS^{+}10]. select [SS19a]. Selected [WC08, Xu08, GZX17, PDD14, XYS17, YWA07, WAS07]. selecting [EAGVBVDS11, MMB^{+}17, PTL^{+}16]. Selection [DLT^{+}16, HJTX17, PB07b, PK08, BV16, BKND16, BFVRC15, CDA09, CWC10, CEB^{+}18, DZL^{+}19, DK19a, GYM14, GLMT15, GMVRGS15, HAJL16, HAA^{+}17, KOOB15, KC18, KTM^{+}09, LFH08a, MSMA19, MABP13, MBMB18, MBC^{+}14, MSM^{+}14, NNvVdA09, PWH18, RR19, SAMS19, SK18, TPV17, XY17, YYC10, YLD13, ZK08, ZLY^{+}13, ZZY^{+}19, ZLW^{+}19, ZFXJ19, ZWL^{+}17]. selective [FSM^{+}19, Jon09, LZC14, WFJ^{+}17, ZLZ^{+}19]. Self [CLL^{+}19, DCCZ18, GCSB19, HHKA14, MZ06, MO15, VD05, WLZ11, XLT^{+}17, XWXC14, AA19, AM15, BFVRC15, Bri16, CEMR19, CSL12, DHV03, DK19b, DYY^{+}19, FMS11, HM16, KF15, KOO12, KYM17, LAM^{+}09, LJML10, ML19a, NSBR07, PB12, RPK08, RVRD10, RS16, RIFR10, SR19b, VH12, WFHT17, XDE^{+}04, XJZ13, YDL09, YWC11, YYZ^{+}17, ZWMT12]. self-adaptable [ML19a]. Self-adaptation [XLT^{+}17, WFHT17]. Self-adaptive [HHKA14, BFVRC15, CEMR19, CLL^{+}19]. Self-balancing [DCCZ18]. self-building [Bri16]. self-caring [KF15]. Self-certified [XWXC14]. self-compensation [DYY^{+}19]. self-configurable [YDL09]. Self-configuring [GCSB19]. self-descripting [Bri16]. self-diagnosis [XJZ13]. Self-healing [MO15, AA19, FMS11, GCSB19]. self-interested [DK19b]. Self-learning [CLL^{+}19]. Self-maintained [MZ06]. self-management [AM15]. self-managing [VH12]. self-optimization [CSL12, YYZ^{+}17]. self-organized [KOO12, KYM17, LAM^{+}09]. self-organizing [HM16, PB12, RIFR10, XDE^{+}04, ZWMT12]. self-powered [SR19b]. self-scaling [RPK08]. self-scheduling [NSBR07, YWC11]. Self-similar [WLZ11, LJML10]. self-similarity [RVRD10]. self-stabilization [CSL12]. self-stabilizing [DHV03]. self-x [RS16]. selfish [LYF^{+}17]. selfishness [ZQLZ12]. Semantic [FHH15, FGP^{+}11, GKP^{+}19, HHWZ08, LLJ18a, LLJ18b, LLZ^{+}17a, LFH^{+}08b, WLDL08, WGY^{+}19, XZZ16b, YZCT17, ZM13, AMRW06, BH09, CWMZ06, CZQ17, CXT^{+}18, DNB19, DHC11, DYW16, FTRA15, GLZ19, HmLGP03, IÁE11, LLWS09, LFZ07, LXL^{+}09, MHK^{+}18, MSST15, PZH^{+}15, PCS^{+}12, SIS19, TLPS18, VvSI07, WZT11, WZ16, WS19a, XLYX11a, XLMH14, YFL18, ZWLY16, Zhu07, ZYL^{+}08, ZZ11, ZX11, DHC13, GH08, GRTX18, OBD^{+}18, WGP^{+}15, ZGST08]. semantic-aware [DYW16]. semantic-based [MSST15]. semantically [CLZX10, FHO^{+}15]. Semantics [FZ08, SD11b, ZL09, Zhu15, ZS19, BGM03, CT16, DDF^{+}17, EdPG^{+}10, FMM08, FM08, KKL09, mLGP03, SRN^{+}15, VSKK09, WDQ^{+}18]. semantics-enabled [FMM08]. semaphores [SSMB15, WKL14]. Semi [IPGCMW18, KT19c, TJ17a, AQRA^{+}18, BTCB16, DLJ15, DH13, YSW19, MBB19]. Semi-asynchronous [IPGCMW18]. semi-automated [BTCB16]. Semi-supervised [KT19c, TJ17a, AQRA^{+}18, DLJ15, DH13, YSW19]. Semisupervised [TJ17b]. sender [AR16]. seniors [WT18]. sensed [Pla08, SPMP11]. sensing [CGOF15, CL19, EZJ^{+}18, GZG^{+}16, LLC^{+}15b, MWL^{+}13, MWL^{+}15, PHY^{+}18, PJW^{+}14, XLHT17, YHJ^{+}14, ZWL^{+}13, ZYLT06]. sensitive [DGW16, LWF^{+}15, PGW06, ZBP07]. Sensitivity [BRK^{+}17, OORVB14]. Sensor [SIST18, AKMZ13, AAF17, ANPR16, BNNH19, BFH17, BBB16, BAT13, CQXW14, CSB^{+}16, CS13, DLJ15, FH13, GBJ19, JNUH17, JBL15, JZMD19, JWZ13, KCBO17, KC18, KDW^{+}17, LL13, LDPZ14, LMO15, MS07, MH07, MO15, NQL^{+}17, OEP^{+}15, PMB15, SCS17a, ŠŽH17, SWS^{+}18, SGCG09, SC07a, WTEG17, WBZ10, WZS^{+}15, XBW^{+}15, XGXH15, YBO10, YKD^{+}15, ZPG10, ZGX11, dCHMJ12]. sensors [DFH10, MH07, SCS17a]. sentence [YHASZ19, YLZ18]. Sentiment [GC18, LWP19, THQ19, KTHA18]. Seoul [WKL^{+}11]. separation [Cla18]. Sequence [BS04, SHH^{+}14, AMHC11, CPS^{+}14, KT19c, LLB04, LMCL19, LS15, MP17, SCRV11, SRF13]. sequences [BWD15, CL14, dFdOSR^{+}19, HSHT14, LS15]. sequencing [KMJ14, MSL^{+}14]. sequential [AUHWJ19, Dut17, MO02b, SK04, SM19a, SLM05, TNIB17]. serial [LCH^{+}06]. serialization [BP03]. series [JLQ^{+}17, LLX^{+}15a, RTMZ13].

SERNOTATE [CHH18]. Server [Lia16, ATKH^{+}17, ACG15, CKOG10, CWL03, DFLNP07, GGS^{+}16, GGC19, HKAC14, KF18, LBdM^{+}16, LGD15, MVML11, PRS01, RGAK15, RO12b, dFMSPSW06, WDT18, ZABP18, ZBZ^{+}18]. Server-side [Lia16]. servers [AAI12, BDP18, GMPT15, KSC12, RJ01, TK10, WLW11]. Service [ADD^{+}05, CR13, FK19, GRQ19a, KTM^{+}09, LWL^{+}06, MN10, ROA^{+}07, RCXS09, RDP10, WBHW08, WL02, AaBT16, AaBT17, AP10, AAHRW04, AMRW06, ACFT15, ACS10, BTCGL17, BV16, CYD^{+}15, CLTT13, CK13, CW11b, CM06, CHH18, CM07a, CLH^{+}08, CPSP17, CM02, CRGR^{+}12, CMS17, CKBB14, DIM18, DFLNP07, DCP^{+}17, DPGA11, DXM^{+}17, EdPG^{+}10, ET09, EAGVBVDS11, EABVGV14, FCY17, FMM08, FN13, FP02, FGG^{+}18, GYM14, GLMT15, GCN09, GKP^{+}09, HAE09, HLZD18, HYL^{+}19, HFTQ13, KKTHL13, KM13, KJŚ^{+}15, KMRT18, LDPZ14, LLX^{+}15a, LWY^{+}16, LGJ17, LDXC13, LFH08a, LZC08, LW13, LLSL15, LFHT15, LLC^{+}15b, MWPL15, MWPX17, MvNK^{+}06, MSL^{+}14, MCC^{+}15, MZW^{+}16, MK15a, MPVT17, MLVBW12, NLFA19, ORdSL13, ORDG15, PSM03, PPC^{+}15, PPBB14, QEB^{+}10, RG18, RBO^{+}02, RHS17, STO17, SBBE07, SFCAV16, SGD15, SAMS19, SVY19, SKJ17, SPSNvS07, TSA^{+}19, TTV08, TZLC15, TLPS18, TWW^{+}19]. service [TPV17, VT15, VBW06, VGN^{+}16, WZZL13, WSL15, WFHT17, WSZ^{+}18, WHW10, WXLM19, XDL^{+}11, XWD^{+}12, YSL^{+}15, YLD13, YT15, YS07, YF13, YCWH07, YLJZ13, ZLY^{+}13, ZDC15, ZM13, ZFT08, ZBZH11, ZHGX16, ZFWJ19, dRL10, vdKEL10, CWZL13, DHC13, FTR15, MCCG11, TKB16]. service-aware [STO17]. service-based [CM06, EAGVBVDS11, GKP^{+}09, SBBE07, SGD15, WFHT17, YT15]. Service-oriented [ROA^{+}07, RDP10, WBHW08, AAHRW04, ACS10, CLTT13, CLH^{+}08, EABVGV14, HFTQ13, KJŚ^{+}15, LFH08a, TTV08, WZZL13, YLJZ13, ZFT08]. Services [HF05, AMBT17a, AMBT17b, ACF^{+}07, ABR^{+}06, ACMM06, AAB^{+}05, BCX15, BAZ18, BHA^{+}15b, Can06, CV07, CSL^{+}18, CPB07, CEMR19, CTY15, CR12, CT12, CSL08, CLL^{+}19, CGH^{+}06, Cuz11, DCY^{+}08, FHH15, FMP10, FKP^{+}02, FAB^{+}07, GCSB19, HFDJ10, HM16, HCD^{+}02, HLB10, Hus15, Jun16, KGGT12, KBB11, LM08, LWYM16, MG09b, MAK18, NAP^{+}07, PSLC11, PRD^{+}13, PGP^{+}10, PCS^{+}12, RBP12, RHS17, SACRGL18, STO17, SDB02, SM04, SPJ14, SFH13, SAM^{+}17, TSL15, VŠC17, WBC^{+}02, WL02, WGG^{+}07, XXX15, YESG^{+}17b, YESG^{+}17a, YJZZ19, ZlC15, ZZ18, ZWF^{+}06, AFPO08, CEH^{+}06, GMS09, MSL^{+}14, PWWR05, WGP^{+}15]. services-based [HFDJ10]. servicing [OK15]. servlet [BPdM06]. SERVmegh [KSK17]. SeShare [HZY^{+}19]. session [JK10]. sessions [AG17a, TAB^{+}06, YLY04]. set [BGM03, BXLJ16, BHBD13, CSMS^{+}19, FJP^{+}05, Kuł14, LHC14, WCR^{+}14, YLWZ18, YY19, vRKS03, TJD^{+}17]. set-neural [YY19]. set-oriented [BGM03]. sets [BZdR^{+}10, LZZ^{+}15, MKKB04, RKS02, YWBS19]. setting [GSWJ19, MML16, WLZ^{+}18]. settings [GSK19, KHL^{+}17a, WW08, ZQR^{+}19]. settlement [KSB^{+}19]. Seventh [BL17]. several [dCPD13]. SGAM [ZLH^{+}15]. SGAS [GEJ^{+}08]. SGI [LL01, LKJ03, LSK04, PIH04]. shadow [LLT19b, ZZD^{+}17]. Shafer [ECP18, JLQ^{+}17]. shale [CNAQ18]. shallow [VLF^{+}13]. shallow-water [VLF^{+}13]. ShanghaiGrid [LWL^{+}06]. Shannon [PSIP16]. shape [QML^{+}17]. shaping [MB15]. shared [BOF15, BB02, BDV02, CFPJ^{+}17, CBPP02, DIK14, Kes04, KC06, LCA^{+}19, LHC14, MVWJ14, MLC04, PCVZ^{+}04, PSLC11, RAFD14, XCL09, YNX^{+}16, ZP06]. shared-space [ZP06]. sharing [ADM06, BGdCCA11, GVK12, HZY^{+}19, LLLJ14, LFWS15, LWB13, LLdA08, OO18, PRP^{+}15, RSKK19, SACRGL18, TYHL12, TC17, TWN07, Tru15, WLW11, WWL^{+}17a, WMC17, WL11b, YCZ^{+}13, ZZ15, ZHM^{+}17, dRC10]. shell [LLY^{+}19a, MO02a]. shift [ZJKL10]. SHMEM [LSK04]. shop [AMTM17]. Shor [dARP17]. Short [WCZ^{+}18, ZGRSC10, LS15, QCB17, ZXX17, ZHX^{+}19]. short-term [QCB17, ZHX^{+}19]. Short-time [ZGRSC10]. Shortest [DT15b, GP07]. should [PRS01]. Shrinker [RMP13b]. Shuffle [CCO15a]. shutdown [ROQL18]. shuttle [LWSZ19]. Sichuan [JW10, MZS^{+}10]. SICSA [LS14]. side [FHH15, Hic18, Lia16]. sided [CS17, DBB^{+}16, KYBV17, LSK04]. sides [LLH^{+}17]. SIESTA [SPH13]. Sigiri [WP12]. sign [KS19c, WF18]. Signal [CSAC19, CL19, KBH15a, LHC14, RP19, RVVPD^{+}17]. signal-extraction [RVVPD^{+}17]. signals [GQR16, MB17, Zhu19]. signature [DXWD16, zGWXT09, LDZ^{+}14a, TJD^{+}17, WXY10, WWC^{+}19, YWL^{+}17a, ZSL^{+}15]. signatures [AYSZ14]. signcryption [LZT12, LMKT13]. significance [AMHC11, HSHT14, OM06a, YZZ^{+}10]. signing [GLL16, HLC19]. Silicon [CSAC19, BG14]. SIM [RMP^{+}13a, FMT16]. SIMD [CXC^{+}18, KL12b, LL16c, PA18, RMW19]. similar [LJML10, WLZ11]. Similarity [DHH^{+}13, RCCH19, AMBT17a, DNB19, DHC11, LXL^{+}09, MMW16, MJZ17, MZJ^{+}19, MPR04, RVRD10, SS17b, XLYX11a, YFL18, YWR^{+}19, YLZ18, ZSWS18, ZZ14, ZLG^{+}19, ZHZ^{+}13]. Similarity-based [DHH^{+}13, SS17b]. SimMon [ZYZC17]. Simple [Cog04, WDT18, HTHW16, Kuł14, MMS17, NIIU17, ZYW^{+}16]. simplicity [RIFR10]. simplified [LPG^{+}14]. Simpósio [DC19a]. simulate [BBSW17, VŠC17]. simulated [DS19, HXY^{+}12, MK15b, WYZ12, WZLL18]. Simulating [CMD11, Eng15, Lyo02, The01, BDY02, EDBS08, SCV^{+}08]. Simulation [Ano02, CDMS15, EN09, KSM^{+}08a, MZS^{+}10, Tur04, vLRF^{+}02, ATVLM14, AML^{+}15, AAV^{+}15, BBdS^{+}17, BM02, CCO15a, CNAQ18, CGN15, CSB^{+}16, CRV15, DFG18a, DBGA16, DVB14, DMR^{+}07, DLK^{+}18, FAPC16, FMT16, FBS16, FRU12, zGWXT09, HMPPT13, HLCW15, ISS^{+}02, IPGCMW18, IBvA^{+}02, JK06, KKS12, KCZ^{+}05, LKPM09, LCT16, LLL^{+}19, MGBC16, MT19a, MHRI14, MT09, Not16a, Ogi02, PCF^{+}17, PIGK16, RHBK11, Sch02, SFH13, SFT15, TSL^{+}19, TRH^{+}02, TSS18, VDPC03, VLF^{+}13, WJLD09, XRD^{+}17, XLY^{+}16, YYC^{+}19, YPLJ11, ZDB^{+}14, ZJS^{+}17, ZLL19, ZYZC17, ZFT08, dARP17, SFN12]. simulation-based [DBGA16]. simulations [AHP^{+}13, ABC^{+}15, AMSR14, BCA^{+}10, BFM^{+}10, BDW14, BDY03, DVD^{+}12, DGJ11, DBR13, EFM17, FBV^{+}13, GQH17, GKS09, GBG^{+}14, HTR10, KDC17, KF11, LW05, LXW^{+}16, LTM^{+}14, MCY^{+}10, MFF04, MT08, MWLS11, Nak02, OKP16, PML^{+}05, QWZ^{+}19, RTPPH12, RCA^{+}11, RDP10, SNK^{+}15, SWB12, SHP14, TGB^{+}10, VLJ17, VL17, WDG^{+}14, XMJ17, YDB^{+}13, ZKJ^{+}07, ZCD^{+}12]. simulator [AFG16, CPD^{+}17, DGR^{+}07, EN16, GMVRGS15, LLRS03, RMP^{+}13a, KHW05]. simulators [DMMA17]. simultaneous [CCC12a, PHCR09, SRF13, SLGL16]. Sina [ZCLL19]. single [GZHF19, GP07, JKZ03, KM03, MCP^{+}12, MA15, MWLS11, OTG^{+}07, SHST13, ZLQ^{+}18]. single-chip [MCP^{+}12]. single-domain [ZLQ^{+}18]. single-hop [JKZ03, MA15]. single-level [KM03]. single-phase [MWLS11]. single-source [GP07]. single-user [SHST13]. singleton [TVCB18]. singular [ZLT^{+}16]. sink [WZS^{+}15]. sink-based [WZS^{+}15]. SINS [CL19]. SINS-aided [CL19]. SIP [ASE^{+}17]. SipaaS [TKB16]. Sistemas [DC19a]. site [YLC11]. sites [GIL17]. situ [BWHS18, LSE^{+}13]. situated [CGN15]. situational [LBOE18]. Sixth [BJ18]. size [VL17, ZWY^{+}19]. sized [FHO^{+}15, LHH^{+}17]. skeletal [GGV14]. skeleton [EK19, KS05]. Skew [WJXZ18, DCJ14, LC18, WBO16]. Skew-aware [WJXZ18]. skip [DFG17]. skyline [DDM16, LRLY17, LLRS19, WYZ^{+}17, ZZL^{+}17a, ZZL^{+}17b]. SLA [ACG15, ACG17, CVK15, DPM17, DRF07, EdPG^{+}10, EBMD13]. SLA-based [DPM17, DRF07]. SLA-driven [EdPG^{+}10]. slander [ECP18]. slant [ZDHJ18]. SLAs [RCKV12]. SLC [AMABS18]. SLC/MLC/TLC [AMABS18]. sleep [XRD^{+}17]. slicing [GM04]. sliding [LJL^{+}17]. Slim [YESG^{+}19]. slip [OKM10]. SLN [ZX11]. slopes [ZWZ^{+}18]. slot [VO15]. Small [ZZC^{+}17, CFTT17, Del08, IQOvdG13, LYF^{+}17, SZR16]. small-footprint [SZR16]. small-world [Del08]. smart [ABMMR19, ANCA19, BC16, CLW^{+}19, CL16, DFG18a, DZC16, EZJ^{+}18, EK19, GSWJ19, KN19, MBMB18, MCG18, MAS^{+}14, ORdSL13, QGZL18, RLC16, SS17a, ŠŽH17, SWZ^{+}18, XZJ11, XYS17, MZG19, YYWQ19, ZP19]. smart/micro [KN19]. SmartGridRPC [BDG^{+}10]. smartphones [ZQR^{+}19, CMCAA17]. Smith [RGB^{+}15, ZDX12]. smoothing [TZLC15]. SMP [HR06, KNT^{+}01, NO02, ON01, ON02, TMAG03, YWC11]. SMPs [BB04, LPC^{+}14, SSB^{+}14]. SMPSs [BHL^{+}09]. SMT [PAdS^{+}17]. SMT-based [PAdS^{+}17]. Snapshot [CS09]. snapshots [GCO^{+}14]. snippets [XLYX11a]. SoC [ZAB^{+}19]. Social [CM18, DFG^{+}18b, LLC^{+}15b, OZI19, QZH16, WLWX14, AQRA^{+}18, AAHA18, AAQAR^{+}17, AMP^{+}18, BA18, BA19, CCK^{+}17, CWYX17, CDF^{+}17, DA19, DGM18, FLG19, GZG^{+}16, God12, GGC19, HLF^{+}17, HCS18, HLW^{+}19, Jun16, KT19c, KWXY18, LJPP16, LCM^{+}17, LGJ17, LL19a, MTGZ17, MLZ19, MS17a, MZA19, PWC^{+}14, PLJ18, PDCA17, PCD15, PGW^{+}08, PRP^{+}15, QMK12, RSPV17, RNAD19, SR19c, TC17, VS19, VRDTB^{+}16, WHXzL15, WWL^{+}17a, WZL^{+}17a, WMC17, XBK17, XBCW19, XZH^{+}16, XLL^{+}18, YG19, ZSWS18, ZACG16, ZHGX16, ZFW^{+}17, WLWX16]. social-based [PCD15, PGW^{+}08]. society [XXY^{+}16]. Socio [SZ11, ZX11].

Socket [KW01]. Sockets [KKJH03]. SOF [DH13]. soft [BDW14, EZJ^{+}18, PSM03, TTPJ16, GCWE15]. Software [DHM14, HY12, TSA^{+}19, Akt18a, ACS10, BLCC19, BPL^{+}19, BDH15, CBPP02, CJZ^{+}15, CLL^{+}19, CJ12, CLRB15, DVL13, FO18, GGHR16, GKM19, HPVRPF14, HCD^{+}02, HLG17, JK10, KTR11, KFS^{+}06, LLC^{+}15a, LKKL16, MST^{+}05, MJ11, MÍM19, MK12, MD02, MM10, MGR02, NTK08, PRS16, QH10, RBO^{+}02, RVVPD^{+}17, SKK01, SPLO06, SGD^{+}18, SRTG^{+}07, VEJD17, WSP17, WMvP^{+}09, XWFH08, YLL^{+}18, ZDR^{+}18, ZTGW17, vLRF^{+}02, Par02]. Software-based [DHM14]. software-defined [FO18, HLG17, ZTGW17]. software-intensive [GGHR16]. SOGC [JM19]. soil [WYY^{+}19]. solar [ADSV16, JZL14, MAVG16]. Solaris [YL01]. Solomon [CSWB11, KCS07]. solution [ASE^{+}17, BJ01, BEQOR17, Bok12, BLDW16, CNAQ18, CSTV06, CS16, DPS07, GQJL18, He19, HSHT14, HC07, JN03, JL10, LHHJ18, LTZ^{+}19, LZZ^{+}15, MSK19, NDT^{+}16, PCsHL18, SGD15, SPJ14, YHK09]. solution-based [BLDW16]. Solutions [ZQH12, BDH15, CG01, WJJM17, ZKA07]. solve [AMTM17, CNKJ18, SSIH19, YT19]. solver [BDE^{+}19, BZB17, BHPS14, CL18, DVD^{+}12, DdB01, FYKW15, GPS^{+}07, GB07, HKB07, KMJ^{+}17, LXRJ13, MCCD18, MSB17, MJD15, PS10, SSK11]. solvers [AAC^{+}15, ABF^{+}17, ABC19, ADF^{+}19, BdL06, ČSMK17, CSMK19, KHVK17, MQOQOH01, MB14, MAK18, NO02, Nak02, RHBK11, SK09]. solves [LLH^{+}17]. Solving [ABV05, BDR^{+}17, vSB06, CW07, JPWH02, KD07, LAC^{+}08, LWK15, LZZ^{+}15, PIAH12, RLVRGÁ14, SDB02, SD15, WBD^{+}19, WLLL16, ZS17, ZLT^{+}16]. SOM [YYZ^{+}17]. Some [ZQK15, CG01]. songs [RP19]. SonicMQ [MP05]. sophisticated [JVPI18]. SOR [KC13]. sort [KVGH11, PSHL11, HTBR12, HTBR16, PPdSTB17]. sorted [WBD^{+}19]. sorting [LBH07, NSN^{+}17, PSHL11, SJVR15]. sound [CMT13, PPMH15, WWY19, ZHJ19]. source [BZB17, BDP^{+}14, FE18, GBJ19, GP07, HSL19, JBL15, LHZS19, Nob08, PPC^{+}15, PPR19, TTL06, TLM17, WBC^{+}17, YGW17, YWA07, YKD^{+}15, CGK^{+}07, MM10]. source-based [JBL15]. source-location [YKD^{+}15]. source-to-source [FE18, TLM17]. sources [GD07, PHY^{+}18, YHYY19, ZSL^{+}15]. southwest [JW10, ZZYW10]. Space [GFL04, MJL01, ADK^{+}16, BYN^{+}17, CPEA18, CZ11, CXW17, GKS^{+}07, GCWE15, HKS^{+}12, JCVU15, KA09, KAP13, LCT16, LLF08, LWLZ11, LBH07, MB15, PL15, TJD^{+}17, WBD^{+}03, XXY^{+}16, ZP06, ZDB^{+}14, ZM13, ZJS11]. space-filling [LBH07]. Space-time [GFL04, MJL01]. Spaces [PPdSTB17, FGP^{+}11]. spacial [DZC16]. Spallation [CGK^{+}07]. Spam [IZXM09, LL18]. spammer [ZHC^{+}18]. spamming [WWL^{+}17b]. spanning [DLM13, NJ15]. Spar [vRKS03]. Spark [DLT^{+}16, GC19, GRW^{+}19, LXP18, RCCH19, SZT18, TVCB19, WWL^{+}17a, WZLL18]. Sparse [ALKD16, PMG19, Akt18b, ABC19, ADF^{+}19, BdL06, BHPS14, CLF^{+}17, CXC^{+}18, CNP^{+}15, ER12, FJZ^{+}14, GWW17, GW15, HZL19, JSS07, LLH^{+}17, MUKY18, MNL15, NA15, OAS^{+}15, PHCR09, PLR^{+}14, SAD13, SLB08, TDM^{+}15, TLPS18, VFG11, WZ04, WGQ^{+}18, XYSW18, YZR14, ZDG^{+}14]. spatial [ASE^{+}17, CLW^{+}15, HLL^{+}15, Jun16, KHHC13, SB19a, WCA08, ZMYA18, Zhu18]. spatial-temporal [ZMYA18, Zhu18]. Spatio [PLJ18]. Spatio-temporal [PLJ18]. SPD [YT15]. speaker [LLLS18]. Spearman [XYER16]. SPEC [GPW05, MvWL^{+}10]. Special [AS19, AHP^{+}13, Ang07, Ano02, AM07, BA04, BL17, BHD13, BM04, Ber07, BKZ^{+}13, BDB^{+}13, BL09a, BL09b, BL11a, BL11b, BL13b, BL13a, CWZL13, CCCW13, CČJ^{+}16, dOCPFJ13, CLTT13, CR08, CL08, CC09, CW11a, CR13, CL13, Che19, CKRO13, CAG^{+}13, CS09, CS06, CMT13, CM07b, CS13, DN19, DRZ13, DRS^{+}13, DVL13, DLM13, DH13, DKJ16, Du18c, EL01, EBMD13, EH18, ESG17, ETR^{+}13, Fed13, FK19, FN13, Fox01, Fox05, FG06, FZ07, FS07, FZ08, Fox17a, GG07, GM10, GTGT11, GvHKK11, GZX17, GMF01, GHPR05, HL13, HYQ17, HQoS11, HF05, HdV13, HTBR12, HMPPT13, HFTQ13, Hus15, JJGL13, JX06, KS02, KM13, KR06, Kni06, KB12, KWXY18, Lee09, LBW14, LBS15, LBT16, LBT17, LBFS17, LXRJ13, LMKT13, LB19, LV12, LDXC13, LW13, MWL^{+}13, MS13, Man08, MSP^{+}13, Mar05]. Special [MFG^{+}13, MISV13, MYS19, MLY10, MN10, MLA^{+}08, NPM19, Nar05, Nel05, NSSAK13, ODS^{+}13, OKG18, OEP^{+}15, OM06b, PLY13, Par02, PRD^{+}13, PHGK10, PW05, Pie08, PB07b, PK08, Puf13, Qiu11, QFT14, QLL10, QLS13, RMP^{+}13a, RHRB13, RK01, RBP12, RTMZ13, Run10, SHT11, SN06, SCNH07, SANB08, SRdS09, SF10, SRF13, STS19, SD11b, TM01, Tho07, TH19, TP14, TH10, TWB13, TFDA07, TSS18, TBH^{+}18, Tur04, UA18, Ur07, VCW13, WAS07, WAD12, WZZL13, WC08, WCLC13, WD07, Wis02, Xha18, XZ09, XLWZ11, XBXS13, XW13, Xu08, XJZ13, XYS17, YLD13, YLR^{+}13, YLJZ13, ZWL^{+}13, ZLY^{+}13, ZLN^{+}13, Zha08, ZYH09, ZQH12, ZYH12, ZHZ^{+}13, ZL09, vdS06b, AF14, DC19a, GWD15, HF17, LL13, MH18, PDD14, PCC17, RHT13, WR17, WCWB19, WDGK15, XBCW19, XXY^{+}16, ZRB19, ZZ17, ZC19, BM12]. Special [DDE^{+}12, HTW14, SFN12, VK12, WDM14]. specialization [DAB09b]. specialized [BP17, MPR04]. Species [CCC^{+}16]. specific [MHH16, RO12a, RO12b, ZS01, ZYZ16, ZLC17b]. Specification [BPB08, GF07, AAW^{+}02, BVGVEA11, BCC^{+}05, CWZL13, Cog03, FGG^{+}18, HM04, KCG^{+}19, MYDM06, MPHL03, PS05, YGL05, YP10]. specifications [AAP13, BBG17]. Specifying [HL13, MLL^{+}11, VH12]. Specmaster [WJP14]. Spectral [HCKF15, AT01, CSBL12, HKB07, PPP10, RZL^{+}19, YSW19]. spectral-based [RZL^{+}19]. spectre [LBJ^{+}19]. spectrometry [WJP14]. spectrum [CNKJ18, DK19a, EZJ^{+}18, LCMY13, TYHL12, WRDZ13, YCZ^{+}13]. speculation [CGR19, MGI17, PSJM13]. speculation-friendly [CGR19]. speculative [LZW17b, PSJM13, WZLL18]. speed [DPK10, DA15, IC19, LZG^{+}19, MB16, UGM18, WLF19, ZGS17, ZKJ^{+}07]. Speeding [LTM^{+}14, MT09, PIAH12]. speedup [GR14, TWB13, TWB13].

Speedup-Test [TWB13]. SPH [ZLL19]. SPH-based [ZLL19]. sphere [BKA19, JKV^{+}15]. spherical [PZ17, SEF^{+}14]. SPICE [LWY^{+}16]. spider [MS17a, FSPC^{+}02]. spilling [WLL14]. SPIRAL [BFK^{+}17]. splitting [MLWA19]. SPMD [AAW^{+}02, LG08, RRR04]. SpMV [GW15, CBIGL19]. spoke [ZDL19]. sponge [ZXW19]. spoofing [CSL^{+}19]. sports [Che18, Yu18]. spot [VKM^{+}09, TKB16]. Spotting [ZHC^{+}18]. spread [HLW^{+}19, RCA^{+}11]. Spreading [TLWZ14, JB19, LWY^{+}17]. spreadsheets [WOH^{+}13]. SPRINT [MSM^{+}14, PSM^{+}11]. SQL [CLW^{+}18]. SQORE [UAW09]. squares [ABV05, DMC^{+}18, ZF18]. squares-based [ZF18]. SR [SCLK15]. SSD [ABFL17, LHH^{+}17]. SSDs [AMABS18]. SSE [AB01, VŠ11]. SSR [DEF08]. stability [DA15]. stability-aware [DA15]. stabilization [CSL12, CHM15]. Stabilizing [BCM^{+}07, DHV03, KA16]. stable [LM08]. STAC [FNBS16]. STAC-A2 [FNBS16]. stack [AdCPdSD17, Gog11, GE06, YLL^{+}18]. StackSync [SACRGL18]. staging [ZJS^{+}17]. stale [BL04]. stale-value [BL04]. Stampede [KKW^{+}14]. Standard [SKD^{+}04, BDB^{+}13, CPSP17, LZT12, ZSL^{+}15]. standardizing [SKNH09]. Standards [GBG^{+}14, ET09, JKZ03, MRJ^{+}14]. Standards-based [GBG^{+}14, ET09, MRJ^{+}14]. STAPL [TTD^{+}05]. star [KA16, ZQK15]. StarPU [ATNW11]. start [OKW18, TWW^{+}19, WSWL12, RM03]. start-up [OKW18]. starvation [WS09]. State [MRS08, BKZ^{+}13, DZL^{+}19, DHC13, HJH19, IHB15, JOK^{+}18, KA09, KT19a, LWG^{+}15, LWS19, MG09a, MT19a, YP10, ZQD^{+}17, ZLW^{+}18]. state- [YP10]. state-of-the-art [BKZ^{+}13]. state-owned [HJH19]. state-space [KA09]. stateful [MLG15]. states [CY07, TBK^{+}15, XRD^{+}17]. Static [NJZZ19, Ano06, BFR05, CA06, GM04, KBB17, KMA04, LOKW^{+}10, SKK02]. statically [STWSP12]. stations [LHZX19]. Statistical [PRNM19, AMHC11, HSHT14, KF18, SA19, THQ19, TWB13, WRLS12]. statistically [DZ13, PPMH15]. statistics [EJD15, TZYL13, WCA08]. status [Dik07]. STBC [LLQL14]. stealing [ADK^{+}16, CZG16, GLM^{+}16, GMMT17, VB16]. Stealthy [WOH^{+}13]. Steering [WW08, CKC09, MMMP01, MP03]. steganography [PMG19]. Steiner [LWK15]. Stencil [EFY17, BP17, GEBA17, GBFP09, LBFS17, NPL19, NDL17, PRG15, RPRG17, RIWS17, RWK17, SRM^{+}15, VFAD17]. stencil-based [RIWS17, RWK17]. stencil-code [LBFS17]. stencils [TPGC15]. step [CL19, CLS14, Hun15, IS10, MZK16, PPdSTB17]. Stepwise [HvNJB15, HAA^{+}17]. Stepwise-refinement [HvNJB15]. StgDomain [QH10]. STILL [AJM12]. STMs [ASP19]. Stochastic [FMP10, CMD11, DPS07, DLZ16, EB10, HKS19, JLH^{+}16, LLRS03, LS05, MKSS16, RTPPH12, SR19a, SS17a, SB17, SHP14, TLF17, WLL03b, XWFH08, XWD^{+}12, ZLZ^{+}19]. stock [DFC12]. Stokes [FBV^{+}13, DdB01, GSV03, HKB07]. Stone [RSM01]. stopping [HM03]. Storage [AS19, AV07, AAE^{+}09, BGGL07, BD08, BRWB06, CLH^{+}16, CCL^{+}17, CCW^{+}15, CSWB11, DZL^{+}17b, DT17, DXZ^{+}16, ERZ^{+}11, GQJL18, GCWE15, GLS^{+}19, HMFK15, HGT14, HP11, HYX05, HKG08, HG11, HHPL16, ID18, JL10, JLL18, LZZ^{+}17, LZW^{+}16, LCW^{+}19, MLG15, MB18, PLC^{+}19, PWMX16, PWMX17, PK17, PPR19, QGZL18, RM19, RCC17, SIST18, SGJ^{+}17, SFCAV16, SWW^{+}16, SCLK15, WLZ^{+}18, Xha18, XGC19, XDJL18, XGXH15, XWPM19, YDL09, YXL17, YSC^{+}17, YZCT17, Yu18, YYL^{+}12, ZNT^{+}16, ZBZ^{+}18, ZFJ16]. storage-based [Yu18]. store [KM03]. stores [ZHZ^{+}13, ZFXJ19]. storing [ZZC^{+}17]. Storm [BUVS10]. Straight [NA15]. Straight-line [NA15]. Strassen [DS04]. Strategies [OGA^{+}01, SRdS09, SF10, VSK17, ABC19, AZF^{+}12, BGGS14, BGV^{+}01, BD04, BDV02, CWC10, CHZ10, CHZ12, DT01, Fer13, GS08, GRGP12, GMPT15, LHL10, LFHT15, LCMY13, MBP16, MCAB^{+}02, PLC^{+}19, PCF^{+}17, RVVPD^{+}17, SM11, SSIH19, YOBS16, ZABP18]. strategy [BGdCCA11, BPL^{+}19, CSL^{+}19, CMW02, CZL^{+}17, CAKH17, DXG13, DRS^{+}13, DS07, DC19b, FCY17, GDJ16, HKS19, HBKM06, JML^{+}16, JK13, KC18, LLC^{+}15a, LCW^{+}19, Li19, LCYJ08, LNCY11, LHH^{+}17, MYY18, MDB^{+}17, MB18, MW18, PMAL14, PGL^{+}17, PLW^{+}18, PSV19, RM03, SV09, SCLL19, SBDP15, TYHL12, TL19, WDG^{+}14, WWX^{+}19, XDJL18, XWH^{+}17, YCL11, YLC11, YYZH19, YYL^{+}12, ZLZ15, ZLH^{+}15, ZZY^{+}19, ZLZ^{+}19, dOOO^{+}12]. strategy-proof [ZLH^{+}15]. Stream [MY17, RS11, dRADFG17, CLNR18, LBOE18, LSJ16, LMDP19, MYY18, QXXZ16, RHD^{+}16, SCS17b, TJ17a, ZWL^{+}19]. Stream-based [MY17, MYY18]. streaming [ABR^{+}06, CA06, DJM12, FAB^{+}07, IHB15, JK10, LLRS19, MABP13, MBP16, TCBR11]. streams [BMPP17, DZM^{+}15, EPA15, HMPPT13, LOSJ17, PF12, TJ17b, TVCB18]. strength [JSPE15]. strike [LMS18]. stripe [LHH^{+}17]. Strong [Pun01, SA19, MZS^{+}10, MRS08, SVY19, AYSZ14]. strongly [Rav16, RSM01]. Structural [SVS^{+}08, SSZ14, MRY^{+}16, MJD15, MJD17, XLZD13, Xu19]. structural-connected [MRY^{+}16]. Structure [SCGZ19, BPL12, BDTdS13, CEM^{+}17, DPS16, DGL^{+}12, GLM^{+}16, HC07, JDH^{+}18, JHCH19, JWY^{+}17, LXYC17, LLY^{+}19a, LS05, MWL^{+}13, MJL01, MB14, QLLS15, RGL^{+}15, TKHA13, WHX19, YC19b, ZMJZ10]. Structured [CMB06, CZ11, GVC10, LB11, SGM18]. structures [CWYX17, DDF^{+}17, GS04b, MISV13, SL14, SER15, Wan18a, vRS05]. student [ZLW^{+}18]. Studied [HLW^{+}19]. studies [ABB^{+}15, EMB11, KCZ^{+}05, LOKW^{+}10, SZT19]. Study [GZHF19, GHLS19, HML19, Li19, LL19a, LHB^{+}19, PCsHL18, TCP^{+}05, XZZ^{+}16a, ZSL^{+}10, ZZYW10, ZGL19, BdL06, BY12, Bok12, CMCAA17, CHZ10, CEB^{+}18, DYY^{+}19, DT01, EGGA^{+}04, EMS11, EDB^{+}14, FPHZ19, GSB^{+}12, GKSR14, GFG^{+}09, GPP^{+}18, GRS^{+}17, HKS^{+}12, HPVRPF14, HWY^{+}17, JJGL13, KF15, KOK14, LBTE14, LLN^{+}14, LL18, LFX^{+}08, MCP^{+}12, NR08, NJ15, Pan20, PRC^{+}14, PSHL11, PBF15, PNJP19, RCC17, RNAD19, RTMZ13, RVVPD^{+}17, RGL^{+}15, RMCHMG15, SBC15, SCBH09, SvDO15, SE01, SLC^{+}18, SLC^{+}19, VJ19, WWL^{+}15, WWL^{+}17a, WTN07, YYWQ19, ZXW19, ZL19, ZCW^{+}18, dARP17, vRS05, FPHZ20]. Studying [NCWD^{+}04, ZZD^{+}17]. style [PW12]. sub [FTR15]. sub-ontologies [FTR15]. Subcarrier [HJTX17]. subcircuit [HLO^{+}16]. subdominant [RR11]. subgrade [ZCW^{+}18]. subgraph [DOJ^{+}19, ZLZ^{+}17]. submission [BWW^{+}08, CBB^{+}19, MHRI14, ZXRZ19]. submissions [BAC^{+}15]. subproblem [MB16]. subroutines [Cog04]. subscribe [BBPV05, MWPL15, MWPX17, TKK^{+}11]. subscriber [TKK^{+}11]. subscriber-defined [TKK^{+}11]. subset [Bok12, CS16, WLLL15, WLLL16]. subset-sum [Bok12, CS16, WLLL15, WLLL16]. substitute [PPMH15]. substrate [BCD^{+}02]. subsumes [BBC16]. subsystem [MO02a]. subtypes [HL13]. subway [ZTZ^{+}18]. suffix [WBO16]. suggestion [XLYX11b]. suitability [PB19b, ZCW^{+}18]. suitable [SKB^{+}17]. suite [DS02, GMT07, GPW03, MM10, MvWL^{+}10, SPQ^{+}17]. suites [GPW05, KG19]. sum [Bok12, CS16, WLLL15, WLLL16]. summarization [SL18, YHASZ19]. summarizing [CZ19]. summary [LLZ^{+}17a]. Sunway [CLF^{+}19, DLK^{+}18]. Super [EEK^{+}04, BBSW17, LCW^{+}19, XYSW18, YC19b]. super-resolution [XYSW18]. supercomputer [CLF^{+}19, DAC^{+}18, EDB^{+}14, FGC06, GKS09, JKL19, KSB^{+}19, LXW^{+}16, MV16, PIH04, SNEP14, SB18]. supercomputers [JOK^{+}18, LZW^{+}16, LGL^{+}17, PSL^{+}16, Roj19, RGL^{+}15]. supercomputing [HCC^{+}15]. superlinear [GR14]. Supernodal [ZDG^{+}14]. superposition [JVMN19]. supervised [AQRA^{+}18, DLJ15, DH13, KT19c, LML^{+}18, TJ17a, YSW19]. supervision [LZG^{+}19]. supply [CHX^{+}19, DSO^{+}01, HAJL16, HLL^{+}15, LS19b]. supplying [MABP13, MBP16]. Support [GED^{+}18, WCCL05, AHB^{+}10, ACMA07, BBCG02, BPL^{+}19, BP03, CC10, CRC15a, CRC^{+}15b, CWL03, CCC12b, CGK^{+}07, DIK14, DVL13, DHH^{+}13, FP02, God12, GMS09, HGB^{+}08, JMF09, KSPM12, LLY^{+}19b, mLGP03, LFZ07, MCG^{+}08, MS19, NAP^{+}07, NDP^{+}05, OSK^{+}01, PCD15, RMCA12, RCXS09, SKK01, SO16, SE01, SWD^{+}17, VRDTB^{+}16, WJ12, WZJD13, WBB^{+}07, YWA07]. Supported [SNM15, XZ09, DGL^{+}12, RCM12]. Supporting [ABB^{+}15, CGOF15, DFPT06, GDD^{+}04, GBD16, LK03, LCT16, LWB13, MMG03, SGG07, Cuz11, ET09, GKPT13, HAA^{+}07, JK10, KJ19b, KA11, PLY13, PC17a, WLDL08, ZHZ^{+}13, CWZL13]. Supportive [ABMMR19]. supports [KL12b, LYL07]. SURF [HPVRPF14]. surface [DCD^{+}14]. surfaces [DG11]. surgical [TWZ^{+}19]. surplus [RCKV12]. surrogate [JM19]. surveillance [ABK^{+}18, Qi17]. Survey [PWS19, AMABS18, BHKW12, DSMM^{+}15, DDF^{+}15, EJD17, GTA10, HKA19b, LAL02, LCH^{+}06, MG09a, MK15a, MJ11, Mit17a, Mit17b, Mit17c, Mit19, MM19, MSP^{+}19, RRBB11, RLZ15, RB17, SK17, Sod05, VSK17, VS19, WGP^{+}15, WLLZ18, XTB17, YJZZ19]. survivability [ET15, MAS^{+}14]. suspected [LLT19b]. sustainability [SvDO15]. Sustainable [WHXB19, BPL^{+}19, ZP19]. SVD [KYBV17]. SVDD [XLG19]. SVM [LL19b]. SW [PL15]. swap [DHM14]. swapping [BBW19]. swarm [dCPD13, DBH^{+}17, EN19, KHL17b, RK15, WZYG19, XDE^{+}04, ZHT08, ZT09, ZS17]. SWARP [PBSB04]. Sweep [YBC^{+}07, AAE^{+}09, ISO^{+}14, RMCN^{+}07, YK10]. SweGrid [GEJ^{+}08]. SWIMM [RGB^{+}15]. SwinDeW [LCYJ08]. SwinDeW-G [LCYJ08]. swirling [SPZ^{+}10]. Swiss [KBH^{+}15b]. switch [LXP^{+}12]. switched [CHM15, CKRO13, MOK04]. switched/optical [CKRO13]. switching [MVWJ14]. SX [MAH^{+}02, OCC^{+}05]. SX-6 [OCC^{+}05]. Sybil [AQRA^{+}18]. SybilTrap [AQRA^{+}18]. symbolic [FSPC^{+}02, SLC^{+}18, SLC^{+}19]. symmetric [AYN^{+}14, BDR^{+}17, BIK^{+}11, OAS^{+}15, YDS^{+}14, YTD17]. symmetrical [ZJL15]. Symposium [GJ17, GZX17, Run10, WCWB19]. SymS [ZJL15]. SYNASC [FB16]. synchronisation [WBM^{+}10]. synchronization [ASS19, BHH09, CS17, DVB14, DKJ13, JK13, JOK^{+}18, LLH^{+}17, MS05, NN07, PCT04, RCA^{+}12, SACRGL18, ZTGW17]. synchronization-free [LLH^{+}17]. synchronize [FJ05]. Synchronous [GDD^{+}04, Kes04, PSRR14, dRRdCRR16, YB12]. synchrotron [ZWW14]. synergistic [ESZ09]. synthesis [TLM17]. Synthesized [YY19]. synthetic [FBV^{+}17]. System [AS15, AFR09, GEJ^{+}08, GED^{+}18, LZG^{+}19, PXY^{+}07, XZ09, Zho06, ZBC^{+}07, ACJ10, AOK19, AMBT17b, AAC^{+}15, ABC19, ABK^{+}18, ANCA19, Ang08, ADF^{+}19, ASG^{+}08, BCI^{+}18, BHJ^{+}16, BFM^{+}06, BRWB06, BAS07, BB19, BAT13, Cha03, CZWH07, CJZ^{+}15, CZY^{+}18, CHX^{+}19, CLS14, CLRB15, CLX^{+}12, DL10, DLK^{+}18, DZZL19, DT17, DZM^{+}15, DMMA17, DCA17, EEK^{+}04, FPHZ19, FPC15, FWU^{+}04, GQJL18, GSR^{+}19, GLS^{+}19, GHB^{+}06, HDDG09, HLA^{+}18, HXY^{+}12, HCD^{+}18, HK01, HJM^{+}11, HYX05, HKG08, HG11, HH19, HY12, HON04, ISS^{+}02, ID18, IT03, IBvA^{+}02, JOC^{+}15, JDB16, JTD^{+}19, JLHH14, JK10, KGP^{+}19, KKDS19, Kar16, KM19, KBB17, KL02, KM03, KDG^{+}08, KCKC15, KAP13, KSM15, KKL09, KN19, KCZ^{+}05, KB18, LLRS03, LM08, LLWS09, LWC12, LLL15, LCC^{+}18, LCW^{+}19, LLLS18, LMDP19, LWS19, LAB^{+}06, MS17a, MS17b, MS18, MSST15, MMW^{+}12, MWRK18, MHRI14, NNvVdA09]. system [NSSAK13, NSSAK16, NS19, PB12, Pac16, PSG03, PN19, PBD^{+}15, PGF19, PPP10, PGO^{+}04, PGW^{+}08, RGAK15, RW10, RSTV07, RG17, RSTV05, SACJ04, SNB^{+}01, SM19b, STL^{+}15, SZR16, TYHL12, TTL06, TKA^{+}02, TMS^{+}12, TMAG03, VYKM19, VGL06, VZB19, WKT08, WLDL08, WXY10, WR17, WJKS18, WLL03a, WT18, XHZ12, XBB13, XTLG08, XLL^{+}12, Yan19b, YYWQ19, YL01, ZABP18, ZH08, ZEB10, ZZ18, ZL12, ZHGX16, ZZL^{+}18, ACD02, Bai17, PA08, WYW^{+}17, WKL14, Pan20]. System- [AS15]. system-aware [BFM^{+}06]. system-level [KAP13]. Systematic [ATİ17, FG16, RWK17, ABS16, FVRM15, GRQ19b, WW19]. systemic [BGV^{+}01]. Systems [AS19, FK19, FG06, Fox10, Fox17b, HTW14, Man08, MN10, OM06b, PDD14, RK01, SNM15, Ur07, XLWZ11, Zha08, ZC19, AFGL09, AM15, ALZR11, AML^{+}15, AGMR05, AC06, Ano06, APHB16, BDR^{+}17, BRK^{+}17, BBPV05, BFR05, BGdCCA11, BEDK18, BB02, BCM^{+}07, BKH08, BBM19, BKND16, BDV02, Bri16, BDP^{+}14, BLSP11, CCCW13, CGBNM17, CKOG10, CGIP16, CLTT13, CLYC16, CBPP02, CY07, CWC10, CHM15, CLT^{+}16, CCW^{+}15, CLZ^{+}17, CM06, CPXA06, CSTV06, CGN15, CN16, CWM18, CCT15, CKD^{+}19, CEM^{+}17, CDP17, DD17, Dab09a, DBGA16, DC19a, DMR^{+}07, DFPT06, DLH01, DZW^{+}11, DZL^{+}17a, DZL^{+}17b, DKJ13, DKJ16, DvNM^{+}11b, DL07, DXZ^{+}16, EGGA^{+}04, EBGS01, EB05, EJF^{+}16, EFA^{+}17, FE18, Fec12, FG16, FAPC16, FVRM15, FD01, FMT16, FN13, FBV^{+}13, Fox17a, FJG^{+}13, FM08, GGHR16, GTFA13, GSB^{+}12]. systems [GMMT17, God12, GPVCdBRO12, GPP^{+}18, GOLL17, GCL08, HKVW16, HmLGP03, HTR10, HPD^{+}15, HWY^{+}17, HCK^{+}08, IOOH12, JAU19, JAA08, JL10, JSR19b, JSS07, Jon09, KNT^{+}01, KT19a, KSN16, KAL07, KF01, KL02, KSG11, KHW05, KR15, KKK^{+}19, KSS^{+}17, KRS11, KHZ^{+}15, KD07, LLMM19, LLD19, LBTE14, LLKC08, LX08, LZW13, LDPZ14, LZW^{+}16, LYF^{+}17, Lia16, LNKZ08, LZC09, LNCY11, LTM^{+}14, LHH^{+}17, LBDS15, LRS15, LTK17, LCH^{+}06, LLQL14, LDS^{+}08, MWPL15, MBP16, MGBC16, MG09a, MSP^{+}13, MJ11, MP17, Men03, MSB17, MME13, MSK19, MWW10, MvWL^{+}10, MV16, NLYZ12, NR08, OM06a, PVR^{+}09, PLC^{+}19, PWMX16, PWMX17, PC14, PRG15, PCJ17, PPdSTB17, PT12, PQP13, QB12, RE03, RS16, RMCA12, RHT13, RHZ^{+}17, RCA^{+}12, RG17, RHBK11, RHL^{+}18, RCT03, SRS16, SJB14, SK09, SZMF19, SJISVR17, SAD13, SLV12, SDH^{+}17, SLD^{+}12]. systems [SBC15, SARL13, SCGZ19, SCS17b, SKJ17, SFH13, SFT15, SW09, SO16, SD15, SSMB15, STWSP12, SS07, SZT19, TYL^{+}15, TLF17, TL19, TKK^{+}11, TVCB18, TWN07, TW07, VLMPS^{+}18, VDPC03, VH12, WS09, WAD12, WML^{+}19, WCC04, WST^{+}17, WTN07, XPS^{+}15, Xha18, XWFH08, XWX^{+}17, XGC19, XCHY13, XPWF15, XBXS13, XXLL17, XLYL17, XBM14, XLL^{+}15, XLY^{+}16, XWPM19, YTF^{+}01, YWY^{+}10, YCL11, YGG14, YZW^{+}15, YHH13, YZR14, YLL^{+}18, YYL^{+}12, ZLKK17, ZWT^{+}18, ZTZ^{+}18, ZQZ^{+}16, ZDC^{+}09, ZFJ16, ZCXH17, ZQW^{+}17, ZJL15, Boe12, CR08, HF17, Pie08, VK12].

T3E [LSK04, PSG03]. T3E-600 [LSK04]. table [GCWE15, MA15, WTN07, ZQW^{+}17]. tables [CCG^{+}08]. tabling [AR19]. Tabu [YPLJ11, EN19]. Tackling [SKS^{+}08]. tag [God12, XBXS13]. tag-based [God12]. Tagging [TLPS18]. tags [YL16]. TaihuLight [CLF^{+}19, DLK^{+}18]. tail [QZYZ16]. tailed [GM17]. Tailor [STWSP12]. Tailor-made [STWSP12]. Tailoring [CKRO13]. tails [RVRD10]. taking [RSPV17]. Tamil [PB19b]. Taming [LYF^{+}17]. tandem [WJP14]. TAP [ZFXJ19]. target [Boe12, PWH18, ZHT08, MIGA18]. Targeted [RSPV17]. Targeting [DAC12, HZH^{+}19, JKM^{+}17, MST13]. tariff [KN19]. Task [ABC^{+}16, MSG10, RSN^{+}19, ZABP18, ABF^{+}17, ABC19, AHM06, ATNW11, BEDK18, CLT^{+}16, CGS15, CFTT17, COdO^{+}11, EB18, GDJ16, HZH^{+}19, Hun15, KQR^{+}17, KO06, KZY15, KZY^{+}18, KR04, LDPZ14, LYF^{+}17, LQL^{+}15, LSL^{+}17, MS17a, MS18, MSP^{+}13, MB14, NFG19, PCGE18, PSS^{+}18, PBF15, PV15, RR15, RR19, SAB15, SPJ14, SKJ17, STL^{+}15, TKZQ17, TYTY15, TLF17, TFG^{+}12, TJF14, XLHT17, Yos06, ZJS^{+}17, ZWT^{+}18, ZCZ^{+}19, ZCH^{+}18, dSGD14, HR06]. Task-based [ABC^{+}16, BEDK18, LSL^{+}17, MB14, PSS^{+}18, STL^{+}15, TFG^{+}12]. task-centric [PBF15]. task-parallel [ABF^{+}17, ABC19, CGS15]. TaskLocalRandom [PPMH15]. tasks [AJY^{+}15, BM08, BKSM^{+}15, KR11, MÖO17, NPTT06, PB12, PPMH15, PRV11, RF15, SRS16, SR19a, Sha15, ZS17, NB12]. TASUS [CČJ^{+}16]. Tat [LPY^{+}08]. TAU [SM03]. Taverna [CMD11, OGA^{+}06, TMF^{+}10, ZGST08]. taxonomy [CY08, NNvVdA09, RRBB11, RB17]. TB [EME19]. TB-LMI [EME19]. TCAM [WHX19]. TCAM-based [WHX19]. TCM [DZL^{+}19]. TCP [KW01, NIIU17]. TCP-Socket [KW01]. TeaLeaf [MMSG17]. tealeaves [LLY^{+}19b]. team [NFG19]. Teams [HR06]. technical [XS19]. technique [Cog04, EB18, EPB14, EMEY14, JLT06, KC15, LYS18, MIGA18, MNL15, Nam19, PYKL16, PVCS18, PB19b, RS07, SVB19, WO02, WWL^{+}17b, ZO14, ZHJ19]. Techniques [LB19, NNON02, SRdS09, AI17, AMABS18, ANH^{+}19, BLSP11, CGST17, CP14, EL01, HPS05, KBG^{+}09, Li18, LLdA08, MAVG16, MBC^{+}14, MJ11, MFG^{+}13, Mit17a, Mit17b, Mit17c, Mit19, MM19, OO18, PKB03, PGP^{+}10, ROQL18, RLZ15, RCM12, SM19a, VJ19, WJH06, WMvP^{+}09, Xha18]. technological [YLLC18]. Technologies [Ang07, CS06, Fox17b, HTW14, Nar05, STS19, SNM15, VK12, ZBEM18, ZC19, BSC^{+}15, CY15, COC18, DR15, DKJ16, Fox17a, GRTX18, LLL16, PPST09, QD17, RBP12, RHT13, RS13, SRM13a, SFH13, Sod07, VRSJ15, VSB^{+}15, XADLC15, XYS17, Zen19b, ZBE17]. Technology [Ber07, Che19, ZYH09, Boc19, BG04, Che18, DCY^{+}08, Du18b, EDB^{+}14, HM16, KJ19a, Kin04, LWL15, LLY19c, MCY^{+}07, MST^{+}05, NK19, WT18, Xu19, ZDC^{+}09, ZYH12]. tectonic [LOKW^{+}10]. telecommunication [AKW04]. telecoms [NTK08]. telehealth [PBD^{+}15]. telemonitoring [LCC^{+}18]. telerehabilitation [PBD^{+}15]. Telescope [RVVPD^{+}17]. temperature [CCC12a]. template [SC19]. templates [KCB09, LH14, MWL^{+}13]. Temporal [WGQ^{+}18, ASE^{+}17, CL01, CY07, DLH01, DZC16, LNCY11, MS05, PLJ18, ZMYA18, Zhu18]. tenancy [TSL15]. tenant [DIM18, SMFM18, VGN^{+}16]. tensor [ZF18]. TensorFlow [KSM^{+}19]. TENT [Sch02]. Terabit [LGP19]. TeraGrid [CGK^{+}07, BDI^{+}07, CM07a, CGK^{+}07, WCA08]. terascale [KHW05]. term [PLZ14, QZYZ16, QCB17, ZXX17, ZHX^{+}19]. terms [ECP18]. terrain [Str11]. territories [DFG18a]. terrorism [MZA19]. terrorism-related [MZA19]. test [BLA^{+}14, GMT07, SKR17, SSZ14, ZCW^{+}18, TWB13]. Testbed [BNNH19, IUCH^{+}17, SWP17]. Testing [Low17, Ur07, ABS16, BBA18, CL10, CTY15, CLL14, DLH01, EFG^{+}03, EHSU07, PSM^{+}11, SVS^{+}08, SSZ14, VGL16, YSL^{+}15]. tests [KB18]. text [CZ15a, HAK19, HZL^{+}16, LLLS18, PLZ14, SMM18, THM^{+}11]. text-independent [LLLS18]. textual [LHXY08, LFX^{+}08]. texture [KL19, VYKM19, WCH^{+}07]. texture-based [KL19]. Tflop [GKS09]. Tflop/s [GKS09]. TFLOPS [SLM^{+}10]. TGS [SB19a]. Thangka [HLY18]. theft [VRDTB^{+}16]. their [BDV02, EQW^{+}18, FLYL16, SSB^{+}14, ZlC15]. theoretic [BXLJ16, BJC17, CG10, ECP18, FCY17]. theoretic-based [ECP18]. Theoretical [DYY^{+}19]. theory [BCCM16, BEQOR13, CHPvdG07, CQXW14, CLW^{+}15, FXX16, Fox12, GRQ19b, Hun15, JLQ^{+}17, JDH^{+}18, JHCH19, KKW^{+}14, LZG^{+}19, NNvVdA09, RKS02, SLT^{+}06, WWS^{+}12, YY19]. therapy [PBD^{+}15]. Thermal [CC15, MO02a, TKZQ17, ACIC^{+}13, IRB19]. Thermal-aware [CC15, TKZQ17, ACIC^{+}13, IRB19]. thermal-hydraulic [MO02a]. theta [BG17, HLA^{+}18, Kri19]. theta-joins [BG17]. thin [BYN^{+}17, MBP16, PIH04]. Things [IAH^{+}15, PCJ17, AA19, CMCAA17, DFG18a, DZW^{+}11, ABMMR19, AA19, AD15, BKS18, CHX^{+}19, CLH19, EH18, GED^{+}18, KJ19a, KC18, KG19, MK15b, MB15, OKO18, SMH^{+}19, SS15b, SRN^{+}15, ZlC15, ZZY^{+}15, dMd^{+}17]. thinking [HAN19, LMH^{+}14]. thinning [JdM12]. Third [Mar05]. thousand [RMP^{+}13a]. thousand-core [RMP^{+}13a]. thread [BDH15, CMMB13, CDN15, DBH^{+}17, LZW17b, MGI17, OTT19, RAFD14, RO12a, RO12b, SSU18, TPGC15]. thread-aware [RAFD14]. thread-block [TPGC15]. thread-level [MGI17, OTT19]. threaded [BIK^{+}11, EFG^{+}03, EHSU07, TMAG03, ZJL15]. threading [QB12, YA04, Tan12]. threads [Bou06, FBV^{+}13, PSM03, ZDR^{+}18]. ThreadScope [WT10]. threat [LMDP19]. Three [JdM12, Ogi02, Wan18b, ZHJ19, Boe12, CLS14, HLC19, JN03, LDZ14b, Li19, MABP13]. Three-dimensional [JdM12, Ogi02, Wan18b, ZHJ19, JN03, MABP13]. three-party [HLC19]. three-phase [Li19]. three-point [LDZ14b]. Threshold [KR15, ZCC^{+}06, AR16, zGWXT09, KW11]. Threshold-based [KR15, ZCC^{+}06]. thresholding [LYL^{+}19]. throughput [EDB^{+}14, EB14, FMT16, JOC^{+}15, Kri05, LCYJ08, MS07, PGF19, QSX^{+}17, SKA^{+}14, SAM^{+}17, SVN12]. Thrust2D [SGM18]. Tianhe [LXW^{+}16]. Tianhe-2 [LXW^{+}16]. TIB [MP05]. TIB/RV [MP05]. ticket [MS19]. tier [HLHC12, LXK^{+}19]. tiered [GLS^{+}19, PRS16]. tightly [SV09]. Tile [WFKS18, DFLL14, HLYD12]. Tile/line [WFKS18]. tiled [BLKD08]. tiling [GKS^{+}07, GFL04, GBFP09, KKGO04, RPRG17, WFKS18]. Time [AdCPdSD17, ACCM17, Fox17b, HTW14, Tur04, VK12, ACC^{+}12, AT01, AWR17, AA19, ABK^{+}18, BVGVEA11, BLA^{+}14, Bri16, BMPP17, BJC17, CDMS15, CY07, CHX^{+}19, CN16, Cuz11, DFG^{+}18b, DVB14, DLH01, DJ19, EN09, EPA15, EAGVBVDS11, EABVGV14, FRKS12, FRdOR^{+}19, FO18, FLB^{+}05, FAB^{+}07, Fox17a, FOTW04, GGS^{+}16, GFL04, GCZ^{+}17, HZHP09, HPS12, JLQ^{+}17, JTD^{+}19, JOK^{+}18, KOO12, KHM^{+}11a, Kal11, KGK17, KD15, KvGS^{+}14, KBB11, KZY15, KSR14, KMRT18, KTZ^{+}18, KWK05, LL10, LH05, LLX^{+}15a, LS19a, LWB13, LZC09, LLC^{+}15b, LTK17, LCH^{+}06, MWPX17, MGBC16, MSP^{+}13, MJL01, MGI17, MQOQOH01, MFF04, MCCD18, MSP^{+}19, MÖO17, Not16a, OSK^{+}01, OKW18, PB12, PSM03, PZH^{+}15, PSW11, Puf13, PRU14, RS16, RF15, RHT13, RTMZ13, RVVPD^{+}17, RK15, SKS^{+}08, SSM04, SC07a, SK18, SPS17, SZR16, TJD^{+}17, TLM17]. time [TY15, TBTZ18, VHBB03, WYZ^{+}17, WCZ^{+}18, XLY^{+}16, XHD^{+}19, YJL12, ZTM12, ZXX17, ZTZ^{+}18, ZG04, ZGRSC10, NDP^{+}05, SKD^{+}04]. time-aware [DFG^{+}18b]. time-critical [LL10, MWPX17]. time-dependent [CN16, PB12]. time-efficient [TJD^{+}17]. time-evolving [DJ19]. time-independent [CDMS15]. time-limited [KTZ^{+}18]. time-multiplexing [GCZ^{+}17]. time-series [JLQ^{+}17]. time-triggered [EABVGV14]. Timeliness [ZFXJ19, LWW06]. Timeliness-aware [ZFXJ19]. Timely [CXW17, VO15]. times [MV16]. timing [vEGW06]. tire [ZGL19]. Tit [LPY^{+}08]. Tit-for-Tat [LPY^{+}08]. Titan [KSB^{+}19]. TLBs [Mit17a]. TLC [AMABS18]. TM [Jac02]. Tmall [ZYZ16]. Tmall-specific [ZYZ16]. TNO [DS02]. TOAST [RPRG17]. today [DH15, LZWD^{+}15]. Toeplitz [ABV05, PV04]. together [ADM06]. token [DHV03]. token-based [DHV03]. tolerance [ADM06, BV11, CJZ^{+}15, ET15, GG19, LFHT15, PYKL16, PGK11, XPWF15, XTLG08]. Tolerant [NDP^{+}05, ACJ10, ALL^{+}15, AAE^{+}09, BF07, BZD16, BHBD13, Fec12, FD01, KAL07, LHT^{+}09, OKBO19, PGL^{+}17, RGCC15, SZD19, WDW^{+}15, XW13, XBW^{+}15, ZJS11]. tomography [SBC15, WKB^{+}19]. tomography-based [WKB^{+}19]. tomorrow [DH15, LZWD^{+}15]. TomusBlobs [CTAB16]. Tongue [LYL^{+}19]. tool [AAV^{+}15, DHH^{+}13, FJP^{+}05, HCG07, JK06, KSR14, LCC^{+}03, OTT19, PSS^{+}19, PPR19, RMG^{+}10, SL14, TF03, VDPC03]. Toolkit [Jac02, ACG18, BM02, DKMV07, LPC^{+}14, PTCN07, SHG^{+}07, SCV^{+}08, TVCB19, ZYZC17, MTD^{+}02]. Toolkits [QEB^{+}10]. Tools [BJ18, GM10, ABF^{+}10, AGMR05, EHSU07, Ger05, GMT07, GvHKK11, Hoh06, KAMB19, LLMK18, LGdVH13, MWW10, MM10, PKB03, ZYL^{+}08]. toolset [BBGA03, GWW^{+}10, KvGS^{+}14]. tooth [ZDHJ18]. Top [XLL^{+}18, LFZ07]. Top- [XLL^{+}18, LFZ07]. TOP100 [EEK^{+}04]. ToPe [JKM^{+}17]. Topic [GLD17, GC18, ZSZ15, ZLQ^{+}18, HLF^{+}17, TLPS18, ZWW^{+}18]. Topics [Man08, LL13, QLL10]. topographic [HYL^{+}19]. topographical [ATSAK15]. TopoLens [HYL^{+}19]. topological [ZQK15]. topologies [dCPD13, PMAL14, PGL^{+}17]. Topology [RH07, VKM^{+}09, BSZ09, CKRO13, HRR^{+}11, NIIU17, PS19a, QLLS15, SWLJ17]. topology-based [NIIU17, SWLJ17]. Tor [HYLG15, WWS^{+}12]. tori [AVS^{+}19]. torus [AMVOSGAC17, LGL16a]. torus-based [LGL16a]. Total [CCM^{+}17, TY15]. tour [LLF08]. tourism [GSWJ19, WZT11]. tourist [GSWJ19]. tournaments [RLVRGÁ14]. trace [EJD17, FMT16, MSN^{+}19, CLR18]. trace-driven [FMT16]. traceability [CHX^{+}19]. traces [BHH09, CLR18, CDMS15, EJD15]. Tracing [CWM18, MGBC16, MD19, VEJD17]. track [FB16, GYS^{+}17]. trackability [ALYD17]. tracker [LWQS19]. Tracking [ATSAK15, CFV^{+}08, FM08, TL19, CCCW13, DvNM^{+}11a, JyLdZ^{+}18, PWH18, RK15, WBZ10, ZLH^{+}18]. traction [ZGL19]. trade [AAC^{+}15, ANTZ09, DD17, HBKM06, KAL07, XLYL17]. trade-off [AAC^{+}15, DD17, HBKM06, KAL07, XLYL17]. tradeoffs [MJL01]. Trading [YLLZ09, DDF16]. Traffic [HLG17, BJC17, FBV^{+}17, GHMX13, IZXM09, KMA04, LZL17a, LJML10, MRP^{+}18, MOK04, MA15, SWLJ17, SSC^{+}16, VO15, WMA07, WXY10, WF18, XHD^{+}19, ZFW^{+}17]. trails [KDG^{+}08]. trained [SNEP14]. Training [AMRT14, CCLP19, KSM^{+}19]. trajectory [PdCMdS^{+}12]. transaction [CTY15, CKOG10, LCYJ08, MS17a, MS17b, MS18, MK12, SE01]. transaction-intensive [LCYJ08]. transactional [BLL^{+}19, DVL13, GTFA13, TSKM18, WS09, WMvP^{+}09, XJAJ18, YHH13, ZDR^{+}18]. transactions [CTY15, MISV13, QLF^{+}06, YCWH07]. transceiver [NS19]. Transfer [ZLH^{+}18, AC06, AC08, DPK10, KKL06, OKBO19, PCsHL18, TZK16, YYCH10, ZGS17, ZTZ^{+}18]. transfers [MLVBW12]. transform [LZJ^{+}18, PJW^{+}14, SP16, ZJT^{+}19]. transformation [CCC12b, CC15, Cuz11, LHC14, SKK01, TXY^{+}16, WLP^{+}17, WCZ^{+}18]. transformational [vWAH^{+}02]. Transformations [OKW15, CSC^{+}17, Dut17, GKS^{+}07]. transformed [BY12, WLL14]. transforms [HP11, SEF^{+}14]. transgenic [SZD19]. transient [BG04, GZHF19]. transistor [GZHF19, JVMN19]. transition [RVD^{+}12]. translating [IS10]. translation [CDN15, LXP18, SD03]. transmission [ASWR12, Dra17, DA15, HLHC12, KBDA19, SM19b, YYZH19, YZ19]. transmission-cost [HLHC12]. transmissions [DZ13]. transparency [GMS09, SK04]. Transparent [KFS^{+}06, CJZZ10, DPST06, MD02, LWY^{+}16]. Transparent-Desktop [LWY^{+}16]. transport [CLL^{+}18, RMCHMG15, VLF^{+}13, ZDL19]. transportation [DBGA16, Zhu19]. transporting [DGW16]. Transpose [AS15, TDM^{+}15]. Transposing [KS04]. transposition [GW19]. transshipment [LXP^{+}12]. travel [KT19c, LWSZ19]. travelers [MCWL06]. traversal [MNL15]. treasures [BG04]. Tree [SB19a, ART14, BMS^{+}09, BBB16, CGR19, DLM13, FH13, GP07, LGY17, LCW^{+}17, LWK15, MLYL17, MWL18, MLWA19, NJ15, SLM04, YKA^{+}19, ZJKL10, ZLLL11, GBD16]. trees [ADK^{+}17, ESGQ^{+}11, HW14, KMBR19, Li04, MLWA19, SLM05, SN16]. trend [CDdW17, WYY^{+}19]. Trends [PCJ17, PT12, SG19, TFDA07, CY15, DS17, ESG17, PB19a, SRM13a]. Triana [CGH^{+}06, TWSM05]. Triangle [HBD18]. triangles [KW19b, ZSWS18]. triangular [CKL17, GRS^{+}17, LLH^{+}17, MSK19]. triangulated [DLW19]. triangulation [CCW04, CCW06, LCW^{+}17, YIN19]. triangulations [DG11, PZ17]. TRIBLER [PGW^{+}08]. tridiagonal [SGR19, TYL^{+}15, VLMPS^{+}18]. Tridiagonalization [YDS^{+}14]. Trie [KW19a]. triggered [EABVGV14, TL19]. Trojan [HYGF19]. trophies [BG04]. Trust [CLMM12, Cla18, GYM14, KJŚ^{+}15, NPM19, SS18, WTEG17, WRC09, WZZL13, ZCS06, AFGL09, CWYX17, DMRS15, ECP18, GGHR16, HAAWA^{+}16, MMBP12, OEP^{+}15, TLWZ14, WAD12, WZL^{+}17a, XADLC15, YZW^{+}15, YLJZ13, ZW09]. trust-aware [DMRS15, HAAWA^{+}16]. Trust-based [KJŚ^{+}15, WRC09, WZZL13, ZCS06]. Trust-oriented [GYM14]. Trusted [QMK12, ARPPM17, XLL^{+}12]. trusting [CM06]. trustworthiness [HAvI13, MCXP15, WXSH19]. trustworthiness-based [WXSH19]. trustworthy [WFS^{+}19]. truthful [GP07]. TSHMEM [LGLA15]. TSP [YT19]. tsunami [SPZ^{+}10, CZL12]. tubing [cWsThY19]. Tunable [ABC^{+}08b]. tuning [BVGVEAFG11, BLDW16, CGMJ^{+}19, FE17, HWY^{+}17, KAM11, KCG^{+}19, Lan17, LGG16, MSCS18, MNL15, PSICU18, RG19, Sør13, TSL^{+}19, MCSML07]. tunnel [JVMN19]. Tunneling [PZZ08, PZZ10]. tuples [MZ06, vRS05]. Turbo [IC19]. turbulent [RR11]. turning [HCKF15]. Twenty [BJ18]. twisted [WLQL16]. Twitter [AAQAR^{+}17, AG17b, LL18, MZA19, WWL^{+}17b]. Two [Pac16, WBZ10, BdL06, CL19, DvdS06, DNB19, DZ13, EMB11, FIO15, GSB^{+}12, HLHC12, HML20, LLKC08, LLH19, LCJ14, MJD15, RVD^{+}12, RBDI17, SCBH09, TBK06, Tru15, WLLL15, ZZC15]. two-dimensional [GSB^{+}12, TBK06]. two-hop [DZ13]. two-layer [Tru15]. Two-level [Pac16, WBZ10, DNB19, LLKC08, MJD15]. two-list [WLLL15]. two-parameter [HML20, LLH19]. two-party [FIO15, ZZC15]. two-step [CL19]. two-tier [HLHC12]. Type [CG01, WL11b, FVLS15, GE08]. Type-safe [WL11b]. types [Pun01, VJ19]. typhoon [ZCLL19]. typing [RR01].

u [WKL^{+}11]. u-City [WKL^{+}11]. U.K [WAS07, WC08]. U.K. [Xu08]. UAV [YYZH19]. Ubiquitous [MCY^{+}10, HAE09, KR15, LCC^{+}18, LDS^{+}08, MZK16, XCHY13]. UDP [GKM19]. UFS [HBKM06]. Uintah [MB14, SBS19]. ultra [KW19a]. UltraScan [MRJ^{+}14]. ultrasound [PLL17]. UMM [YGL05]. unaware [DFPT06]. uncertain [LLRS19, RGX^{+}17, RTMZ13, ZZL^{+}17b]. uncertainty [BCF12, LAC^{+}08, WHXB19, vdKEL10]. uncontrolled [CAC^{+}08]. unconventional [FS18]. uncooperative [BCK^{+}09]. undergraduate [MTVF14]. underperforming [CSL^{+}18]. underperforming-aware [CSL^{+}18]. Understanding [GGHR16, JKL^{+}17]. underwater [MS07, SWS^{+}18]. unicast [CQXW14]. UNICORE [Erw02]. Unified [BBB^{+}14, ATNW11, DvNM^{+}11a, LHC14, PLC^{+}19, SCGZ19, ZWW14, GDMT^{+}12]. uniform [Bac03, LLL^{+}19, LLYL09, WP12]. unintended [Kin04]. union [STTW18]. union-find [STTW18]. unique [HMM^{+}09]. unit [ACC^{+}12, ADF^{+}13, BDW14, CGIP16, CP14, DCJ12, DCJ14, DG11, DT15b, GWW17, GGV14, KC13, LKPM09, LDZ^{+}15, MCB14, MPSGD14, OLG^{+}15, PSRR14, RCR^{+}15, RK15, SPZ^{+}10, SAD13, lSsCY17, Str11, TZKH12, VDL^{+}15, VCW13, WCZX16, LSXL17]. unit-accelerated [BDW14, CP14, MCB14, RK15]. unit-based [DCJ12, DG11, DT15b]. units [ATVLM14, ABDP15, AUHWJ19, BHQOS15, CC15, JdM12, LPH09, LLH^{+}15, MAS16, MAdS^{+}10, PSCK^{+}15, RMP^{+}13a, RCA^{+}11, SPMP11, SHST13, SEF^{+}14, WJT^{+}14, ZO14, ZDG^{+}14]. units/multi [SEF^{+}14]. units/multi-core [SEF^{+}14]. UNITY [DD16]. Universal [IKP19]. universally [YL16]. universe [LFZ^{+}17, SHH^{+}14]. unknown [WXY10, Zhe16]. unlabeled [HZL^{+}16]. unload [YZ10, ZYL10]. unmanned [ABK^{+}18, LZH^{+}15]. unmixing [SPMP11]. Unobtrusive [MGS19]. unpacking [TNH15, TNI16]. unreliable [BPdM06]. unrolling [HBKM06, KKGO04]. unsharp [PJW^{+}14]. unstable [RR11]. unstructured [Fer13, Fer15, FYKW15, LW05, LDXC13, LAM^{+}09, NNH^{+}14, NO02, Nak02, RLMG16, SGD^{+}18, VDPC03]. unsupervised [LML^{+}18]. untrusted [ATKH^{+}17]. Unveiling [AAC^{+}15]. UPC [MTK16]. UPCBLAS [GDMT^{+}12]. upcoming [BDG08]. update [FTRA15, VDPC03, WLFX17, WHX19]. updates [KTR11]. updating [YC19b]. UPGMA [LLH^{+}15]. upgrade [BUVS10]. upland [SZD19]. uplink [SHST13]. upload [MME13]. upon [CR12]. upper [PSIP16]. urban [BH05, MZG19, NLFA19, SBB^{+}15, XZH^{+}16, ZL19]. URL [LWY^{+}17]. Usability [SLB08, HYL^{+}19, KBH^{+}15b]. Usage [DRF07, AHH14, BDP^{+}14, CSL^{+}18, CBB^{+}19, FVLS15, JMF09, KTB17, PSL^{+}16, SLV12, SS18, dRC10, vLDA07]. Use [Kri19, CZ15b, FGC06, FBC10, JC07, KSK19, MFG^{+}13, PRS01, SdSL18, WST^{+}17]. use-case [WST^{+}17]. used [ZCW^{+}18]. User [Hoh06, MH18, OK15, SK04, SWZ^{+}18, TH19, ZQR^{+}19, AaBT16, AFGL09, ATKH^{+}17, BKM^{+}07a, BT18, CZWH07, CCSS10, CHZ10, CHZ12, DGM18, DJ19, HHWZ08, HCS18, JSPE15, JJGL13, KKJH03, KFS^{+}06, KDW^{+}17, MH07, MML16, OZI19, OTT19, RSC^{+}15, SHST13, Sod07, SZT19, TBTZ18, XY17, YBX^{+}17, ZWY^{+}19, dRC10, vHKT^{+}11]. user-assisted [OTT19]. user-cloudlet [YBX^{+}17]. user-friendliness [BKM^{+}07a]. user-interactive [CZWH07, HHWZ08]. user-level [CCSS10, KKJH03]. users [AAQAR^{+}17, FLYL16, GYS^{+}17, HSM14, HCD^{+}18, MLZ19, MDX14, MH07, XDP18, YAA07, ZACG16]. uses [YWL^{+}17b]. Using [AG17b, CLL14, CNP^{+}15, CFP^{+}03, DKMM14, JMF09, KW01, LLB04, LFZ07, PRD^{+}13, PLR^{+}14, PFC^{+}09, SHG^{+}07, SWD^{+}17, SS15c, TRW07, WJT^{+}14, WLR05, XYER16, YTF^{+}01, ZBP06, ZSWS18, ZBP07, ANH16, ATVLM14, ALKD16, AB01, ASE^{+}17, AD02, AMHC11, ASWR12, ATKH^{+}17, AR16, ATSAK15, ART14, ABF^{+}17, AC02, And13, ARPPM17, BDR^{+}17, BYN^{+}17, BHL^{+}09, BA18, BA19, BV16, BCCM16, BCM^{+}07, BdL06, BAZ09, BAG17, BBA18, BYT^{+}12, CGOF15, CRC^{+}15b, CSBL12, CJX^{+}19, CW07, CH04, Cla18, CSB^{+}16, CBBCD08, CPSP17, COdO^{+}11, CDN15, CDP17, CMD11, DD16, DMC^{+}18, DLZ16, DPK10, DFLL14, Dra17, DK19b, EN19, EB18, EZJ^{+}18, ERZ^{+}11, FG16, FJG^{+}13, GQH17, GRSB09, GRQ19a, GG09, GMMT17, GGV14, GQR16, Guo19, GAS19, GKM19, HDDG09, HAK19, HZL^{+}16, HP11, HAA^{+}17, HLB10, HLO^{+}16, JVPI18, JZMD19]. using [JVMN19, KS19a, KGP^{+}19, KA09, KL19, KMBR19, KB17, KMJ^{+}17, KHM^{+}11b, KKW^{+}14, KHF^{+}17, KW19b, KN19, LLD19, LW05, LYS18, LKKL16, LSH^{+}16, LGQ^{+}17, LWYZ19, LWSZ19, LS05, LTM^{+}14, LLQL14, LL19c, LGD15, MSMA19, MMW16, MTGZ17, MZJ^{+}19, MRS^{+}10, MSL^{+}14, MS17a, MS17b, MS18, MBB19, MAVG16, MMSG17, MBC^{+}14, MS10, MCB14, MRH14, MSM^{+}14, MJ19, MFC18, MvWL^{+}10, MT09, NO02, NIIU17, NNK^{+}07, NCWD^{+}04, NRR15, NSN^{+}17, NJZZ19, Ogi02, PN19, PWWR05, PDY14, PIAH12, PGdCJ^{+}18, PSV19, PPP10, PPR19, PB19b, PS19a, PCD15, PV15, PXY^{+}07, RVRD10, RS11, RTMZ13, RCA^{+}11, RVVPD^{+}17, RCCH19, RCLSK16, RSTV05, RK15, SM02, SJ19, SP16, SZT18, SNEP14, SS19a, SS19b, SAdB^{+}16, SM19a, SPJ14, SNB^{+}01, SWB12, Slo06, SM19b, SA19, SS19c, SC19, SVN12, SRR19, TMF^{+}10, TTPJ16]. using [TSKM18, VŠ11, VJ19, VFAD17, WBD^{+}19, WGZL06, WCA08, WBM^{+}10, XXLL17, YGW17, YAA07, YWC11, YLC11, Yan19a, YR15, YYZ^{+}17, YESG^{+}19, YWBS19, ZGL19, ZYZ16, ZWW^{+}18, ZXW16b, ZWW14, ZZL^{+}18]. Utility [LPSF11, CL07, JZL15, OISS07, PC14, TAB^{+}06]. Utilization [KCKC15, KC15, TK10, ZXW19]. utilizing [MvWvM^{+}17, Roj19, ZYH12].

V [WKL14]. V2 [MAH^{+}02]. V2G [BT18]. V8 [MGI17]. VAED [MPVT17]. validated [AFG16]. Validation [BZB17, Dut17, CY08, RGAK15, SC07b, vdABST10]. valuable [HLW^{+}19]. Value [CKOG10, BL04, LZW^{+}16, Mit17c, SAMS19, WSRM12, ZLT^{+}16, ZHZ^{+}13, ZFXJ19, DCK12]. Value-at-Risk [DCK12]. VANET [BB19, TXZ^{+}17, XZH^{+}17]. VANETs [MZW^{+}16]. vanishing [LCJ14]. Variability [GEBA17, KS19a, YC19b]. Variable [CMMS17, LBDS15, KYBV17, LHH^{+}17, MFG^{+}13, ZYL10]. variable-sized [LHH^{+}17]. variables [CMT13]. variance [JK13, SWD^{+}15]. variety [Zha19]. various [CKL17, SK18]. VASP [SSV19]. vault [SC19]. Vector [ALKD16, BF07, LL16c, Akt18b, CXC^{+}18, ER12, FVLS15, FJZ^{+}14, GWW17, GLM^{+}16, God12, GW15, GSK19, HGW18, JLT06, LLY^{+}19b, LWLZ11, MJZ17, NA15, OCC^{+}05, OAS^{+}15, PLR^{+}14, SAD13, TDM^{+}15, VFG11, VRDTB^{+}16, WJ12, WZJD13, CCO^{+}15b]. Vectorization [AdCPdSD17, BTG06]. Vectorizing [RLMG16]. vehicle [JZJW15, VZB19, WW19, WXSH19, ZY12, ZHJ19]. vehicle-to-vehicle [WXSH19]. vehicles [ABK^{+}18, LZH^{+}15, ZY12]. vehicular [IHB15, NLFA19, QWW^{+}16, SWLJ17, WXSH19, YBZ^{+}15]. velocity [ZMJZ10]. Velskii [BBB16]. Verifiable [LWW^{+}19, TC17, LZY^{+}16, LJW^{+}17]. verification [AA19, CY07, CY08, Cog03, Cog04, DLX^{+}16, KN01, LKKL16, PSW11, Sin10, SE01, YNX^{+}16]. Verified [KN01]. verifiers [AYSZ14]. verify [SWW^{+}16]. Verifying [RG18, WFS^{+}19, LSW07]. Versatile [QWW^{+}16, SPSNvS07]. version [BCD^{+}10, DXWD16, BF07]. versions [RSM01]. versus [HLCW15, Kri13, MP05, Pla08, TZK16]. Vertex [ANH16]. vertical [LXYC17, MJ19, PS19b]. Very [FAM^{+}18, BZdR^{+}10, DP19, YWBS19]. Vese [YHJ^{+}14]. VFS [PLL17]. VI [BBCG02]. via [ANPR16, DXHL17, ET15, EBMD13, GBXL17, HTHW16, HMFK15, HW16, LH14, LYL^{+}19, LPY^{+}08, MHLC^{+}05, MB14, MWL18, OKW18, QZDJ16, SWLJ17, SİM^{+}07, TZLC15, WML^{+}19, WW08, WLL14, XWX^{+}17, YXLZ16]. vibration [Xu19, Zhu19]. victim [WFJ^{+}17]. Video [ZYY^{+}19, CDF^{+}17, DZM^{+}15, IHB15, KCS07, PMG19, TSBR10, WYZ12, WGQ^{+}18, ZYZ^{+}12, ZJT^{+}19, ZHM^{+}17]. view [CWZL13, HLX^{+}16, KSM^{+}08b, LLF08, MML^{+}17, NDT^{+}16, VRSJ15, ZWLY16, ZCS06]. viewshed [DLW19]. VIGO [MTA^{+}07]. violation [BA18]. violations [KMRT18, RC09, WKZL19]. VIRGO [JLHH14]. Virtual [BP03, CKSC10, EN09, GBB^{+}15, LTK17, MW18, RIP18, SGV12, WLP^{+}17, ZS01, ZWF^{+}06, ZYLY18, AVS^{+}19, AFT01, AMAB17, AMB^{+}17, Bao19, BB12, BB15, Boc19, BDF15, BAZ09, CSMB15, CLR18, CG10, CCL^{+}17, CH04, CFV^{+}08, CPE^{+}19, DFC12, DXM^{+}17, DCA17, EDB^{+}14, EB14, EMS15, GRSB09, GPW03, GE06, GCPS^{+}14, HKS19, HG11, IRB19, JvAB^{+}15, KD10, KTB17, KBB11, KCKC15, KMG^{+}18, KBT^{+}14, KF18, LLL15, LHLH16, LMDP19, LSMVML15, LWL^{+}19, MS17c, MST15, MVML11, MRS^{+}09, PLY13, PCH^{+}08, PCB^{+}18, RGCC15, RMP13b, SJB14, SYMA17, TB12, VGL16, WKT08, WDT18, XHCL15, XTB17, XLQL18, XXY^{+}16, YBZ^{+}15, YYC^{+}19, ZXW16a, ZYN^{+}07, ZLZ15, ZLH^{+}15, ZBP07, ZZZX19, ZWH^{+}17, BBGA03, GKP^{+}19, GGR^{+}10, KKJH03, WL02]. virtualization [AKK^{+}07, Che18, EdPG^{+}10, MMG^{+}18, QZDJ16, RSC^{+}15, SIRP17]. virtualization-based [QZDJ16]. virtualized [ABtGT^{+}12, ACG18, CJZZ10, JCJ17, LJHL10, QLS13, RGAK15, RHZ^{+}17, TZG^{+}19, WTL^{+}16]. Virtualizing [WSP17]. virus [MJL01]. virus-structure [MJL01]. visibility [Str11]. vision [Dik07]. Visual [BLA^{+}14, OTO18, PDY14, BCI^{+}18, CP14, LWQS19, MYDM06, PSS^{+}18, Qi17, WBC^{+}17, LSY^{+}12]. Visualization [ZLL19, ASWR12, BDI^{+}07, BDY03, BMPS07, CMD17, FCT^{+}02, GAW09, KSM^{+}08a, LS19a, MCY^{+}10, PSLC11, PGO^{+}04, SLV12, WBHW08, ZH16]. Visualizing [SHH^{+}14, WT10, vLDA07]. visually [SHT^{+}17]. Viterbi [LDZ14b, Ree01]. VLab [NAP^{+}07]. VLCC [FAM^{+}18]. VLDB [PB07b, PK08, PB07b, PK08]. VLIW [GSG06, HBKM06, KB06, KBE07, KL12b, LHC14, LYL07, LLYL09]. VM [ZYFZ19]. VMBackup [ZXW16a]. VMI [MPVT17]. VMI-assisted [MPVT17]. vocabulary [mLGP03]. Voice [GRGP12, KBDA19, PCL17, YJL12]. void [SWS^{+}18]. VoIP [PCL17]. volatility [DCJ14, DLZ16, ZHX^{+}19]. volcano [LLT19b]. voltage [FPHZ19, Pan20, AMAB17]. volume [MHRI14, NNH^{+}14, PSG03, WJ09, YSC^{+}17, ZHX^{+}19]. volumetric [VŠ11]. voluminous [BMPP17, SMS^{+}19]. volunteer [AMGCC17, SNGR18]. volunteered [CR12, ZCLL19]. Voronoi [WCR^{+}14]. Vortals [HK07]. VOs [AAE^{+}09]. voter [ZTM12]. voting [BV16, BF07]. VP9 [ZJT^{+}19]. VPP [ISKvW02]. VR [Che18]. VRSA [WWC^{+}19]. VRSA-based [WWC^{+}19]. vs [ASP19]. VSIPL [ASS^{+}05]. vSwitch [TZG^{+}19]. vSwitch-enabled [TZG^{+}19]. vulnerabilities [GWHJL19, WLLZ18]. vulnerability [BLCC19, LKKL16, OTT19]. vulnerable [QKSJ07].

Waals [BDTdS13]. WaGe2007 [CL08]. WaGe2008 [CC09]. wait [IR11]. waiting [MV16]. walk [ANPR16, Li04]. walkabouts [GBJ19]. walking [ZWZ^{+}18]. wall [SSV19]. WAN [ZW17]. warehouse [PdCMdS^{+}12]. warehouses [MRS03]. warning [YWLQ18]. Water [YHYY19, DSO^{+}01, VLF^{+}13, VAC^{+}07, YGW17, ZXW19]. Waterman [RGB^{+}15, ZDX12]. watermarking [HAK19]. Waters [KB18]. Watershed [RHD^{+}16]. Watershed-ng [RHD^{+}16]. wave [OFR^{+}17]. waveform [RSTV05]. wavefront [GdMK^{+}18]. wavelength [WBD^{+}03, YSWZ17]. Wavelet [RHBK11, ZJT^{+}19]. Wavelet-adaptive [RHBK11]. waves [SPZ^{+}10]. way [DGW16, GCZ^{+}17, Mar19, Pre01, dRC10]. WBAN [XXCY19]. WdCM [LZC08]. Weak [JKZ03]. weakly [BF07]. wear [GHLS19]. weather [MAVG16, VCW13, ZBC^{+}07, ZDC^{+}09, ZCD^{+}12]. Weaver [BYT^{+}12]. WEb [FK19, DZW^{+}11, DHC13, HBH02, MSL^{+}14, WGP^{+}15, CPSP17, GMPT15, IPGCMW18, KMBR19, LFHT15, MLZ19, MCY^{+}10, MM19, QCB17, RG18, SS15a, TK10, WLDL08, YYK^{+}19, YLEB14, vHKT^{+}11, SAM^{+}17, ATKH^{+}17, AC02, And13, AHH14, ADD^{+}05, AAI12, BvIF10, BAZ18, CTY15, CRC^{+}15b, CWL03, CLZX10, CW07, CDL08, CHZ10, CHZ12, DCY^{+}08, DWC09, DK19b, ET15, FÁBE11, FHH15, FN13, FMP10, FP09, GH08, GRTX18, GMS09, HFDJ10, HKAC14, HF05, KGGT12, KG19, KSC12, LVN^{+}12, MMMP01, MK15a, MG09b, MWL^{+}15, MN10, MVML11, OBD^{+}18, PWWR05, PZZ08, PZZ10, PYF02, PFC^{+}09, PCS^{+}12, QEB^{+}10, RGAK15, RRWS08, SBBE07, SDB02, SAMS19, SZL09, TLPS18, TWW^{+}19, UAW09, VCP16, WBC^{+}02, WBC^{+}17, XLYX11a, XLYX11b, YSL^{+}15, YF13, ZX09, ZGST08, ZL09, Zic12]. WEb-based [FK19, WLDL08, YLEB14, vHKT^{+}11, MN10, AC02, CRC^{+}15b, CW07, LVN^{+}12, MMMP01, RRWS08, WBC^{+}17]. Web-cloud [KG19]. Web-inclusive [DWC09]. Web-of-things [DZW^{+}11]. web-portal [MCY^{+}10]. Weight [PLW^{+}18, JSR^{+}19a, ON01, Qia19]. Weight-based [PLW^{+}18]. Weighted [BF07, And13, GDJ16, Rua15, XWX^{+}17, YYWQ19]. weighting [PLZ14]. weightless [DFC12]. weights [JyLdZ^{+}18]. Weka [SS19c]. Weka4WS [TTV08]. well [PCsHL18]. well-known [PCsHL18]. Wenchuan [JW10, PWJ10]. WETICE [DR15]. Whale [KAA19]. Where [LMS18, LSS15]. Which [MS17c, PRS01]. whole [BK05, MSV^{+}10]. Wi [MB17]. Wi-Fi [MB17]. wide [BMA03, GEJ^{+}08, RLS^{+}09, RMP13b, WWL^{+}15, XPBS11, BBGA03]. wide-area [BMA03, XPBS11]. wideband [FO18]. widely [PGW06]. width [ZLF19]. Wiedemann [SAD13]. WiFi [GQR16]. wiki [BM10, DCEK15]. wikis [DSMM^{+}15]. will [FMS11]. WiMAX [TKHA13]. window [BBC16, LJL^{+}17, BBC16]. window-based [LJL^{+}17]. window-LFU [BBC16]. Window-LRFU [BBC16]. windows [CMCAA17, QB12, KF01, KBT^{+}14, XBB13]. WinGrid [MT09]. Wings [KDG^{+}08]. Wings/Pegasus [KDG^{+}08]. Winograd [DS04]. wire [Li19]. Wireless [AM07, BOB13, CSB^{+}16, HJTX17, Not16b, ŠŽH17, AKMZ13, AAF17, BNNH19, BFH17, BBB16, BAT13, CQXW14, CLH13, CGKW13, CFTT17, DLJ15, DFH10, DMA13, DZ13, DA15, DCA17, FH13, GHMX13, GBJ19, HZC^{+}14, JNUH17, JBL15, JWZ13, KCBO17, KBH15a, KC18, KDW^{+}17, LL13, LDPZ14, LMO15, MDX14, MLRR09, MO15, OEP^{+}15, OKBO19, PY19, PS19b, RS13, SGCG09, SC07a, SG18, SM19c, VT15, WTEG17, WBZ10, WRLS12, WYQ^{+}13, WZS^{+}15, WMC17, XBW^{+}15, XJZ13, XGXH15, YBO10, YHHS16, YKD^{+}15, YCWH07, ZPG10, ZL12, ZGX11, dCHMJ12]. wisdom [KJ19a]. WISP [BBG17]. within [ACC^{+}15, BPB08, BHA^{+}15b, CPE^{+}19, DvdS06, PPC^{+}15, YDB^{+}13]. without [Hey19, HM03, ON02, SBS19]. witnessing [LBJ^{+}19]. WLAN [LWZ^{+}19, YYZ^{+}17]. WLANs [XHW^{+}19]. WMSC2010 [CR13]. WMSNs [VO15]. Wolf [KB17]. word [GSG06, HLW^{+}19, YFL18, YLZ18]. word-interleaved [GSG06]. words [HLW^{+}19, XLYX11a]. Work [ADK^{+}16, GLM^{+}16, STTW18, CZG16, CZ19, DKKL06, FMS11, FRU12, GMMT17, MTHK14, MRS^{+}09, SRM^{+}15, TSBR10, VB16]. Work-efficient [STTW18]. work-stealing [CZG16, VB16]. workday [LZC08]. worker [ACIC^{+}13, CAG^{+}13, PRV11, VCP16]. Workflow [CL08, CC09, CW11a, CR13, FG06, GAS19, KKM^{+}06, ABC^{+}08a, AMS17, BBG17, CEM^{+}08, CY07, CY08, CM06, CGH^{+}06, CKBB14, Cyb06, DRS^{+}13, DCFC08, EQW^{+}18, FMMD13, GFG^{+}09, GHB^{+}06, HKS19, HZHP09, HAA^{+}07, HWZX08, JOC^{+}15, JBL16, JZZL06, KS19b, KCW09, KCKC15, KTM^{+}09, LPS^{+}09, LX08, LS19a, LZC09, LWLZ11, LNCY11, LZBF17, LAB^{+}06, LGD15, MTT15, MDB^{+}17, MWHW16, OGA^{+}06, OKP16, PLY13, PVR^{+}09, QLD^{+}11, RHRB13, RHuR^{+}19, RCXS09, RC09, RCLSK16, RRWS08, SWU08, SD11a, SPBL06, SRL^{+}14, SW11, TKB09, WKT08, WRC09, WL11a, WZZL13, WCLC13, XZHW09, YPLJ11, YYL^{+}12, ZWL^{+}15, ZCZ^{+}19, ZFT08, dSGD14, CR08]. workflow-based [RCLSK16]. workflows [BML08, BPB08, BYT^{+}12, CLTT13, CMD11, DCG11, DKKL06, DT17, DYW16, GAE^{+}06, HPHB^{+}15, Hoh06, JBL16, KB17, LPSF11, LGL^{+}17, LGL16b, LCYJ08, LZC08, MWJ^{+}10, MMW^{+}12, MYDM06, MCD^{+}15, NAK^{+}15, ODS^{+}13, RB17, Slo06, SGD^{+}18, TMF^{+}10, TCBR^{+}10, TCBR11, TC12, WGG^{+}07, XLYL17, XWH^{+}17, YCQF19, ZJS^{+}17, ZH15, dAAVS12, dOOO^{+}12]. Working [GG07]. Workload [BDV02, HWL18, SCC^{+}10, ZF14, DKMV07, HHKA14, KHW05, LLT^{+}19a, MFG^{+}13, PCF^{+}17, PGC^{+}19, SW09]. workload-aware [PCF^{+}17, PGC^{+}19]. workloads [BPT^{+}16, GGS^{+}16, KMBR19, MOF15, MCC16, PSL^{+}16, PCGE18, RGAK15, WQS^{+}16]. Workshop [Ang07, BL17, BJ18, CL08, CC09, CW11a, CR13, CS06, DDE^{+}12, DMD16, Kni06, Mar05, PB07b, PK08, QFG14, ZZ16, DC19a, Fox17a, MKO^{+}17a, OKG18, BL17, CR08, FK19, Qiu11, QFT14, Tho07, TH10]. Workshops [WDGK15, GWD15]. workspace [CBHTE11]. workstations [RCLSK16]. WorkWays [NAK^{+}15]. world [Del08, DvNM^{+}11b, FBH^{+}01, GL19, HSRN11, HM03, LSL^{+}17, RLS^{+}09, SIOS02, BBGA03, ZC19]. world-wide [RLS^{+}09, BBGA03]. WorldCIST [ZC19]. Worm [CWXW16]. wormhole [ZLC17b]. worms [GGC19]. Worst [HPS12, LLN^{+}14]. Worst-case [HPS12, LLN^{+}14]. Wound [LLZ^{+}17b]. WPAN [CLH^{+}11]. write [MWRK18]. Writing [GBFP09]. written [MLVB05]. WS [GMS09, HLZD18, PWWR05, XDL^{+}11]. WS-BPEL [HLZD18]. WS-CDL [XDL^{+}11]. WS-GAF [PWWR05]. WS-Naming [GMS09]. WSCAD [DC19a]. WSGE2006 [CR08]. WSN [KC18]. WSNs [Dra17, MMB^{+}17]. WSPE [RGV09]. WSRF [Slo06]. WSRP [YWA07]. Wudi [WYY^{+}19]. Wuyi [LLY^{+}19b].

x [RS16, Ros06, SBC15]. X-Folders [Ros06]. X-ray [SBC15]. X.509 [BFG14]. X10 [MRH14]. x86 [CL16, CL16]. x86-Android [CL16]. XC [CKD^{+}19, DAC^{+}18, MWRK18]. XC-40 [DAC^{+}18]. XC40 [Cla18, HLA^{+}18, HCD^{+}18]. Xdraw [DLW19]. XE6 [KB18]. XE6/XK7 [KB18]. Xen [RGAK15, RHZ^{+}17]. Xen-based [RGAK15]. Xeon [CBIGL19, GdMK^{+}18, KKW^{+}14, MTK16, RMW19, SSK11]. XeonPhi [CCO^{+}15b]. XK7 [KB18]. XML [AFPO08, CT11b, DXG13, SW12]. XMT [BB13, BÇG14]. XMT-2 [BB13, BÇG14]. XpressSpace [ZDB^{+}14]. XSEDE13 [WDM14]. XSEDE16 [DN19]. XSL [CCC12b]. XSS [GGC19, GSK19]. XtreemFS [HCK^{+}08]. XVIII [DC19a].

YARN [LL16b]. yoking [SZXG19]. yoking-proof [SZXG19]. Yunnan [MZS^{+}10].

Z [WFKS18]. Z-order [WFKS18]. zero [CAKH17, MVWJ14, MZK16]. zeroday [CAKH17]. zone [KABD07, SM19b]. zone-based [KABD07, SM19b]. ZOOM [CBBCD08]. Zorilla [DvNM^{+}11b]. Zynq [ZAB^{+}19].

References

Acosta:2016:PFA

[AA16] A. Acosta and F. Almeida. The particle filter algorithm: parallel implementations and performance analysis over Android mobile devices. Concurrency and Computation: Practice and Experience, 28(3):788–801, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aktas:2019:PAR

[AA19] Mehmet S. Aktas and Merve Astekin. Provenance aware runtime verification of things for self-healing Internet of Things applications. Concurrency and Computation: Practice and Experience, 31(3):e4263:1–e4263:??, February 10, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Antonioletti:2005:DIG

[AAB^{+}05] Mario Antonioletti, Malcolm Atkinson, Rob Baxter, Andrew Borley, Neil P. Chue Hong, Brian Collins, Neil Hardman, Alastair C. Hume, Alan Knox, Mike Jackson, Amy Krause, Simon Laws, James Magowan, Norman W. Paton, Dave Pearson, Tom Sugden, Paul Watson, and Martin Westhead. The design and implementation of Grid database services in OGSA-DAI. Concurrency and Computation: Practice and Experience, 17(2–4):357–376, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abo-alian:2016:KDB

[AaBT16] Alshaimaa Abo-alian, Nagwa L. Badr, and M. F. Tolba. Keystroke dynamics-based user authentication service for cloud computing. Concurrency and Computation: Practice and Experience, 28(9):2567–2585, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abo-alian:2017:ISR

[AaBT17] Alshaimaa Abo-alian, Nagwa L. Badr, and M. F. Tolba. Integrity as a service for replicated data on the cloud. Concurrency and Computation: Practice and Experience, 29(4): ??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aliaga:2015:UPE

[AAC^{+}15] José I. Aliaga, Hartwig Anzt, Maribel Castillo, Juan C. Fernández, Germán León, Joaquín Pérez, and Enrique S. Quintana-Ortí. Unveiling the performance-energy trade-off in iterative linear system solvers for multithreaded processors. Concurrency and Computation: Practice and Experience, 27 (4):885–904, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ayyub:2009:FTE

[AAE^{+}09] Shahaan Ayyub, David Abramson, Colin Enticott, Slavisa Garic, and Jefferson Tan. Fault-tolerant execution of large parameter sweep applications across multiple VOs with storage constraints. Concurrency and Computation: Practice and Experience, 21(3):377–392, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Andrade:2007:AAC

[AAF^{+}07] Diego Andrade, Manuel Arenaz, Basilio B. Fraguela, Juan Touriño, and Ramón Doallo. Automated and accurate cache behavior analysis for codes with irregular access patterns.

Concurrency and Computation: Practice and Experience, 19 (18):2407–2423, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alghamdi:2017:SAB

[AAF17] Bandar Ali Alghamdi, Marwane Ayaida, and Hacene Fouchal. Scheduling approaches in beacon-enabled mode for wireless sensor networks. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Almasoud:2018:EAI

[AAHA18] Suliman K. Almasoud, Ahmad Almogren, Mohammad Mehedi Hassan, and Iehab Alrassan. An efficient approach of improving privacy and security in online social networks. Concurrency and Computation: Practice and Experience, 30(5):??, March 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1002/cpe.4272.

Al-Ali:2004:AQS

[AAHRW04] Rashid Al-Ali, Abdelhakim Hafid, Omer Rana, and David Walker. An approach for quality of service adaptation in service-oriented Grids. Concurrency and Computation: Practice and Experience, 16(5):401–412, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Azzedin:2012:HIO

[AAI12] Farag Azzedin and Khalid Al-Issa. Handling input/output operations in Web servers: issues and countermeasures. Concurrency and Computation: Practice and Experience, 24(7): 699–710, May 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Acosta:2013:HLS

[AAP13] Alejandro Acosta, Francisco Almeida, and Ignacio Peláez. High-level specifications for automatically generating parallel code. Concurrency and Computation: Practice and Experience, 25(7):989–1012, May 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alrubaian:2017:RBC

[AAQAR^{+}17] Majed Alrubaian, Muhammad Al-Qurishi, Mabrook Al-Rakhami, Mohammad Mehedi Hassan, and Atif Alamri. Reputation-based credibility analysis of Twitter social network users. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arroqui:2015:JGT

[AAV^{+}15] M. Arroqui, J. Rodriguez Alvarez, H. Vazquez, C. Machado, C. Mateos, and A. Zunino. JASAG: a gridification tool for agricultural simulation applications. Concurrency and Computation: Practice and Experience, 27(17):4716–4740, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Allan:2002:CCS

[AAW^{+}02] Benjamin A. Allan, Robert C. Armstrong, Alicia P. Wolfe, Jaideep Ray, David E. Bernholdt, and James A. Kohl. The CCA core specification in a distributed memory SPMD framework. Concurrency and Computation: Practice and Experience, 14(5):323–345, April 25, 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/93513494/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
93513494{\&}PLACEBO=IE.pdf.

Aberdeen:2001:EFM

[AB01] Douglas Aberdeen and Jonathan Baxter. Emmerald: a fast matrix-matrix multiply using Intel’s SSE instructions. Concurrency and Computation: Practice and Experience, 13(2): 103–119, February 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/77004416/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=77004416&PLACEBO=IE.pdf.

Amaral:2015:SBS

[ABB^{+}15] Rafael Amaral, Rosa M. Badia, Ignacio Blanquer, Ricardo Braga-Neto, Leonardo Candela, Donatella Castelli, Christina Flann, Renato De Giovanni, William A. Gray, Andrew Jones, Daniele Lezzi, Pasquale Pagano, Vanderlei Perez-Canhos, Francisco Quevedo, Roger Rafanell, Vinod Rebello, Mariane S. Sousa-Baena, and Erik Torres. Supporting biodiversity studies with the EUBrazilOpenBio Hybrid Data Infrastructure. Concurrency and Computation: Practice and Experience, 27(2):376–394, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Allen:2008:TIG

[ABC^{+}08a] Gabrielle Allen, Philip Bogden, Gerry Creager, Chirag Dekate, Carola Jesch, Hartmut Kaiser, Jon McLaren, Will Perrie, Gregory W. Stone, and Xiongping Zhang. Towards an integrated GIS-based coastal forecast workflow. Concurrency and Computation: Practice and Experience, 20(14): 1637–1651, September 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amar:2008:TSG

[ABC^{+}08b] A. Amar, R. Bolze, Y. Caniou, E. Caron, A. Chis, F. Desprez, B. Depardon, J.-S. Gay, G. Le Mahec, and D. Loureiro. Tunable scheduling in a GridRPC framework. Concurrency and Computation: Practice and Experience, 20(9):1051–1069, June 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aliaga:2015:CMS

[ABC^{+}15] José I. Aliaga, José M. Badía, Maribel Castillo, Davor Davidović, Rafael Mayo, and Enrique S. Quintana-Ortí. Out-of-core macromolecular simulations on multithreaded architectures. Concurrency and Computation: Practice and Experience, 27(6):1540–1550, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agullo:2016:TBF

[ABC^{+}16] Emmanuel Agullo, Berenger Bramas, Olivier Coulaud, Eric Darve, Matthias Messner, and Toru Takahashi. Task-based FMM for heterogeneous architectures. Concurrency and Computation: Practice and Experience, 28(9):2608–2629, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aliaga:2019:EAS

[ABC19] José I. Aliaga, María Barreda, and Asunción Castaño. Energy-aware strategies for task-parallel sparse linear system solvers. Concurrency and Computation: Practice and Experience, 31(6):e4633:1–e4633:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amar:2009:RGA

[ABDO09] Lior Amar, Amnon Barak, Zvi Drezner, and Michael Okun. Randomized gossip algorithms for maintaining a distributed bulletin board with guaranteed age properties. Concurrency and Computation: Practice and Experience, 21(15):1907– 1927, October 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agosta:2015:OPP

[ABDP15] Giovanni Agosta, Alessandro Barenghi, Alessandro Di Federico, and Gerardo Pelosi. OpenCL performance portability for general-purpose computation on graphics processor units: an exploration on cryptographic primitives. Concurrency and Computation: Practice and Experience, 27(14): 3633–3660, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aupy:2013:RES

[ABDR13] Guillaume Aupy, Anne Benoit, Fanny Dufossé, and Yves Robert. Reclaiming the energy of a schedule: models and algorithms. Concurrency and Computation: Practice and Experience, 25(11):1505–1523, August 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Adhianto:2010:HOT

[ABF^{+}10] L. Adhianto, S. Banerjee, M. Fagan, M. Krentel, G. Marin, J. Mellor-Crummey, and N. R. Tallent. HPCT OOLKIT: tools for performance analysis of optimized parallel programs. Concurrency and Computation: Practice and Experience, 22 (6):685–701, April 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aliaga:2017:CTP

[ABF^{+}17] José I. Aliaga, María Barreda, Goran Flegar, Matthias Bollhöfer, and Enrique S. Quintana-Ortí. Communication in task-parallel ILU-preconditioned CG solvers using MPI + OmpSs. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alencar:2017:DSN

[ABFL17] Namom Alencar, Angelo Brayner, Jose Aguiar Filho, and Hamilton Lopes. DaC scan: a novel scan operator for exploiting SSD internal parallelism. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ashworth:2005:HTC

[ABG^{+}05] Mike Ashworth, Ian J. Bush, Martyn F. Guest, Andrew G. Sunderland, Stephen Booth, Joachim Hein, Lorna Smith, Kevin Stratford, and Alessandro Curioni. HPCx: towards capability computing. Concurrency and Computation: Practice and Experience, 17(10):1329–1361, August 25, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ayuso:2013:GBA

[ABG^{+}13] F. Ayuso, G. Botella, C. García, M. Prieto, and F. Tirado. GPU-based acceleration of bio-inspired motion estimation model. Concurrency and Computation: Practice and Experience, 25(8):1037–1056, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amanatiadis:2018:RTS

[ABK^{+}18] Angelos Amanatiadis, Loukas Bampis, Evangelos G. Karakasis, Antonios Gasteratos, and Georgios Sirakoulis. Realtime surveillance detection system for medium-altitude longendurance unmanned aerial vehicles. Concurrency and Computation: Practice and Experience, 30(7):??, April 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4145.

Alencar:2019:DJJ

[ABMdAMF19] Namom Alencar, Angelo Brayner, José Maria Monteiro, and José de Aguiar Moraes Filho. DaC–Join: a join operator for improving database performance on modern hardware. Concurrency and Computation: Practice and Experience, 31(17): e5180:1–e5180:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdel-Basset:2019:ITS

[ABMMR19] Mohamed Abdel-Basset, Gunasekaran Manogaran, Mai Mohamed, and Ehab Rushdy. Internet of Things in smart education environment: Supportive framework in the decisionmaking process. Concurrency and Computation: Practice and Experience, 31(10):e4515:1–e4515:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amoretti:2006:DGS

[ABR^{+}06] M. Amoretti, R. Bertolazzi, M. Reggiani, F. Zanichelli, and G. Conte. Designing Grid services for multimedia streaming in an e-learning environment. Concurrency and Computation: Practice and Experience, 18(8):911–923, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arora:2016:SRA

[ABS16] Vinay Arora, Rajesh Bhatia, and Maninder Singh. A systematic review of approaches for testing concurrent programs. Concurrency and Computation: Practice and Experience, 28 (5):1572–1611, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Afgan:2012:RMD

[ABtGT^{+}12] Enis Afgan, Dannon Baker, the Galaxy Team, Anton Nekrutenko, and James Taylor. A reference model for deploying applications in virtualized environments. Concurrency and Computation: Practice and Experience, 24(12): 1349–1361, August 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alonso:2005:SBT

[ABV05] P. Alonso, J. M. Badía, and A. M. Vidal. Solving the block-Toeplitz least-squares problem in parallel. Concurrency and Computation: Practice and Experience, 17(1):49–67, January 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aloisio:2002:WBA

[AC02] Giovanni Aloisio and Massimo Cafaro. Web-based access to the Grid using the Grid Resource Broker portal. Concurrency and Computation: Practice and Experience, 14(13–15):1145– 1160, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anglano:2006:PAH

[AC06] Cosimo Anglano and Massimo Canonico. Performance analysis of high-performance file-transfer systems for Grid applications. Concurrency and Computation: Practice and Experience, 18(8):807–816, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anglano:2008:FMH

[AC08] Cosimo Anglano and Massimo Canonico. The File Mover: high-performance data transfer for the grid. Concurrency and Computation: Practice and Experience, 20(1):99–123, January 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

An:2009:DPC

[AC09] Xiangdong An and Nick Cercone. Distributed parallel compilation of MSBNs. Concurrency and Computation: Practice and Experience, 21(12):1607–1631, August 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aloisio:2007:GRB

[ACC^{+}07] G. Aloisio, M. Cafaro, G. Carteni, I. Epicoco, S. Fiore, D. Lezzi, M. Mirto, and S. Mocavero. The Grid Resource Broker portal. Concurrency and Computation: Practice and Experience, 19(12):1663–1670, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdelkhalek:2012:FSM

[ACC^{+}12] Rached Abdelkhalek, Henri Calandra, Olivier Coulaud, Guillaume Latu, and Jean Roman. Fast seismic modeling and reverse time migration on a graphics processing unit cluster. Concurrency and Computation: Practice and Experience, 24 (7):739–750, May 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Afgan:2015:ECB

[ACC^{+}15] Enis Afgan, Nate Coraor, John Chilton, Dannon Baker, James Taylor, and The Galaxy Team. Enabling cloud bursting for life sciences within Galaxy. Concurrency and Computation: Practice and Experience, 27(16):4330–4343, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Artes:2017:TAG

[ACCM17] Tomàs Artés, Andrés Cencerrado, Ana Cortés, and Tomàs Margalef. Time aware genetic algorithm for forest fire propagation prediction: exploiting multi-core platforms. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arnold:2002:ING

[ACD02] Dorian C. Arnold, Henri Casanova, and Jack Dongarra. Innovations of the NetSolve Grid Computing System. Concurrency and Computation: Practice and Experience, 14(13–15): 1457–1479, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alameda:2007:OGC

[ACF^{+}07] Jay Alameda, Marcus Christie, Geoffrey Fox, Joe Futrelle, Dennis Gannon, Mihael Hategan, Gopi Kandaswamy, Gregor von Laszewski, Mehmet A. Nacar, Marlon Pierce, Eric Roberts, Charles Severance, and Mary Thomas. The Open Grid Computing Environments collaboration: portlets and services for science gateways. Concurrency and Computation: Practice and Experience, 19(6):921–942, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ananthakrishnan:2015:GPS

[ACFT15] Rachana Ananthakrishnan, Kyle Chard, Ian Foster, and Steven Tuecke. Globus platform-as-a-service for collaborative science applications. Concurrency and Computation: Practice and Experience, 27(2):290–305, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anglano:2015:FEF

[ACG15] Cosimo Anglano, Massimo Canonico, and Marco Guazzone. FC2Q: exploiting fuzzy control in server consolidation for cloud applications with SLA constraints. Concurrency and Computation: Practice and Experience, 27(17):4491–4514, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anglano:2017:FFC

[ACG17] Cosimo Anglano, Massimo Canonico, and Marco Guazzone. FCMS: a fuzzy controller for CPU and memory consolidation under SLA constraints. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anglano:2018:PFT

[ACG18] Cosimo Anglano, Massimo Canonico, and Marco Guazzone. Prometheus: a flexible toolkit for the experimentation with virtualized infrastructures. Concurrency and Computation: Practice and Experience, 30(11):??, June 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4400.

Andonie:2006:ECI

[ACGG06] R. Andonie, A. T. Chronopoulos, D. Grosu, and H. Galmeanu. An efficient concurrent implementation of a neural network algorithm. Concurrency and Computation: Practice and Experience, 18(12):1559–1573, October 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arnaldo:2013:BTA

[ACIC^{+}13] Ignacio Arnaldo, Alfredo Cuesta-Infante, J. Manuel Colmenar, José L. Risco-Martín, and José L. Ayala. Boosting the 3D thermal-aware floorplanning problem through a master-worker parallel MOEA. Concurrency and Computation: Practice and Experience, 25(8):1089–1103, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abbes:2010:DFT

[ACJ10] Heithem Abbes, Christophe Cérin, and Mohamed Jemni. A decentralized and fault-tolerant Desktop Grid system for distributed applications. Concurrency and Computation: Practice and Experience, 22(3):261–277, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akram:2007:CPI

[ACMA07] Asif Akram, Dharmesh Chohan, David Meredith, and Rob Allan. CCLRC Portal infrastructure to support research facilities. Concurrency and Computation: Practice and Experience, 19(6):751–766, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Andreozzi:2006:AQE

[ACMM06] Sergio Andreozzi, Paolo Ciancarini, Danilo Montesi, and Rocco Moretti. An approach to the quantitative evaluation of Grid services. Concurrency and Computation: Practice and Experience, 18(8):827–836, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Antos:2010:EAB

[ACS10] Dimitrios Antos, Costas Courcoubetis, and George D. Stamoulis. Economic aspects of building software for serviceoriented architectures. Concurrency and Computation: Practice and Experience, 22(14):2012–2035, September 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Afsahi:2002:ECU

[AD02] Ahmad Afsahi and Nikitas J. Dimopoulos. Efficient communication using message prediction for clusters of multiprocessors. Concurrency and Computation: Practice and Experience, 14(10):859–883, August 25, 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/97517967/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=97517967{\&}PLACEBO=
IE.pdf.

Asadpour:2015:SPP

[AD15] Mahdi Asadpour and Mohammad Torabi Dashti. Scalable, privacy preserving radio-frequency identification protocol for the Internet of Things. Concurrency and Computation: Practice and Experience, 27(8):1932–1950, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alves:2017:APS

[AdCPdSD17] Maicon Melo Alves, Reynam da Cruz Pestana, Rodrigo Alves Prado da Silva, and Lúcia M. A. Drummond. Accelerating pre-stack Kirchhoff time migration by manual vectorization. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Atkinson:2005:WSG

[ADD^{+}05] Malcolm Atkinson, David DeRoure, Alistair Dunlop, Geoffrey Fox, Peter Henderson, Tony Hey, Norman Paton, Steven Newhouse, Savas Parastatidis, Anne Trefethen, Paul Watson, and Jim Webber. Web Service Grids: an evolutionary approach. Concurrency and Computation: Practice and Experience, 17(2–4):377–389, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amor:2013:EGP

[ADF^{+}13] Margarita Amor, Ramón Doallo, Basilio B. Fraguela, José R. Herrero, Enrique S. Quintana-Ortí, and Robert Strzodka. Editorials: Graphics processing unit computing and exploitation of hardware accelerators. Concurrency and Computation: Practice and Experience, 25(8):1104–1106, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anzt:2019:APB

[ADF^{+}19] Hartwig Anzt, Jack Dongarra, Goran Flegar, Nicholas J. Higham, and Enrique S. Quintana-Ortí. Adaptive precision in block-Jacobi preconditioning for iterative sparse linear system solvers. Concurrency and Computation: Practice and Experience, 31(6):e4460:1–e4460:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alonso:2014:EPE

[ADI^{+}14] Pedro Alonso, Manuel F. Dolz, Francisco D. Igual, Rafael Mayo, and Enrique S. Quintana-Ortí. Enhancing performance and energy consumption of runtime schedulers for dense linear algebra. Concurrency and Computation: Practice and Experience, 26(15):2591–2611, October 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arafat:2016:WSG

[ADK^{+}16] Humayun Arafat, James Dinan, Sriram Krishnamoorthy, Pavan Balaji, and P. Sadayappan. Work stealing for GPUaccelerated parallel programs in a global address space framework. Concurrency and Computation: Practice and Experience, 28(13):3637–3654, September 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Avros:2017:BDT

[ADK^{+}17] R. Avros, V. Dudka, B. Křena, Z. Letko, H. Pluháčková, S. Ur, T. Vojnar, and Z. Volkovich. Boosted decision trees for behaviour mining of concurrent programmes. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Antoniu:2006:HBT

[ADM06] G. Antoniu, J.-F. Deverge, and S. Monnet. How to bring together fault tolerance and data consistency to enable Grid data sharing. Concurrency and Computation: Practice and Experience, 18(13):1705–1723, November 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alonso:2014:MPE

[ADMQO14] Pedro Alonso, Manuel F. Dolz, Rafael Mayo, and Enrique S. Quintana-Ortí. Modeling power and energy consumption of dense matrix factorizations on multicore processors. Concurrency and Computation: Practice and Experience, 26(17): 2743–2757, December 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Almeida:2019:GPD

[AdSCdR^{+}19] André Luís Barroso Almeida, Leonardo de Souza Cimino, José Estevão Eugênio de Resende, Lucas Henrique Moreira Silva, Samuel Queiroz Souza Rocha, Guilherme Aparecido Gregorio, Gustavo Silva Paiva, Saul Delabrida, Haroldo Gambini Santos, Marco Antonio Moreira de Carvalho, Andre Luiz Lins Aquino, and Joubert de Castro Lima. A generalpurpose distributed computing Java middleware. Concurrency and Computation: Practice and Experience, 31(7): e4967:1–e4967:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amato:2016:DAA

[ADSV16] Alba Amato, Beniamino Di Martino, Marco Scialdone, and Salvatore Venticinque. Distributed architecture for agentsbased energy negotiation in solar powered micro-grids. Concurrency and Computation: Practice and Experience, 28(4): 1275–1290, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amaral:2014:ESI

[AF14] J. N. Amaral and A. J. Field. Editorial: A special issue from the International Conference on Performance Engineering 2013. Concurrency and Computation: Practice and Experience, 26(12):1947–1948, August 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Armstrong:2005:PCS

[AFG^{+}05] Christopher Armstrong, Rupert W. Ford, John R. Gurd, Mikel Luján, Kenneth R. Mayes, and Graham D. Riley. Performance control of scientific coupled models in Grid environments. Concurrency and Computation: Practice and Experience, 17(2–4):259–295, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anjos:2016:BSM

[AFG16] Julio C. S. Anjos, Gilles Fedak, and Claudio F. R. Geyer. BIGhybrid: a simulator for MapReduce applications in hybrid distributed infrastructures validated with the Grid5000 experimental platform. Concurrency and Computation: Practice and Experience, 28(8):2416–2439, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agudo:2009:CAC

[AFGL09] Isaac Agudo, Carmen Fernandez-Gago, and Javier Lopez. Concurrent access control for multi-user and multi-processor systems based on trust relationships. Concurrency and Computation: Practice and Experience, 21(10):1389–1403, July 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aktas:2008:XMS

[AFPO08] Mehmet S. Aktas, Geoffrey C. Fox, Marlon Pierce, and Sangyoon Oh. XML Metadata Services. Concurrency and Computation: Practice and Experience, 20(7):801–823, May 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Armstrong:2009:CIE

[AFR09] C. W. Armstrong, R. W. Ford, and G. D. Riley. Coupling integrated Earth System Model components with BFG2. Concurrency and Computation: Practice and Experience, 21(6): 767–791, April 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aridor:2001:DIV

[AFT01] Yariv Aridor, Michael Factor, and Avi Teperman. A distributed implementation of a virtual machine for Java. Concurrency and Computation: Practice and Experience, 13(3): 221–244, March 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/78003113/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=78003113&PLACEBO=IE.pdf.

Alshareef:2017:RCM

[AG17a] Hazzaa Naif Alshareef and Dan Grigoras. Robust cloud management of MANET checkpoint sessions. Concurrency and Computation: Practice and Experience, 29(2):??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alshareef:2017:UTM

[AG17b] Hazzaa Naif Alshareef and Dan Grigoras. Using Twitter and the mobile cloud for delivering medical help in emergencies. Concurrency and Computation: Practice and Experience, 29 (24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Almeida:2005:PHS

[AGMR05] F. Almeida, D. Gonzalez, L. M. Moreno, and C. Rodriguez. Pipelines on heterogeneous systems: models and tools. Concurrency and Computation: Practice and Experience, 17(9):

1173–1195, August 10, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agarwal:2010:DCC

[AHB^{+}10] Deborah A. Agarwal, Marty Humphrey, Norm F. Beekwilder, Keith R. Jackson, Monte M. Goode, and Catharine van Ingen. A data-centered collaboration portal to support global carbon-flux analysis. Concurrency and Computation: Practice and Experience, 22(17):2323–2334, December 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ashraf:2014:EAD

[AHH14] Jamshaid Ashraf, Omar Khadeer Hussain, and Farookh Khadeer Hussain. Empirical analysis of domain ontology usage on the Web: eCommerce domain in focus. Concurrency and Computation: Practice and Experience, 26(5):1157–1184, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anzt:2015:EAM

[AHK^{+}15] Hartwig Anzt, Blake Haugen, Jakub Kurzak, Piotr Luszczek, and Jack Dongarra. Experiences in autotuning matrix multiplication for energy minimization on GPUs. Concurrency and Computation: Practice and Experience, 27(17):5096–5113, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alonso:2006:GGM

[AHM06] J. M. Alonso, V. Hernández, and G. Moltó. GMarte: Grid middleware to abstract remote task execution. Concurrency and Computation: Practice and Experience, 18(15):2021– 2036, December 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agarwal:2013:SIP

[AHP^{+}13] Pratul K. Agarwal, Scott Hampton, Jeffrey Poznanovic, Arvind Ramanthan, Sadaf R. Alam, and Paul S. Crozier. Special issue papers: Performance modeling of microsecond scale biological molecular dynamics simulations on heterogeneous architectures. Concurrency and Computation: Practice and Experience, 25(10):1356–1375, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abawajy:2017:EAT

[AI17] Jemal H. Abawajy and Rafiqul Islam. Editorial: Applications and techniques in information and network security. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aiash:2015:FAA

[Aia15] Mahdi Aiash. A formal analysis of authentication protocols for mobile devices in next generation networks. Concurrency and Computation: Practice and Experience, 27(12): 2938–2953, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Airaj:2017:ECD

[Air17] Mohammed Airaj. Enable cloud DevOps approach for industry and higher education. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Jaroodi:2012:MSE

[AJM12] Jameela Al-Jaroodi and Nader Mohamed. Middleware is STILL Everywhere!!! Concurrency and Computation: Practice and Experience, 24(16):1919–1926, November 2012. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Jaroodi:2005:JJO

[AJMJS05] Jameela Al-Jaroodi, Nader Mohamed, Hong Jiang, and David Swanson. JOPI: a Java object-passing interface. Concurrency and Computation: Practice and Experience, 17(7–8):775–795, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Althebyan:2015:EMR

[AJY^{+}15] Qutaibah Althebyan, Yaser Jararweh, Qussai Yaseen, Omar AlQudah, and Mahmoud Al-Ayyoub. Evaluating map reduce tasks scheduling algorithms over cloud computing infrastructure. Concurrency and Computation: Practice and Experience, 27(18):5686–5699, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aleksy:2001:ASB

[AK01] M. Aleksy and A. Korthaus. Access to SAP’s Business Framework from Java-based applications. Concurrency and Computation: Practice and Experience, 13(7):621–641, June 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/83002172/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
83002172&PLACEBO=IE.pdf.

Almulla:2013:CKE

[AKG13] M. Almulla, A. Kanso, and M. Ghebleh. A concurrent key exchange protocol based on commuting matrices. Concurrency and Computation: Practice and Experience, 25(5):743–751, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Adamski:2007:SPE

[AKK^{+}07] Marcin Adamski, Michal Kulczewski, Krzysztof Kurowski, Jarek Nabrzyski, and Alastair Hume. Security and performance enhancements to OGSA-DAI for Grid data virtualization. Concurrency and Computation: Practice and Experience, 19(16):2171–2182, November 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Armstrong:2006:CCM

[AKM^{+}06] Rob Armstrong, Gary Kumfert, Lois Curfman McInnes, Steven Parker, Ben Allan, Matt Sottile, Thomas Epperly, and Tamara Dahlgren. The CCA component model for highperformance scientific computing. Concurrency and Computation: Practice and Experience, 18(2):215–229, February 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdallah:2013:RBW

[AKMZ13] Nesrine Ouled Abdallah, Hatem Hadj Kacem, Mohamed Mosbah, and Akka Zemmari. Randomized broadcasting in wireless mobile sensor networks. Concurrency and Computation: Practice and Experience, 25(2):203–217, February 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aktas:2018:HCC

[Akt18a] Mehmet S. Aktas. Hybrid cloud computing monitoring software architecture. Concurrency and Computation: Practice and Experience, 30(21):e4694:1–e4694:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aktemur:2018:SMV

[Akt18b] Baris Aktemur. A sparse matrix–vector multiplication method with low preprocessing cost. Concurrency and Computation: Practice and Experience, 30(21):e4701:1–e4701:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Assuncao:2004:GAC

[AKW04] M. D. Assunção, F. L. Koch, and C. B. Westphall. Grids of agents for computer and telecommunication network management. Concurrency and Computation: Practice and Experience, 16(5):413–424, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Atallah:2004:ALA

[AL04] Mikhail J. Atallah and Stefano Lonardi. Augmenting LZ-77 with authentication and integrity assurance capabilities. Concurrency and Computation: Practice and Experience, 16 (11):1063–1076, September 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdelfattah:2016:POS

[ALKD16] Ahmad Abdelfattah, Hatem Ltaief, David Keyes, and Jack Dongarra. Performance optimization of sparse matrix-vector multiplication for multi-component PDE-based applications using GPUs. Concurrency and Computation: Practice and Experience, 28(12):3447–3465, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

An:2015:CCS

[ALL^{+}15] Ying An, Xi Luo, Yao Liu, Jiawei Huang, Geyong Min, and Jianxin Wang. A congestion control scheme based on probabilistic packet acceptance and drop in delay-tolerant networks. Concurrency and Computation: Practice and Experience, 27(13):3429–3445, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amaris:2019:GAS

[ALMT19] Marcos Amaris, Giorgio Lucarelli, Clément Mommessin, and Denis Trystram. Generic algorithms for scheduling applications on heterogeneous platforms. Concurrency and Computation: Practice and Experience, 31(15):e4647:1–e4647:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Andonov:2005:DPL

[ALVY05] Rumen Andonov, Dominique Lavenier, Philippe Veber, and Nicola Yanev. Dynamic programming for LR-PCR segmentation of bacterium genomes. Concurrency and Computation: Practice and Experience, 17(14):1657–1668, December 10, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdelhafidi:2017:LOC

[ALYD17] Z. Abdelhafidi, N. Lagraa, M. B. Yagoubi, and M. Djoudi. Low overhead communication-induced checkpointing protocols ensuring rollback-dependency trackability property. Concurrency and Computation: Practice and Experience, 29(21): ??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alcaraz:2011:SSF

[ALZR11] Cristina Alcaraz, Javier Lopez, Jianying Zhou, and Rodrigo Roman. Secure SCADA framework for the protection of energy control systems. Concurrency and Computation: Practice and Experience, 23(12):1431–1442, August 25, 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aksit:2001:DED

[AM01] Mehmet Aksit and Francesco Marcelloni. Deferring elimination of design alternatives in object-oriented methods. Concurrency and Computation: Practice and Experience, 13(14):1247–1279, December 10, 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/88511595/START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?ID=88511595&PLACEBO=IE.
pdf.

Awan:2007:SIP

[AM07] Irfan Awan and Geyong Min. Special issue: Performance analysis and enhancements of wireless networks. Concurrency and Computation: Practice and Experience, 19(8):1115–1118, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alaya:2015:FOB

[AM15] Mahdi Ben Alaya and Thierry Monteil. FRAMESELF: an ontology-based framework for the self-management of machine-to-machine systems. Concurrency and Computation: Practice and Experience, 27(6):1412–1426, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arroba:2017:DVF

[AMAB17] Patricia Arroba, José M. Moya, José L. Ayala, and Rajkumar Buyya. Dynamic Voltage and Frequency Scaling-aware dynamic consolidation of virtual machines for energy efficient cloud data centers. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alsalibi:2018:STA

[AMABS18] Ahmed Izzat Alsalibi, Sparsh Mittal, Mohammed Azmi Al-Betar, and Putra Bin Sumari. A survey of techniques for architecting SLC/MLC/TLC hybrid Flash memory-based SSDs. Concurrency and Computation: Practice and Experience, 30(13):??, July 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4420.

Auler:2017:HIP

[AMB^{+}17] Rafael Auler, Carlos Eduardo Millani, Alexandre Brisighello, Alisson Linhares, and Edson Borin. Handling IoT platform heterogeneity with COISA, a compact OpenISA virtual platform. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Afify:2017:ESM

[AMBT17a] Y. M. Afify, I. F. Moawad, N. L. Badr, and M. F. Tolba. Enhanced similarity measure for personalized cloud services recommendation. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Afify:2017:PRS

[AMBT17b] Yasmine M. Afify, Ibrahim F. Moawad, Nagwa L. Badr, and Mohamed F. Tolba. A personalized recommender system for SaaS services. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alonso-Monsalve:2017:NVC

[AMGCC17] Saúl Alonso-Monsalve, Félix García-Carballeira, and Alejandro Calderón. A new volunteer computing model for dataintensive applications. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agrawal:2011:PPS

[AMHC11] Ankit Agrawal, Sanchit Misra, Daniel Honbo, and Alok Choudhary. Parallel pairwise statistical significance estimation of local sequence alignment using Message Passing Interface library. Concurrency and Computation: Practice and Experience, 23(17):2269–2279, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alexandru:2015:LES

[AML^{+}15] Mihai Alexandru, Thierry Monteil, Petr Lorenz, Fabio Coccetti, and Hervé Aubert. Large electromagnetic simulation by hybrid approach on large-scale parallel computing systems. Concurrency and Computation: Practice and Experience, 27 (13):3184–3204, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amato:2018:CHS

[AMP^{+}18] Flora Amato, Vincenzo Moscato, Antonio Picariello, Francesco Piccialli, and Giancarlo Sperlí. Centrality in heterogeneous social networks for lurkers detection: an approach based on hypergraphs. Concurrency and Computation: Practice and Experience, 30(3):??, February 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4188.

Akli:2014:TEO

[AMRT14] Linda Akli, Samuel L. Moore, Lorna I. Rivera, and Patricia J. Teller. Training, education, and outreach — raising the bar. Concurrency and Computation: Practice and Experience, 26(13):2336–2343, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ali:2006:RBS

[AMRW06] Ali Shaikh Ali, Shalil Majithia, Omer F. Rana, and David W. Walker. Reputation-based semantic service discovery. Concurrency and Computation: Practice and Experience, 18(8): 817–826, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arunarani:2017:FSC

[AMS17] A. R. Arunarani, D. Manjula, and Vijayan Sugumaran. FF-BAT: a security and cost-aware workflow scheduling approach combining firefly and bat algorithms. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arellano:2014:PCS

[AMSR14] C. Arellano, S. F. Mingaleev, E. S. Sokolov, and A. Richter. The power of circuit simulations for designing photonic integrated circuits. Concurrency and Computation: Practice and Experience, 26(15):2518–2529, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aldini:2015:DRM

[AMSS15] Alessandro Aldini, Fabio Martinelli, Andrea Saracino, and Daniele Sgandurra. Detection of repackaged mobile applications through a collaborative approach. Concurrency and Computation: Practice and Experience, 27(11):2818–2838, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alekseeva:2017:PMC

[AMTM17] Ekaterina Alekseeva, Mohand Mezmaz, Daniel Tuyttens, and Nouredine Melab. Parallel multi-core hyper-heuristic GRASP to solve permutation flow-shop problem. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Andujar-Munoz:2017:ASA

[AMVOSGAC17] Francisco-José Andújar-Muñoz, Juan-Antonio Villar-Ortiz, José-Luis Sánchez-García, and Francisco-José Alfaro-Cortés. Applying search algorithms to obtain the optimal configuration of nDT torus nodes. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Attal:2019:PDC

[AMZ19] Jean-Philippe Attal, Maria Malek, and Marc Zolghadri. Parallel and distributed core label propagation with graph coloring. Concurrency and Computation: Practice and Experience, 31(2):e4355:1–e4355:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Andrade:2019:DAC

[ANCA19] Ermeson Andrade, Bruno Nogueira, Gustavo Callou, and Gabriel Alves. Dependability analysis of a cyber-physical system for smart environments. Concurrency and Computation: Practice and Experience, 31(1):e4739:1–e4739:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anderson:2013:CAW

[And13] Nicole Anderson. Context-aware Web search using dynamically weighted information fusion. Concurrency and Computation: Practice and Experience, 25(5):672–685, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anglano:2007:SIS

[Ang07] Cosimo Anglano. Special issue: Second International Workshop on Emerging Technologies for Next-generation GRID (ETNGRID 2005). Concurrency and Computation: Practice and Experience, 19(9):1251–1252, June 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anglano:2008:SSP

[Ang08] Cosimo Anglano. Segregation and scheduling for P2P applications with the interceptor middleware system. Concurrency and Computation: Practice and Experience, 20(6):601–624, April 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abawajy:2016:VRI

[ANH16] Jemal Abawajy, Mohd Izuan Hafez Ninggal, and Tutut Herawan. Vertex re-identification attack using neighbourhoodpair properties. Concurrency and Computation: Practice and Experience, 28(10):2906–2919, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Azhir:2019:DND

[ANH^{+}19] Elham Azhir, Nima Jafari Navimipour, Mehdi Hosseinzadeh, Arash Sharifi, and Aso Darwesh. Deterministic and nondeterministic query optimization techniques in the cloud computing. Concurrency and Computation: Practice and Experience, 31(17):e5240:1–e5240:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Astsatryan:2017:EOM

[ANK^{+}17] Hrachya Astsatryan, Wahi Narsisian, Aram Kocharyan, Georges Da Costa, Albert Hankel, and Ariel Oleksiak. Energy optimization methodology for e-infrastructure providers. Concurrency and Computation: Practice and Experience, 29 (10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2002:SIA

[Ano02] Anonymous. Special issue: APEC cooperation for earthquake simulation. Concurrency and Computation: Practice and Experience, 14(6–7):369–370, May/June 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/94515742/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=94515742{\&}PLACEBO=
IE.pdf.

Anonymous:2006:EBR

[Ano06] Anonymous. Erratum: Baraglia R, Ferrini R, Ritrovato P., A static mapping heuristics to map parallel applications to heterogeneous computing systems. Concurrency Computation: Practice and Experience 2005; 17: 1579–1605. Concurrency and Computation: Practice and Experience, 18(5):547, April 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/fulltext/110546382/PDFSTART. See

[BFR05].

Anonymous:2012:R

[Ano12] Anonymous. Retraction. Concurrency and Computation: Practice and Experience, 24(15):1817, October 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2013:N

[Ano13] Anonymous. Note. Concurrency and Computation: Practice and Experience, 25(2):i–ii, February 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIa

[Ano14a] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(1):i–ii, January 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIb

[Ano14b] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(2):i–ii, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIc

[Ano14c] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(3):i–ii, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IId

[Ano14d] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(4):i–ii, March 25, 2014.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIe

[Ano14e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(5):i–ii, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIf

[Ano14f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(6):i–ii, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIg

[Ano14g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(7):i–ii, May 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIh

[Ano14h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(8):i–ii, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIi

[Ano14i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(9):i–ii, June 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIj

[Ano14j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(10):i–ii, July 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIk

[Ano14k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(11):i–ii, August 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIl

[Ano14l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(12):i–ii, August 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIm

[Ano14m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(13):i–ii, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIn

[Ano14n] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(14):i–ii, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIo

[Ano14o] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(15):i–ii, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIp

[Ano14p] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(16):i–ii, November 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIq

[Ano14q] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(17):i–ii, December 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2014:IIr

[Ano14r] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 26(18):i–ii, December 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:EAP

[Ano15a] Anonymous. Editorials: Advances in parallel and distributed computing and communications. Concurrency and Computation: Practice and Experience, 27(13):3400–3402, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:ENa

[Ano15b] Anonymous. Erratum: Note. Concurrency and Computation: Practice and Experience, 27(9):2369, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:ENb

[Ano15c] Anonymous. Erratum: Note. Concurrency and Computation: Practice and Experience, 27(9):2502, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:EN

[Ano15d] Anonymous. Erratum: Note. Concurrency and Computation: Practice and Experience, 27(13):3399, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIa

[Ano15e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(1):i–ii, January 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIb

[Ano15f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(2):i–ii, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIc

[Ano15g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(3):i–ii, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IId

[Ano15h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(4):i–ii, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIe

[Ano15i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(5):i–ii, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIf

[Ano15j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(6):i–ii, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIg

[Ano15k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(7):i–ii, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIh

[Ano15l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(8):i–ii, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIi

[Ano15m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(9):i–ii, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIj

[Ano15n] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(10):i–ii, July 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIk

[Ano15o] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(11):i–ii, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIl

[Ano15p] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(12):i–ii, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIm

[Ano15q] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(13):i–ii, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIn

[Ano15r] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(14):i–ii, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIo

[Ano15s] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(15):i–ii, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIp

[Ano15t] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(16):i–ii, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIq

[Ano15u] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(17):i–ii, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2015:IIr

[Ano15v] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 27(18):i–ii, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIa

[Ano16a] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(1):1–2, January 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIb

[Ano16b] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(2):185–186, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIc

[Ano16c] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(3):579–580, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IId

[Ano16d] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(4):945–946, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIe

[Ano16e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(5):1397–1398, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIf

[Ano16f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(6):1693–1694, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIg

[Ano16g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(8):2317–2318, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIh

[Ano16h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(9):2565–2566, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIi

[Ano16i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(10):2751–2752, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIj

[Ano16j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(11):3021–3022, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIk

[Ano16k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(12):3255–3256, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIl

[Ano16l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(13):3523–3524, September 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIm

[Ano16m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(14):3707–3708, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIn

[Ano16n] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(15):3943–3944, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIo

[Ano16o] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(16):4209–4210, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIp

[Ano16p] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(17):4365–4366, December 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2016:IIq

[Ano16q] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 28(18):4461–4462, December 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIa

[Ano17a] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(1):??, January 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIb

[Ano17b] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(2):??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIc

[Ano17c] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IId

[Ano17d] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIe

[Ano17e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIf

[Ano17f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIg

[Ano17g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIh

[Ano17h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIi

[Ano17i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIj

[Ano17j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIk

[Ano17k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIl

[Ano17l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIm

[Ano17m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIn

[Ano17n] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIo

[Ano17o] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIp

[Ano17p] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIq

[Ano17q] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIr

[Ano17r] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIs

[Ano17s] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIt

[Ano17t] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIu

[Ano17u] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIv

[Ano17v] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIw

[Ano17w] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2017:IIx

[Ano17x] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IIa

[Ano18a] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IIb

[Ano18b] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(2):??, January 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IIc

[Ano18c] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(17):e4320:1–e4320:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IId

[Ano18d] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(18):e4321:1–e4321:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IIe

[Ano18e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(19):e4322:1–e4322:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IIf

[Ano18f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(20):e4323:1–e4323:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IIg

[Ano18g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(21):e4324:1–e4324:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IIh

[Ano18h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(22):e4325:1–e4325:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IIi

[Ano18i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(23):e4326:1–e4326:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2018:IIj

[Ano18j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 30(24):e4327:1–e4327:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIa

[Ano19a] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(1):e4804:1–e4804:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIb

[Ano19b] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(2):e4805:1–e4805:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIc

[Ano19c] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(3):e4806:1–e4806:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IId

[Ano19d] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(4):e4807:1–e4807:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIe

[Ano19e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(5):e4808:1–e4808:??, March 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIf

[Ano19f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(6):e4809:1–e4809:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIg

[Ano19g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(7):e4810:1–e4810:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIh

[Ano19h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(8):e4811:1–e4811:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIi

[Ano19i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(9):e4812:1–e4812:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIj

[Ano19j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(10):e4813:1–e4813:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIk

[Ano19k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(11):e4814:1–e4814:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIl

[Ano19l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(12):e4815:1–e4815:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIm

[Ano19m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(13):e4816:1–e4816:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIn

[Ano19n] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(14):e4817:1–e4817:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIo

[Ano19o] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(15):e4818:1–e4818:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIp

[Ano19p] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(16):e4819:1–e4819:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIq

[Ano19q] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(17):e4820:1–e4820:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIr

[Ano19r] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(18):e4821:1–e4821:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIs

[Ano19s] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(19):e4822:1–e4822:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIt

[Ano19t] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(20):e4823:1–e4823:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIu

[Ano19u] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(21):e4824:1–e4824:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIv

[Ano19v] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(22):e4825:1–e4825:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIw

[Ano19w] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(23):e4826:1–e4826:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2019:IIx

[Ano19x] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 31(24):e4827:1–e4827:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Angelopoulos:2016:ECS

[ANPR16] C. M. Angelopoulos, S. Nikoletseas, D. Patroumpa, and C. Raptopoulos. Efficient collection of sensor data via a new accelerated random walk. Concurrency and Computation: Practice and Experience, 28(6):1796–1811, April 25, 2016.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Asplund:2009:MET

[ANTZ09] Mikael Asplund, Simin Nadjm-Tehrani, and Klemen Zagar. Middleware extensions that trade consistency for availability. Concurrency and Computation: Practice and Experience, 21 (9):1181–1203, June 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

AbdElhalim:2019:DFC

[AOK19] Eman AbdElhalim, Marwa Obayya, and Sherif Kishk. Distributed fog-to-cloud computing system: a minority game approach. Concurrency and Computation: Practice and Experience, 31(15):e5162:1–e5162:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Azatchi:2006:IGA

[AP06] Hezi Azatchi and Erez Petrank. Integrating generations with advanced reference counting garbage collectors. Concurrency and Computation: Practice and Experience, 18(9):959–995, August 10, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aktas:2010:HPH

[AP10] Mehmet S. Aktas and Marlon Pierce. High-performance hybrid information service architecture. Concurrency and Computation: Practice and Experience, 22(15):2095–2123, October 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Antoine:2016:GAL

[APHB16] Maeva Antoine, Laurent Pellegrino, Fabrice Huet, and Françoise Baude. A generic API for load balancing in distributed systems for big data management. Concurrency and Computation: Practice and Experience, 28(8):2440–2456, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Qurishi:2018:SGB

[AQRA^{+}18] Muhammad Al-Qurishi, Sk Md Mizanur Rahman, Atif Alamri, Mohamed A. Mostafa, Majed Al-Rubaian, M. Shamim Hossain, and B. B. Gupta. SybilTrap: a graph-based semi-supervised Sybil defense scheme for online social networks. Concurrency and Computation: Practice and Experience, 30(5):??, March 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4276.

Alam:2016:ATB

[AR16] Taj Alam and Zahid Raza. An adaptive threshold based hybrid load balancing scheme with sender and receiver initiated approach using random information exchange. Concurrency and Computation: Practice and Experience, 28(9):2729–2746, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Areias:2019:MDL

[AR19] Miguel Areias and Ricardo Rocha. Multi-dimensional lockfree arrays for multithreaded mode-directed tabling in Prolog. Concurrency and Computation: Practice and Experience, 31 (5):e4491:1–e4491:??, March 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Atamli-Reineh:2017:FAP

[ARPPM17] Ahmad Atamli-Reineh, Andrew Paverd, Giuseppe Petracca, and Andrew Martin. A framework for application partitioning using trusted execution environments. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aldinucci:2014:DTB

[ART14] Marco Aldinucci, Salvatore Ruggieri, and Massimo Torquati. Decision tree building on multi-core using FastFlow. Concurrency and Computation: Practice and Experience, 26(3): 800–820, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arzberger:2017:ROE

[Arz17] Peter Arzberger. A reflection on the origins, evolution, and future of PRAGMA. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Sadi:2015:FOF

[AS15] Jehad A. Al-Sadi. Factor-Optical-Factor Exchanges Method: a new load balancing method for Extended Optical Transpose Interconnection System-n-Cube networks. Concurrency and Computation: Practice and Experience, 27(13):3415–3428, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amirjanov:2017:SDA

[AS17] Adil Amirjanov and Konstantin Sobolev. Scheduling of directed acyclic graphs by a genetic algorithm with a repairing mechanism. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Acacio:2019:FSI

[AS19] Manuel E. Acacio and Julio Sahuquillo. Foreword to the special issue on processors, interconnects, storage, and caches for exascale systems. Concurrency and Computation: Practice and Experience, 31(21):e5408:1–e5408:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahri:2019:FAA

[ASB^{+}19] Hassan Qahur Al Mahri, Leonie Simpson, Harry Bartlett, Ed Dawson, and Kenneth Koon-Ho Wong. Fault analysis of AEZ. Concurrency and Computation: Practice and Experience, 31(23):e4785:1–e4785:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aburumman:2017:SRC

[ASE^{+}17] Ala’ Aburumman, Wei Jye Seo, Christian Esposito, Aniello Castiglione, Rafiqul Islam, and Kim-Kwang Raymond Choo. A secure and resilient cross-domain SIP solution for MANETs using dynamic clustering and joint spatial and temporal redundancy. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aydin:2008:BAG

[ASG^{+}08] Galip Aydin, Ahmet Sayar, Harshawardhan Gadgil, Mehmet S. Aktas, Geoffrey C. Fox, Sunghoon Ko, Hasan Bulut, and Marlon E. Pierce. Building and applying geographical information system Grids. Concurrency and Computation: Practice and Experience, 20(14):1653–1695, September 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anand:2019:SPP

[ASP19] Anshu S. Anand, R. K. Shyamasundar, and Sathya Peri. STMs in practice: Partial rollback vs pure abort mechanisms. Concurrency and Computation: Practice and Experience, 31 (5):e4465:1–e4465:??, March 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alford:2005:IIJ

[ASS^{+}05] Torey Alford, Vijay P. Shah, Anthony Skjellum, Nicholas H. Younan, and Clayborne D. Taylor. inAspect: interfacing Java and VSIPL applications. Concurrency and Computation: Practice and Experience, 17(7–8):919–940, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akal:2008:TRG

[ASS08] Fuat Akal, Heiko Schuldt, and Hans-Jörg Schek. Toward replication in grids for digital libraries with freshness and correctness guarantees. Concurrency and Computation: Practice and Experience, 20(17):1981–1993, December 10, 2008. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anand:2019:DFL

[ASS19] Anshu S Anand, Akash Srivastava, and R. K. Shyamasundar. A deadlock-free lock-based synchronization for GPUs. Concurrency and Computation: Practice and Experience, 31(7): e4991:1–e4991:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Saidi:2012:DTM

[ASWR12] Asma Al-Saidi, David W. Walker, and Omer F. Rana. On-demand transmission model for remote visualization using image-based rendering. Concurrency and Computation: Practice and Experience, 24(18):2328–2345, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Achalakul:2001:RTM

[AT01] Tiranee Achalakul and Stephen Taylor. Real-time multispectral image fusion. Concurrency and Computation: Practice and Experience, 13(12):1063–1081, October 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/85512010/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=85512010&PLACEBO=IE.
pdf.

Aupy:2018:PDA

[AT18] Guillaume Aupy and Xueyan Tang. Parallel and distributed algorithms. Concurrency and Computation: Practice and Experience, 30(17):e4663:1–e4663:??, September 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Adhinugraha:2014:FRN

[ATI14] Kiki Maulana Adhinugraha, David Taniar, and Maria Indrawan. Finding reverse nearest neighbors by region. Concurrency and Computation: Practice and Experience, 26(5): 1142–1156, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arkin:2017:SAD

[ATİ17] Ethem Arkin, Bedir Tekinerdogan, and Kayhan M. İmre. Systematic approach for deriving feasible mappings of parallel algorithms to parallel computing platforms. Concurrency and Computation: Practice and Experience, 29(1):??, January 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Tariq:2017:SFP

[ATKH^{+}17] Abdullah Al-Tariq, Abu Raihan Mostofa Kamal, Md. Abdul Hamid, M. Abdullah-Al-Wadud, Mohammad Mehedi Hassan, and Sk Md. Mizanur Rahman. A scalable framework for protecting user identity and access pattern in untrusted Web server using forward secrecy, public key encryption and Bloom filter. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Augonnet:2011:SUP

[ATNW11] Cédric Augonnet, Samuel Thibault, Raymond Namyst, and Pierre-André Wacrenier. StarPU: a unified platform for task scheduling on heterogeneous multicore architectures. Concurrency and Computation: Practice and Experience, 23(2): 187–198, February 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alamri:2015:TMO

[ATSAK15] Sultan Alamri, David Taniar, Maytham Safarb, and Haidar Al-Khalidi. Tracking moving objects using topographical indexing. Concurrency and Computation: Practice and Experience, 27(8):1951–1965, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abbas-Turki:2014:PDG

[ATVLM14] L. A. Abbas-Turki, S. Vialle, B. Lapeyre, and P. Mercier. Pricing derivatives on graphics processing units using Monte Carlo simulation. Concurrency and Computation: Practice and Experience, 26(9):1679–1697, June 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Atre:2019:DSP

[AUHWJ19] Rohit Atre, Zia Ul-Huda, Felix Wolf, and Ali Jannesari. Dissecting sequential programs for parallelization — an approach based on computational units. Concurrency and Computation: Practice and Experience, 31(5):e4770:1–e4770:??, March 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Andersen:2007:DAA

[AV07] Rasmus Andersen and Brian Vinter. Direct application access to Grid storage. Concurrency and Computation: Practice and Experience, 19(9):1287–1298, June 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aldred:2009:DCM

[AvdADtH09] Lachlan Aldred, Wil M. P. van der Aalst, Marlon Dumas, and Arthur H. M. ter Hofstede. Dimensions of coupling in middleware. Concurrency and Computation: Practice and Experience, 21(18):2233–2269, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Andujar:2019:CVD

[AVS^{+}19] Francisco J. Andújar, Juan A. Villar, José L. Sánchez, Francisco J. Alfaro, José Duato, and Holger Fröning. Constructing virtual 5-dimensional tori out of lower-dimensional network cards. Concurrency and Computation: Practice and Experience, 31(2):e4361:1–e4361:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agrawal:2017:ART

[AWR17] Bikash Agrawal, Tomasz Wiktorski, and Chunming Rong. Adaptive real-time anomaly detection in cloud infrastructures. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aktulga:2014:ISS

[AYN^{+}14] Hasan Metin Aktulga, Chao Yang, Esmond G. Ng, Pieter Maris, and James P. Vary. Improving the scalability of a symmetric iterative eigensolver for multi-core platforms. Concurrency and Computation: Practice and Experience, 26(16): 2631–2651, November 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Au:2014:SMV

[AYSZ14] Man Ho Au, Guomin Yang, Willy Susilo, and Yunmei Zhang. (Strong) multidesignated verifiers signatures secure against rogue key attack. Concurrency and Computation: Practice and Experience, 26(8):1574–1592, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Apavatjrut:2012:EEA

[AZF^{+}12] Anya Apavatjrut, Wassim Znaidi, Antoine Fraboulet, Claire Goursaud, Katia Jaffrès-Runser, Cédric Lauradoux, and Marine Minier. Energy efficient authentication strategies for network coding. Concurrency and Computation: Practice and Experience, 24(10):1086–1107, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bader:2004:SIH

[BA04] David A. Bader and Srinivas Aluru. Special issue: High performance computational biology. Concurrency and Computation: Practice and Experience, 16(9):817–821, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baeth:2018:ACP

[BA18] Mohamed Jehad Baeth and Mehmet S. Aktas. An approach to custom privacy policy violation detection problems using big social provenance data. Concurrency and Computation: Practice and Experience, 30(21):e4690:1–e4690:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baeth:2019:DMS

[BA19] Mohamed Jehad Baeth and Mehmet S. Aktas. Detecting misinformation in social networks using provenance data. Concurrency and Computation: Practice and Experience, 31(3): e4793:1–e4793:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bacon:2003:KJD

[Bac03] David F. Bacon. Kava: a Java dialect with a uniform object model for lightweight classes. Concurrency and Computation: Practice and Experience, 15(3–5):185–206, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brookes:2015:GFI

[BAC^{+}15] Emre H. Brookes, Nadeem Anjum, Joseph E. Curtis, Suresh Marru, Raminder Singh, and Marlon Pierce. The GenApp framework integrated with Airavata for managed compute resource submissions. Concurrency and Computation: Practice and Experience, 27(16):4292–4303, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barbera:2011:GPR

[BAD^{+}11] R. Barbera, G. Andronico, G. Donvito, A. Falzone, J. J. Keijser, G. La Rocca, L. Milanesi, G. P. Maggi, and S. Vicario. A grid portal with robot certificates for bioinformatics phylogenetic analyses. Concurrency and Computation: Practice and Experience, 23(3):246–255, March 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bruel:2017:ACC

[BAG17] Pedro Bruel, Marcos Amarís, and Alfredo Goldman. Autotuning CUDA compiler parameters for heterogeneous applications using the OpenTuner framework. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buyya:2002:EMR

[BAGS02] Rajkumar Buyya, David Abramson, Jonathan Giddy, and Heinz Stockinger. Economic models for resource management and scheduling in Grid computing. Concurrency and Computation: Practice and Experience, 14(13–15):1507–1542, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bai:2017:MAI

[Bai17] Yun Bai. Modeling analysis of Intelligent Manufacturing System based on SDN. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bao:2019:EVH

[Bao19] Dapeng Bao. Exploration of virtual human running based on CCD algorithm. Concurrency and Computation: Practice and Experience, 31(10):e4901:1–e4901:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boukerche:2007:TBC

[BAS07] Azzedine Boukerche and Raed A. Al-Shaikh. Towards building a conflict-free mobile distributed file system. Concurrency and Computation: Practice and Experience, 19(8):1237–1250, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brust:2013:MHL

[BAT13] Matthias R. Brust, Mustafa İlhan Akbaş, and Damla Turgut. Multi-hop localization system for environmental monitoring in wireless sensor and actor networks. Concurrency and Computation: Practice and Experience, 25(5):701–717, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Binder:2011:FEP

[BAVM11] Walter Binder, Danilo Ansaloni, Alex Villazón, and Philippe Moret. Flexible and efficient profiling with aspect-oriented programming. Concurrency and Computation: Practice and Experience, 23(15):1749–1773, October 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boukerche:2009:DVE

[BAZ09] Azzedine Boukerche, Abdulaziz Al Hamidi, and Ming Zhang. Design of a virtual environment aided by a model-based formal approach using DEVS. Concurrency and Computation: Practice and Experience, 21(11):1422–1436, August 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brambilla:2018:OWA

[BAZ18] Giacomo Brambilla, Michele Amoretti, and Francesco Zanichelli. An open Web application framework for peer-to-peer location-based services. Concurrency and Computation: Practice and Experience, 30(20):e4254:1–e4254:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benkner:2002:EPP

[BB02] Siegfried Benkner and Thomas Brandes. Efficient parallel programming on scalable shared memory systems with High Performance Fortran. Concurrency and Computation: Practice and Experience, 14(8–9):789–803, July/ August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016125/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016125{\&}PLACEBO=IE.pdf.

Benkner:2004:CDP

[BB04] Siegfried Benkner and Thomas Brandes. Compiling dataparallel programs for clusters of SMPs. Concurrency and Computation: Practice and Experience, 16(2–3):111–132, February/March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beloglazov:2012:OOD

[BB12] Anton Beloglazov and Rajkumar Buyya. Optimal online deterministic algorithms and adaptive heuristics for energy and performance efficient dynamic consolidation of virtual machines in Cloud data centers. Concurrency and Computation: Practice and Experience, 24(13):1397–1420, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bokhari:2013:CCX

[BB13] Shahid H. Bokhari and Saniyah S. Bokhari. A comparison of the Cray XMT and XMT-2. Concurrency and Computation: Practice and Experience, 25(15):2123–2139, October 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beloglazov:2015:ONF

[BB15] Anton Beloglazov and Rajkumar Buyya. OpenStack Neat: a framework for dynamic and energy-efficient consolidation of virtual machines in OpenStack clouds. Concurrency and Computation: Practice and Experience, 27(5):1310–1333, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouksani:2019:RDP

[BB19] Walid Bouksani and Boucif Amar Bensaber. RIN: a dynamic pseudonym change system for privacy in VANET. Concurrency and Computation: Practice and Experience, 31(24): e4719:1–e4719:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Budiardja:2018:ALR

[BBA18] Reuben Budiardja, Timothy Bouvet, and Galen Arnold. Application-level regression testing framework using Jenkins. Concurrency and Computation: Practice and Experience, 30 (1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bosilca:2014:UMA

[BBB^{+}14] George Bosilca, Aurélien Bouteiller, Elisabeth Brunet, Franck Cappello, Jack Dongarra, Amina Guermouche, Thomas Herault, Yves Robert, Frédéric Vivien, and Dounia Zaidouni.

Unified model for assessing checkpointing protocols at extreme-scale. Concurrency and Computation: Practice and Experience, 26(17):2772–2791, December 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boumerzoug:2016:LKM

[BBB16] Hayette Boumerzoug, Boucif Amar Bensaber, and Ismail Biskri. A lightweight key management scheme based on an Adelson-Velskii and Landis tree and elliptic curve cryptography for wireless sensor networks. Concurrency and Computation: Practice and Experience, 28(6):1831–1847, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bai:2016:WLC

[BBC16] Sen Bai, Xin Bai, and Xiangjiu Che. Window-LRFU: a cache replacement policy subsumes the LRU and window-LFU policies. Concurrency and Computation: Practice and Experience, 28(9):2670–2684, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Begel:2002:AVA

[BBCG02] Andrew Begel, Philip Buonadonna, David E. Culler, and David Gay. An analysis of VI Architecture primitives in support of parallel and distributed communication. Concurrency and Computation: Practice and Experience, 14(1):55–76, January 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/91014115/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
91014115&PLACEBO=IE.pdf.

Bouteiller:2010:RML

[BBD10] Aurelien Bouteiller, George Bosilca, and Jack Dongarra. Redesigning the message logging model for high performance. Concurrency and Computation: Practice and Experience, 22 (16):2196–2211, November 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bez:2017:PEE

[BBdS^{+}17] Jean Luca Bez, Eliezer E. Bernart, Fernando F. dos Santos, Lucas Mello Schnorr, and Philippe Olivier Alexandre Navaux. Performance and energy efficiency analysis of HPC physics simulation applications in a cluster of ARM processors. Concurrency and Computation: Practice and Experience, 29(22): ??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bastos:2017:WPB

[BBG17] Bruno F. Bastos, Regina Maria M. Braga, and Antonio Tadeu A. Gomes. WISP: a pattern-based approach to the interchange of scientific workflow specifications. Concurrency and Computation: Practice and Experience, 29(2):??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buyya:2003:VLT

[BBGA03] Rajkumar Buyya, Kim Branson, Jon Giddy, and David Abramson. The Virtual Laboratory: a toolset to enable distributed molecular modelling for drug design on the World-Wide Grid. Concurrency and Computation: Practice and Experience, 15(1):1–25, January 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhatele:2011:OCC

[BBK11] Abhinav Bhatelé, Eric Bohm, and Laxmikant V. Kalé. Optimizing communication for Charm++ applications by reducing network contention. Concurrency and Computation: Practice and Experience, 23(2):211–222, February 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boeres:2019:NAH

[BBM19] Cristina Boeres, Cristiana Bentes, and Edward D. Moreno. New advances in high-performance computing systems. Concurrency and Computation: Practice and Experience, 31(18): e5172:1–e5172:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baldoni:2005:MPS

[BBPV05] R. Baldoni, R. Beraldi, S. Tucci Piergiovanni, and A. Virgillito. On the modelling of publish/subscribe communication systems. Concurrency and Computation: Practice and Experience, 17(12):1471–1495, October 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Burger:2017:MMS

[BBSW17] Michael Burger, Christian Bischof, Christian Schröppel, and Jens Wackerfuß. Methods to model and simulate super carbon nanotubes of higher order. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brown:2019:LMR

[BBW19] Nick Brown, Michael Bareford, and Michèle Weiland. Leveraging MPI RMA to optimize halo-swapping communications in MONC on Cray machines. Concurrency and Computation: Practice and Experience, 31(16):e5008:1–e5008:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bao:2016:LPP

[BC16] Haiyong Bao and Le Chen. A lightweight privacy-preserving scheme with data integrity for smart grid communications. Concurrency and Computation: Practice and Experience, 28 (4):1094–1110, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bekas:2010:ESC

[BCA^{+}10] C. Bekas, A. Curioni, P. Arbenz, C. Flaig, G. H. Van Lenthe, R. Müller, and A. J. Wirth. Extreme scalability challenges in micro-finite element simulations of human bone. Concurrency and Computation: Practice and Experience, 22(16): 2282–2296, November 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brightwell:2005:ASM

[BCC^{+}05] Ron Brightwell, William Camp, Benjamin Cole, Erik DeBenedictis, Robert Leland, James Tomkins, and Arthur B. Maccabe. Architectural specification for massively parallel computers: an experience and measurement-based approach. Concurrency and Computation: Practice and Experience, 17 (10):1271–1316, August 25, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bellettini:2016:DCM

[BCCM16] Carlo Bellettini, Matteo Camilli, Lorenzo Capra, and Mattia Monga. Distributed CTL model checking using MapReduce: theory and practice. Concurrency and Computation: Practice and Experience, 28(11):3025–3041, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barker:2002:DMC

[BCD^{+}02] Kevin Barker, Nikos Chrisochoides, Jeffrey Dobbelaere, Démian Nave, and Keshav Pingali. Date movement and control substrate for parallel adaptive applications. Concurrency and Computation: Practice and Experience, 14(2):77–101, February 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/91513539/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
91513539&PLACEBO=IE.pdf.

Barrett:2010:CVH

[BCD^{+}10] R. F. Barrett, T. H. F. Chan, E. F. D’Azevedo, E. F. Jaeger, K. Wong, and R. Y. Wong. Complex version of high performance computing LINPACK benchmark (HPL). Concurrency and Computation: Practice and Experience, 22(5):573– 587, April 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Belmonte:2006:ERD

[BCdlCT06] M. V. Belmonte, R. Conejo, J. L. Perez de-la Cruz, and F. Triguero. Experiments on robustness and deception in a coalition formation model. Concurrency and Computation: Practice and Experience, 18(4):371–386, April 10, 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bekas:2012:LCD

[BCF12] C. Bekas, A. Curioni, and I. Fedulova. Low-cost data uncertainty quantification. Concurrency and Computation: Practice and Experience, 24(8):908–920, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bokhari:2014:MMM

[BÇG14] Shahid H. Bokhari, Ümit V. Çatalyürek, and Metin N. Gurcan. Massively multithreaded maxflow for image segmentation on the Cray XMT-2. Concurrency and Computation: Practice and Experience, 26(18):2836–2855, December 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barrachina:2009:ECM

[BCI^{+}09] Sergio Barrachina, Maribel Castillo, Francisco D. Igual, Rafael Mayo, Enrique S. Quintana-Ortí, and Gregorio Quintana-Ortí. Exploiting the capabilities of modern GPUs for dense matrix computations. Concurrency and Computation: Practice and Experience, 21(18):2457–2477, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bampis:2018:LBV

[BCI^{+}18] Loukas Bampis, Savvas Chatzichristofis, Chryssanthi Iakovidou, Angelos Amanatiadis, Yiannis Boutalis, and Antonios Gasteratos. A LoCATe-based visual place recognition system for mobile robotics and GPGPUs. Concurrency and Computation: Practice and Experience, 30(7):??, April 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4146.

Baker:2009:AGC

[BCK^{+}09] J. Baker, A. Cunei, T. Kalibera, F. Pizlo, and J. Vitek. Accurate garbage collection in uncooperative environments revisited. Concurrency and Computation: Practice and Experience, 21(12):1572–1606, August 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brebner:2005:MBA

[BCM^{+}05] Paul Brebner, Emmanuel Cecchet, Julie Marguerite, Petr Tůma, Octavian Ciuhandu, Bruno Dufour, Lieven Eeckhout, Stéphane Frénot, Arvind S. Krishna, John Murphy, and Clark Verbrugge. Middleware benchmarking: approaches, results, experiences. Concurrency and Computation: Practice and Experience, 17(15):1799–1805, December 25, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benner:2007:SLS

[BCM^{+}07] Peter Benner, Maribel Castillo, Rafael Mayo, Enrique S. Quintana-Ortí, and Gregorio Quintana-Ortí. Stabilizing large-scale generalized systems on parallel computers using multithreading and message-passing. Concurrency and Computation: Practice and Experience, 19(4):531–542, March 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benoit:2015:NCR

[BCM15] A. Benoit, L.-C. Canon, and L. Marchal. Non-clairvoyant reduction algorithms for heterogeneous platforms. Concurrency and Computation: Practice and Experience, 27(6):1612–1624, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barolli:2015:EAC

[BCX15] Leonard Barolli, Xiaofeng Chen, and Fatos Xhafa. Editorials: Advances on cloud services and cloud computing. Concurrency and Computation: Practice and Experience, 27(8): 1985–1987, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boukerche:2004:PED

[BD04] Azzedine Boukerche and Caron Dzermajko. Performance evaluation of Data Distribution Management strategies. Concurrency and Computation: Practice and Experience, 16(15): 1545–1573, December 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barga:2008:ACE

[BD08] Roger S. Barga and Luciano A. Digiampietri. Automatic capture and efficient storage of e-Science experiment provenance. Concurrency and Computation: Practice and Experience, 20 (5):419–429, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bland:2013:SIP

[BDB^{+}13] Wesley Bland, Peng Du, Aurelien Bouteiller, Thomas Herault, George Bosilca, and Jack J. Dongarra. Special issue papers: Extending the scope of the Checkpoint-on-Failure protocol for forward recovery in standard MPI. Concurrency and Computation: Practice and Experience, 25(17): 2381–2393, December 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benner:2019:GAM

[BDE^{+}19] Peter Benner, Ernesto Dufrechou, Pablo Ezzatti, Alfredo Remón, and Jens Saak. A GPU-aware mixed-precision solver for low-rank algebraic Riccati equations. Concurrency and Computation: Practice and Experience, 31(6):e4462:1– e4462:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bona:2015:DVH

[BDF15] L. C. E. Bona, E. P. Duarte, Jr., and K. V. O. Fonseca. A distributed virtual hypercube algorithm for maintaining scalable and dynamic network overlays. Concurrency and Computation: Practice and Experience, 27(7):1658–1678, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bassi:2008:TNA

[BDG08] Alessandro Bassi, Spyros Denazis, and Pierpaolo Giacomin. Towards a Noah’s ark for the upcoming data deluge. Concurrency and Computation: Practice and Experience, 20(17): 2061–2074, December 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brady:2010:SNR

[BDG^{+}10] Thomas Brady, Jack Dongarra, Michele Guidolin, Alexey Lastovetsky, and Keith Seymour. SmartGridRPC: The new RPC model for high performance Grid computing. Concurrency and Computation: Practice and Experience, 22(18): 2467–2487, December 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buhr:2015:HPT

[BDH15] Peter A. Buhr, David Dice, and Wim H. Hesselink. Highperformance N-thread software solutions for mutual exclusion. Concurrency and Computation: Practice and Experience, 27(3):651–701, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buhr:2016:DME

[BDH16] Peter A. Buhr, David Dice, and Wim H. Hesselink. Dekker’s mutual exclusion algorithm made RW-safe. Concurrency and Computation: Practice and Experience, 28(1):144–165, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buhr:2018:HCM

[BDH18] Peter A. Buhr, David Dice, and Wim H. Hesselink. Highcontention mutual exclusion by elevator algorithms. Concurrency and Computation: Practice and Experience, 30(18): e4475:1–e4475:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Binns:2007:ECA

[BDI^{+}07] Justin Binns, Jonathan DiCarlo, Joseph A. Insley, Ti Leggett, Cory Lueninghoener, John-Paul Navarro, and Michael E. Papka. Enabling community access to TeraGrid visualization resources. Concurrency and Computation: Practice and Experience, 19(6):783–794, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Berglund:2006:SFU

[BdL06] G. Z. M. Berglund and S. W. de Leeuw. A study into the feasibility of using two parallel sparse direct solvers for the Helmholtz equation on Linux clusters. Concurrency and Computation: Practice and Experience, 18(7):749–769, June 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barak:2015:RGA

[BDL^{+}15] Amnon Barak, Zvi Drezner, Ely Levy, Matthias Lieber, and Amnon Shiloh. Resilient gossip algorithms for collecting online management information in exascale clusters. Concurrency and Computation: Practice and Experience, 27(17): 4797–4818, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buravlev:2018:EEL

[BDM18] Vitaly Buravlev, Rocco De Nicola, and Claudio Antares Mezzina. Evaluating the efficiency of Linda implementations. Concurrency and Computation: Practice and Experience, 30(8):??, April 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4381.

Buyya:2005:NID

[BDMM^{+}05] Rajkumar Buyya, Susumu Date, Yuko Mizuno-Matsumoto, Srikumar Venugopal, and David Abramson. Neuroscience instrumentation and distributed analysis of brain activity data: a case for eScience on global Grids. Concurrency and Computation: Practice and Experience, 17(15):1783–1798, December 25, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Browne:2014:COS

[BDP^{+}14] James C. Browne, Robert L. DeLeon, Abani K. Patra, William L. Barth, John Hammond, Matthew D. Jones, Thomas R. Furlani, Barry I. Schneider, Steven M. Gallo, Amin Ghadersohi, Ryan J. Gentner, Jeffrey T. Palmer, Nikolay Simakov, Martins Innus, Andrew E. Bruno, Joseph P. White, Cynthia D. Cornelius, Thomas Yearke, Kyle Marcus, Gregor von Laszewski, and Fugang Wang. Comprehensive, open-source resource usage measurement and analysis for HPC systems. Concurrency and Computation: Practice and Experience, 26(13):2191–2209, September 10, 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bujari:2018:OCA

[BDP18] Armir Bujari, Luigi De Giovanni, and Claudio E. Palazzi. Optimal configuration of active and backup servers for augmented reality cooperative games. Concurrency and Computation: Practice and Experience, 30(20):e4454:1–e4454:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baboulin:2017:SDS

[BDR^{+}17] Marc Baboulin, Jack Dongarra, Adrien Rémy, Stanimire Tomov, and Ichitaro Yamazaki. Solving dense symmetric indefinite systems using GPUs. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bonachea:2001:HPF

[BDT01] Dan Bonachea, Phillip Dickens, and Rajeev Thakur. High-performance file I/O in Java: Existing approaches and bulk I/O extensions. Concurrency and Computation: Practice and Experience, 13(8–9):713–736, July/ August 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/84503223/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
84503223&PLACEBO=IE.pdf.

Bonetti:2013:EVW

[BDTdS13] Daniel R. F. Bonetti, Alexandre C. B. Delbem, Gonzalo Travieso, and Paulo Sergio L. de Souza. Enhanced van der Waals calculations in genetic algorithms for protein structure prediction. Concurrency and Computation: Practice and Experience, 25(15):2170–2186, October 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Briguglio:2002:WDS

[BDV02] Sergio Briguglio, Beniamino Di Martino, and Gregorio Vlad. Workload decomposition strategies for hierarchical distributed-shared memory parallel systems and their implementation with integration of high-level parallel languages. Concurrency and Computation: Practice and Experience, 14(11):933–956, August 25, 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/97519208/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=97519208{\&}PLACEBO=
IE.pdf.

Betz:2014:IEH

[BDW14] Robin M. Betz, Nathan A. DeBardeleben, and Ross C. Walker. An investigation of the effects of hard and soft errors on graphics processing unit-accelerated molecular dynamics simulations. Concurrency and Computation: Practice and Experience, 26(13):2134–2140, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boryczko:2002:PIF

[BDY02] Krzysztof Boryczko, Witold Dzwinel, and David A. Yuen. Parallel implementation of the fluid particle model for simulating complex fluids in the mesoscale. Concurrency and Computation: Practice and Experience, 14(2):137–161, February 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/91513537/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
91513537&PLACEBO=IE.pdf.

Boryczko:2003:CRH

[BDY03] Krzysztof Boryczko, Witold Dzwinel, and David A. Yuen. Clustering revealed in high-resolution simulations and visualization of multi-resolution features in fluid-particle models. Concurrency and Computation: Practice and Experience, 15 (2):101–116, February 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bao:2019:RBM

[BDZ19] Danwen Bao, Zhiwei Di, and Tianxuan Zhang. A reliabilitybased method for optimizing airport collection and distribution network. Concurrency and Computation: Practice and Experience, 31(10):e4733:1–e4733:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beaumont:2018:FAA

[BEDK18] Olivier Beaumont, Lionel Eyraud-Dubois, and Suraj Kumar. Fast approximation algorithms for task-based runtime systems. Concurrency and Computation: Practice and Experience, 30(17):e4502:1–e4502:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benner:2013:MIC

[BEQOR13] Peter Benner, Pablo Ezzatti, Enrique S. Quintana-Ortí, and Alfredo Remón. Matrix inversion on CPU–GPU platforms with applications in control theory. Concurrency and Computation: Practice and Experience, 25(8):1170–1182, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benner:2017:EGH

[BEQOR17] Peter Benner, Pablo Ezzatti, Enrique S. Quintana-Ortí, and Alfredo Remón. Extending the Gauss–Huard method for the solution of Lyapunov matrix equations and matrix inversion. Concurrency and Computation: Practice and Experience, 29 (9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bernholdt:2007:SIC

[Ber07] David E. Bernholdt. Special issue: Component and framework technology in high-performance and scientific computing. Concurrency and Computation: Practice and Experience, 19(5):571–572, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barreto:2007:VVW

[BF07] João Barreto and Paulo Ferreira. Version Vector Weighted Voting protocol: efficient and fault-tolerant commitment for weakly connected replicas. Concurrency and Computation: Practice and Experience, 19(17):2271–2283, December 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boichat:2001:OC

[BFG01] Romain Boichat, Svend Frølund, and Rachid Guerraoui. Open consensus. Concurrency and Computation: Practice and Experience, 13(14):1215–1245, December 10, 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/88511596/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=88511596&PLACEBO=IE.
pdf.

Basney:2014:CFX

[BFG14] Jim Basney, Terry Fleury, and Jeff Gaynor. CILogon: a federated X.509 certification authority for cyberinfrastructure logon. Concurrency and Computation: Practice and Experience, 26(13):2225–2239, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Blanchard:2017:EAM

[BFH17] Frédéric Blanchard, Hacène Fouchal, and Michel Herbin. Energy and activity monitoring over wireless sensor networks. Concurrency and Computation: Practice and Experience, 29 (23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bolten:2017:ADA

[BFK^{+}17] Matthias Bolten, Franz Franchetti, Paul H. J. Kelly, Christian Lengauer, and Marcus Mohr. Algebraic description and automatic generation of multigrid methods in SPIRAL. Concurrency and Computation: Practice and Experience, 29(17): ??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balaji:2010:GSD

[BFL^{+}10] P. Balaji, W. Feng, H. Lin, J. Archuleta, S. Matsuoka, A. Warren, J. Setubal, Ewing Lusk, R. Thakur, I. Foster, D. S. Katz, S. Jha, K. Shinpaugh, S. Coghlan, and D. Reed. Global-scale distributed I/O with ParaMEDIC. Concurrency and Computation: Practice and Experience, 22(16):2266– 2281, November 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boeres:2006:EPS

[BFM^{+}06] C. Boeres, A. A. Fonseca, H. A. Mendes, L. T. Menezes, N. T. Moura, J. A. Silva, B. A. Vianna, and V. E. F. Rebello. An EasyGrid portal for scheduling system-aware applications on computational Grids. Concurrency and Computation: Practice and Experience, 18(6):553–566, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bernaschi:2010:FHP

[BFM^{+}10] Massimo Bernaschi, Massimiliano Fatica, Simone Melchionna, Sauro Succi, and Efthimios Kaxiras. A flexible high-performance Lattice Boltzmann GPU code for the simulations of fluid flows in complex geometries. Concurrency and Computation: Practice and Experience, 22(1):1–14, January 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baraglia:2005:SMH

[BFR05] Ranieri Baraglia, Renato Ferrini, and Pierluigi Ritrovato. A static mapping heuristics to map parallel applications to heterogeneous computing systems. Concurrency and Computation: Practice and Experience, 17(13):1579–1605, November 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See erratum [Ano06].

Biberstein:2007:CAA

[BFU07] Marina Biberstein, Eitan Farchi, and Shmuel Ur. Choosing among alternative pasts. Concurrency and Computation:

Practice and Experience, 19(3):341–353, March 10, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boton-Fernandez:2015:ISA

[BFVRC15] María Botón-Fernández, Miguel A. Vega-Rodríguez, and Francisco Prieto Castrillo. Intelligent self-adaptive resources selection for grid applications. Concurrency and Computation: Practice and Experience, 27(14):3539–3560, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brebner:2004:JEB

[BG04] Paul Brebner and Jeffrey Gosper. The J2EE ECperf benchmark results: transient trophies or technology treasures? Concurrency and Computation: Practice and Experience, 16 (10):1023–1036, August 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bartolini:2014:EES

[BG14] Sandro Bartolini and José M. García. Editorial: Exploiting silicon photonics for energy-efficient heterogeneous parallel architectures. Concurrency and Computation: Practice and Experience, 26(15):2489–2491, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bellas:2017:GPT

[BG17] Christos Bellas and Anastasios Gounaris. GPU processing of theta-joins. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barcellos:2011:CSP

[BGdCCA11] Marinho Pilla Barcellos, Luciano Paschoal Gaspary, Weverton Luis da Costa Cordeiro, and Rodolfo Stoffel Antunes. A conservative strategy to protect P2P file sharing systems from pollution attacks. Concurrency and Computation: Practice and Experience, 23(1):117–141, January 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baboulin:2007:DPS

[BGGL07] Marc Baboulin, Luc Giraud, Serge Gratton, and Julien Langou. A distributed packed storage for large dense parallel in-core calculations. Concurrency and Computation: Practice and Experience, 19(4):483–502, March 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balakrishnan:2014:PCS

[BGGS14] Narayanaswamy Balakrishnan, Jakadeesan Gopinatha, Dhrubajyoti Goswami, and Latha Shanker. Parallel computing strategies in the analysis of the inhibiting effect of price limits on futures prices. Concurrency and Computation: Practice and Experience, 26(9):1666–1678, June 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bertino:2003:SOM

[BGM03] Elisa Bertino, Giovanna Guerrini, and Isabella Merlo. A setoriented method definition language for object databases and its semantics. Concurrency and Computation: Practice and Experience, 15(14):1275–1335, December 10, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Banicescu:2001:EIA

[BGV^{+}01] Ioana Banicescu, Sheikh Ghafoor, Vijay Velusamy, Samuel H. Russ, and Mark Bilderback. Experiences from integrating algorithmic and systemic load balancing strategies. Concurrency and Computation: Practice and Experience, 13(2): 121–139, February 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/77004419/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=77004419&PLACEBO=IE.pdf.

Buhr:2005:CUL

[BH05] Peter A. Buhr and Ashif S. Harji. Concurrent urban legends. Concurrency and Computation: Practice and Experience, 17 (9):1133–1172, August 10, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Babik:2009:ODL

[BH09] Marian Babík and Ladislav Hluchý. Optimizing description logic reasoning for the large-scale semantic repositories. Concurrency and Computation: Practice and Experience, 21(5): 635–650, April 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balaji:2016:EPM

[BH16] Pavan Balaji and Zhiyi Huang. Editorials: Programming models and applications for multicores and manycores. Concurrency and Computation: Practice and Experience, 28(2): 453–454, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Babamir:2015:PPD

[BHA15a] Seyed Morteza Babamir, Elmira Hassanzade, and Mona Azimpour. Predicting potential deadlocks in multithreaded programs. Concurrency and Computation: Practice and Experience, 27(17):5261–5287, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brown:2015:FGF

[BHA^{+}15b] Kevin Brown, Michael Hayes, David Allison, Miriam A. M. Capretz, Margaret Sazio, and Rupinder Mann. Fine-grained filtering to provide access control for data providing services within collaborative environments. Concurrency and Computation: Practice and Experience, 27(6):1445–1466, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouteiller:2013:CSC

[BHBD13] Aurelien Bouteiller, Thomas Herault, George Bosilca, and Jack J. Dongarra. Correlated set coordination in fault tolerant message logging protocols for many-core clusters. Concurrency and Computation: Practice and Experience, 25(4): 572–585, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barrett:2013:SIP

[BHD13] Enda Barrett, Enda Howley, and Jim Duggan. Special issue papers: Applying reinforcement learning towards automating resource allocation and application scalability in the cloud. Concurrency and Computation: Practice and Experience, 25 (12):1656–1674, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Biberstein:2009:CSC

[BHH09] M. Biberstein, Y. Harel, and A. Heilper. Clock synchronization in Cell/B.E. traces. Concurrency and Computation:

Practice and Experience, 21(14):1760–1774, September 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Behrends:2016:HGE

[BHJ^{+}16] Reimer Behrends, Kevin Hammond, Vladimir Janjic, Alexander Konovalov, Steve Linton, Hans-Wolfgang Loidl, Patrick Maier, and Phil Trinder. HPC-GAP: engineering a 21stcentury high-performance computer algebra system. Concurrency and Computation: Practice and Experience, 28(13): 3606–3636, September 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buchty:2012:SHA

[BHKW12] Rainer Buchty, Vincent Heuveline, Wolfgang Karl, and Jan-Philipp Weiss. A survey on hardware-aware and heterogeneous computing on multicore processors and accelerators. Concurrency and Computation: Practice and Experience, 24 (7):663–675, May 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Badia:2009:PDB

[BHL^{+}09] Rosa M. Badia, José R. Herrero, Jesús Labarta, Josep M. Pérez, Enrique S. Quintana-Ortí, and Gregorio Quintana-Ortí. Parallelizing dense and banded linear algebra libraries using SMPSs. Concurrency and Computation: Practice and Experience, 21(18):2438–2456, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barhen:2012:CFC

[BHM^{+}12] J. Barhen, T. Humble, P. Mitra, N. Imam, B. Schleck, C. Kotas, and M. Traweek. Concurrent FFT computing on multicore processors. Concurrency and Computation: Practice and Experience, 24(1):29–44, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bungartz:2014:PBS

[BHPS14] H.-J. Bungartz, A. Heinecke, D. Pflüger, and S. Schraufstetter. Parallelizing a Black–Scholes solver based on finite elements and sparse grids. Concurrency and Computation: Practice and Experience, 26(9):1640–1653, June 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bientinesi:2015:EPC

[BHQOS15] Paolo Bientinesi, José R. Herrero, Enrique S. Quintana-Ortí, and Robert Strzodka. Editorials: Parallel computing on graphics processing units and heterogeneous platforms. Concurrency and Computation: Practice and Experience, 27(6): 1525–1527, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bishop:2005:EIJ

[BHW05] Judith Bishop, R. Nigel Horspool, and Basil Worrall. Experience in integrating Java with C# and .NET. Concurrency and Computation: Practice and Experience, 17(5–6):663–680, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bientinesi:2011:CFS

[BIK^{+}11] Paolo Bientinesi, Francisco D. Igual, Daniel Kressner, Matthias Petschow, and Enrique S. Quintana-Ortí. Condensed forms for the symmetric eigenvalue problem on multithreaded architectures. Concurrency and Computation: Practice and Experience, 23(7):694–707, May 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bank:2001:NPD

[BJ01] Randolph E. Bank and Peter K. Jimack. A new parallel domain decomposition method for the adaptive finite element solution of elliptic partial differential equations. Concurrency and Computation: Practice and Experience, 13(5):327–350, April 25, 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/78505738/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
78505738&PLACEBO=IE.pdf.

Barbosa:2018:TSI

[BJ18] Jorge G. Barbosa and Emmanuel Jeannot. The Twenty Sixth International Heterogeneity in Computing Workshop (HCW) and to the Fifteenth International Workshop on Algorithms, Models and Tools for Parallel Computing on Heterogeneous Platforms (HeteroPar). Concurrency and Computation: Practice and Experience, 30(23):e5007:1–e5007:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bui:2017:GTA

[BJC17] Khac-Hoai Nam Bui, Jai E. Jung, and David Camacho. Game theoretic approach on Real-time decision making for IoTbased traffic light control. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bahrami:2019:HKP

[BJD^{+}19] Pooneh Nikkhah Bahrami, Hamid H. S. Javadi, Tooska Dargahi, Ali Dehghantanha, and Kim-Kwang Raymond Choo. A hierarchical key pre-distribution scheme for fog networks. Concurrency and Computation: Practice and Experience, 31(22):e4776:1–e4776:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Budimlic:2005:CAW

[BK05] Zoran Budimlić and Ken Kennedy. Compiling almost-whole Java programs. Concurrency and Computation: Practice and Experience, 17(5–6):573–587, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balamurugan:2019:RRP

[BKA19] Kavitha Balamurugan, Chitra K, and Jawahar A. Reconfigurable routing protocol with optical sphere in FSO MANET. Concurrency and Computation: Practice and Experience, 31 (14):e4874:1–e4874:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beyler:2009:MPF

[BKCP09] Jean Christophe Beyler, Michael Klemm, Philippe Clauss, and Michael Philippsen. A meta-predictor framework for prefetching in object-based DSMs. Concurrency and Computation: Practice and Experience, 21(14):1789–1803, September 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Binzenhofer:2008:DAS

[BKH08] Andreas Binzenhöfer, Gerald Kunzmann, and Robert Henjes. Design and analysis of a scalable algorithm to monitor chord-based P2P systems at runtime. Concurrency and Computation: Practice and Experience, 20(6):625–641, April 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brown:2009:ASC

[BKLH09] Sheldon Brown, Kristen Kho, Kwangyoon Lee, and Erik Hill. Accelerating the scalable city. Concurrency and Computation: Practice and Experience, 21(17):2187–2198, December 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bogdanski:2007:FUF

[BKM^{+}07a] Maciej Bogdanski, Michal Kosiedowski, Cezary Mazurek, Marzena Rabiega, and Małgorzata Wolniewicz. Flexibility and user-friendliness of Grid portals: the PROGRESS approach. Concurrency and Computation: Practice and Experience, 19(6):827–838, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bogdanski:2007:PGC

[BKM^{+}07b] Maciej Bogdanski, Michal Kosiedowski, Cezary Mazurek, Marzena Rabiega, and Małgorzata Wolniewicz. Plugging grids into computing portals: the PROGRESS grid resource broker plug-in mechanism. Concurrency and Computation: Practice and Experience, 19(12):1653–1661, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boito:2016:ASA

[BKND16] Francieli Zanon Boito, Rodrigo Virote Kassick, Philippe O. A. Navaux, and Yves Denneulin. Automatic I/O scheduling algorithm selection for parallel file systems. Concurrency and Computation: Practice and Experience, 28(8):2457–2472, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bharti:2018:CAS

[BKS18] Monika Bharti, Rajesh Kumar, and Sharad Saxena. Contextaware search optimization framework on the Internet of Things. Concurrency and Computation: Practice and Experience, 30(14):??, July 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4426.

Bleuse:2015:SIT

[BKSM^{+}15] Raphael Bleuse, Safia Kedad-Sidhoum, Florence Monna, Grégory Mounié, and Denis Trystram. Scheduling independent tasks on multi-cores with GPU accelerators. Concurrency and Computation: Practice and Experience, 27(6): 1625–1638, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Belloum:2019:BDO

[BKWM19] Adam S. Z. Belloum, Spiros Koulouzis, Tomasz Wiktorski, and Andrea Manieri. Bridging the demand and the offer in data science. Concurrency and Computation: Practice and Experience, 31(17):e5200:1–e5200:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bilal:2013:SIP

[BKZ^{+}13] Kashif Bilal, Samee U. Khan, Limin Zhang, Hongxiang Li, Khizar Hayat, Sajjad A. Madani, Nasro Min-Allah, Lizhe Wang, Dan Chen, Majid Iqbal, Cheng-Zhong Xu, and Albert Y. Zomaya. Special issue papers: Quantitative comparisons of the state-of-the-art data center architectures. Concurrency and Computation: Practice and Experience, 25(12): 1771–1783, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Burrows:2004:FSV

[BL04] Michael Burrows and K. Rustan M. Leino. Finding stalevalue errors in concurrent programs. Concurrency and Computation: Practice and Experience, 16(12):1161–1172, October 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouge:2009:SIEa

[BL09a] Luc Bougé and Christian Lengauer. Special issue: Euro-Par 2007. Concurrency and Computation: Practice and Experience, 21(7):855–857, May 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouge:2009:SIEb

[BL09b] Luc Bougé and Christian Lengauer. Special issue: Euro-Par 2008. Concurrency and Computation: Practice and Experience, 21(14):1757–1759, September 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouge:2011:SIEa

[BL11a] Luc Bougé and Christian Lengauer. Special issue: Euro-Par 2009. Concurrency and Computation: Practice and Experience, 23(2):143–144, February 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouge:2011:SIEb

[BL11b] Luc Bougé and Christian Lengauer. Special issue: Euro-Par 2010. Concurrency and Computation: Practice and Experience, 23(17):2137–2139, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouge:2013:ESIb

[BL13a] Luc Bougé and Christian Lengauer. Editorial: Special issue: Euro-Par 2012. Concurrency and Computation: Practice and Experience, 25(17):2343–2344, December 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouge:2013:ESIa

[BL13b] Luc Bougé and Christian Lengauer. Editorials: Special issue: Euro-Par 2011. Concurrency and Computation: Practice and Experience, 25(4):510–512, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balaji:2017:EFS

[BL17] Pavan Balaji and Kai-Cheung Leung. Editorial: Foreword to the special issue of the Workshop on the Seventh International Workshop on Programming Models and Applications for Multicores and Manycores (PMAM 2016). Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Blanton:2014:JFV

[BLA^{+}14] Ethan Blanton, Demian Lessa, Puneet Arora, Lukasz Ziarek, and Bharat Jayaraman. JI.FI: Visual test and debug queries for hard real-time. Concurrency and Computation: Practice and Experience, 26(14):2456–2487, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ban:2019:PED

[BLCC19] Xinbo Ban, Shigang Liu, Chao Chen, and Caslon Chua. A performance evaluation of deep-learnt features for software vulnerability detection. Concurrency and Computation: Practice and Experience, 31(19):e5103:1–e5103:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brugger:2016:PTH

[BLDW16] Christian Brugger, Gongda Liu, Christian De Schryver, and Norbert Wehn. Precision-tuning and hybrid pricer for closed-form solution-based Heston calibration. Concurrency and Computation: Practice and Experience, 28(3):892–904, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buttari:2008:PTQ

[BLKD08] Alfredo Buttari, Julien Langou, Jakub Kurzak, and Jack Dongarra. Parallel tiled QR factorization for multicore architectures. Concurrency and Computation: Practice and Experience, 20(13):1573–1590, September 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Berenyi:2018:TSP

[BLL18] Dániel Berényi, András Leitereg, and Gábor Lehel. Towards scalable pattern-based optimization for dense linear algebra. Concurrency and Computation: Practice and Experience, 30(22):e4696:1–e4696:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bewong:2019:RPM

[BLL^{+}19] Michael Bewong, Jixue Liu, Lin Liu, Jiuyong Li, and Kim-Kwang Raymond Choo. A relative privacy model for effective privacy preservation in transactional data. Concurrency and Computation: Practice and Experience, 31(23):e4923:1– e4923:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bruneo:2011:PAJ

[BLSP11] Dario Bruneo, Francesco Longo, Marco Scarpa, and Antonio Puliafito. Performance analysis of job dissemination techniques in Grid systems. Concurrency and Computation:

Practice and Experience, 23(11):1213–1235, August 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bellatreche:2019:ACB

[BLXE19] Ladjel Bellatreche, Carson Leung, Yinglong Xia, and Didier El Baz. Advances in cloud and big data computing. Concurrency and Computation: Practice and Experience, 31 (2):e5053:1–e5053:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buyya:2002:GTM

[BM02] Rajkumar Buyya and Manzur Murshed. GridSim: a toolkit for the modeling and simulation of distributed resource management and scheduling for Grid computing. Concurrency and Computation: Practice and Experience, 14(13–15):1175– 1220, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bella:2004:SIC

[BM04] Giampaolo Bella and Ronaldo Menezes. Special issue: Computer security. Concurrency and Computation: Practice and Experience, 16(11):1061, September 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bernat:2007:ICP

[BM07] Andrew R. Bernat and Barton P. Miller. Incremental callpath profiling. Concurrency and Computation: Practice and Experience, 19(11):1533–1547, August 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bittencourt:2008:POA

[BM08] Luiz F. Bittencourt and Edmundo R. M. Madeira. A performance-oriented adaptive scheduler for dependent tasks on grids. Concurrency and Computation: Practice and Experience, 20(9):1029–1049, June 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bartlang:2010:FCR

[BM10] Udo Bartlang and Jörg P. Müller. A flexible content repository to enable a peer-to-peer-based wiki. Concurrency and Computation: Practice and Experience, 22(7):831–871, May 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Badica:2012:PSI

[BM12] Costin Bādicā and Ronaldo Menezes. Preface to the Special Issue on Intelligent Distributed Computing. Concurrency and Computation: Practice and Experience, 24(6):555–557, May 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bellatreche:2016:EBD

[BM16] Ladjel Bellatreche and Mukesh Mohania. Editorials: Big data analytics and knowledge discovery. Concurrency and Computation: Practice and Experience, 28(15):3945–3947, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhatia:2003:SCM

[BMA03] Karan Bhatia, Keith Marzullo, and Lorenzo Alvisi. Scalable causal message logging for wide-area environments. Concurrency and Computation: Practice and Experience, 15(10): 873–889, August 25, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bowers:2008:PCO

[BML08] Shawn Bowers, Timothy M. McPhillips, and Bertram Ludäscher. Provenance in collection-oriented scientific workflows. Concurrency and Computation: Practice and Experience, 20(5):519–529, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Budgaga:2017:FSR

[BMPP17] Walid Budgaga, Matthew Malensek, Sangmi Lee Pallickara, and Shrideep Pallickara. A framework for scalable real-time anomaly detection over voluminous, geospatial data streams. Concurrency and Computation: Practice and Experience, 29 (12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brooke:2007:ILV

[BMPS07] J. M. Brooke, J. Marsh, S. Pettifer, and L. S. Sastry. The importance of locality in the visualization of large datasets.

Concurrency and Computation: Practice and Experience, 19 (2):195–205, February 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Binder:2009:MCC

[BMS^{+}09] Walter Binder, Adina Mosincat, Samuel Spycher, Ion Constantinescu, and Boi Faltings. Multiversion concurrency control for the generalized search tree. Concurrency and Computation: Practice and Experience, 21(12):1547–1571, August 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baldoni:2003:CRP

[BMV03] R. Baldoni, C. Marchetti, and L. Verde. CORBA request portable interceptors: analysis and applications. Concurrency and Computation: Practice and Experience, 15(6):551–579, May 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhuvaneswari:2019:IEE

[BN19] P. Bhuvaneswari and L. Nithyanandan. Improving energy efficiency in LTE-A networks with the reduction of failure rate in eNB components. Concurrency and Computation: Practice and Experience, 31(14):e4916:1–e4916:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baskaran:2019:TEL

[BNNH19] Annie Gilda Roselin Arockia Baskaran, Priyadarsi Nanda, Surya Nepal, and Sean He. Testbed evaluation of lightweight authentication protocol (LAUP) for 6LoWPAN wireless sensor networks. Concurrency and Computation: Practice and Experience, 31(23):e4868:1–e4868:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ben-Othman:2013:IHN

[BOB13] Jalel Ben-Othman and Yesica I. Saavedra Benitez. IBC-HWMP: a novel secure identity-based cryptography-based scheme for Hybrid Wireless Mesh Protocol for IEEE 802.11s. Concurrency and Computation: Practice and Experience, 25 (5):686–700, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bochen:2019:RIP

[Boc19] Zhang Bochen. Research on the integration of photographic images and photography art based on 3D virtual reality technology. Concurrency and Computation: Practice and Experience, 31(10):e4749:1–e4749:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boehmer:2012:TTC

[Boe12] Wolfgang Boehmer. Toward a target and coupling function of three different Information Security Management Systems. Concurrency and Computation: Practice and Experience, 24 (15):1708–1725, October 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balkir:2015:EGD

[BOF15] Atilla Soner Balkir, Huseyin Oktay, and Ian Foster. Estimating graph distance and centrality on shared nothing architectures. Concurrency and Computation: Practice and Experience, 27(14):3587–3613, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bokhari:2012:PSS

[Bok12] Saniyah S. Bokhari. Parallel solution of the subset-sum problem: an empirical study. Concurrency and Computation: Practice and Experience, 24(18):2241–2254, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boussinot:2006:FMC

[Bou06] Frédéric Boussinot. FairThreads: mixing cooperative and preemptive threads in C. Concurrency and Computation: Practice and Experience, 18(5):445–469, April 25, 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bourne:2013:RLB

[Bou13] Philip E. Bourne. The reaming of life: based on the 2010 Jim Gray eScience Award Lecture. Concurrency and Computation: Practice and Experience, 25(4):445–453, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Breg:2003:JVM

[BP03] Fabian Breg and Constantine D. Polychronopoulos. Java Virtual Machine support for object serialization. Concurrency and Computation: Practice and Experience, 15(3–5): 263–275, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balaiah:2017:EGM

[BP17] Thanasekhar Balaiah and Ranjani Parthasarathi. Exploiting GPU memory hierarchy for accelerating a specialized stencil computation. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brandic:2008:SPE

[BPB08] Ivona Brandic, Sabri Pllana, and Siegfried Benkner. Specification, planning, and execution of QoS-aware Grid workflows within the Amadeus environment. Concurrency and Computation: Practice and Experience, 20(4):331–345, March 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buis:2006:PCF

[BPD06] Samuel Buis, Andrea Piacentini, and Damien Déclat. PALM: a computational framework for assembling high-performance computing applications. Concurrency and Computation: Practice and Experience, 18(2):231–245, February 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bartoli:2006:RFP

[BPdM06] Alberto Bartoli, Milan Prica, and Etienne Antoniutti di Muro. A replication framework for program-to-program interaction across unreliable networks and its implementation in a servlet container. Concurrency and Computation: Practice and Experience, 18(7):701–724, June 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benitez:2012:PHC

[BPL12] César Manuel Vargas Benitez, Rafael Stubs Parpinelli, and Heitor Silvério Lopes. Parallelism, hybridism and coevolution in a multi-level ABC-GA approach for the protein structure prediction problem. Concurrency and Computation: Practice and Experience, 24(6):635–646, May 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Belgin:2019:DDS

[BPL^{+}19] Mehmet Belgin, Tyler A. Perini, Fang (Cherry) Liu, Nuyun Zhang, Semir Sarajlic, Andre McNeill, Paul Manno, and Neil C. Bright. A data-driven support strategy for a sustainable research software repository. Concurrency and Computation: Practice and Experience, 31(20):e5338:1–e5338:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Braganca:2019:RDM

[BPS19] Ricardo Bragança, Filipe Portela, and Manuel Santos. A regression data mining approach in lean production. Concurrency and Computation: Practice and Experience, 31(22): e4449:1–e4449:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Breslow:2016:CCH

[BPT^{+}16] Alex D. Breslow, Leo Porter, Ananta Tiwari, Michael Laurenzano, Laura Carrington, Dean M. Tullsen, and Allan E. Snavely. The case for colocation of high performance computing workloads. Concurrency and Computation: Practice and Experience, 28(2):232–251, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boeres:2004:ETF

[BR04] Cristina Boeres and Vinod E. F. Rebello. EasyGrid: towards a framework for the automatic Grid enabling of legacy MPI applications. Concurrency and Computation: Practice and Experience, 16(5):425–432, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barnes:2010:POD

[BR10] F. R. M. Barnes and C. G. Ritson. Process-oriented device driver development. Concurrency and Computation: Practice and Experience, 22(8):995–1006, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Burke:2016:CPM

[BRCV16] Neil Burke, Andrew Rau-Chaplin, and Blesson Varghese. Computing probable maximum loss in catastrophe reinsurance portfolios on multi-core and many-core architectures. Concurrency and Computation: Practice and Experience, 28 (3):836–847, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brinkschulte:2016:ADS

[Bri16] Uwe Brinkschulte. An artificial DNA for self-descripting and self-building embedded real-time systems. Concurrency and Computation: Practice and Experience, 28(14):3711–3729, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bachniak:2017:SAH

[BRK^{+}17] Daniel Bachniak, Lukasz Rauch, Dariusz Król, Jakub Liput, Renata Słota, Jacek Kitowski, and Maciej Pietrzyk. Sensitivity analysis on HPC systems with Scalarm platform. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bokhari:2006:EAM

[BRWB06] Shahid Bokhari, Benjamin Rutt, Pete Wyckoff, and Paul Buerger. Experimental analysis of a mass storage system. Concurrency and Computation: Practice and Experience, 18 (15):1929–1950, December 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bokhari:2004:SAC

[BS04] Shahid H. Bokhari and Jon R. Sauer. Sequence alignment on the Cray MTA-2. Concurrency and Computation: Practice and Experience, 16(9):823–839, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bokhari:2010:EPM

[BS10] Shahid Bokhari and Joel Saltz. Exploring the performance of massively multithreaded architectures. Concurrency and Computation: Practice and Experience, 22(5):588–616, April 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bull:2003:BJA

[BSB^{+}03] J. M. Bull, L. A. Smith, C. Ball, L. Pottage, and R. Freeman. Benchmarking Java against C and Fortran for scientific applications. Concurrency and Computation: Practice and Experience, 15(3–5):417–430, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Becciani:2015:SGT

[BSC^{+}15] Ugo Becciani, Eva Sciacca, Alessandro Costa, Piero Massimino, Costantino Pistagna, Simone Riggi, Fabio Vitello, Catia Petta, Marilena Bandieramonte, and Mel Krokos. Science gateway technologies for the astrophysics community. Concurrency and Computation: Practice and Experience, 27(2): 306–327, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhuiyan:2011:POF

[BSP11] Mohammad A. Bhuiyan, Melissa C. Smith, and Vivek K. Pallipuram. Performance, optimization, and fitness: Connecting applications to architectures. Concurrency and Computation: Practice and Experience, 23(10):1066–1100, July 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brown:2009:GEF

[BSZ09] A. Brown, M. Swany, and J. Zurawski. A general encoding framework for representing network measurement and topology data. Concurrency and Computation: Practice and Experience, 21(8):1069–1086, June 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boukerche:2004:RME

[BT04] Azzedine Boukerche and Terry Tuck. RF-MVTC: an efficient risk-free multiversion concurrency control algorithm. Concurrency and Computation: Practice and Experience, 16(13): 1291–1311, November 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Braeken:2018:AAA

[BT18] A. Braeken and Abdellah Touhafi. AAA — autonomous anonymous user authentication and its application in V2G. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4303.

Behadada:2016:BDB

[BTCB16] Omar Behadada, Marcello Trovati, Ma Chikh, and Nik Bessis. Big data-based extraction of fuzzy partition rules for heart arrhythmia detection: a semi-automated approach. Concurrency and Computation: Practice and Experience, 28 (2):360–373, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balouek-Thomert:2017:NCB

[BTCGL17] Daniel Balouek-Thomert, Eddy Caron, Pascal Gallard, and Laurent Lefèvre. Nu@ge: a container-based cloud computing service federation. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bik:2006:MVF

[BTG06] Aart J. C. Bik, Xinmin Tian, and Milind B. Girkar. Multimedia vectorization of floating-point MIN/MAX reductions. Concurrency and Computation: Practice and Experience, 18 (9):997–1007, August 10, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brightwell:2010:PER

[BUVS10] Ron Brightwell, Keith D. Underwood, Courtenay Vaughan, and Joel Stevenson. Performance evaluation of the Red Storm dual-core upgrade. Concurrency and Computation: Practice and Experience, 22(2):175–190, February 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bertolli:2011:FTD

[BV11] C. Bertolli and M. Vanneschi. Fault tolerance for data parallel programs. Concurrency and Computation: Practice and Experience, 23(6):595–632, April 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baranwal:2016:CSS

[BV16] Gaurav Baranwal and Deo Prakash Vidyarthi. A cloud service selection model using improved ranked voting method. Concurrency and Computation: Practice and Experience, 28(13):

3540–3567, September 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Basanta-Val:2011:ECM

[BVGVEA11] P. Basanta-Val, M. García-Valls, and I. Estévez-Ayres. Extending the concurrency model of the real-time specification for Java. Concurrency and Computation: Practice and Experience, 23(14):1623–1645, September 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Basanta-Val:2011:FTM

[BVGVEAFG11] Pablo Basanta-Val, Marisol García-Valls, Iria Estévez-Ayres, and Jorge Fernández-González. Fine tuning of the multiplexing facilities of Java’s Remote Method Invocation. Concurrency and Computation: Practice and Experience, 23(11): 1236–1260, August 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benito:2019:NMA

[BVIB19] Mariano Benito, Enrique Vallejo, Cruz Izu, and Ramón Beivide. Non-minimal adaptive routing based on explicit congestion notifications. Concurrency and Computation: Practice and Experience, 31(2):e4440:1–e4440:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beran:2010:SPG

[BvIF10] Bora Beran, Catharine van Ingen, and Dennis Robert Fatland. SciScope: a participatory geoscientific Web application. Concurrency and Computation: Practice and Experience, 22 (17):2300–2312, December 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Berljafa:2015:OSE

[BWD15] Mario Berljafa, Daniel Wortmann, and Edoardo Di Napoli. An optimized and scalable eigensolver for sequences of eigenvalue problems. Concurrency and Computation: Practice and Experience, 27(4):905–922, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bruseke:2014:PPB

[BWEB14] Frank Brüseke, Henning Wachsmuth, Gregor Engels, and Steffen Becker. PBlaman: performance blame analysis based on Palladio contracts. Concurrency and Computation: Practice and Experience, 26(12):1975–2004, August 25, 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brown:2018:SDA

[BWHS18] Nick Brown, Michèle Weiland, Adrian Hill, and Ben Shipway. In situ data analytics for highly scalable cloud modelling on Cray machines. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bruin:2008:JSG

[BWW^{+}08] R. P. Bruin, T. O. H. White, A. M. Walker, K. F. Austen, M. T. Dove, R. P. Tyer, P. A. Couch, I. T. Todorov, and M. O. Blanchard. Job submission to grid computing environments. Concurrency and Computation: Practice and Experience, 20 (11):1329–1340, August 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bi:2016:ARD

[BXLJ16] Zhongqin Bi, Feifei Xu, Jingsheng Lei, and Teng Jiang. Attribute reduction in decision-theoretic rough set model based on minimum decision cost. Concurrency and Computation: Practice and Experience, 28(15):4125–4143, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bu:2017:EHP

[BXQ17] Kai Bu, Bin Xiao, and Yi Qian. Editorial: High performance and security in cloud computing. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Blattner:2012:PSC

[BY12] Timothy Blattner and Shiming Yang. Performance study on CUDA GPUs for parallelizing the local ensemble transformed Kalman filter algorithm. Concurrency and Computation: Practice and Experience, 24(2):167–177, February 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bai:2019:HAF

[BYDC19] Xu Bai, Jiajia Yang, Qiong Dai, and Zhaolin Chen. A hybrid ARM-FPGA cluster for cryptographic algorithm acceleration. Concurrency and Computation: Practice and Experience, 31(24):e5257:1–e5257:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Badawy:2017:OTC

[BYN^{+}17] Abdel-Hameed A. Badawy, Gabriel Yessin, Vikram Narayana, David Mayhew, and Tarek El-Ghazawi. Optimizing thin client caches for mobile cloud computing: Design space exploration using genetic algorithms. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bui:2012:SDS

[BYT^{+}12] Peter Bui, Li Yu, Andrew Thrasher, Rory Carmichael, Irena Lanc, Patrick Donnelly, and Douglas Thain. Scripting distributed scientific workflows using Weaver. Concurrency and Computation: Practice and Experience, 24(15):1685–1707, October 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bricteux:2017:VSO

[BZB17] L. Bricteux, S. Zeoli, and N. Bourgeois. Validation and scalability of an open source parallel flow solver. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bo:2016:ETK

[BZD16] Yang Bo, Mingwu Zhang, and Jun-Qiang Du. An errortolerant keyword search scheme based on public-key encryption in secure cloud computing. Concurrency and Computation: Practice and Experience, 28(4):1083–1093, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Branco:2010:MVL

[BZdR^{+}10] Miguel Branco, Ed Zaluska, David de Roure, Mario Lassnig, and Vincent Garonne. Managing very large distributed data sets on a data grid. Concurrency and Computation: Practice and Experience, 22(11):1338–1364, August 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2006:SRA

[CA06] Liang Chen and Gagan Agrawal. A static resource allocation framework for Grid-based streaming applications. Concurrency and Computation: Practice and Experience, 18(6):653– 666, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carter:2008:EUC

[CAC^{+}08] Donna M. Carter, Ramazan Aygun, Glenn Cox, Mary Ellen Weisskopf, and Letha Etzkorn. The effect of uncontrolled concurrency on model checking. Concurrency and Computation: Practice and Experience, 20(12):1419–1438, August 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Capodieci:2015:CAD

[CAC15] Nicola Capodieci, Emanuel Federico Alsina, and Giacomo Cabri. Context-awareness in the deregulated electric energy market: an agent-based approach. Concurrency and Computation: Practice and Experience, 27(6):1513–1524, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carrion:2011:PIF

[CACC11] I. Marín Carrión, E. Arias Antúnez, M. M. Artigao Castillo, and J. J. Miralles Canals. Parallel implementations of the False Nearest Neighbors method for distributed memory architectures. Concurrency and Computation: Practice and Experience, 23(1):1–16, January 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chakroun:2018:HLL

[CAD^{+}18] Imen Chakroun, Tom Vander Aa, Bruno De Fraine, Tom Haber, Pascal Costanza, and Roel Wuyts. A high-level library for multidimensional arrays programming in computational science. Concurrency and Computation: Practice and Experience, 30(7):??, April 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4376.

Christoforou:2013:SIP

[CAG^{+}13] Evgenia Christoforou, Antonio Fernández Anta, Chryssis Georgiou, Miguel A. Mosteiro, and Angel Sánchez. Special issue papers: Applying the dynamics of evolution to achieve reliability in master–worker computing. Concurrency and Computation: Practice and Experience, 25(17):2363–2380, December 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chowdhury:2017:MHS

[CAKH17] M. U. Chowdhury, J. H. Abawajy, A. V. Kelarev, and T. Hochin. Multilayer hybrid strategy for phishing email zero-day filtering. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cannataro:2006:NGG

[Can06] Mario Cannataro. Next-generation Grids: requirements and knowledge-based services. Concurrency and Computation: Practice and Experience, 18(8):887–898, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cruz:2019:MGR

[CBB^{+}19] Rommel A. Q. Cruz, Cristiana Bentes, Bernardo Breder, Eduardo Vasconcellos, Esteban Clua, Pablo M. C. de Carvalho, and Lúcia M. A. Drummond. Maximizing the GPU resource usage by reordering concurrent kernels submission. Concurrency and Computation: Practice and Experience, 31(18): e4409:1–e4409:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cohen-Boulakia:2008:APC

[CBBCD08] Sarah Cohen-Boulakia, Olivier Biton, Shirley Cohen, and Susan Davidson. Addressing the provenance challenge using ZOOM. Concurrency and Computation: Practice and Experience, 20(5):497–506, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carmichael:2011:BIC

[CBHTE11] Rory Carmichael, Patrick Braga-Henebry, Douglas Thain, and Scott Emrich. Biocompute 2.0: an improved collaborative workspace for data intensive bio-science. Concurrency and Computation: Practice and Experience, 23(17): 2305–2314, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coronado-Barrientos:2019:ANF

[CBIGL19] E. Coronado-Barrientos, G. Indalecio, and A. García-Loureiro. AXC: a new format to perform the SpMV oriented to Intel Xeon Phi architecture in OpenCL. Concurrency and Computation: Practice and Experience, 31(1):e4864:1– e4864:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coulson:2004:ARM

[CBP^{+}04] Geoff Coulson, Gordon Blair, Nikos Parlavantzas, Wai Kit Yeung, and Wei Cai. Applying the reflective middleware approach in Grid computing. Concurrency and Computation: Practice and Experience, 16(5):433–440, April 25, 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chapman:2002:APU

[CBPP02] B. Chapman, F. Bregier, A. Patil, and A. Prabhakar. Achieving performance under OpenMP on ccNUMA and software distributed shared memory systems. Concurrency and Computation: Practice and Experience, 14(8–9):713–739, July/ August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016122/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016122{\&}PLACEBO=IE.pdf.

Choi:2011:BLS

[CBQ^{+}11] Jong Youl Choi, Seung-Hee Bae, Judy Qiu, Bin Chen, and David Wild. Browsing large-scale cheminformatics data with dimension reduction. Concurrency and Computation: Practice and Experience, 23(17):2315–2325, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2009:SII

[CC09] Jinjun Chen and Massimo Cafaro. Special issue: 3rd International Workshop on Workflow Management and Applications in Grid Environments (WaGe2008). Concurrency and Computation: Practice and Experience, 21(16):1961–1964, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cardoso:2010:MSO

[CC10] M. C. Cardoso and F. M. Costa. MPI support on opportunistic grids based on the InteGrade middleware. Concurrency and Computation: Practice and Experience, 22(3):343–357, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Camata:2013:PIP

[CC13] Jose J. Camata and Alvaro L. G. A. Coutinho. Parallel implementation and performance analysis of a linear octree finite element mesh generation scheme. Concurrency and Computation: Practice and Experience, 25(6):826–842, April 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2015:TAC

[CC15] Chia Jung Chen and Rong Guey Chang. Thermal-aware code transformation across functional units. Concurrency and Computation: Practice and Experience, 27(3):594–609, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2012:MLS

[CCC12a] Chih-Yuan Chen, Jhong-Yi Ciou, and Rong-Guey Chang. Multi-level simultaneous multithreading scheduling to reduce the temperature of register files. Concurrency and Computation: Practice and Experience, 24(12):1296–1316, August 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2012:CSE

[CCC12b] Ching-Mao Chen, Fu-Shun Chu, and Peng-Sheng Chen. Compiler support for effective XSL transformation. Concurrency and Computation: Practice and Experience, 24(14):1572– 1593, September 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Candela:2016:SDM

[CCC^{+}16] Leonardo Candela, Donatella Castelli, Gianpaolo Coro, Pasquale Pagano, and Fabio Sinibaldi. Species distribution modeling in the cloud. Concurrency and Computation: Practice and Experience, 28(4):1056–1079, March 25, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2006:RMC

[CCCC06] Jiannong Cao, Alvin T. S. Chan, Stephen C. F. Chan, and Nick K. C. Cheung. A robust monitor construct with runtime fault detection. Concurrency and Computation: Practice and Experience, 18(5):471–500, April 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2013:SIPb

[CCCW13] Weihua Cao, Gang Chen, Xin Chen, and Min Wu. Special issue papers: Optimal tracking agent: a new framework of reinforcement learning for multiagent systems. Concurrency and Computation: Practice and Experience, 25(14): 2002–2015, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chiola:2008:OFT

[CCG^{+}08] Giovanni Chiola, Gennaro Cordasco, Luisa Gargano, Alberto Negro, and Vittorio Scarano. Optimizing the finger tables in Chord-like DHTs. Concurrency and Computation: Practice and Experience, 20(6):643–657, April 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carretero:2016:EHA

[CČJ^{+}16] Jesus Carretero, Raimondas Čiegis, Emmanuel Jeannot, Laurent Lefevre, Gudula Rünger, Domenico Talia, and Julius Žilinskas. Editorials: HeteroPar 2014, APCIE 2014, and TASUS 2014 special issue. Concurrency and Computation: Practice and Experience, 28(3):748–749, March 10, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chard:2017:SCD

[CCK^{+}17] Kyle Chard, Simon Caton, Kai Kugler, Omer Rana, and Daniel S. Katz. A social content delivery network for e-Science. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2017:MLF

[CCL^{+}17] Xian Chen, Wenzhi Chen, Zhongyong Lu, Yu Zhang, Rui Chang, Mohammad Mehedi Hassan, Abdulhameed Alelaiwi, and Yang Xiang. MBSA: a lightweight and flexible storage architecture for virtual machines. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2019:CTB

[CCLP19] Dehua Chen, Nannan Che, Jiajin Le, and Qiao Pan. A cotraining based entity recognition approach for cross-disease clinical documents. Concurrency and Computation: Practice and Experience, 31(23):e4505:1–e4505:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Costea:2017:TOO

[CCM^{+}17] Mihail Costea, Radu-Ioan Ciobanu, Radu-Corneliu Marin, Ciprian Dobre, Constandinos X. Mavromoustakis, George Mastorakis, and Fatos Xhafa. Total order in opportunistic networks. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cassagnes:2015:SDA

[CCO15a] Aurelien Cassagnes, Yu Chen, and Hirotada Ohashi. Shuffle up and deal: accelerating GPGPU Monte Carlo simulation with application to option pricing. Concurrency and Computation: Practice and Experience, 27(17):5203–5213, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chimeh:2015:CVP

[CCO^{+}15b] Mozhgan Chimeh, Paul Cockshott, Susanne B. Oehler, Ashkan Tousimojarad, and Tian Xu. Compiling Vector Pascal to the XeonPhi. Concurrency and Computation: Practice and Experience, 27(17):5060–5075, December 10, 2015.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coro:2015:PEN

[CCP^{+}15] Gianpaolo Coro, Leonardo Candela, Pasquale Pagano, Angela Italiano, and Loredana Liccardo. Parallelizing the execution of native data mining algorithms for computational biology. Concurrency and Computation: Practice and Experience, 27(17):4630–4644, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2014:TSC

[CCS14] Yan Cui, Yu Chen, and Yuanchun Shi. Towards scalability collapse behavior on multicores. Concurrency and Computation: Practice and Experience, 26(2):336–359, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2010:DDC

[CCSS10] Po-Cheng Chen, Jyh-Biau Chang, Yen-Liang Su, and Ce-Kuen Shieh. On-demand data co-allocation with user-level cache for grids. Concurrency and Computation: Practice and Experience, 22(18):2488–2513, December 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cohen:2015:CMD

[CCT15] Johanne Cohen, Daniel Cordeiro, and Denis Trystram. Coordination mechanisms for decentralized parallel systems. Concurrency and Computation: Practice and Experience, 27(5): 1255–1272, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cohen:2011:MOS

[CCTW11] Johanne Cohen, Daniel Cordeiro, Denis Trystram, and Frédéric Wagner. Multi-organization scheduling approximation algorithms. Concurrency and Computation: Practice and Experience, 23(17):2220–2234, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2004:EPI

[CCW04] Min-Bin Chen, Tyng-Ruey Chuang, and Jan-Jan Wu. Efficient parallel implementations of near Delaunay triangulation with High Performance Fortran. Concurrency and Computation: Practice and Experience, 16(12):1143–1159, October 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2006:PDC

[CCW06] Min-Bin Chen, Tyng-Ruey Chuang, and Jan-Jan Wu. Parallel divide-and-conquer scheme for 2D Delaunay triangulation. Concurrency and Computation: Practice and Experience, 18 (12):1595–1612, October 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2015:AAM

[CCW^{+}15] Yuxia Cheng, Wenzhi Chen, Zonghui Wang, Xinjie Yu, and Yang Xiang. AMC: an adaptive multi-level cache algorithm in hybrid storage systems. Concurrency and Computation: Practice and Experience, 27(16):4230–4246, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cafaro:2009:DPS

[CDA09] M. Cafaro, Vincenzo De Bene, and G. Aloisio. Deterministic parallel selection algorithms on coarse-grained multicomputers. Concurrency and Computation: Practice and Experience, 21(18):2336–2354, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2017:RTP

[CDdW17] Xue Gang Chen, Sheng Duan, and Lu da Wang. Research on trend prediction and evaluation of network public opinion. Concurrency and Computation: Practice and Experience, 29 (24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2017:VRA

[CDF^{+}17] Laizhong Cui, Linyong Dong, Xianghua Fu, Zhenkun Wen, Nan Lu, and Guanjing Zhang. A video recommendation algorithm based on the combination of video content and social network. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Crivelli:2015:CGE

[CDH^{+}15] Silvia Crivelli, Rion Dooley, Raquell Holmes, Stephen Mock, and The WeFold Community. Creating a gateway that enables large-scale science coopetition. Concurrency and Computation: Practice and Experience, 27(2):446–457, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cohen:2008:PNN

[CDL08] Jeremy Cohen, John Darlington, and William Lee. Payment and negotiation for the next generation Grid and Web. Concurrency and Computation: Practice and Experience, 20(3): 239–251, March 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Casanova:2015:SMA

[CDMS15] Henri Casanova, Frédéric Desprez, George S. Markomanolis, and Frédéric Suter. Simulation of MPI applications with timeindependent traces. Concurrency and Computation: Practice and Experience, 27(5):1145–1168, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cruz:2015:CAT

[CDN15] Eduardo H. M. Cruz, Matthias Diener, and Philippe O. A. Navaux. Communication-aware thread mapping using the translation lookaside buffer. Concurrency and Computation: Practice and Experience, 27(17):4970–4992, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cuomo:2017:CSP

[CDP17] Salvatore Cuomo, Pasquale De Michele, and Monica Pragliola. A computational scheme to predict dynamics in IoT systems by using particle filter. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2019:RCP

[cDrLyC^{+}19] Hong chao Ding, Min rong Lian, Xue ying Chen, Jia ming Liu, Zu chang Zhong, Yu fan Zhang, and Min yao Zhou. Research on the correlation of port logistics and regional economic growth base on gray relational analysis method. Concurrency and Computation: Practice and Experience, 31(10): e4744:1–e4744:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cuka:2018:ENC

[CEB^{+}18] Miralda Cuka, Donald Elmazi, Kevin Bylykbashi, Evjola Spaho, Makoto Ikeda, and Leonard Barolli. Effect of node centrality for IoT device selection in opportunistic networks: a comparison study. Concurrency and Computation: Practice and Experience, 30(21):e4790:1–e4790:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cierniak:2005:ORP

[CEG^{+}05] Michal Cierniak, Marsha Eng, Neal Glew, Brian Lewis, and James Stichnoth. The Open Runtime Platform: a flexible high-performance managed runtime environment. Concurrency and Computation: Practice and Experience, 17(5–6): 617–637, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2006:ISN

[CEH^{+}06] Zhongtang Cai, Greg Eisenhauer, Qi He, Vibhore Kumar, Karsten Schwan, and Matthew Wolf. IQ-Services: networkaware middleware for interactive large-data applications. Concurrency and Computation: Practice and Experience, 18 (6):635–652, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cafaro:2008:GRB

[CEM^{+}08] M. Cafaro, I. Epicoco, M. Mirto, D. Lezzi, and G. Aloisio. The Grid Resource Broker workflow engine. Concurrency and Computation: Practice and Experience, 20(15):1725–1739, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cruz:2017:CSG

[CEM^{+}17] Francisco Cruz, Antonio Espinosa, Juan C. Moure, Jesus Cerquides, Juan A. Rodriguez-Aguilar, Kim Svensson, and Sarvapali D. Ramchurn. Coalition structure generation problems: optimization and parallelization of the IDP algorithm in multicore systems. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cardinale:2019:FAP

[CEMR19] Yudith Cardinale, Joyce El Haddad, Maude Manouvrier, and Marta Rukoz. Fuzzy ACID properties for self-adaptive composite cloud services execution. Concurrency and Computation: Practice and Experience, 31(2):e4360:1–e4360:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coll:2003:UMN

[CFP^{+}03] Salvador Coll, Eitan Frachtenberg, Fabrizio Petrini, Adolfy Hoisie, and Leonid Gurvits. Using multirail networks in highperformance clusters. Concurrency and Computation: Practice and Experience, 15(7–8):625–651, June/July 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cebrian:2017:DPS

[CFPJ^{+}17] Juan M. Cebrián, Ricardo Fernández-Pascual, Alexandra Jimborean, Manuel E. Acacio, and Alberto Ros. A dedicated private-shared cache design for scalable multiprocessors. Concurrency and Computation: Practice and Experience, 29(2): ??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Comito:2017:EAT

[CFTT17] Carmela Comito, Deborah Falcone, Domenico Talia, and Paolo Trunfio. Energy-aware task allocation for small devices in wireless networks. Concurrency and Computation: Practice and Experience, 29(1):??, January 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Clifford:2008:TPV

[CFV^{+}08] Ben Clifford, Ian Foster, Jens-S. Voeckler, Michael Wilde, and Yong Zhao. Tracking provenance in a virtual data grid. Concurrency and Computation: Practice and Experience, 20 (5):565–575, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coglio:2001:TSJ

[CG01] Alessandro Coglio and Allen Goldberg. Type safety in the JVM: some problems in Java 2 SDK 1.2 and proposed solutions. Concurrency and Computation: Practice and Experience, 13(13):1153–1171, November 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/88011336/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=88011336&PLACEBO=IE.
pdf.

Carroll:2010:FVO

[CG10] Thomas E. Carroll and Daniel Grosu. Formation of virtual organizations in grids: a game-theoretic approach. Concurrency and Computation: Practice and Experience, 22(14): 1972–1989, September 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coulson:2006:CBM

[CGB^{+}06] Geoff Coulson, Paul Grace, Gordon Blair, Wei Cai, Chris Cooper, David Duce, Laurent Mathy, Wai Kit Yeung, Barry Porter, Musbah Sagar, and Wei Li. A component-based middleware framework for configurable and reconfigurable Grid computing. Concurrency and Computation: Practice and Experience, 18(8):865–874, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carretero:2017:EAA

[CGBNM17] Jesus Carretero, Javier Garcia-Blas, Koji Nakano, and Peter Mueller. Editorial: Algorithms and applications towards the convergence of high-end data-intensive and computing systems. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cuenca:2017:GIB

[CGGH17] J. Cuenca, L. P. García, D. Giménez, and F. J. Herrera. Guided installation of basic linear algebra routines in a cluster with manycore components. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Churches:2006:PSD

[CGH^{+}06] David Churches, Gabor Gombas, Andrew Harrison, Jason Maassen, Craig Robinson, Matthew Shields, Ian Taylor, and Ian Wang. Programming scientific and distributed workflow with Triana services. Concurrency and Computation: Practice and Experience, 18(10):1021–1037, August 25, 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chowdhury:2017:RHD

[CGI17] Mozammel Chowdhury, Junbin Gao, and Rafiqul Islam. Robust human detection and localization in security applications. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cerotti:2016:MAP

[CGIP16] D. Cerotti, M. Gribaudo, M. Iacono, and P. Piazzolla. Modeling and analysis of performances for concurrent multithread applications on multicore and graphics processing unit systems. Concurrency and Computation: Practice and Experience, 28(2):438–452, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cobb:2007:NST

[CGK^{+}07] John W. Cobb, Al Geist, James A. Kohl, Stephen D. Miller, Peter F. Peterson, Gregory G. Pike, Michael A. Reuter, Tom Swain, Sudharshan S. Vazhkudai, and Nithya N. Vijayakumar. The Neutron Science TeraGrid Gateway: a TeraGrid science gateway to support the Spallation Neutron Source. Concurrency and Computation: Practice and Experience, 19 (6):809–826, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cockshott:2014:ALB

[CGK14] P. Cockshott, Y. Gdura, and P. Keir. Array languages and the N-body problem. Concurrency and Computation: Practice and Experience, 26(4):935–951, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Calore:2016:PPA

[CGK^{+}16] Enrico Calore, Alessandro Gabbana, Jiri Kraus, Sebastiano Fabio Schifano, and Raffaele Tripiccione. Performance and portability of accelerated lattice Boltzmann applications with OpenACC. Concurrency and Computation: Practice and Experience, 28(12):3485–3502, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Christmann:2013:FEE

[CGKW13] Dennis Christmann, Reinhard Gotzhein, Marc Krämer, and Martin Winkler. Flexible and energy-efficient duty cycling in wireless networks with MacZ. Concurrency and Computation: Practice and Experience, 25(2):218–233, February 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Castro:2019:EBT

[CGMJ^{+}19] F. M. Castro, N. Guil, M. J. Marín-Jiménez, J. Pérez-Serrano, and M. Ujaldón. Energy-based tuning of convolutional neural networks on multi-GPUs. Concurrency and Computation: Practice and Experience, 31(21):e4786:1–e4786:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cicirelli:2015:EEM

[CGN15] Franco Cicirelli, Andrea Giordano, and Libero Nigro. Efficient environment management for distributed simulation of large-scale situated multi-agent systems. Concurrency and Computation: Practice and Experience, 27(3):610–632, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carreno:2015:SPS

[CGOF15] Pablo Carreño, Francisco J. Gutierrez, Sergio F. Ochoa, and Giancarlo Fortino. Supporting personal security using participatory sensing. Concurrency and Computation: Practice and Experience, 27(10):2531–2546, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Crain:2019:SFB

[CGR19] Tyler Crain, Vincent Gramoli, and Michel Raynal. A speculation-friendly binary search tree. Concurrency and Computation: Practice and Experience, 31(4):e4883:1– e4883:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chow:2015:PPO

[CGS15] Jonathan Chow, Nasser Giacaman, and Oliver Sinnen. Pipeline pattern in an object-oriented, task-parallel environment. Concurrency and Computation: Practice and Experience, 27(5):1273–1291, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Calore:2017:EDT

[CGST17] Enrico Calore, Alessandro Gabbana, Sebastiano Fabio Schifano, and Raffaele Tripiccione. Evaluation of DVFS techniques on modern HPC processors and accelerators for energy-aware applications. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chui:2004:ADR

[CH04] Yim-Pan Chui and Pheng-Ann Heng. Attitude dead reckoning in a collaborative virtual environment using cumulative polynomial extrapolation of quaternions. Concurrency and Computation: Practice and Experience, 16(15):1575–1599, December 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2019:RER

[CH19] Hong Chen and WenZhe Hu. Research on exchange rate passthrough effect based on artificial intelligence approach. Concurrency and Computation: Practice and Experience, 31(9): e4986:1–e4986:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chang:2003:OOS

[Cha03] Moon Seok Chang. Optimizing operating system performance for CC-NUMA architectures. Concurrency and Computation: Practice and Experience, 15(14):1257–1274, December 10, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2018:AVC

[Che18] Yuan Cheng. Application of VR computer image technology in 3D virtualization of sports human science. Concurrency and Computation: Practice and Experience, 30(24):e4934:1– e4934:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2019:SII

[Che19] Jinjun Chen. Special issue: 2016 International Conference on Information Technology in Medicine and Education (ITME2016). Concurrency and Computation: Practice and Experience, 31(23):e5434:1–e5434:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chotipant:2018:SAA

[CHH18] Supannada Chotipant, Farookh Khadeer Hussain, and Omar Khadeer Hussain. SERNOTATE: an automated approach for business service description annotation for efficient service retrieval and composition. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carrion:2015:IGI

[CHL15] Ismael Marín Carrión, Eduardo Huedo, and Ignacio M. Llorente. Interoperating grid infrastructures with the Grid-Way metascheduler. Concurrency and Computation: Practice and Experience, 27(9):2278–2290, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2015:RSD

[CHM15] Lanping Chen, Zhengzhi Han, and Zhenghua Ma. Robust stabilization design for large-scale parameterized nonlinear switched systems. Concurrency and Computation: Practice and Experience, 27(12):3065–3074, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choi:2001:PPF

[Cho01] Jaeyoung Choi. PoLAPACK: parallel factorization routines with algorithmic blocking. Concurrency and Computation: Practice and Experience, 13(12):1033–1047, October 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/85512006/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
85512006&PLACEBO=IE.pdf.

Canon:2017:CCC

[CHP17] L.-C. Canon, P.-C. Héam, and L. Philippe. Controlling the correlation of cost matrices to assess scheduling algorithm performance on heterogeneous platforms. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chan:2007:CCT

[CHPvdG07] Ernie Chan, Marcel Heimlich, Avi Purkayastha, and Robert van de Geijn. Collective communication: theory, practice, and experience. Concurrency and Computation: Practice and Experience, 19(13):1749–1783, September 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2019:DIR

[CHX^{+}19] Jinbo Chen, Ye Huang, Pengxiao Xia, Yuying Zhang, and Yu Zhong. Design and implementation of real-time traceability monitoring system for agricultural products supply chain under Internet of Things architecture. Concurrency and Computation: Practice and Experience, 31(10):e4766:1–e4766:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Constantiou:2010:RFS

[CHZ10] Ioanna Constantiou, Natascha Hoebel, and Roberto V. Zicari. Reputation, framing strategies and user’s choice of content on the Web: an empirical study. Concurrency and Computation: Practice and Experience, 22(7):872–889, May 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Constantiou:2012:HDF

[CHZ12] Ioanna Constantiou, Natascha Hoebel, and Roberto V. Zicari. How do framing strategies influence the user’s choice of content on the Web? Concurrency and Computation: Practice and Experience, 24(17):2207–2220, December 10, 2012.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ciobanu:2012:FSA

[CJ12] Gabriel Ciobanu and Cālin Juravle. Flexible software architecture and language for mobile agents. Concurrency and Computation: Practice and Experience, 24(6):559–571, May 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chang:2018:TMP

[CJC^{+}18] Rui Chang, Liehui Jiang, Wenzhi Chen, Hongqi He, Shuiqiao Yang, Hang Jiang, Wei Liu, and Yong Liu. Towards a multilayered permission-based access control for extending Android security. Concurrency and Computation: Practice and Experience, 30(5):??, March 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4180.

Chang:2019:IHA

[CJX^{+}19] Rui Chang, Liehui Jiang, Yaobin Xie, Hongqi He, Danmin Chen, and Lu Ren. Implementing a hardware-assisted memory management mechanism for ARM platforms using the B method. Concurrency and Computation: Practice and Experience, 31(21):e4659:1–e4659:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2015:LSF

[CJZ^{+}15] Gang Chen, Hai Jin, Deqing Zou, Bing Bing Zhou, and Weizhong Qiang. A lightweight software fault-tolerance system in the cloud environment. Concurrency and Computation: Practice and Experience, 27(12):2982–2998, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2010:BDT

[CJZZ10] Ge Cheng, Hai Jin, Deqing Zou, and Xinwen Zhang. Building dynamic and transparent integrity measurement and protection for virtualized platform in cloud computing. Concurrency and Computation: Practice and Experience, 22(13): 1893–1910, September 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chalouf:2013:SAD

[CK13] M. A. Chalouf and F. Krief. A secured, automated, and dynamic end-to-end service level negotiation. Concurrency and Computation: Practice and Experience, 25(2):180–202, February 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cushing:2014:AWS

[CKBB14] Reginald Cushing, Spiros Koulouzis, Adam Belloum, and Marian Bubak. Applying workflow as a service paradigm to application farming. Concurrency and Computation: Practice and Experience, 26(6):1297–1312, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chou:2009:MAB

[CKC09] Yu-Cheng Chou, David Ko, and Harry H. Cheng. Mobile agent-based computational steering for distributed applications. Concurrency and Computation: Practice and Experience, 21(18):2377–2399, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cook:2019:EPC

[CKD^{+}19] Brandon Cook, Thorsten Kurth, Jack Deslippe, Pierre Carrier, Nick Hill, and Nathan Wichmann. Eigensolver performance comparison on Cray XC systems. Concurrency and Computation: Practice and Experience, 31(16):e4997:1– e4997:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Charara:2017:FDT

[CKL17] Ali Charara, David Keyes, and Hatem Ltaief. A framework for dense triangular matrix kernels on various manycore architectures. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Conti:2019:CCE

[CKL19] Mauro Conti, Pallavi Kaliyar, and Chhagan Lal. CENSOR: Cloud-enabled secure IoT architecture over SDN paradigm. Concurrency and Computation: Practice and Experience, 31 (8):e4978:1–e4978:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Curbera:2006:IBA

[CKNW06] Francisco Curbera, Rania Khalaf, William A. Nagy, and Sanjiva Weerawarana. Implementing BPEL4WS: the architecture of a BPEL4WS implementation. Concurrency and Computation: Practice and Experience, 18(10):1219–1228, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cavendish:2010:MVA

[CKOG10] D. Cavendish, H. Koide, Y. Oie, and M. Gerla. A Mean Value Analysis approach to transaction performance evaluation of multi-server systems. Concurrency and Computation: Practice and Experience, 22(10):1267–1285, July 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Christodoulopoulos:2013:SIP

[CKRO13] Kostas Christodoulopoulos, Kostas Katrinis, Marco Ruffini, and Donal O’Mahony. Special issue papers: Tailoring the network to the problem: topology configuration in hybrid electronic packet switched/optical circuit switched interconnects. Concurrency and Computation: Practice and Experience, 25 (17):2412–2432, December 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choi:2010:DDR

[CKSC10] Hyung Won Choi, Hukeun Kwak, Andrew Sohn, and Kyusik Chung. DRIVE — Dispatching Requests Indirectly through Virtual Environment. Concurrency and Computation: Practice and Experience, 22(4):398–418, March 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cabo:2001:TOO

[CL01] M. J. Aramburu Cabo and R. Berlanga Llavori. A temporal object-oriented model for digital libraries of documents. Concurrency and Computation: Practice and Experience, 13(11): 987–1011, September 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/85012150/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=85012150&PLACEBO=IE.pdf.

Chunlin:2007:OAD

[CL07] Li Chunlin and Li Layuan. An optimization approach for decentralized QoS-based scheduling based on utility and pricing in Grid computing. Concurrency and Computation: Practice and Experience, 19(1):107–128, January 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2008:SIS

[CL08] Jinjun Chen and Yonghwan Lee. Special issue: Second International Workshop on Workflow Management and Applications in Grid Environments (WaGe2007). Concurrency and Computation: Practice and Experience, 20(15):1721–1723, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carver:2010:DRT

[CL10] Richard H. Carver and Yu Lei. Distributed reachability testing of concurrent programs. Concurrency and Computation: Practice and Experience, 22(18):2445–2466, December 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2013:ESIb

[CL13] Jinjun Chen and Jianxun Liu. Editorial: Special issue: 2011 international conference on cloud and green computing (CGC2011). Concurrency and Computation: Practice and Experience, 25(18):2433–2434, December 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chang:2014:LCG

[CL14] Zu-Ling Chang and Dandan Li. On the linear complexity of generalized cyclotomic binary sequences of length 2pq. Concurrency and Computation: Practice and Experience, 26(8): 1520–1530, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choi:2016:XAP

[CL16] Min Choi and Seung-Ho Lim. x86-Android performance improvement for x86 smart mobile devices. Concurrency and Computation: Practice and Experience, 28(10):2770–2780, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carracciuolo:2018:INL

[CL18] Luisa Carracciuolo and Marco Lapegna. Implementation of a non-linear solver on heterogeneous architectures. Concurrency and Computation: Practice and Experience, 30(24): e4903:1–e4903:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2019:SAT

[CL19] Junbing Cheng and Dengao Li. A SINS-aided two-step fast acquisition method for GNSS signal based on compressive sensing. Concurrency and Computation: Practice and Experience, 31(24):e5369:1–e5369:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Clarke:2018:TSC

[Cla18] Sam Clarke. Trust separation on the Cray XC40 using PBS Pro. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2017:EHQ

[CLF^{+}17] Xuhao Chen, Pingfan Li, Jianbin Fang, Tao Tang, Zhiying Wang, and Canqun Yang. Efficient and high-quality sparse graph coloring on GPUs. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2019:IOD

[CLF^{+}19] Yuedan Chen, Kenli Li, Xiongwei Fei, Zhe Quan, and Keqin Li. Implementation and optimization of a data protecting model on the Sunway TaihuLight supercomputer with heterogeneous many-core processors. Concurrency and Computation: Practice and Experience, 31(21):e4758:1–e4758:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chu:2008:SOG

[CLH^{+}08] Xingchen Chu, Andrew Lonie, Peter Harris, S. Randall Thomas, and Rajkumar Buyya. A service-oriented Grid environment for integration of distributed kidney models and resources. Concurrency and Computation: Practice and Experience, 20(9):1095–1111, June 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2011:ADH

[CLH^{+}11] Rong-Jian Chen, Jun-Jian Lin, Su-Min Hung, Jui-Lin Lai, and Shi-Jinn Horng. Architecture design of high-efficient and non-memory AES crypto-core for WPAN. Concurrency and Computation: Practice and Experience, 23(12):1332–1347, August 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chou:2013:IEE

[CLH13] Li-Der Chou, David Chunhu Li, and Wei-Yong Hong. Improving energy-efficient communications with a battery lifetimeaware mechanism in IEEE802.16e wireless networks. Concurrency and Computation: Practice and Experience, 25(1): 94–111, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2016:RPR

[CLH^{+}16] Zhenhua Chen, Shundong Li, Qiong Huang, Yilei Wang, and Sufang Zhou. A restricted proxy re-encryption with keyword search for fine-grained data access control in cloud storage. Concurrency and Computation: Practice and Experience, 28 (10):2858–2876, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2019:EPC

[CLH19] Xi Chen, Xin Li, and Zhenzhen Huang. Exploration of power consumption monitoring based on Internet of Things. Concurrency and Computation: Practice and Experience, 31(10): e4737:1–e4737:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chancelier:2014:UPN

[CLL14] Jean-Philippe Chancelier, Bernard Lapeyre, and Jérôme Lelong. Using Premia and Nsp for constructing a risk management benchmark for testing parallel architecture. Concurrency and Computation: Practice and Experience, 26(9): 1654–1665, June 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2018:NCR

[CLL^{+}18] Jun Cai, Jian-Zhen Luo, Yan Liu, Wenguo Wei, and Fangyuan Lei. A network community restructuring mechanism for transport efficiency improvement in scale-free complex networks. Concurrency and Computation: Practice and Experience, 30(5):??, March 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4273.

Chen:2019:SLS

[CLL^{+}19] Xing Chen, Junxin Lin, Bing Lin, Tao Xiang, Ying Zhang, and Gang Huang. Self-learning and self-adaptive resource allocation for cloud-based software services. Concurrency and Computation: Practice and Experience, 31(23):e4463:1– e4463:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carchiolo:2012:TAP

[CLMM12] Vincenza Carchiolo, Alessandro Longheu, Michele Malgeri, and Giuseppe Mangioni. Trust assessment: a personalized, distributed, and secure approach. Concurrency and Computation: Practice and Experience, 24(6):605–617, May 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cardellini:2018:OOD

[CLNR18] Valeria Cardellini, Francesco Lo Presti, Matteo Nardelli, and Gabriele Russo Russo. Optimal operator deployment and replication for elastic distributed data stream processing. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4334.

Caruana:2017:GRA

[CLQ^{+}17] Godwin Caruana, Maozhen Li, Man Qi, Mukhtaj Khan, and Omer Rana. gSched: a resource aware Hadoop scheduler for heterogeneous cloud computing environments. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carlini:2018:MDG

[CLR18] Emanuele Carlini, Alessandro Lulli, and Laura Ricci. Model driven generation of mobility traces for distributed virtual environments with TRACE. Concurrency and Computation: Practice and Experience, 30(20):e4235:1–e4235:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Clauss:2015:NSS

[CLRB15] Carsten Clauss, Stefan Lankes, Pablo Reble, and Thomas Bemmerl. New system software for parallel programming models on the Intel SCC many-core processor. Concurrency and Computation: Practice and Experience, 27(9):2235–2259, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chetsa:2014:TSB

[CLS14] Ghislain Landry Tsafack Chetsa, Laurent Lefevrem, and Patricia Stolf. A three step blind approach for improving high performance computing systems’ energy performance. Concurrency and Computation: Practice and Experience, 26(15): 2612–2629, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2016:DAT

[CLT^{+}16] Junjie Chen, Kenli Li, Zhuo Tang, Chubo Liu, Yan Wang, and Keqin Li. Data-aware task scheduling on heterogeneous hybrid memory multiprocessor systems. Concurrency and Computation: Practice and Experience, 28(17):4443–4459, December 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cesario:2013:SIP

[CLTT13] Eugenio Cesario, Marco Lackovic, Domenico Talia, and Paolo Trunfio. Special issue papers: Programming knowledge discovery workflows in service-oriented distributed systems. Concurrency and Computation: Practice and Experience, 25 (10):1482–1504, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2015:AEI

[CLW^{+}15] Guanghai Cui, Mingchu Li, Zhen Wang, Jiankang Ren, Dong Jiao, and Jianhua Ma. Analysis and evaluation of incentive mechanisms in P2P networks: a spatial evolutionary game theory perspective. Concurrency and Computation: Practice and Experience, 27(12):3044–3064, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2018:QGB

[CLW^{+}18] Ling Chen, Yan Lin, Jingchang Wang, Heqing Huang, Donghui Chen, and Yong Wu. Query grouping-based multiquery optimization framework for interactive SQL query engines on Hadoop. Concurrency and Computation: Practice and Experience, 30(19):e4676:1–e4676:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2019:ADP

[CLW^{+}19] Hui Cao, Shubo Liu, Longfei Wu, Zhitao Guan, and Xiaojiang Du. Achieving differential privacy against non-intrusive load monitoring in smart grid: a fog computing approach. Concurrency and Computation: Practice and Experience, 31(22): e4528:1–e4528:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2007:PIP

[CLX07] Jiannong Cao, Fred Bojin Liu, and Cheng-Zhong Xu. P2PGrid: integrating P2P networks into the Grid environment. Concurrency and Computation: Practice and Experience, 19(7):1023–1046, May 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2012:QBF

[CLX^{+}12] Yongrui Cui, Mingchu Li, Yang Xiang, Yizhi Ren, and Silvio Cesare. A QoS-based fine-grained reputation system in the grid environment. Concurrency and Computation: Practice and Experience, 24(17):1990–2006, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chang:2016:APC

[CLYC16] Chih-Hung Chang, Chih-Wei Lu, Chao-Tung Yang, and Tzu-Chieh Chang. An approach of performance comparisons with OpenMP and CUDA parallel programming on multicore systems. Concurrency and Computation: Practice and Experience, 28(16):4230–4245, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2017:ISK

[CLZ^{+}17] Longwang Cheng, Wei Li, Li Zhou, Chunsheng Zhu, Jibo Wei, and Yantao Guo. Increasing secret key capacity of OFDM systems: a geometric program approach. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2010:AMS

[CLZX10] Xue Chen, Xiangfeng Luo, Shun Xiang Zhang, and Zheng Xu. Analysis and modeling of the semantically associated network on the Web. Concurrency and Computation: Practice and Experience, 22(7):767–787, May 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coulson:2002:QSD

[CM02] Geoff Coulson and Oveeyen Moonian. A quality of service driven concurrency framework for object-based middleware. Concurrency and Computation: Practice and Experience, 14 (4):241–259, April 10, 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/93513492/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=93513492{\&}PLACEBO=IE.pdf.

Cahoon:2005:RAE

[CM05] Brendon Cahoon and Kathryn S. McKinley. Recurrence analysis for effective array prefetching in Java. Concurrency and Computation: Practice and Experience, 17(5–6):589–616, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chivers:2006:RSB

[CM06] Howard Chivers and John McDermid. Refactoring servicebased systems: how to avoid trusting a workflow service. Concurrency and Computation: Practice and Experience, 18 (10):1255–1275, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Christie:2007:LPT

[CM07a] Marcus Christie and Suresh Marru. The LEAD Portal: a TeraGrid gateway and application service architecture. Concurrency and Computation: Practice and Experience, 19(6): 767–781, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cunha:2007:SIP

[CM07b] José C. Cunha and Pedro D. Medeiros. Special issue: Parallel and Distributed Computing (EuroPar 2005). Concurrency and Computation: Practice and Experience, 19(17): 2183–2184, December 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cuomo:2018:SND

[CM18] Salvatore Cuomo and Francesco Maiorano. Social network data analysis and mining applications for the Internet of Data. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4527.

Cicerre:2006:SPE

[CMB06] F. R. L. Cicerre, E. R. M. Madeira, and L. E. Buzato. Structured process execution middleware for Grid computing. Concurrency and Computation: Practice and Experience, 18(6): 581–594, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cahyani:2017:FDA

[CMCAA17] Niken Dwi Wahyu Cahyani, Ben Martini, Kim-Kwang Raymond Choo, and Akbp Muhammad Nuh Al-Azhar. Forensic data acquisition from cloud-of-things devices: windows Smartphones as a case study. Concurrency and Computation:

Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Curcin:2011:STW

[CMD11] Vasa Curcin, Paolo Missier, and David De Roure. Simulating Taverna workflows using stochastic process algebras. Concurrency and Computation: Practice and Experience, 23(16): 1920–1935, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Czech:2017:DCD

[CMD17] Wojciech Czech, Wojciech Mielczarek, and Witold Dzwinel. Distributed computing of distance-based graph invariants for analysis and visualization of complex networks. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Caceres:2010:PRR

[CML^{+}10] E. N. Cáceres, H. Mongelli, L. Loureiro, C. Nishibe, and S. W. Song. Performance results of running parallel applications on the InteGrade. Concurrency and Computation: Practice and Experience, 22(3):375–393, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chakroun:2013:RTD

[CMMB13] I. Chakroun, M. Mezmaz, N. Melab, and A. Bendjoudi. Reducing thread divergence in a GPU-accelerated branch-and-bound algorithm. Concurrency and Computation: Practice and Experience, 25(8):1121–1136, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2017:VME

[CMMS17] Jiageng Chen, Rashed Mazumder, Atsuko Miyaji, and Chunhua Su. Variable message encryption through blockcipher compression function. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Caromel:2008:POM

[CMPT08] Denis Caromel, Luis Mateu, Guillaume Pothier, and Éric Tanter. Parallel object monitors. Concurrency and Computation: Practice and Experience, 20(12):1387–1417, August 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cunha:2017:CCD

[CMS17] M. Cunha, N. C. Mendonça, and A. Sampaio. Cloud Crawler: a declarative performance evaluation environment for infrastructure-as-a-service clouds. Concurrency and Computation: Practice and Experience, 29(1):??, January 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cugler:2013:SIP

[CMT13] Daniel Cintra Cugler, Claudia Bauzer Medeiros, and Luís Felipe Toledo. Special issue papers: An architecture for retrieval of animal sound recordings based on context variables. Concurrency and Computation: Practice and Experience, 25(16): 2310–2326, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cardenas-Montes:2017:PEN

[CMVRRVGI17] Miguel Cárdenas-Montes, Miguel A. Vega-Rodríguez, Juan José Rodríguez-Vázquez, and Antonio Gómez-Iglesias. Parallel evaluation of nonseparable functions by evolutionary algorithms on GPU. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cain:2002:CBS

[CMW02] Harold W. Cain, Barton P. Miller, and Brian J. N. Wylie. A callgraph-based search strategy for automated performance diagnosis. Concurrency and Computation: Practice and Experience, 14(3):203–217, March 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/93513484/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
93513484{\&}PLACEBO=IE.pdf.

Clarke:2002:DCE

[CN02] Jerry A. Clarke and Raju R. Namburu. A distributed computing environment for interdisciplinary applications. Concurrency and Computation: Practice and Experience, 14(13–15):

1161–1174, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cicirelli:2016:CCF

[CN16] Franco Cicirelli and Libero Nigro. Control centric framework for model continuity in time-dependent multi-agent systems. Concurrency and Computation: Practice and Experience, 28 (12):3333–3356, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2018:MGS

[CNAQ18] Tianluo Chen, Yang Ning, Amit Amritkar, and Guan Qin. Multi-GPU solution to the lattice Boltzmann method: an application in multiscale digital rock simulation for shale formation. Concurrency and Computation: Practice and Experience, 30(19):e4530:1–e4530:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2013:PED

[CNG13] Shiping Chen, Alex Ng, and Paul Greenfield. A performance evaluation of distributed database architectures. Concurrency and Computation: Practice and Experience, 25(11): 1524–1546, August 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chang:2018:PSP

[CNKJ18] J. Chang, K. B. Nakshatrala, M. G. Knepley, and L. Johnsson. A performance spectrum for parallel computational frameworks that solve PDEs. Concurrency and Computation: Practice and Experience, 30(11):??, June 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4401.

Chretien:2015:USP

[CNP^{+}15] Stéphane Chrétien, Jean-Marc Nicod, Laurent Philippe, Veronika Rehn-Sonigo, and Lamiel Toch. Using a sparse promoting method in linear programming approximations to schedule parallel jobs. Concurrency and Computation: Practice and Experience, 27(14):3561–3586, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Crescenzi:2009:CBB

[CNPP09] P. Crescenzi, C. Nocentini, A. Pietracaprina, and G. Pucci. On the connectivity of Bluetooth-based ad hoc networks. Concurrency and Computation: Practice and Experience, 21 (7):875–887, May 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choi:2018:IAS

[COC18] Chang Choi, Marek R. Ogiela, and Hsing-Chung Chen. Intelligent approaches for security technologies. Concurrency and Computation: Practice and Experience, 30(3):??, February 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4408.

Coutinho:2011:MTC

[COdO^{+}11] Fábio Coutinho, Eduardo Ogasawara, Daniel de Oliveira, Vanessa Braganholo, Alexandre A. B. Lima, Alberto M. R. Dávila, and Marta Mattoso. Many task computing for orthologous genes identification in protozoan genomes using Hydra. Concurrency and Computation: Practice and Experience, 23(17):2326–2337, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coglio:2003:IOS

[Cog03] Alessandro Coglio. Improving the official specification of Java bytecode verification. Concurrency and Computation: Practice and Experience, 15(2):155–179, February 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coglio:2004:SVT

[Cog04] Alessandro Coglio. Simple verification technique for complex Java bytecode subroutines. Concurrency and Computation: Practice and Experience, 16(7):647–670, June 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chessa:2014:GPU

[CP14] Manuela Chessa and Giulia Pasquale. Graphics processing unit-accelerated techniques for bio-inspired computation in the primary visual cortex. Concurrency and Computation: Practice and Experience, 26(10):1799–1818, July 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cardenas:2007:MCC

[CPB07] Yonny Cardenas, Jean-Marc Pierson, and Lionel Brunie. Management of a cooperative cache in grids with grid cache services. Concurrency and Computation: Practice and Experience, 19(16):2141–2155, November 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chilipirea:2017:SON

[CPD^{+}17] Cristian Chilipirea, Andreea-Cristian Petre, Ciprian Dobre, Florin Pop, and George Suciu. A simulator for opportunistic networks. Concurrency and Computation: Practice and Experience, 29(2):??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coro:2019:RVE

[CPE^{+}19] Gianpaolo Coro, Marco Palma, Anton Ellenbroek, Giancarlo Panichi, Thiviya Nair, and Pasquale Pagano. Reconstructing 3D virtual environments within a collaborative einfrastructure. Concurrency and Computation: Practice and Experience, 31(11):e5028:1–e5028:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cafaro:2018:PSS

[CPEA18] Massimo Cafaro, Marco Pulimeno, Italo Epicoco, and Giovanni Aloisio. Parallel space saving on multi- and many-core processors. Concurrency and Computation: Practice and Experience, 30(7):??, April 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4160.

Carneiro:2016:MIP

[CPG^{+}16] Davide Carneiro, André Pimenta, Sérgio Gonçalves, José Neves, and Paulo Novais. Monitoring and improving performance in human-computer interaction. Concurrency and Computation: Practice and Experience, 28(4):1291–1309, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Couger:2014:ELS

[CPS^{+}14] M. Brian Couger, Lenore Pipes, Fabio Squina, Rolf Prade, Adam Siepel, Robert Palermo, Michael G. Katze, Christopher E. Mason, and Philip D. Blood. Enabling large-scale next-generation sequence assembly with Blacklight. Concurrency and Computation: Practice and Experience, 26(13): 2157–2166, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coro:2017:CCD

[CPSP17] Gianpaolo Coro, Giancarlo Panichi, Paolo Scarponi, and Pasquale Pagano. Cloud computing in a distributed einfrastructure using the web processing service standard. Concurrency and Computation: Practice and Experience, 29 (18):??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chronopoulos:2006:DLS

[CPXA06] Anthony T. Chronopoulos, Satish Penmatsa, Jianhua Xu, and Siraj Ali. Distributed loop-scheduling schemes for heterogeneous computer systems. Concurrency and Computation: Practice and Experience, 18(7):771–785, June 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2014:OWU

[CQXW14] Zhide Chen, Cheng Qiao, Li Xu, and Wei Wu. Optimizing wireless unicast and multicast sensor networks on the basis of evolutionary game theory. Concurrency and Computation: Practice and Experience, 26(5):1130–1141, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2008:SIF

[CR08] Jinjun Chen and Omer F. Rana. Special issue: First International Workshop on Workflow Systems in Grid Environments (WSGE2006). Concurrency and Computation: Practice and Experience, 20(4):311–313, March 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Caton:2012:TAM

[CR12] Simon Caton and Omer Rana. Towards autonomic management for Cloud services based upon volunteered resources.

Concurrency and Computation: Practice and Experience, 24 (9):992–1014, June 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2013:ESIa

[CR13] Jinjun Chen and Rajiv Ranjan. Editorial: Special issue: Second International Workshop on Workflow Management in Service and Cloud Computing (WMSC2010). Concurrency and Computation: Practice and Experience, 25(13): 1813–1815, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Caton:2009:DIP

[CRB09] Simon J. Caton, Omer F. Rana, and Bruce G. Batchelor. Distributed image processing over an adaptive Campus Grid. Concurrency and Computation: Practice and Experience, 21 (3):321–336, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Calheiros:2017:EIB

[CRB^{+}17] Rodrigo N. Calheiros, Kotagiri Ramamohanarao, Rajkumar Buyya, Christopher Leckie, and Steve Versteeg. On the effectiveness of isolation-based anomaly detection in cloud data centers. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carlini:2015:ICP

[CRC15a] Emanuele Carlini, Laura Ricci, and Massimo Coppola. Integrating centralized and peer-to-peer architectures to support interest management in massively multiplayer on-line games. Concurrency and Computation: Practice and Experience, 27(13):3362–3382, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Casarino:2015:GWB

[CRC^{+}15b] C. Casarino, G. Russo, G. C. Candiano, G. La Rocca, R. Barbera, G. Borasi, S. Guatelli, C. Messa, G. Passaro, and M. C. Gilardi. A GEANT4 Web-based application to support Intra-Operative Electron Radiotherapy using the European grid infrastructure. Concurrency and Computation: Practice and Experience, 27(2):458–472, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Caminero:2009:PEA

[CRCC09] A. Caminero, O. Rana, B. Caminero, and C. Carrión. Performance evaluation of an autonomic network-aware metascheduler for Grids. Concurrency and Computation: Practice and Experience, 21(13):1692–1708, September 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cruz:2012:DBM

[CRGR^{+}12] Jesús Martínez Cruz, Adrián Romero-Garcés, Juan Pedro Bandera Rubio, Rebeca Marfil Robles, and Antonio Bandera Rubio. A DDS-based middleware for quality-of-service and high-performance networked robotics. Concurrency and Computation: Practice and Experience, 24(16):1940–1952, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cuomo:2015:PPP

[CRV15] Antonio Cuomo, Massimiliano Rak, and Umberto Villano. Planting parallel program simulation on the cloud. Concurrency and Computation: Practice and Experience, 27(6): 1467–1482, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Corsaro:2006:SIF

[CS06] Angelo Corsaro and Corrado Santoro. Special issue: First International Workshop on Emerging Technologies for Nextgeneration GRID (ETNGRID 2004). Concurrency and Computation: Practice and Experience, 18(8):803–805, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cirne:2009:SIB

[CS09] Walfredo Cirne and Bruno Schulze. Special issue: The best of CCGrid’2007: a snapshot of an ‘adolescent’ area. Concurrency and Computation: Practice and Experience, 21(3): 257–263, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cuzzocrea:2013:SIP

[CS13] Alfredo Cuzzocrea and Domenico Saccà. Special issue papers: Exploiting compression and approximation paradigms for effective and efficient online analytical processing over sensor network readings in data grid environments. Concurrency and Computation: Practice and Experience, 25(14): 2016–2035, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cholia:2015:NPE

[CS15] Shreyas Cholia and Terence Sun. The NEWT platform: an extensible plugin framework for creating ReSTful HPC APIs. Concurrency and Computation: Practice and Experience, 27 (16):4304–4317, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Curtis:2016:ESS

[CS16] V. V. Curtis and C. A. A. Sanches. An efficient solution to the subset-sum problem on GPU. Concurrency and Computation: Practice and Experience, 28(1):95–113, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Christgau:2017:EOS

[CS17] Steffen Christgau and Bettina Schnor. Exploring one-sided communication and synchronization on a non-cache-coherent many-core architecture. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Crespo:2019:SPN

[CSAC19] Juan-Jose Crespo, José L. Sánchez, and Francisco J. Alfaro-Cortés. Silicon photonic networks: Signal loss and power challenges. Concurrency and Computation: Practice and Experience, 31(21):e4777:1–e4777:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Climent:2016:WSN

[CSB^{+}16] S. Climent, A. Sanchez, S. Blanc, J. V. Capella, and R. Ors. Wireless sensor network with energy harvesting: modeling and simulation based on a practical architecture using real radiation levels. Concurrency and Computation: Practice and Experience, 28(6):1812–1830, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Casas:2012:EOS

[CSBL12] Marc Casas, Harald Servat, Rosa M. Badia, and Jesús Labarta. Extracting the optimal sampling frequency of applications using spectral analysis. Concurrency and Computation: Practice and Experience, 24(3):237–259, March 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Caamano:2017:FRP

[CSC^{+}17] Juan Manuel Martinez Caamaño, Manuel Selva, Philippe Clauss, Artyom Baloian, and Willy Wolff. Full runtime polyhedral optimizing loop transformations with the generation, instantiation, and scheduling of code-bones. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chadwick:2008:CAC

[CSL08] David W. Chadwick, Linying Su, and Romain Laborde. Coordinating access control in grid services. Concurrency and Computation: Practice and Experience, 20(9):1071–1094, June 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Champrasert:2012:ESO

[CSL12] Paskorn Champrasert, Junichi Suzuki, and Chonho Lee. Exploring self-optimization and self-stabilization properties in bio-inspired autonomic cloud applications. Concurrency and Computation: Practice and Experience, 24(9):1015–1034, June 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2018:PUA

[CSL^{+}18] Yuanlong Cao, Fei Song, Guoliang Luo, Yugen Yi, Wenle Wang, Ilsun You, and Hao Wang. (P U)^{2}M^{2}: a potentially underperforming-aware path usage management mechanism for secure MPTCP-based multipathing services. Concurrency and Computation: Practice and Experience, 30 (3):??, February 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4191.

Cai:2019:CPG

[CSL^{+}19] Guorong Cai, Songzhi Su, Chengcai Leng, Jipeng Wu, Yundong Wu, and Shaozi Li. Cover patches: a general feature extraction strategy for spoofing detection. Concurrency and Computation: Practice and Experience, 31(23):e4641:1– e4641:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Caballer:2015:PDC

[CSMB15] Miguel Caballer, Damián Segrelles, Germán Moltó, and Ignacio Blanquer. A platform to deploy customized scientific virtual infrastructures on the cloud. Concurrency and Computation: Practice and Experience, 27(16):4318–4329, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ciegis:2017:PSF

[ČSMK17] Raimondas Čiegis, Vadimas Starikovičius, Svetozar Margenov, and Rima Kriauzienė. Parallel solvers for fractional power diffusion problems. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ciegis:2019:SAD

[CSMK19] Raimondas Ciegis, Vadimas Starikovicius, Svetozar Margenov, and Rima Kriauziene. Scalability analysis of different parallel solvers for 3D fractional power diffusion problems. Concurrency and Computation: Practice and Experience, 31(19): e5163:1–e5163:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carastan-Santos:2019:HCG

[CSMS^{+}19] Danilo Carastan-Santos, David C. Martins, Jr., Siang W. Song, Luiz C. S. Rozante, and Raphael Y. de Camargo. A hybrid CPU–GPU–MIC algorithm for minimal hitting set enumeration. Concurrency and Computation: Practice and Experience, 31(18):e5087:1–e5087:??, September 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coutinho:2013:PDG

[CSPM13] Bruno Coutinho, Diogo Sampaio, Fernando M. Q. Pereira, and Wagner Meira, Jr. Profiling divergences in GPU applications. Concurrency and Computation: Practice and Experience, 25(6):775–789, April 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chugunov:2006:PIM

[CSTV06] V. Chugunov, D. Svyatski, E. Tyrtyshnikov, and Yu. Vassilevski. Parallel iterative multilevel solution of mixed finite element systems for scalar equations. Concurrency and Computation: Practice and Experience, 18(5):501–518, April 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Curry:2011:GRS

[CSWB11] Matthew L. Curry, Anthony Skjellum, H. Lee Ward, and Ron Brightwell. Gibraltar: a Reed–Solomon coding library for storage applications on programmable graphics processors. Concurrency and Computation: Practice and Experience, 23 (18):2477–2495, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cafaro:2011:FFI

[CT11a] Massimo Cafaro and Piergiulio Tempesta. Finding frequent items in parallel. Concurrency and Computation: Practice and Experience, 23(15):1774–1788, October 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Comito:2011:PSM

[CT11b] Carmela Comito and Domenico Talia. P2P schema-mapping over network-bound XML data. Concurrency and Computation: Practice and Experience, 23(9):985–1009, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cesario:2012:DDM

[CT12] Eugenio Cesario and Domenico Talia. Distributed data mining patterns and services: an architecture and experiments. Concurrency and Computation: Practice and Experience, 24 (15):1751–1774, October 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ciobanu:2016:CMS

[CT16] Gabriel Ciobanu and Eneia Nicolae Todoran. Correct metric semantics for a language inspired by DNA computing. Concurrency and Computation: Practice and Experience, 28(11): 3042–3060, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Costan:2016:TSD

[CTAB16] Alexandru Costan, Radu Tudoran, Gabriel Antoniu, and Goetz Brasche. TomusBlobs: scalable data-intensive processing on Azure clouds. Concurrency and Computation: Practice and Experience, 28(4):950–976, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Casado:2015:EEA

[CTY15] Rubén Casado, Javier Tuya, and Muhammad Younas. Evaluating the effectiveness of the abstract transaction model in testing Web services transactions. Concurrency and Computation: Practice and Experience, 27(4):765–781, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cuzzocrea:2011:FMS

[Cuz11] A. Cuzzocrea. A framework for modeling and supporting data transformation services over data and knowledge grids with real-time bound constraints. Concurrency and Computation: Practice and Experience, 23(5):436–457, April 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cannataro:2007:MAP

[CV07] Mario Cannataro and Pierangelo Veltri. MS-Analyzer: preprocessing and data mining services for proteomics applications on the Grid. Concurrency and Computation: Practice and Experience, 19(15):2047–2066, October 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chhetri:2015:APB

[CVK15] M. Baruwal Chhetri, Q. Bao. Vo, and R. Kowalczyk. AutoSLAM — a policy-based framework for automated SLA establishment in cloud environments. Concurrency and Computation: Practice and Experience, 27(9):2413–2442, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chrabakh:2007:GPG

[CW07] Wahid Chrabakh and Rich Wolski. The GridSAT portal: a Grid Web-based portal for solving satisfiability problems using the national cyberinfrastructure. Concurrency and Computation: Practice and Experience, 19(6):795–808, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chetlur:2009:CIP

[CW09] Malolan Chetlur and Philip A. Wilsey. Causality information and proactive cancellation mechanisms. Concurrency and Computation: Practice and Experience, 21(11):1483–1503, August 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2011:SIF

[CW11a] Jinjun Chen and Lizhe Wang. Special issue: Fourth International Workshop on Workflow Management (ICWM2009). Concurrency and Computation: Practice and Experience, 23 (16):1853–1856, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2011:DCM

[CW11b] Ying-Ying Chen and Tie-Jun Wu. A decentralized coordination method for optimal load redistribution in heterogeneous service grids. Concurrency and Computation: Practice and Experience, 23(6):633–645, April 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2010:DPS

[CWC10] Ling-Jyh Chen, Bo-Chun Wang, and Kuan-Ta Chen. The design of puzzle selection strategies for GWAP systems. Concurrency and Computation: Practice and Experience, 22(7): 890–908, May 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2003:PJC

[CWL03] Ge Chen, Cho-Li Wang, and Francis C. M. Lau. p-Jigsaw: a cluster-based Web server with cooperative caching support.

Concurrency and Computation: Practice and Experience, 15 (7–8):681–705, June/July 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ciolczyk:2018:TLS

[CWM18] Michal Ciolczyk, Mariusz Wojakowski, and Maciej Malawski. Tracing of large-scale actor systems. Concurrency and Computation: Practice and Experience, 30(22):e4637:1–e4637:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2015:NAD

[CWMW15] Yuzhu Cheng, Weiping Wang, Geyong Min, and Jianxin Wang. A new approach to designing firewall based on multidimensional matrix. Concurrency and Computation: Practice and Experience, 27(12):3075–3088, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2006:DSI

[CWMZ06] Huajun Chen, Zhaohui Wu, Yuxin Mao, and Guozhou Zheng. DartGrid: a semantic infrastructure for building database Grid applications. Concurrency and Computation: Practice and Experience, 18(14):1811–1828, December 15, 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:IPA

[cWsThY19] Ti chun Wang, Chang sheng Tong, and Sheng hu Yao. An improved prediction algorithm of seamless tubing corrosion based on an extension neural network. Concurrency and Computation: Practice and Experience, 31(12):e4801:1– e4801:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2016:WPM

[CWXW16] Zhide Chen, Meng Wang, Li Xu, and Wei Wu. Worm propagation model in mobile network. Concurrency and Computation: Practice and Experience, 28(4):1134–1144, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2017:MDF

[CWYX17] Shuhong Chen, Guojun Wang, Guofeng Yan, and Dongqing Xie. Multi-dimensional fuzzy trust evaluation for mobile social networks based on dynamic community structures. Concurrency and Computation: Practice and Experience, 29(7): ??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2013:SIPa

[CWZL13] Jian Cao, Jie Wang, Haiyan Zhao, and Minglu Li. Special issue papers: An event view specification approach for Supporting Service process collaboration. Concurrency and Computation: Practice and Experience, 25(13):1943–1966, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2018:ESC

[CXC^{+}18] Xinhai Chen, Peizhen Xie, Lihua Chi, Jie Liu, and Chunye Gong. An efficient SIMD compression format for sparse matrix–vector multiplication. Concurrency and Computation: Practice and Experience, 30(16):e4800:1–e4800:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2015:PMP

[CXPL15] Peng Chen, Yunni Xia, Shanchen Pang, and Jia Li. A probabilistic model for performance analysis of cloud infrastructures. Concurrency and Computation: Practice and Experience, 27(17):4784–4796, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2018:DDP

[CXT^{+}18] Keyang Cheng, Fangjie Xu, Fei Tao, Man Qi, and Maozhen Li. Data-driven pedestrian re-identification based on hierarchical semantic representation. Concurrency and Computation: Practice and Experience, 30(23):e4403:1–e4403:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2017:TAS

[CXW17] Xiaoquan Chen, Rui Xue, and Chuankun Wu. Timely address space rerandomization for resisting code reuse attacks. Concurrency and Computation: Practice and Experience, 29(16): ??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2007:MSB

[CY07] Jinjun Chen and Yun Yang. Multiple states based temporal consistency for dynamic verification of fixed-time constraints in Grid workflow systems. Concurrency and Computation: Practice and Experience, 19(7):965–982, May 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2008:TGW

[CY08] Jinjun Chen and Yun Yang. A taxonomy of grid workflow verification and validation. Concurrency and Computation: Practice and Experience, 20(4):347–360, March 25, 2008. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Casado:2015:ETT

[CY15] Rubén Casado and Muhammad Younas. Emerging trends and technologies in big data processing. Concurrency and Computation: Practice and Experience, 27(8):2078–2091, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cybok:2006:GWI

[Cyb06] Dieter Cybok. A Grid workflow infrastructure. Concurrency and Computation: Practice and Experience, 18(10): 1243–1254, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2015:EBD

[CYD^{+}15] Jiuxin Cao, Bin Yu, Fang Dong, Xiangying Zhu, and Shuai Xu. Entropy-based denial-of-service attack detection in cloud data center. Concurrency and Computation: Practice and Experience, 27(18):5623–5639, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2011:BDR

[CZ11] Xue Chen and Junfeng Zhang. Building decentralized resource space on a structured P2P network. Concurrency and Computation: Practice and Experience, 23(9):866–879, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2015:AIQ

[CZ15a] Jingqiang Chen and Hai Zhuge. Aimed information quantity in text. Concurrency and Computation: Practice and Experience, 27(15):3982–4000, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2015:OMU

[CZ15b] Rong Cheng and Fangguo Zhang. Obfuscation for multi-use re-encryption and its application in cloud computing. Concurrency and Computation: Practice and Experience, 27(8): 2170–2190, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2019:AGR

[CZ19] Jingqiang Chen and Hai Zhuge. Automatic generation of related work through summarizing citations. Concurrency and Computation: Practice and Experience, 31(3):e4261:1– e4261:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2016:ADA

[CZG16] Quan Chen, Long Zheng, and Minyi Guo. Adaptive demandaware work-stealing in multi-programmed multi-core architectures. Concurrency and Computation: Practice and Experience, 28(2):455–471, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2019:ROC

[CZJY19] Wenting Chen, Qian Zhang, Maozhu Jin, and Jie Yang. Research on online consumer behavior and psychology under the background of big data. Concurrency and Computation: Practice and Experience, 31(10):e4852:1–e4852:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chu:2012:TMP

[CZL12] Xiaowen Chu, Kaiyong Zhao, and Zongpeng Li. Tsunami: massively parallel homomorphic hashing on many-core GPUs. Concurrency and Computation: Practice and Experience, 24 (17):2028–2039, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2017:APS

[CZL^{+}17] Zhenjie Chen, Chen Zhou, Manchun Li, A xing Zhu, Qiuhao Huang, and Yuzhe Pian. Adaptable parallel strategy to extract polygons from massive classified images on multi-core clusters. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chadwick:2008:PMA

[CZO^{+}08] David Chadwick, Gansen Zhao, Sassa Otenko, Romain Laborde, Linying Su, and Tuan Anh Nguyen. PERMIS: a modular authorization infrastructure. Concurrency and Computation: Practice and Experience, 20(11):1341–1357, August 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2017:ADD

[CZQ17] Keyang Cheng, Yongzhao Zhan, and Man Qi. AL-DDCNN: a distributed crossing semantic gap learning for person reidentification. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2007:URM

[CZWH07] Wei Chen, Qingtian Zeng, Liu Wenyin, and Tianyong Hao. A user reputation model for a user-interactive question answering system. Concurrency and Computation: Practice and Experience, 19(15):2091–2103, October 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2018:QES

[CZY^{+}18] Ling Chen, Yuliang Zhao, Yi Yang, Mingqi Lv, Donghui Chen, Yong Wu, and Jingchang Wang. A query execution scheduling scheme for Impala system. Concurrency and Computation: Practice and Experience, 30(8):??, April 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4392.

Dung:2015:SSA

[DA15] Le The Dung and Beongku An. SCRRM: a stability-aware cooperative routing scheme for reliable high-speed data transmission in multi-rate mobile ad hoc wireless networks. Concurrency and Computation: Practice and Experience, 27(10): 2588–2601, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dashdorj:2019:HLE

[DA19] Zolzaya Dashdorj and Erdenebaatar Altangerel. High-level event identification in social media. Concurrency and Computation: Practice and Experience, 31(3):e4668:1–e4668:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deAlencar:2012:PPS

[dAAVS12] João Marcelo U. de Alencar, Rossana M. C. Andrade, Windson Viana, and Bruno Schulze. P2PScheMe: a P2P scheduling mechanism for workflows in grid computing. Concurrency and Computation: Practice and Experience, 24(13): 1478–1496, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dabrowski:2009:RGC

[Dab09a] Christopher Dabrowski. Reliability in grid computing systems. Concurrency and Computation: Practice and Experience, 21(8):927–959, June 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Djoudi:2009:CAA

[DAB09b] L. Djoudi, J.-T. Acquaviva, and D. Barthou. Compositional approach applied to loop specialization. Concurrency and Computation: Practice and Experience, 21(1):71–84, January 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deAssuncao:2008:ICI

[dABV08] Marcos Dias de Assunção, Rajkumar Buyya, and Srikumar Venugopal. InterGrid: a case for internetworking islands of Grids. Concurrency and Computation: Practice and Experience, 20(8):997–1024, June 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Diaz:2012:TCB

[DAC12] Daniel Diaz, Salvador Abreu, and Philippe Codognet. Targeting the Cell Broadband Engine for constraint-based local search. Concurrency and Computation: Practice and Experience, 24(6):647–660, May 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Doerfler:2018:ENC

[DAC^{+}18] Douglas Doerfler, Brian Austin, Brandon Cook, Jack Deslippe, Krishna Kandalla, and Peter Mendygral. Evaluating the networking characteristics of the Cray XC-40 Intel Knights Landing-based Cori supercomputer at NERSC. Concurrency and Computation: Practice and Experience, 30(1): ??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomes:2011:AEE

[dAGC11] Raphael de Aquino Gomes and Fábio M. Costa. An approach to enhance the efficiency of opportunistic grids. Concurrency and Computation: Practice and Experience, 23(17): 2092–2106, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dorre:2015:MOM

[DAL15] Jens Dörre, Sven Apel, and Christian Lengauer. Modeling and optimizing MapReduce programs. Concurrency and Computation: Practice and Experience, 27(7):1734–1766, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Damevski:2011:OEC

[Dam11] Kostadin Damevski. Offline enforcement of contracts for high-performance computing. Concurrency and Computation: Practice and Experience, 23(13):1465–1473, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deAvila:2017:QCS

[dARP17] Anderson B. de Avila, Renata H. S. Reiser, and Mauricio L. Pilla. Quantum computing simulation through reduction and decomposition optimizations with a case study of Shor’s algorithm. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dinan:2016:IEM

[DBB^{+}16] James Dinan, Pavan Balaji, Darius Buntinas, David Goodell, William Gropp, and Rajeev Thakur. An implementation and evaluation of the MPI 3.0 one-sided communication interface. Concurrency and Computation: Practice and Experience, 28(17):4385–4404, December 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeGrande:2016:MDS

[DBGA16] Robson E. De Grande, Azzedine Boukerche, Shichao Guan, and Noura Aljeri. A modular distributed simulation-based architecture for intelligent transportation systems. Concurrency and Computation: Practice and Experience, 28(12): 3409–3426, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Djenouri:2017:RTD

[DBH^{+}17] Youcef Djenouri, Ahcene Bendjoudi, Zineb Habbas, Malika Mehdi, and Djamel Djenouri. Reducing thread divergence in GPU-based bees swarm optimization applied to association rule mining. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeGrande:2013:DRA

[DBR13] Robson Eduardo De Grande, Azzedine Boukerche, and Hussam Ramadan. Distributed re-arrangement scheme for balancing computational load and minimizing communication delays in HLA-based simulations. Concurrency and Computation: Practice and Experience, 25(5):626–648, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeRose:2019:FSI

[DC19a] César A. F. De Rose and Márcio Castro. Foreword to the special issue of the workshop on high performance computing systems (XVIII Simpósio em Sistemas Computacionais de Alto Desempenho, WSCAD 2017). Concurrency and Computation: Practice and Experience, 31(18):e5319:1–e5319:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dongsheng:2019:DCM

[DC19b] Wang Dongsheng and Huang Chuanhe. Distributed cache memory data migration strategy based on cloud computing. Concurrency and Computation: Practice and Experience, 31 (10):e4828:1–e4828:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dutra:2017:EVS

[DCA17] Diego L. C. Dutra, Edilson C. Corrêa, and Claudio L. Amorim. An efficient virtual system clock for the wireless Raspberry Pi computer platform. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2018:SBR

[DCCZ18] Caichang Ding, Yiqin Chen, Lin Chen, and Baorong Zhong. Self-balancing robot bionic intelligence multi-dimensional decision-making evaluation algorithm. Concurrency and Computation: Practice and Experience, 30(22):e4649:1– e4649:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DAgostino:2014:CAM

[DCD^{+}14] Daniele D’Agostino, Andrea Clematis, Sergio Decherchi, Walter Rocchia, Luciano Milanesi, and Ivan Merelli. CUDA accelerated molecular surface generation. Concurrency and Computation: Practice and Experience, 26(10):1819–1831, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Davoust:2015:PPP

[DCEK15] Alan Davoust, Alexander Craig, Babak Esfandiari, and Vincent Kazmierski. P2Pedia: a peer-to-peer wiki for decentralized collaboration. Concurrency and Computation: Practice and Experience, 27(11):2778–2795, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dou:2008:CRD

[DCFC08] Wanchun Dou, Jinjun Chen, Shaokun Fan, and S. C. Chueng. A context- and role-driven scientific workflow development pattern. Concurrency and Computation: Practice and Experience, 20(15):1741–1757, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DAgostino:2011:PIE

[DCG11] D. D’Agostino, A. Clematis, and V. Gianuzzi. Parallel isosurface extraction for 3D data analysis workflows in distributed environments. Concurrency and Computation: Practice and Experience, 23(11):1284–1310, August 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duncan:2015:OIA

[DCG15] Adrian Duncan, Sadie Creese, and Michael Goldsmith. An overview of insider attacks in cloud computing. Concurrency and Computation: Practice and Experience, 27(12): 2964–2981, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deCamargo:2006:CBP

[dCGKG06] Raphael Y. de Camargo, Andrei Goldchleger, Fabio Kon, and Alfredo Goldman. Checkpointing BSP parallel applications on the InteGrade Grid middleware. Concurrency and Computation: Practice and Experience, 18(6):567–579, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

delCid:2012:AML

[dCHMJ12] Pedro Javier del Cid, Danny Hughes, Sam Michiels, and Wouter Joosen. Applying a metadata level for concurrency in wireless sensor networks. Concurrency and Computation: Practice and Experience, 24(16):1953–1962, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dang:2012:EGP

[DCJ12] Duy Minh Dang, Christina C. Christara, and Kenneth R. Jackson. An efficient graphics processing unit-based parallel algorithm for pricing multi-asset American options. Concurrency and Computation: Practice and Experience, 24(8):

849–866, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dang:2014:GPU

[DCJ14] Duy Minh Dang, Christina C. Christara, and Kenneth R. Jackson. Graphics processing unit pricing of exotic crosscurrency interest rate derivatives with a foreign exchange volatility skew model. Concurrency and Computation: Practice and Experience, 26(9):1609–1625, June 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dixon:2012:AVR

[DCK12] M. F. Dixon, J. Chong, and K. Keutzer. Accelerating Valueat-Risk estimation on highly parallel architectures. Concurrency and Computation: Practice and Experience, 24(8):895– 907, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dikbayir:2018:FMR

[DÇK^{+}18] Doga Dikbayir, Enis Berk Çoban, Ilker Kesen, Deniz Yuret, and Didem Unat. Fast multidimensional reduction and broadcast operations on GPU for machine learning. Concurrency and Computation: Practice and Experience, 30(21):e4691:1– e4691:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2017:NAS

[DCP^{+}17] Zhijun Ding, Yuchen Chen, MeiQin Pan, Xiaolun Li, and Pengwei Wang. Network-aware service composition in mobile environment. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Campos:2013:EAM

[dCPD13] Arion de Campos, Jr., Aurora T. R. Pozo, and Elias P. Duarte, Jr. Evaluation of asynchronous multi-swarm particle optimization on several topologies. Concurrency and Computation: Practice and Experience, 25(8):1057–1071, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deCamargo:2011:MGA

[dCRS11] Raphael Y. de Camargo, Luiz Rozante, and Siang W. Song. A multi-GPU algorithm for large-scale neuronal networks. Concurrency and Computation: Practice and Experience, 23(6): 556–572, April 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Di:2008:DGD

[DCY^{+}08] Liping Di, Aijun Chen, Wenli Yang, Yang Liu, Yaxing Wei, Piyush Mehrotra, Chaumin Hu, and Dean Williams. The development of a geospatial data Grid by integrating OGC Web services with Globus-based Grid technology. Concurrency and Computation: Practice and Experience, 20(14): 1617–1635, September 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

De:2016:MDM

[DD16] Suddhasil De and Sohini De. Modeling decoupled mobile cloud computing using Mobile UNITY. Concurrency and Computation: Practice and Experience, 28(10):2811–2855, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DAmato:2017:QBM

[DD17] André D’Amato and Mario Dantas. A QoC-based model for performance and QoE trade-off in distributed systems. Concurrency and Computation: Practice and Experience, 29(18): ??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DiasdaCunha:2001:PNS

[DdB01] Rudnei Dias da Cunha and Álvaro Luiz de Bortoli. A parallel Navier–Stokes solver for the rotating flow problem. Concurrency and Computation: Practice and Experience, 13(3): 163–180, March 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/78003112/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=78003112&PLACEBO=IE.pdf.

Dixon:2012:SIW

[DDE^{+}12] M. F. Dixon, D. Daly, M. Eleftheriou, J. Moreira, and K. Ryu. Special Issue for the Workshop on High Performance Computational Finance. Concurrency and Computation: Practice and Experience, 24(8):833–836, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Donfack:2015:SRD

[DDF^{+}15] Simplice Donfack, Jack Dongarra, Mathieu Faverge, Mark Gates, Jakub Kurzak, Piotr Luszczek, and Ichitaro Yamazaki. A survey of recent developments in parallel implementations of Gaussian elimination. Concurrency and Computation: Practice and Experience, 27(5):1292–1309, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DarosFreitas:2016:ITA

[DDF16] Fabio Daros Freitas, Christian Daros Freitas, and Alberto Ferreira De Souza. Intelligent trading architecture. Concurrency and Computation: Practice and Experience, 28(3):929– 943, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dolz:2017:ESI

[DDF^{+}17] Manuel F. Dolz, David Del Rio Astorga, Javier Fernández, Massimo Torquati, José Daniel García, Félix García-Carballeira, and Marco Danelutto. Enabling semantics to improve detection of data races and misuses of lock-free data structures. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeMatteis:2016:CSQ

[DDM16] Tiziano De Matteis, Salvatore Di Girolamo, and Gabriele Mencagli. Continuous skyline queries on multicore architectures. Concurrency and Computation: Practice and Experience, 28(12):3503–3522, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DaSilva:2006:APQ

[DDP^{+}06] V. F. V. Da Silva, M. L. Dutra, F. Porto, B. Schulze, A. C. Barbosa, and J. C. de Oliveira. An adaptive parallel query processing middleware for the Grid. Concurrency and Computation: Practice and Experience, 18(6):621–634, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2006:NNO

[DDX^{+}06] Zhiqun Deng, Guanzhong Dai, Ting Xie, Daowu Zhou, Dejun Mu, and Xinjia Zhang. A new node organization mechanism of a campus Grid and performance analysis. Concurrency and Computation: Practice and Experience, 18(14): 1787–1798, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Desertot:2007:TAA

[DED07] Mikael Desertot, Clement Escoffier, and Didier Donsez. Towards an autonomic approach for edge computing. Concurrency and Computation: Practice and Experience, 19(14): 1901–1916, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Di:2008:LCI

[DEF08] Pengfei Di, Johannes Eickhold, and Thomas Fuhrmann. Linyphi: creating IPv6 mesh networks with SSR. Concurrency and Computation: Practice and Experience, 20(6):675– 691, April 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DellAmico:2008:NMD

[Del08] Matteo Dell’Amico. Neighbourhood maps: decentralized ranking in small-world P2P networks. Concurrency and Computation: Practice and Experience, 20(6):659–674, April 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Denko:2007:MMH

[Den07] M. K. Denko. Mobility management in hybrid ad-hoc networks and the Internet environment. Concurrency and Computation: Practice and Experience, 19(8):1193–1206, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeSouza:2012:FLP

[DFC12] Alberto F. De Souza, Fabio Daros Freitas, and André Gustavo Coelho de Almeida. Fast learning and predicting of stock returns with virtual generalized random access memory weightless neural networks. Concurrency and Computation: Practice and Experience, 24(8):921–933, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Figueiredo:2019:MOP

[dFdOSR^{+}19] Marco Antonio C. de Figueiredo, Jr., Edans F. de Oliveira Sandes, Genaina N. Rodrigues, George L. M. Teodoro, and Alba Cristina M. A. de Melo. MASA-OpenCL: Parallel pruned comparison of long DNA sequences with OpenCL. Concurrency and Computation: Practice and Experience, 31(11): e5039:1–e5039:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dick:2017:SLM

[DFG17] Ian Dick, Alan Fekete, and Vincent Gramoli. A skip list for multicore. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DAngelo:2018:DHS

[DFG18a] Gabriele D’Angelo, Stefano Ferretti, and Vittorio Ghini. Distributed hybrid simulation of the Internet of things and smart territories. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4370.

DeMaio:2018:SMM

[DFG^{+}18b] Carmen De Maio, Giuseppe Fenza, Mariacristina Gallo, Vincenzo Loia, and Mimmo Parente. Social media marketing through time-aware collaborative filtering. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dessart:2010:DDW

[DFH10] Nathalie Dessart, Hacène Fouchal, and Philippe Hunel. Distributed diagnosis over wireless sensors networks. Concurrency and Computation: Practice and Experience, 22(10): 1240–1251, July 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dongarra:2014:ANA

[DFLL14] Jack Dongarra, Mathieu Faverge, Hatem Ltaief, and Piotr Luszczek. Achieving numerical accuracy and high performance using recursive tile LU factorization with partial pivoting. Concurrency and Computation: Practice and Experience, 26(7):1408–1431, May 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Del-Fabbro:2007:DSM

[DFLNP07] Bruno Del-Fabbro, David Laiymani, Jean-Marc Nicod, and Laurent Philippe. DTM: a service for managing data persistency and data replication in network-enabled server environments. Concurrency and Computation: Practice and Experience, 19(16):2125–2140, November 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sampaio:2006:MMP

[dFMSPSW06] Sandra de F. Mendes Sampaio, Norman W. Paton, Jim Smith, and Paul Watson. Measuring and modelling the performance of a parallel ODMG compliant object database server. Concurrency and Computation: Practice and Experience, 18(1):63–109, January 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DiStefano:2006:SRR

[DFPT06] Antonella Di Stefano, Marco Fargetta, Giuseppe Pappalardo, and Emiliano Tramontana. Supporting resource reservation and allocation for unaware applications in Grid systems. Concurrency and Computation: Practice and Experience, 18(8): 851–863, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Davidovic:2018:ACF

[DFTHD18] Davor Davidović, Diego Fabregat-Traver, Markus Höhnerbach, and Edoardo Di Napoli. Accelerating the computation of FLAPW methods on heterogeneous architectures. Concurrency and Computation: Practice and Experience, 30(24): e4905:1–e4905:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dias:2011:GPU

[DG11] Sérgio E. D. Dias and Abel J. P. Gomes. Graphics processing unit-based triangulations of Blinn molecular surfaces. Concurrency and Computation: Practice and Experience, 23(17): 2280–2291, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeRoure:2010:TOS

[DGA^{+}10] David De Roure, Carole Goble, Sergejs Aleksejevs, Sean Bechhofer, Jiten Bhagat, Don Cruickshank, Paul Fisher, Duncan Hull, Danius Michaelides, David Newman, Rob Procter, Yuwei Lin, and Meik Poschen. Towards open science: the myExperiment approach. Concurrency and Computation: Practice and Experience, 22(17):2335–2353, December 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Davies:2011:SPM

[DGJ11] Christopher J. Davies, David Gubbins, and Peter K. Jimack. Scalability of pseudospectral methods for geodynamo simulations. Concurrency and Computation: Practice and Experience, 23(1):38–56, January 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2012:MSN

[DGL^{+}12] Zeliu Ding, Deke Guo, Xue Liu, Xueshan Luo, and Guihai Chen. A MapReduce-supported network structure for data centers. Concurrency and Computation: Practice and Experience, 24(12):1271–1295, August 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeSalve:2018:PAU

[DGM18] Andrea De Salve, Barbara Guidi, and Paolo Mori. Predicting the availability of users’ devices in decentralized online social networks. Concurrency and Computation: Practice and Experience, 30(20):e4390:1–e4390:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Diaz:2007:CBN

[DGR^{+}07] Manuel Díaz, Daniel Garrido, Sergio Romero, Bartolomé Rubio, Enrique Soler, and José M. Troya. A component-based nuclear power plant simulator kernel. Concurrency and Computation: Practice and Experience, 19(5):593–607, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2016:MWT

[DGW16] Gang Deng, Zhenghu Gong, and Hong Wang. Make-way: transporting latency-sensitive flows nonblockingly in oversubscription data center networks. Concurrency and Computation: Practice and Experience, 28(14):3803–3813, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2013:SIP

[DH13] Hai Dong and Farookh Khadeer Hussain. Special issue papers: SOF: a semi-supervised ontology-learning-based focused crawler. Concurrency and Computation: Practice and Experience, 25(12):1755–1770, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dooley:2015:RBT

[DH15] Rion Dooley and Matthew R. Hanlon. Recipes 2.0: building for today and tomorrow. Concurrency and Computation: Practice and Experience, 27(2):258–270, February 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2011:CAS

[DHC11] Hai Dong, Farookh Khadeer Hussain, and Elizabeth Chang. A context-aware semantic similarity model for ontology environments. Concurrency and Computation: Practice and Experience, 23(5):505–524, April 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2013:SWS

[DHC13] Hai Dong, Farookh Khadeer Hussain, and Elizabeth Chang. Semantic Web Service matchmakers: state of the art and challenges. Concurrency and Computation: Practice and Experience, 25(7):961–988, May 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2013:KSB

[DHH^{+}13] Wei Ding, Chung-Hsing Hsu, Oscar Hernandez, Barbara Chapman, and Richard Graham. KLONOS: Similarity-based planning tool support for porting scientific applications. Concurrency and Computation: Practice and Experience, 25(8): 1072–1088, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dice:2014:SBC

[DHM14] Dave Dice, Danny Hendler, and Ilya Mirsky. Software-based contention management for efficient compare-and-swap operations. Concurrency and Computation: Practice and Experience, 26(14):2386–2404, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Datta:2003:SST

[DHV03] A. K. Datta, R. Hadid, and V. Villain. A self-stabilizing token-based k-out-of-ℓ exclusion algorithm. Concurrency and Computation: Practice and Experience, 15(11–12):1069– 1091, September 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dikaiakos:2007:GBV

[Dik07] Marios D. Dikaiakos. Grid benchmarking: vision, challenges, and current status. Concurrency and Computation: Practice and Experience, 19(1):89–105, January 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dharanipragada:2014:GMR

[DIK14] Janakiram Dharanipragada, Geeta Iyer, and Sriram Kailasam. Generate-map-reduce: an extension to map-reduce to support shared data and recursive computations. Concurrency and Computation: Practice and Experience, 26(2):561–585, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Das:2018:KDP

[DIM18] Anwesha Das, Arun Iyengar, and Frank Mueller. Key-ValueServe+: Design and performance analysis of a multitenant data grid as a cloud service. Concurrency and Computation: Practice and Experience, 30(14):??, July 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4424.

Diniz:2009:GE

[Din09] Pedro C. Diniz. Guest editorial. Concurrency and Computation: Practice and Experience, 21(1):1–3, January 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2019:BPP

[DJ19] Yunqi Dong and Wenjun Jiang. Brand purchase prediction based on time-evolving user behaviors in e-commerce. Concurrency and Computation: Practice and Experience, 31 (1):e4882:1–e4882:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

David:2012:SSA

[DJM12] Tudor David, Mathias Jacquelin, and Loris Marchal. Scheduling streaming applications on a complex multicore platform. Concurrency and Computation: Practice and Experience, 24 (15):1726–1750, October 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deitos:2009:PSL

[DK09] Rafael Deitos and Florian Kerschbaum. Parallelizing secure linear programming. Concurrency and Computation: Practice and Experience, 21(10):1321–1350, July 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devapriya:2019:ESA

[DK19a] S. Deva priya and N. Kannan. Enhanced spectrum aggregation based frequency-band selection routing protocol for cognitive radio ad-hoc networks. Concurrency and Computation: Practice and Experience, 31(14):e4911:1–e4911:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drias:2019:DWI

[DK19b] Yassine Drias and Samir Kechid. Dynamic Web information foraging using self-interested agents: Application to scientific citations network. Concurrency and Computation: Practice and Experience, 31(22):e4342:1–e4342:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drira:2013:ETO

[DKJ13] Khalil Drira, Ahmed Hadj Kacem, and Mohamed Jmaiel. Editorial: Towards the optimal synchronization granularity for dynamic scheduling of pipelined computations on heterogeneous computing systems. Concurrency and Computation: Practice and Experience, 25(2):159–160, February 2013. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drira:2016:ESI

[DKJ16] Khalil Drira, Ahmed Hadj Kacem, and Mohamed Jmaiel. Errata: Special issue editorial: New technologies of distributed systems. Concurrency and Computation: Practice and Experience, 28(9):2749, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deelman:2006:WMW

[DKKL06] Ewa Deelman, Tevfik Kosar, Carl Kesselman, and Miron Livny. What makes workflows work in an opportunistic environment? Concurrency and Computation: Practice and Experience, 18(10):1187–1199, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Diesendruck:2014:ULI

[DKMM14] Liana Diesendruck, Rob Kooper, Luigi Marini, and Kenton McHenry. Using Lucene to index and search the digitized 1940 US Census. Concurrency and Computation: Practice and Experience, 26(13):2167–2177, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dolkas:2007:BAG

[DKMV07] Konstantinos Dolkas, Dimosthenis Kyriazis, Andreas Menychtas, and Theodora Varvarigou. e-Business applications on the Grid: a toolkit for centralized workload prediction and access. Concurrency and Computation: Practice and Experience, 19(6):867–883, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drozdowski:2007:MID

[DL07] M. Drozdowski and M. Lawenda. Multi-installment divisible load processing in heterogeneous distributed systems. Concurrency and Computation: Practice and Experience, 19(17): 2237–2253, December 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dickens:2010:HPI

[DL10] Phillip M. Dickens and Jeremy Logan. A high performance implementation of MPI-IO for a Lustre file system environment. Concurrency and Computation: Practice and Experience, 22(11):1433–1449, August 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2017:ERA

[DL17] Fang Dong and Junzhou Luo. Editorials: Recent advances in big data analysis and application. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dietrich:2001:MTO

[DLH01] F. Dietrich, X. Logean, and J.-P. Hubaux. Modeling and testing object-oriented distributed systems with linear-time temporal logic. Concurrency and Computation: Practice and Experience, 13(5):385–420, April 25, 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/78505737/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
78505737&PLACEBO=IE.pdf.

Dai:2015:DMN

[DLJ15] Hongjun Dai, Huabo Liu, and Zhiping Jia. Dynamic malicious node detection with semi-supervised multivariate classification in cognitive wireless sensor networks. Concurrency and Computation: Practice and Experience, 27(12):2910– 2923, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2018:IMD

[DLK^{+}18] Wenqian Dong, Kenli Li, Letian Kang, Zhe Quan, and Keqin Li. Implementing molecular dynamics simulation on the Sunway TaihuLight system with heterogeneous many-core processors. Concurrency and Computation: Practice and Experience, 30(16):??, August 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4468.

Do:2013:SIP

[DLM13] Hiep-Thuan Do, Sébastien Limet, and Emmanuel Melin. Special issue papers: A scalable parallel minimum spanning tree algorithm for catchment basin delimitation in large digital elevation models. Concurrency and Computation: Practice and Experience, 25(10):1394–1409, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dongarra:2003:LBP

[DLP03] Jack J. Dongarra, Piotr Luszczek, and Antoine Petitet. The LINPACK Benchmark: past, present and future. Concurrency and Computation: Practice and Experience, 15(9):803– 820, August 10, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dainotti:2007:SPE

[DLPV07] Alberto Dainotti, Salvatore Loreto, Antonio Pescapé, and Giorgio Ventre. SCTP performance evaluation over heterogeneous networks. Concurrency and Computation: Practice and Experience, 19(8):1207–1218, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duan:2016:SRA

[DLT^{+}16] Mingxing Duan, Kenli Li, Zhuo Tang, Guoqing Xiao, and Keqin Li. Selection and replacement algorithms for memory performance improvement in Spark. Concurrency and Computation: Practice and Experience, 28(8):2473–2486, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dou:2019:EAT

[DLW19] Wanfeng Dou, Yanan Li, and Yanli Wang. An equal-area triangulated partition method for parallel Xdraw viewshed analysis. Concurrency and Computation: Practice and Experience, 31(17):e5216:1–e5216:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2016:MBP

[DLX^{+}16] Chuang Deng, Yang Liu, Lixiong Xu, Jie Yang, Junyong Liu, Siguang Li, and Maozhen Li. A MapReduce-based parallel K-means clustering for large-scale CIM data verification. Concurrency and Computation: Practice and Experience, 28 (11):3096–3114, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dixon:2016:PEF

[DLZ16] Matthew Dixon, Jörg Lotze, and Mohammad Zubair. A portable, extensible and fast stochastic volatility model calibration using multi and many-core processors. Concurrency and Computation: Practice and Experience, 28(3):866–877, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2015:ERR

[DM15] Fang Dong and Alisha Malloy. Editorials: Recent research advances in cloud computing and big data. Concurrency and Computation: Practice and Experience, 27(18):5574–5576, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Djouama:2013:PEL

[DMA13] Amir Djouama, Lynda Mokdad, and Marwen Abdennebi. Performance evaluation of lifetime-driven admission control for infrastructure-less clustered wireless networks. Concurrency and Computation: Practice and Experience, 25(5):718– 727, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2018:AMD

[DMC^{+}18] Zengyu Ding, Gang Mei, Salvatore Cuomo, Hong Tian, and Nengxiong Xu. Accelerating multi-dimensional interpolation using moving least-squares on the GPU. Concurrency and Computation: Practice and Experience, 30(24):e4904:1– e4904:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dixon:2016:EWH

[DMD16] Matthew Dixon, Jose Moreira, and David Daly. Editorials: Workshop on high-performance computational finance. Concurrency and Computation: Practice and Experience, 28(3): 834–835, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deTacioPereiraGomes:2017:CSM

[dMd^{+}17] Berto de Tácio Pereira Gomes, Luiz Carlos Melo Muniz, Francisco José da Silva e Silva, Luis Eduardo Talavera Ríos, and Markus Endler. A comprehensive and scalable middleware for Ambient Assisted Living based on cloud computing and Internet of Things. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dolgobrodov:2007:SAN

[DMM^{+}07] S. D. Dolgobrodov, R. Marshall, P. Moore, R. Bittern, R. J. C. Steele, and A. Cuschieri. e-science and artificial neural networks in cancer management. Concurrency and Computation: Practice and Experience, 19(2):251–263, February 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duenha:2017:EPD

[DMMA17] Liana Duenha, Guilherme Madalozzo, Fernando Gehm Moraes, and Rodolfo Azevedo. Exploiting performance, dynamic power and energy scaling in full-system simulators. Concurrency and Computation: Practice and Experience, 29 (22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DiMartino:2007:CSS

[DMR^{+}07] Beniamino Di Martino, Emilio Mancini, Massimiliano Rak, Roberto Torella, and Umberto Villano. Cluster systems and simulation: from benchmarking to off-line performance prediction. Concurrency and Computation: Practice and Experience, 19(11):1549–1562, August 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeMeo:2015:AOT

[DMRS15] Pasquale De Meo, Fabrizio Messina, Domenico Rosaci, and Giuseppe M. L. Sarné. An agent-oriented, trust-aware approach to improve the QoS in dynamic grid federations. Concurrency and Computation: Practice and Experience, 27(17): 5411–5435, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DiRienzo:2010:CCC

[DMW^{+}10] A. DiRienzo, John A. Medeiros, M. Whitlock, E. Wages, and J. Highfield. Concepts for computer center power management. Concurrency and Computation: Practice and Experience, 22(2):145–156, February 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dahan:2019:SIX

[DN19] Maytal Dahan and Paul Navrátil. Special issue XSEDE16 and PEARC17 — practice and experience in advanced research computing. Concurrency and Computation: Practice and Experience, 31(16):e5325:1–e5325:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Djemaa:2019:ESS

[DNB19] Raoudha Ben Djemaa, Hajer Nabli, and Ikram Amous Ben Amor. Enhanced semantic similarity measure based on two-level retrieval model. Concurrency and Computation: Practice and Experience, 31(15):e5135:1–e5135:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Castro:2013:SIP

[dOCPFJ13] Pablo de Oliveira Castro, Eric Petit, Asma Farjallah, and William Jalby. Special issue papers: Adaptive sampling for performance characterization of application kernels. Concurrency and Computation: Practice and Experience, 25(17): 2345–2362, December 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2019:ESS

[DOJ^{+}19] Xiaofeng Ding, Yangling Ou, Jianhong Jia, Hai Jin, and Jixue Liu. Efficient subgraph search on large anonymized graphs. Concurrency and Computation: Practice and Experience, 31(23):e4511:1–e4511:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deOliveira:2012:APE

[dOOO^{+}12] Daniel de Oliveira, Eduardo Ogasawara, Kary Ocaña, Fernanda Baião, and Marta Mattoso. An adaptive parallel execution strategy for cloud-based scientific workflows. Concurrency and Computation: Practice and Experience, 24(13): 1531–1550, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

dAuriol:2006:CPP

[dP06] Brian J. d’Auriol and Jesus Pajaro. A comprehensive peerto-peer characterization model. Concurrency and Computation: Practice and Experience, 18(15):2037–2061, December 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Doallo:2014:EMC

[DP14] Ramón Doallo and Óscar Plata. Editorials: Multicore cache hierarchies: design and programmability issues. Concurrency and Computation: Practice and Experience, 26(6):1326–1327, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Do:2019:LLC

[DP19] Thanh-Nghi Do and François Poulet. Latent-lSVM classification of very high-dimensional and large-scale multi-class datasets. Concurrency and Computation: Practice and Experience, 31(2):e4224:1–e4224:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Djemame:2011:BRA

[DPGA11] K. Djemame, J. Padgett, I. Gourlay, and D. Armstrong. Brokering of risk-aware service level agreements in grids. Concurrency and Computation: Practice and Experience, 23(13): 1558–1582, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Docan:2010:EHS

[DPK10] Ciprian Docan, Manish Parashar, and Scott Klasky. Enabling high-speed asynchronous data extraction and transfer using DART. Concurrency and Computation: Practice and Experience, 22(9):1181–1204, June 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dib:2017:SBP

[DPM17] Djawida Dib, Nikos Parlavantzas, and Christine Morin. SLAbased PaaS profit optimization. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Denis:2003:APE

[DPP03] Alexandre Denis, Christian Pérez, and Thierry Priol. Achieving portable and efficient parallel CORBA objects. Concurrency and Computation: Practice and Experience, 15(10): 891–909, August 25, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duro:2019:MAP

[DPP^{+}19] José Duro, Jose A. Pascual, Salvador Petit, Julio Sahuquillo, and María E. Gómez. Modeling and analysis of the performance of exascale photonic networks. Concurrency and Computation: Practice and Experience, 31(21):e4773:1–e4773:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Distefano:2007:GBA

[DPS07] S. Distefano, A. Puliafito, and M. Scarpa. A Grid-based algorithm for the solution of non-Markovian stochastic Petri nets. Concurrency and Computation: Practice and Experience, 19(9):1353–1370, June 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Diamantini:2016:EDO

[DPS16] Claudia Diamantini, Domenico Potena, and Emanuele Storti. Extended drill-down operator: Digging into the structure of performance indicators. Concurrency and Computation: Practice and Experience, 28(15):3948–3968, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DiStefano:2006:TIA

[DPST06] Antonella Di Stefano, Giuseppe Pappalardo, Corrado Santoro, and Emiliano Tramontana. The transparent implementation of agent communication contexts. Concurrency and Computation: Practice and Experience, 18(4):387–407, April 10, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drira:2015:EIC

[DR15] Khalil Drira and Sumitra Reddy. Editorials: International conference on enabling technologies: Infrastructure for collaborative enterprises (WETICE). Concurrency and Computation: Practice and Experience, 27(6):1375–1377, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drabble:2015:DBC

[Dra15] Brian Drabble. Dependency-based collaborative design: a comparison of modeling methods. Concurrency and Computation: Practice and Experience, 27(11):2686–2705, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dragomir:2017:REE

[Dra17] Dan Dragomir. Reducing EMF exposure from WSNs using transmission scheduling. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Astorga:2017:GPP

[dRADFG17] David del Rio Astorga, Manuel F. Dolz, Javier Fernández, and J. Daniel García. A generic parallel pattern interface for stream and data processing. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

dosReis:2010:CPU

[dRC10] Valéria Quadros dos Reis and Renato Cerqueira. Controlling processing usage at user level: a way to make resource sharing more flexible. Concurrency and Computation: Practice and Experience, 22(3):278–294, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dumitrescu:2007:USB

[DRF07] Catalin L. Dumitrescu, Ioan Raicu, and Ian Foster. Usage SLA-based scheduling in Grids. Concurrency and Computation: Practice and Experience, 19(7):945–963, May 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

dOrazio:2010:ACS

[dRL10] Laurent d’Orazio, Claudia Roncancio, and Cyril Labbé. Adaptable cache service and application to grid caching. Concurrency and Computation: Practice and Experience, 22(9): 1118–1137, June 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Righi:2016:JAP

[dRRdCRR16] Rodrigo da Rosa Righi, Cristiano André da Costa, Vinicius Facco Rodrigues, and Gustavo Rostirolla. Joint-analysis of performance and energy consumption when enabling cloud elasticity for synchronous HPC applications. Concurrency and Computation: Practice and Experience, 28(5):1548–1571, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2013:SIP

[DRS^{+}13] Kefeng Deng, Kaijun Ren, Junqiang Song, Dong Yuan, Yang Xiang, and Jinjun Chen. Special issue papers: A clustering based coscheduling strategy for efficient scientific workflow execution in cloud computing. Concurrency and Computation: Practice and Experience, 25(18):2523–2539, December 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dasari:2013:SIP

[DRZ13] Naga Shailaja Dasari, Desh Ranjan, and Mohammad Zubair. Special issue papers: High-performance implementation of planted motif problem on multicore and GPU. Concurrency and Computation: Practice and Experience, 25(10): 1340–1355, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Denissen:2002:FPB

[DS02] Will Denissen and Henk J. Sips. Finding performance bugs with the TNO HPF benchmark suite. Concurrency and Computation: Practice and Experience, 14(8–9):691–712, July/ August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016124/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016124{\&}PLACEBO=IE.pdf.

Desprez:2004:IMP

[DS04] F. Desprez and F. Suter. Impact of mixed-parallelism on parallel implementations of the Strassen and Winograd matrix multiplication algorithms. Concurrency and Computation: Practice and Experience, 16(8):771–797, July 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DiStefano:2007:PPD

[DS07] Antonella Di Stefano and Corrado Santoro. A peer-to-peer decentralized strategy for resource management in computational Grids. Concurrency and Computation: Practice and Experience, 19(9):1271–1286, June 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dressler:2015:AAE

[DS15] Sebastian Dreßler and Thomas Steinke. An automated approach for estimating the memory footprint of nonlinear data objects. Concurrency and Computation: Practice and Experience, 27(6):1564–1574, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2017:ENT

[DS17] Fang Dong and Jun Shen. Editorial: New trends and innovative methods in cloud computing and big data. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devi:2019:HMO

[DS19] S. Gayathri Devi and M. Sabrigiriraj. A hybrid multiobjective firefly and simulated annealing based algorithm for big data classification. Concurrency and Computation: Practice and Experience, 31(14):e4985:1–e4985:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

daSilva:2014:CFT

[dSGD14] Rafael Ferreira da Silva, Tristan Glatard, and Frédéric Desprez. Controlling fairness and task granularity in distributed, online, non-clairvoyant workflow executions. Concurrency and Computation: Practice and Experience, 26(14):2347– 2366, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2018:ACC

[DSH18] Fang Dong, Jun Shen, and Qiang He. Advances in cloud computing and big data analytics. Concurrency and Computation: Practice and Experience, 30(20):e4960:1–e4960:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Davoust:2015:DWS

[DSMM^{+}15] Alan Davoust, Hala Skaf-Molli, Pascal Molli, Babak Esfandiari, and Khaled Aslan. Distributed wikis: a survey. Concurrency and Computation: Practice and Experience, 27(11): 2751–2777, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Damas:2001:PDW

[DSO^{+}01] M. Damas, M. Salmerón, J. Ortega, G. Olivares, and H. Pomares. Parallel dynamic water supply scheduling in a cluster of computers. Concurrency and Computation: Practice and Experience, 13(15):1281–1302, December 25, 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/89016297/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
89016297&PLACEBO=IE.pdf.

DeFalco:2011:DEA

[DST11] I. De Falco, U. Scafuri, and E. Tarantino. A distributed evolutionary approach for multisite mapping on grids. Concurrency and Computation: Practice and Experience, 23(10):

1146–1168, July 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dubey:2001:RSP

[DT01] Anshu Dubey and Daniele Tessera. Redistribution strategies for portable parallel FFT: a case study. Concurrency and Computation: Practice and Experience, 13(3):209–220, March 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/78003114/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
78003114&PLACEBO=IE.pdf.

Dorn:2015:ARA

[DT15a] Christoph Dorn and Richard N. Taylor. Analyzing runtime adaptability of collaboration patterns. Concurrency and Computation: Practice and Experience, 27(11):2725–2750, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dundjerski:2015:GPU

[DT15b] Dejan Dundjerski and Milo Tomašević. Graphical processing unit-based parallelization of the Open Shortest Path First and Border Gateway Protocol routing protocols. Concurrency and Computation: Practice and Experience, 27(1):237–251, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Donnelly:2017:BPP

[DT17] Patrick Donnelly and Douglas Thain. Balancing push and pull in Confuga, an active storage cluster file system for scientific workflows. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2018:RID

[Du18a] Hongchun Du. Research on image de-disturbing algorithm based on dark channel prior and anisotropic Gaussian filtering. Concurrency and Computation: Practice and Experience, 30(24):e4933:1–e4933:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2018:MPR

[Du18b] Meiyan Du. Mobile payment recognition technology based on face detection algorithm. Concurrency and Computation: Practice and Experience, 30(22):e4655:1–e4655:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2018:SII

[Du18c] Zhenyu Du. Special issue on 2017 International Conference of Intelligence Computation and Evolutionary Computation (ICEC2017). Concurrency and Computation: Practice and Experience, 30(22):e4672:1–e4672:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dutta:2017:VPT

[Dut17] Sudakshina Dutta. Validation of parallelizing transformations of sequential programs. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeMunck:2014:RCT

[DVB14] S. De Munck, K. Vanmechelen, and J. Broeckhove. Revisiting conservative time synchronization protocols in parallel and distributed simulation. Concurrency and Computation: Practice and Experience, 26(2):468–490, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Daley:2012:OMB

[DVD^{+}12] Christopher Daley, Marcos Vanella, Anshu Dubey, Klaus Weide, and Elias Balaras. Optimization of multigrid based elliptic solver for large scale simulations in the FLASH code. Concurrency and Computation: Practice and Experience, 24 (18):2346–2361, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dijkstra:2006:ITO

[DvdS06] Fokke Dijkstra and Aad J. van der Steen. Integration of two ocean models within Cactus. Concurrency and Computation: Practice and Experience, 18(2):193–202, February 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dias:2013:SIP

[DVL13] Ricardo J. Dias, Tiago M. Vale, and João M. Lourenço. Special issue papers: Efficient support for in-place metadata in Java software transactional memory. Concurrency and Computation: Practice and Experience, 25(17):2394–2411, December 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drost:2011:JUR

[DvNM^{+}11a] Niels Drost, Rob V. van Nieuwpoort, Jason Maassen, Frank Seinstra, and Henri E. Bal. JEL: unified resource tracking for parallel and distributed applications. Concurrency and Computation: Practice and Experience, 23(1):17–37, January 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drost:2011:ZPP

[DvNM^{+}11b] Niels Drost, Rob V. van Nieuwpoort, Jason Maassen, Frank J. Seinstra, and Henri E. Bal. Zorilla: a peer-to-peer middleware for real-world distributed systems. Concurrency and Computation: Practice and Experience, 23(13):1506–1521, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dillon:2009:ALM

[DWC09] Tharam S. Dillon, Chen Wu, and Elizabeth Chang. An abstract layered model for Web-inclusive distributed computing leading to enhancing GRIDSpace with Web 2.0. Concurrency and Computation: Practice and Experience, 21(5):605–634, April 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2015:CHP

[DWC^{+}15] Xinnan Dong, Mei Wen, Jun Chai, Xing Cai, Mandan Zhao, and Chunyuan Zhang. Communication-hiding programming for clusters with multi-coprocessor nodes. Concurrency and Computation: Practice and Experience, 27(16):4172–4185, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2019:CCB

[DWG19] Fang Dong, Chenshu Wu, and Shangce Gao. Cloud computing-based big data processing and intelligent analytics. Concurrency and Computation: Practice and Experience, 31(24):e5531:1–e5531:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2013:LNX

[DXG13] Zhi-Hong Deng, Yong-Qing Xiang, and Ning Gao. LAF: a new XML encoding and indexing strategy for keyword-based XML search. Concurrency and Computation: Practice and Experience, 25(11):1604–1621, August 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2017:SCS

[DXHL17] Yong Ding, Ning Xiong, En He, and Kezan Li. Secure communication scheme analysis via complex networks. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dou:2017:EAV

[DXM^{+}17] Wanchun Dou, Xiaolong Xu, Shunmei Meng, Xuyun Zhang, Chunhua Hu, Shui Yu, and Jian Yang. An energy-aware virtual machine scheduling method for service QoS enhancement in clouds over big data. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2016:AIF

[DXWC16] Jiang Deng, Chunxiang Xu, Huai Wu, and Jie Chen. Analysis and improvement of a fair remote retrieval protocol for private medical records. Concurrency and Computation: Practice and Experience, 28(6):1918–1929, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2016:NCS

[DXWD16] Jiang Deng, Chunxiang Xu, Huai Wu, and Liju Dong. A new certificateless signature with enhanced security and aggregation version. Concurrency and Computation: Practice and Experience, 28(4):1124–1133, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duan:2016:MCA

[DXZ^{+}16] Hancong Duan, Xiaoke Xiang, Yulu Zhang, Geyong Min, Wenhan Zhan, and Pengcheng Lv. A multi-channel architecture for metadata management in cloud storage systems by binding CPU-cores to disks. Concurrency and Computation: Practice and Experience, 28(16):4215–4229, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2016:SAD

[DYW16] Naiqiao Du, Xiaojun Ye, and Jianmin Wang. A semanticaware data generator for ETL workflows. Concurrency and Computation: Practice and Experience, 28(4):1016–1040, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2019:TES

[DYY^{+}19] Jinyu Du, Jiayu Yi, Zhongming Yang, Yi Yu, Bo Tu, and Xiangchao An. Theoretical and experimental study on the astigmatism self-compensation in multi-disk laser resonator based on exhaustive method. Concurrency and Computation: Practice and Experience, 31(12):e4698:1–e4698:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2013:QAS

[DZ13] Qinghe Du and Chao Zhang. Queuing analyses and statistically bounded delay control for two-hop green wireless relay transmissions. Concurrency and Computation: Practice and Experience, 25(9):1050–1063, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2016:EPP

[DZC16] Xiaolei Dong, Jun Zhou, and Zhenfu Cao. Efficient privacypreserving temporal and spacial data aggregation for smart grid communications. Concurrency and Computation: Practice and Experience, 28(4):1145–1160, March 25, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Docan:2015:AED

[DZJ^{+}15] Ciprian Docan, Fan Zhang, Tong Jin, Hoang Bui, Qian Sun, Julian Cummings, Norbert Podhorszki, Scott Klasky, and Manish Parashar. ActiveSpaces: Exploring dynamic code deployment for extreme scale data processing. Concurrency and Computation: Practice and Experience, 27(14):3724–3745, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2017:EAA

[DZL^{+}17a] Fang Dong, Junxue Zhang, Junzhou Luo, Dian Shen, and Jiahui Jin. Enabling application-aware flexible graph partition mechanism for parallel graph processing systems. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2017:TFS

[DZL^{+}17b] Fang Dong, Pengcheng Zhou, Zijian Liu, Dian Shen, Zhuqing Xu, and Junzhou Luo. Towards a fast and secure design for enterprise-oriented cloud storage systems. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dai:2019:MLF

[DZL^{+}19] Liang Dai, Jia Zhang, Candong Li, Changen Zhou, and Shaozi Li. Multi-label feature selection with application to TCM state identification. Concurrency and Computation: Practice and Experience, 31(23):e4634:1–e4634:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duan:2015:HPD

[DZM^{+}15] Hancong Duan, Wenhan Zhan, Geyong Min, Hui Guo, and Shengmei Luo. A high-performance distributed file system for large-scale concurrent HD video streams. Concurrency and Computation: Practice and Experience, 27(13):3510–3522, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dillon:2011:WTF

[DZW^{+}11] Tharam S. Dillon, Hai Zhuge, Chen Wu, Jaipal Singh, and Elizabeth Chang. Web-of-things framework for cyber– physical systems. Concurrency and Computation: Practice and Experience, 23(9):905–923, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2019:CHP

[DZZL19] Meiya Dong, Jumin Zhao, Biaokai Zhu, and Zhaobin Liu. CHAMELEON: Hides privacy in cloud IoT system by LSB and CSE. Concurrency and Computation: Practice and Experience, 31(24):e5505:1–e5505:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Erguler:2012:PAI

[EA12] Imran Erguler and Emin Anarim. Practical attacks and improvements to an efficient radio frequency identification authentication protocol. Concurrency and Computation: Practice and Experience, 24(17):2069–2080, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Estevez-Ayres:2014:CSS

[EABVGV14] Iria Estévez-Ayres, Pablo Basanta-Val, and Marisol García-Valls. Composing and scheduling service-oriented applications in time-triggered distributed real-time Java environments. Concurrency and Computation: Practice and Experience, 26(1):152–193, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Estevez-Ayres:2011:HAS

[EAGVBVDS11] Iria Estévez-Ayres, Marisol García-Valls, Pablo Basanta-Val, and Jorge Díez-Sánchez. A hybrid approach for selecting service-based real-time composition algorithms in heterogeneous environments. Concurrency and Computation: Practice and Experience, 23(15):1816–1851, October 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eugster:2005:OOP

[EB05] Patrick Th. Eugster and Sebastien Baehni. Object-oriented programming in peer-to-peer systems. Concurrency and Computation: Practice and Experience, 17(7–8):1053–1078, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Escheikh:2010:OML

[EB10] M. Escheikh and K. Barkaoui. Opportunistic MAC layer design with stochastic Petri Nets for multimedia ad hoc networks. Concurrency and Computation: Practice and Experience, 22(10):1308–1324, July 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ellingson:2014:HTV

[EB14] Sally R. Ellingson and Jerome Baudry. High-throughput virtual molecular docking with AutoDockCloud. Concurrency and Computation: Practice and Experience, 26(4):907–916, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ebadifard:2018:PBT

[EB18] Fatemeh Ebadifard and Seyed Morteza Babamir. A PSObased task scheduling algorithm improved using a loadbalancing technique for the cloud computing environment. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4368.

Eugster:2001:EMP

[EBGS01] Patrick Th. Eugster, Romain Boichat, Rachid Guerraoui, and Joe Sventek. Effective multicast programming in large scale distributed systems. Concurrency and Computation: Practice and Experience, 13(6):421–447, May 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/79503112/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
79503112&PLACEBO=IE.pdf.

Emeakaroha:2013:SIP

[EBMD13] Vincent C. Emeakaroha, Ivona Brandic, Michael Maurer, and Schahram Dustdar. Special issue papers: Cloud resource provisioning and SLA enforcement via LoM2HiS framework. Concurrency and Computation: Practice and Experience, 25 (10):1462–1481, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Esposito:2018:ITB

[ECP18] Christian Esposito, Aniello Castiglione, and Francesco Palmieri. Information theoretic-based detection and removal of slander and/or false-praise attacks for robust trust management with Dempster–Shafer combination of linguistic fuzzy terms. Concurrency and Computation: Practice and Experience, 30(3):??, February 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4302.

Ellingson:2014:AVH

[EDB^{+}14] Sally R. Ellingson, Sivanesan Dakshanamurthy, Milton Brown, Jeremy C. Smith, and Jerome Baudry. Accelerating virtual high-throughput ligand docking: current technology and case study on a petascale supercomputer. Concurrency and Computation: Practice and Experience, 26(6):1268–1277, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ernst-Desmulier:2008:PFS

[EDBS08] J.-B. Ernst-Desmulier, Julien Bourgeois, and François Spies. P2PPerf: a framework for simulating and optimizing peer-to-peer-distributed computing applications. Concurrency and Computation: Practice and Experience, 20(6):693–712, April 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ejarque:2010:ESV

[EdPG^{+}10] Jorge Ejarque, Marc de Palol, Íñigo Goiri, Ferran Julià, Jordi Guitart, Rosa M. Badia, and Jordi Torres. Exploiting semantics and virtualization for SLA-driven resource allocation in service providers. Concurrency and Computation: Practice and Experience, 22(5):541–572, April 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

ElMaghraoui:2009:MIM

[EDSV09] K. El Maghraoui, Travis J. Desell, Boleslaw K. Szymanski, and Carlos A. Varela. Malleable iterative MPI applications. Concurrency and Computation: Practice and Experience, 21 (3):393–413, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Edmundsson:2004:DET

[EEK^{+}04] Niklas Edmundsson, Erik Elmroth, Bo Kågström, Markus Mårtensson, Mats Nylén, Åke Sandgren, and Mattias Wadenstein. Design and evaluation of a TOP100 Linux Super Cluster system. Concurrency and Computation: Practice and Experience, 16(8):735–750, July 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ezugwu:2017:NNB

[EFA^{+}17] Absalom E. Ezugwu, Marc E. Frincu, Aderemi O. Adewumi, Seyed M. Buhari, and Sahalu B. Junaidu. Neural networkbased multi-agent approach for scheduling in distributed systems. Concurrency and Computation: Practice and Experience, 29(1):??, January 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Edelstein:2003:FTM

[EFG^{+}03] Orit Edelstein, Eitan Farchi, Evgeny Goldin, Yarden Nir, Gil Ratsaby, and Shmuel Ur. Framework for testing multithreaded Java programs. Concurrency and Computation: Practice and Experience, 15(3–5):485–499, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ertl:2017:DOE

[EFM17] Christoph Ertl, Jérôme Frisch, and Ralf-Peter Mundani. Design and optimisation of an efficient HDF5 I/O Kernel for massive parallel fluid flow simulations. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Engwer:2017:SCP

[EFY17] C. Engwer, R. D. Falgout, and U. M. Yang. Stencil computations for PDE-based applications with examples from DUNE and hypre. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

El-Ghazawi:2004:PSJ

[EGGA^{+}04] Tarek El-Ghazawi, Kris Gaj, Nikitas Alexandridis, Frederic Vroman, Nguyen Nguyen, Jacek R. Radzikowski, Preeyapong Samipagdi, and Suboh A. Suboh. A performance study of job management systems. Concurrency and Computation: Practice and Experience, 16(13):1229–1246, November 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Enokido:2018:SIC

[EH18] Tomoya Enokido and Omar Khadeer Hussain. Special issue on Cloud of Things applications. Concurrency and Computation: Practice and Experience, 30(21):e4879:1–e4879:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eytani:2007:TFB

[EHSU07] Yaniv Eytani, Klaus Havelund, Scott D. Stoller, and Shmuel Ur. Towards a framework and a benchmark for testing tools for multi-threaded programs. Concurrency and Computation: Practice and Experience, 19(3):267–279, March 10, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ezzati-Jivan:2015:CDM

[EJD15] Naser Ezzati-Jivan and Michel R. Dagenais. Cube data model for multilevel statistics computation of live execution traces. Concurrency and Computation: Practice and Experience, 27 (5):1069–1091, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ezzati-Jivan:2017:MSN

[EJD17] Naser Ezzati-Jivan and Michel R. Dagenais. Multi-scale navigation of large trace data: a survey. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ezugwu:2016:SMC

[EJF^{+}16] Absalom E. Ezugwu, Sahalu B. Junaidu, Marc E. Frincu, Seyed M. Buhari, and Afolayan A. Obiniyi. Scheduling multicomponent applications with mobile agents in heterogeneous distributed systems. Concurrency and Computation: Practice and Experience, 28(5):1462–1479, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ernstsson:2019:ESC

[EK19] August Ernstsson and Christoph Kessler. Extending smart containers for data locality-aware skeleton programming. Concurrency and Computation: Practice and Experience, 31 (5):e5003:1–e5003:??, March 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eisenbach:2001:SIF

[EL01] Susan Eisenbach and Gary T. Leavens. Special issue: formal techniques for Java programs. Concurrency and Computation: Practice and Experience, 13(13):1121–1123, November 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/88011339/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
88011339&PLACEBO=IE.pdf.

Eicker:2016:DPA

[ELM^{+}16] Norbert Eicker, Thomas Lippert, Thomas Moschny, Estela Suarez, and for the Deep project. The DEEP Project: An alternative approach to heterogeneous cluster-computing in the many-core era. Concurrency and Computation: Practice and Experience, 28(8):2394–2411, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elsayed:2011:PCR

[EMB11] Ibrahim Elsayed, Gregory Madey, and Peter Brezany. Portals for collaborative research communities: two distinguished case studies. Concurrency and Computation: Practice and Experience, 23(3):269–278, March 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

El-Moursy:2019:ATL

[EME19] A. A. El-Moursy and A. S. Elhelw. Adaptive TB-LMI: an efficient memory controller and scheduler design. Concurrency and Computation: Practice and Experience, 31(7):e5011:1– e5011:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

El-Moursy:2014:HAH

[EMEY14] Ali A. El-Moursy, Hanan ElAzhary, and Akmal Younis. Highaccuracy hierarchical parallel technique for hidden Markov model-based 3D magnetic resonance image brain segmentation. Concurrency and Computation: Practice and Experience, 26(1):194–216, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

El-Moursy:2011:IPA

[EMS11] Ali A. El-Moursy and Fadi N. Sibai. Image processing applications performance study on Cell BE and Blue Gene/L. Concurrency and Computation: Practice and Experience, 23 (4):351–371, March 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Evoy:2015:ADP

[EMS15] Giacomo Mc Evoy, Antonio R. Mury, and Bruno Schulze. An analysis of definition and placement of virtual machines for high performance applications on Clouds. Concurrency and Computation: Practice and Experience, 27(7):1789–1814, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

ElSaddik:2009:GED

[EN09] Abdulmotaleb El Saddik and Mirela Sechi Moretti Annoni Notare. Guest editorial: Distributed simulation, virtual environments and real-time applications. Concurrency and Computation: Practice and Experience, 21(11):1419–1421, August 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Engelmann:2016:NDR

[EN16] Christian Engelmann and Thomas Naughton. A new deadlock resolution protocol and message matching algorithm for the extreme-scale simulator. Concurrency and Computation: Practice and Experience, 28(12):3369–3389, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ebadi:2019:EAM

[EN19] Yalda Ebadi and Nima Jafari Navimipour. An energy-aware method for data replication in the cloud environments using a Tabu search and particle swarm optimization algorithm. Concurrency and Computation: Practice and Experience, 31 (1):e4757:1–e4757:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

English:2015:SME

[Eng15] Rosanne English. Simulating and modelling the effectiveness of graphical password intersection attacks. Concurrency and Computation: Practice and Experience, 27(12):3089–3107, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Escobar:2019:EAL

[EOD^{+}19] Juan José Escobar, Julio Ortega, Antonio Francisco Díaz, Jesús González, and Miguel Damas. Energy-aware load balancing of parallel evolutionary algorithms with heavy fitness functions in heterogeneous CPU–GPU architectures. Concurrency and Computation: Practice and Experience, 31(6): e4688:1–e4688:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ericson:2015:FRR

[EPA15] Kathleen Ericson, Shrideep Pallickara, and Charles W. Anderson. Failure-resilient real-time processing of health streams. Concurrency and Computation: Practice and Experience, 27(7):1695–1717, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eichinger:2014:DMD

[EPB14] Frank Eichinger, Victor Pankratius, and Klemens Böhm. Data mining for defects in multicore applications: an entropybased call-graph technique. Concurrency and Computation: Practice and Experience, 26(1):1–20, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ezzatti:2019:PAC

[EQORS19] Pablo Ezzatti, Enrique S. Quintana-Ortí, Alfredo Remón, and Jens Saak. Power-aware computing. Concurrency and Computation: Practice and Experience, 31(6):e5034:1–e5034:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Emmanuel:2018:COH

[EQW^{+}18] Bugingo Emmanuel, Yingsheng Qin, Juntao Wang, Defu Zhang, and Wei Zheng. Cost optimization heuristics for deadline constrained workflow scheduling on clouds and their comparative evaluation. Concurrency and Computation: Practice and Experience, 30(20):e4762:1–e4762:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

ElZein:2012:GOC

[ER12] Ahmed H. El Zein and Alistair P. Rendell. Generating optimal CUDA sparse matrix–vector product implementations for evolving GPU hardware. Concurrency and Computation: Practice and Experience, 24(1):3–13, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Erwin:2002:UGC

[Erw02] Dietmar W. Erwin. UNICORE — a Grid computing environment. Concurrency and Computation: Practice and Experience, 14(13–15):1395–1410, November/December 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eyers:2011:CLS

[ERZ^{+}11] David M. Eyers, Ramani Routray, Rui Zhang, Douglas Willcocks, and Peter Pietzuch. Configuring large-scale storage using a middleware with machine learning. Concurrency and Computation: Practice and Experience, 23(17):2063–2077, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Evoy:2011:PDE

[ESG11] Giacomo V. Mc Evoy, Bruno Schulze, and Eduardo L. M. Garcia. Performance and deployment evaluation of a parallel application on a private Cloud. Concurrency and Computation: Practice and Experience, 23(17):2048–2062, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Escudero-Sahuquillo:2017:ESI

[ESG17] Jesús Escudero-Sahuquillo and Pedro Javier Garcia. Editorial: Special issue on trends in high-performance interconnection networks in the exascale and big-data era. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Escudero-Sahuquillo:2011:CEQ

[ESGQ^{+}11] J. Escudero-Sahuquillo, P. J. Garcia, F. J. Quiles, J. Flich, and J. Duato. Cost-effective queue schemes for reducing head-of-line blocking in fat-trees. Concurrency and Computation: Practice and Experience, 23(17):2235–2248, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elghirani:2009:ISR

[ESZ09] Ali Elghirani, Riky Subrata, and Albert Y. Zomaya. Intelligent scheduling and replication: a synergistic approach. Concurrency and Computation: Practice and Experience, 21(3): 357–376, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elmroth:2009:SBG

[ET09] Erik Elmroth and Johan Tordsson. A standards-based Grid resource brokering service supporting advance reservations, coallocation, and cross-Grid interoperability. Concurrency and Computation: Practice and Experience, 21(18):2298– 2335, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Edstrom:2015:ISR

[ET15] John Edstrom and Eli Tilevich. Improving the survivability of RESTful Web applications via declarative fault tolerance. Concurrency and Computation: Practice and Experience, 27 (12):3108–3125, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Exposito:2013:SIP

[ETR^{+}13] Roberto R. Expósito, Guillermo L. Taboada, Sabela Ramos, Juan Touriño, and Ramón Doallo. Special issue papers: General-purpose computation on GPUs for high performance cloud computing. Concurrency and Computation: Practice and Experience, 25(12):1628–1642, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Exposito:2015:LLJ

[ETR^{+}15] Roberto R. Expósito, Guillermo L. Taboada, Sabela Ramos, Juan Touriño, and Ramón Doallo. Low-latency Java communication devices on RDMA-enabled networks. Concurrency and Computation: Practice and Experience, 27(17): 4852–4879, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Emami:2018:SCS

[EZJ^{+}18] Mohammad Emami, Houman Zarrabi, Mohammad Reza Jabbarpour, Masumeh Sadat Taheri, and Jason J. Jung. A soft cooperative spectrum sensing in the presence of most destructive smart PUEA using energy detector. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4524.

Faraji:2018:DCG

[FA18] Iman Faraji and Ahmad Afsahi. Design considerations for GPU-aware collective communications in MPI. Concurrency and Computation: Practice and Experience, 30(17):e4667:1– e4667:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2007:MRT

[FAB^{+}07] Geoffrey Fox, Galip Aydin, Hasan Bulut, Harshawardhan Gadgil, Shrideep Pallickara, Marlon Pierce, and Wenjun Wu. Management of real-time streaming data Grid services. Concurrency and Computation: Practice and Experience, 19(7): 983–998, May 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fabra:2011:DPD

[FÁBE11] J. Fabra, P. Álvarez, J. A. Bañares, and J. Ezpeleta. DENEB: a platform for the development and execution of interoperable dynamic Web processes. Concurrency and Computation: Practice and Experience, 23(18):2421–2451, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fayyaz:2018:VLC

[FAM^{+}18] Mohsin Fayyaz, Khurram Aziz, Ghulam Mujtaba, Ahmad Fayyaz, and Samee U. Khan. Very low computational complexity (VLCC) architecture for optical interconnect in data center networks. Concurrency and Computation: Practice and Experience, 30(10):??, May 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4395.

Ficco:2016:HBF

[FAPC16] Massimo Ficco, Giovanni Avolio, Francesco Palmieri, and Aniello Castiglione. An HLA-based framework for simulation of large-scale critical systems. Concurrency and Computation: Practice and Experience, 28(2):400–419, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Frincu:2016:EDC

[FB16] Marc Frincu and Károly Bósa. Editorials: Distributed computing track at SYNASC 2014. Concurrency and Computation: Practice and Experience, 28(11):3023–3024, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Finger:2010:RUP

[FBC10] Marcelo Finger, Germano C. Bezerra, and Danilo R. Conde. Resource use pattern analysis for predicting resource availability in opportunistic grids. Concurrency and Computation: Practice and Experience, 22(3):295–313, March 10, 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fahringer:2001:DPA

[FBH^{+}01] T. Fahringer, P. Blaha, A. Hössinger, J. Luitz, E. Mehofer, H. Moritsch, and B. Scholz. Development and performance analysis of real-world applications for distributed and parallel architectures. Concurrency and Computation: Practice and Experience, 13(10):841–868, August 25, 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/85007182/START;   http:

//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
85007182&PLACEBO=IE.pdf.

Fu:2016:IPD

[FBS16] Desheng Fu, Matthias Becker, and Helena Szczerbicka. Improving the performance of distributed discrete event simulation by exchange of conditional look-ahead. Concurrency and Computation: Practice and Experience, 28(12):3357–3368, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fournier:2013:MTP

[FBV^{+}13] Yvan Fournier, Jerome Bonelle, Pascal Vezolle, Jerry Heyman, Bruce D’Amora, Karen Magerlein, John Magerlein, Gordon Braudaway, Charles Moulinec, and Andrew Sunderland. Multiple threads and parallel challenges for large simulations to accelerate a general Navier–Stokes CFD code on massively parallel systems. Concurrency and Computation: Practice and Experience, 25(6):843–861, April 25, 2013. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fuentes:2017:SST

[FBV^{+}17] Pablo Fuentes, Mariano Benito, Enrique Vallejo, José Luis Bosque, Ramón Beivide, Andreea Anghel, Germán Rodríguez, Mitch Gusat, Cyriel Minkenberg, and Mateo Valero. A scalable synthetic traffic model of Graph500 for computer networks analysis. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Francisco:2012:MQL

[FBYO12] Alexandre P. Francisco, Ricardo Baeza-Yates, and Arlindo L. Oliveira. Mining query log graphs towards a query folksonomy. Concurrency and Computation: Practice and Experience, 24(17):2179–2192, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fujishiro:2002:PVG

[FCT^{+}02] Issei Fujishiro, Li Chen, Yuriko Takeshima, Hiroko Nakamura, and Yasuko Suzuki. Parallel visualization of gigabyte datasets in GeoFEM. Concurrency and Computation: Practice and Experience, 14(6–7):521–530, May/ June 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/94515741/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
94515741{\&}PLACEBO=IE.pdf.

Fan:2017:GTM

[FCY17] Guisheng Fan, Liqiong Chen, and Huiqun Yu. A game theoretic method to model and analyze attack-defense strategy of resource service in cloud application. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fletcher:2001:FTH

[FD01] Martyn Fletcher and S. Misbah Deen. Fault-tolerant holonic manufacturing systems. Concurrency and Computation: Practice and Experience, 13(1):43–70, January 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/77004400/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
77004400&PLACEBO=IE.pdf.

Falch:2017:MLB

[FE17] Thomas L. Falch and Anne C. Elster. Machine learning-based auto-tuning for enhanced performance portability of OpenCL applications. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Falch:2018:ILS

[FE18] Thomas L. Falch and Anne C. Elster. ImageCL: Language and source-to-source compiler for performance portability, load balancing, and scalability prediction on heterogeneous systems. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4384.

Fechner:2012:MFM

[Fec12] Bernhard Fechner. A multilevel fault model for integrated parallel fault-tolerant systems. Concurrency and Computation: Practice and Experience, 24(7):687–698, May 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fernandez:2003:LRE

[FED03] Manel Fernández, Roger Espasa, and Saumya Debray. Load redundancy elimination on executable code. Concurrency and Computation: Practice and Experience, 15(10):979–997, August 25, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fedak:2013:ESI

[Fed13] Gilles Fedak. Editorials: Special issue: MapReduce and its applications. Concurrency and Computation: Practice and Experience, 25(1):1, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferenbaugh:2013:CGS

[Fer13] Charles R. Ferenbaugh. A comparison of GPU strategies for unstructured mesh physics. Concurrency and Computation: Practice and Experience, 25(11):1547–1558, August 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferenbaugh:2015:PUM

[Fer15] Charles R. Ferenbaugh. PENNANT: an unstructured mesh mini-app for advanced architecture research. Concurrency and Computation: Practice and Experience, 27(17):4555– 4572, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2006:SIW

[FG06] Geoffrey C. Fox and Dennis Gannon. Special issue: Workflow in Grid systems. Concurrency and Computation: Practice and Experience, 18(10):1009–1019, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feldotto:2016:SEP

[FG16] Matthias Feldotto and Kalman Graffi. Systematic evaluation of peer-to-peer systems using PeerfactSim.KOM. Concurrency and Computation: Practice and Experience, 28(5): 1655–1677, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feng:2006:APA

[FGC06] Xizhou Feng, Rong Ge, and Kirk W. Cameron. The Argus prototype: aggregate use of load modules as a high-density supercomputer. Concurrency and Computation: Practice and Experience, 18(15):1975–1987, December 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Furfaro:2018:CCA

[FGG^{+}18] Angelo Furfaro, Teresa Gallo, Alfredo Garro, Domenico Saccà, and Andrea Tundis. Cybersecurity compliance analysis as a service: Requirements specification and application scenarios. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4289.

Futrelle:2011:SMS

[FGP^{+}11] Joe Futrelle, Jeff Gaynor, Joel Plutchak, James D. Myers, Robert E. McGrath, Peter Bajcsy, Jason Kastner, Kailash Kotwani, Jong Sung Lee, Luigi Marini, Rob Kooper, Terry McLaren, and Yong Liu. Semantic middleware for e-Science knowledge spaces. Concurrency and Computation: Practice and Experience, 23(17):2107–2117, December 10, 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Figiela:2018:PEH

[FGZ^{+}18] Kamil Figiela, Adam Gajek, Adam Zima, Beata Obrok, and Maciej Malawski. Performance evaluation of heterogeneous cloud functions. Concurrency and Computation: Practice and Experience, 30(23):e4792:1–e4792:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2001:ECC

[FH01] Geoffrey C. Fox and Anthony J. G. Hey. Editorial: Concurrency and Computation: Practice and Experience. Concurrency and Computation: Practice and Experience, 13(1):1–2, January 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/77004401/START;   http:

//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
77004401&PLACEBO=IE.pdf.

Fouchal:2013:DBA

[FH13] Hacène Fouchal and Zineb Habbas. Distributed backtracking algorithm based on tree decomposition over wireless sensor networks. Concurrency and Computation: Practice and Experience, 25(5):728–742, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2015:SCS

[FHH15] Haolong Fan, Farookh Khadeer Hussain, and Omar Khadeer Hussain. Semantic client-side approach for Web personalization of SaaS-based cloud services. Concurrency and Computation: Practice and Experience, 27(8):2144–2169, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fabra:2015:IGC

[FHO^{+}15] Javier Fabra, Sergio Hernández, Estefanía Otero, Juan C. Vidal, Manuel Lama, and Pedro Álvarez. Integration of grid, cluster and cloud resources to semantically annotate a largesized repository of learning objects. Concurrency and Computation: Practice and Experience, 27(17):4603–4629, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Farash:2015:PSE

[FIO15] Mohammad Sabzinejad Farash, Sk Hafizul Islam, and Mohammad S. Obaidat. A provably secure and efficient twoparty password-based explicit authenticated key exchange protocol resistance to password guessing attacks. Concurrency and Computation: Practice and Experience, 27(17): 4897–4913, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fahringer:2005:JNP

[FJ05] Thomas Fahringer and Alexandru Jugravu. JavaSymphony: a new programming paradigm to control and synchronize locality, parallelism and load balancing for parallel and distributed computing. Concurrency and Computation: Practice and Experience, 17(7–8):1005–1025, June/July 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Furlani:2013:PMA

[FJG^{+}13] Thomas R. Furlani, Matthew D. Jones, Steven M. Gallo, Andrew E. Bruno, Charng-Da Lu, Amin Ghadersohi, Ryan J. Gentner, Abani Patra, Robert L. DeLeon, Gregor von Laszewski, Fugang Wang, and Ann Zimmerman. Performance metrics and auditing framework using application kernels for high-performance computer systems. Concurrency and Computation: Practice and Experience, 25(7):918–931, May 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fahringer:2005:ATS

[FJP^{+}05] Thomas Fahringer, Alexandru Jugravu, Sabri Pllana, Radu Prodan, Clovis Seragiotto, Jr., and Hong-Linh Truong. ASKALON: a tool set for cluster and Grid computing. Concurrency and Computation: Practice and Experience, 17(2– 4):143–169, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feng:2014:SBS

[FJZ^{+}14] Xiaowen Feng, Hai Jin, Ran Zheng, Zhiyuan Shao, and Lei Zhu. A segment-based sparse matrix–vector multiplication on CUDA. Concurrency and Computation: Practice and Experience, 26(1):271–286, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Foschini:2019:FSI

[FK19] Luca Foschini and Hyunbum Kim. Foreword to the special issue on the 2017 edition of the Workshop on Performance Evaluation of communications in DIstributed Systems and WEb-based Service Architectures (PEDISWESA 2017). Concurrency and Computation: Practice and Experience, 31(24): e5429:1–e5429:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2002:GSE

[FKP^{+}02] Geoffrey Fox, Sung-Hoon Ko, Marlon Pierce, Ozgur Balsoy, Jake Kim, Sangmi Lee, Kangseok Kim, Sangyoon Oh, Xi Rao, Mustafa Varank, Hasan Bulut, Gurhan Gunduz, Xiaohong Qiu, Shrideep Pallickara, Ahmet Uyar, and Choonhan Youn.

Grid services for earthquake science. Concurrency and Computation: Practice and Experience, 14(6–7):371–393, May/ June 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/94515743/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
94515743{\&}PLACEBO=IE.pdf.

Florez:2005:LMM

[FLB^{+}05] German Florez, Zhen Liu, Susan M. Bridges, Anthony Skjellum, and Rayford B. Vaughn. Lightweight monitoring of MPI programs in real time. Concurrency and Computation: Practice and Experience, 17(13):1547–1578, November 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fu:2019:EEM

[FLG19] Xuefeng Fu, Xiangfeng Luo, and Yike Guo. Entity emotion mining in social media environment. Concurrency and Computation: Practice and Experience, 31(20):e5336:1–e5336:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2014:RRS

[FLL^{+}14] Kai Fan, Jie Li, Hui Li, Xiaohui Liang, Xuemin (Sherman) Shen, and Yintang Yang. RSEL: revocable secure efficient lightweight RFID authentication scheme. Concurrency and Computation: Practice and Experience, 26(5):1084–1096, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferreira:2002:PCI

[FLMRC02] A. Ferreira, I. Guérin Lassous, K. Marcus, and A. Rau-Chaplin. Parallel computation on interval graphs: algorithms and experiments. Concurrency and Computation: Practice and Experience, 14(11):885–910, August 25, 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/97519210/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
97519210{\&}PLACEBO=IE.pdf.

Fei:2016:PPA

[FLYL16] Xiongwei Fei, Kenli Li, Wangdong Yang, and Keqin Li. Practical parallel AES algorithms on cloud for massive users and their performance evaluation. Concurrency and Computation: Practice and Experience, 28(16):4246–4263, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Futrelle:2008:TPS

[FM08] Joe Futrelle and James Myers. Tracking provenance semantics in heterogeneous execution systems. Concurrency and Computation: Practice and Experience, 20(5):555–564, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fang:2008:PAS

[FMM08] W. Fang, S. Miles, and L. Moreau. Performance analysis of a semantics-enabled service registry. Concurrency and Computation: Practice and Experience, 20(3):207–223, March 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Frey:2013:MEW

[FMMD13] Jeremy G. Frey, Andrew Milsted, Danius Michaelides, and David De Roure. MyExperimentalScience, extending the ‘workflow’. Concurrency and Computation: Practice and Experience, 25(4):481–496, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fourneau:2010:SBP

[FMP10] Jean-Michel Fourneau, Lynda Mokdad, and Nihal Pekergin. Stochastic bounds for performance evaluation of Web services. Concurrency and Computation: Practice and Experience, 22(10):1286–1307, July 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Frew:2008:ACR

[FMS08] James Frew, Dominic Metzger, and Peter Slaughter. Automatic capture and reconstruction of computational provenance. Concurrency and Computation: Practice and Experience, 20(5):485–496, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferro:2011:PAI

[FMS11] Mariza Ferro, Antonio Roberto Mury, and Bruno Schulze. A proposal to apply inductive logic programming to self-healing problem in grid computing: How will it work? Concurrency and Computation: Practice and Experience, 23(17): 2118–2135, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferretti:2015:MCH

[FMS15] Marco Ferretti, Mirto Musci, and Luigi Santangelo. MPI– CMS: a hybrid parallel approach to geometrical motif search in proteins. Concurrency and Computation: Practice and Experience, 27(18):5500–5516, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Forshaw:2016:HST

[FMT16] M. Forshaw, A. S. McGough, and N. Thomas. HTC-Sim: a trace-driven simulation framework for energy consumption in high-throughput computing systems. Concurrency and Computation: Practice and Experience, 28(12):3260–3290, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fouchal:2013:ESI

[FN13] Hacène Fouchal and Mirela Sechi Moretti Annoni Notare. Editorial: Special issue: Performance evaluation of communications in distributed systems and Web based service architectures. Concurrency and Computation: Practice and Experience, 25(5):605–607, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fiksman:2016:SAI

[FNBS16] Evgeny Fiksman, Andrey Nikolaev, Ilya Burylov, and Sania Salahuddin. STAC-A2 on Intel architecture: evolving from scalar code to heterogeneous applications. Concurrency and Computation: Practice and Experience, 28(3):878–891, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Funasaka:2017:ALL

[FNI17] Shunji Funasaka, Koji Nakano, and Yasuaki Ito. Adaptive loss-less data compression method optimized for GPU decompression. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fisne:2018:DIR

[FO18] Alparslan Fisne and Adnan Ozsoy. Design and implementation of real-time wideband software-defined radio applications with GPGPUs. Concurrency and Computation: Practice and Experience, 30(21):e4791:1–e4791:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fursin:2004:FAM

[FOTW04] G. Fursin, M. F. P. O’Boyle, O. Temam, and G. Watts. A fast and accurate method for determining a lower bound on execution time. Concurrency and Computation: Practice and Experience, 16(2–3):271–292, February/March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2001:SIA

[Fox01] Geoffrey Fox. Special issue: ACM 2000 Java Grande Conference. Concurrency and Computation: Practice and Experience, 13(8–9):643–644, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/84503221/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=84503221&PLACEBO=IE.
pdf.

Fox:2005:SIA

[Fox05] Geoffrey Fox. Special issue: ACM 2002 Java Grande– ISCOPE Conference. Concurrency and Computation: Practice and Experience, 17(5–6):419–422, April/May 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2010:EEM

[Fox10] Geoffrey Fox. Editorial for economic models and algorithms for Grid systems. Concurrency and Computation: Practice and Experience, 22(14):1971, September 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2012:ATP

[Fox12] Geoffrey Fox. Advanced theory and practice for high performance computing and communications. Concurrency and Computation: Practice and Experience, 24(4):341, March 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2017:ESI

[Fox17a] Geoffrey Fox. Editorial: Special issue on 12th international workshop on Java technologies for real-time and embedded systems (JTRES2014). Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2017:EJT

[Fox17b] Geoffrey Fox. Editorials: Java Technologies for Real-Time and Embedded Systems (JTRES2013). Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2002:ESS

[FP02] Geoffrey C. Fox and Shrideep Pallickara. An event service to support Grid computational environments. Concurrency and Computation: Practice and Experience, 14(13–15):1097– 1127, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2009:GCW

[FP09] Geoffrey Fox and Marlon Pierce. Grids challenged by a Web 2.0 and multicore sandwich. Concurrency and Computation: Practice and Experience, 21(3):265–280, March 10, 2009. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ficco:2015:MSR

[FPC15] Massimo Ficco, Francesco Palmieri, and Aniello Castiglione. Modeling security requirements for cloud-based system development. Concurrency and Computation: Practice and Experience, 27(8):2107–2124, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2019:RME

[FPHZ19] Linjing Fan, Xuchao Pan, Zhengxiang Huang, and Xudong Zu. Retracted: The mechanism and experimental study on the interference of high voltage lines to navigation system. Concurrency and Computation: Practice and Experience, 31 (12):e4718:1–e4718:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See retraction notice [Pan20].

Fan:2020:RSM

[FPHZ20] Linjing Fan, Xuchao Pan, Zhengxiang Huang, and Xudong Zu. Retracted: Study on measurement reliability based on Liu estimator. Concurrency and Computation: Practice and Experience, 31(12):e5560:1–e5560:??, January 25, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [HML19].

Fox:2005:TEP

[FPR05] Geoffrey Fox, Shrideep Pallickara, and Xi Rao. Towards enabling peer-to-peer Grids. Concurrency and Computation: Practice and Experience, 17(7–8):1109–1131, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Felber:2002:ACC

[FR02] Pascal Felber and Michael K. Reiter. Advanced concurrency control in Java. Concurrency and Computation: Practice and Experience, 14(4):261–285, April 10, 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/93513491/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
93513491{\&}PLACEBO=IE.pdf.

Ford:2006:GGC

[FRB^{+}06] R. W. Ford, G. D. Riley, M. K. Bane, C. W. Armstrong, and T. L. Freeman. GCF: a general coupling framework. Concurrency and Computation: Practice and Experience, 18 (2):163–181, February 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Filgueiras:2019:RJP

[FRdOR^{+}19] Tatiana Pereira Filgueiras, Leonardo M. Rodrigues, Luciana de Oliveira Rech, Luciana Moreira Sá de Souza, and Hylson Vescovi Netto. RT-JADE: a preemptive real-time scheduling middleware for mobile agents. Concurrency and Computation: Practice and Experience, 31(13):e5061:1–e5061:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fey:2012:RRT

[FRKS12] Dietmar Fey, Marc Reichenbach, Marcus Komann, and Ralf Seidler. Realizing real-time centroid detection of multiple objects with marching pixels algorithms on programmable customizing hardware. Concurrency and Computation: Practice and Experience, 24(16):1821–1839, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fugini:2012:RAW

[FRU12] Mariagrazia Fugini, Claudia Raibulet, and Luigi Ubezio. Risk assessment in work environments: modeling and simulation. Concurrency and Computation: Practice and Experience, 24 (18):2381–2403, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2007:SIP

[FS07] Geoffrey C. Fox and Xiaoping Sun. Special issue: Progress of the Knowledge Grid. Concurrency and Computation: Practice and Experience, 19(15):2011–2012, October 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Froning:2018:HUC

[FS18] Holger Fröning and Federico Silla. Heterogeneous and unconventional cluster architectures and applications. Concurrency and Computation: Practice and Experience, 30(17):e4661:1– e4661:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2019:BPA

[FSG19] Xing Fan, Oliver Sinnen, and Nasser Giacaman. Balancing parallelization and asynchronization in event-driven programs with OpenMP. Concurrency and Computation: Practice and Experience, 31(4):e4959:1–e4959:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

FeFreitas:2019:PRB

[FSM^{+}19] Mateus F.e Freitas, Daniel X. Sousa, Wellington S. Martins, Thierson C. Rosa, Rodrigo M. Silva, and Marcos A. Gonçalves. Parallel rule-based selective sampling and ondemand learning to rank. Concurrency and Computation: Practice and Experience, 31(18):e4464:1–e4464:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fahringer:2002:SAS

[FSPC^{+}02] T. Fahringer, K. Sowa-Piekło, P. Czerwiński, P. Brezany, M. Bubak, R. Koppler, and R. Wismüller. SPiDER — an advanced symbolic debugger for Fortran 90/HPF programs. Concurrency and Computation: Practice and Experience, 14 (2):103–136, February 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/91513538/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=91513538&PLACEBO=IE.pdf.

Felea:2006:DLB

[FT06] Violeta Felea and Bernard Toursel. Dynamic load-balancing mechanism for distributed Java applications. Concurrency and Computation: Practice and Experience, 18(3):305–331, March 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Flahive:2015:OSO

[FTR15] Andrew Flahive, David Taniar, and Wenny Rahayu. Ontology as a Service (OaaS): extending sub-ontologies on the cloud. Concurrency and Computation: Practice and Experience, 27(8):2028–2040, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Flahive:2015:MOU

[FTRA15] Andrew Flahive, David Taniar, Wenny Rahayu, and Bernady O. Apduhan. A methodology for ontology update in the semantic grid environment. Concurrency and Computation: Practice and Experience, 27(4):782–808, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Franz:2015:EEH

[FTT15] Wayne Franz, Parimala Thulasiraman, and Ruppa K. Thulasiram. Exploration/exploitation of a hybrid-enhanced MPSO–GA algorithm on a fused CPU–GPU architecture. Concurrency and Computation: Practice and Experience, 27 (4):973–993, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fang:2015:EVD

[FVLS15] Jianbin Fang, Ana Lucia Varbanescu, Xiangke Liao, and Henk Sips. Evaluating vector data type usage in OpenCL kernels. Concurrency and Computation: Practice and Experience, 27(17):4586–4602, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Figueroa:2015:SLR

[FVRM15] Cristhian Figueroa, Iacopo Vagliano, Oscar Rodríguez Rocha, and Maurizio Morisio. A systematic literature review of Linked Data-based recommender systems. Concurrency and Computation: Practice and Experience, 27(17):4659–4684, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2004:GMC

[FWU^{+}04] Geoffrey Fox, Wenjun Wu, Ahmet Uyar, Hasan Bulut, and Shrideep Pallickara. Global multimedia collaboration system. Concurrency and Computation: Practice and Experience, 16 (5):441–447, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fang:2016:SRR

[FXX16] He Fang, Li Xu, and Liang Xiao. Secure routing and resource allocation based on game theory in cooperative cognitive radio networks. Concurrency and Computation: Practice and Experience, 28(10):2958–2977, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fu:2015:FPS

[FYKW15] Zhisong Fu, Sergey Yakovlev, Robert M. Kirby, and Ross T. Whitaker. Fast parallel solver for the levelset equations on unstructured meshes. Concurrency and Computation: Practice and Experience, 27(7):1639–1657, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2007:SIA

[FZ07] Geoffrey C. Fox and Hai Zhuge. Special issue: Autonomous Grid computing. Concurrency and Computation: Practice and Experience, 19(7):943–944, May 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2008:SII

[FZ08] Geoffrey C. Fox and Hai Zhuge. Special issue: The 2nd International Conference on Semantics, Knowledge and Grid. Concurrency and Computation: Practice and Experience, 20 (7):751–752, May 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goscinski:2008:PPH

[GA08] Wojtek James Goscinski and David Abramson. Parallel programming on a high-performance application-runtime. Concurrency and Computation: Practice and Experience, 20(18): 2141–2177, December 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grant:2009:IEE

[GA09] Ryan E. Grant and Ahmad Afsahi. Improving energy efficiency of asymmetric chip multithreaded multiprocessors through reduced OS noise scheduling. Concurrency and Computation: Practice and Experience, 21(18):2355–2376, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guimaraes:2019:OIQ

[GAB19] Antonio Guimarães, Diego F. Aranha, and Edson Borin. Optimized implementation of QC–MDPC code-based cryptography. Concurrency and Computation: Practice and Experience, 31(18):e5089:1–e5089:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Graham:2006:CCM

[GAE^{+}06] Greg Graham, Anzar Afaq, David Evans, Gerald Guglielmo, Eric Wicklund, and Peter Love. Contextual constraint modeling in Grid application workflows. Concurrency and Computation: Practice and Experience, 18(10):1277–1292, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goldman:2017:ECA

[GAM17] Alfredo Goldman, Luciana Arantes, and Edward Moreno. Editorial: Computer architecture and high performance computing. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2019:WSU

[GAS19] Swati Gupta, Isha Agarwal, and Ravi Shankar Singh. Workflow scheduling using Jaya algorithm in cloud. Concurrency and Computation: Practice and Experience, 31(17):e5251:1– e5251:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grimstead:2009:RRA

[GAW09] I. J. Grimstead, N. J. Avis, and D. W. Walker. RAVE: the resource-aware visualization environment. Concurrency and Computation: Practice and Experience, 21(4):415–448, March 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goodyer:2007:PSI

[GB07] C. E. Goodyer and M. Berzins. Parallelization and scalability issues of a multilevel elastohydrodynamic lubrication solver. Concurrency and Computation: Practice and Experience, 19 (4):369–396, March 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gordienko:2015:ISG

[GBB^{+}15] Yuri Gordienko, Lev Bekenov, Olexandra Baskova, Olexandr Gatsenko, Elena Zasimchuk, and Sergii Stirenko. IMP Science Gateway: from the portal to the hub of virtual experimental labs in e-science and multiscale courses in e-learning. Concurrency and Computation: Practice and Experience, 27 (16):4451–4464, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2016:SMR

[GBD16] Yunfeng Gu, Azzedine Boukerche, and Robson E. De Grande. Supporting multidimensional range queries in Hierarchically Distributed Tree. Concurrency and Computation: Practice and Experience, 28(6):1848–1869, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2009:WPS

[GBFP09] Jia Guo, Ganesh Bikshandi, Basilio B. Fraguela, and David Padua. Writing productive stencil codes with overlapped tiling. Concurrency and Computation: Practice and Experience, 21(1):25–39, January 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grunzke:2014:SBM

[GBG^{+}14] Richard Grunzke, Sebastian Breuers, Sandra Gesing, Sonja Herres-Pawlis, Martin Kruse, Dirk Blunk, Luis de la Garza, Lars Packschies, Patrick Schäfer, Charlotta Schärfe, Tobias Schlemmer, Thomas Steinke, Bernd Schuller, Ralph Müller-Pfefferkorn, René Jäkel, Wolfgang E. Nagel, Malcolm Atkinson, and Jens Krüger. Standards-based metadata management for molecular simulations. Concurrency and Computation: Practice and Experience, 26(10):1744–1759, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2019:PWC

[GBJ19] Chen Gu, Matthew Bradbury, and Arshad Jhumka. Phantom walkabouts: a customisable source location privacy aware routing protocol for wireless sensor networks. Concurrency and Computation: Practice and Experience, 31(20):e5304:1– e5304:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomes:2015:EBN

[GBMM15] Antonio Tadeu A. Gomes, Bruno F. Bastos, Vivian Medeiros, and Vinicius M. Moreira. Experiences of the Brazilian national high-performance computing network on the rapid prototyping of science gateways. Concurrency and Computation: Practice and Experience, 27(2):271–289, February 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gidron:2001:DCA

[GBSHA01] Yoad Gidron, Israel Ben-Shaul, Ophir Holder, and Yariv Aridor. Dynamic configuration of access control for mobile components in FarGo. Concurrency and Computation: Practice and Experience, 13(1):5–22, January 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/77004397/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
77004397&PLACEBO=IE.pdf.

Gou:2017:DAB

[GBXL17] Gaopeng Gou, Quan Bai, Gang Xiong, and Zhenzhen Li. Discovering abnormal behaviors via HTTP header fields measurement. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2018:BST

[GC18] Juncai Guo and Xue Chen. Bias-sentiment-topic model for microblog sentiment analysis. Concurrency and Computation: Practice and Experience, 30(13):??, July 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4417.

Ghasemi:2019:AAS

[GC19] Ehsan Ghasemi and Paul Chow. Accelerating Apache Spark with FPGAs. Concurrency and Computation: Practice and Experience, 31(2):e4222:1–e4222:??, January 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grosu:2008:CLB

[GCL08] D. Grosu, A. T. Chronopoulos, and M. Y. Leung. Cooperative load balancing in distributed systems. Concurrency and Computation: Practice and Experience, 20(16):1953–1976, November 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gorton:2009:DIH

[GCN09] Ian Gorton, Daniel Chavarria, and Jarek Nieplocha. Design and implementation of a high-performance CCA event service. Concurrency and Computation: Practice and Experience, 21(9):1159–1179, June 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gai:2014:FCN

[GCO^{+}14] Jiading Gai, Dong Ju Choi, David O’Neal, Mao Ye, and Robert S. Sinkovits. Fast construction of nanosecond level snapshots of financial markets. Concurrency and Computation: Practice and Experience, 26(13):2149–2156, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guerrero:2014:TEE

[GCPS^{+}14] Ginés D. Guerrero, Juan M. Cebrián, Horacio Pérez-Sánchez, José M. García, Manuel Ujaldón, and José M. Cecilia. Toward energy efficiency in heterogeneous processors: findings on virtual screening methods. Concurrency and Computation: Practice and Experience, 26(10):1832–1846, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gill:2019:RSC

[GCSB19] Sukhpal Singh Gill, Inderveer Chana, Maninder Singh, and Rajkumar Buyya. RADAR: Self-configuring and self-healing in resource management for enhancing quality of cloud services. Concurrency and Computation: Practice and Experience, 31(1):e4834:1–e4834:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ge:2015:HSH

[GCWE15] Jingguo Ge, Zhi Chen, Yulei Wu, and Yuepeng E. H-SOFT: a heuristic storage space optimisation algorithm for flow table of OpenFlow. Concurrency and Computation: Practice and Experience, 27(13):3497–3509, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2017:RSI

[GCZ^{+}17] Chonglin Gu, Shi Chen, Jiangtao Zhang, Hejiao Huang, and Xiaohua Jia. Reservation schemes for IaaS cloud broker: a time-multiplexing way for different rental time. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grosu:2006:ARG

[GD06] D. Grosu and A. Das. Auctioning resources in Grids: model and protocols. Concurrency and Computation: Practice and Experience, 18(15):1909–1927, December 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gores:2007:DDS

[GD07] Jürgen Göres and Stefan Dessloch. Discovering data sources in a dynamic Grid environment. Concurrency and Computation: Practice and Experience, 19(16):2109–2124, November 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gores:2008:MMF

[GD08] Jürgen Göres and Stefan Dessloch. A metadata management framework for dynamic information integration. Concurrency and Computation: Practice and Experience, 20(17): 2009–2023, December 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gourlay:2004:SBS

[GDD^{+}04] I. Gourlay, P. M. Dew, K. Djemame, J. F. Snowdon, and G. Russell. Supporting Bulk Synchronous Parallelism with a high-bandwidth optical interconnect. Concurrency and Computation: Practice and Experience, 16(13):1247–1270, November 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ge:2016:TSS

[GDJ16] Dajie Ge, Zhijun Ding, and Hongfei Ji. A task scheduling strategy based on weighted round robin for distributed crawler. Concurrency and Computation: Practice and Experience, 28(11):3202–3212, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomes:2018:CCE

[GdMK^{+}18] Jeremias Gomes, Alba C. M. A. de Melo, Jun Kong, Tahsin Kurc, Joel H. Saltz, and George Teodoro. Cooperative and out-of-core execution of the irregular wavefront propagation pattern on hybrid machines with Intel(R) Xeon Phi^{T M}. Concurrency and Computation: Practice and Experience, 30(14):??, July 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4425.

Gonzalez-Dominguez:2012:ULP

[GDMT^{+}12] Jorge González-Domínguez, María J. Martín, Guillermo L. Taboada, Juan Touriño, Ramón Doallo, Damián A. Mallón, and Brian Wibecan. UPCBLAS: a library for parallel matrix computations in Unified Parallel C. Concurrency and Computation: Practice and Experience, 24(14):1645–1667, September 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gregg:2006:OCC

[GE06] David Gregg and M. Anton Ertl. Optimizing code-copying JIT compilers for virtual stack machines. Concurrency and Computation: Practice and Experience, 18(11):1465–1484, September 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gravvanis:2008:JMB

[GE08] George A. Gravvanis and Victor N. Epitropou. Java multithreading-based parallel approximate arrow-type inverses. Concurrency and Computation: Practice and Experience, 20(10):1151–1172, July 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grebhahn:2017:VSC

[GEBA17] Alexander Grebhahn, Christian Engwer, Matthias Bolten, and Sven Apel. Variability of stencil computations for porous media. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Giannino:2018:PDS

[GED^{+}18] Francesco Giannino, Serena Esposito, Marcello Diano, Salvatore Cuomo, and Gerardo Toraldo. A predictive Decision Support System (DSS) for a microalgae production plant based on Internet of Things paradigm. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4476.

Gardfjall:2008:SGW

[GEJ^{+}08] Peter Gardfjäll, Erik Elmroth, Lennart Johnsson, Olle Mulmo, and Thomas Sandholm. Scalable Grid-wide capacity allocation with the SweGrid Accounting System (SGAS). Concurrency and Computation: Practice and Experience, 20 (18):2089–2122, December 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gerndt:2005:APA

[Ger05] M. Gerndt. Automatic performance analysis tools for the Grid. Concurrency and Computation: Practice and Experience, 17(2–4):99–115, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gerndt:2007:SDP

[GF07] Michael Gerndt and Karl Fürlinger. Specification and detection of performance problems with ASL. Concurrency and Computation: Practice and Experience, 19(11):1451–1464, August 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gabriel:2010:TPP

[GFBR10] Edgar Gabriel, Saber Feki, Katharina Benkert, and Michael M. Resch. Towards performance portability through runtime adaptation for high-performance computing applications. Concurrency and Computation: Practice and Experience, 22 (16):2230–2246, November 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goderis:2009:BWD

[GFG^{+}09] Antoon Goderis, Paul Fisher, Andrew Gibson, Franck Tanoh, Katy Wolstencroft, David De Roure, and Carole Goble. Benchmarking workflow discovery: a case study from bioinformatics. Concurrency and Computation: Practice and Experience, 21(16):2052–2069, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Griebl:2004:STM

[GFL04] Martin Griebl, Peter Faber, and Christian Lengauer. Spacetime mapping and tiling: a helpful combination. Concurrency and Computation: Practice and Experience, 16(2–3):221–246, February/March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gerndt:2007:SIE

[GG07] Michael Gerndt and John Gurd. Special issue: European– American Working Group on Automatic Performance Analysis (APART). Concurrency and Computation: Practice and Experience, 19(11):1447–1449, August 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Giannoutakis:2009:DIP

[GG09] Konstantinos M. Giannoutakis and George A. Gravvanis. Design and implementation of parallel approximate inverse classes using OpenMP. Concurrency and Computation: Practice and Experience, 21(2):115–131, February 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2019:NKP

[GG19] Litao Guo and Xiaofeng Guo. A new kind of parameter for fault tolerance of graphs. Concurrency and Computation: Practice and Experience, 31(10):e4787:1–e4787:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2019:CSJ

[GGC19] Shashank Gupta, B. B. Gupta, and Pooja Chaudhary. A client-server JavaScript code rewriting-based framework to detect the XSS worms from online social network. Concurrency and Computation: Practice and Experience, 31(21): e4646:1–e4646:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Garcia-Guirado:2014:MRD

[GGFPGB14] Antonio García-Guirado, Ricardo Fernández-Pascual, José M. García, and Sandro Bartolini. Managing resources dynamically in hybrid photonic-electronic networks-on-chip. Concurrency and Computation: Practice and Experience, 26(15): 2530–2550, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gallege:2016:UTS

[GGHR16] Lahiru S. Gallege, Dimuthu U. Gamage, James H. Hill, and Rajeev R. Raje. Understanding the trust of software-intensive distributed systems. Concurrency and Computation: Practice and Experience, 28(1):114–143, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomez:2011:HRP

[GGLD11] Crispín Gómez, María E. Gómez, Pedro López, and José Duato. How to reduce packet dropping in a bufferless NoC. Concurrency and Computation: Practice and Experience, 23(1): 86–99, January 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Glasscoe:2010:AEF

[GGR^{+}10] Margaret T. Glasscoe, Robert A. Granat, John B. Rundle, Paul B. Rundle, Andrea Donnellan, and Louise H. Kellogg. Analysis of emergent fault element behavior in Virtual California. Concurrency and Computation: Practice and Experience, 22(12):1665–1683, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Georgakoudis:2016:MMF

[GGS^{+}16] Giorgis Georgakoudis, Charles J. Gillan, Ahmed Sayed, Ivor Spence, Richard Faloon, and Dimitrios S. Nikolopoulos. Methods and metrics for fair server assessment under realtime financial workloads. Concurrency and Computation: Practice and Experience, 28(3):916–928, March 10, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goli:2014:BCU

[GGV14] Mehdi Goli and Horacio González-Vélez. N-body computations using skeletal frameworks on multicore CPU/graphics processing unit architectures: an empirical performance evaluation. Concurrency and Computation: Practice and Experience, 26(4):972–986, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Golbeck:2008:SWA

[GH08] Jennifer Golbeck and James Hendler. A Semantic Web approach to the provenance challenge. Concurrency and Computation: Practice and Experience, 20(5):431–439, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guan:2006:GFG

[GHB^{+}06] Zhijie Guan, Francisco Hernandez, Purushotham Bangalore, Jeff Gray, Anthony Skjellum, Vijay Velusamy, and Yin Liu. Grid-Flow: a Grid-enabled scientific workflow system with a Petri-net-based interface. Concurrency and Computation: Practice and Experience, 18(10):1115–1140, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2019:SWM

[GHLS19] Liqiang Gao, Xiong Hu, Wei Liu, and Dejian Sun. Study on wear mechanism of gear pump end plate of combined harvester based on CFD fluid analysis. Concurrency and Computation: Practice and Experience, 31(10):e4711:1–e4711:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2013:AMP

[GHMX13] Jianliang Gao, Jia Hu, Geyong Min, and Li Xu. Analysis of the MAC protocol in low rate wireless personal area networks with bursty ON–OFF traffic. Concurrency and Computation: Practice and Experience, 25(1):85–93, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gurd:2005:SIG

[GHPR05] John Gurd, Tony Hey, Juri Papay, and Graham Riley. Special issue: Grid performance. Concurrency and Computation: Practice and Experience, 17(2–4):95–98, February/ April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gribaudo:2017:IBM

[GIL17] Marco Gribaudo, Mauro Iacono, and Alexander H. Levis. An IoT-based monitoring approach for cultural heritage sites: The Matera case. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomez-Iglesias:2010:GBM

[GIVRC^{+}10] Antonio Gómez-Iglesias, Miguel A. Vega-Rodríguez, Francisco Castejón, Enrique Morales-Ramos, Miguel Cárdenas-Montes, and José M. Reynolds. Grid-based metaheuristics to improve a nuclear fusion device. Concurrency and Computation: Practice and Experience, 22(11):1476–1493, August 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grosu:2017:EIS

[GJ17] Daniel Grosu and Hai Jin. Editorials: The 14th International Symposium on Parallel and Distributed Computing. Concurrency and Computation: Practice and Experience, 29(2): ??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goldchleger:2004:IOO

[GKG^{+}04] Andrei Goldchleger, Fabio Kon, Alfredo Goldman, Marcelo Finger, and Germano Capistrano Bezerra. InteGrade: objectoriented Grid middleware leveraging the idle computing power of desktop machines. Concurrency and Computation: Practice and Experience, 16(5):449–459, April 25, 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Granat:2009:PER

[GKK09] R. Granat, B. Kågström, and D. Kressner. Parallel eigenvalue reordering in real Schur forms. Concurrency and Computation: Practice and Experience, 21(9):1225–1250, June 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Geldhill:2008:MSM

[GKM^{+}08] Robert Geldhill, Sarah Kent, Andrew Milsted, Richard Chapman, Jonathan W. Essex, and Jeremy G. Frey. e-Malaria: the schools Malaria project. Concurrency and Computation: Practice and Experience, 20(3):225–238, March 10, 2008. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gurusamy:2019:DMU

[GKM19] UmaMaheswari Gurusamy, Hariharan K, and Manikandan Msk. Detection and mitigation of UDP flooding attack in a multicontroller software defined network using secure flow management model. Concurrency and Computation: Practice and Experience, 31(20):e5326:1–e5326:??, October 25, 2019.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gutierrez:2009:ARD

[GKP^{+}09] Miguel Esteban Gutiérrez, Isao Kojima, Said Mirza Pahlevi, Óscar Corcho, and Asunción Gómez-Pérez. Accessing RDF(S) data resources in service-based Grid infrastructures. Concurrency and Computation: Practice and Experience, 21 (8):1029–1051, June 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gec:2019:SAM

[GKP^{+}19] Sandi Gec, Dragi Kimovski, Uros Pascinski, Radu Prodan, and Vlado Stankovski. Semantic approach for multi-objective optimisation of the ENTICE distributed Virtual Machine and container images repository. Concurrency and Computation: Practice and Experience, 31(3):e4264:1–e4264:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gidenstam:2013:SGC

[GKPT13] Anders Gidenstam, Boris Koldehofe, Marina Papatriantafilou, and Philippas Tsigas. Scalable group communication supporting configurable levels of consistency. Concurrency and Computation: Practice and Experience, 25(5): 649–671, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2007:ESS

[GKS^{+}07] Xiaoyang Gao, Sriram Krishnamoorthy, Swarup Kumar Sahoo, Chi-Chung Lam, Gerald Baumgartner, J. Ramanujam, and P. Sadayappan. Efficient search-space pruning for integrated fusion and tiling transformations. Concurrency and Computation: Practice and Experience, 19(18):2425–2443, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Germann:2009:TMD

[GKS09] Timothy C. Germann, Kai Kadau, and Sriram Swaminarayan. 369 Tflop/s molecular dynamics simulations on the petaflop hybrid supercomputer ‘Roadrunner’. Concurrency and Computation: Practice and Experience, 21(17):

2143–2159, December 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gmeiner:2014:PMH

[GKSR14] Björn Gmeiner, Harald Köstler, Markus Stürmer, and Ulrich Rüde. Parallel multigrid on hierarchical hybrid grids: a performance study on current high performance computing clusters. Concurrency and Computation: Practice and Experience, 26(1):217–240, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Garcia:2019:HLP

[GL19] J. Daniel Garcia and Diego R. Llanos. High-level parallel programming in a heterogeneous world. Concurrency and Computation: Practice and Experience, 31(5):e5052:1–e5052:??, March 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gassend:2004:IAI

[GLC^{+}04] Blaise Gassend, Daihyun Lim, Dwaine Clarke, Marten van Dijk, and Srinivas Devadas. Identification and authentication of integrated circuits. Concurrency and Computation: Practice and Experience, 16(11):1077–1098, September 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Govindaraju:2007:DII

[GLC07] Madhusudhan Govindaraju, Michael J. Lewis, and Kenneth Chiu. Design and implementation issues for distributed CCA framework interoperability. Concurrency and Computation: Practice and Experience, 19(5):651–666, April 10, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2017:TMC

[GLD17] Wenming Guo, Lihong Liang, and Tianlang Deng. Topic mining for call centers based on A-LDA and distributed computing. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ge:2016:DSR

[GLL16] Yongyong Ge, Yannan Li, and Zhusong Liu. Delegation of signing rights for emerging 5G networks. Concurrency and Computation: Practice and Experience, 28(4):1193–1203, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gmys:2016:WSP

[GLM^{+}16] Jan Gmys, Rudi Leroy, Mohand Mezmaz, Nouredine Melab, and Daniel Tuyttens. Work stealing with private integer– vector–matrix data structure for multi-core branch-and-bound algorithms. Concurrency and Computation: Practice and Experience, 28(18):4463–4484, December 25, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghezzi:2015:PDD

[GLMT15] Carlo Ghezzi, Valerio Panzica La Manna, Alfredo Motta, and Giordano Tamburrelli. Performance-driven dynamic service selection. Concurrency and Computation: Practice and Experience, 27(3):633–650, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2019:ACP

[GLS^{+}19] Rong Gu, Chongjie Li, Peng Shu, Chunfeng Yuan, and Yihua Huang. Adaptive cache policy scheduling for big data applications on distributed tiered storage system. Concurrency and Computation: Practice and Experience, 31(15):e5138:1– e5138:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gong:2019:ADL

[GLZ19] Yunlu Gong, Nannan Lu, and Jiajian Zhang. Application of deep learning fusion algorithm in natural language processing in emotional semantic analysis. Concurrency and Computation: Practice and Experience, 31(10):e4779:1–e4779:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goswami:2004:PAS

[GM04] D. Goswami and R. Mall. A parallel algorithm for static slicing of concurrent programs. Concurrency and Computation:

Practice and Experience, 16(8):751–769, July 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gerndt:2010:SIS

[GM10] Michael Gerndt and Barton Miller. Special issue: Scalable tools for high-end computing. Concurrency and Computation: Practice and Experience, 22(6):683–684, April 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gil:2017:FLT

[GM17] Joseph (Yossi) Gil and Cristina Monni. Fitting long-tailed distribution to empirical data. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guerrini:2001:SIO

[GMF01] Giovanna Guerrini, Isabella Merlo, and Elena Ferrari. Special issue: Object-oriented databases. Concurrency and Computation: Practice and Experience, 13(11):929–931, September 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/85012146/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
85012146&PLACEBO=IE.pdf.

Gmys:2017:IBP

[GMMT17] J. Gmys, M. Mezmaz, N. Melab, and D. Tuyttens. IVM-based parallel branch-and-bound using hierarchical work stealing on multi-GPU systems. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Giunta:2015:PQS

[GMPT15] Rosario Giunta, Fabrizio Messina, Giuseppe Pappalardo, and Emiliano Tramontana. Providing QoS strategies and cloudintegration to web servers by means of aspects. Concurrency and Computation: Practice and Experience, 27(6):1498–1512, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grimshaw:2009:WNL

[GMS09] Andrew Grimshaw, Mark Morgan, and Karolina Sarnowska. WS-Naming: location migration, replication, and failure transparency support for Web Services. Concurrency and Computation: Practice and Experience, 21(8):1013–1028, June 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gerndt:2007:TSP

[GMT07] Michael Gerndt, Bernd Mohr, and Jesper Larsson Träff. A test suite for parallel performance analysis tools. Concurrency and Computation: Practice and Experience, 19(11): 1465–1480, August 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomez-Martin:2015:PEA

[GMVRGS15] César Gómez-Martín, Miguel A. Vega-Rodríguez, and José-Luis González-Sánchez. Performance and energy aware scheduling simulator for HPC: evaluating different resource selection methods. Concurrency and Computation: Practice and Experience, 27(17):5436–5459, December 10, 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gerndt:2010:APA

[GO10] M. Gerndt and M. Ott. Automatic performance analysis with periscope. Concurrency and Computation: Practice and Experience, 22(6):736–748, April 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guler:2018:ECM

[GÖ18] Berkin Güler and Öznur Özkasap. Efficient checkpointing mechanisms for primary-backup replication on the cloud. Concurrency and Computation: Practice and Experience, 30(21):e4707:1–e4707:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Godoy:2012:OCS

[God12] Daniela Godoy. One-class support vector machines for personalized tag-based resource classification in social bookmarking systems. Concurrency and Computation: Practice and Experience, 24(17):2193–2206, December 10, 2012. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goglin:2011:NAC

[Gog11] Brice Goglin. NIC-assisted cache-efficient receive stack for message passing over Ethernet. Concurrency and Computation: Practice and Experience, 23(2):199–210, February 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gopalakrishna:2017:RCP

[GOLL17] Aravind Kota Gopalakrishna, Tanir Ozcelebi, Johan J. Lukkien, and Antonio Liotta. Relevance in cyber-physical systems with humans in the loop. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guala:2007:ETM

[GP07] L. Gualà and G. Proietti. Efficient truthful mechanisms for the single-source shortest paths tree problem. Concurrency and Computation: Practice and Experience, 19(17): 2285–2297, December 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gonzalez:2018:MOC

[GPP^{+}18] Patricia González, David R. Penas, Xoan C. Pardo, Julio R. Banga, and Ramón Doallo. Multimethod optimization in the cloud: a case-study in systems biology modelling. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4488.

Goossens:2017:CMC

[GPPR17] Bernard Goossens, David Parello, Katarzyna Porada, and Djallal Rahmoune. Computing on many cores. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gel:2007:CFN

[GPS^{+}07] A. Gel, S. Pannala, M. Syamlal, T. J. O’Brien, and E. S. Gel. Comparison of frameworks for a next-generation multiphase flow solver, MFIX: a group decision-making exercise. Concurrency and Computation: Practice and Experience, 19(5): 609–624, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gracia:2009:PPH

[GPV09] Luis Gracia and Carlos Perez-Vidal. Parallelized and pipelined hardware implementation of computationally expensive prediction filters. Concurrency and Computation: Practice and Experience, 21(18):2478–2490, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gonzalez-Pardo:2012:CID

[GPVCdBRO12] Antonio González-Pardo, Pablo Varona, David Camacho, and Francisco de Borja Rodriguez Ortiz. Communication by identity discrimination in bio-inspired multi-agent systems. Concurrency and Computation: Practice and Experience, 24(6): 589–603, May 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gregg:2003:PID

[GPW03] David Gregg, James Power, and John Waldron. Platform independent dynamic Java virtual machine analysis: the Java Grande Forum benchmark suite. Concurrency and Computation: Practice and Experience, 15(3–5):459–484, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gregg:2005:MLC

[GPW05] David Gregg, James Power, and John Waldron. A methodlevel comparison of the Java Grande and SPEC JVM98 benchmark suites. Concurrency and Computation: Practice and Experience, 17(7–8):757–773, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gutierrez:2004:DPB

[GPZ04] Eladio Gutiérrez, Oscar Plata, and Emilio L. Zapata. Data partitioning-based parallel irregular reductions. Concurrency and Computation: Practice and Experience, 16(2–3):155–172, February/March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goldman:2004:MPJ

[GQ04] Alfredo Goldman and Carlos Queiroz. A model for parallel job scheduling on dynamical computer Grids. Concurrency and Computation: Practice and Experience, 16(5):461–468, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gai:2017:SCI

[GQH17] Keke Gai, Meikang Qiu, and Houcine Hassan. Secure cyber incident analytics framework using Monte Carlo simulations for financial cybersecurity insurance in cloud computing. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2018:DMS

[GQJL18] Yang Gao, Heng Qi, Yingwei Jin, and Keqiu Li. A decisionmaking solution for cloud storage system. Concurrency and Computation: Practice and Experience, 30(20):e4717:1– e4717:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2016:AAA

[GQR16] Yu Gu, Lianghu Quan, and Fuji Ren. AAH: accurate activity recognition of human beings using WiFi signals. Concurrency and Computation: Practice and Experience, 28(14): 3910–3926, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Greenfield:2013:ABQ

[GR13] Paul Greenfield and Uwe Roehm. Answering biological questions by querying k-mer databases. Concurrency and Computation: Practice and Experience, 25(4):497–509, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gusev:2014:SSR

[GR14] Marjan Gusev and Sasko Ristov. A superlinear speedup region for matrix multiplication. Concurrency and Computation: Practice and Experience, 26(11):1847–1868, August 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Garcia:2012:VIC

[GRGP12] Víctor Manuel Álvarez García, M. Puerto Paule Ruiz, Moisés Riestra González, and Juan Ramón Pérez Pérez. Voice interactive classroom: best practices and design strategies. Concurrency and Computation: Practice and Experience, 24 (16):1963–1973, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghomi:2019:SLB

[GRQ19a] Einollah Jafarnejad Ghomi, Amir Masood Rahmani, and Nooruldeen Nasih Qader. Service load balancing, scheduling, and logistics optimization in cloud manufacturing by using genetic algorithm. Concurrency and Computation: Practice and Experience, 31(20):e5329:1–e5329:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghomi:2019:AQT

[GRQ19b] Einollah Jafarnejad Ghomi, Amir Masoud Rahmani, and Nooruldeen Nasih Qader. Applying queue theory for modeling of cloud computing: a systematic review. Concurrency and Computation: Practice and Experience, 31(17):e5186:1– e5186:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomez:2006:SCM

[GRS06] Juan Carlos Gomez, Vernon Rego, and V. S. Sunderam. Scheduling communication in multithreaded programs: experimental results. Concurrency and Computation: Practice and Experience, 18(1):1–28, January 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grebhahn:2017:PIM

[GRS^{+}17] Alexander Grebhahn, Carmen Rodrigo, Norbert Siegmund, Francisco J. Gaspar, and Sven Apel. Performance-influence models of multigrid methods: a case study on triangular grids. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gemmill:2009:CDA

[GRSB09] Jill Gemmill, John-Paul Robinson, Tom Scavo, and Purushotham Bangalore. Cross-domain authorization for federated virtual organizations using the myVocs collaboration environment. Concurrency and Computation: Practice and Experience, 21(4):509–532, March 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Greco:2018:ISW

[GRTX18] Luca Greco, Pierluigi Ritrovato, Thanassis Tiropanis, and Fatos Xhafa. IoT and Semantic Web technologies for event detection in natural disasters. Concurrency and Computation: Practice and Experience, 30(21):e4789:1–e4789:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gittens:2019:AAS

[GRW^{+}19] Alex Gittens, Kai Rothauge, Shusen Wang, Michael W. Mahoney, Jey Kottalam, Lisa Gerhardt, Prabhat, Michael Ringenburg, and Kristyn Maschhoff. Alchemist: an Apache Spark ↔ MPI interface. Concurrency and Computation: Practice and Experience, 31(16):e5026:1–e5026:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomoluch:2004:PEM

[GS04a] J. Gomoluch and M. Schroeder. Performance evaluation of market-based resource allocation for Grid computing. Concurrency and Computation: Practice and Experience, 16(5): 469–475, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gundersen:2004:DSJ

[GS04b] Geir Gundersen and Trond Steihaug. Data structures in Java for matrix computations. Concurrency and Computation: Practice and Experience, 16(8):799–815, July 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gaeta:2008:EFB

[GS08] Rossano Gaeta and Matteo Sereno. On the evaluation of flooding-based search strategies in peer-to-peer networks. Concurrency and Computation: Practice and Experience, 20 (6):713–734, April 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gillan:2012:CIT

[GSB^{+}12] C. J. Gillan, T. Steinke, J. Bock, S. Borchert, I. Spence, and N. S. Scott. Comparing the implementation of twodimensional numerical quadrature on GPU, FPGA and Clear-Speed systems to study electron scattering by atoms. Concurrency and Computation: Practice and Experience, 24(1): 84–95, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gibert:2006:ISC

[GSG06] Enric Gibert, Jesús Sánchez, and Antonio González. Instruction scheduling for a clustered VLIW processor with a wordinterleaved cache. Concurrency and Computation: Practice and Experience, 18(11):1391–1411, September 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2019:SHS

[GSK19] Shashank Gupta, Monika Sharma, and Dhrub Kumar. SEC-H5: Secure and efficient integration of settings of enhanced HTML5 XSS vector defensive framework on edge network of fog nodes. Concurrency and Computation: Practice and Experience, 31(17):e5188:1–e5188:??, September 10, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomes:2019:MID

[GSR^{+}19] Demis Gomes, Guto Leoni Santos, Daniel Rosendo, Glauco Gonçalves, Andre Moreira, Judith Kelner, Djamel Sadok, and Patricia Takako Endo. Measuring the impact of data center failures on a cloud-based emergency medical call system. Concurrency and Computation: Practice and Experience, 31 (15):e5156:1–e5156:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Graham:2003:PIM

[GSV03] Ivan G. Graham, Alastair Spence, and Eero Vainikko. Parallel iterative methods for Navier–Stokes equations and application to eigenvalue computation. Concurrency and Computation: Practice and Experience, 15(11–12):1151–1168, September 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2019:KAS

[GSWJ19] Tao Gu, Lijun Song, Hua Wang, and Maozhu Jin. Key analysis of smart tourism project setting and tourists’ satisfaction degree based on data mining. Concurrency and Computation: Practice and Experience, 31(10):e4755:1–e4755:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guitart:2010:SPM

[GTA10] Jordi Guitart, Jordi Torres, and Eduard Ayguadé. A survey on performance management for Internet applications. Concurrency and Computation: Practice and Experience, 22(1): 68–106, January 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gaona:2013:DEE

[GTFA13] E. Gaona, R. Titos, J. Fernández, and M. E. Acacio. On the design of energy-efficient hardware transactional memory systems. Concurrency and Computation: Practice and Experience, 25(6):862–880, April 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gervasi:2011:SIC

[GTGT11] Osvaldo Gervasi, C. J. Kenneth Tan, Marina Gavrilova, and David Taniar. Special issue on Computational Science and Its Applications. Concurrency and Computation: Practice and Experience, 23(5):435, April 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Groep:2006:CRD

[GTL06] D. Groep, J. Templon, and C. Loomis. Crunching real data on the Grid: practice and experience with the European Data-Grid. Concurrency and Computation: Practice and Experience, 18(9):925–940, August 10, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2019:PAI

[Guo19] Xu Guo. Performance analysis of Israeli–Jalfon’s algorithm using probabilistic model checking. Concurrency and Computation: Practice and Experience, 31(9):e4973:1–e4973:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gonzalez-Velez:2010:ASP

[GVC10] Horacio González-Vélez and Murray Cole. Adaptive structured parallelism for distributed heterogeneous architectures: a methodological approach with pipelines and farms. Concurrency and Computation: Practice and Experience, 22(15): 2073–2094, October 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grelck:2012:AAO

[GvDHS12] Clemens Grelck, Tim van Deurzen, Stephan Herhut, and Sven-Bodo Scholz. Asynchronous adaptive optimisation for generic data-parallel array programming. Concurrency and Computation: Practice and Experience, 24(5):499–516, April 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gesing:2011:SIP

[GvHKK11] Sandra Gesing, Jano van Hemert, Peter Kacsuk, and Oliver Kohlbacher. Special issue: Portals for life sciences — providing intuitive access to bioinformatic tools. Concurrency and Computation: Practice and Experience, 23(3):223–234, March 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomes:2012:PEM

[GVK12] Eduardo R. Gomes, Quoc Bao Vo, and Ryszard Kowalczyk. Pure exchange markets for resource sharing in federated clouds. Concurrency and Computation: Practice and Experience, 24(9):977–991, June 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guzek:2014:HMP

[GVP^{+}14] Mateusz Guzek, Sébastien Varrette, Valentin Plugaru, Johnatan E. Pecero, and Pascal Bouvry. A holistic model of the performance and the energy efficiency of hypervisors in a high-performance computing environment. Concurrency and Computation: Practice and Experience, 26(15):2569– 2590, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2015:ACA

[GW15] Ping Guo and Liqiang Wang. Accurate cross-architecture performance modeling for sparse matrix-vector multiplication (SpMV) on GPUs. Concurrency and Computation: Practice and Experience, 27(13):3281–3294, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gustavson:2019:APM

[GW19] Fred G. Gustavson and David W. Walker. Algorithms for inplace matrix transposition. Concurrency and Computation: Practice and Experience, 31(13):e5071:1–e5071:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gunarathne:2011:CCP

[GWC^{+}11] Thilina Gunarathne, Tak-Lon Wu, Jong Youl Choi, Seung-Hee Bae, and Judy Qiu. Cloud computing paradigms for pleasingly parallel biomedical applications. Concurrency and Computation: Practice and Experience, 23(17):2338–2354, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gesing:2015:ESG

[GWD15] Sandra Gesing and Nancy Wilkins-Diehr. Editorials: Science gateway workshops 2014 special issue conference publications. Concurrency and Computation: Practice and Experience, 27 (16):4247–4251, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Given-Wilson:2019:ASF

[GWHJL19] Thomas Given-Wilson, Annelie Heuser, Nisrine Jafri, and Axel Legay. An automated and scalable formal process for detecting fault injection vulnerabilities in binaries. Concurrency and Computation: Practice and Experience, 31(23):e4794:1– e4794:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guerrero:2014:PCM

[GWVP^{+}14] Ginés D. Guerrero, Richard M. Wallace, José L. Vázquez-Poletti, José M. Cecilia, José M. García, Daniel Mozos, and Horacio Pérez-Sánchez. A performance/cost model for a CUDA drug discovery application on physical and public cloud infrastructures. Concurrency and Computation: Practice and Experience, 26(10):1787–1798, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Geimer:2010:SPT

[GWW^{+}10] Markus Geimer, Felix Wolf, Brian J. N. Wylie, Erika Ábrahám, Daniel Becker, and Bernd Mohr. The Scalasca performance toolset architecture. Concurrency and Computation: Practice and Experience, 22(6):702–719, April 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2017:NMG

[GWW17] Jiaquan Gao, Yu Wang, and Jun Wang. A novel multigraphics processing unit parallel optimization framework for the sparse matrix–vector multiplication. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Garbervetsky:2011:QDM

[GYB^{+}11] Diego Garbervetsky, Sergio Yovine, Víctor Braberman, Martín Rouaux, and Alejandro Taboada. Quantitative dynamic-memory analysis for Java. Concurrency and Computation: Practice and Experience, 23(14):1665–1678, September 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2014:TOQ

[GYM14] Hao Gao, Jun Yan, and Yi Mu. Trust-oriented QoS-aware composite service selection based on genetic algorithms. Concurrency and Computation: Practice and Experience, 26(2): 500–515, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2016:PPC

[GYP^{+}16] KunYin Guo, Ke Yu, ShanChen Pang, Dan Yang, Jun Huang, YunNi Xia, Xin Luo, and Jia Li. On the performance and power consumption analysis of elastic clouds. Concurrency and Computation: Practice and Experience, 28(17):4367– 4384, December 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2017:NAT

[GYS^{+}17] Xiaodan Gu, Ming Yang, Congcong Shi, Zhen Ling, and Junzhou Luo. A novel attack to track users based on the behavior patterns. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2016:RSN

[GZG^{+}16] Kun Gao, Yiwei Zhu, Songjie Gong, Hengsong Tan, and Guangyu Zhou. Research on social network discovery algorithm in pervasive sensing environment. Concurrency and Computation: Practice and Experience, 28(15):4093–4106, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2019:SSE

[GZHF19] Cheng Gao, Rui Zhang, Jiaoying Huang, and Chengcheng Fu. Study of single event transient induced by heavy-ion in NMOS transistor and CMOS inverter. Concurrency and Computation: Practice and Experience, 31(12):e4802:1–e4802:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grosu:2017:ESI

[GZX17] Daniel Grosu, Sheng Zheng, and Li Xu. Editorial: Special issue on selected papers from the 15th International Symposium on Parallel and Distributed Computing. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2007:WPS

[HAA^{+}07] Lican Huang, Asif Akram, Rob Allan, David W. Walker, Omer F. Rana, and Yan Huang. A workflow portal supporting multi-language interoperation and optimization. Concurrency and Computation: Practice and Experience, 19(12): 1583–1595, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huda:2017:FMF

[HAA^{+}17] Shamsul Huda, Jemal Abawajy, Mali Abdollahian, Rafiqul Islam, and John Yearwood. A fast malware feature selection approach using a hybrid of multi-linear and stepwise binary logistic regression. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hassan:2016:QTA

[HAAWA^{+}16] Mohammad Mehedi Hassan, Mohammad Abdullah-Al-Wadud, Ahmad Almogren, Sk Md. Mizanur Rahman, Abdulhameed Alelaiwi, Atif Alamri, and Md. Abdul Hamid. QoS and trust-aware coalition formation game in data-intensive cloud federations. Concurrency and Computation: Practice and Experience, 28(10):2889–2905, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hossain:2009:BIF

[HAE09] M. Shamim Hossain, A. Alamri, and Abdulmotaleb El Saddik. A biologically inspired framework for multimedia service management in a ubiquitous environment. Concurrency and Computation: Practice and Experience, 21(11):1450–1466, August 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2016:PSA

[HAJL16] Xuefeng He, Xingzheng Ai, Yuewu Jing, and Yuanyuan Liu. Partner selection of agricultural products supply chain based on data mining. Concurrency and Computation: Practice and Experience, 28(4):1246–1256, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hajiali:2019:PPA

[HAK19] Mahdi Hajiali, Maryam Amirmazlaghani, and Hossain Kordestani. Preventing phishing attacks using text and image watermarking. Concurrency and Computation: Practice and Experience, 31(13):e5083:1–e5083:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haghparast:2019:FLC

[HAN19] Mahboobeh Haghparast, Noorhidawati Abdullah, and Fariza Hanum Nasaruddin. Fog learning for cultivating critical thinking in information seeking process. Concurrency and Computation: Practice and Experience, 31(8):e5002:1–e5002:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hassanein:2017:BGH

[Has17] Maha Amin Hassanein. Block Gauss–Huard algorithm with column pivoting on a hybrid CPU–GPU architecture. Concurrency and Computation: Practice and Experience, 29(4): ??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hunter:2013:AQT

[HAvI13] Jane Hunter, Abdulmonem Alabri, and Catharine van Ingen. Assessing the quality and trustworthiness of citizen science data. Concurrency and Computation: Practice and Experience, 25(4):454–466, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hesselink:2018:FME

[HBD18] Wim H. Hesselink, Peter A. Buhr, and David Dice. Fast mutual exclusion by the Triangle algorithm. Concurrency and Computation: Practice and Experience, 30(4):??, February 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4183.

Hernandez:2006:GGA

[HBG^{+}06] Francisco Hernández, Purushotham Bangalore, Jeff Gray, Zhijie Guan, and Kevin Reilly. GAUGE: Grid Automation and Generative Environment. Concurrency and Computation: Practice and Experience, 18(10):1293–1316, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haupt:2002:MCW

[HBH02] Tomasz Haupt, Purushotham Bangalore, and Gregory Henley. Mississippi Computational Web Portal. Concurrency and Computation: Practice and Experience, 14(13–15):1275– 1287, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Heydemann:2006:UGT

[HBKM06] K. Heydemann, F. Bodin, P. M. W. Knijnenburg, and L. Morin. UFS: a global trade-off strategy for loop unrolling for VLIW architectures. Concurrency and Computation: Practice and Experience, 18(11):1413–1434, September 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hughes:2007:EPI

[HC07] Martyn D. Hughes and Ke Chen. Efficient parallelization of the iterative solution of a coupled fluid–structure interaction problem. Concurrency and Computation: Practice and Experience, 19(10):1423–1445, July 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hernandez-Castro:2016:MLE

[HCBRM16] Carlos Javier Hernández-Castro, David F. Barrero, and María D. R-Moreno. Machine learning and empathy: the Civil Rights CAPTCHA. Concurrency and Computation: Practice and Experience, 28(4):1310–1323, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2015:QIM

[HCC^{+}15] Yonghong Hu, Xuebin Chi, Debbi Chen, David K. Kahaner, and David A. Yuen. A quantitative index for measuring the development of supercomputing. Concurrency and Computation: Practice and Experience, 27(17):4685–4703, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Houstis:2002:MRP

[HCD^{+}02] E. N. Houstis, A. C. Catlin, N. Dhanjani, J. R. Rice, N. Ramakrishnan, and V. Verykios. MyPYTHIA: a recommendation portal for scientific software and services. Concurrency and Computation: Practice and Experience, 14(13–15):1481– 1505, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2018:PNU

[HCD^{+}18] Yun He, Brandon Cook, Jack Deslippe, Brian Friesen, Richard Gerber, Rebecca Hartman-Baker, Alice Koniges, Thorsten Kurth, Stephen Leak, Woo-Sun Yang, Zhengji Zhao, Eddie Baron, and Peter Hauschildt. Preparing NERSC users for Cori, a Cray XC40 system with Intel many integrated cores. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Howell:2007:CNT

[HCG07] F. W. Howell, R. C. Cannon, and N. H. Goddard. Catalyzer: a novel tool for integrating, managing and publishing heterogeneous bioscience data. Concurrency and Computation: Practice and Experience, 19(2):207–221, February 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hupfeld:2008:XAC

[HCK^{+}08] Felix Hupfeld, Toni Cortes, Björn Kolbeck, Jan Stender, Erich Focht, Matthias Hess, Jesus Malo, Jonathan Marti, and Eugenio Cesario. The XtreemFS architecture — a case for object-based file systems in Grids. Concurrency and Computation: Practice and Experience, 20(17):2049–2060, December 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hunger:2015:STB

[HCKF15] Lars Hunger, Biagio Cosenza, Stefan Kimeswenger, and Thomas Fahringer. Spectral turning bands for efficient Gaussian random fields generation on GPUs and accelerators. Concurrency and Computation: Practice and Experience, 27(16): 4122–4136, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hong:2018:CBM

[HCS18] Taekeun Hong, Chang Choi, and Juhyun Shin. CNN-based malicious user detection in social networks. Concurrency and Computation: Practice and Experience, 30(2):??, January 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hadjidoukas:2009:HPF

[HDDG09] P. E. Hadjidoukas, V. V. Dimakopoulos, M. Delakis, and C. Garcia. A high-performance face detection system using OpenMP. Concurrency and Computation: Practice and Experience, 21(15):1819–1837, October 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hidalgo:2013:ESI

[HdV13] J. Ignacio Hidalgo and Francisco Fernández de Vega. Editorials: Special issue on parallel architectures and bioinspired algorithm: guest editors message. Concurrency and Computation: Practice and Experience, 25(8):1013–1014, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2017:RNP

[HDX^{+}17] Bohan He, Ligang Dong, Tijie Xu, Shuocheng Fei, Huafei Zhang, and Weiming Wang. Research on network programming language and policy conflicts for SDN. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2019:DCG

[He19] Ming He. Distinguish computer generated and digital images: a CNN solution. Concurrency and Computation: Practice and Experience, 31(12):e4788:1–e4788:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Heye:2019:SDL

[Hey19] Alexander Heye. Scaling deep learning without increasing batchsize. Concurrency and Computation: Practice and Experience, 31(16):e5147:1–e5147:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hey:2005:SIG

[HF05] Tony Hey and Geoffrey Fox. Special issue: Grids and Web services for e-science. Concurrency and Computation: Practice and Experience, 17(2–4):317–322, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hodon:2017:EFS

[HF17] Michal Hodon and Hacene Fouchal. Editorial: Foreword to the special issue of the International Conference on Innovative Network Systems and Applications held under the Federated Conference on Computer Science and Information Systems. Concurrency and Computation: Practice and Experience, 29 (23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Halima:2010:LSM

[HFDJ10] Riadh Ben Halima, Emna Fki, Khalil Drira, and Mohamed Jmaiel. A large-scale monitoring and measurement campaign for Web services-based applications. Concurrency and Computation: Practice and Experience, 22(10):1207–1222, July 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hayardeny:2007:DDC

[HFF07] Amiram Hayardeny, Shachar Fienblit, and Eitan Farchi. Distributed desk checking. Concurrency and Computation: Practice and Experience, 19(3):295–309, March 10, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hofmann:2017:PAK

[HFR^{+}17] Johannes Hofmann, Dietmar Fey, Michael Riedmann, Jan Eitzinger, Georg Hager, and Gerhard Wellein. Performance analysis of the Kahan-enhanced scalar product on current multi-core and many-core processors. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2013:SIP

[HFTQ13] Keman Huang, Yushun Fan, Wei Tan, and Minghui Qian. Special issue papers: BSNet: a network-based framework for service-oriented business ecosystem management. Concurrency and Computation: Practice and Experience, 25(13): 1861–1878, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2011:DIE

[HG11] H. Howie Huang and Andrew S. Grimshaw. Design, implementation and evaluation of a virtual storage system. Concurrency and Computation: Practice and Experience, 23(4): 311–331, March 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hughes:2008:ERM

[HGB^{+}08] Danny Hughes, Phil Greenwood, Gordon Blair, Geoff Coulson, Paul Grace, Florian Pappenberger, Paul Smith, and Keith Beven. An experiment with reflective middleware to support grid-based flood monitoring. Concurrency and Computation: Practice and Experience, 20(11):1303–1316, August 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hart:2014:NCA

[HGT14] David L. Hart, Pam Gillman, and Erich Thanhardt. National Center for Atmospheric Research storage accounting and analysis possibilities. Concurrency and Computation: Practice and Experience, 26(13):2210–2224, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2018:EDM

[HGW18] Guixia He, Jiaquan Gao, and Jun Wang. Efficient dense matrix–vector multiplication on GPU. Concurrency and Computation: Practice and Experience, 30(19):e4705:1– e4705:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2019:IDR

[HH19] Yan Huang and Yadong Huang. Informatization design of raw material purchase and payment for feed processing enterprises under ERP system environment. Concurrency and Computation: Practice and Experience, 31(10):e4941:1–e4941:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Herbst:2014:SAW

[HHKA14] Nikolas Roman Herbst, Nikolaus Huber, Samuel Kounev, and Erich Amrehn. Self-adaptive workload classification and forecasting for proactive resource provisioning. Concurrency and Computation: Practice and Experience, 26(12):2053–2078, August 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hwang:2016:FMN

[HHPL16] Gwan-Hwan Hwang, Wei-Sian Huang, Jenn-Zjone Peng, and Yu-Wei Lin. Fulfilling mutual nonrepudiation for cloud storage. Concurrency and Computation: Practice and Experience, 28(3):583–599, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hao:2008:SPU

[HHWZ08] Tianyong Hao, Dawei Hu, Liu Wenyin, and Qingtian Zeng. Semantic patterns for user-interactive question answering. Concurrency and Computation: Practice and Experience, 20 (7):783–799, May 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hicks:2018:IBC

[Hic18] Bryce Hicks. Improving I/O bandwidth with Cray DVS client-side caching. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hackenberg:2012:PAM

[HJB12] Daniel Hackenberg, Guido Juckeland, and Holger Brunst. Performance analysis of multi-level parallelism: inter-node, intra-node and hardware accelerators. Concurrency and Computation: Practice and Experience, 24(1):62–72, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2019:MAF

[HJH19] Mei-Ling He, Qijie Jiang, and Zheng Hong. Mathematical analysis of financial decentralization and economic efficiency in both state-owned and private enterprises. Concurrency and Computation: Practice and Experience, 31(10):e4750:1– e4750:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hinkelmann:2011:CPA

[HJM^{+}11] Markus Hinkelmann, Andreas Jakoby, Nina Moebius, Tiark Rompf, and Peer Stechert. A cryptographically t-private auction system. Concurrency and Computation: Practice and Experience, 23(12):1399–1413, August 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2017:SAC

[HJTX17] Shuanglin Huang, Aixia Jing, Jianjun Tan, and Jian Xu. Subcarrier allocation and cooperative partner selection based on Nash bargaining game for physical layer security in OFDM wireless networks. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haidar:2019:IPC

[HJV^{+}19] Azzam Haidar, Heike Jagode, Phil Vaccaro, Asim YarKhan, Stanimire Tomov, and Jack Dongarra. Investigating power capping toward energy-efficient scientific applications. Concurrency and Computation: Practice and Experience, 31(6): e4485:1–e4485:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hillston:2001:PIL

[HK01] Jane Hillston and Leïla Kloul. Performance investigation of an on-line auction system. Concurrency and Computation: Practice and Experience, 13(1):23–41, January 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/77004402/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
77004402&PLACEBO=IE.pdf.

Hornung:2002:MAC

[HK02] Richard D. Hornung and Scott R. Kohn. Managing application complexity in the SAMRAI object-oriented framework. Concurrency and Computation: Practice and Experience, 14 (5):347–368, April 25, 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/93513504/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=93513504{\&}PLACEBO=IE.pdf.

Haupt:2007:CGV

[HK07] Tomasz Haupt and Anand Kalyanasundaram. Cooperative Grid Vortals. Concurrency and Computation: Practice and Experience, 19(12):1671–1681, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hassan:2015:DEE

[HKA^{+}15] Mohammad Mehedi Hassan, S. M. Kamruzzaman, Atif Alamri, Ahmad Almogren, Abdulhameed Alelaiwi, Mohammed Alnuem, Md. Manowarul Islam, and Md. Abdur Razzaque. Design of an energy-efficient and reliable data delivery mechanism for mobile ad hoc networks: a cross-layer approach. Concurrency and Computation: Practice and Experience, 27(10):2637–2655, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Handa:2019:DCE

[HKA19a] Rohit Handa, C. Rama Krishna, and Naveen Aggarwal. Document clustering for efficient and secure information retrieval from cloud. Concurrency and Computation: Practice and Experience, 31(15):e5127:1–e5127:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Handa:2019:SES

[HKA19b] Rohit Handa, C. Rama Krishna, and Naveen Aggarwal. Searchable encryption: a survey on privacy-preserving search schemes on encrypted outsourced data. Concurrency and Computation: Practice and Experience, 31(17):e5201:1– e5201:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hashemian:2014:CSW

[HKAC14] Raoufehsadat Hashemian, Diwakar Krishnamurthy, Martin Arlitt, and Niklas Carlsson. Characterizing the scalability of a Web application on a multi-core server. Concurrency and Computation: Practice and Experience, 26(12):2027–2052, August 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hamman:2007:PSS

[HKB07] C. W. Hamman, R. M. Kirby, and M. Berzins. Parallelization and scalability of a spectral element channel flow solver for incompressible Navier–Stokes equations. Concurrency and Computation: Practice and Experience, 19(10):1403–1422, July 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2008:AFB

[HKG08] H. Howie Huang, John F. Karpovich, and Andrew S. Grimshaw. Analyzing the feasibility of building a new mass storage system on distributed resources. Concurrency and Computation: Practice and Experience, 20(10):1131–1150, July 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hartmann:2008:AEL

[HKRR08] O. Hartmann, M. Kühnemann, T. Rauber, and G. Rünger. An adaptive extension library for improving collective communication operations. Concurrency and Computation: Practice and Experience, 20(10):1173–1194, July 2008. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hammond:2012:PCG

[HKS^{+}12] Jeff R. Hammond, Sriram Krishnamoorthy, Sameer Shende, Nichols A. Romero, and Allen D. Malony. Performance characterization of global address space applications: a case study with NWChem. Concurrency and Computation: Practice and Experience, 24(2):135–154, February 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haidri:2019:CED

[HKS19] R. A. Haidri, C. P. Katti, and P. C. Saxena. Cost-effective deadline-aware stochastic scheduling strategy for workflow applications on virtual machines in cloud computing. Concurrency and Computation: Practice and Experience, 31(7): e5006:1–e5006:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hager:2016:EPPa

[HKVW16] Georg Hager, Darren Kerbyson, Abhinav Vishnu, and Gerhard Wellein. Editorials: Performance and power for highly parallel systems. Concurrency and Computation: Practice and Experience, 28(2):187–188, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haustein:2006:JDJ

[HL06] Max Haustein and Klaus-Peter Löhr. JAC: declarative Java concurrency. Concurrency and Computation: Practice and Experience, 18(5):519–546, April 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haddad:2013:SIP

[HL13] Ghaith Haddad and Gary T. Leavens. Special issue papers: Specifying subtypes in Safety Critical Java programs. Concurrency and Computation: Practice and Experience, 25(16):

2290–2306, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hejun:2019:OAI

[HL19] Zhu Hejun and Zhu Liehuang. Online and automatic identification of encryption network behaviors in big data environment. Concurrency and Computation: Practice and Experience, 31(12):e4849:1–e4849:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Harms:2018:TRI

[HLA^{+}18] Kevin Harms, Ti Leggett, Ben Allen, Susan Coghlan, Mark Fahey, Carissa Holohan, Gordon McPheeters, and Paul Rich. Theta: Rapid installation and acceptance of an XC40 KNL system. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Humphrey:2010:PCC

[HLB10] Marty Humphrey, Jie Li, and Norm Beekwilder. Publication and consumption of caBIG data services using .NET. Concurrency and Computation: Practice and Experience, 22(17): 2313–2322, December 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2019:BBF

[HLC19] Hui Huang, Kuan-Ching Li, and Xiaofeng Chen. Blockchainbased fair three-party contract signing protocol for fog computing. Concurrency and Computation: Practice and Experience, 31(22):e4469:1–e4469:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2015:PSH

[HLCW15] Yonghong Hu, Qiang Li, Zongyan Cao, and Jue Wang. Parallel simulation of high-dimensional American option pricing based on CPU versus MIC. Concurrency and Computation: Practice and Experience, 27(5):1110–1121, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2017:TCB

[HLF^{+}17] Chaobo He, Hanchao Li, Xiang Fei, Atiao Yang, Yong Tang, and Jia Zhu. A topic community-based method for friend recommendation in large-scale online social networks. Concurrency and Computation: Practice and Experience, 29(6): ??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2017:TSD

[HLG17] Huawei Huang, Peng Li, and Song Guo. Traffic scheduling for deep packet inspection in software-defined networks. Concurrency and Computation: Practice and Experience, 29(16): ??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2012:TTF

[HLHC12] Chia-Cheng Hu, Chin-Feng Lai, Yueh-Min Huang, and Han-Chieh Chao. A two-tier framework for transmission-cost minimization of high-performance communication applications. Concurrency and Computation: Practice and Experience, 24 (4):383–395, March 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2015:CBC

[HLL^{+}15] Hao Hu, Tao Lin, Yan Y. Liu, Shaowen Wang, and Luis F. Rodríguez. CyberGIS–BioScope: a cyberinfrastructure-based spatial decision-making environment for biomass-to-biofuel supply chain optimization. Concurrency and Computation: Practice and Experience, 27(16):4437–4450, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hung:2016:EBP

[HLO^{+}16] Che-Lun Hung, Chun-Yuan Lin, Chia-Shin Ou, Yuan-Hong Tseng, Po-Yen Hung, Ship-Peng Li, and Chun-Ting Fu. Efficient bit-parallel subcircuit extraction using CUDA. Concurrency and Computation: Practice and Experience, 28(16): 4326–4338, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2019:EWV

[HLW^{+}19] Xiangyu Hu, Lemin Li, Tingmin Wu, Xiaoxiang Ai, Jie Gu, and Sheng Wen. Every word is valuable: Studied influence of negative words that spread during election period in social media. Concurrency and Computation: Practice and Experience, 31(21):e4525:1–e4525:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2016:III

[HLX^{+}16] Chunjing Han, Zhenyu Li, Gaogang Xie, Steve Uhlig, Yulei Wu, Liangxiong Li, Jingguo Ge, and Yunjie Liu. Insights into the issue in IPv6 adoption: a view from the Chinese IPv6 application mix. Concurrency and Computation: Practice and Experience, 28(3):616–630, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2018:NQA

[HLY18] Wenjin Hu, Zhongmin Liu, and Yuqi Ye. A new quality assessment for Thangka image inpainting. Concurrency and Computation: Practice and Experience, 30(22):e4671:1–e4671:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haidar:2012:ADS

[HLYD12] Azzam Haidar, Hatem Ltaief, Asim YarKhan, and Jack Dongarra. Analysis of dynamically scheduled tile algorithms for dense linear algebra on multicore architectures. Concurrency and Computation: Practice and Experience, 24(3):305–321, March 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2018:PDM

[HLZD18] Qiang Hu, Minghua Liu, Zhen Zhao, and Junwei Du. A path detecting method to analyze the interactive compatibility of service processes based on WS-BPEL. Concurrency and Computation: Practice and Experience, 30(19):e4699:1– e4699:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hudson:2003:SCG

[HM03] Richard L. Hudson and J. Eliot B. Moss. Sapphire: copying garbage collection without stopping the world. Concurrency and Computation: Practice and Experience, 15(3–5): 223–261, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huband:2004:PPD

[HM04] Simon Huband and Chris McDonald. Parallel program debugging by specification. Concurrency and Computation: Practice and Experience, 16(6):551–585, May 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Habib:2012:ONP

[HM12] Sami J. Habib and Paulvanna N. Marimuthu. Optimizing network performance and carbon offset through opportunistic reclustering. Concurrency and Computation: Practice and Experience, 24(16):1927–1939, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hou:2016:CCS

[HM16] Fu Hou and Xinjun Mao. Cross-clouds services autonomic management approach based on self-organizing multi-agent technology. Concurrency and Computation: Practice and Experience, 28(11):3213–3237, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hajnal:2015:RSM

[HMFK15] Ákos Hajnal, István Márton, Zoltán Farkas, and Péter Kacsuk. Remote storage management in science gateways via data bridging. Concurrency and Computation: Practice and Experience, 27(16):4398–4411, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2019:SMR

[HML19] Jiewu Huang, Lie Ma, and Rong Li. Study on measurement reliability based on Liu estimator. Concurrency and Computation: Practice and Experience, 31(12):e4709:1–e4709:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See retraction notice [FPHZ20].

Huang:2020:REP

[HML20] Jiewu Huang, Lie Ma, and Rong Li. Retracted: Evaluation of the predictive performance of the principal component two-parameter estimator. Concurrency and Computation: Practice and Experience, 31(12):e5562:1–e5562:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [LLH19].

Halper:2003:FIS

[HmLGP03] Michael Halper, Li min Liu, James Geller, and Yehoshua Perl. Frameworks for incorporating semantic relationships into object-oriented database systems. Concurrency and Computation: Practice and Experience, 15(15):1337–1362, December 25, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haglin:2009:FAP

[HMM^{+}09] D. J. Haglin, K. R. Mayes, A. M. Manning, J. Feo, J. R. Gurd, M. Elliot, and J. A. Keane. Factors affecting the performance of parallel mining of minimal unique itemsets on diverse architectures. Concurrency and Computation: Practice and Experience, 21(9):1131–1158, June 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hill:2013:SIP

[HMPPT13] David R. C. Hill, Claude Mazel, Jonathan Passerat-Palmbach, and Mamadou K. Traore. Special issue papers: Distribution of random streams for simulation practitioners. Concurrency and Computation: Practice and Experience, 25 (10):1427–1442, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hoare:2010:FGC

[Hoa10] Tony Hoare. Fine-grain concurrency. Concurrency and Computation: Practice and Experience, 22(8):912–934, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hoheisel:2006:UTL

[Hoh06] Andreas Hoheisel. User tools and languages for graph-based Grid workflows. Concurrency and Computation: Practice and Experience, 18(10):1101–1113, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hyoudou:2004:PCS

[HON04] Kazuki Hyoudou, Ryota Ozaki, and Yasuichi Nakayama. A PC cluster system employing IEEE 1394. Concurrency and Computation: Practice and Experience, 16(10):989–1003, August 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hou:2012:I

[Hou12] Zhixiang Hou. (ICICTA 2012). Concurrency and Computation: Practice and Experience, 24(11):1167–1168, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hawick:2011:HSL

[HP11] K. A. Hawick and D. P. Playne. Hypercubic storage layout and transforms in arbitrary dimensions using GPUs and CUDA. Concurrency and Computation: Practice and Experience, 23(10):1027–1050, July 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hong:2015:DHD

[HPD^{+}15] Bin Hong, Fuyang Peng, Bo Deng, Yazhou Hu, and Dongxia Wang. DAC-Hmm: detecting anomaly in cloud systems with hidden Markov models. Concurrency and Computation: Practice and Experience, 27(18):5749–5764, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Herres-Pawlis:2015:QCM

[HPHB^{+}15] Sonja Herres-Pawlis, Alexander Hoffmann, Ákos Balaskó, Peter Kacsuk, Georg Birkenheuer, André Brinkmann, Luis de la Garza, Jens Krüger, Sandra Gesing, Richard Grunzke, Gabor Terstyansky, and Noam Weingarten. Quantum chemical meta-workflows in MoSGrid. Concurrency and Computation: Practice and Experience, 27(2):344–357, February 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hong:2018:CMI

[HPK^{+}18] Jiwon Hong, Sanghyun Park, Sang-Wook Kim, Dongphil Kim, and Wonho Kim. Classifying malwares for identification of author groups. Concurrency and Computation: Practice and Experience, 30(3):??, February 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4197.

Hey:2005:RPE

[HPS05] A. J. G. Hey, J. Papay, and M. Surridge. The role of performance engineering techniques in the context of the Grid. Concurrency and Computation: Practice and Experience, 17(2– 4):297–316, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huber:2012:WCE

[HPS12] Benedikt Huber, Wolfgang Puffitsch, and Martin Schoeberl. Worst-case execution time analysis-driven object cache design. Concurrency and Computation: Practice and Experience, 24(8):753–771, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hidalgo-Paniagua:2014:CSP

[HPVRPF14] Alejandro Hidalgo-Paniagua, Miguel A. Vega-Rodríguez, Nieves Pavón, and Joaquín Ferruz. A comparative study of parallel software SURF implementations. Concurrency and Computation: Practice and Experience, 26(17):2758–2771, December 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Herrero:2011:SIG

[HQoS11] José R. Herrero, Enrique S. Quintana-ortí, and Robert Strzodka. Special issue: GPU computing. Concurrency and Computation: Practice and Experience, 23(7):667–668, May 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hippold:2006:TPT

[HR06] Judith Hippold and Gudula Rünger. Task Pool Teams: a hybrid programming environment for irregular algorithms on SMP clusters. Concurrency and Computation: Practice and Experience, 18(12):1575–1594, October 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hofmann:2018:FAA

[HR18] Michael Hofmann and Gudula Rünger. Flexible all-to-all data redistribution methods for grid-based particle codes. Concurrency and Computation: Practice and Experience, 30(13):??, July 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4421.

Hoefler:2011:SPT

[HRR^{+}11] Torsten Hoefler, Rolf Rabenseifner, Hubert Ritzdorf, Bronis R. de Supinski, Rajeev Thakur, and Jesper Larsson Träff. The scalable process topology interface of MPI 2.2. Concurrency and Computation: Practice and Experience, 23(4):293– 310, March 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Holland:2008:PPC

[HSBMR08] David A. Holland, Margo I. Seltzer, Uri Braun, and Kiran-Kumar Muniswamy-Reddy. PASSing the provenance challenge. Concurrency and Computation: Practice and Experience, 20(5):531–540, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hosny:2014:CBP

[HSHT14] Ahmad M. Hosny, Howida A. Shedeed, Ashraf S. Hussein, and Mohamed F. Tolba. Cloud-based parallel solution for estimating statistical significance of megabyte-scale DNA sequences. Concurrency and Computation: Practice and Experience, 26(1):118–133, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hatef:2018:HHI

[HSJ^{+}18] Mohammad Amin Hatef, Vahid Shaker, Mohammad Reza Jabbarpour, Jason Jung, and Houman Zarrabi. HIDCC: a hybrid intrusion detection approach in cloud computing. Concurrency and Computation: Practice and Experience, 30(3):??, February 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4171.

He:2019:CMO

[HSL19] Huaiwen He, Hong Shen, and Dieyan Liang. Cost-minimizing online algorithm for Internet green data centers on multisource energy. Concurrency and Computation: Practice and Experience, 31(21):e5044:1–e5044:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hanlon:2014:PRI

[HSM14] Matthew Hanlon, Warren Smith, and Stephen Mock. Providing resource information to users of a national computing center. Concurrency and Computation: Practice and Experience, 26(13):2292–2302, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hogan:2011:BNW

[HSRN11] James M. Hogan, Jiro Sumitomo, Paul Roe, and Felicity Newell. Biomashups: the new world of exploratory bioinformatics? Concurrency and Computation: Practice and Experience, 23(11):1169–1178, August 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hancock:2019:JEO

[HSV^{+}19] David Y. Hancock, Craig A. Stewart, Matthew Vaughn, Jeremy Fischer, John Michael Lowe, George Turner, Tyson L. Swetnam, Tyler K. Chafin, Enis Afgan, Marlon E. Pierce, and Winona Snapp-Childs. Jetstream — early operations performance, adoption, and impacts. Concurrency and Computation: Practice and Experience, 31(16):e4683:1–e4683:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Heinecke:2015:COM

[HT15] Alexander Heinecke and Carsten Trinitis. Cache-oblivious matrix algorithms in the age of multicores and many cores. Concurrency and Computation: Practice and Experience, 27 (9):2215–2234, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Higuera-Toledano:2012:ISI

[HTBR12] M. Teresa Higuera-Toledano, Uwe Brinkschulte, and Achim Rettberg. Introduction to the special issue: SORT 2010. Concurrency and Computation: Practice and Experience, 24(16): 1819–1820, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Higuera-Toledano:2016:ES

[HTBR16] M. Teresa Higuera-Toledano, Uwe Brinkschulte, and Achim Rettberg. Editorials: SORT 2014. Concurrency and Computation: Practice and Experience, 28(14):3709–3710, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hager:2016:EPPb

[HTHW16] Georg Hager, Jan Treibig, Johannes Habich, and Gerhard Wellein. Exploring performance and power properties of modern multi-core chips via simple machine models. Concurrency and Computation: Practice and Experience, 28(2):189–210, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Higuera-Toledano:2005:IMM

[HTI05] M. Teresa Higuera-Toledano and Valérie Issarny. Improving the memory management performance of RTSJ. Concurrency and Computation: Practice and Experience, 17(5–6):715–737, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hayes:2010:LSN

[HTR10] T. J. Hayes, K. F. Tiampo, and J. B. Rundle. Large-scale numerical simulations of earthquake fault systems: illuminating the role of dilatational gravity in earthquake nucleation. Concurrency and Computation: Practice and Experience, 22(12): 1644–1652, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Higuera-Toledano:2014:EIS

[HTW14] M. Teresa Higuera-Toledano and Andy Wellings. Editorials: Introduction to the Special Issue on Java Technologies for Real-Time and Embedded Systems: JTRES 2012. Concurrency and Computation: Practice and Experience, 26(14): 2405–2406, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2019:TEA

[HTWW19] DongXing Huang, Yong Tang, Yi Wang, and ShuNing Wei. Toward efficient and accurate function-call graph matching of binary codes. Concurrency and Computation: Practice and Experience, 31(21):e4871:1–e4871:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hunold:2015:OST

[Hun15] Sascha Hunold. One step toward bridging the gap between theory and practice in moldable task scheduling with precedence constraints. Concurrency and Computation: Practice and Experience, 27(4):1010–1026, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hussain:2015:ESI

[Hus15] Farookh Khadeer Hussain. Editorials: Special issue on intelligent e-services. Concurrency and Computation: Practice and Experience, 27(4):763–764, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hanlon:2015:AAP

[HVM^{+}15] Matthew R. Hanlon, Matthew Vaughn, Stephen Mock, Rion Dooley, Walter Moreira, Joe Stubbs, Chris Town, Jason Miller, Vivek Krishnakumar, Erik Ferlanti, and Eleanor Pence. Araport: an application platform for data discovery. Concurrency and Computation: Practice and Experience, 27 (16):4412–4422, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hijma:2015:SRP

[HvNJB15] P. Hijma, R. V. van Nieuwpoort, C. J. H. Jacobs, and H. E. Bal. Stepwise-refinement for performance: a methodology for many-core programming. Concurrency and Computation: Practice and Experience, 27(17):4515–4554, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Howard:2014:RRB

[HW14] Philip W. Howard and Jonathan Walpole. Relativistic red– black trees. Concurrency and Computation: Practice and Experience, 26(16):2684–2712, November 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hung:2016:GBP

[HW16] Che-Lun Hung and Yuan-Huai Wu. GPU-based parallel fuzzy c-mean clustering model via genetic algorithm. Concurrency and Computation: Practice and Experience, 28(16): 4277–4290, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ho:2018:WPB

[HWL18] Li-Yung Ho, Jan-Jan Wu, and Pangfeng Liu. Workload prediction and balance for distributed reachability processing for large-scale attribute graphs. Concurrency and Computation: Practice and Experience, 30(6):??, March 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4344.

He:2016:EGK

[HWQ^{+}16] Shuangyu He, Qianhong Wu, Bo Qin, Jianwei Liu, and Yan Li. Efficient group key management for secure big data in predictable large-scale networks. Concurrency and Computation: Practice and Experience, 28(4):1174–1192, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2003:OOD

[HWR03] Yan Huang, David W. Walker, and Omer F. Rana. Objectoriented distributed computing based on remote class reference. Concurrency and Computation: Practice and Experience, 15(1):79–91, January 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2017:EST

[HWY^{+}17] Xiangdong Huang, Jianmin Wang, Philip S. Yu, Jian Bai, and Jinrui Zhang. An experimental study on tuning the consistency of NoSQL systems. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2015:ODP

[HWZ^{+}15] Xiangdong Huang, Jianmin Wang, Yu Zhong, Shaoxu Song, and Philip S. Yu. Optimizing data partition for scaling out NoSQL cluster. Concurrency and Computation: Practice and Experience, 27(18):5793–5809, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2008:PMA

[HWZX08] Yu Huang, Hanpin Wang, Wen Zhao, and Chunxiang Xu. A practical method to analyze workflow logic models. Concurrency and Computation: Practice and Experience, 20(1):

41–60, January 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2012:EHD

[HXY^{+}12] Guoliang He, Naixue Xiong, Laurence T. Yang, Tai hoon Kim, Ching Hsien Hsu, Yuanxiang Li, and Ting Hu. Evolvable hardware design based on a novel simulated annealing in an embedded system. Concurrency and Computation: Practice and Experience, 24(4):354–370, March 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hwang:2012:SPC

[HY12] Sun Myung Hwang and Sang-Soo Yeo. Software process certification system based on K-model for high-performance software engineering. Concurrency and Computation: Practice and Experience, 24(4):396–406, March 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2019:PMD

[HYGF19] Jiaoying Huang, Daming Yang, Cheng Gao, and Chengcheng Fu. A PCA and Mahalanobis distance-based detection method for logical hardware Trojan. Concurrency and Computation: Practice and Experience, 31(12):e4724:1–e4724:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2019:TBC

[HYL^{+}19] Hao Hu, Dandong Yin, Yan Y. Liu, Jeff Terstriep, Xingchen Hong, Jeff Wendel, and Shaowen Wang. TopoLens: Building a CyberGIS community data service for enhancing the usability of high-resolution national topographic datasets. Concurrency and Computation: Practice and Experience, 31(16): e4682:1–e4682:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2015:NAC

[HYLG15] Gaofeng He, Ming Yang, Junzhou Luo, and Xiaodan Gu. A novel application classification attack against Tor. Concurrency and Computation: Practice and Experience, 27(18): 5640–5661, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hassan:2017:ESI

[HYQ17] Houcine Hassan, Laurence T. Yang, and Meikang Qiu. Editorial: Special issue on Big Data Security and Intelligent Data in Clouds (BDS-IDC). Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hua:2005:CJE

[HYX05] Han Hua, Dai Yafei, and Li Xiaoming. CSFS: a Java enabled network file storage system. Concurrency and Computation: Practice and Experience, 17(7–8):991–1003, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2014:ERI

[HZC^{+}14] Qi Han, Yinghui Zhang, Xiaofeng Chen, Hui Li, and Jiaxiang Quan. Efficient and robust identity-based handoff authentication for EAP-based wireless networks. Concurrency and Computation: Practice and Experience, 26(8):1561–1573, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2019:TMD

[HZH^{+}19] Lansheng Han, Man Zhou, Shuxia Han, Wenjing Jia, Changhua Sun, and Cai Fu. Targeting malware discrimination based on reversed association task. Concurrency and Computation: Practice and Experience, 31(23):e4922:1– e4922:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2009:NTC

[HZHP09] Jing He, Yanchun Zhang, Guangyan Huang, and Chaoyi Pang. A novel time computation model based on algorithm complexity for data intensive scientific workflow design and scheduling. Concurrency and Computation: Practice and Experience, 21(16):2070–2083, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2016:BTC

[HZL^{+}16] Jiayu Han, Wanli Zuo, Lu Liu, Yuanbo Xu, and Tao Peng. Building text classifiers using positive, unlabeled and ‘outdated’ examples. Concurrency and Computation: Practice and Experience, 28(13):3691–3706, September 10, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2019:OLS

[HZL19] Wenwu He, Fumin Zou, and Quan Liang. Online learning with sparse labels. Concurrency and Computation: Practice and Experience, 31(23):e4480:1–e4480:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2019:SSC

[HZY^{+}19] Longxia Huang, Gongxuan Zhang, Shui Yu, Anmin Fu, and John Yearwood. SeShare: Secure cloud data sharing based on blockchain and public auditing. Concurrency and Computation: Practice and Experience, 31(22):e4359:1–e4359:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ibanez:2011:CDF

[IÁBE11] María José Ibáñez, Pedro Álvarez, José Ángel Bañares, and Joaquín Ezpeleta. Control and data flow compatibility in the interaction between dynamic business processes. Concurrency and Computation: Practice and Experience, 23(1): 57–85, January 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ibanez:2011:ABP

[IÁE11] María José Ibáñez, Pedro Álvarez, and Joaquín Ezpeleta. Analyzing behavioral properties of semantic business processes with parametric data. Concurrency and Computation: Practice and Experience, 23(6):525–555, April 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ikram:2015:AIT

[IAH^{+}15] Ahsan Ikram, Ashiq Anjum, Richard Hill, Nick Antonopoulos, Lu Liu, and Stelios Sotiriadis. Approaching the Internet of Things (IoT): a modelling, analysis and abstraction framework. Concurrency and Computation: Practice and Experience, 27(8):1966–1984, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iskra:2002:PCE

[IBvA^{+}02] K. A. Iskra, R. G. Belleman, G. D. van Albada, J. Santoso, P. M. A. Sloot, H. E. Bal, H. J. W. Spoelder, and M. Bubak. The Polder Computing Environment: a system for interactive distributed simulation. Concurrency and Computation: Practice and Experience, 14(13–15):1313–1335, November/ December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ilango:2019:NAM

[IC19] Poonguzhali Ilango and Arun Chokkalingam. A novel architecture of modified turbo codes with an area efficient high speed interleaver. Concurrency and Computation: Practice and Experience, 31(14):e5067:1–e5067:??, July 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Inacio:2018:CGP

[ID18] Eduardo C. Inacio and Mario A. R. Dantas. A coarse-grained page cache aware multivariate analytical model for the storage performance of a parallel file system. Concurrency and Computation: Practice and Experience, 30(8):??, April 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4389.

Idris:2015:CAS

[IHA^{+}15] Muhammad Idris, Shujaat Hussain, Maqbool Ali, Arsen Abdulali, Muhammad Hameed Siddiqi, Byeong Ho Kang, and Sungyoung Lee. Context-aware scheduling in MapReduce: a compact review. Concurrency and Computation: Practice and Experience, 27(17):5332–5349, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ikeda:2015:POL

[IHB15] Makoto Ikeda, Taiki Honda, and Leonard Barolli. Performance of optimized link state routing protocol for video streaming application in vehicular ad-hoc networks cloud computing. Concurrency and Computation: Practice and Experience, 27(8):2054–2063, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Innocent:2019:UGG

[IKP19] A. Anasuya Threse Innocent, Sangeeta K, and G. Prakash. Universal gates on garbled circuit construction. Concurrency and Computation: Practice and Experience, 31(17):e5236:1– e5236:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ino:2012:CMG

[IOOH12] Fumihiko Ino, Akihiro Ogita, Kentaro Oita, and Kenichi Hagihara. Cooperative multitasking for GPU-accelerated grid systems. Concurrency and Computation: Practice and Experience, 24(1):96–107, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iosup:2011:PGP

[Ios11] Alexandru Iosup. POGGI: generating puzzle instances for online games on grid infrastructures. Concurrency and Computation: Practice and Experience, 23(2):158–171, February 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Inostrosa-Psijas:2018:SAA

[IPGCMW18] Alonso Inostrosa-Psijas, Veronica Gil-Costa, Mauricio Marin, and Gabriel Wainer. Semi-asynchronous approximate parallel DEVS simulation of web search engines. Concurrency and Computation: Practice and Experience, 30(7):??, April 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4149.

Igual:2013:SAB

[IQOvdG13] Francisco D. Igual, Gregorio Quintana-Ortí, and Robert van de Geijn. Scheduling algorithms-by-blocks on small clusters. Concurrency and Computation: Practice and Experience, 25(3):367–384, March 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Imbs:2011:LCC

[IR11] Damien Imbs and Michel Raynal. A liveness condition for concurrent objects: x-wait-freedom. Concurrency and Computation: Practice and Experience, 23(17):2154–2166, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ilager:2019:EET

[IRB19] Shashikant Ilager, Kotagiri Ramamohanarao, and Rajkumar Buyya. ETAS: Energy and thermal-aware dynamic virtual machine consolidation in cloud data center with proactive hotspot mitigation. Concurrency and Computation: Practice and Experience, 31(17):e5221:1–e5221:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ifill:2010:STR

[IS10] W. Ifill and S. Schneider. A step towards refining and translating B control annotations to Handel-C. Concurrency and Computation: Practice and Experience, 22(8):1023–1048, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iwashita:2002:VFD

[ISKvW02] Hidetoshi Iwashita, Naoki Sueyasu, Sachio Kamiya, and Matthijs van Waveren. VPP Fortran and the design of HPF/JA extensions. Concurrency and Computation: Practice and Experience, 14(8–9):575–588, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016131/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016131{\&}PLACEBO=IE.pdf.

Ino:2014:PSA

[ISO^{+}14] Fumihiko Ino, Kentaro Shigeoka, Tomohiro Okuyama, Masaya Motokubota, and Kenichi Hagihara. A parallel scheme for accelerating parameter sweep applications on a GPU. Concurrency and Computation: Practice and Experience, 26(2):516–531, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iizuka:2002:PSS

[ISS^{+}02] Mikio Iizuka, Daigo Sekita, Hisashi Suito, Mamoru Hyodo, Kazuro Hirahara, David Place, Peter Mora, Osamu Hazama, and Hiroshi Okuda. Parallel simulation system for earthquake generation: fault analysis modules and parallel coupling analysis. Concurrency and Computation: Practice and Experience, 14(6–7):499–519, May/ June 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/94515740/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
94515740{\&}PLACEBO=IE.pdf.

Isaila:2003:CFP

[IT03] Florin Isaila and Walter F. Tichy. Clusterfile: a flexible physical layout parallel file system. Concurrency and Computation: Practice and Experience, 15(7–8):653–679, June/July 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ichikawa:2009:OPA

[ITK09] Shuichi Ichikawa, Sho Takahashi, and Yuu Kawai. Optimizing process allocation of parallel programs for heterogeneous clusters. Concurrency and Computation: Practice and Experience, 21(4):475–507, March 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ichikawa:2017:PEI

[IUCH^{+}17] Kohei Ichikawa, Pongsakorn U-Chupala, Che Huang, Chawanat Nakasan, Te-Lung Liu, Jo-Yu Chang, Li-Chi Ku, Whey-Fone Tsai, Jason Haga, Hiroaki Yamanaka, Eiji Kawai, Yoshiyuki Kido, Susumu Date, Shinji Shimojo, Philip Papadopoulos, Mauricio Tsugawa, Matthew Collins, Kyuho Jeong, Renato Figueiredo, and Jose Fortes. PRAGMA-ENT: An international SDN testbed for cyberinfrastructure in the Pacific Rim. Concurrency and Computation: Practice and Experience, 29 (13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Islam:2009:SFN

[IZXM09] Rafiqul Islam, Wanlei Zhou, Yang Xiang, and Abdun Naser Mahmood. Spam filtering for network traffic security on a multi-core environment. Concurrency and Computation: Practice and Experience, 21(10):1307–1320, July 2009. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Javadi:2008:CAM

[JAA08] Bahman Javadi, Jemal H. Abawajy, and Mohammad K. Akbari. A comprehensive analytical model of interconnection networks in large-scale cluster systems. Concurrency and Computation: Practice and Experience, 20(1):75–97, January 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jackson:2002:PPI

[Jac02] Keith R. Jackson. pyGlobus: a Python interface to the Globus Toolkit TM. Concurrency and Computation: Practice and Experience, 14(13–15):1075–1083, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jarzebowska:2019:AGR

[JAU19] Elzbieta Jarzebowska, Krzysztof Augustynek, and Andrzej Urba’s. Automated generation of reference dynamical models for constrained robotic systems in the presence of friction and damping effects. Concurrency and Computation: Practice and Experience, 31(22):e4452:1–e4452:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Juan:2019:MIS

[JB19] Li Juan and Chen Bin. Modeling for information spreading basing on dynamics. Concurrency and Computation: Practice and Experience, 31(9):e4679:1–e4679:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jhumka:2015:FSB

[JBL15] Arshad Jhumka, Matthew Bradbury, and Matthew Leeke. Fake source-based source location privacy in wireless sensor networks. Concurrency and Computation: Practice and Experience, 27(12):2999–3020, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jander:2016:DMW

[JBL16] Kai Jander, Lars Braubach, and Winfried Lamersdorf. Distributed monitoring and workflow management for goaloriented workflows. Concurrency and Computation: Practice and Experience, 28(4):1324–1335, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaimes:2007:MNP

[JC07] Antonio López Jaimes and Carlos A. Coello Coello. MR-MOGA: a new parallel multi-objective evolutionary algorithm based on the use of multiple resolutions. Concurrency and Computation: Practice and Experience, 19(4):397–441, March 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeong:2017:ENP

[JCJ17] Jinkyu Jeong, Dong Hoon Choi, and Heeseung Jo. Enhancing network I/O performance for a virtualized Hadoop cluster. Concurrency and Computation: Practice and Experience, 29 (8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jha:2013:DCP

[JCK^{+}13] Shantenu Jha, Murray Cole, Daniel S. Katz, Manish Parashar, Omer Rana, and Jon Weissman. Distributed computing practice for large-scale science and engineering applications. Concurrency and Computation: Practice and Experience, 25(11):1559–1585, August 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaeger:2015:FGD

[JCP15] Julien Jaeger, Patrick Carribault, and Marc Pérache. Finegrain data management directory for OpenMP 4.0 and OpenACC. Concurrency and Computation: Practice and Experience, 27(6):1528–1539, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jahr:2015:FNP

[JCVU15] Ralf Jahr, Horia Calborean, Lucian Vintan, and Theo Ungerer. Finding near-perfect parameters for hardware and code optimizations with automatic multi-objective design space explorations. Concurrency and Computation: Practice and Experience, 27(9):2196–2214, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaradat:2016:TAD

[JDB16] Ward Jaradat, Alan Dearle, and Adam Barker. Towards an autonomous decentralized orchestration system. Concurrency and Computation: Practice and Experience, 28(11): 3164–3179, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jagode:2018:EDP

[JDH^{+}18] Heike Jagode, Anthony Danalis, Reazul Hoque, Mathieu Faverge, and Jack Dongarra. Evaluation of dataflow programming models for electronic structure theory. Concurrency and Computation: Practice and Experience, 30(17):e4490:1– e4490:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jimenez:2012:TDT

[JdM12] J. Jiménez and J. Ruiz de Miras. Three-dimensional thinning algorithms on graphics processing units and multicore CPUs. Concurrency and Computation: Practice and Experience, 24(14):1551–1571, September 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jarvis:2008:PPC

[JFI^{+}08] S. A. Jarvis, B. P. Foley, P. J. Isitt, D. P. Spooner, D. Rueckert, and G. R. Nudd. Performance prediction for a code with data-dependent runtimes. Concurrency and Computation: Practice and Experience, 20(3):195–206, March 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ji:2019:NSC

[JHCH19] Jian Ji, Chuanchao Huang, Ya Cao, and Sen Hu. The network structure of Chinese finance market through the method of complex network and random matrix theory. Concurrency and Computation: Practice and Experience, 31(9):e4877:1– e4877:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2013:SIP

[JJGL13] Honglei Jiang, Hai Jin, Song Guo, and Xiaofei Liao. Special issue papers: A measurement-based study on user management in private BitTorrent communities. Concurrency and Computation: Practice and Experience, 25(14):2052–2066, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Joppich:2006:MTS

[JK06] W. Joppich and M. Kürschner. MpCCI — a tool for the simulation of coupled applications. Concurrency and Computation: Practice and Experience, 18(2):183–192, February 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jung:2010:OCN

[JK10] Gi-Hoon Jung and Soon-Ju Kang. An optimized concurrent network system software supporting effective multi-session streaming in a multimedia appliance. Concurrency and Computation: Practice and Experience, 22(18):2401–2420, December 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jones:2013:CSH

[JK13] Terry Jones and Gregory A. Koenig. Clock synchronization in high-end computing environments: a strategy for minimizing clock variance at runtime. Concurrency and Computation: Practice and Experience, 25(6):881–897, April 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jocksch:2019:OAA

[JKD19] Andreas Jocksch, Matthias Kraushaar, and David Daverio. Optimized all-to-all communication on multicore architectures applied to FFTs with pencil decomposition. Concurrency and Computation: Practice and Experience, 31(16): e4964:1–e4964:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jha:2017:IDD

[JKL^{+}17] Shantenu Jha, Daniel S. Katz, Andre Luckow, Neil Chue Hong, Omer Rana, and Yogesh Simmhan. Introducing distributed dynamic data-intensive (D3) science: Understanding applications and infrastructure. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jacobsen:2019:MCS

[JKL19] Douglas M. Jacobsen, Randy Kleinman, and Harold Longley. Managing a Cray supercomputer as a git branch. Concurrency and Computation: Practice and Experience, 31(16): e5092:1–e5092:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jan:2017:ITF

[JKM^{+}17] Bilal Jan, Fiaz Gul Khan, Bartolomeo Montrucchio, Anthony Theodore Chronopoulos, Shahaboddin Shamshirband, and Abdul Nasir Khan. Introducing ToPe–FFT: An OpenCLbased FFT library targeting GPUs. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jozsa:2015:PSD

[JKV^{+}15] Csaba M. Józsa, Géza Kolumbán, Antonio M. Vidal, Francisco J. Martínez-Zaldívar, and Alberto González. Parallel sphere detector algorithm providing optimal MIMO detection on massively parallel architectures. Concurrency and Computation: Practice and Experience, 27(17):4993–5018, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jurdzinski:2003:WCS

[JKZ03] Tomasz Jurdziński, Mirosław Kutyłowski, and Jan Zatopiański. Weak communication in single-hop radio networks: adjusting algorithms to industrial standards. Concurrency and Computation: Practice and Experience, 15(11–12):1117–1131, September 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jin:2010:OMO

[JL10] Yingwei Jin and Keqiu Li. An optimal multimedia object allocation solution in multi-powermode storage systems. Concurrency and Computation: Practice and Experience, 22(13): 1852–1873, September 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Junior:2007:TAC

[JLCA07] F. H. Carvalho Junior, R. D. Lins, R. C. Corrêa, and G. A. Araújo. Towards an architecture for component-oriented parallel programming. Concurrency and Computation: Practice and Experience, 19(5):697–719, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jin:2016:GGA

[JLH^{+}16] Jing Jin, Siyan Lai, Su Hu, Jing Lin, and Xiaola Lin. GPUSGD: A GPU-accelerated stochastic gradient descent algorithm for matrix factorization. Concurrency and Computation: Practice and Experience, 28(14):3844–3865, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2014:PAM

[JLHH14] Yunliang Jiang, Yong Liu, Wenliang Huang, and Lican Huang. Performance analysis of a mobile agent prototype system based on VIRGO P2P protocols. Concurrency and Computation: Practice and Experience, 26(2):447–467, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jin:2018:CSE

[JLL18] Jiahui Jin, Yunhao Li, and Junzhou Luo. Cooperative storage by exploiting graph-based data placement algorithm for edge computing environment. Concurrency and Computation: Practice and Experience, 30(20):e4914:1–e4914:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jafari:2017:MCD

[JLQ^{+}17] Hossein Jafari, Xiangfang Li, Lijun Qian, Alexander Aved, and Timothy Kroecker. Multisensor change detection on the basis of big time-series data and Dempster–Shafer theory. Concurrency and Computation: Practice and Experience, 29 (17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jalby:2006:EMO

[JLT06] William Jalby, Christophe Lemuet, and Sid-Ahmed-Ali Touati. An efficient memory operations optimization technique for vector loops on Itanium 2 processors. Concurrency and Computation: Practice and Experience, 18(11): 1485–1508, September 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Junqueira:2007:FDD

[JM07] Flavio Junqueira and Keith Marzullo. A framework for the design of dependent-failure algorithms. Concurrency and Computation: Practice and Experience, 19(17):2255–2269, December 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeevan:2019:SIS

[JM19] A. N. Gnana Jeevan and M. A. Maluk Mohamed. SOGC: Implementing surrogate object garbage collector management for a Mobile Cloud Environment. Concurrency and Computation: Practice and Experience, 31(4):e4943:1–e4943:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jha:2009:UCP

[JMF09] Shantenu Jha, Andre Merzky, and Geoffrey Fox. Using clouds to provide grids with higher levels of abstraction and explicit support for usage modes. Concurrency and Computation: Practice and Experience, 21(8):1087–1108, June 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2016:NPD

[JML^{+}16] Wenbin Jiang, Hongyan Mei, Feng Lu, Hai Jin, Laurence T. Yang, Bin Luo, and Ye Chi. A novel parallel deblocking filtering strategy for HEVC/H.265 based on GPU. Concurrency and Computation: Practice and Experience, 28(16): 4264–4276, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jimack:2003:PAN

[JN03] P. K. Jimack and S. A. Nadeem. Parallel application of a novel domain decomposition preconditioner for the adaptive finiteelement solution of three-dimensional convection-dominated PDEs. Concurrency and Computation: Practice and Experience, 15(10):939–956, August 25, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jan:2017:PPB

[JNUH17] Mian Jan, Priyadarsi Nanda, Muhammad Usman, and Xiangjian He. PAWN: a payload-based mutual authentication scheme for wireless sensor networks. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jain:2015:FDW

[JOC^{+}15] Anubhav Jain, Shyue Ping Ong, Wei Chen, Bharat Medasani, Xiaohui Qu, Michael Kocher, Miriam Brafman, Guido Petretto, Gian-Marco Rignanese, Geoffroy Hautier, Daniel Gunter, and Kristin A. Persson. FireWorks: a dynamic workflow system designed for high-throughput applications. Concurrency and Computation: Practice and Experience, 27(17): 5037–5059, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jones:2018:EST

[JOK^{+}18] Terry Jones, George Ostrouchov, Gregory A. Koenig, Oscar H. Mondragon, and Patrick G. Bridges. An evaluation of the state of time synchronization on leadership class supercomputers. Concurrency and Computation: Practice and Experience, 30(4):??, February 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4341.

Jones:2009:NAS

[Jon09] William M. Jones. Network-aware selective job checkpoint and migration to enhance co-allocation in multi-cluster systems. Concurrency and Computation: Practice and Experience, 21(13):1672–1691, September 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Joshi:2005:DFP

[Jos05] Rushikesh K. Joshi. Distributed filter processes. Concurrency and Computation: Practice and Experience, 17(12): 1497–1506, October 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Junior:2017:PEA

[JPS17] Manoel Baptista Silva Junior, Jairo Panetta, and Stephan Stephany. Portability with efficiency of the advection of BRAMS between multi-core and many-core architectures. Concurrency and Computation: Practice and Experience, 29 (22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Johnson:2002:CBP

[JPWH02] Chris Johnson, Steve Parker, David Weinstein, and Sean Heffernan. Component-based, problem-solving environments for large-scale scientific computing. Concurrency and Computation: Practice and Experience, 14(13–15):1337–1349, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ji:2015:SMD

[JQL^{+}15] Changqing Ji, Wenyu Qu, Zhiyang Li, Yujie Xu, Yuanyuan Li, and Junfeng Wu. Scalable multi-dimensional RNN query processing. Concurrency and Computation: Practice and Experience, 27(16):4156–4171, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2008:MMI

[JQSP08] Nanyan Jiang, Andres Quiroz, Cristina Schmidt, and Manish Parashar. Meteor: a middleware infrastructure for contentbased decoupled interactions in pervasive grid environments. Concurrency and Computation: Practice and Experience, 20 (12):1455–1484, August 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2019:SOP

[JR19] Zhihua Jiang and Dongning Rao. Scalable and optimal planning based on Pregel. Concurrency and Computation: Practice and Experience, 31(7):e4966:1–e4966:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeong:2016:ECC

[JRHJ16] Hwa-Young Jeong, Omer F. Rana, Ching-Hsien Hsu, and Young-Sik Jeong. Editorials: Cloud computing for mobile environments. Concurrency and Computation: Practice and Experience, 28(10):2753–2755, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

John:2017:ABA

[JSG17] John C. John, Shamik Sural, and Arobinda Gupta. Attributebased access control management for multicloud collaboration. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jia:2015:EUI

[JSPE15] Adele Lu Jia, Boudewijn Schoon, Johan A. Pouwelse, and Dick H. J. Epema. Estimating user interaction strength in distributed online networks. Concurrency and Computation: Practice and Experience, 27(17):5311–5331, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jahan:2019:LWF

[JSR^{+}19a] Mosarrat Jahan, Suranga Seneviratne, Partha Sarathi Roy, Kouichi Sakurai, Aruna Seneviratne, and Sanjay Jha. Light weight and fine-grained access mechanism for secure access to outsourced data*. Concurrency and Computation: Practice and Experience, 31(23):e4736:1–e4736:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Joao:2019:PHA

[JSR19b] Mario João, Jr., Alexandre C. Sena, and Vinod E. F. Rebello. On the parallelization of Hirschberg’s algorithm for multicore and many-core systems. Concurrency and Computation: Practice and Experience, 31(18):e5174:1–e5174:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jones:2007:CBI

[JSS07] J. Jones, M. Sosonkina, and Y. Saad. Component-based iterative methods for sparse linear systems. Concurrency and Computation: Practice and Experience, 19(5):625–635, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jia:2019:RTH

[JTD^{+}19] Jing Jia, Geng Tu, Xin Deng, Chuchu Zhao, and Wenlong Yi. Real-time hand gestures system based on leap motion. Concurrency and Computation: Practice and Experience, 31(10): e4898:1–e4898:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jung:2016:EGR

[Jun16] Jason J. Jung. Exploiting geotagged resources for spatial clustering on social network services. Concurrency and Computation: Practice and Experience, 28(4):1356–1367, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaghoori:2015:MIG

[JvAB^{+}15] Mohammad Mahdi Jaghoori, Allard J. van Altena, Boris Bleijlevens, Sara Ramezani, Juan Luis Font, and Silvia D. Olabarriaga. A multi-infrastructure gateway for virtual drug screening. Concurrency and Computation: Practice and Experience, 27(16):4478–4490, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jossy:2019:CMD

[JVMN19] A. Maria Jossy, T. Vigneswaran, S. Malarvizhi, and K. K. Nagarajan. Characterization and modeling of dual material double gate tunnel field effect transistor using superposition approximation method. Concurrency and Computation: Practice and Experience, 31(14):e4860:1–e4860:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2018:MAS

[JVPI18] M. H. Jiang, Otto W. Visser, I. S. W. B. Prasetya, and Alexandru Iosup. A mirroring architecture for sophisticated mobile games using computation-offloading. Concurrency and Computation: Practice and Experience, 30(17):e4494:1– e4494:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2010:SMR

[JW10] Changsheng Jiang and Zhongliang Wu. Seismic moment release before the May 12, 2008, Wenchuan earthquake in Sichuan of southwest China. Concurrency and Computation: Practice and Experience, 22(12):1784–1795, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jin:2017:CGN

[JWW17] Yong Jin, Jialiang Weng, and Zhaoba Wang. Conjugate gradient neural network-based online recognition of glass defects. Concurrency and Computation: Practice and Experience, 29 (3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jones:2005:PPP

[JWY^{+}05] P. W. Jones, P. H. Worley, Y. Yoshida, J. B. White III, and J. Levesque. Practical performance portability in the Parallel Ocean Program (POP). Concurrency and Computation: Practice and Experience, 17(10):1317–1327, August 25, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2017:SCS

[JWY^{+}17] Jiaojiao Jiang, Sheng Wen, Shui Yu, Yang Xiang, Wanlei Zhou, and Houcine Hassan. The structure of communities in scale-free networks. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jumira:2013:EEB

[JWZ13] Oswald Jumira, Riaan Wolhuter, and Sherali Zeadally. Energy-efficient beaconless geographic routing in energy harvested wireless sensor networks. Concurrency and Computation: Practice and Experience, 25(1):58–84, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jin:2006:SIG

[JX06] Hai Jin and Zhiwei Xu. Special issue: Grid and Cooperative Computing (GCC2004). Concurrency and Computation: Practice and Experience, 18(14):1725–1728, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jin:2018:CTC

[JyLdZ^{+}18] Yong Jin, Hong ying Li, Dan dan Zhang, Jing Wu, and Maozhen Li. Compressive tracking combined with sample weights and adaptive learning factor. Concurrency and Computation: Practice and Experience, 30(23):e4398:1–e4398:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2015:LMF

[JZJW15] Jianchun Jiang, Suhua Zeng, Yanmei Jing, and Kailong Wang. A lightweight middleware framework for vehicle networking application. Concurrency and Computation: Practice and Experience, 27(17):4914–4937, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jing:2014:SSE

[JZL14] Chao Jing, Yanmin Zhu, and Minglu Li. SEED: solar energyaware efficient scheduling for data centers. Concurrency and Computation: Practice and Experience, 26(18):2811–2835, December 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jing:2015:CSA

[JZL15] Chao Jing, Yanmin Zhu, and Minglu Li. Customer satisfaction-aware scheduling for utility maximization on geodistributed data centers. Concurrency and Computation: Practice and Experience, 27(5):1334–1354, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2019:IQR

[JZMD19] Yu Jiang, Hua Zhu, Reza Malekian, and Cong Ding. An improved quantitative recurrence analysis using artificial intelligence based image processing applied to sensor measurements. Concurrency and Computation: Practice and Experience, 31(10):e4858:1–e4858:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jin:2006:CWB

[JZZL06] Hai Jin, Ran Zheng, Qin Zhang, and Ying Li. Components and workflow based Grid programming environment for integrated image-processing applications. Concurrency and Computation: Practice and Experience, 18(14):1857–1869, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kalantari:2009:GPP

[KA09] Mohammad Kalantari and Mohammad Kazem Akbari. Grid performance prediction using state-space model. Concurrency and Computation: Practice and Experience, 21(9):1109–1130, June 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kurniawan:2011:IIS

[KA11] Donny Kurniawan and David Abramson. ISENGARD: an infrastructure for supporting e-science and grid application development. Concurrency and Computation: Practice and Experience, 23(4):390–414, March 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karaata:2016:OIS

[KA16] Mehmet Hakan Karaata and Thamer Alsulaiman. An optimal inherently stabilizing routing algorithm for star P2P overlay networks. Concurrency and Computation: Practice and Experience, 28(17):4405–4428, December 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khalil:2019:DWO

[KAA19] Yasser Khalil, Mohammad Alshayeji, and Imtiaz Ahmad. Distributed Whale Optimization Algorithm based on MapReduce. Concurrency and Computation: Practice and Experience, 31(1):e4872:1–e4872:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Koyama:2007:EEZ

[KABD07] A. Koyama, J. Arai, L. Barolli, and A. Durresi. EZRP: an enhanced zone-based routing protocol for ad-hoc networks. Concurrency and Computation: Practice and Experience, 19 (8):1157–1170, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kacsuk:2011:PGP

[Kac11] Peter Kacsuk. P-GRADE portal family for grid infrastructures. Concurrency and Computation: Practice and Experience, 23(3):235–245, March 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Katsaros:2007:PET

[KAL07] Panagiotis Katsaros, Lefteris Angelis, and Constantine Lazos. Performance and effectiveness trade-off for checkpointing in fault-tolerant distributed systems. Concurrency and Computation: Practice and Experience, 19(1):37–63, January 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kalibera:2011:RRT

[Kal11] Tomas Kalibera. Replicating real-time garbage collector. Concurrency and Computation: Practice and Experience, 23 (14):1646–1664, September 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krishnamurthy:2011:TAH

[KAM11] Diwakar Krishnamurthy, Mehrnoush Alemzadeh, and Mahmood Moussavi. Towards automated HPC scheduler configuration tuning. Concurrency and Computation: Practice and Experience, 23(15):1723–1748, October 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kwack:2019:RAC

[KAMB19] JaeHyuk Kwack, Galen Arnold, Celso Mendes, and Gregory H. Bauer. Roofline analysis with Cray performance analysis tools (CrayPat) and roofline-based performance projections for a future architecture. Concurrency and Computation: Practice and Experience, 31(16):e4963:1–e4963:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Knorreck:2013:FSL

[KAP13] Daniel Knorreck, Ludovic Apvrille, and Renaud Pacalet. Formal system-level design space exploration. Concurrency and Computation: Practice and Experience, 25(2):250–264, February 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karakostas:2014:HPE

[Kar14a] Bill Karakostas. A high performance engine for concurrent complex event processing. Concurrency and Computation:

Practice and Experience, 26(2):491–499, February 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karakostas:2014:SAC

[Kar14b] Bill Karakostas. A scalable architecture for concurrent online auctions. Concurrency and Computation: Practice and Experience, 26(3):841–850, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karpowicz:2016:EEC

[Kar16] Michał P. Karpowicz. Energy-efficient CPU frequency control for the Linux system. Concurrency and Computation: Practice and Experience, 28(2):420–437, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kessler:2006:OIC

[KB06] Christoph Kessler and Andrzej Bednarski. Optimal integrated code generation for VLIW architectures. Concurrency and Computation: Practice and Experience, 18(11): 1353–1390, September 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krall:2012:SIC

[KB12] Andreas Krall and Gergö Barany. Special issue: Compilers for Parallel Computing (CPC 2010). Concurrency and Computation: Practice and Experience, 24(5):443–444, April 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kurdziel:2013:FED

[KB13] Marcin Kurdziel and Krzysztof Boryczko. Finding exemplars in dense data with affinity propagation on clusters of GPUs. Concurrency and Computation: Practice and Experience, 25 (8):1137–1152, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khalili:2017:OSW

[KB17] Azade Khalili and Seyed Morteza Babamir. Optimal scheduling workflows in cloud computing environment using Paretobased Grey Wolf Optimizer. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kwack:2018:HHB

[KB18] JaeHyuk Kwack and Gregory H. Bauer. HPCG and HPGMG benchmark tests on multiple program, multiple data (MPMD) mode on Blue Waters — a Cray XE6/XK7 hybrid system. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2011:PAP

[KBB11] Kyong Hoon Kim, Anton Beloglazov, and Rajkumar Buyya. Power-aware provisioning of virtual machines for real-time Cloud services. Concurrency and Computation: Practice and Experience, 23(13):1491–1505, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2017:DCS

[KBB17] Nirmal Kaur, Savina Bansal, and Rakesh Kumar Bansal. Duplication-controlled static energy-efficient scheduling on multiprocessor computing system. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kanthavel:2019:DKB

[KBDA19] R. Kanthavel, J. Banumathi, R. Dhaya, and Fahad Algarni. On demand knowledge-based memory-aided rebroadcast algorithm for multimedia voice data transmission in MANET. Concurrency and Computation: Practice and Experience, 31 (14):e5016:1–e5016:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kessler:2007:CGS

[KBE07] Christoph Kessler, Andrzej Bednarski, and Mattias Eriksson. Classification and generation of schedules for VLIW processors. Concurrency and Computation: Practice and Experience, 19(18):2369–2389, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Klemm:2009:RTM

[KBG^{+}09] Michael Klemm, Matthias Bezold, Stefan Gabriel, Ronald Veldema, and Michael Philippsen. Reparallelization techniques for migrating OpenMP codes in computational grids.

Concurrency and Computation: Practice and Experience, 21 (3):281–299, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kong:2015:RWS

[KBH15a] Liang Kong, Gavin Bauer, and John Hale. Robust wireless signal indoor localization. Concurrency and Computation: Practice and Experience, 27(11):2839–2850, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kunszt:2015:ISG

[KBH^{+}15b] Peter Kunszt, Lorenz Blum, Béla Hullár, Emanuel Schmid, Adam Srebniak, Witold Wolski, Bernd Rinn, Franz-Josef Elmer, Chandrasekhar Ramakrishnan, Andreas Quandt, and Lars Malmström. iPortal: the Swiss grid proteomics portal: Requirements and new features based on experience and usability considerations. Concurrency and Computation: Practice and Experience, 27(2):433–445, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kiss:2014:LSV

[KBT^{+}14] Tamas Kiss, Peter Borsody, Gabor Terstyanszky, Stephen Winter, Pamela Greenwell, Sharron McEldowney, and Hans Heindl. Large-scale virtual screening experiments on Windows Azure-based cloud resources. Concurrency and Computation: Practice and Experience, 26(10):1760–1770, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Klemm:2007:JIO

[KBVP07] Michael Klemm, Matthias Bezold, Ronald Veldema, and Michael Philippsen. JaMP: an implementation of OpenMP for a Java DSM. Concurrency and Computation: Practice and Experience, 19(18):2333–2352, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krawezik:2006:PCM

[KC06] Géraud Krawezik and Franck Cappello. Performance comparison of MPI and OpenMP on shared memory multiprocessors. Concurrency and Computation: Practice and Experience, 18 (1):29–61, January 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Konstantinidis:2013:GPU

[KC13] Elias Konstantinidis and Yiannis Cotronis. Graphics processing unit acceleration of the red/black SOR method. Concurrency and Computation: Practice and Experience, 25(8): 1107–1120, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kansal:2015:ABC

[KC15] Nidhi Jain Kansal and Inderveer Chana. Artificial bee colony based energy-aware resource utilization technique for cloud computing. Concurrency and Computation: Practice and Experience, 27(5):1207–1225, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2018:MDF

[KC18] Shishupal Kumar and Vijay Kumar Chaurasiya. A multisensor data fusion strategy for path selection in Internet-of-Things oriented wireless sensor network (WSN). Concurrency and Computation: Practice and Experience, 30(18):e4477:1– e4477:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2009:IPO

[KCB09] Minhaj Ahmad Khan, H.-P. Charles, and D. Barthou. Improving performance of optimized kernels through fast instantiations of templates. Concurrency and Computation: Practice and Experience, 21(1):59–70, January 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2017:MLR

[KCBO17] Hyunbum Kim, Jorge A. Cobb, and Jalel Ben-Othman. Maximizing the lifetime of reinforced barriers in wireless sensor networks. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaijun:2019:DBC

[KCD19] Wu Kaijun, Yu Chao, and Wang Dicong. The dynamics behaviors of Chay neuron model under different parameters. Concurrency and Computation: Practice and Experience, 31 (12):e4836:1–e4836:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumaraswamy:2019:DKS

[KCG^{+}19] Madhura Kumaraswamy, Anamika Chowdhury, Michael Gerndt, Zakaria Bendifallah, Othman Bouizi, Uldis Locans, Lubomír Ríha, Ondrej Vysocký, Martin Beseda, and Jan Zapletal. Domain knowledge specification for energy tuning. Concurrency and Computation: Practice and Experience, 31 (6):e4650:1–e4650:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2015:UWM

[KCKC15] Huioon Kim, Kyungwon Chun, Hyounggyu Kim, and Youngjoo Chung. Utilization of workflow management system for virtual machine instance management on cloud. Concurrency and Computation: Practice and Experience, 27(17): 5350–5373, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kang:2007:ARS

[KCS07] Kyungtae Kang, Yongwoo Cho, and Heonshik Shin. Augmenting Reed–Solomon coding with retransmission for error recovery in 3G video broadcasts. Concurrency and Computation: Practice and Experience, 19(8):1171–1181, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kelly:2009:LCW

[KCW09] Peter M. Kelly, Paul D. Coddington, and Andrew L. Wendelborn. Lambda calculus as a workflow model. Concurrency and Computation: Practice and Experience, 21(16): 1999–2017, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kurc:2005:SDA

[KCZ^{+}05] Tahsin Kurc, Umit Catalyurek, Xi Zhang, Joel Saltz, Ryan Martino, Mary Wheeler, Małgorzata Peszyńska, Alan Sussman, Christian Hansen, Mrinal Sen, Roustam Seifoullaev, Paul Stoffa, Carlos Torres-Verdin, and Manish Parashar. A simulation and data analysis system for large-scale, datadriven oil reservoir simulation studies. Concurrency and Computation: Practice and Experience, 17(11):1441–1467, September 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kurzak:2007:IMP

[KD07] Jakub Kurzak and Jack Dongarra. Implementation of mixed precision in solving systems of linear equations on the Cell processor. Concurrency and Computation: Practice and Experience, 19(10):1371–1385, July 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kane:2010:CSI

[KD10] Kevin Kane and Blair Dillaway. Cyclotron: a secure, isolated, virtual cycle-scavenging grid in the enterprise. Concurrency and Computation: Practice and Experience, 22(3):241–260, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kavanagh:2015:EMB

[KD15] Richard Kavanagh and Karim Djemame. An economic market for the brokering of time and budget guarantees. Concurrency and Computation: Practice and Experience, 27(3): 546–580, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kavanagh:2019:RAE

[KD19] Richard Kavanagh and Karim Djemame. Rapid and accurate energy models through calibration with IPMI and RAPL. Concurrency and Computation: Practice and Experience, 31 (13):e5124:1–e5124:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kajiyama:2017:HPC

[KDC17] Tamito Kajiyama, Davide D’Alimonte, and José C. Cunha. A high-performance computing framework for Monte Carlo ocean color simulations. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2008:PTW

[KDG^{+}08] Jihie Kim, Ewa Deelman, Yolanda Gil, Gaurang Mehta, and Varun Ratnakar. Provenance trails in the Wings/Pegasus system. Concurrency and Computation: Practice and Experience, 20(5):587–597, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumari:2017:DSU

[KDW^{+}17] Saru Kumari, Ashok Kumar Das, Mohammad Wazid, Xiong Li, Fan Wu, Kim-Kwang Raymond Choo, and Muhammad Khurram Khan. On the design of a secure user authentication and key agreement scheme for wireless sensor networks. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kessler:2004:MDS

[Kes04] Christoph W. Kessler. Managing distributed shared arrays in a bulk-synchronous parallel programming environment. Concurrency and Computation: Practice and Experience, 16(2– 3):133–153, February/March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kavas:2001:CWN

[KF01] Avi Kavas and Dror G. Feitelson. Comparing Windows NT, Linux, and QNX as the basis for cluster systems. Concurrency and Computation: Practice and Experience, 13(15):1303– 1332, December 25, 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/89016296/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=89016296&PLACEBO=IE.pdf.

Kruger:2011:IPS

[KF11] J. Krüger and G. Fels. Ion permeation simulations by Gromacs — an example of high performance molecular dynamics. Concurrency and Computation: Practice and Experience, 23 (3):279–291, March 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kadirvel:2015:TSC

[KF15] Selvi Kadirvel and José A. B. Fortes. Towards self-caring MapReduce: a study of performance penalties under faults. Concurrency and Computation: Practice and Experience, 27 (9):2310–2328, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kontoudis:2018:SAV

[KF18] Dimitris Kontoudis and Panayotis Fouliras. A statistical approach to virtual server resource management. Concurrency and Computation: Practice and Experience, 30 (4):??, February 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4335.

Krzywicki:2018:CAB

[KFD18] Daniel Krzywicki, Lukasz Faber, and Roman Debski. Concurrent agent-based evolutionary computations as adaptive dataflows. Concurrency and Computation: Practice and Experience, 30(22):e4702:1–e4702:??, November 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Klous:2006:TAG

[KFS^{+}06] Sander Klous, Jaime Frey, Se-Chang Son, Douglas Thain, Alain Roy, Miron Livny, and Jo van den Brand. Transparent access to Grid resources for user software. Concurrency and Computation: Practice and Experience, 18(7):787–801, June 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2019:WCA

[KG19] P. C. Prabhu Kumar and G. Geetha. Web-cloud architecture levels and optimized MQTT and COAP protocol suites for Web of Things. Concurrency and Computation: Practice and Experience, 31(12):e4867:1–e4867:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kehagias:2012:OBM

[KGGT12] Dionysios D. Kehagias, Konstantinos M. Giannoutakis, George A. Gravvanis, and Dimitrios Tzovaras. An ontologybased mechanism for automatic categorization of Web services. Concurrency and Computation: Practice and Experience, 24(3):214–236, March 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kalkov:2017:EPP

[KGK17] Igor Kalkov, Alexandru Gurghian, and Stefan Kowalewski. Explicit prioritization of parallel intent broadcasts in realtime Android. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2019:IAM

[KGP^{+}19] Deepa K., Radhamani G., Vinod P., Mohammad Shojafar, Neeraj Kumar, and Mauro Conti. Identification of Android malware using refined system calls. Concurrency and Computation: Practice and Experience, 31(20):e5311:1–e5311:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kato:2012:MGA

[KH12] Kimikazu Kato and Tikara Hosino. Multi-GPU algorithm for k-nearest neighbor problem. Concurrency and Computation: Practice and Experience, 24(1):45–53, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kouge:2017:ADH

[KHF^{+}17] Hiroaki Kouge, Takumi Honda, Toru Fujita, Yasuaki Ito, Koji Nakano, and Jacir L. Bordim. Accelerating digital halftoning using the local exhaustive search on the GPU. Concurrency and Computation: Practice and Experience, 29(2):??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2013:HOB

[KHHC13] Yong-Ki Kim, Amina Hossain, Al-Amin Hossain, and Jae-Woo Chang. Hilbert-order based spatial cloaking algorithm in road network. Concurrency and Computation: Practice and Experience, 25(1):143–158, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2017:OHP

[KHL^{+}17a] Mukhtaj Khan, Zhengwen Huang, Maozhen Li, Gareth A. Taylor, and Mushtaq Khan. Optimizing Hadoop parameter settings with gene expression programming guided PSO. Concurrency and Computation: Practice and Experience, 29(3): ??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2017:PSO

[KHL17b] Jae Kwon Kim, Young Shin Han, and Jong Sik Lee. Particle swarm optimization–deep belief network–based rare class prediction model for highly class imbalance problem. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kalibera:2011:FRT

[KHM^{+}11a] Tomas Kalibera, Jeff Hagelberg, Petr Maj, Filip Pizlo, Ben Titzer, and Jan Vitek. A family of real-time Java benchmarks. Concurrency and Computation: Practice and Experience, 23(14):1679–1700, September 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2011:ELA

[KHM^{+}11b] Joohyun Kim, Wei Huang, Sharath Maddineni, Fareed Aboul-ela, and Shantenu Jha. Energy landscape analysis for regulatory RNA finding using scalable distributed cyberinfrastructure. Concurrency and Computation: Practice and Experience, 23(17):2292–2304, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kuckuk:2017:TGE

[KHVK17] Sebastian Kuckuk, Gundolf Haase, Diego A. Vasco, and Harald Köstler. Towards generating efficient flow solvers with the ExaStencils approach. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kerbyson:2005:PCB

[KHW05] Darren J. Kerbyson, Adolfy Hoisie, and Harvey J. Wasserman. A performance comparison between the Earth Simulator and other terascale systems on a characteristic ASCI workload. Concurrency and Computation: Practice and Experience, 17(10):1219–1238, August 25, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krichen:2015:DRD

[KHZ^{+}15] Fatma Krichen, Brahim Hamid, Bechir Zalila, Mohamed Jmaiel, and Bernard Coulette. Development of reconfigurable distributed embedded systems with a model-driven approach.

Concurrency and Computation: Practice and Experience, 27 (6):1391–1411, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Katsinis:2006:PPM

[KHZN06] Constantine Katsinis, Diana Hecht, Ming Zhu, and Harsha Narravula. The performance of parallel matrix algorithms on a broadcast-based architecture. Concurrency and Computation: Practice and Experience, 18(3):271–303, March 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Keen:2003:CCP

[KIM^{+}03] Aaron W. Keen, Takashi Ishihara, Justin T. Maris, Tiejun Li, Eugene F. Fodor, and Ronald A. Olsson. A comparison of concurrent programming and cooperative multithreading. Concurrency and Computation: Practice and Experience, 15 (1):27–53, January 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

King:2004:SMM

[Kin04] Roger (Buzz) King. Security Maintenance Mediation: a technology for preventing unintended security breaches. Concurrency and Computation: Practice and Experience, 16(1): 49–60, January 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ke:2019:AIT

[KJ19a] Ge Ke and Qijie Jiang. Application of Internet of Things technology in the construction of wisdom museum. Concurrency and Computation: Practice and Experience, 31(10):e4680:1– e4680:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kishanlal:2019:EIM

[KJ19b] M. Samayaraj Murali Kishanlal and A. Jawahar. Enabling internal methods in passive optical network architecture for next generation OFDM–PON supporting radio-over-fiber. Concurrency and Computation: Practice and Experience, 31 (14):e5095:1–e5095:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kolaczek:2015:TBS

[KJŚ^{+}15] Grzegorz Kołaczek, Krzysztof Juszczyszyn, Paweł Światek, Adam Grzech, Patryk Schauer, Paweł Stelmach, and Łukasz Falas. Trust-based security-level evaluation method for dynamic service-oriented environments. Concurrency and Computation: Practice and Experience, 27(18):5700–5718, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kanimozhi:2019:SCB

[KKDS19] S. Kanimozhi, A. Kannan, K. Suganya Devi, and K Selvamani. Secure cloud-based e-learning system with access control and group key mechanism. Concurrency and Computation: Practice and Experience, 31(12):e4841:1–e4841:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Knijnenburg:2004:ECM

[KKGO04] P. M. W. Knijnenburg, T. Kisuki, K. Gallivan, and M. F. P. O’Boyle. The effect of cache models on iterative compilation for combined tiling and unrolling. Concurrency and Computation: Practice and Experience, 16(2–3):247–270, February/ March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2003:DIU

[KKJH03] Jin-Soo Kim, Kangho Kim, Sung-In Jung, and Soonhoi Ha. Design and implementation of a user-level Sockets layer over Virtual Interface Architecture. Concurrency and Computation: Practice and Experience, 15(7–8):727–749, June/July 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kang:2010:ARI

[KKK10] Byung-Seok Kang, Han-Shik Kim, and In-Young Ko. AODV-RIP: improved security in mobile ad hoc networks through route investigation procedure. Concurrency and Computation: Practice and Experience, 22(7):816–830, May 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khurshid:2019:SCD

[KKK^{+}19] Anum Khurshid, Abdul Nasir Khan, Fiaz Gul Khan, Mazhar Ali, Junaid Shuja, and Atta ur Rehman Khan. Secure-CamFlow: a device-oriented security model to assist information flow control systems in cloud environments for IoTs. Concurrency and Computation: Practice and Experience, 31 (8):e4729:1–e4729:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kosar:2006:BRE

[KKL06] T. Kosar, G. Kola, and M. Livny. Building reliable and efficient data transfer and processing pipelines. Concurrency and Computation: Practice and Experience, 18(6):609–620, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kuhn:2009:DFS

[KKL09] Michael Kuhn, Julian Martin Kunkel, and Thomas Ludwig. Dynamic file system semantics to enable metadata optimizations in PVFS. Concurrency and Computation: Practice and Experience, 21(14):1775–1788, September 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kosiedowski:2006:WAG

[KKM^{+}06] Michal Kosiedowski, Krzysztof Kurowski, Cezary Mazurek, Jarek Nabrzyski, and Juliusz Pukacki. Workflow applications in GridLab and PROGRESS projects. Concurrency and Computation: Practice and Experience, 18(10):1141–1154, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kotenko:2012:ABS

[KKS12] Igor Kotenko, Alexey Konovalov, and Andrey Shorov. Agentbased simulation of cooperative defence against botnets. Concurrency and Computation: Practice and Experience, 24(6): 573–588, May 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kolodziej:2013:ESO

[KKT13] Joanna Kołodziej, Samee Ullah Khan, and El-Ghazali Talbi. Editorials: Scalable optimization in grid, cloud, and intelligent network computing – foreword. Concurrency and Computation: Practice and Experience, 25(12):1719–1721, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kanso:2013:ACG

[KKTHL13] A. Kanso, F. Khendek, M. Toeroe, and A. Hamou-Lhadj. Automatic configuration generation for service high availability with load balancing. Concurrency and Computation: Practice and Experience, 25(2):265–287, February 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kretsis:2013:IES

[KKV13] A. Kretsis, P. Kokkinos, and E. A. Varvarigos. Implementing and evaluating scheduling policies in gLite middleware. Concurrency and Computation: Practice and Experience, 25(3): 349–366, March 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Koesterke:2014:DBN

[KKW^{+}14] Lars Koesterke, James E. Koltes, Nathan T. Weeks, Kent Milfeld, Matthew W. Vaughn, James M. Reecy, and Dan Stanzione. Discovery of biological networks using an optimized partial correlation coefficient with information theory algorithm on Stampede’s Xeon and Xeon Phi processors. Concurrency and Computation: Practice and Experience, 26 (13):2178–2190, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kolodziej:2015:EEG

[KKWZ15] Joanna Kołodziej, Samee Ullah Khan, Lizhe Wang, and Albert Y. Zomaya. Energy efficient genetic-based schedulers in computational grids. Concurrency and Computation: Practice and Experience, 27(4):809–829, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kazi:2002:DAS

[KL02] Iffat H. Kazi and David J. Lilja. Dynamically adapting to system load and program behavior in multiprogrammed multiprocessor systems. Concurrency and Computation: Practice and Experience, 14(12):957–985, October 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/98516161/START;   http:

//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
98516161{\&}PLACEBO=IE.pdf.

Kessler:2012:OCP

[KL12a] C. Kessler and W. Löwe. Optimized composition of performance-aware parallel components. Concurrency and Computation: Practice and Experience, 24(5):481–498, April 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kuan:2012:CSV

[KL12b] Chi-Bang Kuan and Jenq Kuen Lee. Compiler supports for VLIW DSP processors with SIMD intrinsics. Concurrency and Computation: Practice and Experience, 24(5):517–532, April 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kandavalli:2019:DIC

[KL19] Michael Angelo Kandavalli and S. Abraham Lincon. Design and implementation of colour texture-based multiple object detection using morphological gradient approach. Concurrency and Computation: Practice and Experience, 31(14): e4980:1–e4980:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kurzak:2010:SDL

[KLDB10] Jakub Kurzak, Hatem Ltaief, Jack Dongarra, and Rosa M. Badia. Scheduling dense linear algebra operations on multicore processors. Concurrency and Computation: Practice and Experience, 22(1):15–44, January 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kotowski:2008:PQP

[KLP^{+}08] Nelson Kotowski, Alexandre A. B. Lima, Esther Pacitti, Patrick Valduriez, and Marta Mattoso. Parallel query processing for OLAP in grids. Concurrency and Computation: Practice and Experience, 20(17):2039–2048, December 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kermarrec:2003:PHA

[KM03] Anne-Marie Kermarrec and Christine Morin. HA-PSLS: a highly available parallel single-level store system. Concurrency and Computation: Practice and Experience, 15(10): 911–937, August 25, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2013:SIP

[KM13] Khaled M. Khan and Qutaibah Malluhi. Special issue papers: Role of contextual properties in enterprise service migration to cloud computing. Concurrency and Computation: Practice and Experience, 25(18):2455–2470, December 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karthick:2019:FAI

[KM19] T Karthick and M Manikandan. Fog assisted IoT based medical cyber system for cardiovascular diseases affected patients. Concurrency and Computation: Practice and Experience, 31 (12):e4861:1–e4861:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kouvatsos:2004:OAB

[KMA04] D. D. Kouvatsos, I. M. Mkwawa, and I. Awan. One-to-all broadcasting scheme for arbitrary static interconnection networks with bursty background traffic. Concurrency and Computation: Practice and Experience, 16(13):1365–1383, November 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kardani-Moghaddam:2019:PAD

[KMBR19] Sara Kardani-Moghaddam, Rajkumar Buyya, and Kotagiri Ramamohanarao. Performance anomaly detection using isolation-trees in heterogeneous workloads of web applications in computing clouds. Concurrency and Computation: Practice and Experience, 31(20):e5306:1–e5306:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kimovski:2018:DEE

[KMG^{+}18] Dragi Kimovski, Attila Marosi, Sandi Gec, Nishant Saurabh, Attila Kertesz, Gabor Kecskemeti, Vlado Stankovski, and Radu Prodan. Distributed environment for efficient virtual machine image management in federated Cloud architectures. Concurrency and Computation: Practice and Experience, 30 (20):e4220:1–e4220:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2014:ANG

[KMJ14] Joohyun Kim, Sharath Maddineni, and Shantenu Jha. Advancing next-generation sequencing data analytics with scalable distributed infrastructure. Concurrency and Computation: Practice and Experience, 26(4):894–906, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2017:OMF

[KMJ^{+}17] Fiaz Gul Khan, Bartolomeo Montrucchio, Bilal Jan, Abdul Nasir Khan, Waqas Jadoon, Shahaboddin Shamshirband, Anthony Theodore Chronopoulos, and Iftikhar Ahmed Khan. An optimized magnetostatic field solver on GPU using open computing language. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kraemer:2018:RNR

[KMRT18] Alessandro Kraemer, Carlos Maziero, Olivier Richard, and Denis Trystram. Reducing the number of response time service level objective violations by a cloud-HPC convergence scheduler. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4352.

Klein:2001:VLB

[KN01] Gerwin Klein and Tobias Nipkow. Verified lightweight bytecode verification. Concurrency and Computation: Practice and Experience, 13(13):1133–1151, November 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/88011337/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=88011337&PLACEBO=IE.
pdf.

Kumar:2019:DII

[KN19] G. S. Satheesh Kumar and C. Nagarajan. Design and implementation of intelligent energy control for tariff management using multi-agent system in smart/micro grid. Concurrency and Computation: Practice and Experience, 31(12):e4846:1– e4846:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Knijnenburg:2006:SII

[Kni06] Peter M. W. Knijnenburg. Special issue: 10th International Workshop on Compilers for Parallel Computers (CPC 2003). Concurrency and Computation: Practice and Experience, 18 (11):1333–1334, September 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaieda:2001:AME

[KNT^{+}01] Akihiro Kaieda, Yasuichi Nakayama, Atsuhiro Tanaka, Takashi Horikawa, Toshiyasu Kurasugi, and Issei Kino. Analysis and measurement of the effect of kernel locks in SMP systems. Concurrency and Computation: Practice and Experience, 13(2):141–152, February 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/77004420/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=77004420&PLACEBO=IE.
pdf.

Koide:2006:NTS

[KO06] Hiroshi Koide and Yuji Oie. A new task scheduling method for distributed programs that require memory management. Concurrency and Computation: Practice and Experience, 18 (9):941–958, August 10, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kambites:2001:OLI

[KOB01] M. E. Kambites, J. Obdržálek, and J. M. Bull. An OpenMPlike interface for parallel programming in Java. Concurrency and Computation: Practice and Experience, 13(8–9): 793–814, July/August 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/84503220/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=84503220&PLACEBO=IE.pdf.

Kertesz:2014:CSB

[KOK14] Attila Kertesz, Ferenc Otvos, and Peter Kacsuk. A case study for biochemical application porting in European grids and clouds. Concurrency and Computation: Practice and Experience, 26(10):1730–1743, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kakuda:2012:MRT

[KOO12] Yoshiaki Kakuda, Tomoyuki Ohta, and Ryotaro Oda. A methodology for real-time self-organized autonomous clustering in mobile ad hoc networks. Concurrency and Computation: Practice and Experience, 24(16):1840–1859, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kimovski:2015:LCP

[KOOB15] Dragi Kimovski, Julio Ortega, Andrés Ortiz, and Raúl Baños. Leveraging cooperation for parallel multi-objective feature selection in high-dimensional EEG data. Concurrency and Computation: Practice and Experience, 27(18):5476–5499, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kromer:2018:AQO

[KPNS18] Pavel Krömer, Jan Platoš, Jana Nowaková, and Václav Snášel. An acceleration of quasigroup operations by residue arithmetic. Concurrency and Computation: Practice and Experience, 30(2):??, January 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kromer:2014:DPD

[KPS14] Pavel Krömer, Jan Platoš, and Václav Snášel. Data parallel density-based genetic clustering on CUDA architecture. Concurrency and Computation: Practice and Experience, 26(5): 1097–1112, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2017:TES

[KQR^{+}17] Jik-Soo Kim, Bui Quang, Seungwoo Rho, Seoyoung Kim, Sangwan Kim, Vincent Breton, and Soonwook Hwang. Towards effective scheduling policies for many-task applications: Practice and experience based on HTCaaS. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Korch:2004:CTP

[KR04] Matthias Korch and Thomas Rauber. A comparison of task pools for dynamic load balancing of irregular algorithms. Concurrency and Computation: Practice and Experience, 16 (1):1–47, January 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kielmann:2006:SIA

[KR06] Thilo Kielmann and Wilson Rivera. Special issue: Adaptive Grid middleware. Concurrency and Computation: Practice and Experience, 18(13):1657–1658, November 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kunis:2011:OLB

[KR11] R. Kunis and G. Rünger. Optimizing layer-based scheduling algorithms for parallel tasks with dependencies. Concurrency and Computation: Practice and Experience, 23(8):827–849, June 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khabou:2015:TBC

[KR15] Nesrine Khabou and Ismael Bouassida Rodriguez. Thresholdbased context analysis approach for ubiquitous systems. Concurrency and Computation: Practice and Experience, 27(6): 1378–1390, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krishnan:2005:GGB

[Kri05]  Arun Krishnan.  GridBLAST: a Globus-based high-

throughput implementation of BLAST in a Grid computing framework. Concurrency and Computation: Practice and Experience, 17(13):1607–1623, November 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kristensen:2013:RBC

[Kri13] Bent Bruun Kristensen. Rendezvous-based collaboration between autonomous entities: centric versus associative. Concurrency and Computation: Practice and Experience, 25(3): 289–308, March 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kristiansen:2019:UEA

[Kri19] Alex Kristiansen. Use of the ERD for administrative monitoring of Theta. Concurrency and Computation: Practice and Experience, 31(16):e5099:1–e5099:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Korch:2011:SLE

[KRS11] Matthias Korch, Thomas Rauber, and Carsten Scholtes. Scalability and locality of extrapolation methods on large parallel systems. Concurrency and Computation: Practice and Experience, 23(15):1789–1815, October 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaczmarski:2017:PAC

[KRW17] Krzysztof Kaczmarski, Paweł Rzażewski, and Albert Wolant. Parallel algorithms constructing the cell graph. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kennedy:2002:SIH

[KS02] K. Kennedy and Y. Seo. Special issue: High Performance Fortran. Concurrency and Computation: Practice and Experience, 14(8–9):551–553, July/August 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/95016134/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=95016134{\&}PLACEBO=
IE.pdf.

Kennedy:2004:TFC

[KS04] Andrew Kennedy and Don Syme. Transposing F to C: expressivity of parametric polymorphism in an object-oriented language. Concurrency and Computation: Practice and Experience, 16(7):707–733, June 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kuchen:2005:FFP

[KS05] H. Kuchen and J. Striegnitz. Features from functional programming for a C++ skeleton library. Concurrency and Computation: Practice and Experience, 17(7–8):739–756, June/ July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2019:EMR

[KS19a] Akila K and Chitrakala S. An efficient method to resolve intraclass variability using highly refined HOG description model for human action recognition. Concurrency and Computation: Practice and Experience, 31(12):e4856:1–e4856:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kalra:2019:MCW

[KS19b] Mala Kalra and Sarbjeet Singh. Multi-criteria workflow scheduling on clouds under deadline and budget constraints. Concurrency and Computation: Practice and Experience, 31(17):e5193:1–e5193:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kohler:2019:FSE

[KS19c] Martin Köhler and Jens Saak. Frequency scaling and energy efficiency regarding the Gauss–Jordan elimination scheme with application to the matrix-sign-function on Open-POWER 8. Concurrency and Computation: Practice and Experience, 31(6):e4504:1–e4504:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kurte:2019:PAO

[KSB^{+}19] Kuldeep Kurte, Jibonananda Sanyal, Anne Berres, Dalton Lunga, Mark Coletti, Hsiuhan Lexie Yang, Daniel Graves, Benjamin Liebersohn, and Amy Rose. Performance analysis and optimization for scalable deployment of deep learning models for country-scale settlement mapping on Titan supercomputer. Concurrency and Computation: Practice and Experience, 31(20):e5305:1–e5305:??, October 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kwak:2012:DIB

[KSC12] Hukeun Kwak, Andrew Sohn, and Kyusik Chung. Dynamic information-based scalable hashing on a cluster of Web cache servers. Concurrency and Computation: Practice and Experience, 24(3):322–340, March 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kegel:2011:CPM

[KSG11] Philipp Kegel, Maraike Schellmann, and Sergei Gorlatch. Comparing programming models for medical imaging on multi-core systems. Concurrency and Computation: Practice and Experience, 23(10):1051–1065, July 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2017:SFG

[KSK17] Ashok Kumar, Anju Sharma, and Rajesh Kumar. SERVmegh: framework for green cloud. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khalid:2019:IUF

[KSK19] Adnan Khalid, Muhammad Shahbaz, and Imran Ahmed Khan. Intelligent use of fog devices in edge-cloud paradigm to assist in E-polling. Concurrency and Computation: Practice and Experience, 31(8):e4957:1–e4957:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kharche:2008:SCE

[KSM^{+}08a] Sanjay Kharche, Gunnar Seemann, Lee Margetts, Joanna Leng, Arun V. Holden, and Henggui Zhang. Simulation of clinical electrophysiology in 3D human atria: a highperformance computing and high-performance visualization application. Concurrency and Computation: Practice and Experience, 20(11):1317–1328, August 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krenek:2008:GJP

[KSM^{+}08b] Aleš Křenek, Jiří Sitera, Ludĕk Matyska, František Dvořák, Miloš Mulač, Miroslav Ruda, and Zdenĕk Salvet. gLite Job Provenance — a job-centric view. Concurrency and Computation: Practice and Experience, 20(5):453–462, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kowalik:2015:IGH

[KSM15] Grzegorz Tomasz Kowalik, Jennifer Anne Steeden, and Vivek Muthurangu. Implementation of a generalized heterogeneous image reconstruction system for clinical magnetic resonance. Concurrency and Computation: Practice and Experience, 27 (6):1603–1611, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kurth:2019:TSP

[KSM^{+}19] Thorsten Kurth, Mikhail Smorkalov, Peter Mendygral, Srinivas Sridharan, and Amrita Mathuriya. TensorFlow at scale: Performance and productivity analysis of distributed training with Horovod, MLSL, and Cray PE ML. Concurrency and Computation: Practice and Experience, 31(16):e4989:1– e4989:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karlsen:2016:DIC

[KSN16] R. Karlsen, D. Sundby, and J. Nordbotten. Development of image collection representations for intelligent distributed systems. Concurrency and Computation: Practice and Experience, 28(4):1336–1355, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kloh:2012:BCS

[KSPM12] Henrique Kloh, Bruno Schulze, Raquel Pinto, and Antonio Mury. A bi-criteria scheduling process with CoS support on grids and clouds. Concurrency and Computation: Practice and Experience, 24(13):1443–1460, September 10, 2012. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Korsholm:2014:RTJ

[KSR14] Stephan E. Korsholm, Hans Søndergaard, and Anders P. Ravn. A real-time Java tool chain for resource constrained platforms. Concurrency and Computation: Practice and Experience, 26(14):2407–2431, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kienberger:2017:PHC

[KSS^{+}17] Julian Kienberger, Stefan Schmidhuber, Christian Saad, Stefan Kuntz, and Bernhard Bauer. Parallelizing highly complex engine management systems. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kanithan:2019:NEB

[KT19a] S. Kanithan and G. M. Tamilselvan. A novel enhanced bat optimization based energy efficiency and imperfect channel state information in cooperative MIMO–AF systems. Concurrency and Computation: Practice and Experience, 31(14): e4839:1–e4839:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kodabagi:2019:CBA

[KT19b] Mallikarjun M. Kodabagi and Ahelam Tikotikar. Clusteringbased approach for medical data classification. Concurrency and Computation: Practice and Experience, 31(14):e5079:1– e5079:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2019:PSS

[KT19c] N. Suresh Kumar and M. Thangamani. Parallel Semisupervised enhanced fuzzy Co-Clustering (PSEFC) and Rapid Association Rule Mining (RARM) based frequent route mining algorithm for travel sequence recommendation on big social media. Concurrency and Computation: Practice and Experience, 31(14):e4837:1–e4837:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2004:RMF

[KTB04] Raj Kumar, Vanish Talwar, and Sujoy Basu. A resource management framework for interactive Grids. Concurrency and Computation: Practice and Experience, 16(5):489–501, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khosravi:2017:OVM

[KTB17] Atefeh Khosravi, Adel Nadjaran Toosi, and Rajkumar Buyya. Online virtual machine migration for renewable energy usage maximization in geographically distributed cloud data centers. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karakus:2018:EDL

[KTHA18] Betül Ay Karakus, Muhammed Talo, Ibrahim Riza Hallaç, and Galip Aydin. Evaluating deep learning models for sentiment classification. Concurrency and Computation: Practice and Experience, 30(21):e4783:1–e4783:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kyriazis:2009:SSW

[KTM^{+}09] Dimosthenis Kyriazis, Konstantinos Tserpes, Andreas Menychtas, Ioannis Sarantidis, and Theodora Varvarigou. Service selection and workflow mapping for Grids: an approach exploiting quality-of-service information. Concurrency and Computation: Practice and Experience, 21(6):739–766, April 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2011:DSU

[KTR11] Dong Kwan Kim, Eli Tilevich, and Calvin J. Ribbens. Dynamic software updates for parallel high-performance applications. Concurrency and Computation: Practice and Experience, 23(4):415–434, March 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2018:EPS

[KTZ^{+}18] Naveen Kumar, Shailesh Tiwari, Zhigao Zheng, Krishn K. Mishra, and Arun Kumar Sangaiah. An efficient and provably secure time-limited key management scheme for outsourced data. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4498.

Kulakowski:2014:CVE

[Kuł14] Konrad Kułakowski. A concurrent van Emde Boas array as a fast and simple concurrent dynamic set alternative. Concurrency and Computation: Practice and Experience, 26(2): 360–379, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kashyap:2012:SAS

[KV12] Rekha Kashyap and Deo Prakash Vidyarthi. Security-aware scheduling model for computational grid. Concurrency and Computation: Practice and Experience, 24(12):1377–1391, August 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kolonias:2011:DIE

[KVGH11] Vasileios Kolonias, Artemios G. Voyiatzis, George Goulas, and Efthymios Housos. Design and implementation of an efficient integer count sort in CUDA GPUs. Concurrency and Computation: Practice and Experience, 23(18):2365–2381, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kersten:2014:RRA

[KvGS^{+}14] Rody W. J. Kersten, Bernard E. van Gastel, Olha Shkaravska, Manuel Montenegro, and Marko C. J. D. van Eekelen. ResAna: a resource analysis toolset for (real-time) JAVA. Concurrency and Computation: Practice and Experience, 26 (14):2432–2455, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krintz:2001:UJC

[KW01] Chandra Krintz and Rich Wolski. Using JavaNws to compare C and Java TCP-Socket performance. Concurrency and Computation: Practice and Experience, 13(8–9):815–839, July/ August 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/84503224/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
84503224&PLACEBO=IE.pdf.

Kim:2011:AFP

[KW11] MinSeong Kim and Andy Wellings. Applying fixed-priority preemptive scheduling with preemption threshold to asynchronous event handling in the RTSJ. Concurrency and Computation: Practice and Experience, 23(14):1609–1622, September 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Korch:2018:AEO

[KW18] Matthias Korch and Tim Werner. Accelerating explicit ODE methods on GPUs by kernel fusion. Concurrency and Computation: Practice and Experience, 30(18):e4470:1–e4470:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaczmarski:2019:GRT

[KW19a] Krzysztof Kaczmarski and Albert Wolant. GPU R-Trie: Dictionary with ultra fast lookup. Concurrency and Computation: Practice and Experience, 31(19):e5027:1–e5027:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krestenitis:2019:MCR

[KW19b] Konstantinos Krestenitis and Tobias Weinzierl. A multi-core ready discrete element method with triangles using dynamically adaptive multiscale grids. Concurrency and Computation: Practice and Experience, 31(19):e4935:1–e4935:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kwon:2005:RJH

[KWK05] Jagun Kwon, Andy Wellings, and Steve King. Ravenscar– Java: a high-integrity profile for real-time Java. Concurrency and Computation: Practice and Experience, 17(5–6):681–713, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Keahey:2004:FGA

[KWL^{+}04] K. Keahey, V. Welch, S. Lang, B. Liu, and S. Meder. Finegrained authorization for job execution in the Grid: design and implementation. Concurrency and Computation: Practice and Experience, 16(5):477–488, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kutylowski:2018:SIS

[KWXY18] Miroslaw Kutylowski, Yu Wang, Shouhuai Xu, and Laurence T. Yang. Special issue on social network security and privacy. Concurrency and Computation: Practice and Experience, 30(5):??, March 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4414.

Kudo:2017:PAO

[KYBV17] Shuhei Kudo, Yusaku Yamamoto, Martin Bečka, and Marian Vajtersic. Performance analysis and optimization of the parallel one-sided block Jacobi SVD algorithm with dynamic ordering and variable blocking. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khani:2017:SOL

[KYM17] Hadi Khani, Nasser Yazdani, and Siamak Mohammadi. A self-organized load balancing mechanism for cloud computing. Concurrency and Computation: Practice and Experience, 29 (4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kong:2015:NBM

[KZY15] Yan Kong, Minjie Zhang, and Dayong Ye. A negotiationbased method for task allocation with time constraints in open grid environments. Concurrency and Computation: Practice and Experience, 27(3):735–761, March 10, 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kong:2018:IAB

[KZY^{+}18] Yan Kong, Minjie Zhang, Dayong Ye, Jinxiu Zhu, and Junho Choi. An intelligent agent-based method for task allocation in competitive cloud environments. Concurrency and Computation: Practice and Experience, 30(3):??, February 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4178.

Ludascher:2006:SWM

[LAB^{+}06] Bertram Ludäscher, Ilkay Altintas, Chad Berkley, Dan Higgins, Efrat Jaeger, Matthew Jones, Edward A. Lee, Jing Tao, and Yang Zhao. Scientific workflow management and the Kepler system. Concurrency and Computation: Practice and Experience, 18(10):1039–1065, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lei:2008:GEP

[LAC^{+}08] Zhou Lei, Gabrielle Allen, Promita Chakraborty, Dayong Huang, John Lewis, Xin Li, and Christopher D. White. A Grid-enabled problem-solving environment for advanced reservoir uncertainty analysis. Concurrency and Computation: Practice and Experience, 20(18):2123–2140, December 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lobosco:2002:JHP

[LAL02] M. Lobosco, C. Amorim, and O. Loques. Java for highperformance network-based computing: a survey. Concurrency and Computation: Practice and Experience, 14(1):1–31, January 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/91014114/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
91014114&PLACEBO=IE.pdf.

Liu:2009:ERD

[LAM^{+}09] Lu Liu, Nick Antonopoulos, Stephen Mackin, Jie Xu, and Duncan Russell. Efficient resource discovery in self-organized unstructured peer-to-peer networks. Concurrency and Computation: Practice and Experience, 21(2):159–183, February 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lang:2017:DAT

[Lan17] Jens Lang. Data-aware tuning of scientific applications with model-based autotuning. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luitjens:2011:SPR

[LB11] J. Luitjens and M. Berzins. Scalable parallel regridding algorithms for block-structured adaptive mesh refinement. Concurrency and Computation: Practice and Experience, 23(13): 1522–1537, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:FSI

[LB19] Gang Li and Lynn Batten. Foreword to the special issue of the 8th International Conference on Applications and Techniques for Information Security (ATIS 2017). Concurrency and Computation: Practice and Experience, 31(24):e5411:1– e5411:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Leite:2016:PAS

[LBdM^{+}16] Alessandro Ferreira Leite, Azzedine Boukerche, Alba Cristina Magalhaes Alves de Melo, Christine Eisenbeis, Claude Tadonki, and Célia Ghedini Ralha. Power-aware server consolidation for federated clouds. Concurrency and Computation: Practice and Experience, 28(12):3427–3444, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Longo:2015:VOC

[LBDS15] Francesco Longo, Dario Bruneo, Salvatore Distefano, and Marco Scarpa. Variable operating conditions in distributed systems: modeling and evaluation. Concurrency and Computation: Practice and Experience, 27(10):2506–2530, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lengauer:2017:ESIb

[LBFS17] Christian Lengauer, Matthias Bolten, Robert Falgout, and Olaf Schenk. Editorial: Special issue: Advanced stencilcode engineering. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luitjens:2007:PSF

[LBH07] J. Luitjens, M. Berzins, and T. Henderson. Parallel spacefilling curve generation through sorting. Concurrency and Computation: Practice and Experience, 19(10):1387–1402, July 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Larrea:2019:WWS

[LBJ^{+}19] Verónica G. Vergara Larrea, Michael J. Brim, Wayne Joubert, Swen Boehm, Matthew Baker, Oscar Hernandez, Sarp Oral, James Simmons, and Don Maxwell. Are we witnessing the spectre of an HPC meltdown? Concurrency and Computation: Practice and Experience, 31(16):e5020:1–e5020:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lachhab:2018:PEL

[LBOE18] Fadwa Lachhab, Mohamed Bakhouya, Radouane Ouladsine, and Mohammed Essaaidi. Performance evaluation of linked stream data processing engines for situational awareness applications. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4380.

Lengauer:2015:ESI

[LBS15] Christian Lengauer, Luc Bougé, and Fernando Silva. Editorials: Special issue: Euro-Par 2014. Concurrency and Computation: Practice and Experience, 27(16):4073–4074, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lengauer:2016:ESI

[LBT16] Christian Lengauer, Luc Bougé, and Jesper Larsson Träff. Editorials: Special issue: Euro-Par 2015. Concurrency and Computation: Practice and Experience, 28(12):3445–3446, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lengauer:2017:ESIa

[LBT17] Christian Lengauer, Luc Bougé, and Denis Trystram. Editorial: Special issue: Euro-Par 2016. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lanc:2014:ABA

[LBTE14] Irena Lanc, Peter Bui, Douglas Thain, and Scott Emrich. Adapting bioinformatics applications for heterogeneous systems: a case study. Concurrency and Computation: Practice and Experience, 26(4):866–877, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2016:HAA

[LBV16] Kevin Lee, Georg Buss, and Daniel Veit. A heuristic approach for the allocation of resources in large-scale computing infrastructures. Concurrency and Computation: Practice and Experience, 28(5):1527–1547, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lengauer:2014:ESI

[LBW14] Christian Lengauer, Luc Bougé, and Felix Wolf. Editorials: Special issue: Euro-Par 2013. Concurrency and Computation: Practice and Experience, 26(14):2345–2346, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2016:HBH

[LBY^{+}16] Chang Liu, Nick Beaugeard, Chi Yang, Xuyun Zhang, and Jinjun Chen. HKE-BC: hierarchical key exchange for secure scheduling and auditing of big data in cloud computing. Concurrency and Computation: Practice and Experience, 28(3): 646–660, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2009:EGA

[LC09] Zhi-Jie Li and Chun-Tian Cheng. An evolutionary game algorithm for grid resource allocation under bounded rationality. Concurrency and Computation: Practice and Experience, 21(9):1205–1223, June 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luan:2017:EAQ

[LC17] Hua Luan and Lei Chang. An evaluation of analytical queries on CPUs and coupled GPUs. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2018:GSI

[LC18] Zhengyu Liu and Zhiqiang Chang. Geometric skew image classification algorithm based on fusion adaptive reasoning. Concurrency and Computation: Practice and Experience, 30(22):e4674:1–e4674:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lant:2019:ESM

[LCA^{+}19] Joshua Lant, Caroline Concatto, Andrew Attwood, Jose A. Pascual, Mike Ashworth, Javier Navaridas, Mikel Luján, and John Goodacre. Enabling shared memory communication in networks of MPSoCs. Concurrency and Computation: Practice and Experience, 31(21):e4774:1–e4774:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luecke:2003:MCT

[LCC^{+}03] Glenn Luecke, Hua Chen, James Coyle, Jim Hoekstra, Marina Kraeva, and Yan Zou. MPI-CHECK: a tool for checking Fortran 90 MPI programs. Concurrency and Computation: Practice and Experience, 15(2):93–100, February 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2018:DSE

[LCC^{+}18] Guan-Chen Li, Chin-Ling Chen, Hsing-Chung Chen, Feng Lin, and Cheng Gu. Design of a secure and effective medical cyber-physical system for ubiquitous telemonitoring pregnancy. Concurrency and Computation: Practice and Experience, 30(2):??, January 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lim:2005:CCH

[LCFkL05] Sang Boem Lim, Bryan Carpenter, Geoffrey Fox, and Han ku Lee. Collective communication for the HPJava programming language. Concurrency and Computation: Practice and Experience, 17(7–8):867–894, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luecke:2006:SSD

[LCH^{+}06] Glenn R. Luecke, James Coyle, Jim Hoekstra, Marina Kraeva, Ying Li, Olga Taborskaia, and Yanmei Wang. A survey of systems for detecting serial run-time errors. Concurrency and Computation: Practice and Experience, 18(15):1885–1907, December 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lugang:2014:CCM

[LCJ14] Zhao Lugang, Wu Chengke, and Ning Jifeng. A camera calibration method based on two orthogonal vanishing points. Concurrency and Computation: Practice and Experience, 26 (5):1185–1199, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2012:DHP

[LCM12] Yong-Hwan Lee, Young-Sung Cho, and Sangook Moon. Design of a high precision logarithmic converter in a binary floating point divider. Concurrency and Computation: Practice and Experience, 24(4):342–353, March 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:CCF

[LCM^{+}17] Jenny S. Li, Li-Chiou Chen, John V. Monaco, Pranjal Singh, and Charles C. Tappert. A comparison of classifiers and features for authorship authentication of social networking messages. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2013:ASA

[LCMY13] Yanhui Lu, Huijin Cao, Xiaomin Mu, and Shouyi Yang. Adaptive spectrum access strategies in the context of spectrum fragmentation in cognitive radio networks. Concurrency and Computation: Practice and Experience, 25(9):1101–1112, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:SEE

[LCT16] Xiaosong Li, Wentong Cai, and Stephen John Turner. Supporting efficient execution of continuous space agent-based simulation on GPU. Concurrency and Computation: Practice and Experience, 28(12):3313–3332, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2017:APD

[LCW^{+}17] Jiaxiang Lin, Riqing Chen, Liping Wu, Zhaogang Shu, and Changcai Yang. Adaptive parallel Delaunay triangulation construction with dynamic pruned binary tree model in Cloud. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:RDO

[LCW^{+}19] Pengfei Li, Hongyun Chen, Shangshang Wang, Zhipeng Xu, and Zhenyu Lei. Research on dynamic optimal control strategy of distributed super capacitor energy storage system based on convolution neural network. Concurrency and Computation: Practice and Experience, 31(10):e4873:1–e4873:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2008:TMS

[LCYJ08] Ke Liu, Jinjun Chen, Yun Yang, and Hai Jin. A throughput maximization strategy for scheduling transaction-intensive workflows on SwinDeW-G. Concurrency and Computation: Practice and Experience, 20(15):1807–1820, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2014:EET

[LDPZ14] Wei Li, Flávia C. Delicato, Paulo F. Pires, and Albert Y. Zomaya. Energy-efficient task allocation with quality of service provisioning for concurrent applications in multifunctional wireless sensor network systems. Concurrency and Computation: Practice and Experience, 26(11):1869–1888, August 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lupu:2008:AAM

[LDS^{+}08] E. Lupu, N. Dulay, M. Sloman, J. Sventek, S. Heeps, S. Strowes, K. Twidle, S.-L. Keoh, and A. Schaeffer-Filho. AMUSE: autonomic management of ubiquitous e-Health systems. Concurrency and Computation: Practice and Experience, 20(3):277–295, March 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2013:SIP

[LDXC13] Wenmin Lin, Wanchun Dou, Zhanyang Xu, and Jinjun Chen. Special issue papers: A QoS-aware service discovery method for elastic cloud computing in an unstructured peer-to-peer network. Concurrency and Computation: Practice and Experience, 25(13):1843–1860, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2014:PSC

[LDZ^{+}14a] Jiguo Li, Haiting Du, Yichen Zhang, Tao Li, and Yuexin Zhang. Provably secure certificate-based key-insulated signature scheme. Concurrency and Computation: Practice and Experience, 26(8):1546–1560, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2014:EPI

[LDZ14b] Rongchun Li, Yong Dou, and Dan Zou. Efficient parallel implementation of three-point Viterbi decoding algorithm on CPU, GPU, and FPGA. Concurrency and Computation: Practice and Experience, 26(3):821–840, March 10, 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:EGP

[LDZ^{+}15] Rongchun Li, Yong Dou, Dan Zou, Shi Wang, and Ying Zhang. Efficient graphics processing unit based layered decoders for quasicyclic low-density parity-check codes. Concurrency and Computation: Practice and Experience, 27(1): 29–46, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2019:AAP

[LDZ^{+}19] Bo Liu, Ming Ding, Tianqing Zhu, Yong Xiang, and Wanlei Zhou. Adversaries or allies? Privacy and deep learning in big data era. Concurrency and Computation: Practice and Experience, 31(19):e5102:1–e5102:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2009:ESI

[Lee09] C. A. Lee. Editorial: a special issue from the Open Grid Forum. Concurrency and Computation: Practice and Experience, 21(8):923–925, June 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Leymann:2006:CGT

[Ley06] Frank Leymann. Choreography for the Grid: towards fitting BPEL to the resource framework. Concurrency and Computation: Practice and Experience, 18(10):1201–1217, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lang:2015:CMS

[LF15] Fabian Lang and Andreas Fink. Collaborative machine scheduling: Challenges of individually optimizing behavior. Concurrency and Computation: Practice and Experience, 27 (11):2869–2888, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Limmer:2017:CCP

[LF17] Steffen Limmer and Dietmar Fey. Comparison of common parallel architectures for the execution of the island model and the global parallelization of evolutionary algorithms. Concurrency and Computation: Practice and Experience, 29 (9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lujan:2005:CNR

[LFG05] Mikel Luján, T. L. Freeman, and John R. Gurd. On the conditions necessary for removing abstraction penalties in O O L A L A. Concurrency and Computation: Practice and Experience, 17(7–8):839–866, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2008:SOB

[LFH08a] Bo Liu, Yushun Fan, and Shuangxi Huang. A service-oriented business performance evaluation model and the performanceaware service selection method. Concurrency and Computation: Practice and Experience, 20(15):1821–1836, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2008:SRS

[LFH^{+}08b] Xiangfeng Luo, Ning Fang, Bo Hu, Kai Yan, and Huizhe Xiao. Semantic representation of scientific documents for the e-science Knowledge Grid. Concurrency and Computation: Practice and Experience, 20(7):839–862, May 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2015:FAT

[LFHT15] Yi Liu, Yushun Fan, Keman Huang, and Wei Tan. Failure analysis and tolerance strategies in web service ecosystems. Concurrency and Computation: Practice and Experience, 27 (5):1355–1374, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lage-Freitas:2017:CRM

[LFPP17] André Lage-Freitas, Nikos Parlavantzas, and Jean Pazat. Cloud resource management driven by profit augmentation. Concurrency and Computation: Practice and Experience, 29 (4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2019:AFA

[LFW19] Wenqi Liu, MingYu Fan, and Guangwei Wang. Adaptive flow abnormity identification based on information entropy. Concurrency and Computation: Practice and Experience, 31(10): e4713:1–e4713:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liang:2015:CPA

[LFWS15] Kaitai Liang, Liming Fang, Duncan S. Wong, and Willy Susilo. A ciphertext-policy attribute-based proxy reencryption scheme for data sharing in public clouds. Concurrency and Computation: Practice and Experience, 27(8): 2004–2027, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2008:ESE

[LFX^{+}08] Xiangfeng Luo, Ning Fang, Weimin Xu, Sheng Yu, Kai Yan, and Huizhe Xiao. Experimental study on the extraction and distribution of textual domain keywords. Concurrency and Computation: Practice and Experience, 20(16):1917–1932, November 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2007:USL

[LFZ07] Jie Liu, Liang Feng, and Hai Zhuge. Using semantic links to support top-K join queries in peer-to-peer networks. Concurrency and Computation: Practice and Experience, 19(15):

2031–2046, October 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:PCL

[LFZ^{+}17] Cong Li, Yuejian Fang, Xing Zhang, Cancan Jin, Qingni Shen, and Zhonghai Wu. A practical construction for large universe hierarchical attribute-based encryption. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liebrock:2008:MMS

[LG08] L. M. Liebrock and S. P. Goudy. Methodology for modelling SPMD hybrid parallel computation. Concurrency and Computation: Practice and Experience, 20(8):903–940, June 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lindsay:2013:BGM

[LGCJ^{+}13] A. M. Lindsay, M. Galloway-Carson, C. R. Johnson, D. P. Bunde, and V. J. Leung. Backfilling with guarantees made as jobs arrive. Concurrency and Computation: Practice and Experience, 25(4):513–523, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lushbough:2015:LSD

[LGD15] Carol M. Lushbough, Etienne Z. Gnimpieba, and Rion Dooley. Life science data analysis workflow development using the bioextract server leveraging the iPlant collaborative cyberinfrastructure. Concurrency and Computation: Practice and Experience, 27(2):408–419, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lanchares:2013:RBC

[LGdVH13] Juan Lanchares, Oscar Garnica, Francisco Fernández de Vega, and J. Ignacio Hidalgo. A review of bioinspired computer-aided design tools for hardware design. Concurrency and Computation: Practice and Experience, 25(8): 1015–1036, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lujan:2005:EJA

[LGFM05] Mikel Luján, John R. Gurd, T. L. Freeman, and José Miguel. Elimination of Java array bounds checks in the presence of indirection. Concurrency and Computation: Practice and Experience, 17(5–6):489–514, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2016:MBM

[LGG16] Weifeng Liu, Michael Gerndt, and Bin Gong. Model-based MPI-IO tuning with Periscope tuning framework. Concurrency and Computation: Practice and Experience, 28(1):3–20, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:ISC

[LGJ17] Zhichao Li, Siyun Gan, and Ru Jia. The impact of social capitals on service quality of Chinese educational institutions: a multilevel analysis. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lawande:2016:NGM

[LGL16a] Abhijeet G. Lawande, Alan D. George, and Herman Lam. Novo-G#: a multidimensional torus-based reconfigurable cluster for molecular dynamics. Concurrency and Computation: Practice and Experience, 28(8):2374–2393, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2016:OOS

[LGL16b] Bing Lin, Wenzhong Guo, and Xiuyan Lin. Online optimization scheduling for scientific workflows with deadline constraint on hybrid clouds. Concurrency and Computation: Practice and Experience, 28(11):3079–3095, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liao:2017:FAF

[LGL^{+}17] Jianwei Liao, Balazs Gerofi, Guo-Yuan Lien, Takemasa Miyoshi, Seiya Nishizawa, Hirofumi Tomita, Wei-Keng Liao, Alok Choudhary, and Yutaka Ishikawa. A flexible I/O arbitration framework for netCDF-based big data processing workflows on high-end supercomputers. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lam:2015:LLP

[LGLA15] Bryant C. Lam, Alan D. George, Herman Lam, and Vikas Aggarwal. Low-level PGAS computing on many-core processors with TSHMEM. Concurrency and Computation: Practice and Experience, 27(17):5288–5310, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2019:TES

[LGP19] Wai-Kong Lee, Bok-Min Goi, and Raphael C.-W. Phan. Terabit encryption in a second: Performance evaluation of block ciphers in GPU with Kepler, Maxwell, and Pascal architectures. Concurrency and Computation: Practice and Experience, 31(11):e5048:1–e5048:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:AHD

[LGQ^{+}17] Yibin Li, Keke Gai, Meikang Qiu, Wenyun Dai, and Meiqin Liu. Adaptive human detection approach using FPGA-based parallel architecture in reconfigurable hardware. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2012:LCM

[LGQS12] Ting Liu, Xiaohong Guan, Yu Qu, and Yanan Sun. A layered classification for malicious function identification and malware detection. Concurrency and Computation: Practice and Experience, 24(11):1169–1179, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:BCD

[LGY17] Guozhi Li, Songtao Guo, and Yuanyuan Yang. Blocking costdriven multicast scheduling in fat-tree data center networks. Concurrency and Computation: Practice and Experience, 29 (19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lhotak:2005:RTE

[LH05] Ondřej Lhoták and Laurie Hendren. Run-time evaluation of opportunities for object inlining in Java. Concurrency and Computation: Practice and Experience, 17(5–6):515–537, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2014:ARM

[LH14] Cai Li and Jiankun Hu. Attacks via record multiplicity on cancelable biometrics templates. Concurrency and Computation: Practice and Experience, 26(8):1593–1605, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:COM

[LH17] Chenyang Li and Kejing He. CBMR: an optimized MapReduce for item-based collaborative filtering recommendation algorithm with empirical analysis. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2019:SPP

[LHB^{+}19] Xia Lin, Qingjia Huang, Fei Bai, Dun Jin, Wei Lu, and Hongbing Tao. Study on the posts and performance evaluation model of inpatients in public hospitals based on .NET platform DRGs payment mode. Concurrency and Computation: Practice and Experience, 31(10):e4869:1–e4869:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2015:AMA

[LHBW15] Paul T. Lin, Michael A. Heroux, Richard F. Barrett, and Alan B. Williams. Assessing a mini-application as a performance proxy for a finite element method engineering application. Concurrency and Computation: Practice and Experience, 27(17):5374–5389, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liao:2007:OOP

[LHC^{+}07] Chunhua Liao, Oscar Hernandez, Barbara Chapman, Wenguang Chen, and Weimin Zheng. OpenUH: an optimizing, portable OpenMP compiler. Concurrency and Computation: Practice and Experience, 19(18):2317–2332, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2014:IST

[LHC14] Cheng-Yu Lee, Min-Chin Hung, and Rong-Guey Chang. Instruction scheduling and transformation for a VLIW unified reduced instruction set computer/digital signal processor processor with shared register architecture. Concurrency and Computation: Practice and Experience, 26(1):134–151, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2017:VSS

[LHH^{+}17] Yan Liu, Xin Huang, Yizi Huang, Shaofeng Geng, Xin Peng, and Renfa Li. A variable-sized stripe level data layout strategy for HDD/SSD hybrid parallel file systems. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2018:GAE

[LHHJ18] Siguang Li, Zhengwen Huang, Liangxiu Han, and Changjun Jiang. A genetic algorithm enhanced automatic data flow management solution for facilitating data intensive applications in the cloud. Concurrency and Computation: Practice and Experience, 30(16):e4844:1–e4844:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Leal:2010:PBS

[LHL10] Katia Leal, Eduardo Huedo, and Ignacio M. Llorente. Performance-based scheduling strategies for HTC applications in complex federated grids. Concurrency and Computation: Practice and Experience, 22(11):1416–1432, August 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2016:TVD

[LHLH16] Zhenyu Liu, Jingyuan Hu, Yang Li, and Yi Huang. Towards virtual dataspaces for material scientific data cloud. Concurrency and Computation: Practice and Experience, 28(6):

1737–1750, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Litke:2009:FTP

[LHT^{+}09] Antonios Litke, Dimitrios Halkos, Konstantinos Tserpes, Dimosthenis Kyriazis, and Theodora Varvarigou. Fault tolerant and prioritized scheduling in OGSA-based mobile Grids. Concurrency and Computation: Practice and Experience, 21(4): 533–556, March 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2008:DTK

[LHXY08] Xiangfeng Luo, Qingliang Hu, Weimin Xu, and Zhian Yu. Discovery of textual knowledge flow based on the management of knowledge maps. Concurrency and Computation: Practice and Experience, 20(15):1791–1806, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2019:ANI

[LHZS19] Yuying Liu, Yonggang Huang, Jiao Zhang, and Hualei Shen. Anti-noise image source identification. Concurrency and Computation: Practice and Experience, 31(19):e5104:1– e5104:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2019:RCA

[LHZX19] Liangxu Liu, Zhenghua Hu, Chaolan Zhou, and Guohu Xu. Research on the clustering algorithm of the bicycle stations based on OPTICS. Concurrency and Computation: Practice and Experience, 31(10):e4876:1–e4876:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2004:PER

[Li04] Keqin Li. Performance evaluation of a random-walk-based algorithm for embedding dynamically evolving trees in hypercubic networks. Concurrency and Computation: Practice and Experience, 16(13):1327–1351, November 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:EIP

[Li17] Maozhen Li. Editorial: Internet of People. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2018:HPD

[Li18] Maozhen Li. High performance deep learning techniques for big data analytics. Concurrency and Computation: Practice and Experience, 30(23):e5032:1–e5032:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:SCS

[Li19] Delu Li. Study on control strategy and application of a new harmonic detection method in three-phase four-wire APF. Concurrency and Computation: Practice and Experience, 31 (10):e4853:1–e4853:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liao:2016:SSP

[Lia16] Jianwei Liao. Server-side prefetching in distributed file systems. Concurrency and Computation: Practice and Experience, 28(2):294–310, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liao:2010:TVD

[LJHL10] Xiaofei Liao, Hai Jin, Liting Hu, and Haikun Liu. Towards virtualized desktop environment. Concurrency and Computation: Practice and Experience, 22(4):419–440, March 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2017:SWB

[LJL^{+}17] Yang Liu, Weizhe Jing, Youbo Liu, Lin Lv, Man Qi, and Yang Xiang. A sliding window-based dynamic load balancing for heterogeneous Hadoop clusters. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2010:PMA

[LJML10] Lei Liu, Xiaolong Jin, Geyong Min, and Keqiu Li. Performance modelling and analysis of Deficit Round Robin scheduling scheme with self-similar traffic. Concurrency and Computation: Practice and Experience, 22(13):1911–1926, September 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Leung:2016:PSN

[LJPP16] Carson Kai-Sang Leung, Fan Jiang, Adam G. M. Pazdor, and Aaron M. Peddle. Parallel social network mining for interesting ‘following’ patterns. Concurrency and Computation: Practice and Experience, 28(15):3994–4012, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2017:OOA

[LJW^{+}17] Zechao Liu, Zoe L. Jiang, Xuan Wang, Xinyi Huang, S. M. Yiu, and Kunihiko Sadakane. Offline/online attribute-based encryption with verifiable outsourced decryption. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lawlor:2003:SDP

[LK03] Orion S. Lawlor and Laxmikant V. Kalé. Supporting dynamic parallel object arrays. Concurrency and Computation: Practice and Experience, 15(3–5):371–393, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luecke:2003:CPM

[LKJ03] Glenn R. Luecke, Marina Kraeva, and Lili Ju. Comparing the performance of MPICH with Cray’s MPI and with SGI’s MPI. Concurrency and Computation: Practice and Experience, 15(9):779–802, August 10, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:CSV

[LKKL16] Hongzhe Li, Hyuckmin Kwon, Jonghoon Kwon, and Heejo Lee. CLORIFI: software vulnerability discovery using code clone verification. Concurrency and Computation: Practice and Experience, 28(6):1900–1917, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2009:PSF

[LKPM09] Hong Li, Allison Kolpas, Linda Petzold, and Jeff Moehlis. Parallel simulation for a fish schooling model on a generalpurpose graphics processing unit. Concurrency and Computation: Practice and Experience, 21(6):725–737, April 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luecke:2004:PSM

[LKYS04] Glenn R. Luecke, Marina Kraeva, Jing Yuan, and Silvia Spanoyannis. Performance and scalability of MPI on PC clusters. Concurrency and Computation: Practice and Experience, 16(1):79–107, January 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luecke:2001:SPO

[LL01] Glenn R. Luecke and Wei-Hua Lin. Scalability and performance of OpenMP and MPI on a 128-processor SGI Origin 2000. Concurrency and Computation: Practice and Experience, 13(10):905–928, August 25, 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/85007180/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=85007180&PLACEBO=IE.
pdf.

Lee:2005:GPH

[LL05] Lie-Quan Lee and Andrew Lumsdaine. Generic programming for high-performance scientific applications. Concurrency and Computation: Practice and Experience, 17(7–8): 941–965, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2010:STC

[LL10] Victor C. S. Lee and Kai Liu. Scheduling time-critical requests for multiple data objects in on-demand broadcast. Concurrency and Computation: Practice and Experience, 22 (15):2124–2143, October 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2013:EPS

[LL13] Keqiu Li and Kai Lin. Editorials: Preface of special issue: advanced topics on wireless sensor networks. Concurrency and Computation: Practice and Experience, 25(1):55–57, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:ERA

[LL15] Zhiyang Li and Keqiu Li. Editorials: Recent advances in parallel computing and distributed network. Concurrency and Computation: Practice and Experience, 27(16):4137–4139, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:AAO

[LL16a] Shuo Li and James Lin. Accelerating Asian option pricing on many-core architectures. Concurrency and Computation: Practice and Experience, 28(3):848–865, March 10, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2016:PEJ

[LL16b] Jia-Chun Lin and Ming-Chang Lee. Performance evaluation of job schedulers on Hadoop YARN. Concurrency and Computation: Practice and Experience, 28(9):2711–2728, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2016:VDF

[LL16c] Yu-Te Lin and Jenq-Kuen Lee. Vector data flow analysis for SIMD optimizations on OpenCL programs. Concurrency and Computation: Practice and Experience, 28(5):1629–1654, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2018:CSC

[LL18] Chaoliang Li and Shigang Liu. A comparative study of the class imbalance problem in Twitter spam detection. Concurrency and Computation: Practice and Experience, 30(5):??, March 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4281.

Li:2019:SPI

[LL19a] Xiaofei Li and Da Li. Study on the public intention for environmental protection based on social network. Concurrency and Computation: Practice and Experience, 31(10):e4742:1– e4742:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2019:CIQ

[LL19b] Yaqin Liu and XinXing Luo. C2C interaction quality identification with SVM based on Kano model. Concurrency and Computation: Practice and Experience, 31(12):e4689:1– e4689:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2019:IPE

[LL19c] Le Luo and Yi Liu. Improving parallel efficiency for asynchronous graph analytics using Gauss–Seidel-based matrix computation. Concurrency and Computation: Practice and Experience, 31(17):e5267:1–e5267:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2004:UHA

[LLB04] Yuheng Li, Mario Lauria, and Ralf Bundschuh. Using hybrid alignment for iterative sequence database searches. Concurrency and Computation: Practice and Experience, 16(9):841– 853, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:CFR

[LLC^{+}15a] Yan Li, Wenxin Li, Honghui Chen, Deke Guo, Tian Zhang, and Ting Qu. Congestion-free routing strategy in software defined data center networks. Concurrency and Computation: Practice and Experience, 27(18):5735–5748, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2015:SSE

[LLC^{+}15b] Jin Liu, Juan Li, Xiaohui Cui, Xiaoguang Niu, Xiaoping Sun, and Jing Zhou. Social sensing enhanced time estimation for bus service. Concurrency and Computation: Practice and Experience, 27(15):3961–3981, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lakiotakis:2019:INM

[LLD19] Emmanouil Lakiotakis, Christos Liaskos, and Xenofontas Dimitropoulos. Improving networked music performance systems using application-network collaboration. Concurrency and Computation: Practice and Experience, 31(24):e4730:1– e4730:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lobosco:2008:ERT

[LLdA08] Marcelo Lobosco, Orlando Loques, and Claudio L. de Amorim. On the effectiveness of runtime techniques to reduce memory sharing overheads in distributed Java implementations. Concurrency and Computation: Practice and Experience, 20(13): 1509–1538, September 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2008:RSV

[LLF08] Jin Liu, Xiang Li, and Liang Feng. Resource space view tour mechanism. Concurrency and Computation: Practice and Experience, 20(7):863–883, May 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:RMB

[LLG^{+}15] Ruixuan Li, Huaqing Li, Xiwu Gu, Yuhua Li, Wei Ye, and Xiaopu Ma. Role mining based on cardinality constraints. Concurrency and Computation: Practice and Experience, 27 (12):3126–3144, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2009:ARL

[LLH^{+}09] Huan Li, Xin Li, Dawei Hu, Tianyong Hao, Liu Wenyin, and Xiaoping Chen. Adaptation rule learning for case-based reasoning. Concurrency and Computation: Practice and Experience, 21(5):673–689, April 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2015:GUH

[LLH^{+}15] Yu-Shiang Lin, Chun-Yuan Lin, Che-Lun Hung, Yeh-Ching Chung, and Kual-Zheng Lee. GPU–UPGMA: highperformance computing for UPGMA algorithm based on graphics processing units. Concurrency and Computation: Practice and Experience, 27(13):3403–3414, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2017:FSF

[LLH^{+}17] Weifeng Liu, Ang Li, Jonathan D. Hogg, Iain S. Duff, and Brian Vinter. Fast synchronization-free algorithms for parallel sparse triangular solves with multiple right-hand sides. Concurrency and Computation: Practice and Experience, 29 (21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2018:EGB

[LLH^{+}18] Zhanxian Liu, Rongke Liu, Yi Hou, Hao Peng, and Ling Zhao. Efficient GPU-based implementation for decoding non-binary LDPC codes with layered and flooding schedules. Concurrency and Computation: Practice and Experience, 30(16):??, August 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4442.

Li:2019:EPP

[LLH19] Rong Li, Fei Li, and Jiewu Huang. Evaluation of the predictive performance of the principal component two-parameter estimator. Concurrency and Computation: Practice and Experience, 31(12):e4710:1–e4710:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See retraction notice [HML20].

Li:2018:SEDa

[LLJ18a] Siguang Li, Maozhen Li, and Changjun Jiang. Semantic enhanced deep learning for image classification. Concurrency and Computation: Practice and Experience, 30 (16):??, August 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4720.

Li:2018:SEDb

[LLJ18b] Siguang Li, Maozhen Li, and Changjun Jiang. Semantic enhanced deep learning for image classification. Concurrency and Computation: Practice and Experience, 30(23):e4388:1– e4388:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2008:EFS

[LLKC08] Byoung-Hoon Lee, Sung-Hwa Lim, Jai-Hoon Kim, and We-Duke Cho. Efficient and fair scheduling for two-level information broadcasting systems. Concurrency and Computation: Practice and Experience, 20(18):2179–2200, December 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:CRE

[LLL15] Bo Li, Jianxin Li, and Lu Liu. CloudMon: a resource-efficient IaaS cloud monitoring system based on networked intrusion detection system virtual appliances. Concurrency and Computation: Practice and Experience, 27(8):1861–1885, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2016:EBD

[LLL16] Xiangfeng Luo, Yunhuai Liu, and Qing Li. Editorials: Big data-related technologies and applications. Concurrency and Computation: Practice and Experience, 28(15):4036–4037, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2019:GAB

[LLL^{+}19] Hong Liu, Yan Li, Wenhao Li, Dianjie Lu, and Guijuan Zhang. A grouping approach based on non-uniform binary grid partitioning for crowd evacuation simulation. Concurrency and Computation: Practice and Experience, 31(23): e4493:1–e4493:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2014:EES

[LLLJ14] Jingwei Li, Jin Li, Zheli Liu, and Chunfu Jia. Enabling efficient and secure data sharing in cloud computing. Concurrency and Computation: Practice and Experience, 26(5): 1052–1066, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2018:MBT

[LLLS18] Jung-Chun Liu, Fang-Yie Leu, Guan-Liang Lin, and Heru Susanto. An MFCC-based text-independent speaker identification system for access control. Concurrency and Computation: Practice and Experience, 30(2):??, January 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2016:BDA

[LLLyL16] Xia Lin, Yao Li, Yuping Lu, and Nai yong Li. Big data analysis on the coordination optimization of protection schemes and automatic devices for distribution network. Concurrency and Computation: Practice and Experience, 28(15): 4053–4066, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2019:PPP

[LLM19] Sharon X. Lee, Kaleb L. Leemaqz, and Geoffrey J. McLachlan. PPEM: Privacy-preserving EM learning for mixture models. Concurrency and Computation: Practice and Experience, 31(24):e5208:1–e5208:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Laccetti:2018:MAT

[LLMK18] Giuliano Laccetti, Marco Lapegna, Raffaele Montella, and Sokol Kosta. Models, algorithms, and tools for highly heterogeneous computing environments. Concurrency and Computation: Practice and Experience, 30(24):e4950:1–e4950:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Laccetti:2019:AAH

[LLMM19] Giuliano Laccetti, Marco Lapegna, Valeria Mele, and Raffaele Montella. An adaptive algorithm for high-dimensional integrals on heterogeneous CPU–GPU systems. Concurrency and Computation: Practice and Experience, 31(19):e4945:1– e4945:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

LeBeux:2014:OCC

[LLN^{+}14] Sébastien Le Beux, Hui Li, Gabriela Nicolescu, Jelena Trajkovic, and Ian O’Connor. Optical crossbars on chip, a comparative study based on worst-case losses. Concurrency and Computation: Practice and Experience, 26(15):2492–2503, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2014:BMI

[LLQL14] Minggang Luo, Liping Li, Guobing Qian, and Jianqi Lu. A blind modulation identification algorithm for STBC systems using multidimensional ICA. Concurrency and Computation: Practice and Experience, 26(8):1490–1505, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Langlais:2003:PAQ

[LLRS03] M. Langlais, G. Latu, J. Roman, and P. Silan. Performance analysis and qualitative results of an efficient parallel stochastic simulator for a marine host-parasite system. Concurrency and Computation: Practice and Experience, 15(11–12):1133– 1150, September 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2019:PUS

[LLRS19] Jun Liu, Xiaoyong Li, Kaijun Ren, and Junqiang Song. Parallelizing uncertain skyline computation against n-of-N data streaming model. Concurrency and Computation: Practice and Experience, 31(4):e4848:1–e4848:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2015:AMS

[LLSL15] Zhusong Liu, Zhike Li, Sheng Sun, and Yunfa Li. An auditing mechanism for the security of service in grid. Concurrency and Computation: Practice and Experience, 27(8):2041–2053, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lees:2009:APC

[LLT09] Michael Lees, Brian Logan, and Georgios Theodoropoulos. Analysing probabilistically constrained optimism. Concurrency and Computation: Practice and Experience, 21(11): 1467–1482, August 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2014:HAC

[LLT^{+}14] Qing Liu, Jeremy Logan, Yuan Tian, Hasan Abbasi, Norbert Podhorszki, Jong Youl Choi, Scott Klasky, Roselyne Tchoua, Jay Lofstead, Ron Oldfield, Manish Parashar, Nagiza Samatova, Karsten Schwan, Arie Shoshani, Matthew Wolf, Kesheng Wu, and Weikuan Yu. Hello ADIOS: the challenges and lessons of developing leadership class I/O frameworks. Concurrency and Computation: Practice and Experience, 26(7): 1453–1473, May 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:MRW

[LLT^{+}19a] Xiaoling Li, Xiaoyong Li, Yusong Tan, Hao Zhu, and Shuang Tan. Multi-resource workload mapping with minimum cost in cloud environment. Concurrency and Computation: Practice and Experience, 31(15):e5167:1–e5167:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:SLV

[LLT19b] Xiaoyi Li, Jiaqi Li, and Xiaolin Tian. A suspected lunar volcano identification algorithm based on convolutional neural network by the rapid location of light shadow impacts. Concurrency and Computation: Practice and Experience, 31(10): e4727:1–e4727:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2009:NSM

[LLWS09] Qing Li, Na Li, Liping Wang, and Xiaoping Sun. A new semantic model with applications in a multimedia database system. Concurrency and Computation: Practice and Experience, 21(5):691–704, April 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:TSR

[LLX^{+}15a] Jia Li, Xin Luo, Yunni Xia, Yakai Han, and Qingsheng Zhu. A time series and reduction-based model for modeling and QoS prediction of service compositions. Concurrency and Computation: Practice and Experience, 27(1):146–163, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2015:OHE

[LLX15b] Yang Liu, Xiangfeng Luo, and Junyu Xuan. Online hot event discovery based on Association Link Network. Concurrency and Computation: Practice and Experience, 27(15): 4001–4014, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:IPM

[LLY^{+}19a] Bin Li, Jian Li, Shilin Yan, Linfeng Ju, and Xu He. Impacts of porous material fluid bulk properties on noise attenuation performance of cylinder shell structure based on finite element model. Concurrency and Computation: Practice and Experience, 31(10):e4714:1–e4714:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2019:ACW

[LLY^{+}19b] Li-Hui Lin, Cheng-Hsuan Li, Sheng Yang, Shao-Zi Li, and Yi Wei. Automated classification of Wuyi rock tealeaves based on support vector machine. Concurrency and Computation: Practice and Experience, 31(23):e4519:1–e4519:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2019:RMP

[LLY19c] Tingting Luo, Guangyao Li, and Naijiang Yu. Research on manufacturing productivity based on improved genetic algorithms under Internet information technology. Concurrency and Computation: Practice and Experience, 31(10):e4859:1– e4859:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2009:LGG

[LLYL09] Chia-Han Lu, Yung-Chia Lin, Yi-Ping You, and Jenq-Kuen Lee. LC-GRFA: global register file assignment with local consciousness for VLIW DSP processors with non-uniform register files. Concurrency and Computation: Practice and Experience, 21(1):101–114, January 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2017:SSA

[LLZ^{+}17a] Weidong Liu, Xiangfeng Luo, Jun Zhang, Ruirong Xue, and Richard Yi Da Xu. Semantic summary automatic generation in news event. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2017:WIC

[LLZ^{+}17b] Huimin Lu, Bin Li, Junwu Zhu, Yujie Li, Yun Li, Xing Xu, Li He, Xin Li, Jianru Li, and Seiichi Serikawa. Wound intensity correction and segmentation with convolutional neural networks. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2007:KEA

[LM07] L. Li and A. D. Malony. Knowledge engineering for automatic parallel performance diagnosis. Concurrency and Computation: Practice and Experience, 19(11):1497–1515, August 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2008:PLC

[LM08] Yonghwan Lee and Dugki Min. A peak load control-based orchestration system for stable execution of hybrid services. Concurrency and Computation: Practice and Experience, 20 (4):361–376, March 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2019:CMR

[LMCL19] Dazhen Lin, Ben Ma, Donglin Cao, and Shaozi Li. Chinese microblog rumor detection based on deep sequence context. Concurrency and Computation: Practice and Experience, 31(23):e4508:1–e4508:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lopez:2019:TMT

[LMDP19] Martin Andreoni Lopez, Diogo M. F. Mattos, Otto Carlos M. B. Duarte, and Guy Pujolle. Toward a monitoring and threat detection system based on stream processing as a virtual network function for big data. Concurrency and Computation: Practice and Experience, 31(20):e5344:1–e5344:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lathrop:2019:BPM

[LME^{+}19] Scott Lathrop, Celso Mendes, Jeremy Enos, Brett Bode, Gregory Bauer, Roberto Sisneros, and William Kramer. Best practices for management and operation of large HPC installations. Concurrency and Computation: Practice and Experience, 31(16):e5069:1–e5069:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2019:AMR

[LMGZ19] Rong Lu, Fengshan Ma, Jie Guo, and Haijun Zhao. Analysis and monitoring of roadway deformation mechanisms in nickel mine, China. Concurrency and Computation: Practice and Experience, 31(10):e4832:1–e4832:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Landau:2014:IIC

[LMH^{+}14] Rubin Landau, Greg Mulder, Raquell Holmes, Sofya Borinskaya, NamHwa Kang, and Cristian Bordeianu. IN-STANCES: incorporating computational scientific thinking advances into education and science courses. Concurrency and Computation: Practice and Experience, 26(13):2316– 2328, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2013:SIPb

[LMKT13] Fagen Li, Fahad T. Bin Muhaya, Muhammad Khurram Khan, and Tsuyoshi Takagi. Special issue papers: Lattice-based signcryption. Concurrency and Computation: Practice and Experience, 25(14):2112–2122, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2018:EDE

[LML^{+}18] Jingxin Liu, Hongying Meng, Maozhen Li, Fan Zhang, Rui Qin, and Asoke K. Nandi. Emotion detection from EEG recordings based on supervised and unsupervised dimension reduction. Concurrency and Computation: Practice and Experience, 30(23):e4446:1–e4446:??, December 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

LoRe:2015:SRN

[LMO15] Giuseppe Lo Re, Fabrizio Milazzo, and Marco Ortolani. Secure random number generation in wireless sensor networks. Concurrency and Computation: Practice and Experience, 27 (15):3842–3862, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lucchese:2010:MTP

[LMOT10] C. Lucchese, C. Mastroianni, S. Orlando, and D. Talia. Mining@home: toward a public-resource computing framework for distributed data mining. Concurrency and Computation: Practice and Experience, 22(5):658–682, April 10, 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Limet:2018:HCS

[LMS18] Sébastien Limet, Alessio Merlo, and Luca Spalazzi. HPC & Co strike back: Where are distributed paradigms heading toward? Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4431.

Liu:2018:MBP

[LMX^{+}18] Yang Liu, Chenxiao Ma, Lixiong Xu, Xiaodong Shen, Maozhen Li, and Pengcheng Li. MapReduce-based parallel GEP algorithm for efficient function mining in big data applications. Concurrency and Computation: Practice and Experience, 30(16):e4379:1–e4379:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:ENA

[LNBL17] Gang Li, Wenjia Niu, Lynn Batten, and Jiqiang Liu. Editorial: New advances in securing cyberspace and curbing crowdturfing. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2011:PST

[LNCY11] Xiao Liu, Zhiwei Ni, Jinjun Chen, and Yun Yang. A probabilistic strategy for temporal constraint management in scientific workflow systems. Concurrency and Computation: Practice and Experience, 23(16):1893–1919, November 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:ESC

[LNG^{+}16] Gang Li, Wenjia Niu, Li Guo, Lynn Batten, Yinlong Liu, and Guoyong Cai. Editorials: Securing cyberspace. Concurrency and Computation: Practice and Experience, 28(6):1870–1871, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liogkas:2008:ERB

[LNKZ08] Nikitas Liogkas, Robert Nelson, Eddie Kohler, and Lixia Zhang. Exploring the robustness of BitTorrent peer-to-peer content distribution systems. Concurrency and Computation: Practice and Experience, 20(2):179–189, February 2008. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Langer:2010:NSQ

[LOKW^{+}10] S. Langer, L. M. Olsen-Kettle, D. K. Weatherley, L. Gross, and H.-B. Mühlhaus. Numerical studies of quasi-static tectonic loading and dynamic rupture of bi-material interfaces. Concurrency and Computation: Practice and Experience, 22 (12):1684–1702, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lettich:2017:MGP

[LOSJ17] Francesco Lettich, Salvatore Orlando, Claudio Silvestri, and Christian S. Jensen. Manycore GPU processing of repeated range queries over streams of moving objects observations. Concurrency and Computation: Practice and Experience, 29 (4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lowe:2017:TL

[Low17] Gavin Lowe. Testing for linearizability. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ludascher:2008:CMM

[LPA^{+}08] Bertram Ludäscher, Norbert Podhorszki, Ilkay Altintas, Shawn Bowers, and Timothy McPhillips. From computation models to models of provenance: the RWS approach. Concurrency and Computation: Practice and Experience, 20(5): 507–518, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lorenzo:2014:HCB

[LPC^{+}14] Oscar G. Lorenzo, Tomás F. Pena, José C. Cabaleiro, Juan C. Pichel, Juan A. Lorenzo, and Francisco F. Rivera. A hardware counter-based toolkit for the analysis of memory accesses in SMPs. Concurrency and Computation: Practice and Experience, 26(6):1328–1341, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2014:HMT

[LPG^{+}14] Yuan Luo, Beth Plale, Zhenhua Guo, Wilfred W. Li, Judy Qiu, and Yiming Sun. Hierarchical MapReduce: towards simplified cross-domain data processing. Concurrency and Computation: Practice and Experience, 26(4):878–893, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Leist:2009:EGP

[LPH09] A. Leist, D. P. Playne, and K. A. Hawick. Exploiting graphical processing units for data-parallel scientific applications. Concurrency and Computation: Practice and Experience, 21 (18):2400–2437, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2009:AWP

[LPS^{+}09] Kevin Lee, Norman W. Paton, Rizos Sakellariou, Ewa Deelman, Alvaro A. A. Fernandes, and Gaurang Mehta. Adaptive workflow processing and execution in Pegasus. Concurrency and Computation: Practice and Experience, 21(16): 1965–1981, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2011:UFA

[LPSF11] Kevin Lee, Norman W. Paton, Rizos Sakellariou, and Alvaro A. A. Fernandes. Utility functions for adaptively executing concurrent workflows. Concurrency and Computation: Practice and Experience, 23(6):646–666, April 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2015:CGE

[LPW15] Yan Liu, Anand Padmanabhan, and Shaowen Wang. Cyber-GIS Gateway for enabling data-rich geospatial research and education. Concurrency and Computation: Practice and Experience, 27(2):395–407, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lian:2008:RIM

[LPY^{+}08] Qiao Lian, Yu Peng, Mao Yang, Zheng Zhang, Yafei Dai, and Xiaoming Li. Robust incentives via multi-level Tit-for-Tat. Concurrency and Computation: Practice and Experience, 20 (2):167–178, February 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2009:SPA

[LQL^{+}09] Zhaobin Liu, Wenyu Qu, Haitao Li, Min Ruan, and Wanlei Zhou. I/O scheduling and performance analysis on multicore platforms. Concurrency and Computation: Practice and Experience, 21(10):1405–1417, July 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2015:RPO

[LQL^{+}15] Zhaobin Liu, Wenyu Qu, Weijiang Liu, Zhiyang Li, and Yujie Xu. Resource preprocessing and optimal task scheduling in cloud computing environments. Concurrency and Computation: Practice and Experience, 27(13):3461–3482, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Leff:2005:EJC

[LR05] Avraham Leff and James T. Rayfield. Enterprise JavaBeans caching in clustered environments. Concurrency and Computation: Practice and Experience, 17(7–8):1027–1051, June/ July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:ESC

[LRLY17] Xiaoyong Li, Kaijun Ren, Xiaoling Li, and Jie Yu. Efficient skyline computation over distributed interval data. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2015:RED

[LRS15] Shilin Lu, Rajiv Ranjan, and Peter Strazdins. Reporting an experience on design and implementation of e-Health systems on Azure cloud. Concurrency and Computation: Practice and Experience, 27(10):2602–2615, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2005:PRS

[LS05] Tong Liu and Bertil Schmidt. Parallel RNA secondary structure prediction using stochastic context-free grammars. Concurrency and Computation: Practice and Experience, 17(14): 1669–1685, December 10, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Loidl:2014:ESM

[LS14] Hans-Wolfgang Loidl and Jeremy Singer. Editorials: SICSA multicore challenge editorial preface. Concurrency and Computation: Practice and Experience, 26(4):929–934, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2015:GFG

[LS15] Yongchao Liu and Bertil Schmidt. GSWABE: faster GPUaccelerated sequence alignment with optimal alignment retrieval for short DNA sequences. Concurrency and Computation: Practice and Experience, 27(4):958–972, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:BSW

[LS19a] Feng Li and Fengguang Song. Building a scientific workflow framework to enable real-time machine learning and visualization. Concurrency and Computation: Practice and Experience, 31(16):e4703:1–e4703:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liang:2019:ICM

[LS19b] Hejun Liang and Lily Sun. Improve cloud manufacturing supply chain note-enterprises optimize combination of the Cuckoo search. Concurrency and Computation: Practice and Experience, 31(10):e4764:1–e4764:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:PPP

[LSC^{+}19] Jie Li, Jinshu Su, Rongmao Chen, Xiaofeng Wang, and Shuhui Chen. Practical privacy-preserving deep packet inspection outsourcing. Concurrency and Computation: Practice and Experience, 31(22):e4435:1–e4435:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lakshminarasimhan:2013:IES

[LSE^{+}13] Sriram Lakshminarasimhan, Neil Shah, Stephane Ethier, Seung-Hoe Ku, C. S. Chang, Scott Klasky, Rob Latham, Rob Ross, and Nagiza F. Samatova. ISABELA for effective in situ compression of scientific data. Concurrency and Computation: Practice and Experience, 25(4):524–540, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:IPG

[LSH^{+}16] Yuanzhe Li, Loren Schwiebert, Eyad Hailat, Jason Mick, and Jeffrey Potoff. Improving performance of GPU code using novel features of the NVIDIA Kepler architecture. Concurrency and Computation: Practice and Experience, 28(13): 3586–3605, September 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2016:RAA

[LSJ16] Yuan Liu, Xuanhua Shi, and Hai Jin. Runtime-aware adaptive scheduling in stream processing. Concurrency and Computation: Practice and Experience, 28(14):3830–3843, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luecke:2004:PSS

[LSK04] Glenn R. Luecke, Silvia Spanoyannis, and Marina Kraeva. The performance and scalability of SHMEM and MPI-2 onesided routines on a SGI Origin 2000 and a Cray T3E-600. Concurrency and Computation: Practice and Experience, 16 (10):1037–1060, August 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lv:2017:TBA

[LSL^{+}17] Huiming Lv, Zhiyuan Shao, Lang Li, Xuanhua Shi, and Hai Jin. A task-based approach for finding SCCs in real-world graphs on external memory. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lucas-Simarro:2015:COV

[LSMVML15] Jose Luis Lucas-Simarro, Rafael Moreno-Vozmediano, Ruben S. Montero, and Ignacio M. Llorente. Cost optimization of virtual infrastructures in dynamic multi-cloud scenarios. Concurrency and Computation: Practice and Experience, 27(9): 2260–2277, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2015:HGP

[LSP15] Fengshun Lu, Junqiang Song, and Yufei Pang. HLog_{n} GP: a parallel computation model for GPU clusters. Concurrency and Computation: Practice and Experience, 27(17): 4880–4896, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Laure:2005:PED

[LSS05] Erwin Laure, Heinz Stockinger, and Kurt Stockinger. Performance engineering in data Grids. Concurrency and Computation: Practice and Experience, 17(2–4):171–191, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Limet:2015:EHP

[LSS15] Sébastien Limet, Waleed W. Smari, and Luca Spalazzi. Editorials: High-performance computing: to boldly go where no human has gone before. Concurrency and Computation: Practice and Experience, 27(13):3145–3165, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shao:2017:RIH

[lSsCY17] Guo lin Shao, Xing shu Chen, and Lu Yang. Research and implementation of a high performance parallel computing digital down converter on graphics processing unit. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Long:2007:MVC

[LSW07] Brad Long, Paul Strooper, and Luke Wildman. A method for verifying concurrent Java components based on an analysis of concurrency failures. Concurrency and Computation: Practice and Experience, 19(3):281–294, March 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lai:2017:PCL

[LSXL17] Siyan Lai, Bo Shao, Ying Xu, and Xiaola Lin. Parallel computations of local PageRank problem based on Graphics Processing Unit. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2012:VCR

[LSY^{+}12] Tian Liu, Shihai Shao, Daixiong Ye, Youxi Tang, and Juan Zhou. Visual Cognitive Radio. Concurrency and Computation: Practice and Experience, 24(11):1252–1260, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:RIS

[LSZ19] Yuxue Li, Lijun Song, and Yucheng Zeng. Research on information security and privacy protection model based on consumer behavior in big data environment. Concurrency and Computation: Practice and Experience, 31(10):e4881:1– e4881:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luckow:2017:HTP

[LTK17] Kasper Søe Luckow, Bent Thomsen, and Stephan Erbs Korsholm. HVM_{T P}: a time predictable and portable Java Virtual Machine for hard real-time embedded systems. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2011:MHS

[LTKF11] Wantao Liu, Brian Tieman, Rajkumar Kettimuthu, and Ian Foster. Moving huge scientific datasets over the Internet. Concurrency and Computation: Practice and Experience, 23 (18):2404–2420, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2017:PQD

[LTL^{+}17] Jen-Gaw Lee, Whey-Fone Tsai, Lung-Cheng Lee, Ching-Yao Lin, Hsi-Ching Lin, and Ben-Jei Tsuang. In-place query driven big data platform: Applications to post processing of environmental monitoring. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2014:SSB

[LTM^{+}14] Xuan Liu, Simon J. E. Taylor, Navonil Mustafee, Jun Wang, Qian Gao, and David Gilbert. Speeding up systems biology simulations of biochemical pathways using condor. Concurrency and Computation: Practice and Experience, 26(17): 2727–2742, December 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:RPS

[LTZ^{+}19] Jingmei Li, Qiao Tian, Fangyuan Zheng, Weifei Wu, and Jiaxiang Wang. Research on parallel solution of GRAPES Helmholtz equation. Concurrency and Computation: Practice and Experience, 31(12):e4479:1–e4479:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lian:2012:SIM

[LV12] Shiguo Lian and Athanasios Vasilakos. Special issue on Multimedia Computing and Management in Cloud Environment. Concurrency and Computation: Practice and Experience, 24 (17):2081–2082, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lagoze:2012:WBR

[LVN^{+}12] Carl Lagoze, Herbert Van de Sompel, Michael Nelson, Simeon Warner, Robert Sanderson, and Pete Johnston. A Web-based resource model for Scholarship 2.0: object reuse & exchange.

Concurrency and Computation: Practice and Experience, 24 (18):2221–2240, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lang:2005:LSD

[LW05] Stefan Lang and Gabriel Wittum. Large-scale density-driven flow simulations using parallel unstructured Grid adaptation and local multigrid methods. Concurrency and Computation: Practice and Experience, 17(11):1415–1440, September 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2013:SIP

[LW13] Dongxi Liu and Shenlu Wang. Special issue papers: Nonlinear order preserving index for encrypted database query in service cloud environments. Concurrency and Computation: Practice and Experience, 25(13):1967–1984, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2013:SLB

[LWB13] Shiyao Lin, Andy Wellings, and Alan Burns. Supporting lock-based multiprocessor resource sharing protocols in realtime programming languages. Concurrency and Computation: Practice and Experience, 25(16):2227–2251, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2012:CCL

[LWC12] Xuefeng Li, Ying Wang, and Xi Chen. Cold chain logistics system based on cloud computing. Concurrency and Computation: Practice and Experience, 24(17):2138–2150, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luckcuck:2017:SCJ

[LWC17] Matt Luckcuck, Andy Wellings, and Ana Cavalcanti. Safety-Critical Java: level 2 in practice. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:BEB

[LWF^{+}15] Xiaoyong Li, Yijie Wang, Yongquan Fu, Xiaoling Li, and Weidong Sun. BLOR: an efficient bandwidth and latency sensitive overlay routing approach for flash data dissemination. Concurrency and Computation: Practice and Experience, 27 (14):3614–3632, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2014:MDS

[LWFL14] Hongliang Li, Xiaohui Wei, Qingwu Fu, and Yuan Luo. MapReduce delay scheduling with deadline constraint. Concurrency and Computation: Practice and Experience, 26(3): 766–778, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:PAB

[LWG^{+}15] Jing Li, Mingying Wu, Jianhua Ge, Chensi Zhang, Florin Pop, and Yingguan Wang. Performance analysis of bidirectional cloud networks with imperfect channel state information. Concurrency and Computation: Practice and Experience, 27(8):2092–2106, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2015:IES

[LWK15] Lei Liu, Hua Wang, and Guohong Kong. An improved EDA for solving Steiner tree problem. Concurrency and Computation: Practice and Experience, 27(13):3483–3496, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2006:SIS

[LWL^{+}06] Minglu Li, Min-You Wu, Ying Li, Jian Cao, Linpeng Huang, Qianni Deng, Xinhua Lin, Changjun Jiang, Weiqin Tong, Yadong Gui, Aoying Zhou, Xinhong Wu, and Shui Jiang. ShanghaiGrid: an Information Service Grid. Concurrency and Computation: Practice and Experience, 18(1):111–135, January 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:EBD

[LWL15] Keqiu Li, Hongyi Wu, and Zhiyang Li. Editorials: Big data cloud and the frontier of computer science and technology.

Concurrency and Computation: Practice and Experience, 27 (18):5719–5721, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:EAH

[LWL17] Kenli Li, Lipo Wang, and Yong Liu. Editorial: Applications in heterogeneous parallel and distributed environment. Concurrency and Computation: Practice and Experience, 29(20): ??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lv:2019:ERV

[LWL^{+}19] Zhipeng Lv, Ming Wu, Lan Liu, Haitao Liu, Zhenhao Song, and Lijing Sun. Energy router with virtual inertia control schemes and concurrency operation. Concurrency and Computation: Practice and Experience, 31(9):e4715:1–e4715:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2011:DOM

[LWLZ11] Jianxun Liu, Yiping Wen, Ting Li, and Xuyun Zhang. A data-operation model based on partial vector space for batch processing in workflow. Concurrency and Computation: Practice and Experience, 23(16):1936–1950, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2019:SCM

[LWP19] Xiao-Qin Liu, Qiu-Lin Wu, and Wen-Tsao Pan. Sentiment classification of micro-blog comments based on Randomforest algorithm. Concurrency and Computation: Practice and Experience, 31(10):e4746:1–e4746:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:CVT

[LWQS19] Weiguang Li, Wang Wei, Han Qiang, and Mingquan Shi. Collaborating visual tracker based on particle filter and correlation filter. Concurrency and Computation: Practice and Experience, 31(12):e4665:1–e4665:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2019:PSD

[LWS19] Zigang Lu, Zhinong Wei, and Yonghui Sun. Power system dynamic state estimation considering correlation of measurement error from PMU and SCADA. Concurrency and Computation: Practice and Experience, 31(10):e4726:1–e4726:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:RCS

[LWSZ19] Zejun Li, Jingjing Wang, Zhongheng Shi, and Yan Zou. Relieve the congestion by shuttle bus in rush hours using aggregation clustering algorithm on group travel pattern. Concurrency and Computation: Practice and Experience, 31(9): e4847:1–e4847:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:BED

[LWT^{+}16] Jianxin Li, Jianfeng Wen, Zhenying Tai, Richong Zhang, and Weiren Yu. Bursty event detection from microblog: a distributed and incremental approach. Concurrency and Computation: Practice and Experience, 28(11):3115–3130, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2006:AFL

[LWW06] Wilfred W. K. Lin, Allan K. Y. Wong, and Richard S. L. Wu. Applying fuzzy logic and genetic algorithms to enhance the efficacy of the PID controller in buffer overflow elimination for better channel response timeliness over the Internet. Concurrency and Computation: Practice and Experience, 18 (7):725–747, June 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:VCM

[LWW^{+}19] Jing Li, Licheng Wang, Lihua Wang, Xianmin Wang, Zhengan Huang, and Jin Li. Verifiable Chebyshev maps-based chaotic encryption schemes with outsourcing computations in the cloud/fog scenarios. Concurrency and Computation: Practice and Experience, 31(22):e4523:1–e4523:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2015:PLP

[LWY15] Chi Lin, Guowei Wu, and Chang Wu Yu. Protecting location privacy and query privacy: a combined clustering approach. Concurrency and Computation: Practice and Experience, 27 (12):3021–3043, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:OTD

[LWY^{+}16] Weimin Li, Bin Wang, Jiaqi Yu, Chong Zhu, Sinuo Xiao, and Jinfang Sheng. The optimization of Transparent-Desktop service mechanism based on SPICE. Concurrency and Computation: Practice and Experience, 28(18):4543–4556, December 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2017:DAM

[LWY^{+}17] Xin Liu, Feng Wang, Yang Yang, Jiuyun Xu, Pingjun Zou, and Yiwen Wang. Defense against malicious URL spreading in micro-blog network with hub nodes. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:DDA

[LWYM16] Xinghua Li, Ermeng Wang, Weidong Yang, and Jianfeng Ma. DALP: a demand-aware location privacy protection scheme in continuous location-based services. Concurrency and Computation: Practice and Experience, 28(4):1219–1236, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:CED

[LWYZ19] Xiaole Li, Hua Wang, Shanwen Yi, and Linbo Zhai. Costefficient disaster backup for multiple data centers using capacity-constrained multicast. Concurrency and Computation: Practice and Experience, 31(17):e5266:1–e5266:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:NMB

[LWZ^{+}17] Dengao Li, Gang Wu, Jumin Zhao, Wenhui Niu, and Shuai Li. A novel method based on MapReduce to extract auxiliary information for GNSS receivers. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:IPA

[LWZ^{+}19] Dengao Li, Zheng Wei, Jumin Zhao, Junbing Cheng, and Zhiying Ma. Integrated positioning algorithm based on GPS/ WLAN. Concurrency and Computation: Practice and Experience, 31(17):e5277:1–e5277:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2008:CSG

[LX08] Zhongwen Li and Yang Xiang. Checkpointing schemes for Grid workflow systems. Concurrency and Computation: Practice and Experience, 20(15):1773–1790, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ludwig:2019:OMT

[LXK^{+}19] Uillian L. Ludwig, Miguel G. Xavier, Dionatrã F. Kirchoff, Ian B. Cezar, and César A. F. De Rose. Optimizing multi-tier application performance with interference and affinity-aware placement algorithms. Concurrency and Computation: Practice and Experience, 31(18):e5098:1–e5098:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2009:GSK

[LXL^{+}09] Xiangfeng Luo, Zheng Xu, Qing Li, Qingliang Hu, Jie Yu, and Xinhuai Tang. Generation of similarity knowledge flow for intelligent browsing based on semantic link networks. Concurrency and Computation: Practice and Experience, 21(16): 2018–2032, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2012:CCI

[LXP^{+}12] Jizi Li, Naixue Xiong, Jong Hyuk Park, Chunling Liu, Shihua Ma, and Linfu Sun. Chain-to-chain inventory transshipment model and robust switch control in CSC networks. Concurrency and Computation: Practice and Experience, 24(4):423– 443, March 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2018:ATJ

[LXP18] Bing Li, Xueli Xiao, and Yi Pan. Automatic translation from Java to Spark. Concurrency and Computation: Practice and Experience, 30(20):e4459:1–e4459:??, October 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2013:SIPa

[LXRJ13] Linfeng Li, Wei Xue, Rajiv Ranjan, and Zhiyan Jin. Special issue papers: A scalable Helmholtz solver in GRAPES over large-scale multicore cluster. Concurrency and Computation: Practice and Experience, 25(12):1722–1737, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:POL

[LXW^{+}16] Dali Li, Chuanfu Xu, Yongxian Wang, Zhifang Song, Min Xiong, Xiang Gao, and Xiaogang Deng. Parallelizing and optimizing large-scale 3D multi-phase flow simulations on the Tianhe-2 supercomputer. Concurrency and Computation: Practice and Experience, 28(5):1678–1692, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2017:PEA

[LXW17] Xiangwen Liu, Qingqing Xie, and Liangmin Wang. Personalized extended (α, k)-anonymity model for privacy-preserving data publishing. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:NCF

[LXYC17] Yun Li, Jie Xu, Yun-Hao Yuan, and Ling Chen. A new closed frequent itemset mining algorithm based on GPU and improved vertical structure. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ltaief:2014:DDE

[LY14] Hatem Ltaief and Rio Yokota. Data-driven execution of fast multipole methods. Concurrency and Computation: Practice and Experience, 26(11):1935–1946, August 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2016:EAS

[LYC16] Duanyang Liu, Xi Yang, and Zhen Cheng. An energy-aware scheduling algorithm for divisible loads in a bus network. Concurrency and Computation: Practice and Experience, 28 (5):1612–1628, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:TBS

[LYF^{+}17] Qingyu Li, Panlong Yang, Xiaochen Fan, Shaojie Tang, Chaocan Xiang, Deke Guo, and Fan Li. Taming the big to small: efficient selfish task allocation in mobile crowdsourcing systems. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2007:PCS

[LYL07] Yung-Chia Lin, Yi-Ping You, and Jenq-Kuen Lee. PALF: compiler supports for irregular register files in clustered VLIW DSP processors. Concurrency and Computation: Practice and Experience, 19(18):2391–2406, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:TIS

[LYL^{+}19] Zuoyong Li, Zhaochai Yu, Weixia Liu, Yong Xu, Daoqiang Zhang, and Yong Cheng. Tongue image segmentation via color decomposition and thresholding. Concurrency and Computation: Practice and Experience, 31(23):e4662:1– e4662:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lyon:2002:SMI

[Lyo02] Douglas Lyon. Simulating multiple inheritance in Java. Concurrency and Computation: Practice and Experience, 14(12): 987–1008, October 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/98516164/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=98516164{\&}PLACEBO=IE.pdf.

Lee:2018:RPT

[LYS18] Kyungroul Lee, Kangbin Yim, and Jung Taek Seo. Ransomware prevention technique using key backup. Concurrency and Computation: Practice and Experience, 30 (3):??, February 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4337.

Liu:2017:DCC

[LZBF17] Li Liu, Miao Zhang, Rajkumar Buyya, and Qi Fan. Deadlineconstrained coevolutionary genetic algorithm for scientific workflow scheduling in cloud computing. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luecke:2002:DDM

[LZC^{+}02] Glenn R. Luecke, Yan Zou, James Coyle, Jim Hoekstra, and Marina Kraeva. Deadlock detection in MPI programs. Concurrency and Computation: Practice and Experience, 14(11): 911–932, August 25, 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/97519209/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=97519209{\&}PLACEBO=IE.pdf.

Liu:2008:WWC

[LZC08] Jianxun Liu, Chunjie Zhou, and Jinjun Chen. WdCM: a workday calendar model for workflows in service grid environments. Concurrency and Computation: Practice and Experience, 20(4):377–392, March 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2009:ITM

[LZC09] Jianxun Liu, Chunjie Zhou, and Jian Cao. An integrated time management model for distributed workflow management systems in Grid environments. Concurrency and Computation: Practice and Experience, 21(16):2084–2098, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2014:PKE

[LZC14] Shengli Liu, Fangguo Zhang, and Kefei Chen. Public-key encryption scheme with selective opening chosen-ciphertext security based on the Decisional Diffie–Hellman assumption. Concurrency and Computation: Practice and Experience, 26 (8):1506–1519, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:SDM

[LZG^{+}19] Kehong Li, Yadong Zhang, Jin Guo, Xiaocheng Ge, and Yuebin Su. System dynamics model for high-speed railway operation safety supervision system based on evolutionary game theory. Concurrency and Computation: Practice and Experience, 31(10):e4743:1–e4743:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2015:FPP

[LZH^{+}15] Shidong Li, Huihua Zhou, Jia Hu, Qing Ai, and Chao Cai. A fast path planning approach for unmanned aerial vehicles. Concurrency and Computation: Practice and Experience, 27 (13):3446–3460, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2018:PHT

[LZJ^{+}18] Maozhen Li, Hanyuan Zhang, Yong Jin, Zhaoba Wang, and Guodong Guo. Parallelizing Hartley transform with Hadoop for fast detection of glass defects. Concurrency and Computation: Practice and Experience, 30(16):e4499:1–e4499:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:TML

[LZL17a] Bo Li, Simin Zhang, and Ke Li. Towards a multi-layers anomaly detection framework for analyzing network traffic. Concurrency and Computation: Practice and Experience, 29 (14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:EEF

[LZL^{+}17b] Guo Li, Dafang Zhang, Yanbiao Li, Jintao Zheng, and Keqin Li. Energy-efficient fuzzy control model for GPU-accelerated packet classification. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2012:ESS

[LZT12] Fagen Li, Mingwu Zhang, and Tsuyoshi Takagi. Efficient signcryption in the standard model. Concurrency and Computation: Practice and Experience, 24(17):1977–1989, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2013:MMC

[LZW13] Fan Li, Danyang Zhang, and Minghui Wang. Multiuser multimedia communication over orthogonal frequency-division multiple access downlink systems. Concurrency and Computation: Practice and Experience, 25(9):1081–1090, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:CKV

[LZW^{+}16] Tonglin Li, Xiaobing Zhou, Ke Wang, Dongfang Zhao, Iman Sadooghi, Zhao Zhang, and Ioan Raicu. A convergence of keyvalue storage systems from clouds to supercomputers. Concurrency and Computation: Practice and Experience, 28(1): 44–69, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:CAO

[LZW^{+}17a] Wenxiang Li, Chunsheng Zhu, Xia Wei, Joel J. P. C. Rodrigues, and Kun Wang. Characteristics analysis and optimization design of entities collaboration for cloud manufacturing. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2017:GGB

[LZW17b] Yuxiang Li, Yinliang Zhao, and Qiangsheng Wu. GbA: a graph-based thread partition approach in speculative multithreading. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lawrence:2015:SGT

[LZWD^{+}15] Katherine A. Lawrence, Michael Zentner, Nancy Wilkins-Diehr, Julie A. Wernert, Marlon Pierce, Suresh Marru, and Scott Michael. Science gateways today and tomorrow: positive perspectives of nearly 5000 members of the research community. Concurrency and Computation: Practice and Experience, 27(16):4252–4268, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2016:BMA

[LZY^{+}16] Fuxiang Li, Fucai Zhou, Heqing Yuan, Zifeng Xu, and Qiang Wang. Bilinear-map accumulator-based verifiable intersection operations on encrypted data in cloud. Concurrency and Computation: Practice and Experience, 28(11):3238–3253, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2015:NSA

[LZZ^{+}15] Jie Lu, Zheng Zheng, Guangquan Zhang, Qing He, and Zhongzhi Shi. A new solution algorithm for solving rule-sets based bilevel decision problems. Concurrency and Computation: Practice and Experience, 27(4):830–854, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Leng:2017:BBR

[LZZ^{+}17] Yonglin Leng, Chen Zhikui, Fangming Zhong, Xiongjiu Li, Yueming Hu, and Chao Yang. BRGP: a balanced RDF graph partitioning algorithm for cloud storage. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Monnerat:2015:ESH

[MA15] Luiz Monnerat and Claudio L. Amorim. An effective singlehop distributed hash table with high lookup performance and low traffic overhead. Concurrency and Computation: Practice and Experience, 27(7):1767–1788, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maamar:2013:EMC

[MABP13] Haifa Raja Maamar, Graciela Roman Alonso, Azzedine Boukerche, and Emil Petriu. Energy management control for supplying partner selection protocol in mobile peer-to-peer three-dimensional streaming. Concurrency and Computation: Practice and Experience, 25(5):752–769, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martins:2010:DMU

[MAdS^{+}10] Felipe S. Martins, Rossana M. Andrade, Aldri L. dos Santos, Bruno Schulze, and José N. de Souza. Detecting misbehaving units on computational grids. Concurrency and Computation: Practice and Experience, 22(3):329–342, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Murai:2002:IEH

[MAH^{+}02] Hitoshi Murai, Takuya Araki, Yasuharu Hayashi, Kenji Suehiro, and Yoshiki Seo. Implementation and evaluation of HPF/SX V2. Concurrency and Computation: Practice and Experience, 14(8–9):603–629, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016132/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016132{\&}PLACEBO=IE.pdf.

Muhtaroglu:2018:DHC

[MAK18] Nitel Muhtaroglu, Ismail Ari, and Birkan Kolcu. Democratization of HPC cloud services with automated parallel solvers and application containers. Concurrency and Computation: Practice and Experience, 30(21):e4782:1–e4782:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malard:2005:RPO

[Mal05] Joël M. Malard. A role for Pareto optimality in mining performance data. Concurrency and Computation: Practice and Experience, 17(1):27–48, January 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mancini:2008:SIH

[Man08] Luigi V. Mancini. Special issue: Hot topics in peer-to-peer systems. Concurrency and Computation: Practice and Experience, 20(6):599, April 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marinescu:2005:SIT

[Mar05] Dan Marinescu. Special issue: Third IEEE International Workshop on High Performance Computational Biology (HiCOMB 2004). Concurrency and Computation: Practice and Experience, 17(14):1625–1626, December 10, 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marchiori:2019:LWC

[Mar19] Massimo Marchiori. Learning the way to the cloud: Big Data Park. Concurrency and Computation: Practice and Experience, 31(2):e4234:1–e4234:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Menasche:2014:ASS

[MAS^{+}14] Daniel Sadoc Menasché, Alberto Avritzer, Sindhu Suresh, Rosa M. Leão, Edmundo de Souza e Silva, Morganna Diniz, Kishor Trivedi, Lucia Happe, and Anne Koziolek. Assessing survivability of smart grid distribution network designs accounting for multiple failures. Concurrency and Computation: Practice and Experience, 26(12):1949–1974, August 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Magoules:2016:GCG

[MAS16] Frédéric Magoulès, Abal-Kassim Cheik Ahamed, and Atsushi Suzuki. Green computing on graphics processing units. Concurrency and Computation: Practice and Experience, 28(16): 4305–4325, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martin:2016:MLT

[MAVG16] R. Martin, R. Aler, J. M. Valls, and I. M. Galvan. Machine learning techniques for daily solar energy prediction and interpolation using numerical weather models. Concurrency and Computation: Practice and Experience, 28(4):1261–1274, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

May:2010:CPA

[May10] D. May. Communicating process architecture for multicores. Concurrency and Computation: Practice and Experience, 22 (8):935–948, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mayhew:2018:S

[May18] David Mayhew. SAHARA. Concurrency and Computation: Practice and Experience, 30(17):e4708:1–e4708:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moritsch:2002:HPN

[MB02] Hans Moritsch and Siegfried Benkner. High-performance numerical pricing methods. Concurrency and Computation: Practice and Experience, 14(8–9):665–678, July/ August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016129/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016129{\&}PLACEBO=IE.pdf.

Manias:2012:CBM

[MB12] Elton Manias and Françoise Baude. A component-based middleware for hybrid grid/cloud computing platforms. Concurrency and Computation: Practice and Experience, 24(13): 1461–1477, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meng:2014:SLS

[MB14] Qingyu Meng and Martin Berzins. Scalable large-scale fluid–structure interaction solvers in the Uintah framework via hybrid task-based parallelism algorithms. Concurrency and Computation: Practice and Experience, 26(7):1388–1407, May 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mustafee:2015:EIT

[MB15] Navonil Mustafee and Nik Bessis. Editorials: The Internet of Things: shaping the new Internet space. Concurrency and Computation: Practice and Experience, 27(8):1815–1818, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moreno:2016:IFC

[MB16] Alexander Moreno and Tucker Balch. Improving financial computation speed with full and subproblem memoization. Concurrency and Computation: Practice and Experience, 28 (3):905–915, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Machaj:2017:IOA

[MB17] Juraj Machaj and Peter Brida. Impact of optimization algorithms on hybrid indoor positioning based on GSM and Wi-Fi signals. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meroufel:2018:OCR

[MB18] Bakhta Meroufel and Ghalem Belalem. Optimization of checkpointing/recovery strategy in cloud computing with adaptive storage management. Concurrency and Computation: Practice and Experience, 30(24):e4906:1–e4906:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maragatharajan:2019:SMU

[MBB19] M. Maragatharajan, C. Balasubramanian, and S. P. Balakannan. A secured MANET using position-based opportunistic routing and SEMI MARKOV process. Concurrency and Computation: Practice and Experience, 31(14):e5047:1– e5047:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martinez:2014:MPP

[MBC^{+}14] D. R. Martínez, V. Blanco, J. C. Cabaleiro, T. F. Pena, and F. F. Rivera. Modeling the performance of parallel applications using model selection techniques. Concurrency and Computation: Practice and Experience, 26(2):586–599, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahmoudi:2018:TSS

[MBMB18] Sidi Ahmed Mahmoudi, Mohammed Amin Belarbi, Saïd Mahmoudi, and Ghalem Belalem. Towards a smart selection of resources in the cloud for low-energy multimedia processing. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4372.

Matossian:2005:AOR

[MBP^{+}05] Vincent Matossian, Viraj Bhat, Manish Parashar, Małgorzata Peszyńska, Mrinal Sen, Paul Stoffa, and Mary F. Wheeler. Autonomic oil reservoir optimization on the Grid. Concurrency and Computation: Practice and Experience, 17(1):1–26, January 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maamar:2016:PEM

[MBP16] Haifa Raja Maamar, Azzedine Boukerche, and Emil Petriu. A performance evaluation of mobility management and multihop supplying partner strategies for 3D streaming systems over thin mobile devices. Concurrency and Computation: Practice and Experience, 28(6):1769–1795, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mellor-Crummey:2002:AOS

[MCAB^{+}02] J. Mellor-Crummey, V. Adve, B. Broom, D. Chavarría-Miranda, R. Fowler, G. Jin, K. Kennedy, and Q. Yi. Advanced optimization strategies in the Rice dHPF compiler. Concurrency and Computation: Practice and Experience, 14(8–9):741–767, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/95016130/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=95016130{\&}PLACEBO=
IE.pdf.

Melab:2014:GPU

[MCB14] N. Melab, I. Chakroun, and A. Bendjoudi. Graphics processing unit-accelerated bounding for branch-and-bound applied to a permutation problem using data access optimization.

Concurrency and Computation: Practice and Experience, 26 (16):2667–2683, November 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Madduri:2015:GGP

[MCC^{+}15] Ravi Madduri, Kyle Chard, Ryan Chard, Lukasz Lacinski, Alex Rodriguez, Dinanath Sulakhe, David Kelly, Utpal Dave, and Ian Foster. The Globus Galaxies platform: delivering science gateways as a service. Concurrency and Computation: Practice and Experience, 27(16):4344–4360, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Morelli:2016:HLA

[MCC16] Davide Morelli, Andrea Canciani, and Antonio Cisternino. A high-level and accurate energy model of parallel and concurrent workloads. Concurrency and Computation: Practice and Experience, 28(3):822–833, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mele:2018:PPT

[MCCD18] Valeria Mele, Emil M. Constantinescu, Luisa Carracciuolo, and Luisa D’Amore. A PETSc parallel-in-time solver based on MGRIT algorithm. Concurrency and Computation: Practice and Experience, 30(24):e4928:1–e4928:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Merlo:2011:QSG

[MCCG11] A. Merlo, A. Clematis, A. Corana, and V. Gianuzzi. Quality of Service on Grid: architectural and methodological issues. Concurrency and Computation: Practice and Experience, 23 (8):745–766, June 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McLennan:2015:HPI

[MCD^{+}15] Michael McLennan, Steven Clark, Ewa Deelman, Mats Rynge, Karan Vahi, Frank McKenna, Derrick Kearney, and Carol Song. HUBzero and Pegasus: integrating scientific workflows into science gateways. Concurrency and Computation: Practice and Experience, 27(2):328–343, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McEwan:2010:GE

[McE10] Alistair A. McEwan. Guest editorial. Concurrency and Computation: Practice and Experience, 22(8):909–911, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Man:2008:DLS

[MCG^{+}08] Billy Yan-Kit Man, Hiu Ning (Angela) Chan, Andrew J. Gallagher, Appu S. Goundan, Aaron W. Keen, and Ronald A. Olsson. A definition of and linguistic support for partial quiescence. Concurrency and Computation: Practice and Experience, 20(8):969–995, June 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mastroianni:2018:ESE

[MCG18] Carlo Mastroianni, Eugenio Cesario, and Andrea Giordano. Efficient and scalable execution of smart city parallel applications. Concurrency and Computation: Practice and Experience, 30(20):e4258:1–e4258:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marker:2012:PMC

[MCP^{+}12] Bryan Marker, Ernie Chan, Jack Poulson, Robert van de Geijn, Rob F. Van der Wijngaart, Timothy G. Mattson, and Theodore E. Kubaska. Programming many-core architectures — a case study: dense matrix computations on the Intel single-chip cloud computer processor. Concurrency and Computation: Practice and Experience, 24(12):1317–1333, August 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Morajko:2007:MMA

[MCSML07] A. Morajko, P. Caymes-Scutari, T. Margalef, and E. Luque. MATE: Monitoring, Analysis and Tuning Environment for parallel/distributed applications. Concurrency and Computation: Practice and Experience, 19(11):1517–1531, August 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2006:GPI

[MCWL06] Tianchi Ma, Lin Chen, Cho-Li Wang, and Francis C. M. Lau. G-PASS: an instance-oriented security infrastructure for Grid travelers. Concurrency and Computation: Practice and Experience, 18(14):1871–1884, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miguel:2015:MDP

[MCXP15] Jorge Miguel, Santi Caballé, Fatos Xhafa, and Josep Prieto. A massive data processing approach for effective trustworthiness in online learning groups. Concurrency and Computation: Practice and Experience, 27(8):1988–2003, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mackenzie:2007:AGC

[MCY^{+}07] Lewis Mackenzie, Paul Cockshott, Viktor Yarmolenko, Ewan Borland, Paul Graham, and Kostas Kavoussanakis. Applying the Grid to 3D capture technology. Concurrency and Computation: Practice and Experience, 19(2):235–249, February 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McLane:2010:UIV

[MCY^{+}10] Jonathan C. McLane, W. Walter Czech, David A. Yuen, Mike R. Knox, Shuo Wang, Jim B. S. Greensky, and Erik O. D. Sevre. Ubiquitous interactive visualization of largescale simulations in geosciences over a Java-based web-portal. Concurrency and Computation: Practice and Experience, 22 (12):1750–1773, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moore:2002:NTC

[MD02] Keith Moore and Jack Dongarra. NetBuild: transparent cross-platform access to computational software libraries. Concurrency and Computation: Practice and Experience, 14(13–15):1445–1456, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Margheritta:2019:LHB

[MD19] Paul Margheritta and Michel R. Dagenais. LTTng–HSA: Bringing LTTng tracing to HSA-based GPU runtimes. Concurrency and Computation: Practice and Experience, 31(17): e5231:1–e5231:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marozzo:2017:DAS

[MDB^{+}17] Fabrizio Marozzo, Francisco Rodrigo Duro, Javier Garcia Blas, Jesus Carretero, Domenico Talia, and Paolo Trunfio. A data-aware scheduling strategy for workflow execution in clouds. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mehrotra:2016:PEA

[MDH^{+}16] Piyush Mehrotra, Jahed Djomehri, Steve Heistand, Robert Hood, Haoqiang Jin, Arthur Lazanoff, Subhash Saini, and Rupak Biswas. Performance evaluation of Amazon Elastic Compute Cloud for NASA high-performance computing applications. Concurrency and Computation: Practice and Experience, 28(4):1041–1055, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Murli:2010:MGD

[MDL^{+}10] A. Murli, L. D’Amore, G. Laccetti, F. Gregoretti, and G. Oliva. A multi-grained distributed implementation of the parallel Block Conjugate Gradient algorithm. Concurrency and Computation: Practice and Experience, 22(15): 2053–2072, October 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohr:2007:FPA

[MDV07] Bernd Mohr, Luiz DeRose, and Jeffrey Vetter. A framework for performance analysis of Co-Array Fortran. Concurrency and Computation: Practice and Experience, 19(17): 2207–2218, December 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marin:2014:MAP

[MDX14] Radu-Corneliu Marin, Ciprian Dobre, and Fatos Xhafa. A methodology for assessing the predictable behaviour of mobile users in wireless networks. Concurrency and Computation: Practice and Experience, 26(5):1215–1230, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohamed:2008:KCA

[ME08] Hashim Mohamed and Dick Epema. KOALA: a co-allocating grid scheduler. Concurrency and Computation: Practice and Experience, 20(16):1851–1876, November 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Menezes:2003:RMO

[Men03] Ronaldo Menezes. Resource management in open Linda systems. Concurrency and Computation: Practice and Experience, 15(13):1233–1256, November 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moore:2018:LLE

[MFC18] Michael Moore, Patrick Farrell, and Bob Cernohous. Lustre Lockahead: Early experience and performance using optimized locking. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McLean:2004:MRT

[MFF04] Thom McLean, Richard Fujimoto, and Brad Fitzgibbons. Middleware for real-time distributed simulations. Concurrency and Computation: Practice and Experience, 16(15): 1483–1501, December 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McGough:2013:SIP

[MFG^{+}13] Andrew Stephen McGough, Matthew Forshaw, Clive Gerrard, Paul Robinson, and Stuart Wheater. Special issue papers: Analysis of power-saving techniques over a large multiuse cluster with variable workload. Concurrency and Computation: Practice and Experience, 25(18):2501–2522, December 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moreton-Fernandez:2019:ARC

[MFGE19] Ana Moreton-Fernandez and Arturo Gonzalez-Escribano. Automatic runtime calculation of communications for dataparallel expressions with periodic conditions. Concurrency and Computation: Practice and Experience, 31(5):e4430:1– e4430:??, March 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maloney:2009:SRC

[MG09a] Andrew Maloney and Andrzej Goscinski. A survey and review of the current state of rollback-recovery for cluster systems. Concurrency and Computation: Practice and Experience, 21 (12):1632–1666, August 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Merrill:2009:PCS

[MG09b] Duane Merrill and Andrew Grimshaw. Profiles for conveying the secure communication requirements of Web services. Concurrency and Computation: Practice and Experience, 21 (8):991–1011, June 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Magoules:2016:RTD

[MGBC16] Frédéric Magoulès, Guillaume Gbikpi-Benissan, and Patrick Callet. Ray-tracing domain decomposition methods for realtime simulation on multi-core and multi-processor systems. Concurrency and Computation: Practice and Experience, 28 (16):4352–4364, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martinsen:2017:CTL

[MGI17] Jan Kasper Martinsen, Håkan Grahn, and Anders Isberg. Combining thread-level speculation and just-in-time compilation in Google’s V8 JavaScript engine. Concurrency and Computation: Practice and Experience, 29(1):??, January 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miles:2008:ECG

[MGM^{+}08] Simon Miles, Paul Groth, Steve Munroe, Sheng Jiang, Thibaut Assandri, and Luc Moreau. Extracting causal graphs from an open provenance data model. Concurrency and Computation: Practice and Experience, 20(5):577–586, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muller:2002:SDE

[MGR02] M. S. Müller, E. Gabriel, and M. M. Resch. A software development environment for Grid computing. Concurrency and Computation: Practice and Experience, 14(13–15):1543– 1551, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mehrabi:2019:PUP

[MGS19] Mostafa Mehrabi, Nasser Giacaman, and Oliver Sinnen. @PT: Unobtrusive parallel programming with Java annotations. Concurrency and Computation: Practice and Experience, 31(1):e4831:1–e4831:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McMullen:2007:CUI

[MH07] Donald F. McMullen and Kianosh Huffman. Connecting users to instruments and sensors: portals as multi-user GUIs for instrument and sensor facilities. Concurrency and Computation: Practice and Experience, 19(12):1621–1631, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Michael:2018:EFS

[MH18] Scott Michael and Yun He. Editorial: Foreword to the special issue of the Cray User Group (CUG 2017). Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mansouri:2016:DSH

[MHH16] Farouk Mansouri, Sylvain Huet, and Dominque Houzet. A domain-specific high-level programming model. Concurrency and Computation: Practice and Experience, 28(3):750–767, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Magoules:2016:EEM

[MHJH16] Frédéric Magoulès, Che-Lun Hung, Hai Jiang, and Jia Hu. Editorial: Embedded multicore computing and applications. Concurrency and Computation: Practice and Experience, 28 (16):4211–4214, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malik:2018:GMG

[MHK^{+}18] Kaleem Razzaq Malik, Muhammad Asif Habib, Shehzad Khalid, Mudassar Ahmad, Mai Alfawair, Awais Ahmad, and Gwanggil Jeon. A generic methodology for geo-related data semantic annotation. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4495.

Malard:2005:PIC

[MHLC^{+}05] J. M. Malard, A. Heredia-Langner, W. R. Cannon, R. Mooney, and D. J. Baxter. Peptide identification via constrained multi-objective optimization: Pareto-based genetic algorithms. Concurrency and Computation: Practice and Experience, 17(14):1687–1704, December 10, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohan:2014:BHV

[MHRI14] Anup Mohan, Thomas Hacker, Gregory P. Rodgers, and Tanzima Islam. Batchsubmit: a high-volume batch submission system for earthquake engineering simulation. Concurrency and Computation: Practice and Experience, 26(13): 2240–2252, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maabreh:2018:MHT

[MIGA18] Majdi Maabreh, Hafez Irshid, Ajay Gupta, and Izzat Alasmadi. A multithreading and hashing technique for indexing Target–Decoy peptides databases. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4371.

Mejia:2019:ARI

[MÍM19] Jezreel Mejía, Freddy Íñiguez, and Mirna Muñoz. Analyzing the requirements to implement a data analysis model for software process improvement. Concurrency and Computation: Practice and Experience, 31(22):e4438:1–e4438:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meawad:2013:SIP

[MISV13] Fadi Meawad, Karthik Iyer, Martin Schoeberl, and Jan Vitek. Special issue papers: Micro-transactions for concurrent data structures. Concurrency and Computation: Practice and Experience, 25(16):2252–2268, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mittal:2017:STA

[Mit17a] Sparsh Mittal. A survey of techniques for architecting TLBs. Concurrency and Computation: Practice and Experience, 29 (10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mittal:2017:STD

[Mit17b] Sparsh Mittal. A survey of techniques for designing and managing CPU register file. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mittal:2017:SVP

[Mit17c] Sparsh Mittal. A survey of value prediction techniques for leveraging value locality. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mittal:2019:STD

[Mit19] Sparsh Mittal. A survey of techniques for dynamic branch prediction. Concurrency and Computation: Practice and Experience, 31(1):e4666:1–e4666:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marzouk:2011:SSC

[MJ11] Soumaya Marzouk and Mohamed Jmaiel. A survey on software checkpointing and mobility techniques in distributed systems. Concurrency and Computation: Practice and Experience, 23(11):1196–1212, August 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohanapriya:2019:PIV

[MJ19] R. Mohanapriya and K. B. Jayanthi. Performance improvement in vertical heterogeneous handoff methodology using CANFIS classification approaches. Concurrency and Computation: Practice and Experience, 31(14):e5062:1–e5062:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miao:2015:NHS

[MJD15] Xinqiang Miao, Xianlong Jin, and Junhong Ding. A new hybrid solver with two-level parallel computing for large-scale structural analysis. Concurrency and Computation: Practice and Experience, 27(14):3661–3675, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miao:2017:AEP

[MJD17] Xinqiang Miao, Xianlong Jin, and Junhong Ding. An approach to enhance the performance of large-scale structural analysis on CPU-MIC heterogeneous clusters. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marinescu:2001:STT

[MJL01] Dan C. Marinescu, Yongchang Ji, and Robert E. Lynch. Space-time tradeoffs for parallel 3D reconstruction algorithms for virus-structure determination. Concurrency and Computation: Practice and Experience, 13(12):1083–1106, October 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/85512007/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
85512007&PLACEBO=IE.pdf.

Miao:2019:DPB

[MJS19] Qiucheng Miao, Weipeng Jing, and Houbing Song. Differential privacy-based location privacy enhancing in edge computing. Concurrency and Computation: Practice and Experience, 31(8):e4735:1–e4735:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2017:NAH

[MJZ17] Youzhong Ma, Shijie Jia, and Yongxin Zhang. A novel approach for high-dimensional vector similarity join query. Concurrency and Computation: Practice and Experience, 29(5): ??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mentis:2012:MCC

[MK12] Anakreon Mentis and Panagiotis Katsaros. Model checking and code generation for transaction processing software. Concurrency and Computation: Practice and Experience, 24(7): 711–722, May 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marmol:2015:RBW

[MK15a] Félix Gómez Mármol and Marcus Quintino Kuhnen. Reputation-based Web service orchestration in cloud computing: a survey. Concurrency and Computation: Practice and Experience, 27(9):2390–2412, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moschakis:2015:TSI

[MK15b] Ioannis A. Moschakis and Helen D. Karatza. Towards scheduling for Internet-of-Things applications on clouds: a simulated annealing approach. Concurrency and Computation: Practice and Experience, 27(8):1886–1899, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meier-Kolthoff:2014:HPI

[MKAKG14] Jan P. Meier-Kolthoff, Alexander F. Auch, Hans-Peter Klenk, and Markus Göker. Highly parallelized inference of large genome-based phylogenies. Concurrency and Computation: Practice and Experience, 26(10):1715–1729, July 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maassen:2001:PAE

[MKB01] Jason Maassen, Thilo Kielmann, and Henri E. Bal. Parallel application experience with replicated method invocation. Concurrency and Computation: Practice and Experience, 13(8–9):681–712, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/84503225/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=84503225&PLACEBO=IE.
pdf.

Maris:2004:CCP

[MKIO04] Justin T. Maris, Aaron W. Keen, Takashi Ishihara, and Ronald A. Olsson. A comparison of concurrent programming and cooperative multithreading under load balancing applications. Concurrency and Computation: Practice and Experience, 16(4):345–369, April 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meier:2004:DIP

[MKKB04] H. Meier, A. Krause, M. Kräutner, and A. Bode. Development and implementation of a parallel algorithm for the fast design of oligonucleotide probe sets for diagnostic DNA microarrays. Concurrency and Computation: Practice and Experience, 16(9):873–893, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maheshwari:2017:ERF

[MKO^{+}17a] Ketan Maheshwari, Daniel Katz, Silvia D. Olabarriaga, Justin Wozniak, and Douglas Thain. Editorials: Report on the first workshop on negative and null results in eScience. Concurrency and Computation: Practice and Experience, 29 (2):??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Montella:2017:ALA

[MKO^{+}17b] Raffaele Montella, Sokol Kosta, David Oro, Javier Vera, Carles Fernández, Carlo Palmieri, Diana Di Luccio, Giulio Giunta, Marco Lapegna, and Giuliano Laccetti. Accelerating Linux and Android applications on low-power devices through remote GPGPU offloading. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Morgan:2016:SPM

[MKSS16] Hannah Morgan, Matthew G. Knepley, Patrick Sanan, and L. Ridgway Scott. A stochastic performance model for pipelined Krylov methods. Concurrency and Computation: Practice and Experience, 28(18):4532–4542, December 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Montella:2015:FIS

[MKX^{+}15] Raffaele Montella, David Kelly, Wei Xiong, Alison Brizius, Joshua Elliott, Ravi Madduri, Ketan Maheshwari, Cheryl Porter, Peter Vilter, Michael Wilde, Meng Zhang, and Ian Foster. FACE-IT: a science gateway for food security research. Concurrency and Computation: Practice and Experience, 27(16):4423–4436, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Manumachu:2019:DSA

[ML19a] Ravi Reddy Manumachu and Alexey L. Lastovetsky. Design of self-adaptable data parallel applications on multicore clusters automatically optimized for performance and energy through load distribution. Concurrency and Computation: Practice and Experience, 31(4):e4958:1–e4958:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mathur:2019:SNC

[ML19b] Saurabh Mathur and Daphne Lopez. A scaled-down neural conversational model for chatbots. Concurrency and Computation: Practice and Experience, 31(10):e4761:1–e4761:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moreau:2008:SIF

[MLA^{+}08] Luc Moreau, Bertram Ludäscher, Ilkay Altintas, Roger S. Barga, Shawn Bowers, Steven Callahan, George Chin, Jr., Ben Clifford, Shirley Cohen, Sarah Cohen-Boulakia, Susan Davidson, Ewa Deelman, Luciano Digiampietri, Ian Foster, Juliana Freire, James Frew, Joe Futrelle, Tara Gibson, Yolanda Gil, Carole Goble, Jennifer Golbeck, Paul Groth, David A. Holland, Sheng Jiang, Jihie Kim, David Koop, Ales Krenek, Timothy McPhillips, Gaurang Mehta, Simon Miles, Dominic Metzger, Steve Munroe, Jim Myers, Beth Plale, Norbert Podhorszki, Varun Ratnakar, Emanuele Santos, Carlos Scheidegger, Karen Schuchardt, Margo Seltzer, Yogesh L. Simmhan, Claudio Silva, Peter Slaughter, Eric Stephan, Robert Stevens, Daniele Turi, Huy Vo, Mike Wilde, Jun Zhao, and Yong Zhao. Special issue: The First Provenance Challenge. Concurrency and Computation: Practice and Experience, 20(5):409–418, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marowka:2004:OOA

[MLC04] Ami Marowka, Zhenying Liu, and Barbara Chapman. OpenMP-oriented applications for distributed shared memory architectures. Concurrency and Computation: Practice and Experience, 16(4):371–384, April 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2015:SSA

[MLG15] Jianwei Ma, Wanyu Liu, and Tristan Glatard. A stateful storage availability and entropy model to control storage distribution on grids. Concurrency and Computation: Practice and Experience, 27(1):164–171, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2003:EOS

[mLGP03] Li min Liu, James Geller, and Yehoshua Perl. Enhancing OODB semantics to support browsing in an OODB vocabulary representation. Concurrency and Computation: Practice and Experience, 15(9):845–869, August 10, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2011:SEP

[MLL^{+}11] Xiaopu Ma, Ruixuan Li, Zhengding Lu, Jianfeng Lu, and Meng Dong. Specifying and enforcing the principle of least privilege in role-based access control. Concurrency and Computation: Practice and Experience, 23(12):1313–1331, August 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Midkiff:2004:CMM

[MLP04] S. P. Midkiff, J. Lee, and D. A. Padua. A compiler for multiple memory models. Concurrency and Computation: Practice and Experience, 16(2–3):197–220, February/March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miranda:2009:ADR

[MLRR09] H. Miranda, S. Leggio, L. Rodrigues, and K. Raatikainen. An algorithm for dissemination and retrieval of information in wireless ad hoc networks. Concurrency and Computation: Practice and Experience, 21(7):889–904, May 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malik:2012:PGU

[MLS^{+}12] Maria Malik, Teng Li, Umar Sharif, Rabia Shahid, Tarek El-Ghazawi, and Greg Newby. Productivity of GPUs under different programming paradigms. Concurrency and Computation: Practice and Experience, 24(2):179–191, February 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2015:DRD

[MLS^{+}15] Tinghuai Ma, Yinhua Lu, Sunyuan Shi, Wei Tian, Xin Wang, and Donghai Guan. Data resource discovery model based on hybrid architecture in data grid environment. Concurrency and Computation: Practice and Experience, 27(3):507–525, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Markidis:2005:IPP

[MLVB05] S. Markidis, G. Lapenta, W. B. VanderHeyden, and Z. Budimlić. Implementation and performance of a particlein-cell code written in Java. Concurrency and Computation: Practice and Experience, 17(7–8):821–837, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Munir:2012:PDT

[MLVBW12] Kashif Munir, Renato Lo Cigno, Pascale Primet Vicat-Blanc, and Michael Welzl. Planning data transfers in grids: a multiservice queueing approach. Concurrency and Computation: Practice and Experience, 24(4):407–422, March 25, 2012. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mu:2019:PTN

[MLWA19] Yashuang Mu, Xiaodong Liu, Lidong Wang, and Aamer Bilal Asghar. A parallel tree node splitting criterion for fuzzy decision trees. Concurrency and Computation: Practice and Experience, 31(17):e5268:1–e5268:??, September 10, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Min:2010:SIA

[MLY10] Geyong Min, Keqiu Li, and Laurence T. Yang. Special issue: Advances in high-performance computing and communications. Concurrency and Computation: Practice and Experience, 22(4):395–397, March 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mu:2017:PCD

[MLYL17] Yashuang Mu, Xiaodong Liu, Zhihao Yang, and Xiaolin Liu. A parallel C4.5 decision tree algorithm based on MapReduce. Concurrency and Computation: Practice and Experience, 29 (8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2019:FIU

[MLZ19] Qichen Ma, Xiangfeng Luo, and Hai Zhuge. Finding influential users of web event in social media. Concurrency and Computation: Practice and Experience, 31(3):e5029:1– e5029:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mucci:2010:OSP

[MM10] Philip J. Mucci and Tushar Mohan. An Open Source performance tools software suite for scientific computing. Concurrency and Computation: Practice and Experience, 22(2): 206–216, February 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Melab:2017:EMM

[MM17] Nouredine Melab and Mohand Mezmaz. Editorials: Multi and many-core computing for parallel metaheuristics. Concurrency and Computation: Practice and Experience, 29(9): ??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mittal:2019:STI

[MM19] Sparsh Mittal and Venkat Mattela. A survey of techniques for improving efficiency of mobile web browsing. Concurrency and Computation: Practice and Experience, 31(15):e5126:1– e5126:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Monnet:2017:DWE

[MMB^{+}17] Quentin Monnet, Lynda Mokdad, Paolo Ballarini, Youcef Hammal, and Jalel Ben-Othman. DoS detection in WSNs: Energy-efficient methods for selecting monitoring nodes. Concurrency and Computation: Practice and Experience, 29(23): ??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marmol:2012:LLF

[MMBP12] Félix Gómez Mármol, Javier G. Marín-Blázquez, and Gregorio Martínez Pérez. LFTM, linguistic fuzzy trust mechanism for distributed networks. Concurrency and Computation: Practice and Experience, 24(17):2007–2027, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meulpolder:2013:PUC

[MME13] M. Meulpolder, L. E. Meester, and D. H. J. Epema. The problem of upload competition in peer-to-peer systems with incentive mechanisms. Concurrency and Computation: Practice and Experience, 25(7):899–917, May 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moreira:2003:SMA

[MMG03] José E. Moreira, Samuel P. Midkiff, and Manish Gupta. Supporting multidimensional arrays in Java. Concurrency and Computation: Practice and Experience, 15(3–5):317–340, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Montella:2018:MBP

[MMG^{+}18] Raffaele Montella, Livia Marcellino, Ardelio Galletti, Diana Di Luccio, Sokol Kosta, Giuliano Laccetti, and Giulio Giunta. Marine bathymetry processing through GPGPU virtualization in high performance cloud computing. Concurrency and Computation: Practice and Experience, 30(24):e4895:1– e4895:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miao:2016:RAS

[MML16] Yinbin Miao, Jianfeng Ma, and Zhiquan Liu. Revocable and anonymous searchable encryption in multi-user setting. Concurrency and Computation: Practice and Experience, 28(4): 1204–1218, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mendez:2017:IAB

[MML^{+}17] Alejandro Pérez Méndez, Gabriel López Millán, Rafael Marín López, David W. Chadwick, and Ioram Schechtman Sette. Integrating an AAA-based federation mechanism for Open-Stack — the CLASSe view. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mann:2001:DEW

[MMMP01] V. Mann, V. Matossian, R. Muralidhar, and M. Parashar. DISCOVER: an environment for Web-based interaction and steering of high-performance scientific applications. Concurrency and Computation: Practice and Experience, 13(8–9): 737–754, July/August 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/84503217/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=84503217&PLACEBO=IE.pdf.

Manca:2016:CQI

[MMO^{+}16] Emanuele Manca, Andrea Manconi, Alessandro Orro, Giuliano Armano, and Luciano Milanesi. CUDA-quicksort: an improved GPU-based implementation of quicksort. Concurrency and Computation: Practice and Experience, 28(1):21– 43, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Massingill:2007:RPE

[MMS07] Berna L. Massingill, Timothy G. Mattson, and Beverly A. Sanders. Reengineering for Parallelism: an entry point into PLPP for legacy applications. Concurrency and Computation: Practice and Experience, 19(4):503–529, March 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mazumder:2017:SAE

[MMS17] Rashed Mazumder, Atsuko Miyaji, and Chunhua Su. A simple authentication encryption scheme. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martineau:2017:APP

[MMSG17] Matthew Martineau, Simon McIntosh-Smith, and Wayne Gaudin. Assessing the performance portability of modern parallel programming models using TeaLeaf. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moser:2001:MGC

[MMSN^{+}01] L. E. Moser, P. M. Melliar-Smith, P. Narasimhan, R. R. Koch, and K. Berket. A multicast group communication protocol, engine, and bridge for CORBA. Concurrency and Computation: Practice and Experience, 13(7):579–603, June 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/83002176/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
83002176&PLACEBO=IE.pdf.

Marinho:2012:PPM

[MMW^{+}12] Anderson Marinho, Leonardo Murta, Cláudia Werner, Vanessa Braganholo, Sérgio Manuel Serra da Cruz, Eduardo Ogasawara, and Marta Mattoso. ProvManager: a provenance management system for scientific workflows. Concurrency and Computation: Practice and Experience, 24(13): 1513–1530, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2016:PSJ

[MMW16] Youzhong Ma, Xiaofeng Meng, and Shaoya Wang. Parallel similarity joins on massive high-dimensional data using MapReduce. Concurrency and Computation: Practice and Experience, 28(1):166–183, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mokdad:2010:SIP

[MN10] Lynda Mokdad and Mirela Sechi Annoni Notare. Special issue: Performance evaluation of communications in distributed systems and Web-based service architectures. Concurrency and Computation: Practice and Experience, 22(10):

1205–1206, July 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Murtojarvi:2015:PTS

[MNL15] Mika Murtojärvi, Olli S. Nevalainen, and Ville Leppänen. Performance tuning and sparse traversal technique for a cellbased fetch length algorithm on a GPU. Concurrency and Computation: Practice and Experience, 27(17):5114–5133, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Matsui:2002:TCA

[MO02a] Hiroaki Matsui and Hiroshi Okuda. Thermal convection analysis in a rotating shell by a parallel finiteelement method — development of a thermal-hydraulic subsystem of GeoFEM. Concurrency and Computation: Practice and Experience, 14(6–7):465–481, May/ June 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/94515744/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
94515744{\&}PLACEBO=IE.pdf.

Minami:2002:OGH

[MO02b] Kazuo Minami and Hiroshi Okuda. Optimization of GeoFEM for high performance sequential computer architectures. Concurrency and Computation: Practice and Experience, 14(6– 7):395–409, May/June 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/94515746/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=94515746{\&}PLACEBO=IE.pdf.

Miyaji:2015:SHW

[MO15] Atsuko Miyaji and Kazumasa Omote. Self-healing wireless sensor networks. Concurrency and Computation: Practice and Experience, 27(10):2547–2568, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maqbool:2015:EAH

[MOF15] Jahanzeb Maqbool, Sangyoon Oh, and Geoffrey C. Fox. Evaluating ARM HPC clusters for scientific workloads. Concurrency and Computation: Practice and Experience, 27(17):

5390–5410, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Min:2004:PCS

[MOK04] Geyong Min and Mohamed Ould-Khaoua. On the performance of circuit-switched networks in the presence of correlated traffic. Concurrency and Computation: Practice and Experience, 16(13):1313–1326, November 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muyan-Ozcelik:2017:MMA

[MÖO17] Pınar Muyan-Özçelik and John D. Owens. Methods for multitasking among real-time embedded compute tasks running on the GPU. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mosleh:2019:NDM

[Mos19] Mohammad Mosleh. A novel design of multiplexer based on nano-scale quantum-dot cellular automata. Concurrency and Computation: Practice and Experience, 31(13):e5070:1– e5070:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mann:2002:EIC

[MP02] Vijay Mann and Manish Parashar. Engineering an interoperable computational collaboratory on the Grid. Concurrency and Computation: Practice and Experience, 14(13–15):1569– 1593, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muralidhar:2003:DOI

[MP03] Rajeev Muralidhar and Manish Parashar. A distributed object infrastructure for interaction and steering. Concurrency and Computation: Practice and Experience, 15(10):957–977, August 25, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Morse:2004:HLA

[MP04] Katherine L. Morse and Mikel D. Petty. High Level Architecture Data Distribution Management migration from DoD 1.3 to IEEE 1516. Concurrency and Computation: Practice and Experience, 16(15):1527–1543, December 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maheshwari:2005:BMO

[MP05] Piyush Maheshwari and Michael Pang. Benchmarking message-oriented middleware: TIB/RV versus SonicMQ. Concurrency and Computation: Practice and Experience, 17 (12):1507–1526, October 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Memeti:2017:COD

[MP17] Suejb Memeti and Sabri Pllana. Combinatorial optimization of DNA sequence analysis on heterogeneous systems. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muller:2003:MSF

[MPHL03] Peter Müller, Arnd Poetzsch-Heffter, and Gary T. Leavens. Modular specification of frame properties in JML. Concurrency and Computation: Practice and Experience, 15(2):117– 154, February 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marongiu:2004:SHD

[MPR04] A. Marongiu, P. Palazzari, and V. Rosato. A specialized hardware device for the protein similarity search. Concurrency and Computation: Practice and Experience, 16(9):917– 931, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mao:2011:OMB

[MPS11] Ming Mao, Yefei Peng, and Michael Spring. Ontology mapping: as a binary classification problem. Concurrency and Computation: Practice and Experience, 23(9):1010–1025, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Merelli:2014:ELA

[MPSGD14] Ivan Merelli, Horacio Pérez-Sánchez, Sandra Gesing, and Daniele D’Agostino. Editorial: Latest advances in distributed, parallel, and graphic processing unit accelerated approaches to computational biology. Concurrency and Computation: Practice and Experience, 26(10):1699–1704, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Messina:2007:DEH

[MPT07] Fabrizio Messina, Giuseppe Pappalardo, and Emiliano Tramontana. Design and evaluation of a high-level Grid communication infrastructure. Concurrency and Computation: Practice and Experience, 19(9):1299–1316, June 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mishra:2017:VVA

[MPVT17] Preeti Mishra, Emmanuel S. Pilli, Vijay Varadharajan, and Udaya Tupakula. VAED: VMI-assisted evasion detection approach for infrastructure as a service cloud. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mayo:2001:PSD

[MQOQOH01] Rafael Mayo, Enrique S. Quintana-Ortí, Gregorio Quintana-Ortí, and Vicente Hernández. Parallel solvers for discretetime algebric Riccati equations. Concurrency and Computation: Practice and Experience, 13(2):153–162, February 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/77004417/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
77004417&PLACEBO=IE.pdf.

Milthorpe:2014:PFI

[MRH14] Josh Milthorpe, Alistair P. Rendell, and Thomas Huber. PGAS-FMM: Implementing a distributed fast multipole method using the X10 programming language. Concurrency and Computation: Practice and Experience, 26(3):712–727, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Memon:2014:AUS

[MRJ^{+}14] Shahbaz Memon, Morris Riedel, Florian Janetzko, Borries Demeler, Gary Gorbet, Suresh Marru, Andrew Grimshaw, Lahiru Gunathilake, Raminder Singh, Norbert Attig, and Thomas Lippert. Advancements of the UltraScan scientific gateway for open standards-based cyberinfrastructures. Concurrency and Computation: Practice and Experience, 26(13): 2280–2291, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malik:2016:NAO

[MRL16] Tania Malik, Vladimir Rychkov, and Alexey Lastovetsky. Network-aware optimization of communications for parallel matrix multiplication on hierarchical HPC platforms. Concurrency and Computation: Practice and Experience, 28(3): 802–821, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Macia:2015:PNB

[MRMC15] H. Macia, M. C. Ruiz, J. A. Mateo, and J. L. Calleja. Petri nets-based model for the analysis of NORIA protocol. Concurrency and Computation: Practice and Experience, 27(17): 4704–4715, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miao:2018:CAN

[MRP^{+}18] Yuantian Miao, Zichan Ruan, Lei Pan, Jun Zhang, and Yang Xiang. Comprehensive analysis of network traffic data. Concurrency and Computation: Practice and Experience, 30(5):??, March 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4181.

Martens:2003:DQS

[MRS03] Holger Märtens, Erhard Rahm, and Thomas Stöhr. Dynamic query scheduling in parallel data warehouses. Concurrency and Computation: Practice and Experience, 15(11–12):1169– 1190, September 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Milanes:2008:SAH

[MRS08] A. Milanés, N. Rodriguez, and B. Schulze. State of the art in heterogeneous strong migration of computations. Concurrency and Computation: Practice and Experience, 20(13): 1485–1508, September 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Murray:2009:EGV

[MRS^{+}09] Norman Murray, Dave Roberts, Anthony Steed, Paul Sharkey, Paul Dickerson, John Rae, and Robin Wolff. Eye gaze in virtual environments: evaluating the need and initial work on implementation. Concurrency and Computation: Practice and Experience, 21(11):1437–1449, August 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Macias:2010:MRG

[MRS^{+}10] M. Macías, O. Rana, G. Smith, J. Guitart, and J. Torres. Maximizing revenue in Grid markets using an economically enhanced resource manager. Concurrency and Computation: Practice and Experience, 22(14):1990–2011, September 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2016:DSC

[MRY^{+}16] Tinghuai Ma, Huan Rong, Changhong Ying, Yuan Tian, Abdullah Al-Dhelaan, and Mznah Al-Rodhaan. Detect structural-connected communities based on BSCHEF in C-DBLP. Concurrency and Computation: Practice and Experience, 28(2):311–330, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Milicia:2005:JTC

[MS05] Giuseppe Milicia and Vladimiro Sassone. Jeeg: temporal constraints for the synchronization of concurrent objects. Concurrency and Computation: Practice and Experience, 17(5– 6):539–572, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mar:2007:DCT

[MS07] Choong Hock Mar and Winston K. G. Seah. DS/CDMA throughput of a multi-hop sensor network in a Rayleigh fading underwater acoustic channel. Concurrency and Computation: Practice and Experience, 19(8):1129–1140, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McEwan:2010:MAA

[MS10] Alistair A. McEwan and Steve Schneider. Modelling and analysis of the AMBA bus using CSP and B. Concurrency and Computation: Practice and Experience, 22(8):949–964, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mach:2013:SIP

[MS13] Werner Mach and Erich Schikuta. Special issue papers: Toward an economic and energy-aware cloud cost model. Concurrency and Computation: Practice and Experience, 25(18): 2471–2487, December 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahato:2017:BTA

[MS17a] Dharmendra Prasad Mahato and Ravi Shankar Singh. Balanced task allocation in the on-demand computing-based transaction processing system using social spider optimization. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahato:2017:LBT

[MS17b] Dharmendra Prasad Mahato and Ravi Shankar Singh. Load balanced transaction scheduling using Honey Bee Optimization considering performability in on-demand computing system. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mann:2017:WBA

[MS17c] Zoltán Ádám Mann and Máté Szabó. Which is the best algorithm for virtual machine placement optimization? Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahato:2018:MAT

[MS18] Dharmendra Prasad Mahato and Ravi Shankar Singh. Maximizing availability for task scheduling in on-demand computing-based transaction processing system using ant colony optimization. Concurrency and Computation: Practice and Experience, 30(11):??, June 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4405.

Mukunthan:2019:MPN

[MS19] M. A. Mukunthan and S. Selvakumar. Multilevel Petri netbased ticket assignment and IT management for improved IT organization support. Concurrency and Computation: Practice and Experience, 31(14):e5297:1–e5297:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Menouer:2017:LPS

[MSB17] Tarek Menouer, Nitin Sukhija, and Le Cun Bertrand. A learning Portfolio solver for optimizing the performance of constraint programming problems on multi-core computing systems. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martinez:2018:EFM

[MSCS18] Andrea Martínez, Anna Sikora, Eduardo César, and Joan Sorribes. Evaluating a formal methodology for dynamic tuning of large-scale parallel applications. Concurrency and Computation: Practice and Experience, 30(4):??, February 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4340.

Majd:2018:PIC

[MSD^{+}18] Amin Majd, Golnaz Sahebi, Masoud Daneshtalab, Juha Plosila, Shahriar Lotfi, and Hannu Tenhunen. Parallel imperialist competitive algorithms. Concurrency and Computation: Practice and Experience, 30(7):??, April 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4393.

Mury:2010:TDM

[MSG10] Antonio R. Mury, Bruno Schulze, and Antônio T. A. Gomes. Task distribution models in grids: towards a profile-based approach. Concurrency and Computation: Practice and Experience, 22(3):358–374, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mikkelsen:2019:PRS

[MSK19] Carl Christian Kjelgaard Mikkelsen, Angelika Beatrix Schwarz, and Lars Karlsson. Parallel robust solution of triangular linear systems. Concurrency and Computation: Practice and Experience, 31(19):e5064:1–e5064:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Madduri:2014:EBG

[MSL^{+}14] Ravi K. Madduri, Dinanath Sulakhe, Lukasz Lacinski, Bo Liu, Alex Rodriguez, Kyle Chard, Utpal J. Dave, and Ian T. Foster. Experiences building Globus Genomics: a nextgeneration sequencing analysis service using Galaxy, Globus, and Amazon Web Services. Concurrency and Computation: Practice and Experience, 26(13):2266–2279, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mitchell:2014:PCF

[MSM^{+}14] Lawrence Mitchell, Terence M. Sloan, Muriel Mewissen, Peter Ghazal, Thorsten Forster, Michal Piotrowski, and Arthur Trew. Parallel classification and feature selection in microarray data using SPRINT. Concurrency and Computation: Practice and Experience, 26(4):854–865, March 25, 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

M:2019:IPM

[MSMA19] Anbarasan M., Prakash S., Anand M., and Antonidoss A. Improving performance in mobile ad hoc networks by reliable path selection routing using RPS-LEACH. Concurrency and Computation: Practice and Experience, 31(7):e4984:1– e4984:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marzulo:2019:DCE

[MSN^{+}19] Leandro A. J. Marzulo, Alexandre C. Sena, Alexandre S. Nery, Cristiana Bentes, Igor M. Coelho, Maria Clicia S. de Castro, Saulo T. Oliveira, Tiago A. O. Alves, and Felipe M. G. França. DTM@GPU: Characterizing and evaluating trace redundancy in GPU. Concurrency and Computation: Practice and Experience, 31(18):e4450:1–e4450:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

March:2013:SIP

[MSP^{+}13] José Luis March, Julio Sahuquillo, Salvador Petit, Houcine Hassan, and José Duato. Special issue papers: Power-aware scheduling with effective task migration for real-time multicore embedded systems. Concurrency and Computation: Practice and Experience, 25(14):1987–2001, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Murphy:2019:SRT

[MSP^{+}19] Jeffrey C Murphy, Bhargav Shivkumar, Amy Pritchard, Grant Iraci, Dhruv Kumar, Sun Hyoung Kim, and Lukasz Ziarek. A survey of real-time capabilities in functional languages and compilers. Concurrency and Computation: Practice and Experience, 31(4):e4902:1–e4902:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McIntosh-Smith:2019:PAF

[MSPDP19] Simon McIntosh-Smith, James Price, Tom Deakin, and Andrei Poenaru. A performance analysis of the first generation of HPC-optimized Arm processors. Concurrency and Computation: Practice and Experience, 31(16):e5110:1–e5110:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Medeiros:2016:GCA

[MSS16] B. Medeiros, R. Silva, and J. L. Sobral. Gaspar: a compositional aspect-oriented approach for cluster applications. Concurrency and Computation: Practice and Experience, 28(8): 2353–2373, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Manno:2015:SBF

[MSST15] Giuliano Manno, Waleed W. Smari, Luca Spalazzi, and Gilberto Taccari. A semantic-based federated cloud system for emergency response. Concurrency and Computation:

Practice and Experience, 27(13):3316–3344, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malony:2005:PTP

[MST^{+}05] A. Malony, S. Shende, N. Trebon, J. Ray, R. Armstrong, C. Rasmussen, and M. Sottile. Performance technology for parallel and distributed component software. Concurrency and Computation: Practice and Experience, 17(2–4):117–141, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miller:2013:RFR

[MST13] Timothy Miller, Nagarjuna Surapaneni, and Radu Teodorescu. Runtime failure rate targeting for energy-efficient reliability in chip microprocessors. Concurrency and Computation: Practice and Experience, 25(6):790–807, April 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Missikoff:2015:OBM

[MST15] Michele Missikoff, Fabrizio Smith, and Francesco Taglino. Ontology building and maintenance in collaborative virtual environments. Concurrency and Computation: Practice and Experience, 27(11):2796–2817, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Montes:2010:FOC

[MSV^{+}10] Jesús Montes, Alberto Sánchez, Julio J. Valdés, María S. Pérez, and Pilar Herrero. Finding order in chaos: a behavior model of the whole grid. Concurrency and Computation: Practice and Experience, 22(11):1386–1415, August 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Minson:2008:DRA

[MT08] R. Minson and G. K. Theodoropoulos. Distributing RePast agent-based simulations with HLA. Concurrency and Computation: Practice and Experience, 20(10):1225–1256, July 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mustafee:2009:SSA

[MT09] Navonil Mustafee and Simon J. E. Taylor. Speeding up simulation applications using WinGrid. Concurrency and Computation: Practice and Experience, 21(11):1504–1523, August 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Masko:2019:AGS

[MT19a] Lukasz Maśko and Marek Tudruj. Application global state monitoring in optimization of parallel event-driven simulation. Concurrency and Computation: Practice and Experience, 31(19):e5015:1–e5015:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maurya:2019:EPB

[MT19b] Ashish Kumar Maurya and Anil Kumar Tripathi. An edge priority-based clustering algorithm for multiprocessor environments. Concurrency and Computation: Practice and Experience, 31(11):e5060:1–e5060:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Matsunaga:2007:SGM

[MTA^{+}07] Andréa M. Matsunaga, Maurício O. Tsugawa, Sumalatha Adabala, Renato J. Figueiredo, Herman Lam, and José A. B. Fortes. Science gateways made easy: the In-VIGO approach. Concurrency and Computation: Practice and Experience, 19 (6):905–919, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mock:2002:PCG

[MTD^{+}02] Stephen Mock, Mary Thomas, Maytal Dahan, Kurt Mueller, Catherine Mills, and Gregor von Lazewski. The Perl Commodity Grid Toolkit. Concurrency and Computation: Practice and Experience, 14(13–15):1085–1095, November/ December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2017:NSN

[MTGZ17] Li Ma, Lixin Tao, Keke Gai, and Yong Zhong. A novel social network access control model using logical authorization language in cloud computing. Concurrency and Computation:

Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Michael:2014:MCB

[MTHK14] Scott Michael, Abhinav Thota, Robert Henschel, and Rich Knepper. Making campus bridging work for researchers: Can campus bridging experts accelerate discovery? Concurrency and Computation: Practice and Experience, 26(13): 2141–2148, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mallon:2016:MUB

[MTK16] Damián A. Mallón, Guillermo L. Taboada, and Lars Koesterke. MPI and UPC broadcast, scatter and gather algorithms in Xeon Phi. Concurrency and Computation: Practice and Experience, 28(8):2322–2340, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

M:2019:EEC

[MTM19] Anand M., Sasikala T., and Anbarasan M. Energy efficient channel aware multipath routing protocol for mobile ad-hoc network. Concurrency and Computation: Practice and Experience, 31(4):e4940:1–e4940:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marozzo:2015:JSB

[MTT15] Fabrizio Marozzo, Domenico Talia, and Paolo Trunfio. JS4Cloud: script-based workflow programming for scalable data analysis on cloud platforms. Concurrency and Computation: Practice and Experience, 27(17):5214–5237, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Molnar:2014:DUG

[MTVF14] Péter Molnár, David M. Toth, and Rachel E. Vincent-Finley. Development of undergraduate and graduate programs in computational science. Concurrency and Computation: Practice and Experience, 26(13):2329–2335, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mumcuyan:2018:OBS

[MUKY18] Aras Mumcuyan, Baran Usta, Kamer Kaya, and Hüsnü Yenigün. Optimally bipartitioning sparse matrices with reordering and parallelization. Concurrency and Computation: Practice and Experience, 30(21):e4687:1–e4687:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Murali:2016:QAF

[MV16] Prakash Murali and Sathish Vadhiyar. Qespera: an adaptive framework for prediction of queue waiting times in supercomputer systems. Concurrency and Computation: Practice and Experience, 28(9):2685–2710, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moreno-Vozmediano:2011:EMW

[MVML11] Rafael Moreno-Vozmediano, Ruben S. Montero, and Ignacio M. Llorente. Elastic management of Web server clusters on distributed virtual infrastructures. Concurrency and Computation: Practice and Experience, 23(13):1474–1490, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maassen:2006:MAD

[MvNK^{+}06] Jason Maassen, Rob V. van Nieuwpoort, Thilo Kielmann, Kees Verstoep, and Mathijs den Burger. Middleware adaptation with the Delphoi service. Concurrency and Computation: Practice and Experience, 18(13):1659–1679, November 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Madarbux:2014:TZL

[MVWJ14] Muhammad Ridwan Madarbux, Anouk Van Laer, Philip M. Watts, and Timothy M. Jones. Towards zero latency photonic switching in shared memory networks. Concurrency and Computation: Practice and Experience, 26(15):2551–2566, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muller:2010:SMA

[MvWL^{+}10] Matthias S. Müller, Matthijs van Waveren, Ron Lieberman, Brian Whitney, Hideki Saito, Kalyan Kumaran, John Baron, William C. Brantley, Chris Parrott, Tom Elken, Huiyu Feng, and Carl Ponder. SPEC MPI2007 — an application benchmark suite for parallel systems using MPI. Concurrency and Computation: Practice and Experience, 22(2):191–205, February 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maassen:2017:CUL

[MvWvM^{+}17] Jason Maassen, Ben van Werkhoven, Maarten van Meersbergen, Henri E. Bal, Michael Kliphuis, Sandra E. Brunnabend, Henk A. Dijkstra, Gerben van Malenstein, Migiel de Vos, Sylvia Kuijpers, Sander Boele, Jules Wolfrat, Nick Hill, David Wallom, Christian Grimm, Dieter Kranzlmüller, Dinesh Ganpathi, Shantenu Jha, Yaakoub El Khamra, Frank O. Bryan, Benjamin Kirtman, and Frank J. Seinstra. On the complexities of utilizing large-scale lightpath-connected distributed cyberinfrastructure. Concurrency and Computation: Practice and Experience, 29(2):??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miao:2018:VMA

[MW18] Xuna Miao and Xiaobo Wu. Virtual machine anomaly detection strategy based on cloud platform operating environment perception. Concurrency and Computation: Practice and Experience, 30(22):e4656:1–e4656:??, November 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Missier:2016:PDD

[MWHW16] Paolo Missier, Simon Woodman, Hugo Hiden, and Paul Watson. Provenance and data differencing for workflow reproducibility analysis. Concurrency and Computation: Practice and Experience, 28(4):995–1015, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mangler:2010:MGS

[MWJ^{+}10] J. Mangler, C. Witzany, O. Jorns, E. Schikuta, H. Wanek, and I. Ul Haq. Mobile gSET — secure business workflows for Mobile-Grid clients. Concurrency and Computation: Practice and Experience, 22(14):2036–2051, September 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2013:SIP

[MWL^{+}13] Yan Ma, Lizhe Wang, Dingsheng Liu, Tao Yuan, Peng Liu, and Wanfeng Zhang. Special issue papers: Distributed data structure templates for data-intensive remote sensing applications. Concurrency and Computation: Practice and Experience, 25(12):1784–1797, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miao:2015:WBS

[MWL^{+}15] Yanhua Miao, Lizhe Wang, Dingsheng Liu, Yan Ma, Wanfeng Zhang, and Lajiao Chen. A Web 2.0-based science gateway for massive remote sensing image processing. Concurrency and Computation: Practice and Experience, 27(9):2489–2501, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mu:2018:FRM

[MWL18] Yashuang Mu, Lidong Wang, and Xiaodong Liu. A fast rank mutual information based decision tree and its implementation via Map-Reduce. Concurrency and Computation: Practice and Experience, 30(10):??, May 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4387.

Myre:2011:PAS

[MWLS11] J. Myre, S. D. C. Walsh, D. Lilja, and M. O. Saar. Performance analysis of single-phase, multiphase, and multicomponent lattice-Boltzmann fluid flow simulations on GPU clusters. Concurrency and Computation: Practice and Experience, 23(4):332–350, March 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2015:GSE

[MWPL15] Xingkong Ma, Yijie Wang, Xiaoqiang Pei, and Xiaoyong Li. A general scalable and elastic matching service for contentbased publish/subscribe systems. Concurrency and Computation: Practice and Experience, 27(1):94–118, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2017:CAP

[MWPX17] Xingkong Ma, Yijie Wang, Xiaoqiang Pei, and Fangliang Xu. A cloud-assisted publish/subscribe service for time-critical dissemination of bulk content. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martin:2018:HWP

[MWRK18] Steven Martin, Cary Whitney, David Rush, and Matthew Kappel. How to write a plugin to export job, power, energy, and system environmental data from your Cray XC system. Concurrency and Computation: Practice and Experience, 30 (1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohr:2010:PMA

[MWW10] B. Mohr, B. J. N. Wylie, and F. Wolf. Performance measurement and analysis tools for extremely scalable systems. Concurrency and Computation: Practice and Experience, 22 (16):2212–2229, November 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2017:SBL

[MY17] Kun Ma and Bo Yang. Stream-based live data replication approach of in-memory cache. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McCann:2006:SEV

[MYDM06] Karen M. McCann, Maurice Yarrow, Adrian DeVivo, and Piyush Mehrotra. ScyFlow: an environment for the visual specification and execution of scientific workflows. Concurrency and Computation: Practice and Experience, 18(10): 1155–1167, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mei:2019:SIM

[MYS19] Lin Mei, Zhiguo Yan, and Jie Shao. Special issue on multimodal information learning and analytics on cross-media big data. Concurrency and Computation: Practice and Experience, 31(9):e5144:1–e5144:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2018:OSB

[MYY18] Kun Ma, Bo Yang, and Ziqiang Yu. Optimization of streambased live data migration strategy in the cloud. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4293.

Mamei:2006:SMD

[MZ06] Marco Mamei and Franco Zambonelli. Self-maintained distributed tuples for field-based coordination in dynamic networks. Concurrency and Computation: Practice and Experience, 18(4):427–443, April 10, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moussaoui:2019:PFD

[MZA19] Mohamed Moussaoui, Montaceur Zaghdoud, and Jalel Akaichi. A possibilistic framework for the detection of terrorism-related Twitter communities in social media. Concurrency and Computation: Practice and Experience, 31(13): e5077:1–e5077:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2019:UCD

[MZG19] Xi’na Ma, Jingyuan Zhao, and Peng Guo. The urban computing on the distribution of inhalable particulate matters to Smart City-based residential groups. Concurrency and Computation: Practice and Experience, 31(9):e4803:1–e4803:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2019:ESJ

[MZJ^{+}19] Youzhong Ma, Ruiling Zhang, Shijie Jia, Yongxin Zhang, and Xiaofeng Meng. An efficient similarity join approach on large-scale high-dimensional data using random projection. Concurrency and Computation: Practice and Experience, 31 (20):e5303:1–e5303:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muhlbauer:2016:PZS

[MZK16] Adam Muhlbauer, Timothy Zelinsky, and Salil S. Kanhere. Platform zero: a step closer to ubiquitous computing. Concurrency and Computation: Practice and Experience, 28(3): 688–706, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2010:SSE

[MZS^{+}10] Hong-Sheng Ma, Yong Zheng, Zhi-Gang Shao, Chang-Sheng Jiang, Long-Quan Zhou, and Guo-Min Zhang. Simulation on seismogenic environment of strong earthquakes in Sichuan– Yunnan region, China. Concurrency and Computation: Practice and Experience, 22(12):1626–1643, August 25, 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mao:2016:GBI

[MZW^{+}16] Yuxin Mao, Ping Zhu, Guiyi Wei, Mohammad Mehedi Hassan, and M. Anwar Hossain. A game-based incentive model for service cooperation in VANETs. Concurrency and Computation: Practice and Experience, 28(3):674–687, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Neves:2015:SLP

[NA15] Samuel Neves and Filipe Araujo. Straight-line programs for fast sparse matrix–vector multiplication. Concurrency and Computation: Practice and Experience, 27(13):3245–3261, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nagel:2010:Pa

[Nag10] Wolfgang E. Nagel. Preface. Concurrency and Computation: Practice and Experience, 22(2):143–144, February 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nakajima:2002:PMI

[Nak02] Kengo Nakajima. Parallel multilevel iterative linear solvers with unstructured adaptive grids for simulations in earth science. Concurrency and Computation: Practice and Experience, 14(6–7):483–498, May/June 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/94515749/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=94515749{\&}PLACEBO=
IE.pdf.

Nguyen:2015:WIS

[NAK^{+}15] Hoang Anh Nguyen, David Abramson, Timoleon Kipouros, Andrew Janke, and Graham Galloway. WorkWays: interacting with scientific workflows. Concurrency and Computation: Practice and Experience, 27(16):4377–4397, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Namasudra:2019:IAB

[Nam19] Suyel Namasudra. An improved attribute-based encryption technique towards the data security in cloud computing. Concurrency and Computation: Practice and Experience, 31(3): e4364:1–e4364:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nacar:2007:VCG

[NAP^{+}07] Mehmet A. Nacar, Mehmet S. Aktas, Marlon Pierce, Zhenyu Lu, Gordon Erlebacher, Dan Kigelman, Evan F. Bollig, Cesar R. S. da Silva, Benny Sowell, and David A. Yuen. VLab: collaborative Grid services and portals to support computational material science. Concurrency and Computation: Practice and Experience, 19(12):1717–1728, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Narasimhan:2005:SIF

[Nar05] Priya Narasimhan. Special issue: Foundations of middleware technologies. Concurrency and Computation: Practice and Experience, 17(12):1469–1470, October 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Netto:2012:CRB

[NB12] Marco A. S. Netto and Rajkumar Buyya. Coordinated rescheduling of Bag-of-Tasks for executions on multiple resource providers. Concurrency and Computation: Practice and Experience, 24(12):1362–1376, August 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Neary:2005:AES

[NC05] Michael O. Neary and Peter Cappello. Advanced eager scheduling for Java-based adaptive parallel computing. Concurrency and Computation: Practice and Experience, 17(7– 8):797–819, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nicholson:2008:DDR

[NCD^{+}08] C. Nicholson, D. G. Cameron, A. T. Doyle, A. P. Millar, and K. Stockinger. Dynamic data replication in LCG 2008. Concurrency and Computation: Practice and Experience, 20 (11):1259–1271, August 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Natrajan:2004:SPF

[NCWD^{+}04] Anand Natrajan, Michael Crowley, Nancy Wilkins-Diehr, Marty A. Humphrey, Anthony D. Fox, Andrew S. Grimshaw, and Charles L. Brooks III. Studying protein folding on the Grid: experiences using CHARMM on NPACI resources under Legion. Concurrency and Computation: Practice and Experience, 16(4):385–397, April 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Natarajan:2017:ADC

[NDL17] Ekanathan Palamadai Natarajan, Maryam Mehri Dehnavi, and Charles Leiserson. Autotuning divide-and-conquer stencil computations. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Narasimhan:2005:MSR

[NDP^{+}05] P. Narasimhan, T. A. Dumitraş, A. M. Paulos, S. M. Pertet, C. F. Reverte, J. G. Slember, and D. Srivastava. MEAD: support for Real-Time Fault-Tolerant CORBA. Concurrency and Computation: Practice and Experience, 17(12):1527–1545, October 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Niu:2016:GVC

[NDT^{+}16] Qingpeng Niu, James Dinan, Sravya Tirukkovalur, Anouar Benali, Jeongnim Kim, Lubos Mitas, Lucas Wagner, and P. Sadayappan. Global-view coefficients: a data management solution for parallel quantum Monte Carlo applications. Concurrency and Computation: Practice and Experience, 28(13): 3655–3671, September 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nelson:2005:SIP

[Nel05] David B. Nelson. Special issue: Performance issues in computer architecture and design. Concurrency and Computation: Practice and Experience, 17(10):1215–1218, August 25, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Neves:2017:EFE

[Nev17] Diogo Telmo Neves. EPIC: a framework to exploit parallelism in irregular codes. Concurrency and Computation: Practice and Experience, 29(2):??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Niu:2019:TTC

[NFG19] Ke Niu, Weining Fang, and Beiyuan Guo. A team task complexity measure for emergency procedures in fully automatic metro. Concurrency and Computation: Practice and Experience, 31(10):e4753:1–e4753:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nakasan:2017:SMO

[NIIU17] Chawanat Nakasan, Kohei Ichikawa, Hajimu Iida, and Putchong Uthayopas. A simple multipath OpenFlow controller using topology-based algorithm for multipath TCP. Concurrency and Computation: Practice and Experience, 29 (13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nishigaki:2018:HIS

[Nis18] Masakatsu Nishigaki. Humanics information security. Concurrency and Computation: Practice and Experience, 30(2): ??, January 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nudd:2005:PBM

[NJ05] G. R. Nudd and S. A. Jarvis. Performance-based middleware for Grid computing. Concurrency and Computation: Practice and Experience, 17(2–4):215–234, February/April 2005.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nyers:2015:CSS

[NJ15] Lehel Nyers and Márk Jelasity. A comparative study of spanning tree and gossip protocols for aggregation. Concurrency and Computation: Practice and Experience, 27(16): 4091–4106, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Njah:2019:DBN

[NJM19] Hasna Njah, Salma Jamoussi, and Walid Mahdi. Deep Bayesian network architecture for Big Data mining. Concurrency and Computation: Practice and Experience, 31(2): e4418:1–e4418:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ng:2019:SMC

[NJZZ19] Chee Keong Ng, Frank Jiang, Leo Yu Zhang, and Wanlei Zhou. Static malware clustering using enhanced deep embedding method. Concurrency and Computation: Practice and Experience, 31(19):e5234:1–e5234:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

N:2019:EEL

[NK19] Suresh Kumar N and Paramasivam K. Energy efficient lowpower full-adder by 65 nm CMOS technology in ALU. Concurrency and Computation: Practice and Experience, 31(12): e4741:1–e4741:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nebbou:2019:ULS

[NLFA19] Tawfiq Nebbou, Mohamed Lehsaini, Hacène Fouchal, and Marwane Ayaida. An urban location service for vehicular area networks. Concurrency and Computation: Practice and Experience, 31(24):e4693:1–e4693:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ning:2012:DCA

[NLYZ12] Huansheng Ning, Hong Liu, Laurence T. Yang, and Yan Zhang. Dual cryptography authentication protocol and its security analysis for radio frequency identification systems. Concurrency and Computation: Practice and Experience, 24 (17):2040–2054, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nagel:2010:Pb

[NM10] Wolfgang E. Nagel and Matthias S. Muller. Preface. Concurrency and Computation: Practice and Experience, 22(16): 2195, November 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nelisse:2003:COB

[NMKB03] Arnold Nelisse, Jason Maassen, Thilo Kielmann, and Henri E. Bal. CCJ: object-based message passing and collective communication in Java. Concurrency and Computation: Practice and Experience, 15(3–5):341–369, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nataraj:2010:FSP

[NMM^{+}10] Aroon Nataraj, Allen D. Malony, Alan Morris, Dorian C. Arnold, and Barton P. Miller. A framework for scalable, parallel performance monitoring. Concurrency and Computation: Practice and Experience, 22(6):720–735, April 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Narasimhan:2001:IJR

[NMMS01] N. Narasimhan, L. E. Moser, and P. M. Melliar-Smith. Interceptors for Java Remote Method Invocation. Concurrency and Computation: Practice and Experience, 13(8–9): 755–774, July/August 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/84503222/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=84503222&PLACEBO=IE.pdf.

Nishiyama:2007:FSQ

[NN07] Hiroyasu Nishiyama and Kei Nakajima. Fast synchronization for quasi-immutable objects. Concurrency and Computation: Practice and Experience, 19(18):2353–2367, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nagy:2014:AUF

[NNH^{+}14] Zoltán Nagy, Csaba Nemes, Antal Hiba, Árpád Csík, András Kiss, Miklós Ruszinkó, and Péter Szolgay. Accelerating unstructured finite volume computations on field-programmable gate arrays. Concurrency and Computation: Practice and Experience, 26(3):615–643, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nassif:2007:JCP

[NNK^{+}07] Lilian Noronha Nassif, José Marcos Nogueira, Ahmed Karmouch, Mohamed Ahmed, and Flávio Vinícius de Andrade. Job completion prediction using case-based reasoning for Grid computing environments. Concurrency and Computation: Practice and Experience, 19(9):1253–1269, June 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nishitani:2002:TCI

[NNON02] Yasunori Nishitani, Kiyoshi Negishi, Hiroshi Ohta, and Eiji Nunohiro. Techniques for compiling and implementing all NAS parallel benchmarks in HPF. Concurrency and Computation: Practice and Experience, 14(8–9):769–787, July/ August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016133/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016133{\&}PLACEBO=IE.pdf.

Natrajan:2002:LGP

[NNTH^{+}02] Anand Natrajan, Anh Nguyen-Tuong, Marty A. Humphrey, Michael Herrick, Brian P. Clarke, and Andrew S. Grimshaw. The Legion Grid Portal. Concurrency and Computation: Practice and Experience, 14(13–15):1365–1394, November/ December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nassif:2009:RSG

[NNvVdA09] Lilian Noronha Nassif, José Marcos Nogueira, and Flá vio Vinícius de Andrade. Resource selection in grid: a taxonomy and a new system based on decision theory, case-based reasoning, and fine-grain policies. Concurrency and Computation: Practice and Experience, 21(3):337–355, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nakajima:2002:PIS

[NO02] Kengo Nakajima and Hiroshi Okuda. Parallel iterative solvers for unstructured grids using a directive/MPI hybrid programming model for the GeoFEM platform on SMP cluster architectures. Concurrency and Computation: Practice and Experience, 14(6–7):411–429, May/ June 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/94515747/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
94515747{\&}PLACEBO=IE.pdf.

Noble:2008:GMY

[Nob08] Michael S. Noble. Getting more from your multicore: exploiting OpenMP from an open-source numerical scripting language. Concurrency and Computation: Practice and Experience, 20(16):1877–1891, November 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Notare:2016:ENA

[Not16a] Mirela Sechi Moretti Annoni Notare. Editorials: New advances in large-scale distributed simulation and real-time applications. Concurrency and Computation: Practice and Experience, 28(12):3257–3259, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Notare:2016:EWM

[Not16b] Mirela Sechi Moretti Annoni Notare. Editorials: Wireless and mobile network performance evaluation. Concurrency and Computation: Practice and Experience, 28(6):1767–1768, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Novotny:2002:GPD

[Nov02] Jason Novotny. The Grid Portal Development Kit. Concurrency and Computation: Practice and Experience, 14(13–15): 1129–1144, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nasciutti:2019:EOR

[NPL19] Thiago Carrijo Nasciutti, Jairo Panetta, and Pedro Pais Lopes. Evaluating optimizations that reduce global memory accesses of stencil computations in GPGPUs. Concurrency and Computation: Practice and Experience, 31(18):e4929:1– e4929:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nanda:2019:ESI

[NPM19] Priyadarsi Nanda, Deepak Puthal, and Saraju P. Mohanty. Editorial to the special issue on recent advances on trust, security and privacy in computing and communications. Concurrency and Computation: Practice and Experience, 31(23): e5431:1–e5431:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Negri:2006:DGT

[NPTT06] A. Negri, A. Poggi, M. Tomaiuolo, and P. Turci. Dynamic Grid tasks composition and distribution through agents. Concurrency and Computation: Practice and Experience, 18(8): 875–885, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Niu:2017:DSD

[NQL^{+}17] Bingxin Niu, Heng Qi, Keqiu Li, Xiulong Liu, and Weilian Xue. Dynamic scheming the duty cycle in the opportunistic routing sensor network. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nussbaum:2008:LES

[NR08] Lucas Nussbaum and Olivier Richard. Lightweight emulation to study peer-to-peer systems. Concurrency and Computation: Practice and Experience, 20(6):735–749, April 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Namasudra:2017:NSA

[NR17] Suyel Namasudra and Pinki Roy. A new secure authentication scheme for cloud computing environment. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Neelima:2015:CCO

[NRR15] B. Neelima, G. Ram Mohana Reddy, and Prakash S. Raghavendra. Communication and computation optimization of concurrent kernels using kernel coalesce on a GPU. Concurrency and Computation: Practice and Experience, 27 (1):47–68, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Novotny:2004:GPF

[NRW04] Jason Novotny, Michael Russell, and Oliver Wehrens. Grid-Sphere: a portal framework for building collaborations. Concurrency and Computation: Practice and Experience, 16(5): 503–513, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nirmala:2019:AEF

[NS19] N. Nirmala and S. Sumathi. An area-efficient FFT processor for the OFDMA transceiver communication system. Concurrency and Computation: Practice and Experience, 31(14): e4875:1–e4875:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nascimento:2007:DDS

[NSBR07] Aline P. Nascimento, Alexandre C. Sena, Cristina Boeres, and Vinod E. F. Rebello. Distributed and dynamic self-scheduling of parallel MPI Grid applications. Concurrency and Computation: Practice and Experience, 19(14):1955–1974, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Neelima:2017:KGA

[NSN^{+}17] B. Neelima, Bharath Shamsundar, Anjjan Narayan, Rithesh Prabhu, and Crystal Gomes. Kepler GPU accelerated recursive sorting using dynamic parallelism. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Niewiadomska-Szynkiewicz:2013:SIP

[NSSAK13] Ewa Niewiadomska-Szynkiewicz, Andrzej Sikora, Piotr Arabas, and Joanna Kołodziej. Special issue papers: Control system for reducing energy consumption in backbone computer network. Concurrency and Computation: Practice and Experience, 25(12):1738–1754, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See erratum [NSSAK16].

Niewiadomska-Szynkiewicz:2016:ECS

[NSSAK16] Ewa Niewiadomska-Szynkiewicz, Andrzej Sikora, Piotr Arabas, and Joanna Kołodziej. Erratum: Control system for reducing energy consumption in backbone computer network. Concurrency and Computation: Practice and Experience, 28 (18):4557, December 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [NSSAK13].

Nystrom:2008:HLD

[NTK08] J. H. Nyström, P. W. Trinder, and D. J. King. High-level distribution for the rapid production of robust telecoms software: comparing C++ and ERLANG. Concurrency and Computation: Practice and Experience, 20(8):941–968, June 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ngubiri:2009:MFP

[NvV09] John Ngubiri and Mario van Vliet. A metric of fairness for parallel job schedulers. Concurrency and Computation: Practice and Experience, 21(12):1525–1546, August 25, 2009. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nakanishi:2011:IHK

[NZKK11] T. Nakanishi, K. Zettsu, Y. Kidawara, and Y. Kiyoki. Interconnection of heterogeneous knowledge bases and its application on Knowledge Grid. Concurrency and Computation: Practice and Experience, 23(9):940–955, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Okuda:2002:OEE

[OA02] Hiroshi Okuda and Norihisa Anan. Optimization of elementby-element FEM in HPF 1.1. Concurrency and Computation: Practice and Experience, 14(8–9):647–663, July/ August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016128/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016128{\&}PLACEBO=IE.pdf.

Oryspayev:2015:PAD

[OAS^{+}15] Dossay Oryspayev, Hasan Metin Aktulga, Masha Sosonkina, Pieter Maris, and James P. Vary. Performance analysis of distributed symmetric sparse matrix vector multiplication algorithm for multi-core architectures. Concurrency and Computation: Practice and Experience, 27(17):5019–5036, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ornelas:2018:PDD

[OBD^{+}18] Tatiane Ornelas, Regina Braga, José Maria N. David, Fernanda Campos, and Gabriella Castro. Provenance data discovery through Semantic Web resources. Concurrency and Computation: Practice and Experience, 30(6):??, March 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4366.

Oliker:2005:PES

[OCC^{+}05] Leonid Oliker, Andrew Canning, Jonathan Carter, John Shalf, David Skinner, Stéphane Ethier, Rupak Biswas, Jahed Djomehri, and Rob Van der Wijngaart. Performance evaluation of the SX-6 vector architecture for scientific computations. Concurrency and Computation: Practice and Experience, 17(1):69–93, January 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oakasha:2001:CMO

[OCS01] H. Oakasha, S. Conrad, and G. Saake. Consistency management in object-oriented databases. Concurrency and Computation: Practice and Experience, 13(11):955–985, September 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/85012149/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
85012149&PLACEBO=IE.pdf.

Ogasawara:2013:SIP

[ODS^{+}13] Eduardo Ogasawara, Jonas Dias, Vitor Silva, Fernando Chirigati, Daniel de Oliveira, Fabio Porto, Patrick Valduriez, and Marta Mattoso. Special issue papers: Chiron: a parallel engine for algebraic scientific workflows. Concurrency and Computation: Practice and Experience, 25(16):2327–2341, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Orgun:2015:ESI

[OEP^{+}15] Mehmet A. Orgun, Atilla Elçi, Josef Pieprzyk, Alexander Chefranov, Rajan Shankaran, and Huaxiong Wang. Editorials: Special issue on trust and security in wireless sensor networks. Concurrency and Computation: Practice and Experience, 27(15):3791–3793, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Otero:2017:PAM

[OFR^{+}17] Beatriz Otero, Jorge Francés, Robert Rodriguez, Otilio Rojas, Freysimar Solano, and Juan Guevara-Jordan. A performance analysis of a mimetic finite difference scheme for acoustic wave propagation on GPU platforms. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oliver:2001:SEE

[OGA^{+}01] José Oliver, Jordi Guitart, Eduard Ayguadé, Nacho Navarro, and Jordi Torres. Strategies for the efficient exploitation of loop-level parallelism in Java. Concurrency and Computation: Practice and Experience, 13(8–9):663–680, July/ August 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/84503218/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
84503218&PLACEBO=IE.pdf.

Oinn:2006:TLC

[OGA^{+}06] Tom Oinn, Mark Greenwood, Matthew Addis, M. Nedim Alpdemir, Justin Ferris, Kevin Glover, Carole Goble, Antoon Goderis, Duncan Hull, Darren Marvin, Peter Li, Phillip Lord, Matthew R. Pocock, Martin Senger, Robert Stevens, Anil Wipat, and Chris Wroe. Taverna: lessons in creating a workflow environment for the life sciences. Concurrency and Computation: Practice and Experience, 18(10):1067–1100, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogino:2002:TDG

[Ogi02] Tatsuki Ogino. Three-dimensional global MHD simulation code for the Earth’s magnetosphere using HPF/JA. Concurrency and Computation: Practice and Experience, 14(8–9): 631–646, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/95016136/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=95016136{\&}PLACEBO=IE.pdf.

Ormandi:2013:GLL

[OHJ13] Róbert Ormándi, István Hegedűs, and Márk Jelasity. Gossip learning with linear models on fully distributed data. Concurrency and Computation: Practice and Experience, 25(4): 556–571, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogawa:2007:GAF

[OISS07] Hirotaka Ogawa, Satoshi Itoh, Tetsuya Sonoda, and Satoshi Sekiguchi. GridASP: an ASP framework for Grid utility computing. Concurrency and Computation: Practice and Experience, 19(6):885–891, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Olsson:2015:UAR

[OK15] Ronald A. Olsson and Aaron W. Keen. User accessible reply capabilities in invoking and servicing operations. Concurrency and Computation: Practice and Experience, 27(17): 5134–5155, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2018:EBI

[OK18] Marek R. Ogiela and Hoon Ko. Editorial: Bio-inspired and cognitive approaches in cryptography and security applications. Concurrency and Computation: Practice and Experience, 30(2):??, January 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oualhaj:2019:MDT

[OKBO19] Omar Ait Oualhaj, Abdellatif Kobbane, and Jalel Ben-Othman. Mobile delay-tolerant networks with energyharvesting and wireless energy transfer cooperation. Concurrency and Computation: Practice and Experience, 31(24): e5112:1–e5112:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Olabarriaga:2018:SIC

[OKG18] Sílvia Delgado Olabarriaga, Dagmar Krefting, and Tristan Glatard. Special isssue of the CCGrid-Life workshop 2017. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4520.

Olsen-Kettle:2010:MDS

[OKM10] L. M. Olsen-Kettle and H. B. Mühlhaus. Mesh dependence and slip complexity in earthquake fault models. Concurrency and Computation: Practice and Experience, 22(12): 1653–1664, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2018:AKB

[OKO18] Marek R. Ogiela, Natalia Krzyworzeka, and Lidia Ogiela. Application of knowledge-based cognitive CAPTCHA in Cloud of Things security. Concurrency and Computation: Practice and Experience, 30(21):e4769:1–e4769:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ostermann:2016:MLS

[OKP16] Simon Ostermann, Gabor Kecskemeti, and Radu Prodan. Multi-layered simulations at the heart of workflow enactment on clouds. Concurrency and Computation: Practice and Experience, 28(11):3180–3201, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Olsson:2015:TER

[OKW15] Ronald A. Olsson, Aaron W. Keen, and Todd Williamson. Transformations for early reply and forward message passing mechanisms. Concurrency and Computation: Practice and Experience, 27(17):4645–4658, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Olsson:2018:RDJ

[OKW18] Ronald A. Olsson, Aaron W. Keen, and Todd Williamson. Reducing distributed JR program start-up time via extending JR’s operation abstraction. Concurrency and Computation: Practice and Experience, 30(14):??, July 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4428.

Ortega:2015:PRH

[OLG^{+}15] Gloria Ortega, Julia Lobera, Inmaculada García, M. Pilar Arroyo, and Ester M. Garzón. Parallel resolution of the 3D Helmholtz equation based on multi-graphics processing unit clusters. Concurrency and Computation: Practice and Experience, 27(13):3205–3219, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ossowski:2006:CSD

[OM06a] Sascha Ossowski and Ronaldo Menezes. On coordination and its significance to distributed and multi-agent systems. Concurrency and Computation: Practice and Experience, 18(4): 359–370, April 10, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ossowski:2006:SIC

[OM06b] Sascha Ossowski and Ronaldo Menezes. Special issue: Coordination models and systems. Concurrency and Computation: Practice and Experience, 18(4):357, April 10, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oguma:2001:LBL

[ON01] Hisashi Oguma and Yasuichi Nakayama. Lesser Bear — a light-weight process library for SMP computers. Concurrency and Computation: Practice and Experience, 13(12): 1107–1120, October 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/85512008/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=85512008&PLACEBO=IE.pdf.

Oguma:2002:LBL

[ON02] Hisashi Oguma and Yasuichi Nakayama. Lesser Bear: a lightweight process library for SMP computers — scheduling mechanism without a lock operation. Concurrency and Computation: Practice and Experience, 14(10):841–857, August 25, 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/97517966/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
97517966{\&}PLACEBO=IE.pdf.

Ogiela:2018:LTC

[OO18] Urszula Ogiela and Lidia Ogiela. Linguistic techniques for cryptographic data sharing algorithms. Concurrency and Computation: Practice and Experience, 30(3):??, February 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4275.

Ortin-Obon:2014:CSO

[OORVB14] Marta Ortín-Obón, Luca Ramini, Víctor Viñals, and Davide Bertozzi. Capturing the sensitivity of optical network quality metrics to its network interface parameters. Concurrency and Computation: Practice and Experience, 26(15): 2504–2517, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Och:2001:CDM

[OOTK01] Christian Och, Richard M. Osborne, John Todd, and Roger King. The COIL data mediator definition language. Concurrency and Computation: Practice and Experience, 13(7):543– 577, June 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/83002175/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
83002175&PLACEBO=IE.pdf.

Olson:2015:LIS

[ORDG15] Andrew M. Olson, Rajeev R. Raje, Barun Devaraju, and Lahiru S. Gallege. Learning improves service discovery. Concurrency and Computation: Practice and Experience, 27(7): 1679–1694, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oliveira:2013:DSM

[ORdSL13] Luís M. L. Oliveira, Joel J. P. C. Rodrigues, Amaro F. de Sousa, and Jaime Lloret. Denial of service mitigation approach for IPv6-enabled smart object networks. Concurrency and Computation: Practice and Experience, 25(1):129–142, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

OBoyle:2009:OPK

[OS09] Michael O’Boyle and Henk Sips. Obituary: Peter Knijnenburg (1961–2007). Concurrency and Computation: Practice and Experience, 21(1):5, January 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

ORyan:2001:EPM

[OSK^{+}01] Carlos O’Ryan, Douglas C. Schmidt, Fred Kuhns, Marina Spivak, Jeff Parsons, Irfan Pyarali, and David L. Levine. Evaluating policies and mechanisms to support distributed real-time applications with CORBA. Concurrency and Computation: Practice and Experience, 13(7):507–541, June 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/83002173/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
83002173&PLACEBO=IE.pdf.

Oleksiak:2007:HES

[OTG^{+}07] Ariel Oleksiak, Alisdair Tullo, Paul Graham, Tomasz Kuczyński, Jarek Nabrzyski, Dawid Szejnfeld, and Terry Sloan. HPC-Europa single point of access as a framework for building science gateways. Concurrency and Computation: Practice and Experience, 19(6):851–866, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2018:VCA

[OTO18] Urszula Ogiela, Makoto Takizawa, and Lidia Ogiela. Visual CAPTCHA application in linguistic cryptography. Concurrency and Computation: Practice and Experience, 30(2): ??, January 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oz:2019:UAT

[OTT19] Isil Oz, Haluk Rahmi Topcuoglu, and Oguz Tosun. A userassisted thread-level vulnerability assessment tool. Concurrency and Computation: Practice and Experience, 31(13): e5085:1–e5085:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ouaftouh:2019:SRU

[OZI19] Sara Ouaftouh, Ahmed Zellou, and Ali Idri. Social recommendation: a user profile clustering-based approach. Concurrency and Computation: Practice and Experience, 31(20):e5330:1– e5330:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Powers:2008:HPP

[PA08] Franklin E. Powers, Jr. and Gita Alaghband. The Hydra Parallel Programming System. Concurrency and Computation: Practice and Experience, 20(1):1–27, January 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perez:2018:ELR

[PA18] Daniel Hugo Cámpora Pérez and Omar Awile. An efficient low-rank Kalman filter for modern SIMD architectures. Concurrency and Computation: Practice and Experience, 30(23): e4483:1–e4483:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pacher:2016:TLE

[Pac16] Mathias Pacher. Two-level extensions of an artificial hormone system. Concurrency and Computation: Practice and Experience, 28(14):3730–3750, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Popov:2017:PHA

[PAC^{+}17] Mihail Popov, Chadi Akel, Yohan Chatelain, William Jalby, and Pablo de Oliveira Castro. Piecewise holistic autotuning of parallel programs with CERE. Concurrency and Computation: Practice and Experience, 29(15):??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pereira:2017:SBC

[PAdS^{+}17] Phillipe Pereira, Higo Albuquerque, Isabela da Silva, Hendrio Marques, Felipe Monteiro, Ricardo Ferreira, and Lucas Cordeiro. SMT-based context-bounded model checking for CUDA programs. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Picone:2015:CGR

[PAM^{+}15] Marco Picone, Michele Amoretti, Marco Martalò, Francesco Zanichelli, and Gianluigi Ferrari. Combining geo-referencing and network coding for distributed large-scale information management. Concurrency and Computation: Practice and Experience, 27(13):3295–3315, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pan:2020:RME

[Pan20] Xuchao Pan. Retracted: The mechanism and experimental study on the interference of high voltage lines to navigation system. Concurrency and Computation: Practice and Experience, 32(2):e5513:1–e5513:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [FPHZ19].

Parashar:2002:SIS

[Par02] Manish Parashar. Special issue: Software architectures for scientific applications. Concurrency and Computation: Practice and Experience, 14(5):321–322, April 25, 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/93513495/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
93513495{\&}PLACEBO=IE.pdf.

Paton:2008:ADM

[Pat08] N. W. Paton. Autonomics and data management. Concurrency and Computation: Practice and Experience, 20(17): 2075–2088, December 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parkin:2007:PCC

[PB07a] M. Parkin and J. M. Brooke. A PDA client for the computational Grid. Concurrency and Computation: Practice and Experience, 19(9):1317–1331, June 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierson:2007:SIS

[PB07b] Jean-Marc Pierson and Lionel Brunie. Special issue: Selection of best papers of the VLDB Data Management in Grids Workshop (VLDB DMG 2006). Concurrency and Computation: Practice and Experience, 19(16):2105–2107, November 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pacher:2012:IES

[PB12] Mathias Pacher and Uwe Brinkschulte. Implementation and evaluation of a self-organizing artificial hormone system to assign time-dependent tasks. Concurrency and Computation: Practice and Experience, 24(16):1879–1902, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pokahr:2016:ECB

[PB16] Alexander Pokahr and Lars Braubach. Elastic componentbased applications in PaaS clouds. Concurrency and Computation: Practice and Experience, 28(4):1368–1384, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

P:2019:AIN

[PB19a] Nivash J P and L. D. Dhinesh Babu. Analyzing the impact of news trends on research publications and scientific collaboration networks. Concurrency and Computation: Practice and Experience, 31(14):e5058:1–e5058:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Poongodi:2019:ACS

[PB19b] S. Poongodi and M. Rajesh Babu. Analysis of crop suitability using clustering technique in Coimbatore region of Tamil Nadu. Concurrency and Computation: Practice and Experience, 31(14):e5294:1–e5294:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pani:2015:CCT

[PBD^{+}15] Danilo Pani, Gianluca Barabino, Alessia Dessì, Selene Uras, and Luigi Raffo. The challenge of collaborative telerehabilitation: conception and evaluation of a telehealth system enhancement for home-therapy follow-up. Concurrency and Computation: Practice and Experience, 27(11):2889–2906, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Podobas:2015:CPS

[PBF15] Artur Podobas, Mats Brorsson, and Karl-Filip Faxén. A comparative performance study of common and popular taskcentric programming frameworks. Concurrency and Computation: Practice and Experience, 27(1):1–28, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pourghaffari:2019:EMA

[PBK19] Ali Pourghaffari, Morteza Barari, and Saeed Sedighian Kashi. An efficient method for allocating resources in a cloud computing environment with a load balancing approach. Concurrency and Computation: Practice and Experience, 31(17): e5285:1–e5285:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pokam:2004:SRP

[PBSB04] Gilles Pokam, Stéphane Bihan, Julien Simonnet, and François Bodin. SWARP: a retargetable preprocessor for multimedia instructions. Concurrency and Computation: Practice and Experience, 16(2–3):303–318, February/March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Penmatsa:2014:CMU

[PC14] Satish Penmatsa and Anthony T. Chronopoulos. Cost minimization in utility computing systems. Concurrency and Computation: Practice and Experience, 26(1):287–307, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Piccialli:2017:LBI

[PC17a] Francesco Piccialli and Angelo Chianese. A location-based IoT platform supporting the cultural heritage domain. Concurrency and Computation: Practice and Experience, 29(11): ??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Plale:2017:EPR

[PC17b] Beth Plale and Maio Chen. Editorial: Pacific Rim Applications and Grid Middleware Assembly (PRAGMA): International clouds for data science. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pickartz:2018:PCV

[PCB^{+}18] Simon Pickartz, Carsten Clauss, Jens Breitbart, Stefan Lankes, and Antonello Monti. Prospects and challenges of virtual machine migration in HPC. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4412.

Plessl:2017:EFS

[PCC17] Christian Plessl, Guojing Cong, and João M. P. Cardoso. Editorials: Foreword to the special issue of the 18th IEEE international conference on computational science and engineering (CSE2015). Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pop:2015:AMS

[PCD15] Florin Pop, Radu-Ioan Ciobanu, and Ciprian Dobre. Adaptive method to support social-based mobile networks using a PageRank approach. Concurrency and Computation: Practice and Experience, 27(8):1900–1912, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Penna:2017:DMW

[PCF^{+}17] Pedro H. Penna, Márcio Castro, Henrique C. Freitas, François Broquedis, and Jean-François Méhaut. Design methodology for workload-aware loop scheduling strategies based on genetic algorithm and simulation. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pelaez:2018:OSD

[PCGE18] Víctor Peláez, Antonio Campos, Daniel F. García, and Joaquín Entrialgo. Online scheduling of deadline-constrained bag-of-task workloads on hybrid clouds. Concurrency and Computation: Practice and Experience, 30(19):e4639:1– e4639:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Periorellis:2008:GIV

[PCH^{+}08] P. Periorellis, N. Cook, H. Hiden, A. Conlin, M. D. Hamilton, J. Wu, J. Bryans, X. Gong, F. Zhu, R. Smith, P. Watson, and A. R. Wright. GOLD infrastructure for virtual organizations. Concurrency and Computation: Practice and Experience, 20 (11):1273–1288, August 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Piccialli:2017:ECH

[PCJ17] Francesco Piccialli, Angelo Chianese, and Jason J. Jung. Editorial: Cultural heritage on Internet of Things (IoT) systems: Trends and challenges. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pokharel:2017:CAV

[PCL17] Shasi Pokharel, Kim-Kwang Raymond Choo, and Jixue Liu. Can Android VoIP voice conversations be decoded? I can eavesdrop on your Android VoIP communication. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pop:2012:BIC

[PCS^{+}12] Cristina Bianca Pop, Viorica Rozina Chifu, Ioan Salomie, Mihaela Dinsoreanu, Tudor David, Vlad Acretoaie, Aliz Nagy, and Ciprian Oprisa. Biologically-inspired clustering of semantic Web services. Birds or ants intelligence? Concurrency and Computation: Practice and Experience, 24(6):619–633, May 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Park:2018:SSI

[PCsHL18] Hyoungwoo Park, Buseung Cho, Il sun Hwang, and Jongsuk Ruth Lee. Study on the SDN-IP-based solution of wellknown bottleneck problems in private sector of national R&E network for big data transfer. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perez:2004:OBA

[PCT04] José A. Pérez, Rafael Corchuelo, and Miguel Toro. An orderbased algorithm for implementing multiparty synchronization. Concurrency and Computation: Practice and Experience, 16(12):1173–1206, October 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Padial-Collins:2004:POC

[PCVZ^{+}04] N. T. Padial-Collins, W. B. VanderHeyden, D. Z. Zhang, E. D. Dendy, and D. Livescu. Parallel operation of CartaBlanca on shared and distributed memory computers. Concurrency and Computation: Practice and Experience, 16 (1):61–77, January 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perrot:2016:OGB

[PDC16] Gilles Perrot, Stéphane Domas, and Raphaël Couturier. An optimized GPU-based 2D convolution implementation. Concurrency and Computation: Practice and Experience, 28(16): 4291–4304, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2017:ADS

[PDCA17] Jian Peng, Sam Detchon, Kim-Kwang Raymond Choo, and Helen Ashman. Astroturfing detection in social media: a binary n-gram–based approach. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Porto:2012:MTD

[PdCMdS^{+}12] Fabio Porto, Ana Maria de C. Moura, Frederico C. da Silva, Adriana Bassini, Daniele C. Palazzi, Maira Poltosi, Luis Eduardo Viveiros de Castro, and L. C. Cameron. A metaphoric trajectory data warehouse for Olympic athlete follow-up. Concurrency and Computation: Practice and Experience, 24 (13):1497–1512, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierson:2014:EIS

[PDD14] Jean-Marc Pierson, Lars Dittmann, and Georges Da Costa. Editorial: Introduction to special issue on selected papers from Energy Efficiency in Large-Scale Distributed Systems 2013 conference. Concurrency and Computation: Practice and Experience, 26(15):2567–2568, October 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pawliczek:2014:VED

[PDY14] Piotr Pawliczek, Witold Dzwinel, and David A. Yuen. Visual exploration of data by using multidimensional scaling on multicore CPU, GPU, and MPI cluster. Concurrency and Computation: Practice and Experience, 26(3):662–682, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pallickara:2012:EHD

[PF12] Shrideep Pallickara and Geoffrey Fox. Enabling hierarchical dissemination of streams in content distribution networks. Concurrency and Computation: Practice and Experience, 24 (14):1594–1606, September 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierce:2009:UWS

[PFC^{+}09] Marlon E. Pierce, Geoffrey C. Fox, Jong Y. Choi, Zhenhua Guo, Xiaoming Gao, and Yu Ma. Using Web 2.0 for scientific applications and scientific communities. Concurrency and Computation: Practice and Experience, 21(5):583–603, April 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Palmieri:2014:DAN

[PFC14] Francesco Palmieri, Ugo Fiore, and Aniello Castiglione. A distributed approach to network anomaly detection based on independent component analysis. Concurrency and Computation: Practice and Experience, 26(5):1113–1129, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pallickara:2005:PPG

[PFU^{+}05] Shrideep Pallickara, Geoffrey Fox, Ahmet Uyar, Hongbin Liu, Xi Rao, David Walker, and Beytullah Yildiz. Performance of a possible Grid message infrastructure. Concurrency and Computation: Practice and Experience, 17(2–4): 193–214, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Paul:2003:PCC

[PGB03] Himadri Sekhar Paul, Arobinda Gupta, and R. Badrinath. Performance comparison of checkpoint and recovery protocols. Concurrency and Computation: Practice and Experience, 15(15):1363–1386, December 25, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Penna:2019:CPE

[PGC^{+}19] Pedro Henrique Penna, Antônio Tadeu A. Gomes, Márcio Castro, Patricia D. M. Plentz, Henrique C. Freitas, François Broquedis, and Jean-François Méhaut. A comprehensive performance evaluation of the BinLPT workload-aware loop scheduler. Concurrency and Computation: Practice and Experience, 31(18):e5170:1–e5170:??, September 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pessoa:2018:GAB

[PGdCJ^{+}18] Tiago Carneiro Pessoa, Jan Gmys, Francisco Heron de Carvalho Júnior, Nouredine Melab, and Daniel Tuyttens. GPUaccelerated backtracking using CUDA Dynamic Parallelism. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4374.

Piernas:2019:LOD

[PGF19] Juan Piernas and Pilar González-Férez. Leveraging OSD+ devices for implementing a high-throughput parallel file system. Concurrency and Computation: Practice and Experience, 31(21):e4778:1–e4778:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pinheiro:2011:AFT

[PGK11] V. G. Pinheiro, A. Goldman, and F. Kon. Adaptive fault tolerance mechanisms for opportunistic environments: a mobile agent approach. Concurrency and Computation: Practice and Experience, 23(17):2078–2091, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Penaranda:2017:FTR

[PGL^{+}17] Roberto Peñaranda, María Engracia Gómez, Pedro López, Ernst Gunnar Gran, and Tor Skeie. A fault-tolerant routing strategy for k-ary n-direct s-indirect topologies based on intermediate nodes. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Porto:2004:CAM

[PGO^{+}04] F. Porto, G. Giraldi, J. C. Oliveira, R. Silva, and B. Schulze. CoDIMS: an adaptable middleware system for scientific visualization in Grids. Concurrency and Computation: Practice and Experience, 16(5):515–522, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierce:2010:QPS

[PGP^{+}10] Marlon E. Pierce, Xiaoming Gao, Sangmi L. Pallickara, Zhenhua Guo, and Geoffrey C. Fox. The Quakesim portal and services: new approaches to science gateway development techniques. Concurrency and Computation: Practice and Experience, 22(12):1732–1749, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Piro:2006:PSR

[PGW06] Rosario M. Piro, Andrea Guarise, and Albert Werbrouck. Price-sensitive resource brokering with the Hybrid Pricing Model and widely overlapping price domains. Concurrency and Computation: Practice and Experience, 18(8):837–850, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pouwelse:2008:TSB

[PGW^{+}08] J. A. Pouwelse, P. Garbacki, J. Wang, A. Bakker, J. Yang, A. Iosup, D. H. J. Epema, M. Reinders, M. R. van Steen, and H. J. Sips. TRIBLER: a social-based peer-to-peer system. Concurrency and Computation: Practice and Experience, 20 (2):127–138, February 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Playne:2012:CGA

[PH12] D. P. Playne and K. A. Hawick. Comparison of GPU architectures for asynchronous communication with finitedifferencing applications. Concurrency and Computation: Practice and Experience, 24(1):73–83, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pichel:2009:IDR

[PHCR09] J. C. Pichel, D. B. Heras, J. C. Cabaleiro, and F. F. Rivera. Increasing data reuse of sparse algebra codes on simultaneous multithreading architectures. Concurrency and Computation: Practice and Experience, 21(15):1838–1856, October 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perez:2010:SIG

[PHGK10] María S. Pérez, Pilar Herrero, Dennis Gannon, and Daniel S. Katz. Special issue: Grid computing, high performance and distributed application. Concurrency and Computation: Practice and Experience, 22(11):1335–1337, August 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Park:2018:GDC

[PHY^{+}18] Jeonghun Park, Seunggye Hwang, Janghoon Yang, Kitae Bae, Hoon Ko, and Dong Ku Kim. Generalized distributed compressive sensing with security challenges for linearly correlated information sources. Concurrency and Computation: Practice and Experience, 30(3):??, February 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4243.

Peinado:2012:SSD

[PIAH12] Jesus Peinado, Jacinto J. Ibáñez, Enrique Arias, and Vicente Hernández. Speeding up solving of differential matrix Riccati equations using GPGPU computing and MATLAB. Concurrency and Computation: Practice and Experience, 24(12): 1334–1348, August 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierce:2008:SIE

[Pie08] Marlon E. Pierce. Special issue editorial introduction: Grids and Geospatial Information Systems. Concurrency and Computation: Practice and Experience, 20(14):1611–1615, September 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pop:2016:EAM

[PIGK16] Florin Pop, Mauro Iacono, Marco Gribaudo, and Joanna Kołodziej. Editorials: Advances in modelling and simulation for big-data applications (AMSBA). Concurrency and Computation: Practice and Experience, 28(2):291–293, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pan:2004:PBC

[PIH04] Yi Pan, Constantinos S. Ierotheou, and Majeed M. Hayat. Parallel bandwidth characteristics calculations for thin avalanche photodiodes on a SGI Origin 2000 supercomputer. Concurrency and Computation: Practice and Experience, 16 (12):1207–1225, October 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Piccialli:2018:DFI

[PJ18] Francesco Piccialli and Jason J. Jung. Data fusion in the Internet of Data. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4700.

Pu:2014:RSI

[PJW^{+}14] Xiaoting Pu, Zhenhong Jia, Liejun Wang, Yingjie Hu, and Jie Yang. The remote sensing image enhancement based on nonsubsampled contourlet transform and unsharp masking. Concurrency and Computation: Practice and Experience, 26 (3):742–747, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierson:2008:SIS

[PK08] Jean-Marc Pierson and Harald Kosch. Special issue: Selection of best papers of the VLDB Data Management in Grids Workshop (VLDB DMG 2007). Concurrency and Computation: Practice and Experience, 20(17):1977–1979, December 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pirahandeh:2017:HPG

[PK17] Mehdi Pirahandeh and Deok-Hwan Kim. High performance GPU-based parity computing scheduler in storage applications. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Papadopoulos:2003:NRT

[PKB03] Philip M. Papadopoulos, Mason J. Katz, and Greg Bruno. NPACI Rocks: tools and techniques for easily deploying manageable Linux clusters. Concurrency and Computation: Practice and Experience, 15(7–8):707–725, June/July 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Phuong:2015:DSE

[PL15] Thi Yen Phuong and Jeong-Gun Lee. Design space exploration of SW beamformer on GPU. Concurrency and Computation: Practice and Experience, 27(7):1718–1733, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Plaza:2008:PPR

[Pla08] Antonio J. Plaza. Parallel processing of remotely sensed hyperspectral imagery: full-pixel versus mixed-pixel classification. Concurrency and Computation: Practice and Experience, 20(13):1539–1572, September 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pascual:2019:EAS

[PLC^{+}19] Jose A. Pascual, Joshua Lant, Caroline Concatto, Andrew Attwood, Javier Navaridas, Mikel Luján, and John Goodacre. On the effects of allocation strategies for exascale computing systems with distributed storage and unified interconnects. Concurrency and Computation: Practice and Experience, 31(21):e4784:1–e4784:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Park:2018:STQ

[PLJ18] Jae-Hong Park, O-Joun Lee, and Jai E. Jung. Spatiotemporal query contextualization for microtext retrieval in social media. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4458.

Peng:2014:ACD

[PLL14] Yali Peng, Fang Liu, and Shigang Liu. Active contours driven by normalized local image fitting energy. Concurrency and Computation: Practice and Experience, 26(5):1200–1214, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Phuong:2017:PPE

[PLL17] Thi Yen Phuong, Deok-Young Lee, and Jeong-Gun Lee. A performance, power, and energy analysis of ultrasound Bmode imaging on a GPU with VFS. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pavan:2019:EEP

[PLM^{+}19] Pablo J. Pavan, Ricardo K. Lorenzoni, Vinícius R. Machado, Jean L. Bez, Edson L. Padoin, Francieli Z. Boito, Philippe O. A. Navaux, and Jean-François Méhaut. Energy efficiency and I/O performance of low-power architectures. Concurrency and Computation: Practice and Experience, 31(18): e4948:1–e4948:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pichel:2014:USI

[PLR^{+}14] Juan C. Pichel, Juan A. Lorenzo, Francisco F. Rivera, Dora B. Heras, and Tomás F. Pena. Using sampled information: is it enough for the sparse matrix–vector product locality optimization? Concurrency and Computation: Practice and Experience, 26(1):98–117, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2018:WBS

[PLW^{+}18] Junjie Peng, Danxu Liu, Yingtao Wang, Ying Zeng, Feng Cheng, and Wenqiang Zhang. Weight-based strategy for an I/O-intensive application at a cloud data center. Concurrency and Computation: Practice and Experience, 30(19):e4648:1– e4648:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pan:2013:SIP

[PLY13] Maolin Pan, MiaoMiao Li, and Yang Yu. Special issue papers: A group-choose algorithm supporting virtual organization creation for workflow deployment in cloud environment. Concurrency and Computation: Practice and Experience, 25 (13):1894–1908, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2014:PTC

[PLZ14] Tao Peng, Lu Liu, and Wanli Zuo. PU text classification enhanced by term frequency-inverse document frequencyimproved weighting. Concurrency and Computation: Practice and Experience, 26(3):728–741, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pascual:2014:FIF

[PMAL14] Jose A. Pascual, Jose Miguel-Alonso, and Jose A. Lozano. A fast implementation of the first fit contiguous partitioning strategy for cubic topologies. Concurrency and Computation: Practice and Experience, 26(17):2792–2810, December 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peralta:2015:RAS

[PMB15] Manuel Peralta, Supratik Mukhopadhyay, and Ramesh Bharadwaj. Reasoning about security in sensor networks.

Concurrency and Computation: Practice and Experience, 27 (15):3816–3841, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Preece:2008:OBA

[PME^{+}08] A. Preece, P. Missier, S. Embury, B. Jin, and M. Greenwood. An ontology-based approach to handling information quality in e-Science. Concurrency and Computation: Practice and Experience, 20(3):253–264, March 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierce:2015:AAD

[PMG^{+}15] Marlon E. Pierce, Suresh Marru, Lahiru Gunathilake, Don Kushan Wijeratne, Raminder Singh, Chathuri Wimalasena, Shameera Ratnayaka, and Sudhakar Pamidighantam. Apache Airavata: design and directions of a science gateway framework. Concurrency and Computation: Practice and Experience, 27(16):4282–4291, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perumal:2019:SDE

[PMG19] Kaliram Perumal, Suganthi Muthusamy, and Gowrison Gengavel. Sparse data encoder and decoder to improve security in video steganography. Concurrency and Computation: Practice and Experience, 31(14):e4971:1–e4971:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parashar:2005:EIC

[PML^{+}05] Manish Parashar, Rajeev Muralidhar, Wonsuck Lee, Dorian Arnold, Jack Dongarra, and Mary Wheeler. Enabling interactive and collaborative oil reservoir simulations on the Grid. Concurrency and Computation: Practice and Experience, 17 (11):1387–1414, September 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Panahi:2019:JQO

[PN19] Vahideh Panahi and Nima Jafari Navimipour. Join query optimization in the distributed database system using an artificial bee colony algorithm and genetic operators. Concurrency and Computation: Practice and Experience, 31(17):e5218:1– e5218:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Potanin:2004:COC

[PNB04] Alex Potanin, James Noble, and Robert Biddle. Checking ownership and confinement. Concurrency and Computation: Practice and Experience, 16(7):671–687, June 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Prema:2019:SPA

[PNJP19] S. Prema, Rupesh Nasre, R. Jehadeesan, and B. K. Panigrahi. A study on popular auto-parallelization frameworks. Concurrency and Computation: Practice and Experience, 31(17): e5168:1–e5168:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parker:2010:PGR

[PNL10] Jay Parker, Charles Norton, and Gregory Lyzenga. Parallel GeoFEST for regional faulted deformation. Concurrency and Computation: Practice and Experience, 22(12):1604–1625, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pascal:2017:CPN

[PP17] Carlos Pascal and Doru Panescu. A colored Petri net model for DisCSP algorithms. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peternier:2014:HPE

[PPBB14] Achille Peternier, Cesare Pautasso, Walter Binder, and Daniele Bonetta. High-performance execution of service compositions: a multicore-aware engine design. Concurrency and Computation: Practice and Experience, 26(1):71–97, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Petcu:2015:CRO

[PPC^{+}15] Dana Petcu, Silviu Panica, Ciprian Crăciun, Marian Neagul, and Calin Şandru. Cloud resource orchestration within an open-source component-based platform as a service. Concurrency and Computation: Practice and Experience, 27(9): 2443–2469, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pisani:2017:CSR

[PPdSTB17] Flávia Pisani, Daniel C. G. Pedronette, Ricardo da S. Torres, and Edson Borin. Contextual Spaces Re-Ranking: accelerating the Re-sort Ranked Lists step on heterogeneous systems. Concurrency and Computation: Practice and Experience, 29 (22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Passerat-Palmbach:2015:TSS

[PPMH15] Jonathan Passerat-Palmbach, Claude Mazel, and David R. C. Hill. TaskLocalRandom: a statistically sound substitute to pseudorandom number generation in parallel Java tasks frameworks. Concurrency and Computation: Practice and Experience, 27(13):3383–3398, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Plaza:2010:PHC

[PPP10] Antonio J. Plaza, Javier Plaza, and Abel Paz. Parallel heterogeneous CBIR system for efficient hyperspectral image retrieval using spectral mixture analysis. Concurrency and Computation: Practice and Experience, 22(9):1138–1159, June 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ponnuswamy:2019:FRS

[PPR19] P. Priya Ponnuswamy, R. Vidhya Priya, and C. P. Shabari Ram. File retrieval and storage in the open source cloud tool using digital bipartite and digit compact prefix indexing method. Concurrency and Computation: Practice and Experience, 31(14):e5307:1–e5307:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Phan:2009:APP

[PPST09] Tu Huy Phan, Enrico Pontelli, Tran Cao Son, and Son Thanh To. Applications of parallel processing technologies in heuristic search planning: methodologies and experiments. Concurrency and Computation: Practice and Experience, 21(15): 1928–1960, October 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pontisso:2013:ADM

[PQP13] Nadège Pontisso, Philippe Quéinnec, and Gérard Padiou. Analysis of distributed multiperiodic systems to achieve consistent data matching. Concurrency and Computation: Practice and Experience, 25(2):234–249, February 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perry:2014:GDB

[PRC^{+}14] Daniel Perry, John Robinson, Stephanie Cruz, Cecilia Aragon, Jeanne Ting Chowning, and Mette Peters. Game design for bioinformatics and cyberinfrastructure learning: a parallel computing case study. Concurrency and Computation: Practice and Experience, 26(13):2303–2315, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Patel:2016:ARB

[PRCV16] Ishan Patel, Andrew Rau-Chaplin, and Blesson Varghese. Accelerating R-based analytics on the cloud. Concurrency and Computation: Practice and Experience, 28(4):977–994, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pedersen:2013:SIP

[PRD^{+}13] Jens Myrup Pedersen, M. Tahir Riaz, Bozydar Dubalski, Damian Ledzinski, Joaquim Celestino Júnior, and Ahmed Patel. Special issue papers: Using latency as a QoS indicator for global cloud computing services. Concurrency and Computation: Practice and Experience, 25(18):2488–2500, December 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Prehofer:2001:FOP

[Pre01] Christian Prehofer. Feature-oriented programming: a new way of object composition. Concurrency and Computation: Practice and Experience, 13(6):465–501, May 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/79503111/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
79503111&PLACEBO=IE.pdf.

Pereira:2015:PSP

[PRG15] Alyson D. Pereira, Luiz Ramos, and Luís F. W. Góes. PSkel: a stencil programming framework for CPU–GPU systems. Concurrency and Computation: Practice and Experience, 27 (17):4938–4953, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Portella:2019:SAA

[PRNM19] Gustavo Portella, Genaina N. Rodrigues, Eduardo Nakano, and Alba C. M. A. Melo. Statistical analysis of Amazon EC2 cloud pricing models. Concurrency and Computation: Practice and Experience, 31(18):e4451:1–e4451:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Punceva:2015:IRS

[PRP^{+}15] Magdalena Punceva, Ivan Rodero, Manish Parashar, Omer. F. Rana, and Ioan Petri. Incentivising resource sharing in social clouds. Concurrency and Computation: Practice and Experience, 27(6):1483–1497, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Puliafito:2001:WPS

[PRS01] Antonio Puliafito, Salvatore Riccobene, and Marco Scarpa. Which paradigm should I use? An analytical comparison of the client-server, remote evaluation and mobile agent paradigms. Concurrency and Computation: Practice and Experience, 13(1):71–94, January 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/77004399/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=77004399&PLACEBO=IE.
pdf.

Palopoli:2016:DMT

[PRS16] Luigi Palopoli, Domenico Rosaci, and Giuseppe M. L. Sarné. A distributed and multi-tiered software architecture for assessing e-Commerce recommendations. Concurrency and Computation: Practice and Experience, 28(18):4507–4531, December 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pascual:2009:CMO

[PRT09] Fanny Pascual, Krzysztof Rzadca, and Denis Trystram. Cooperation in multi-organization scheduling. Concurrency and Computation: Practice and Experience, 21(7):905–921, May 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pyka:2014:RTC

[PRU14] Arthur Pyka, Mathias Rohde, and Sascha Uhrig. A realtime capable coherent data cache for multicores. Concurrency and Computation: Practice and Experience, 26(6):1342–1354, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pineau:2011:EAS

[PRV11] Jean-François Pineau, Yves Robert, and Frédéric Vivien. Energy-aware scheduling of bag-of-tasks applications on master–worker platforms. Concurrency and Computation: Practice and Experience, 23(2):145–157, February 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pechtchanski:2005:ISA

[PS05] Igor Pechtchanski and Vivek Sarkar. Immutability specification and its applications. Concurrency and Computation: Practice and Experience, 17(5–6):639–662, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pozniansky:2007:MEF

[PS07] Eli Pozniansky and Assaf Schuster. MultiRace: efficient onthe-fly data race detection in multithreaded C++ programs. Concurrency and Computation: Practice and Experience, 19 (3):327–340, March 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Paszynski:2010:GGD

[PS10] Maciej Paszyński and Robert Schaefer. Graph grammardriven parallel partial differential equation solver. Concurrency and Computation: Practice and Experience, 22(9): 1063–1097, June 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Porto:2013:EDM

[PS13] Fabio Porto and Bruno Schulze. Editorials: Data management for eScience in Brazil. Concurrency and Computation: Practice and Experience, 25(16):2307–2309, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Poovendran:2019:AEL

[PS19a] R. Poovendran and S. Sumathi. An area-efficient low-power SCM topology for high performance Network-on Chip (NoC) architecture using an optimized routing design. Concurrency and Computation: Practice and Experience, 31(14):e4760:1– e4760:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pradeep:2019:OMA

[PS19b] M. Pradeep and P. Sampath. An optimized multi-attribute vertical handoff approach for heterogeneous wireless networks. Concurrency and Computation: Practice and Experience, 31(20):e5296:1–e5296:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Platania:2017:HBR

[PSC^{+}17] Richard Platania, Shayan Shams, Chui-Hui Chiu, Nayong Kim, Joohyun Kim, and Seung-Jong Park. Hadoop-based replica exchange over heterogeneous distributed cyberinfrastructures. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Porter-Sobieraj:2015:OCP

[PSCK^{+}15] J. Porter-Sobieraj, S. Cygert, D. Kikoła, J. Sikorski, and M. Słodkowski. Optimizing the computation of a parallel 3D finite difference algorithm for graphics processing units. Concurrency and Computation: Practice and Experience, 27(6): 1591–1602, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pan:2003:SHI

[PSG03] Yi Pan, Joseph J. S. Shang, and Minyi Guo. A scalable HPF implementation of a finite-volume computational electromagnetics application on a CRAY T3E parallel system. Concurrency and Computation: Practice and Experience, 15(6):607– 621, May 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peters:2011:FPC

[PSHL11] Hagen Peters, Ole Schulz-Hildebrandt, and Norbert Luttenberger. Fast in-place, comparison-based sorting with CUDA: a study with bitonic sort. Concurrency and Computation: Practice and Experience, 23(7):681–693, May 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perez-Serrano:2018:EBT

[PSICU18] J. Pérez-Serrano, B. Imbernón, J. M. Cecilia, and M. Ujaldón. Energy-based tuning of metaheuristics for molecular docking on multi-GPUs. Concurrency and Computation: Practice and Experience, 30(17):e4684:1–e4684:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Popescu:2016:NUB

[PSIP16] Pantelimon George Popescu, Emil-Ioan Sluşanschi, Voichiţa Iancu, and Florin Pop. A new upper bound for Shannon entropy. A novel approach in modeling of Big Data applications. Concurrency and Computation: Practice and Experience, 28 (2):351–359, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Puiggali:2013:DBS

[PSJM13] Joan Puiggali, Boleslaw K. Szymanski, Teo Jové, and Jose L. Marzo. Dynamic branch speculation in a speculative parallelization architecture for computer clusters. Concurrency and Computation: Practice and Experience, 25(7):932–960, May 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pakin:2016:PUP

[PSL^{+}16] Scott Pakin, Curtis Storlie, Michael Lang, Robert E. Fields III, Eloy E. Romero, Jr., Craig Idler, Sarah Michalak, Hugh Greenberg, Josip Loncaric, Randal Rheinheimer, Gary Grider, and Joanne Wendelberger. Power usage of production supercomputers and production workloads. Concurrency and Computation: Practice and Experience, 28(2):274–290, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Paravati:2011:OSA

[PSLC11] Gianluca Paravati, Andrea Sanna, Fabrizio Lamberti, and Luigi Ciminiera. An open and scalable architecture for delivering 3D shared visualization services to heterogeneous devices. Concurrency and Computation: Practice and Experience, 23(11):1179–1195, August 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pang:2003:PSR

[PSM03] James C. Pang, Gholamali C. Shoja, and Eric G. Manning. Providing soft real-time quality of service guarantees for Java threads. Concurrency and Computation: Practice and Experience, 15(3–5):521–538, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Petrou:2011:OPP

[PSM^{+}11] Savvas Petrou, Terence M. Sloan, Muriel Mewissen, Thorsten Forster, Michal Piotrowski, Bartosz Dobrzelecki, Peter Ghazal, Arthur Trew, and Jon Hill. Optimization of a parallel permutation testing function for the SPRINT R package. Concurrency and Computation: Practice and Experience, 23 (17):2258–2268, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pallipuram:2014:RBP

[PSRR14] Vivek K. Pallipuram, Melissa C. Smith, Nimisha Raut, and Xiaoyu Ren. A regression-based performance prediction framework for synchronous iterative algorithms on general purpose graphical processing unit clusters. Concurrency and Computation: Practice and Experience, 26(2):532–560, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pinto:2018:VPA

[PSS^{+}18] Vinícius Garcia Pinto, Lucas Mello Schnorr, Luka Stanisic, Arnaud Legrand, Samuel Thibault, and Vincent Danjean. A visual performance analysis framework for task-based parallel applications running on hybrid clusters. Concurrency and Computation: Practice and Experience, 30(18):e4472:1– e4472:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Penha:2019:AAD

[PSS^{+}19] Jeronimo C. Penha, Lucas B. Silva, Jansen M. Silva, Kristtopher K. Coelho, Hector P. Baranda, José Augusto M. Nacif, and Ricardo S. Ferreira. ADD: Accelerator Design and deploy — a tool for FPGA high-performance dataflow computing. Concurrency and Computation: Practice and Experience, 31(18):e5096:1–e5096:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pikle:2019:AFE

[PSV19] Nileshchandra K. Pikle, Shailesh R. Sathe, and Arvind Y. Vyavahare. Accelerating the finite element analysis of functionally graded materials using fixed-grid strategy on CUDAenabled GPUs. Concurrency and Computation: Practice and Experience, 31(17):e5207:1–e5207:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pour:2011:MBD

[PSW11] Niusha Hakimi Pour, Paul Strooper, and Andy Wellings. A model-based development approach for the verification of real-time Java code. Concurrency and Computation: Practice and Experience, 23(13):1583–1606, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Polze:2012:TCO

[PT12] Andreas Polze and Peter Tröger. Trends and challenges in operating systems—from parallel computing to cloud computing. Concurrency and Computation: Practice and Experience, 24(7):676–686, May 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Paventhan:2007:MNM

[PTCN07] A. Paventhan, Kenji Takeda, Simon J. Cox, and Denis A. Nicole. MyCoG.NET: a multi-language CoG toolkit. Concurrency and Computation: Practice and Experience, 19(14): 1885–1900, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peraza:2016:PGQ

[PTL^{+}16] Joshua Peraza, Ananta Tiwari, Michael Laurenzano, Laura Carrington, and Allan Snavely. PMaC’s green queue: a framework for selecting energy optimal DVFS configurations in large scale MPI applications. Concurrency and Computation: Practice and Experience, 28(2):211–231, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Puffitsch:2013:SIP

[Puf13] Wolfgang Puffitsch. Special issue papers: Design and analysis of a hard real-time garbage collector for a Java chip multi-processor. Concurrency and Computation: Practice and Experience, 25(16):2269–2289, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Puntigam:2001:STC

[Pun01] Franz Puntigam. Strong types for coordinating active objects. Concurrency and Computation: Practice and Experience, 13 (4):293–326, April 10, 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/78502301/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=78502301&PLACEBO=IE.pdf.

Peinado:2004:PBA

[PV04] Jesús Peinado and Antonio M. Vidal. A parallel Broyden approach to the Toeplitz inverse eigenproblem. Concurrency and Computation: Practice and Experience, 16(6):587–610, May 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Prakash:2015:MAT

[PV15] Shiv Prakash and Deo Prakash Vidyarthi. Maximizing availability for task scheduling in computational grid using genetic algorithm. Concurrency and Computation: Practice and Experience, 27(1):193–210, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parkinson:2018:GGT

[PVCS18] Simon Parkinson, Mauro Vallati, Andrew Crampton, and Shirin Sohrabi. GraphBAD: a general technique for anomaly detection in security information and event management. Concurrency and Computation: Practice and Experience, 30(16):??, August 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4433.

Parashar:2002:CCG

[PvLV^{+}02] Manish Parashar, Gregor von Laszewski, Snigdha Verma, Jarek Gawor, Kate Keahey, and Nell Rehn. A CORBA Commodity Grid Kit. Concurrency and Computation: Practice and Experience, 14(13–15):1057–1074, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pandey:2009:GWE

[PVR^{+}09] Suraj Pandey, William Voorsluys, Mustafizur Rahman, Rajkumar Buyya, James E. Dobson, and Kenneth Chiu. A grid workflow environment for brain imaging analysis on distributed systems. Concurrency and Computation: Practice and Experience, 21(16):2118–2139, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peszynska:2005:SIH

[PW05] M. Peszynska and M. F. Wheeler. Special issue: Highperformance computing in geosciences. Concurrency and Computation: Practice and Experience, 17(11):1363–1364, September 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pages:2012:GCF

[PW12] Gilles Pagès and Benedikt Wilbertz. GPGPUs in computational finance: massive parallel computing for American style options. Concurrency and Computation: Practice and Experience, 24(8):837–848, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Padmanabhan:2014:FCA

[PWC^{+}14] Anand Padmanabhan, Shaowen Wang, Guofeng Cao, Myunghwa Hwang, Zhenhua Zhang, Yizhao Gao, Kiumars Soltani, and Yan Liu. FluMapper: a cyberGIS application for interactive analysis of massive location-based social media. Concurrency and Computation: Practice and Experience, 26(13):2253–2265, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2018:OET

[PWH18] Liangqing Peng, Yanpeng Wu, and Lei Huang. Opto-electric target tracking algorithm based on local feature selection and particle filter optimization. Concurrency and Computation: Practice and Experience, 30(22):e4670:1–e4670:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2010:PSC

[PWJ10] Hanshu Peng, Zhongliang Wu, and Changsheng Jiang. Preseismic changes of noise correlation function (NCF) before the Wenchuan earthquake? Concurrency and Computation: Practice and Experience, 22(12):1774–1783, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pei:2016:RMF

[PWMX16] Xiaoqiang Pei, Yijie Wang, Xingkong Ma, and Fangliang Xu. Repairing multiple failures adaptively with erasure codes in distributed storage systems. Concurrency and Computation: Practice and Experience, 28(5):1437–1461, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pei:2017:DRG

[PWMX17] Xiaoqiang Pei, Yijie Wang, Xingkong Ma, and Fangliang Xu. A decentralized redundancy generation scheme for codes with locality in distributed storage systems. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pham:2019:SSS

[PWS19] Hoang Pham, Jason Woodworth, and Mohsen Amini Salehi. Survey on secure search over encrypted data on the cloud. Concurrency and Computation: Practice and Experience, 31(17):e5284:1–e5284:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parastatidis:2005:WGF

[PWWR05] Savas Parastatidis, Jim Webber, Paul Watson, and Thomas Rischbeck. WS-GAF: a framework for building Grid applications using Web Services. Concurrency and Computation: Practice and Experience, 17(2–4):391–417, February/ April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Price:2007:OIE

[PXY^{+}07] A. R. Price, G. Xue, A. Yool, D. J. Lunt, P. J. Valdes, T. M. Lenton, J. L. Wason, G. E. Pound, and S. J. Cox. Optimization of integrated Earth System Model components using Grid-enabled data management and computation. Concurrency and Computation: Practice and Experience, 19(2): 153–165, February 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pja:2019:SSG

[PY19] Alphonse Pja and Venkatramana Reddy Y. Scalable and secure group key agreement for wireless ad-hoc networks by extending RSA scheme. Concurrency and Computation: Practice and Experience, 31(14):e4969:1–e4969:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierce:2002:GCW

[PYF02] Marlon E. Pierce, Choonhan Youn, and Geoffrey C. Fox. The Gateway computational Web portal. Concurrency and Computation: Practice and Experience, 14(13–15):1411–1426, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Park:2016:DGB

[PYKL16] JiSu Park, HeonChang Yu, Hyongsoon Kim, and Eunyoung Lee. Dynamic group-based fault tolerance technique for reliable resource management in mobile cloud computing. Concurrency and Computation: Practice and Experience, 28(10): 2756–2769, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pande:2011:EMA

[PZ11] Amit Pande and Joseph Zambreno. Efficient mapping and acceleration of AES on custom multi-core architectures. Concurrency and Computation: Practice and Experience, 23(4): 372–389, March 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Prill:2017:CPA

[PZ17] Florian Prill and Günther Zängl. A compact parallel algorithm for spherical Delaunay triangulations. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2015:DNL

[PZH^{+}15] Li Peng, Zhiying Zhang, Qiaoli Huang, Zhixing Huang, and Hai Zhuge. Designing a novel linear-time graph kernel for semantic link network. Concurrency and Computation: Practice and Experience, 27(15):4039–4052, October 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pan:2018:EAR

[PZHS18] Ning Pan, Zhiqiang Zhu, Liangsheng He, and Lei Sun. An efficiency approach for RBAC reconfiguration with minimal roles and perturbation. Concurrency and Computation: Practice and Experience, 30(11):??, June 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4399.

Pan:2019:ERA

[PZO19] Lei Pan, Jun Zhang, and Jonathan Oliver. Editorial: Recent advances in machine learning for cybersecurity. Concurrency and Computation: Practice and Experience, 31(19):e5270:1– e5270:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2008:TEW

[PZZ08] Tao Peng, Changli Zhang, and Wanli Zuo. Tunneling enhanced by Web page content block partition for focused crawling. Concurrency and Computation: Practice and Experience, 20(1):61–74, January 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2010:TEW

[PZZ10] Tao Peng, Changli Zhang, and Wanli Zuo. Tunneling enhanced by Web page content block partition for focused crawling. Concurrency and Computation: Practice and Experience, 22(4):538–539, March 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiu:2012:PWM

[QB12] Judy Qiu and Seung-Hee Bae. Performance of windows multicore systems on threading and MPI. Concurrency and Computation: Practice and Experience, 24(1):14–28, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qu:2017:MIS

[QCB17] Chenhao Qu, Rodrigo Neves Calheiros, and Rajkumar Buyya. Mitigating impact of short-term overload on multicloud web applications through geographical load balancing. Concurrency and Computation: Practice and Experience, 29 (12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qi:2017:ENA

[QD17] Heng Qi and Mianxiong Dong. Editorial: New advances in future network technologies. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qu:2010:TAW

[QEB^{+}10] Yenan Qu, Gordon Erlebacher, Evan Bollig, Julien Lafourcade, and Magali Lapeyre-Mirande. Toolkits for automatic Web service and GUI generation: KWATT. Concurrency and Computation: Practice and Experience, 22(12):1703–1719, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiu:2014:EEC

[QFG14] Judy Qiu, Ian Foster, and Carole Goble. Editorials: Emerging Computational Methods for the Life Sciences Workshop 2012.

Concurrency and Computation: Practice and Experience, 26 (6):1231–1233, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiu:2014:ESI

[QFT14] Judy Qiu, Ian Foster, and Ronald Taylor. Editorials: Special issue for Emerging Computational Methods for the Life Sciences Workshop. Concurrency and Computation: Practice and Experience, 26(4):851–853, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiu:2018:PPS

[QGZL18] Meikang Qiu, Keke Gai, Hui Zhao, and Meiqin Liu. Privacypreserving smart data storage for financial industry in cloud computing. Concurrency and Computation: Practice and Experience, 30(5):??, March 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4278.

Quenette:2010:SSP

[QH10] S. Quenette and L. Hodkinson. StgDomain — scalable parallel domain software components for particle-in-cell finite element methods. Concurrency and Computation: Practice and Experience, 22(12):1593–1603, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qi:2017:FVS

[Qi17] Man Qi. Facilitating visual surveillance with motion detections. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qian:2019:RIW

[Qia19] Guo Qian. Research on index weight of logistics integration based on cloud models. Concurrency and Computation: Practice and Experience, 31(9):e4632:1–e4632:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiu:2011:ECM

[Qiu11] Judy Qiu. Emerging Computational Methods for the Life Sciences Workshop special issue. Concurrency and Computation: Practice and Experience, 23(17):2249, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qu:2007:DMA

[QKSJ07] Wenyu Qu, Masaru Kitsuregawa, Hong Shen, and Yingwei Jin. Distribution of mobile agents in vulnerable networks. Concurrency and Computation: Practice and Experience, 19 (7):1047–1064, May 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Quema:2004:AMG

[QLC04] V. Quéma, R. Lachaize, and E. Cecchet. An asynchronous middleware for Grid resource monitoring. Concurrency and Computation: Practice and Experience, 16(5):523–534, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qi:2011:QAE

[QLD^{+}11] Lianyong Qi, Wenmin Lin, Wanchun Dou, Jian Jiang, and Jinjun Chen. A QoS-aware exception handling method in scientific workflow execution. Concurrency and Computation: Practice and Experience, 23(16):1951–1968, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qi:2006:MCF

[QLF^{+}06] Zhengwei Qi, Minglu Li, Cheng Fu, Dongyu Shi, and Jinyuan You. Membrane Calculus: a formal method for Grid transactions. Concurrency and Computation: Practice and Experience, 18(14):1799–1809, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qu:2010:SIA

[QLL10] Wenyu Qu, Zhaobin Liu, and Kai Lin. Special issue: advanced topics on scalable computing. Concurrency and Computation: Practice and Experience, 22(13):1849–1851, September 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qi:2015:EFH

[QLLS15] Heng Qi, Yang Li, Keqiu Li, and Milos Stojmenovic. An exchanged folded hypercube-based topology structure for interconnection networks. Concurrency and Computation: Practice and Experience, 27(16):4194–4210, November 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Quesnel:2013:SIP

[QLS13] Flavien Quesnel, Adrien Lèbre, and Mario Südholt. Special issue papers: Cooperative and reactive scheduling in large-scale virtualized platforms with DVMS. Concurrency and Computation: Practice and Experience, 25(12):1643–1655, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiao:2019:RAM

[QLZX19] Jian Qiao, Qinghua Lu, Wenbao Zhang, and Xiangyao Xue. The research about micromachining accuracy control based on fuzzy evaluation algorithm. Concurrency and Computation: Practice and Experience, 31(10):e4745:1–e4745:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qureshi:2012:TIE

[QMK12] Basit Qureshi, Geyong Min, and Demetres Kouvatsos. Trusted information exchange in peer-to-peer mobile social networks. Concurrency and Computation: Practice and Experience, 24(17):2055–2068, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qu:2017:CFS

[QML^{+}17] Wenyu Qu, Minmin Ma, Zhiyang Li, Milos Stojmenovic, and Zhaobin Liu. A coarse-to-fine shape decomposition based on critical points. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Quinlan:2004:POO

[QSMK04] Daniel J. Quinlan, Markus Schordan, Brian Miller, and Markus Kowarschik. Parallel object-oriented framework optimization. Concurrency and Computation: Practice and Experience, 16(2–3):293–302, February/March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiao:2017:FAD

[QSX^{+}17] Yuran Qiao, Junzhong Shen, Tao Xiao, Qianming Yang, Mei Wen, and Chunyuan Zhang. FPGA-accelerated deep convolutional neural networks for high throughput and energy efficiency. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qian:2018:FRA

[QSZL18] Kun Qian, Kehua Su, Jialing Zhang, and Yinghua Li. A 3D face registration algorithm based on conformal mapping. Concurrency and Computation: Practice and Experience, 30(22):e4654:1–e4654:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Quan:2017:DEP

[QW17] Zhe Quan and Lei Wu. Design and evaluation of a parallel neighbor algorithm for the disjunctively constrained knapsack problem. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qin:2016:VLK

[QWW^{+}16] Bo Qin, Linxiao Wang, Yujue Wang, Qianhong Wu, Wenchang Shi, and Bin Liang. Versatile lightweight key distribution for big data privacy in vehicular ad hoc networks. Concurrency and Computation: Practice and Experience, 28(10): 2920–2939, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiu:2019:CLB

[QWZ^{+}19] Ruo-Fan Qiu, Hai-Ning Wang, Jian-Feng Zhu, Rong-Qian Chen, Cheng-Xiang Zhu, and Yan-Cheng You. Compressible lattice Boltzmann simulations on high-performance and lowcost GeForce GPU. Concurrency and Computation: Practice and Experience, 31(20):e5341:1–e5341:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiang:2017:DFA

[QXJS17] Weizhong Qiang, Shifan Xin, Hai Jin, and Guozhong Sun. DroidAuditor: a framework for auditing covert communication on Android. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qian:2016:GBH

[QXXZ16] Quan Qian, Mengbo Xie, Chaojie Xiao, and Rui Zhang. Gridbased high performance ensemble classification for evolving data stream. Concurrency and Computation: Practice and Experience, 28(16):4339–4351, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiang:2016:SCF

[QZDJ16] Weizhong Qiang, Kang Zhang, Weiqi Dai, and Hai Jin. Secure cryptographic functions via virtualization-based outsourced computing. Concurrency and Computation: Practice and Experience, 28(11):3149–3163, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiang:2016:ESN

[QZH16] Weizhong Qiang, Xianghan Zheng, and Ching-Hsien Hsu. Editorials: Social network analysis and its application. Concurrency and Computation: Practice and Experience, 28(14): 3798–3802, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qu:2016:ILT

[QZYZ16] Peng Qu, Junsheng Zhang, Changqing Yao, and Wen Zeng. Identifying long tail term from large-scale candidate pairs for big data-oriented patent analysis. Concurrency and Computation: Practice and Experience, 28(15):4194–4208, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rolan:2014:FGT

[RAFD14] Dyer Rolán, Diego Andrade, Basilio B. Fraguela, and Ramón Doallo. A fine-grained thread-aware management policy for shared caches. Concurrency and Computation: Practice and Experience, 26(6):1355–1374, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ravve:2016:ICS

[Rav16] Elena V. Ravve. Incremental computations over strongly distributed databases. Concurrency and Computation: Practice and Experience, 28(11):3061–3076, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rodriguez:2017:TSS

[RB17] Maria Alejandra Rodriguez and Rajkumar Buyya. A taxonomy and survey on scheduling algorithms for scientific workflows in IaaS cloud computing environments. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Riesen:2009:DIL

[RBB^{+}09] Rolf Riesen, Ron Brightwell, Patrick G. Bridges, Trammell Hudson, Arthur B. Maccabe, Patrick M. Widener, and Kurt Ferreira. Designing and implementing lightweight kernels for capability computing. Concurrency and Computation: Practice and Experience, 21(6):793–817, April 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rheinheimer:2002:ACG

[RBBH02] Randal Rheinheimer, Judy I. Beiriger, Hugh P. Bivens, and Steven L. Humphreys. The ASCI Computational Grid: initial deployment. Concurrency and Computation: Practice and Experience, 14(13–15):1351–1363, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rossi:2017:CTP

[RBDI17] Silvia Rossi, Francesco Barile, Sergio Di Martino, and Davide Improta. A comparison of two preference elicitation approaches for museum recommendations. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ranjan:2015:ENR

[RBNG15] Rajiv Ranjan, Rajkumar Buyya, Surya Nepal, and Dimitrios Georgakopulos. Editorials: A note on resource orchestration for cloud computing. Concurrency and Computation: Practice and Experience, 27(9):2370–2372, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raje:2002:QSB

[RBO^{+}02] Rajeev R. Raje, Barrett R. Bryant, Andrew M. Olson, Mikhail Auguston, and Carol Burt. A quality-of-service-based framework for creating distributed heterogeneous software components. Concurrency and Computation: Practice and Experience, 14(12):1009–1034, October 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/98516165/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
98516165{\&}PLACEBO=IE.pdf.

Ranjan:2012:SSA

[RBP12] Rajiv Ranjan, Rajkumar Buyya, and Manish Parashar. Special section on autonomic cloud computing: technologies, services, and applications. Concurrency and Computation: Practice and Experience, 24(9):935–937, June 25, 2012. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raouf:2018:DDR

[RBT18] Ahmed E. Abdel Raouf, Nagwa L. Badr, and Mohamed F. Tolba. Dynamic data reallocation and replication over a cloud environment. Concurrency and Computation: Practice and Experience, 30(13):??, July 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4416.

Ren:2009:OEP

[RC09] Kaijun Ren and Jinjun Chen. Optimizing execution path of scientific workflow by gradual removal of QoS constraint violations in Reverse Order. Concurrency and Computation: Practice and Experience, 21(16):2033–2051, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rocha:2011:ACE

[RCA^{+}11] B. M. Rocha, F. O. Campos, R. M. Amorim, G. Plank, R. W. dos Santos, M. Liebmann, and G. Haase. Accelerating cardiac excitation spread simulations using graphics processing units. Concurrency and Computation: Practice and Experience, 23 (7):708–720, May 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Riakiotakis:2012:TOS

[RCA^{+}12] I. Riakiotakis, F. M. Ciorba, T. Andronikos, G. Papakonstantinou, and A. T. Chronopoulos. Towards the optimal synchronization granularity for dynamic scheduling of pipelined computations on heterogeneous computing systems. Concurrency and Computation: Practice and Experience, 24(18): 2302–2327, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Riley:2003:HPJ

[RCB03] Christopher J. Riley, Siddhartha Chatterjee, and Rupak Biswas. High-performance Java codes for computational fluid dynamics. Concurrency and Computation: Practice and Experience, 15(3–5):395–415, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raman:2004:ACR

[RCB^{+}04] Sanjay Raman, Dwaine Clarke, Matt Burnside, Srinivas Devadas, and Ronald Rivest. Access-controlled resource discovery in pervasive networks. Concurrency and Computation: Practice and Experience, 16(11):1099–1120, September 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rahman:2017:CIH

[RCC17] Nurul Hidayah Ab Rahman, Niken Dwi Wahyu Cahyani, and Kim-Kwang Raymond Choo. Cloud incident handling and forensic-by-design: cloud storage as a case study. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rong:2019:SJH

[RCCH19] Chuitian Rong, Xiaohai Cheng, Ziliang Chen, and Na Huo. Similarity joins for high-dimensional data using Spark. Concurrency and Computation: Practice and Experience, 31(20): e5339:1–e5339:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Richter:2012:ECS

[RCKV12] J. Richter, M. Baruwal Chhetri, R. Kowalczyk, and Q. Bao Vo. Establishing composite SLAs through concurrent QoS negotiation with surplus redistribution. Concurrency and Computation: Practice and Experience, 24(9):938–955, June 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rosciszewski:2016:KNW

[RCLSK16] Paweł Rościszewski, Paweł Czarnul, Rafał Lewandowski, and Marcel Schally-Kacprzak. KernelHive: a new workflow-based framework for multilevel high performance computing using clusters and workstations with CPUs and GPUs. Concurrency and Computation: Practice and Experience, 28(9): 2586–2607, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roschke:2012:ACP

[RCM12] Sebastian Roschke, Feng Cheng, and Christoph Meinel. An alert correlation platform for memory-supported techniques. Concurrency and Computation: Practice and Experience, 24 (10):1123–1136, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rojek:2015:AFM

[RCR^{+}15] Krzysztof Andrzej Rojek, Milosz Ciznicki, Bogdan Rosa, Piotr Kopta, Michal Kulczewski, Krzysztof Kurowski, Zbigniew Pawel Piotrowski, Lukasz Szustak, Damian Karol Wojcik, and Roman Wyrzykowski. Adaptation of fluid model EU-LAG to graphics processing unit architecture. Concurrency and Computation: Practice and Experience, 27(4):937–957, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ruiz:2003:FSB

[RCT03] David Ruiz, Rafael Corchuelo, and Miguel Toro. Fairness in systems based on multiparty interactions. Concurrency and Computation: Practice and Experience, 15(11–12):1093– 1116, September 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2009:BQS

[RCXS09] Kaijun Ren, Jinjun Chen, Nong Xiao, and Junqiang Song. Building Quick Service Query list (QSQL) to support automated service discovery for scientific workflow. Concurrency and Computation: Practice and Experience, 21(16):

2099–2117, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rotaru:2010:SOM

[RDP10] T. Rotaru, M. Dalheimer, and F.-J. Pfreundt. Serviceoriented middleware for financial Monte Carlo simulations on the cell broadband engine. Concurrency and Computation: Practice and Experience, 22(5):643–657, April 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Radestock:2003:CCM

[RE03] Matthias Radestock and Susan Eisenbach. Coordinating components in middleware systems. Concurrency and Computation: Practice and Experience, 15(13):1205–1231, November 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reeve:2001:PVD

[Ree01] J. S. Reeve. A parallel Viterbi decoding algorithm. Concurrency and Computation: Practice and Experience, 13(2): 95–102, February 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/77004418/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=77004418&PLACEBO=IE.pdf.

Rattanatamrong:2015:IRT

[RF15] P. Rattanatamrong and J. A. B. Fortes. Improved real-time scheduling of periodic tasks on multiprocessors. Concurrency and Computation: Practice and Experience, 27(9):2291–2309, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rosenberg:2017:BSC

[RG17] Ishai Rosenberg and Ehud Gudes. Bypassing system calls– based intrusion detection systems. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rai:2018:VCE

[RG18] Gopal N. Rai and G. R. Gangadharan. Verifying compositional equivalence between web service composition graphs. Concurrency and Computation: Practice and Experience, 30(16):??, August 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4434.

Rasch:2019:AGD

[RG19] Ari Rasch and Sergei Gorlatch. ATF: a generic directivebased auto-tuning framework. Concurrency and Computation: Practice and Experience, 31(5):e4423:1–e4423:??, March 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

RahimiZadeh:2015:PEW

[RGAK15] Keyvan RahimiZadeh, Reza Nasiri Gerde, Morteza AnaLoui, and Peyman Kabiri. Performance evaluation of Web server workloads in Xen-based virtualized computer system: analytical modeling and experimental validation. Concurrency and Computation: Practice and Experience, 27(17):4741–4762, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rucci:2015:EAP

[RGB^{+}15] Enzo Rucci, Carlos García, Guillermo Botella, Armando De Giusti, Marcelo Naiouf, and Manuel Prieto-Matías. An energy-aware performance analysis of SWIMM: Smith– Waterman implementation on Intel’s Multicore and Manycore architectures. Concurrency and Computation: Practice and Experience, 27(18):5517–5537, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ricci:2015:ACD

[RGCC15] Laura Ricci, Luca Genovali, Emanuele Carlini, and Massimo Coppola. AOI-cast in distributed virtual environments: an approach based on delay tolerant reverse compass routing. Concurrency and Computation: Practice and Experience, 27 (9):2329–2350, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Romero:2015:DPC

[RGL^{+}15] N. A. Romero, C. Glinsvad, A. H. Larsen, J. Enkovaara, S. Shende, V. A. Morozov, and J. J. Mortensen. Design and performance characterization of electronic structure calculations on massively parallel supercomputers: a case study of GPAW on the Blue Gene/P architecture. Concurrency and Computation: Practice and Experience, 27(1):69–93, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rosinha:2009:WPP

[RGV09] Romulo B. Rosinha, Cláudio F. R. Geyer, and Patrícia Kayser Vargas. WSPE: a peer-to-peer grid programming environment. Concurrency and Computation: Practice and Experience, 21(13):1709–1724, September 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2017:PPU

[RGX^{+}17] Bangbang Ren, Deke Guo, Junjie Xie, Wenxin Li, Bo Yuan, and Yunfei Liu. The packing problem of uncertain multicasts. Concurrency and Computation: Practice and Experience, 29 (16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rostami:2007:TAO

[RH07] Habib Rostami and Jafar Habibi. Topology awareness of overlay P2P networks. Concurrency and Computation: Practice and Experience, 19(7):999–1021, May 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rossinelli:2011:WAS

[RHBK11] Diego Rossinelli, Babak Hejazialhosseini, Michael Bergdorf, and Petros Koumoutsakos. Wavelet-adaptive solvers on multi-core architectures for the simulation of complex systems. Concurrency and Computation: Practice and Experience, 23(2):172–186, February 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rocha:2016:WNE

[RHD^{+}16] Rodrigo Rocha, Bruno Hott, Vinícius Dias, Renato Ferreira, Wagner Meira, and Dorgival Guedes. Watershed-ng: an extensible distributed stream processing framework. Concurrency and Computation: Practice and Experience, 28(8): 2487–2502, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Remick:2013:E

[RHJ13] Carolyn Remick, James Hunt, and Harold Javid. Editorial. Concurrency and Computation: Practice and Experience, 25 (4):443–444, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rousset:2018:NGM

[RHL^{+}18] Alban Rousset, Bénédicte Herrmann, Christophe Lang, Laurent Philippe, and Hadrien Bride. Nested graphs: a model to efficiently distribute multi-agent systems on HPC clusters. Concurrency and Computation: Practice and Experience, 30(7):??, April 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4407.

Rahman:2013:SIP

[RHRB13] Mustafizur Rahman, Rafiul Hassan, Rajiv Ranjan, and Rajkumar Buyya. Special issue papers: Adaptive workflow scheduling for dynamic grid and cloud computing environment. Concurrency and Computation: Practice and Experience, 25(13):1816–1842, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rossi:2017:SOC

[RHS17] C. Rossi, M. H. Heyi, and F. Scullino. A service oriented cloud-based architecture for mobile geolocated emergency services. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ravn:2013:EIS

[RHT13] Anders P. Ravn and M. Teresa Higuera-Toledano. Editorials: Introduction to the special issue on Java technologies for real-time and embedded systems: JTRES2011. Concurrency and Computation: Practice and Experience, 25(16): 2225–2226, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rehman:2019:MOA

[RHuR^{+}19] Attiqa Rehman, Syed S. Hussain, Zia ur Rehman, Seemal Zia, and Shahaboddin Shamshirband. Multi-objective approach of energy efficient workflow scheduling in cloud environments. Concurrency and Computation: Practice and Experience, 31(8):e4949:1–e4949:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2017:DPA

[RHZ^{+}17] Shenyuan Ren, Ligang He, Huanzhou Zhu, Zhuoer Gu, Wei Song, and Jiandong Shang. Developing power-aware scheduling mechanisms for computing systems virtualized by Xen. Concurrency and Computation: Practice and Experience, 29 (3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ripeanu:2010:SSS

[RIFR10] Matei Ripeanu, Adriana Iamnitchi, Ian Foster, and Anne Rogers. In search of simplicity: a self-organizing group communication overlay. Concurrency and Computation: Practice and Experience, 22(7):788–815, May 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ricci:2018:LSC

[RIP18] Laura Ricci, Alexandru Iosup, and Radu Prodan. Large scale cooperative virtual environments. Concurrency and Computation: Practice and Experience, 30(20):e4878:1–e4878:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rojek:2017:EAM

[RIWS17] Krzysztof Rojek, Aleksandar Ilic, Roman Wyrzykowski, and Leonel Sousa. Energy-aware mechanism for stencil-based MP-DATA algorithm with constraints. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rizk:2004:PIC

[Riz04] Nouhad J. Rizk. Parallelization of IBD computation for determining genetic disease maps. Concurrency and Computation: Practice and Experience, 16(9):933–943, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Risch:2001:DDI

[RJ01] Tore Risch and Vanja Josifovski. Distributed data integration by object-oriented mediator servers. Concurrency and Computation: Practice and Experience, 13(11):933–953, September 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/85012147/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
85012147&PLACEBO=IE.pdf.

Rana:2001:SIH

[RK01] Omer F. Rana and David Kotz. Special issue: High performance agent systems. Concurrency and Computation: Practice and Experience, 13(1):3–4, January 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/77004398/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
77004398&PLACEBO=IE.pdf.

Rymut:2015:RTM

[RK15] Boguslaw Rymut and Bogdan Kwolek. Real-time multiview human pose tracking using graphics processing unitaccelerated particle swarm optimization. Concurrency and Computation: Practice and Experience, 27(6):1551–1563, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rauch:2002:ODL

[RKS02] Felix Rauch, Christian Kurmann, and Thomas M. Stricker. Optimizing the distribution of large data sets in theory and practice. Concurrency and Computation: Practice and Experience, 14(3):165–181, March 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/93513485/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=93513485{\&}PLACEBO=
IE.pdf.

Rong:2016:EBD

[RLC16] Chunming Rong, Lu Liu, and Guolong Chen. Editorials: Big data and smart computing: methodology and practice. Concurrency and Computation: Practice and Experience, 28(11): 3077–3078, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rong:2013:EED

[RLDZ13] Chuitian Rong, Wei Lu, Xiaoyong Du, and Xiao Zhang. Efficient and exact duplicate detection on cloud. Concurrency and Computation: Practice and Experience, 25(15): 2187–2206, October 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reguly:2016:VUM

[RLMG16] I Z. Reguly, Endre László, Gihan R. Mudalige, and Mike B. Giles. Vectorizing unstructured mesh computations for manycore architectures. Concurrency and Computation: Practice and Experience, 28(2):557–577, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rozar:2015:TMS

[RLRG15] Fabien Rozar, Guillaume Latu, Jean Roman, and Virginie Grandgirard. Toward memory scalability of GYSELA code for extreme scale computers. Concurrency and Computation: Practice and Experience, 27(4):994–1009, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Riedel:2009:IWW

[RLS^{+}09] M. Riedel, E. Laure, Th. Soddemann, L. Field, J. P. Navarro, J. Casey, M. Litmaath, J. Ph. Baud, B. Koblitz, C. Catlett, D. Skow, C. Zheng, P. M. Papadopoulos, M. Katz, N. Sharma, O. Smirnova, B. Kónya, P. Arzberger, F. Würthwein, A. S. Rana, T. Martin, M. Wan, V. Welch, T. Rimovsky, S. Newhouse, A. Vanni, Y. Tanaka, Y. Tanimura, T. Ikegami, D. Abramson, C. Enticott, G. Jenkins, R. Pordes, N. Sharma, S. Timm, N. Sharma, G. Moont, M. Aggarwal, D. Colling, O. van der Aa, A. Sim, V. Natarajan, A. Shoshani, J. Gu, S. Chen, G. Galang, R. Zappi, L. Magnoni, V. Ciaschini, M. Pace, V. Venturi, M. Marzolla, P. Andreetto, B. Cowles, S. Wang, Y. Saeki, H. Sato, S. Matsuoka, P. Uthayopas, S. Sriprayoonsakul, O. Koeroo, M. Viljoen, L. Pearlman, S. Pickles, David Wallom, G. Moloney, J. Lauret, J. Marsteller, P. Sheldon, S. Pathak, S. De Witt, J. Mencák, J. Jensen, M. Hodges, D. Ross, S. Phatanapherom, G. Netzer, A. R. Gregersen, M. Jones, S. Chen, P. Kacsuk, A. Streit, D. Mallmann, F. Wolf, Th. Lippert, Th. Delaitre, E. Huedo, and N. Geddes. Interoperation of world-wide production e-Science infrastructures. Concurrency and Computation: Practice and Experience, 21(8):961–990, June 10, 2009. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rubio-Largo:2014:DFG

[RLVRGÁ14] Álvaro Rubio-Largo, Miguel A. Vega-Rodríguez, and David L. González-Álvarez. Designing a fine-grained parallel differential evolution with Pareto tournaments for solving an optical networking problem. Concurrency and Computation: Practice and Experience, 26(11):1908–1934, August 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2015:SRT

[RLZ15] Yongli Ren, Gang Li, and Wanlei Zhou. A survey of recommendation techniques based on offline data processing. Concurrency and Computation: Practice and Experience, 27(15): 3915–3942, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roth:2003:DSH

[RM03] Philip C. Roth and Barton P. Miller. Deep Start: a hybrid strategy for automated performance problem searches. Concurrency and Computation: Practice and Experience, 15(11– 12):1027–1046, September 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ropars:2011:AOD

[RM11] Thomas Ropars and Christine Morin. Active optimistic and distributed message logging for message-passing applications. Concurrency and Computation: Practice and Experience, 23 (17):2167–2178, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

R:2019:MBS

[RM19] Gayathiri N. R. and Natarajan A. M. MapReduce-based storage and indexing for big health data. Concurrency and Computation: Practice and Experience, 31(14):e4854:1–e4854:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ravi:2012:CRS

[RMCA12] Vignesh T. Ravi, Wenjing Ma, David Chiu, and Gagan Agrawal. Compiler and runtime support for enabling reduction computations on heterogeneous systems. Concurrency and Computation: Practice and Experience, 24(5):463–480, April 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rubio-Montero:2015:NPJ

[RMCHMG15] A. J. Rubio-Montero, F. Castejón, E. Huedo, and R. Mayo-García. A novel pilot job approach for improving the execution of distributed codes: application to the study of ordering in collisional transport in fusion plasmas. Concurrency and Computation: Practice and Experience, 27(13): 3220–3244, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reyes:2007:PCI

[RMCN^{+}07] S. Reyes, C. Muñoz-Caro, A. Niño, R. M. Badia, and J. M. Cela. Performance of computationally intensive parameter sweep applications on Internet-based Grids of computers: the mapping of molecular potential energy hypersurfaces. Concurrency and Computation: Practice and Experience, 19(4): 463–481, March 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rodriguez:2010:CCA

[RMG^{+}10] Gabriel Rodríguez, María J. Martín, Patricia González, Juan Touriño, and Ramón Doallo. CPPC: a compiler-assisted tool for portable checkpointing of message-passing applications. Concurrency and Computation: Practice and Experience, 22 (6):749–766, April 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raghav:2013:SIP

[RMP^{+}13a] Shivani Raghav, Andrea Marongiu, Christian Pinto, Martino Ruggiero, David Atienza, and Luca Benini. Special issue papers: SIM in G-1 k: a thousand-core simulator running on general-purpose graphical processing units. Concurrency and Computation: Practice and Experience, 25(10):1443–1461, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Riteau:2013:SEL

[RMP13b] Pierre Riteau, Chritine Morin, and Thierry Priol. Shrinker: efficient live migration of virtual clusters over wide area networks. Concurrency and Computation: Practice and Experience, 25(4):541–555, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Robertsen:2019:HPS

[RMW19] Fredrik Robertsén, Keijo Mattila, and Jan Westerholm. Highperformance SIMD implementation of the lattice-Boltzmann method on the Xeon Phi processor. Concurrency and Computation: Practice and Experience, 31(13):e5072:1–e5072:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Riveni:2019:APS

[RNAD19] Mirela Riveni, Tien-Dung Nguyen, Mehmet S. Aktas, and Schahram Dustdar. Application of provenance in social computing: a case study. Concurrency and Computation: Practice and Experience, 31(3):e4894:1–e4894:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramesh:2017:MFM

[RNJM17] Varun Ramesh, Shivanee Nagarajan, Jason J. Jung, and Saswati Mukherjee. Max-flow min-cut algorithm with application to road networks. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roper:2012:ASTa

[RO12a] Matthew D. Roper and Ronald A. Olsson. Applicationspecific thread schedulers for distributed applications. Concurrency and Computation: Practice and Experience, 24(3): 260–280, March 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roper:2012:ASTb

[RO12b] Matthew D. Roper and Ronald A. Olsson. Applicationspecific thread schedulers for Internet server applications. Concurrency and Computation: Practice and Experience, 24 (3):281–304, March 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2007:SOA

[ROA^{+}07] X. Ren, M. Ong, G. Allan, V. Kadirkamanathan, H. A. Thompson, and P. J. Fleming. Service-oriented architecture on the Grid for integrated fault diagnostics. Concurrency and Computation: Practice and Experience, 19(2):223–234, February 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rojek:2019:MLM

[Roj19] Krzysztof Rojek. Machine learning method for energy reduction by utilizing dynamic mixed precision on GPU-based supercomputers. Concurrency and Computation: Practice and Experience, 31(6):e4644:1–e4644:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rais:2018:QIS

[ROQL18] Issam Raïs, Anne-Cécile Orgerie, Martin Quinson, and Laurent Lefèvre. Quantifying the impact of shutdown techniques for energy-efficient data centers. Concurrency and Computation: Practice and Experience, 30(17):e4471:1–e4471:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rossi:2006:XFD

[Ros06] D. Rossi. X-Folders: documents on the move. Concurrency and Computation: Practice and Experience, 18(4):409–425, April 10, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramani:2019:IEG

[RP19] R. Geetha Ramani and K. Priya. Improvised emotion and genre detection for songs through signal processing and genetic algorithm. Concurrency and Computation: Practice and Experience, 31(14):e5065:1–e5065:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rabl:2008:DAS

[RPK08] T. Rabl, M. Pfeffer, and H. Kosch. Dynamic allocation in a self-scaling cluster database. Concurrency and Computation: Practice and Experience, 20(17):2025–2038, December 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rocha:2017:TAT

[RPRG17] Rodrigo C. O. Rocha, Alyson D. Pereira, Luiz Ramos, and Luís F. W. Góes. TOAST: Automatic tiling for iterative stencil computations on GPUs. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rose:2001:JAP

[RR01] Eva Rose and Kristoffer Høgsbro Rose. Java access protection through typing. Concurrency and Computation: Practice and Experience, 13(13):1125–1132, November 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/88011340/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
88011340&PLACEBO=IE.pdf.

Romero:2011:CSU

[RR11] Eloy Romero and Jose E. Roman. Computing subdominant unstable modes of turbulent plasma with a parallel Jacobi– Davidson eigensolver. Concurrency and Computation: Practice and Experience, 23(17):2179–2191, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rauber:2015:MAE

[RR15] Thomas Rauber and Gudula Rünger. Modeling and analyzing the energy consumption of fork-join-based task parallel programs. Concurrency and Computation: Practice and Experience, 27(1):211–236, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rauber:2019:SSP

[RR19] Thomas Rauber and Gudula Rünger. A scheduling selection process for energy-efficient task execution on DVFS processors. Concurrency and Computation: Practice and Experience, 31(19):e5043:1–e5043:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rahman:2011:TSA

[RRBB11] Mustafizur Rahman, Rajiv Ranjan, Rajkumar Buyya, and Boualem Benatallah. A taxonomy and survey on autonomic management of applications in grid computing environments. Concurrency and Computation: Practice and Experience, 23 (16):1990–2019, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rauber:2004:GSP

[RRR04] Thomas Rauber, Robert Reilein, and Gudula Rünger. Group-SPMD programming with orthogonal processor groups. Concurrency and Computation: Practice and Experience, 16(2– 3):173–195, February/March 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roche:2015:CDS

[RRR15] Scott T. Roche, Arnold L. Rosenberg, and Rajmohan Rajaraman. On constructing DAG-schedules with large areas. Concurrency and Computation: Practice and Experience, 27 (16):4107–4121, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rygg:2008:GIE

[RRWS08] Asbjørn Rygg, Paul Roe, On Wong, and Jiro Sumitomo. GPFlow: an intuitive environment for Web-based scientific workflow. Concurrency and Computation: Practice and Experience, 20(4):393–408, March 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rosu:2007:ITO

[RS07] Grigore Roşu and Koushik Sen. An instrumentation technique for online analysis of multithreaded programs. Concurrency and Computation: Practice and Experience, 19(3): 311–325, March 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Repplinger:2011:SPG

[RS11] Michael Repplinger and Philipp Slusallek. Stream processing on GPUs using distributed multimedia middleware. Concurrency and Computation: Practice and Experience, 23(7):669– 680, May 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ravn:2012:SCJ

[RS12] Anders P. Ravn and Martin Schoeberl. Safety-critical Java with cyclic executives on chip-multiprocessors. Concurrency and Computation: Practice and Experience, 24(8):772–788, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2013:EGE

[RS13] Pinyi Ren and Zhou Su. Editorial: Guest editorial: advances in networking technologies for wireless Internet. Concurrency and Computation: Practice and Experience, 25(9):1031–1033, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rammig:2016:OBC

[RS16] Franz Rammig and Katharina Stahl. Online behavior classification for anomaly detection in self-x real-time systems. Concurrency and Computation: Practice and Experience, 28(14): 3751–3772, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reano:2015:IUE

[RSC^{+}15] Carlos Reaño, Federico Silla, Adrián Castelló, Antonio J. Peña, Rafael Mayo, Enrique S. Quintana-Ortí, and José Duato. Improving the user experience of the rCUDA remote GPU virtualization framework. Concurrency and Computation: Practice and Experience, 27(14):3746–3770, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rao:2019:HGK

[RSKK19] R. Velumadhava Rao, K. Selvamani, S. Kanimozhi, and A. Kannan. Hierarchical group key management for secure data sharing in a cloud-based environment. Concurrency and Computation: Practice and Experience, 31(12):e4866:1– e4866:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reeve:2001:PVS

[RSM01] J. S. Reeve, A. D. Scurr, and J. H. Merlin. Parallel versions of Stone’s strongly implicit algorithm. Concurrency and Computation: Practice and Experience, 13(12):1049–1062, October 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/85512009/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
85512009&PLACEBO=IE.pdf.

Rodriguez-Sanchez:2012:OHA

[RSMFE^{+}12] Rafael Rodríguez-Sánchez, José Luis Martínez, Gerardo Fernández-Escribano, José L. Sánchez, José M. Claver, and Pedro Diaz. Optimizing H.264/AVC interprediction on a GPU-based framework. Concurrency and Computation: Practice and Experience, 24(14):1607–1624, September 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rouberte:2019:DDD

[RSN^{+}19] Leandro Rouberte, Alexandre C. Sena, Alexandre S. Nery, Leandro A. J. Marzulo, Tiago A. O. Alves, and Felipe M. G. França. DF–DTM: Dynamic task memoization and reuse in dataflow. Concurrency and Computation: Practice and Experience, 31(18):e4937:1–e4937:??, September 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ravaei:2017:TCD

[RSPV17] Bahman Ravaei, Masoud Sabaei, Hossein Pedram, and Shahrokh Valaee. Targeted content dissemination in mobile social networks taking account of resource limitation. Concurrency and Computation: Practice and Experience, 29(18): ??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roblitz:2006:RRF

[RSR06] T. Röblitz, F. Schintke, and A. Reinefeld. Resource reservations with fuzzy requests. Concurrency and Computation:

Practice and Experience, 18(13):1681–1703, November 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rasmussen:2006:BLG

[RSSM06] C. E. Rasmussen, M. J. Sottile, S. S. Shende, and A. D. Malony. Bridging the language gap in scientific computing: the Chasm approach. Concurrency and Computation: Practice and Experience, 18(2):151–162, February 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roy:2005:FWS

[RSTV05] Indrajit G. Roy, Mrinal K. Sen, and Carlos Torres-Verdin. Full waveform seismic inversion using a distributed system of computers. Concurrency and Computation: Practice and Experience, 17(11):1365–1385, September 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Romei:2007:KGG

[RSTV07] Andrea Romei, Matteo Sciolla, Franco Turini, and Marlis Valentini. KDDML-G: a grid-enabled knowledge discovery system. Concurrency and Computation: Practice and Experience, 19(13):1785–1809, September 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramos:2015:NCS

[RTET15] Sabela Ramos, Guillermo L. Taboada, Roberto R. Expósito, and Juan Touriño. Nonblocking collectives for scalable Java communications. Concurrency and Computation: Practice and Experience, 27(5):1169–1187, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rizvandi:2013:SIP

[RTMZ13] Nikzad Babaii Rizvandi, Javid Taheri, Reza Moraveji, and Albert Y. Zomaya. Special issue papers: A study on using uncertain time series matching algorithms for MapReduce applications. Concurrency and Computation: Practice and Experience, 25(12):1699–1718, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reuillon:2012:PSS

[RTPPH12] Romain Reuillon, Mamadou K. Traore, Jonathan Passerat-Palmbach, and David R. C. Hill. Parallel stochastic simulations with rigorous distribution of pseudo-random numbers with DistMe: Application to life science simulations. Concurrency and Computation: Practice and Experience, 24(7):723– 738, May 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/cpe.1883/abstract.

Ruan:2015:DEW

[Rua15] Chun Ruan. Designing and evaluating weighted delegatable authorizations. Concurrency and Computation: Practice and Experience, 27(15):3877–3891, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rundle:2010:PAS

[Run10] John B. Rundle. Proceedings of the 6th ACES Symposium, May 11–16, 2008, Cairns, Australia: Introduction to the special issue. Concurrency and Computation: Practice and Experience, 22(12):1517–1519, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajan:2012:ADA

[RVD^{+}12] Mahesh Rajan, Courtenay T. Vaughan, Doug W. Doerfler, Richard F. Barrett, Paul T. Lin, Kevin T. Pedretti, and K. Scott Hemmert. Application-driven analysis of two generations of capability computing: the transition to multicore processors. Concurrency and Computation: Practice and Experience, 24(18):2404–2420, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramirez-Velarde:2010:CCH

[RVRD10] Raul V. Ramirez-Velarde and Ramón M. Rodríguez-Dagnino. From commodity computers to high-performance environments: scalability analysis using self-similarity, large deviations and heavy-tails. Concurrency and Computation: Practice and Experience, 22(11):1494–1515, August 10, 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rodriguez-Vazquez:2017:PSS

[RVVPD^{+}17] Juan José Rodríguez-Vázquez, José Luis Vázquez-Poletti, Carlos Delgado, Andrea Bulgarelli, and Miguel Cárdenas-Montes. Performance study of a signal-extraction algorithm using different parallelisation strategies for the Cherenkov Telescope Array’s real-time-analysis software. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ritson:2010:POA

[RW10] C. G. Ritson and P. H. Welch. A process-oriented architecture for complex system modelling. Concurrency and Computation: Practice and Experience, 22(8):965–980, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramakrishnan:2002:PEM

[RWK^{+}02] N. Ramakrishnan, L. T. Watson, D. G. Kafura, C. J. Ribbens, and C. A. Shaffer. Programming environments for multidisciplinary Grid communities. Concurrency and Computation: Practice and Experience, 14(13–15):1241–1273, November/ December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rojek:2017:SAS

[RWK17] Krzysztof Rojek, Roman Wyrzykowski, and Lukasz Kuczynski. Systematic adaptation of stencil-based 3D MPDATA to GPU architectures. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2019:PSP

[RZL^{+}19] Xiaoli Ren, Juan Zhao, Xiaoyong Li, Kaijun Ren, Junqiang Song, and Difu Sun. PAGCM: a scalable parallel spectralbased atmospheric general circulation model. Concurrency and Computation: Practice and Experience, 31(20):e5290:1– e5290:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Subramanian:2019:SAF

[SA19] Nalini Subramanian and J. Andrews. Strong authentication framework using statistical approach for cloud environments. Concurrency and Computation: Practice and Experience, 31 (12):e4870:1–e4870:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sankaran:2015:GAB

[SAB15] Ramanan Sankaran, Jordan Angel, and W. Michael Brown. Genetic algorithm based task reordering to improve the performance of batch scheduled massively parallel scientific applications. Concurrency and Computation: Practice and Experience, 27(17):4763–4783, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Santos:2007:RAA

[SAC^{+}07] Robson Santos, Alisson Andrade, Walfredo Cirne, Francisco Brasileiro, and Nazareno Andrade. Relative autonomous accounting for peer-to-peer Grids. Concurrency and Computation: Practice and Experience, 19(14):1937–1954, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shin:2004:RBI

[SACJ04] Dongwan Shin, Gail-Joon Ahn, Sangrae Cho, and Seunghun Jin. A role-based infrastructure management system: design and implementation. Concurrency and Computation: Practice and Experience, 16(11):1121–1141, September 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sanchez-Artigas:2018:SAB

[SACRGL18] Marc Sánchez-Artigas, Cristian Cotes, Marc R. Rodriguez, and Pedro García-Lopez. StackSync: Attribute-based data sharing in file synchronization services. Concurrency and Computation: Practice and Experience, 30(8):??, April 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4391.

Schmidt:2013:ISM

[SAD13] Bertil Schmidt, Hans Aribowo, and Hoang-Vu Dang. Iterative sparse matrix–vector multiplication for accelerating the block Wiedemann algorithm over GF(2) on multi-graphics processing unit systems. Concurrency and Computation: Practice and Experience, 25(4):586–603, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Serbanescu:2016:DPO

[SAdB^{+}16] V. Serbanescu, K. Azadbakht, F. de Boer, C. Nagarajagowda, and B. Nobakht. A design pattern for optimizations in data intensive applications using ABS and JAVA 8. Concurrency and Computation: Practice and Experience, 28(2):374–385, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Subratie:2017:GDC

[SAM^{+}17] Kensworth C. Subratie, Saumitra Aditya, Srinivas Mahesula, Renato Figueiredo, Cayelan C. Carey, and Paul C. Hanson. GRAPLEr: a distributed collaborative environment for lake ecosystem modeling that integrates overlay networks, highthroughput computing, and WEB services. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Serrai:2019:HDQ

[SAMS19] Walid Serrai, Abdelkrim Abdelli, Lynda Mokdad, and Ashref Serrai. How to deal with QoS value constraints in MCDM based Web service selection. Concurrency and Computation: Practice and Experience, 31(24):e4512:1–e4512:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schulze:2008:SIM

[SANB08] Bruno Schulze, David Abramson, Radha Nandkumar, and Rajkumar Buyya. Special issue: Middleware for Grid Computing. Concurrency and Computation: Practice and Experience, 20(9):1025–1027, June 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sarbazi-Azad:2004:TMR

[SAOKM04] H. Sarbazi-Azad, M. Ould-Khaoua, and L. M. Mackenzie. Towards a more realistic comparative analysis of multicomputer networks. Concurrency and Computation: Practice and Experience, 16(13):1271–1289, November 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soldado:2016:ECM

[SAP16] Fábio Soldado, Fernando Alexandre, and Hervé Paulino. Execution of compound multi-kernel OpenCL computations in multi-CPU/multi-GPU environments. Concurrency and Computation: Practice and Experience, 28(3):768–787, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sfaxi:2013:IFC

[SARL13] Lilia Sfaxi, Takoua Abdellatif, Riadh Robbana, and Yassine Lakhnech. Information flow control of component-based distributed systems. Concurrency and Computation: Practice and Experience, 25(2):161–179, February 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Srivastava:2017:RRA

[SB17] Srishti Srivastava and Ioana Banicescu. Robust resource allocations through performance modeling with stochastic process algebra. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sirbu:2018:DDA

[SB18] Alina Sîrbu and Ozalp Babaoglu. A data-driven approach to modeling power consumption for a hybrid supercomputer. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4410.

Singh:2019:IIG

[SB19a] Hari Singh and Seema Bawa. An improved integrated Grid and MapReduce–Hadoop architecture for spatial data: Hilbert TGS R-Tree-based IGSIM. Concurrency and Computation: Practice and Experience, 31(17):e5202:1–e5202:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sree:2019:SLS

[SB19b] Thankaraja Raja Sree and Somasundaram Mary Saira Bhanu. Secure logging scheme for forensic analysis in cloud. Concurrency and Computation: Practice and Experience, 31 (15):e5143:1–e5143:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sinnott:2015:AUR

[SBB^{+}15] Richard O. Sinnott, Christopher Bayliss, Andrew Bromage, Gerson Galang, Guido Grazioli, Philip Greenwood, Angus Macaulay, Luca Morandini, Ghazal Nogoorani, Marcos Nino-Ruiz, Martin Tomko, Christopher Pettit, Muhammad Sarwar, Robert Stimson, William Voorsluys, and Ivo Widjaja. The Australia urban research gateway. Concurrency and Computation: Practice and Experience, 27(2):358–375, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schmidt:2007:COA

[SBBE07] Rainer Schmidt, Siegfried Benkner, Ivona Brandic, and Gerhard Engelbrecht. Component-oriented application construction for a Web service-based Grid. Concurrency and Computation: Practice and Experience, 19(5):637–650, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Serrano:2015:CSX

[SBC15] Estefania Serrano, Javier Garcia Blas, and Jesus Carretero. A comparative study of an X-ray tomography reconstruction algorithm in accelerated and cloud computing systems. Concurrency and Computation: Practice and Experience, 27(18): 5538–5556, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Silva:2015:MAL

[SBDP15] Juliana M. N. Silva, Cristina Boeres, Lúcia M. A. Drummond, and Artur A. Pessoa. Memory aware load balance strategy on a parallel branch-and-bound application. Concurrency and Computation: Practice and Experience, 27(5): 1122–1144, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shahand:2015:DCN

[SBJ^{+}15] Shayan Shahand, Ammar Benabdelkader, Mohammad Mahdi Jaghoori, Mostapha al Mourabit, Jordi Huguet, Matthan W. A. Caan, Antoine H. C. van Kampen, and Sílvia D. Olabarriaga. A data-centric neuroscience gateway: design, implementation, and experiences. Concurrency and Computation: Practice and Experience, 27(2):489–506, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schulze:2012:MCS

[SBP12] Bruno Schulze, Rajkumar Buyya, and Fabio Porto. Middleware for Clouds and e-Science. Concurrency and Computation: Practice and Experience, 24(13):1393–1396, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sahasrabudhe:2019:NFR

[SBS19] Damodar Sahasrabudhe, Martin Berzins, and John Schmidt. Node failure resiliency for Uintah without checkpointing. Concurrency and Computation: Practice and Experience, 31 (20):e5340:1–e5340:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Speer:2007:ERM

[SC07a] Anh Phan Speer and Ing-Ray Chen. Effect of redundancy on the mean time to failure of wireless sensor networks. Concurrency and Computation: Practice and Experience, 19(8): 1119–1128, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Subramaniam:2007:MEV

[SC07b] Mahadevan Subramaniam and Patrick Conway. A methodology for early validation of cache coherence protocols based on relational databases. Concurrency and Computation: Practice and Experience, 19(3):355–368, March 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sujitha:2019:HSP

[SC19] V. Sujitha and D. Chitra. Highly secure palmprint based biometric template using fuzzy vault. Concurrency and Computation: Practice and Experience, 31(12):e4513:1–e4513:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shafi:2009:CSJ

[SCBH09] Aamir Shafi, Bryan Carpenter, Mark Baker, and Aftab Hussain. A comparative study of Java and C performance in two large-scale parallel applications. Concurrency and Computation: Practice and Experience, 21(15):1882–1906, October 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seelam:2010:WPC

[SCC^{+}10] Seetharami Seelam, I-Hsin Chung, Guojing Cong, Hui-Fang Wen, and David Klepacki. Workload performance characterization of DARPA HPCS benchmarks. Concurrency and Computation: Practice and Experience, 22(4):441–461, March 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2019:NMD

[SCGZ19] Qingguo Shen, Qinghua Chen, Jing Gong, and Feng Zhou. Networking of multi-domain unified communications systems: Structure design and performance evaluation. Concurrency and Computation: Practice and Experience, 31(9):e4677:1– e4677:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schreiber:2002:ISE

[Sch02] Andreas Schreiber. The integrated simulation environment TENT. Concurrency and Computation: Practice and Experience, 14(13–15):1553–1568, November/December 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schirmer:2004:AJP

[Sch04] Norbert Schirmer. Analysing the Java package/access concepts in Isabelle/HOL. Concurrency and Computation: Practice and Experience, 16(7):689–706, June 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2015:DSL

[SCLK15] Rong Sun, Xin Cai, Jingwei Liu, and Kyung Sup Kwak. Distributed SR–LDPC codes over multiple-access relay channel and its applications in cloud storage. Concurrency and Computation: Practice and Experience, 27(8):2064–2077, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shan:2019:NIC

[SCLL19] Chun Shan, Jun Cai, Yan Liu, and JianZhen Luo. Node importance to community based caching strategy for information centric networking. Concurrency and Computation: Practice and Experience, 31(21):e4797:1–e4797:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schulze:2007:SIM

[SCNH07] Bruno Schulze, Geoff Coulson, Radha Nandkumar, and Peter Henderson. Special issue: Middleware for Grid computing: a ‘possible future’. Concurrency and Computation: Practice and Experience, 19(14):1879–1884, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

shree:2019:ERC

[sCR19] S. Raja shree, A. Chilambu Chelvan, and M. Rajesh. An efficient RSA cryptosystem by applying cuckoo search optimization algorithm. Concurrency and Computation: Practice and Experience, 31(12):e4845:1–e4845:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sanchez:2011:SMA

[SCRV11] Friman Sánchez, Felipe Cabarcas, Alex Ramirez, and Mateo Valero. Scalable multicore architectures for long DNA sequence comparison. Concurrency and Computation: Practice and Experience, 23(17):2205–2219, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sarkar:2017:SCK

[SCS17a] Pinaki Sarkar, Morshed Uddin Chowdhury, and Kouichi Sakurai. Secure combinatorial key predistribution scheme for sensor networks by regulating frequencies: magneto optic sensors. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shukla:2017:RIB

[SCS17b] Anshu Shukla, Shilpa Chaturvedi, and Yogesh Simmhan. RIoTBench: An IoT benchmark for distributed stream processing systems. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sulistio:2008:TMS

[SCV^{+}08] Anthony Sulistio, Uros Cibej, Srikumar Venugopal, Borut Robic, and Rajkumar Buyya. A toolkit for modelling and simulating data Grids: an extension to GridSim. Concurrency and Computation: Practice and Experience, 20(13): 1591–1609, September 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seymour:2003:ATF

[SD03] Keith Seymour and Jack Dongarra. Automatic translation of Fortran to JVM bytecode. Concurrency and Computation: Practice and Experience, 15(3–5):207–222, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL http://www.netlib.org/netlib/
utk/people/JackDongarra/PAPERS/f2jreport.pdf.

Singh:2011:IRP

[SD11a] Gurmeet Singh and Ewa Deelman. The interplay of resource provisioning and workflow optimization in scientific applications. Concurrency and Computation: Practice and Experience, 23(16):1969–1989, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2011:SIS

[SD11b] Xiaoping Sun and Xin Dong. Special issue: Semantics, knowledge and Grids. Concurrency and Computation: Practice and Experience, 23(9):863–865, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2015:SAS

[SD15] Fengguang Song and Jack Dongarra. A scalable approach to solving dense linear algebra problems on hybrid CPU–GPU systems. Concurrency and Computation: Practice and Experience, 27(14):3702–3723, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schuchardt:2002:EPS

[SDB02] Karen Schuchardt, Brett Didier, and Gary Black. Ecce — a problem-solving environment’s evolution toward Grid services and a Web architecture. Concurrency and Computation: Practice and Experience, 14(13–15):1221–1239, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schoeberl:2017:SCJ

[SDH^{+}17] Martin Schoeberl, Andreas Engelbredt Dalsgaard, René Rydhof Hansen, Stephan E. Korsholm, Anders P. Ravn, Juan Ricardo Rios Rivas, Tórur Biskopstø Strøm, Hans Søndergaard, Andy Wellings, and Shuai Zhao. Safety-critical Java for embedded systems. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Silva:2016:ARR

[SdOVM16] Vítor Silva, Daniel de Oliveira, Patrick Valduriez, and Marta Mattoso. Analyzing related raw data files through dataflows. Concurrency and Computation: Practice and Experience, 28 (8):2528–2545, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soares:2018:UHD

[SdSL18] Thiago Marques Soares, Rodrigo Weber dos Santos, and Marcelo Lobosco. On the use of HCM to develop a resource allocation algorithm for heterogeneous clusters. Concurrency and Computation: Practice and Experience, 30(24):e4899:1– e4899:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Spelt:2001:CMS

[SE01] David Spelt and Susan Even. Compensation methods to support cooperative applications: a case study in automated verification of schema requirements for an advanced transaction model. Concurrency and Computation: Practice and Experience, 13(11):1013–1032, September 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/85012148/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
85012148&PLACEBO=IE.pdf.

Szydlarski:2014:PSH

[SEF^{+}14] Mikolaj Szydlarski, Pierre Esterie, Joel Falcou, Laura Grigori, and Radek Stompor. Parallel spherical harmonic transforms on heterogeneous architectures (graphics processing units/ multi-core CPUs). Concurrency and Computation: Practice and Experience, 26(3):683–711, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sebastiao:2015:IPA

[SER15] Nuno Sebastião, Gustavo Encarnação, and Nuno Roma. Implementation and performance analysis of efficient index structures for DNA search algorithms in parallel platforms. Concurrency and Computation: Practice and Experience, 27 (9):2351–2368, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schulze:2010:SIA

[SF10] Bruno Schulze and Geoffrey C. Fox. Special issue: Advanced scheduling strategies and Grid programming environments. Concurrency and Computation: Practice and Experience, 22 (3):233–240, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Silla:2016:EHC

[SF16] Federico Silla and Holger Fröning. Editorials: Heterogeneous cluster architectures and applications. Concurrency and Computation: Practice and Experience, 28(8):2319–2321, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Selimi:2016:PED

[SFCAV16] Mennan Selimi, Felix Freitag, Llorenç Cerdà-Alabern, and Luís Veiga. Performance evaluation of a distributed storage service in community network clouds. Concurrency and Computation: Practice and Experience, 28(11):3131–3148, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Smari:2013:EHP

[SFH13] Waleed W. Smari, Sandro Fiore, and David Hill. Editorial: High performance computing and simulation: architectures, systems, algorithms, technologies, services, and applications.

Concurrency and Computation: Practice and Experience, 25 (10):1313–1318, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schmidt:2004:DDB

[SFLS04] Bertil Schmidt, Lin Feng, Amey Laud, and Yusdi Santoso. Development of distributed bioinformatics applications with GMP. Concurrency and Computation: Practice and Experience, 16(9):945–959, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Smari:2012:SIA

[SFN12] Waleed W. Smari, Sandro Fiore, and Mads Nygaard. Special Issue on Advances in High Performance Computing and Simulation. Concurrency and Computation: Practice and Experience, 24(7):661–662, May 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Smari:2015:ERD

[SFT15] Waleed W. Smari, Sandro Fiore, and Carsten Trinitis. Editorials: Recent developments in high-performance computing and simulation: distributed systems, architectures, algorithms, and applications. Concurrency and Computation: Practice and Experience, 27(9):2191–2195, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Senger:2016:EPD

[SG16] Hermes Senger and Claudio Geyer. Editorial: Parallel and distributed computing for Big Data applications. Concurrency and Computation: Practice and Experience, 28(8): 2412–2415, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2018:MOO

[SG18] Lingfan Sun and Lin Gui. Multi-objective optimization algorithm of adaptive entropy wireless network based on embedded DSP. Concurrency and Computation: Practice and Experience, 30(22):e4653:1–e4653:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sahuquillo:2019:THP

[SG19] Jesús Escudero Sahuquillo and Pedro Javier Garcia. Trends in high-performance interconnection networks in the exascale and big-data era (HiPINEB 2017). Concurrency and Computation: Practice and Experience, 31(2):e5041:1–e5041:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Senger:2016:BCS

[SGCA^{+}16] Hermes Senger, Veronica Gil-Costa, Luciana Arantes, Cesar A. C. Marcondes, Mauricio Marín, Liria M. Sato, and Fabrício A. B. da Silva. BSP cost and scalability analysis for MapReduce operations. Concurrency and Computation: Practice and Experience, 28(8):2503–2527, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2009:SKP

[SGCG09] An-Ni Shen, Song Guo, Hung Yu Chien, and Minyi Guo. A scalable key pre-distribution mechanism for large-scale wireless sensor networks. Concurrency and Computation: Practice and Experience, 21(10):1373–1387, July 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sellami:2015:DSB

[SGD15] Mohamed Sellami, Walid Gaaloul, and Bruno Defude. A decentralized and service-based solution for data mediation: the case for data providing service compositions. Concurrency and Computation: Practice and Experience, 27(6):1427–1444, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Smith:2018:MIE

[SGD^{+}18] Cameron W. Smith, Brian Granzow, Gerrett Diamond, Daniel Ibanez, Onkar Sahni, Kenneth E. Jansen, and Mark S. Shephard. In-memory integration of existing software components for parallel adaptive unstructured mesh workflows. Concurrency and Computation: Practice and Experience, 30(18):e4510:1–e4510:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Syed:2007:SSD

[SGG07] Jameel Syed, Moustafa Ghanem, and Yike Guo. Supporting scientific discovery processes in Discovery Net. Concurrency and Computation: Practice and Experience, 19(2):167–179, February 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sapegin:2017:EMS

[SGJ^{+}17] Andrey Sapegin, Marian Gawron, David Jaeger, Feng Cheng, and Christoph Meinel. Evaluation of in-memory storage engine for machine learning analysis of security events. Concurrency and Computation: Practice and Experience, 29(2): ??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Souza:2017:PEE

[SGL^{+}17] S. M. P. C. Souza, R. F. Gonçalves, E. Leonova, R. S. Puttini, and A. C. A. Nascimento. Privacy-ensuring electronic health records in the cloud. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sarkar:2018:TND

[SGM18] Santonu Sarkar, Ajai V George, and Sankar Manoj. Thrust2D: a new design abstraction framework for structured grid class of algorithms. Concurrency and Computation: Practice and Experience, 30(19):e4740:1–e4740:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Spellacy:2019:PAP

[SGR19] Louise Spellacy, Darach Golden, and Ivan Rungger. Performance analysis of a pairwise method for partial inversion of complex block tridiagonal matrices. Concurrency and Computation: Practice and Experience, 31(19):e4918:1–e4918:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schuchardt:2008:ACM

[SGSC08] Karen L. Schuchardt, Tara Gibson, Eric Stephan, and George Chin, Jr. Applying content management to automated provenance capture. Concurrency and Computation: Practice and Experience, 20(5):541–554, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Simao:2012:CER

[SGV12] José Simão, Tiago Garrochinho, and Luís Veiga. A checkpointing-enabled and resource-aware Java Virtual Machine for efficient and robust e-Science applications in grid environments. Concurrency and Computation: Practice and Experience, 24(13):1421–1442, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shah:2015:EER

[Sha15] Sayed Chhattan Shah. Energy efficient and robust allocation of interdependent tasks on mobile ad hoc computational grid. Concurrency and Computation: Practice and Experience, 27 (5):1226–1254, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sodsong:2016:DPB

[SHC^{+}16] Wasuwee Sodsong, Jingun Hong, Seongwook Chung, Yeongkyu Lim, Shin-Dug Kim, and Bernd Burgstaller. Dynamic partitioning-based JPEG decompression on heterogeneous multicore architectures. Concurrency and Computation: Practice and Experience, 28(2):517–536, February 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Severance:2007:USC

[SHG^{+}07] Charles Severance, Joseph Hardin, Glenn Golden, Robert Crouchley, Adrian Fish, Tom Finholt, Beth Kirschner, Jim Eng, and Rob Allan. Using the Sakai collaborative toolkit in e-Research applications. Concurrency and Computation: Practice and Experience, 19(12):1643–1652, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stanberry:2014:VPS

[SHH^{+}14] Larissa Stanberry, Roger Higdon, Winston Haynes, Natali Kolker, William Broomall, Saliya Ekanayake, Adam Hughes, Yang Ruan, Judy Qiu, Eugene Kolker, and Geoffrey Fox. Visualizing the Protein Sequence Universe. Concurrency and Computation: Practice and Experience, 26(6):1313–1325, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sui:2014:DOS

[SHP14] Zhiquan Sui, Neil Harvey, and Shrideep Pallickara. On the distributed orchestration of stochastic discrete event simulations. Concurrency and Computation: Practice and Experience, 26(11):1889–1907, August 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shao:2013:EER

[SHST13] Shihai Shao, Liang Han, Ying Shen, and Youxi Tang. Energyefficient radio remote units placement for single-user uplink in C-RAN. Concurrency and Computation: Practice and Experience, 25(9):1073–1080, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schoeberl:2011:ISI

[SHT11] Martin Schoeberl and M. Teresa Higuera-Toledano. Introduction to the special issue: JTRES 2009. Concurrency and Computation: Practice and Experience, 23(14):1607–1608, September 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schenck:2017:CPH

[SHT^{+}17] Wolfram Schenck, Michael Horst, Tim Tiedemann, Sergius Gaulik, and Ralf Möller. Comparing parallel hardware architectures for visually guided robot navigation. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2007:PPA

[SİM^{+}07] Karan Singh, Engin İpek, Sally A. McKee, Bronis R. de Supinski, Martin Schulz, and Rich Caruana. Predicting parallel application performance via machine learning approaches. Concurrency and Computation: Practice and Experience, 19(17):2219–2235, December 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2010:DVP

[Sin10] S. Singh. Design and verification of peripheral control circuits in Esterel. Concurrency and Computation: Practice and Experience, 22(8):981–994, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seo:2002:HJE

[SIOS02] Yoshiki Seo, Hidetoshi Iwashita, Hiroshi Ohta, and Hitoshi Sakagami. HPF/JA: extensions of High Performance Fortran for accelerating real-world applications. Concurrency and Computation: Practice and Experience, 14(8–9):555– 573, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/95016127/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=95016127{\&}PLACEBO=IE.pdf.

Silla:2017:BRG

[SIRP17] Federico Silla, Sergio Iserte, Carlos Reaño, and Javier Prades. On the benefits of the remote GPU virtualization mechanism: The rCUDA case. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seshadri:2019:DEP

[SIS19] Karthick Seshadri, K. Viswanathan Iyer, and Mercy Shalinie S. Design and evaluation of a parallel document clustering algorithm based on hierarchical latent semantic analysis. Concurrency and Computation: Practice and Experience, 31(13): e5094:1–e5094:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sangat:2018:SDM

[SIST18] Prajwol Sangat, Maria Indrawan-Santiago, and David Taniar. Sensor data management in the cloud: Data storage, data ingestion, and data retrieval. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sakthitharan:2019:EEC

[SJ19] S Sakthitharan and S Jayashri. Establishing an emergency communication network and optimal path using multiple autonomous rover robots. Concurrency and Computation: Practice and Experience, 31(14):e4636:1–e4636:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Salehi:2014:RPB

[SJB14] Mohsen Amini Salehi, Bahman Javadi, and Rajkumar Buyya. Resource provisioning based on preempting virtual machines in distributed systems. Concurrency and Computation: Practice and Experience, 26(2):412–433, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Santander-Jimenez:2017:APP

[SJISVR17] Sergio Santander-Jiménez, Aleksandar Ilic, Leonel Sousa, and Miguel A. Vega-Rodríguez. Accelerating the phylogenetic parsimony function on heterogeneous systems. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sodsong:2017:JPE

[SJPB17] Wasuwee Sodsong, Minyoung Jung, Jinwoo Park, and Bernd Burgstaller. JParEnt: Parallel entropy decoding for JPEG decompression on heterogeneous multicore architectures. Concurrency and Computation: Practice and Experience, 29(15): ??, August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Santander-Jimenez:2015:HAP

[SJVR15] Sergio Santander-Jiménez and Miguel A. Vega-Rodríguez. A hybrid approach to parallelize a fast non-dominated sorting genetic algorithm for phylogenetic inference. Concurrency and Computation: Practice and Experience, 27(3):702–734, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saxena:2018:QCA

[SJW18] Gaurav Saxena, Peter K. Jimack, and Mark A. Walkley. A quasi-cache-aware model for optimal domain partitioning in parallel geometric multigrid. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4328.

Seinstra:2004:UTF

[SK04] F. J. Seinstra and D. Koelma. User transparency: a fully sequential programming model for efficient data parallel image processing. Concurrency and Computation: Practice and Experience, 16(6):611–644, May 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Santos:2008:SDM

[SK08] Nuno Santos and Birger Koblitz. Security in distributed metadata catalogues. Concurrency and Computation: Practice and Experience, 20(17):1995–2007, December 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sancho:2009:OMC

[SK09] Jose Carlos Sancho and Darren J. Kerbyson. Optimizing multiple conjugate gradient solvers for large-scale systems. Concurrency and Computation: Practice and Experience, 21(14): 1804–1818, September 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2017:CFN

[SK17] Vishal Sharma and Rajesh Kumar. Cooperative frameworks and network models for flying ad hoc networks: a survey. Concurrency and Computation: Practice and Experience, 29 (4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stavrinides:2018:ICI

[SK18] Georgios L. Stavrinides and Helen D. Karatza. The impact of checkpointing interval selection on the scheduling performance of real-time fine-grained parallel applications in SaaS clouds under various failure probabilities. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4288.

Spampinato:2014:DBK

[SKA^{+}14] C. Spampinato, I. Kavasidis, M. Aldinucci, C. Pino, D. Giordano, and A. Faro. Discovering biological knowledge by integrating high-throughput data and scientific literature on the cloud. Concurrency and Computation: Practice and Experience, 26(10):1771–1786, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Son:2017:NOC

[SKB^{+}17] Junggab Son, Donghyun Kim, Md Zakirul Alam Bhuiyan, Rasheed Hussain, and Heekuck Oh. A new outsourcing conditional proxy re-encryption suitable for mobile cloud environment. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Skjellum:2004:RTM

[SKD^{+}04] Anthony Skjellum, Arkady Kanevsky, Yoginder S. Dandass, Jerrell Watts, Steve Paavola, Dennis Cottel, Greg Henley, L. Shane Hebert, Zhenqian Cui, Anna Rounbehler, and The Real-Time Message Passing Interface (Mpi and Rt) Forum. The Real-Time Message Passing Interface Standard (MPI/RT-1.1). Concurrency and Computation: Practice and Experience, 16(S1):Si–S322, December 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Skocir:2017:EET

[SKJ17] Pavle Skocir, Mario Kusek, and Gordan Jezic. Energyefficient task allocation for service provisioning in machineto-machine systems. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schonberger:2001:ASM

[SKK01] Siegfried Schönberger, Rudolf K. Keller, and Ismail Khriss. Algorithmic support for model transformation in objectoriented software development. Concurrency and Computation: Practice and Experience, 13(5):351–383, April 25, 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/78505739/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
78505739&PLACEBO=IE.pdf.

Schloegel:2002:PSD

[SKK02] Kirk Schloegel, George Karypis, and Vipin Kumar. Parallel static and dynamic multi-constraint graph partitioning.

Concurrency and Computation: Practice and Experience, 14 (3):219–240, March 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/93513488/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=93513488{\&}PLACEBO=IE.pdf.

Smith:2009:HBP

[SKNH09] Chris Smith, Thilo Kielmann, Steven Newhouse, and Marty Humphrey. The HPC basic profile and SAGA: standardizing compute grid access in the Open Grid Forum. Concurrency and Computation: Practice and Experience, 21(8):1053–1068, June 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sondergaard:2017:CTD

[SKR17] Hans Søndergaard, Stephan E. Korsholm, and Anders P. Ravn. Conformance test development with the Java modeling language. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Scheidegger:2008:TPC

[SKS^{+}08] Carlos Scheidegger, David Koop, Emanuele Santos, Huy Vo, Steven Callahan, Juliana Freire, and Cláudio Silva. Tackling the Provenance Challenge one layer at a time. Concurrency and Computation: Practice and Experience, 20(5):473–483, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seneviratne:2010:CPP

[SL10] Sena Seneviratne and David C. Levy. Cost profile prediction for Grid computing. Concurrency and Computation: Practice and Experience, 22(1):107–142, January 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Salah:2014:PBP

[SL14] Ahmad Salah and Kenli Li. PAR-3D-BLAST: a parallel tool for searching and aligning protein structures. Concurrency and Computation: Practice and Experience, 26(10): 1705–1714, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shou:2018:LDS

[SL18] Zhaoyu Shou and Simin Li. Large dataset summarization with automatic parameter optimization and parallel processing for local outlier detection. Concurrency and Computation: Practice and Experience, 30(16):e4466:1–e4466:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sosonkina:2008:ULS

[SLB08] M. Sosonkina, F. Liu, and R. Bramley. Usability levels for sparse linear algebra components. Concurrency and Computation: Practice and Experience, 20(12):1439–1454, August 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2018:NMS

[SLC^{+}18] Chuan Sun, Wei Liu, Duanfeng Chu, Wushuang Li, Zhenji Lu, and Jianyu Wang. A novel method of symbolic representation in diving data mining: a case study of highways in China. Concurrency and Computation: Practice and Experience, 30(24):e4976:1–e4976:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2019:NMS

[SLC^{+}19] Chuan Sun, Wei Liu, Duanfeng Chu, Wushuang Li, Zhenji Lu, and Jianyu Wang. A novel method of symbolic representation in diving data mining: a case study of highways in China. Concurrency and Computation: Practice and Experience, 31(7):e5115:1–e5115:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seelam:2012:EBS

[SLD^{+}12] Seetharami Seelam, Yanbin Liu, Parijat Dube, Megumi Ito, Deniz Binay, Michael Dawson, Pramod Nagaraja, Graeme Johnson, Liana Fong, Michel Hack, Xiaoqiao Meng, Yuqing Gao, and Li Zhang. Experiences in building and scaling an enterprise application on multicore systems. Concurrency and Computation: Practice and Experience, 24(2):111–123, February 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Servat:2016:DSP

[SLGL16] Harald Servat, Germán Llort, Judit Giménez, and Jesús Labarta. Detailed and simultaneous power and performance analysis. Concurrency and Computation: Practice and Experience, 28(2):252–273, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stamatakis:2004:APF

[SLM04] Alexandros P. Stamatakis, Thomas Ludwig, and Harald Meier. The AxML program family for maximum likelihoodbased phylogenetic tree inference. Concurrency and Computation: Practice and Experience, 16(9):975–988, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stamatakis:2005:RIP

[SLM05] Alexandros Stamatakis, Thomas Ludwig, and Harald Meier. RAxML-II: a program for sequential, parallel and distributed inference of large phylogenetic trees. Concurrency and Computation: Practice and Experience, 17(14):1705–1723, December 10, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stewart:2010:IPS

[SLM^{+}10] Craig A. Stewart, Matthew Link, D. Scott McCaulay, Greg Rodgers, George Turner, David Hancock, Peng Wang, Faisal Saied, Marlon Pierce, Ross Aiken, Matthias S. Mueller, Matthias Jurenz, Matthias Lieber, Jenett Tillotson, and Beth A. Plale. Implementation, performance, and science results from a 30.7 TFLOPS IBM BladeCenter cluster. Concurrency and Computation: Practice and Experience, 22(2): 157–174, February 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Slomiski:2006:UBE

[Slo06] Aleksander Slomiski. On using BPEL extensibility to implement OGSI and WSRF Grid workflows. Concurrency and Computation: Practice and Experience, 18(10):1229–1241, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2006:CTG

[SLT^{+}06] Yao Shi, Francis C. M. Lau, Savio S. H. Tse, Zhi-Hui Du, Rui-Chun Tang, and San-Li Li. Club theory of the Grid. Concurrency and Computation: Practice and Experience, 18(14): 1759–1773, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schnorr:2012:DAR

[SLV12] Lucas Mello Schnorr, Arnaud Legrand, and Jean-Marc Vincent. Detection and analysis of resource usage anomalies in large distributed systems through multi-scale visualization. Concurrency and Computation: Practice and Experience, 24 (15):1792–1816, October 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sakagami:2002:CCP

[SM02] H. Sakagami and T. Mizuno. Compatibility comparison and performance evaluation for Japanese HPF compilers using scientific applications. Concurrency and Computation: Practice and Experience, 14(8–9):679–689, July/ August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016123/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016123{\&}PLACEBO=IE.pdf.

Shende:2003:IAT

[SM03] Sameer Shende and Allen D. Malony. Integration and application of TAU in parallel Java environments. Concurrency and Computation: Practice and Experience, 15(3–5): 501–519, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schulze:2004:GCA

[SM04] B. Schulze and E. R. M. Madeira. Grid computing and active services. Concurrency and Computation: Practice and Experience, 16(5):535–542, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Simon:2009:OAM

[SM09] Tyler A. Simon and James McGalliard. Observation and analysis of the multicore performance impact on scientific applications. Concurrency and Computation: Practice and Experience, 21(17):2213–2231, December 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schulze:2011:MSC

[SM11] Bruno Schulze and James Myers. Middleware strategies for clouds and grids in e-Science. Concurrency and Computation: Practice and Experience, 23(17):2043–2047, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shibu:2019:ARR

[SM19a] N. V. Shibu and P. Malathi. Accurate and reliable reversible data hiding using sequential encoding techniques. Concurrency and Computation: Practice and Experience, 31(14): e4979:1–e4979:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soms:2019:SAD

[SM19b] Nisha Soms and P. Malathi. Secured and anonymous data transmission in Manet environment using zone-based intrusion detection system. Concurrency and Computation: Practice and Experience, 31(12):e4734:1–e4734:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suvitha:2019:NAM

[SM19c] Sundar Suvitha and J. M. Mathana. A novel automated MOALO algorithm aided RF low-noise amplifier design for wireless applications. Concurrency and Computation: Practice and Experience, 31(14):e4915:1–e4915:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Simpson:2007:PIP

[SMBT07] James J. Simpson, Timothy J. McIntire, Jared S. Berg, and Yueh Lung (Ben) Tsou. The Parallel Image Processing Environment (PIPE): automated parallelization of satellite data analyses. Concurrency and Computation: Practice and Experience, 19(1):1–36, January 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sousa:2018:PER

[SMFM18] Flávio R. C. Sousa, Leonardo O. Moreira, José S. Costa Filho, and Javam C. Machado. Predictive elastic replication for multi-tenant databases in the cloud. Concurrency and Computation: Practice and Experience, 30(16):??, August 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4437.

Safa:2019:ORM

[SMH^{+}19] Nader Sohrabi Safa, Carsten Maple, Mahboobeh Haghparast, Tim Watson, and Mehrdad Dianati. An opportunistic resource management model to overcome resource-constraint in the Internet of Things. Concurrency and Computation: Practice and Experience, 31(8):e5014:1–e5014:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seshadri:2018:DPA

[SMM18] Karthick Seshadri, Shalinie S. Mercy, and Sidharth Manohar. A distributed parallel algorithm for inferring hierarchical groups from large-scale text corpuses. Concurrency and Computation: Practice and Experience, 30(11):??, June 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4404.

Shah:2019:SNA

[SMS^{+}19] Naman Shah, Matthew Malensek, Harshil Shah, Shrideep Pallickara, and Sangmi Lee Pallickara. Scalable network analytics for characterization of outbreak influence in voluminous epidemiology datasets. Concurrency and Computation: Practice and Experience, 31(7):e4998:1–e4998:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sivagnanam:2015:EED

[SMY^{+}15] Subhashini Sivagnanam, Amit Majumdar, Kenneth Yoshimoto, Vadim Astakhov, Anita Bandrowski, MaryAnn Martone, and Nicholas. T. Carnevale. Early experiences in developing and managing the neuroscience gateway. Concurrency and Computation: Practice and Experience, 27(2):473–488, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schulze:2006:SIM

[SN06] Bruno Schulze and Radha Nandkumar. Special issue: Middleware for Grid computing. Concurrency and Computation: Practice and Experience, 18(6):549–552, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Strnad:2016:PCC

[SN16] D. Strnad and A. Nerat. Parallel construction of classification trees on a GPU. Concurrency and Computation: Practice and Experience, 28(5):1417–1436, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sheikholeslami:2018:ABR

[SN18] Fereshteh Sheikholeslami and Nima Jafari Navimipour. Auction-based resource allocation mechanisms in the cloud environments: a review of the literature and reflection on future challenges. Concurrency and Computation: Practice and Experience, 30(16):??, August 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4456.

Sim:2001:EUC

[SNB^{+}01] A. Sim, H. Nordberg, L. M. Bernardo, A. Shoshani, and D. Rotem. Experience with using CORBA to implement a file caching coordination system. Concurrency and Computation: Practice and Experience, 13(7):605–619, June 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/83002174/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
83002174&PLACEBO=IE.pdf.

Sanyal:2014:CBE

[SNEP14] Jibonananda Sanyal, Joshua New, Richard E. Edwards, and Lynne Parker. Calibrating building energy models using supercomputer trained machine learning agents. Concurrency and Computation: Practice and Experience, 26(13): 2122–2133, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Subhlok:2018:RPC

[SNGR18] Jaspal Subhlok, Hien Nguyen, Edgar Gabriel, and Mohammad Tanvir Rahman. Resilient parallel computing on volunteer PC grids. Concurrency and Computation: Practice and Experience, 30(18):e4478:1–e4478:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sagharichian:2015:ENC

[SNH15] M. Sagharichian, H. Naderi, and M. Haghjoo. ExPregel: a new computational model for large-scale graph processing. Concurrency and Computation: Practice and Experience, 27(17):4954–4969, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sengupta:2015:AES

[SNK^{+}15] Prasenjit Sengupta, Jimmy Nguyen, Jason Kwan, Padmanabhan K. Menon, Eric M. Heien, and John B. Rundle. Accelerating earthquake simulations on general-purpose graphics processors. Concurrency and Computation: Practice and Experience, 27(17):5460–5471, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Smari:2015:ECC

[SNM15] Waleed W. Smari, Antonio M. Navarro, and William K. Mc-Quay. Editorial: Contemporary computer supported collaboration: Systems, technologies, algorithms, and applications. Concurrency and Computation: Practice and Experience, 27 (11):2681–2685, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soner:2016:NAB

[SO16] Seren Soner and Can Ozturan. A new auction-based scheduler for heterogeneous systems with moldable generic resources support. Concurrency and Computation: Practice and Experience, 28(8):2341–2352, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sodan:2005:LCC

[Sod05] A. C. Sodan. Loosely coordinated coscheduling in the context of other approaches for dynamic job scheduling: a survey. Concurrency and Computation: Practice and Experience, 17 (15):1725–1781, December 25, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soddemann:2007:SGD

[Sod07] Thomas Soddemann. Science gateways to DEISA: user requirements, technologies, and the material sciences and plasma physics gateway. Concurrency and Computation: Practice and Experience, 19(6):839–850, April 25, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sood:2016:FPB

[Soo16] Sandeep K. Sood. Function points-based resource prediction in cloud computing. Concurrency and Computation: Practice and Experience, 28(10):2781–2794, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sorensen:2013:ATL

[Sør13] Hans Henrik Brandenborg Sørensen. Auto-tuning of level 1 and level 2 BLAS for GPUs. Concurrency and Computation: Practice and Experience, 25(8):1183–1198, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sakthithasan:2016:CRC

[SP16] Sripirakas Sakthithasan and Russel Pears. Capturing recurring concepts using discrete Fourier transform. Concurrency and Computation: Practice and Experience, 28(15): 4013–4035, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sparks:2019:EDH

[Spa19] Jonathan Sparks. Enabling Docker for HPC. Concurrency and Computation: Practice and Experience, 31(16):e5018:1– e5018:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sirvent:2006:AGW

[SPBL06] Raül Sirvent, Josep M. Pérez, Rosa M. Badia, and Jesús Labarta. Automatic Grid workflow based on imperative programming languages. Concurrency and Computation: Practice and Experience, 18(10):1169–1186, August 25, 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Simmhan:2008:QCK

[SPG08] Yogesh L. Simmhan, Beth Plale, and Dennis Gannon. Query capabilities of the Karma provenance framework. Concurrency and Computation: Practice and Experience, 20(5):441– 451, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seal:2013:SSM

[SPH13] Sudip K. Seal, Kalyan S. Perumalla, and Steven P. Hirshman. Scaling the SIESTA magnetohydrodynamics equilibrium code. Concurrency and Computation: Practice and Experience, 25(15):2207–2223, October 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shin:2014:DTA

[SPJ14] Kwang Sup Shin, Myoung-Ju Park, and Jae-Yoon Jung. Dynamic task assignment and resource management in cloud services by using bargaining solution. Concurrency and Computation: Practice and Experience, 26(7):1432–1452, May 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Silvestri:2006:TSA

[SPLO06] Fabrizio Silvestri, Diego Puppin, Domenico Laforenza, and Salvatore Orlando. Toward a search architecture for software components. Concurrency and Computation: Practice and Experience, 18(10):1317–1331, August 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sanchez:2011:PUR

[SPMP11] Sergio Sánchez, Abel Paz, Gabriel Martín, and Antonio Plaza. Parallel unmixing of remotely sensed hyperspectral images on commodity graphics processing units. Concurrency and Computation: Practice and Experience, 23(13): 1538–1557, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Souza:2017:CBB

[SPQ^{+}17] Matheus A. Souza, Pedro Henrique Penna, Matheus M. Queiroz, Alyson D. Pereira, Luís Fabricio Wanderley Góes, Henrique C. Freitas, Márcio Castro, Philippe O. A. Navaux, and Jean-François Méhaut. CAP Bench: a benchmark suite for performance and energy evaluation of low-power manycore processors. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schuchardt:2007:PBK

[SPR^{+}07] Karen Schuchardt, Carmen Pancerella, Larry A. Rahn, Brett Didier, Deepti Kodeboyina, David Leahy, James D. Myers, Oluwayemisi O. Oluwole, William Pitz, Branko Ruscic, Jing Song, Gregor von Laszewski, and Christine Yang. Portal-based Knowledge Environment for Collaborative Science. Concurrency and Computation: Practice and Experience, 19(12):1703–1716, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Strom:2017:HLR

[SPS17] Tórur Biskopstø Strøm, Wolfgang Puffitsch, and Martin Schoeberl. Hardware locks for a real-time Java chip multiprocessor. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shojafar:2019:RAC

[SPSB19] Mohammad Shojafar, Zahra Pooranian, Mehdi Sookhak, and Rajkumar Buyya. Recent advances in cloud data centers toward fog data centers. Concurrency and Computation: Practice and Experience, 31(8):e5164:1–e5164:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Szymaniak:2007:ESA

[SPSNvS07] Michal Szymaniak, Guillaume Pierre, Mariana Simons-Nikolova, and Maarten van Steen. Enabling service adaptability with versatile anycast. Concurrency and Computation: Practice and Experience, 19(13):1837–1863, September 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sommer:2009:AEM

[SPW09] Robin Sommer, Vern Paxson, and Nicholas Weaver. An architecture for exploiting multi-core processors to parallelize network intrusion prevention. Concurrency and Computation: Practice and Experience, 21(10):1255–1279, July 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schmidt:2010:MTW

[SPZ^{+}10] Jessica Schmidt, Cécile Piret, Nan Zhang, Benjamin J. Kadlec, David A. Yuen, Yingchun Liu, Grady Barrett Wright, and Erik O. D. Sevre. Modeling of tsunami waves and atmospheric swirling flows with graphics processing unit (GPU) and radial basis functions (RBF). Concurrency and Computation: Practice and Experience, 22(12):1813–1835, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Siddiqa:2019:CNE

[SQS^{+}19] Ayesha Siddiqa, Faisal Fayyaz Qureshi, Munam Ali Shah, Rahat Iqbal, Abdul Wahid, and Victor Chang. CCN: a novel energy efficient greedy routing protocol for green computing. Concurrency and Computation: Practice and Experience, 31(23):e4461:1–e4461:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2017:QIB

[SR17] Krishan Veer Singh and Zahid Raza. A quantum-inspired binary gravitational search algorithm–based job-scheduling model for mobile computational grid. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sajid:2019:EEQ

[SR19a] Mohammad Sajid and Zahid Raza. Energy-efficient quantuminspired stochastic Q-HypE algorithm for batch-of-stochastic-tasks on heterogeneous DVFS-enabled processors. Concurrency and Computation: Practice and Experience, 31(20): e5327:1–e5327:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saravanakumar:2019:MFC

[SR19b] K. Saravanakumar and R. Rajeswari. Microbial fuel cellbased self-powered biosensor for environment monitoring in IoT cloud framework. Concurrency and Computation: Practice and Experience, 31(15):e5165:1–e5165:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Su:2019:RPM

[SR19c] Wangshu Su and Jinzheng Ren. Risk propagation model based on social life and credit activities multilayers fusion network. Concurrency and Computation: Practice and Experience, 31(10):e4732:1–e4732:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Simmhan:2016:ECC

[SRAG16] Yogesh Simmhan, Lavanya Ramakrishnan, Gabriel Antoniu, and Carole Goble. Editorials: Cloud computing for datadriven science and engineering. Concurrency and Computation: Practice and Experience, 28(4):947–949, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schulze:2009:SIA

[SRdS09] Bruno Schulze, Omer Rana, and José Neuman de Souza. Special issue: Advanced strategies in Grid environments — models and techniques for scheduling and programming. Concurrency and Computation: Practice and Experience, 21(13): 1667–1671, September 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sebastiao:2013:SIP

[SRF13] Nuno Sebastião, Nuno Roma, and Paulo Flores. Special issue papers: Configurable and scalable class of high performance hardware accelerators for simultaneous DNA sequence alignment. Concurrency and Computation: Practice and Experience, 25(10):1319–1339, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stanberry:2014:OHP

[SRL^{+}14] Larissa Stanberry, Bhanu Rekepalli, Yuan Liu, Paul Giblock, Roger Higdon, Elizabeth Montague, William Broomall, Natali Kolker, and Eugene Kolker. Optimizing high performance computing workflow for protein functional annotation. Concurrency and Computation: Practice and Experience, 26(13):

2112–2121, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schulze:2013:ELT

[SRM13a] Bruno Schulze, Vinod Rebello, and Jose Moreira. Editorial: Latest trends in computer architectures and parallel and distributed technologies. Concurrency and Computation: Practice and Experience, 25(6):771–774, April 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schutt:2013:MSM

[SRM13b] Thorsten Schütt, Alexander Reinefeld, and Robert Maier. MR-search: massively parallel heuristic search. Concurrency and Computation: Practice and Experience, 25(1):40– 54, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schneible:2015:CEW

[SRM^{+}15] Joseph Schneible, Lubomir Riha, Maria Malik, Tarek El-Ghazawi, and Andrei Alexandru. Communication efficient work distributions in stencil operation based applications. Concurrency and Computation: Practice and Experience, 27 (13):3262–3280, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Su:2015:ASI

[SRN^{+}15] Xiang Su, Jukka Riekki, Jukka K. Nurminen, Johanna Nieminen, and Markus Koskimies. Adding semantics to Internet of Things. Concurrency and Computation: Practice and Experience, 27(8):1844–1860, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suresh:2019:DCN

[SRR19] R. Suresh, A. Nagaraja Rao, and B. Eswara Reddy. Detection and classification of normal and abnormal patterns in mammograms using deep neural network. Concurrency and Computation: Practice and Experience, 31(14):e5293:1–e5293:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sajid:2016:EES

[SRS16] Mohammad Sajid, Zahid Raza, and Mohammad Shahid. Energy-efficient scheduling algorithms for batch-of-tasks (BoT) applications on heterogeneous computing systems. Concurrency and Computation: Practice and Experience, 28 (9):2644–2669, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stokes-Rees:2007:DLG

[SRTG^{+}07] I. Stokes-Rees, A. Tsaregorodtsev, V. Garonne, R. Graciani, M. Sanchez, M. Frank, and J. Closier. Developing LHCb Grid software: experiences and advances. Concurrency and Computation: Practice and Experience, 19(2):133–152, February 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Strohmaier:2007:AMP

[SS07] Erich Strohmaier and Hongzhang Shan. APEX-Map: a parameterized scalable memory access probe for highperformance computing systems. Concurrency and Computation: Practice and Experience, 19(17):2185–2205, December 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seshadri:2015:PGC

[SS15a] Karthick Seshadri and S. Mercy Shalinie. Parallelization of a graph-cut based algorithm for hierarchical clustering of web documents. Concurrency and Computation: Practice and Experience, 27(17):5156–5176, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sonntag:2015:EFM

[SS15b] Sebastian Sonntag and Henna Suomi. Economic feasibility of multipath protocols in mobile Internet of Things applications. Concurrency and Computation: Practice and Experience, 27 (8):1913–1931, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Szalkowski:2015:UDM

[SS15c] Dominik Szałkowski and Przemysław Stpiczyński. Using distributed memory parallel computers and GPU clusters for multidimensional Monte Carlo integration. Concurrency and Computation: Practice and Experience, 27(4):923–936, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sandhu:2017:SGN

[SS17a] Rajinder Sandhu and Sandeep K. Sood. A stochastic game net-based model for effective decision-making in smart environments. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Subramanian:2017:SOS

[SS17b] Sabitha Malli Subramanian and Vijayalakshmi Soundarajan. SC-OCR: similarity-based clustering and optimum cache replacement approach. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shang:2018:TMR

[SS18] Yizi Shang and Ling Shang. Trust model for reliable node allocation based on daily computer usage behavior. Concurrency and Computation: Practice and Experience, 30(6):??, March 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4346.

Sekar:2019:IFF

[SS19a] R. Arun Sekar and S. Sasipriya. Implementation of FIR filter using reversible modified carry select adder. Concurrency and Computation: Practice and Experience, 31(14):e4952:1– e4952:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Selvi:2019:RSP

[SS19b] A. Senthil Selvi and R. Sukumar. Removal of salt and pepper noise from images using hybrid filter (HF) and fuzzy logic noise detector (FLND). Concurrency and Computation: Practice and Experience, 31(12):e4501:1–e4501:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sudarsan:2019:BDK

[SS19c] V. Sudarsan and R. Sugumar. Building a distributed Kmeans model for Weka using remote method invocation (RMI) feature of Java. Concurrency and Computation: Practice and Experience, 31(14):e5313:1–e5313:??, July 25, 2019.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sinkarovs:2014:SCF

[SSB^{+}14] Artjoms Sinkarovs, Sven-Bodo Scholz, Robert Bernecky, Roeland Douma, and Clemens Grelck. SaC/C formulations of the all-pairs N-body problem and their performance on SMPs and GPGPUs. Concurrency and Computation: Practice and Experience, 26(4):952–971, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2016:AIA

[SSC^{+}16] Jianhua Sun, Lingjun She, Hao Chen, Wenyong Zhong, Cheng Chang, Zhiwen Chen, Wentao Li, and Shuna Yao. Automatically identifying apps in mobile traffic. Concurrency and Computation: Practice and Experience, 28(14): 3927–3941, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sakdhnagool:2019:CAC

[SSE19] Putt Sakdhnagool, Amit Sabne, and Rudolf Eigenmann. Comparative analysis of coprocessors. Concurrency and Computation: Practice and Experience, 31(1):e4756:1–e4756:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2019:GBB

[SSIH19] Jingcheng Shen, Kentaro Shigeoka, Fumihiko Ino, and Kenichi Hagihara. GPU-based branch-and-bound method to solve large 0–1 knapsack problems with data-centric strategies. Concurrency and Computation: Practice and Experience, 31(4):e4954:1–e4954:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sibai:2011:PPC

[SSK11] Fadi N. Sibai, Mohammad Saad, and Hashir K. Kidwai. Parallelization and performance comparison of the conjugate gradient equation solver on multicore Cell and Xeon computers. Concurrency and Computation: Practice and Experience, 23(18):2463–2467, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Simmler:2004:RTP

[SSM04] H. Simmler, H. Singpiel, and R. Männer. Real-time primer design for DNA chips. Concurrency and Computation: Practice and Experience, 16(9):855–872, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soriano-Salvador:2015:OSN

[SSMB15] Enrique Soriano-Salvador, Gorka Guardiola Muzquiz, and Francisco J. Ballesteros. Optimistic semaphores with nondeterministic choice operation for heterogeneous manycore systems. Concurrency and Computation: Practice and Experience, 27(14):3676–3701, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soomro:2018:BTB

[SSU18] Pirah Noor Soomro, Muhammad Aditya Sasongko, and Didem Unat. BindMe: a thread binding library with advanced mapping algorithms. Concurrency and Computation: Practice and Experience, 30(21):e4692:1–e4692:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stegailov:2019:VHM

[SSV19] Vladimir Stegailov, Grigory Smirnov, and Vyacheslav Vecher. VASP hits the memory wall: Processors efficiency comparison. Concurrency and Computation: Practice and Experience, 31(19):e5136:1–e5136:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sundriyal:2013:AEE

[SSZ13] Vaibhav Sundriyal, Masha Sosonkina, and Zhao Zhang. Achieving energy efficiency during collective communications. Concurrency and Computation: Practice and Experience, 25 (15):2140–2156, October 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Souza:2014:STM

[SSZ14] Paulo S. L. Souza, Simone R. S. Souza, and Ed Zaluska. Structural testing for message-passing concurrent programs: an extended test model. Concurrency and Computation: Practice and Experience, 26(1):21–50, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stanisic:2015:FPP

[STL^{+}15] Luka Stanisic, Samuel Thibault, Arnaud Legrand, Brice Videau, and Jean-François Méhaut. Faithful performance prediction of a dynamic task-based runtime system for heterogeneous multi-core architectures. Concurrency and Computation: Practice and Experience, 27(16):4075–4090, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sawabe:2017:EQS

[STO17] Anan Sawabe, Kazuya Tsukamoto, and Yuji Oie. Efficient quality of service-aware packet chunking scheme for machineto-machine cloud services. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Strnad:2011:PTV

[Str11] Damjan Strnad. Parallel terrain visibility calculation on the graphics processing unit. Concurrency and Computation: Practice and Experience, 23(18):2452–2462, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2019:FSS

[STS19] Hong Shen, Hui Tian, and Yingpeng Sang. Foreword to the special section on parallel and distributed computing, applications and technologies. Concurrency and Computation: Practice and Experience, 31(21):e5409:1–e5409:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Simsiri:2018:WEP

[STTW18] Natcha Simsiri, Kanat Tangwongsan, Srikanta Tirthapura, and Kun-Lung Wu. Work-efficient parallel union-find. Concurrency and Computation: Practice and Experience, 30 (4):??, February 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4333.

Stilkerich:2012:TMJ

[STWSP12] Michael Stilkerich, Isabella Thomm, Christian Wawersich, and Wolfgang Schröder-Preikschat. Tailor-made JVMs for statically configured embedded systems. Concurrency and Computation: Practice and Experience, 24(8):789–812, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tu:2016:FGA

[sTzNL16] Shan shan Tu, Shao zhang Niu, and Hui Li. A fine-grained access control and revocation scheme on clouds. Concurrency and Computation: Practice and Experience, 28(6):1697–1714, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sanjay:2009:SST

[SV09] H. A. Sanjay and Sathish S. Vadhiyar. A strategy for scheduling tightly coupled parallel applications on clusters. Concurrency and Computation: Practice and Experience, 21(18): 2491–2517, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sakthivelmurugan:2019:MOE

[SVB19] V. Sakthivelmurugan, R. Vimala, and K. R. Aravind Britto. Magnum opus of an efficient hospitality technique for load balancing in cloud environment. Concurrency and Computation: Practice and Experience, 31(14):e5078:1–e5078:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shahand:2015:RSG

[SvDO15] S. Shahand, J. van Duffelen, and S. D. Olabarriaga. Reflections on science gateways sustainability through the business model canvas: case study of a neuroscience gateway. Concurrency and Computation: Practice and Experience, 27(16): 4269–4281, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sundari:2012:LIA

[SVN12] Sivagama Sundari M, Sathish S. Vadhiyar, and Ravi S. Nanjundiah. Large improvements in application throughput of long-running multi-component applications using batch grids. Concurrency and Computation: Practice and Experience, 24 (15):1775–1791, October 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Souza:2008:STC

[SVS^{+}08] S. R. S. Souza, S. R. Vergilio, P. S. L. Souza, A. S. Simão, and A. C. Hausen. Structural testing criteria for message-passing parallel programs. Concurrency and Computation: Practice and Experience, 20(16):1893–1916, November 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

ShanmugaPriya:2019:PAS

[SVY19] S. ShanmugaPriya, A. Valarmathi, and D. Yuvaraj. The personal authentication service and security enhancement for optimal strong password. Concurrency and Computation: Practice and Experience, 31(14):e5009:1–e5009:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soewito:2009:CWM

[SW09] Benfano Soewito and Ning Weng. Concurrent workload mapping for multicore security systems. Concurrency and Computation: Practice and Experience, 21(10):1281–1306, July 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2011:ACA

[SW11] Lili Sun and Hua Wang. Access control and authorization for protecting disseminative information in E-learning workflow. Concurrency and Computation: Practice and Experience, 23 (16):2034–2042, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2012:PBA

[SW12] Lili Sun and Hua Wang. A purpose-based access control in native XML databases. Concurrency and Computation: Practice and Experience, 24(10):1154–1166, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Simon:2012:PAI

[SWB12] Tyler A. Simon, William A. Ward, Jr., and Alan P. Boss. Performance analysis of Intel multiprocessors using astrophysics simulations. Concurrency and Computation: Practice and Experience, 24(2):155–166, February 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Simakov:2015:AKH

[SWD^{+}15] Nikolay A. Simakov, Joseph P. White, Robert L. DeLeon, Amin Ghadersohi, Thomas R. Furlani, Matthew D. Jones, Steven M. Gallo, and Abani K. Patra. Application kernels: HPC resources performance monitoring and variance analysis. Concurrency and Computation: Practice and Experience, 27(17):5238–5260, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stolle:2017:UAR

[SWD^{+}17] Jonas Stolle, Michael Wagner, Jens Doleschal, Felix Schmitt, and Holger Brunst. Using adaptive runtime filtering to support an event-based performance analysis. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stojmenovic:2016:OFC

[SWHL16] Ivan Stojmenovic, Sheng Wen, Xinyi Huang, and Hao Luan. An overview of Fog computing and its security issues. Concurrency and Computation: Practice and Experience, 28(10): 2991–3005, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Skjellum:2001:OOA

[SWL^{+}01] Anthony Skjellum, Diane G. Wooley, Ziyang Lu, Michael Wolf, Purushotham V. Bangalore, Andrew Lumsdaine, Jeffrey M. Squyres, and Brian McCandless. Object-oriented analysis and design of the Message Passing Interface. Concurrency and Computation: Practice and Experience, 13(4): 245–292, April 10, 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/78502300/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=78502300&PLACEBO=IE.pdf.

Shao:2017:RRT

[SWLJ17] Lu Shao, Cheng Wang, Lu Liu, and Changjun Jiang. RTS: road topology-based scheme for traffic condition estimation via vehicular crowdsensing. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Smallen:2017:LSP

[SWP17] Shava Smallen, Nadya Williams, and Philip Papadopoulos. Lightweight scheduling for the PRAGMA cloud testbed. Concurrency and Computation: Practice and Experience, 29(13): ??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shah:2018:PAB

[SWS^{+}18] Mehreen Shah, Zahid Wadud, Arshad Sher, Mahmood Ashraf, Zahoor Ali Khan, and Nadeem Javaid. Position adjustment-based location error-resilient geo-opportunistic routing for void hole avoidance in underwater sensor networks. Concurrency and Computation: Practice and Experience, 30(21):e4772:1–e4772:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schikuta:2008:GWO

[SWU08] Erich Schikuta, Helmut Wanek, and Irfan Ul Haq. Grid workflow optimization regarding dynamically changing resources and conditions. Concurrency and Computation: Practice and Experience, 20(15):1837–1849, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2016:IBS

[SWW^{+}16] Lingwei Song, Jinxia Wei, Licheng Wang, Chenlei Cao, and Xinxin Niu. Identity-based storage management and integrity verify protocol for secure outsourcing in multi-cloud. Concurrency and Computation: Practice and Experience, 28(6): 1930–1945, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2012:IAH

[SWZ12] Xiaoxun Sun, Hua Wang, and Yanchun Zhang. On the identity anonymization of high-dimensional rating data. Concurrency and Computation: Practice and Experience, 24(10): 1108–1122, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Su:2018:UBI

[SWZ^{+}18] Xin Su, Bingying Wang, Xuewu Zhang, Yupeng Wang, and Dongmin Choi. User biometric information-based secure method for smart devices. Concurrency and Computation: Practice and Experience, 30(3):??, February 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4150.

Sohrabi:2017:EEA

[SYMA17] Sahar Sohrabi, Yun Yang, Irene Moser, and Aldeida Aleti. Energy-efficient adaptive virtual machine migration mechanism for private clouds. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2011:CPS

[SZ11] Xiaoqing Shi and Hai Zhuge. Cyber physical socio ecology. Concurrency and Computation: Practice and Experience, 23 (9):972–984, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2018:APA

[SZ18] Yijun Shen and Guanhong Zhang. Adaptive predistortion algorithm based on the multilinear multiplier fusion. Concurrency and Computation: Practice and Experience, 30(22): e4645:1–e4645:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Swamynathan:2008:EFP

[SZA08] Gayatri Swamynathan, Ben Y. Zhao, and Kevin C. Almeroth. Exploring the feasibility of proactive reputations. Concurrency and Computation: Practice and Experience, 20(2):155– 166, February 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2019:ATS

[SZD19] Yina Sun, Fanhe Zhu, and Weimin Dai. Acquisition of transgenic salt-tolerant upland rice based on genetic algorithm and its resistance. Concurrency and Computation: Practice and Experience, 31(10):e4748:1–e4748:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2019:RDP

[SZG^{+}19] Dian Shen, Pengcheng Zhou, Yidan Gao, Xiaolin Guo, and Runqun Xiong. Rendering differential performance preference through intelligent network edge in cloud data centers. Concurrency and Computation: Practice and Experience, 31(24): e5262:1–e5262:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sevcik:2017:WSN

[ŠŽH17] Peter Ševčík, Samuel Žák, and Michal Hodoň. Wireless sensor network for smart power metering. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2009:FMK

[SZL09] Xiaoping Sun, Hai Zhuge, and Qing Li. A framework for the massive knowledge Web. Concurrency and Computation: Practice and Experience, 21(5):705–723, April 10, 2009. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sanislav:2019:RFD

[SZMF19] Teodora Sanislav, Sherali Zeadally, George Dan Mois, and Hacène Fouchal. Reliability, failure detection and prevention in cyber-physical systems (CPSs) with agents. Concurrency and Computation: Practice and Experience, 31(24):e4481:1– e4481:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sudhakar:2016:EIO

[SZR16] Krishna Sudhakar, Yuhong Zhao, and Franz-Josef Rammig. Efficient integration of online model checking into a small-footprint real-time operating system. Concurrency and Computation: Practice and Experience, 28(14):3773–3797, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Samadi:2018:PCB

[SZT18] Yassir Samadi, Mostapha Zbakh, and Claude Tadonki. Performance comparison between Hadoop and Spark frameworks using HiBench benchmarks. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4367.

Sulikowski:2019:MOU

[SZT19] Piotr Sulikowski, Tomasz Zdziebko, and Dominik Turzyński. Modeling online user product interest for recommender systems and ergonomics studies. Concurrency and Computation: Practice and Experience, 31(22):e4301:1–e4301:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2019:ORP

[SZXG19] Da-Zhi Sun, Ze-Guang Zhu, Guang-Quan Xu, and Wei Guo. One-round provably secure yoking-proof for RFID applications. Concurrency and Computation: Practice and Experience, 31(23):e4931:1–e4931:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Talwar:2006:RAR

[TAB^{+}06] Vanish Talwar, Bikash Agarwalla, Sujoy Basu, Raj Kumar, and Klara Nahrstedt. Resource allocation for remote desktop sessions in Utility Grids. Concurrency and Computation: Practice and Experience, 18(6):667–684, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tsuneizumi:2011:SGC

[TAI^{+}11] Isamu Tsuneizumi, Ailixier Aikebaier, Makoto Ikeda, Tomoya Enokido, and Makoto Takizawa. A scalable group communication protocol with hybrid clocks. Concurrency and Computation: Practice and Experience, 23(5):477–490, April 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Theodorou:2016:QME

[TALT16] Vasileios Theodorou, Alberto Abelló, Wolfgang Lehner, and Maik Thiele. Quality measures for ETL processes: from goals to implementation. Concurrency and Computation: Practice and Experience, 28(15):3969–3993, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tajiki:2019:SBR

[TAMC19] Mohammad Mahdi Tajiki, Behzad Akbari, Nader Mokari, and Luca Chiaraviglio. SDN-based resource allocation in MPLS networks: a hybrid approach. Concurrency and Computation: Practice and Experience, 31(8):e4728:1–e4728:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2012:MOI

[Tan12] Peiyi Tang. Measuring the overhead of Intel C++ Concurrent Collections over Threading Building Blocks for Gauss–Jordan elimination. Concurrency and Computation: Practice and Experience, 24(18):2282–2301, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2015:HAC

[Tan15] Zhaohui Tang. Homomorphic authentication codes for network coding. Concurrency and Computation: Practice and Experience, 27(15):3892–3911, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tracy:2012:APL

[TB12] Daniel J. Tracy and Sheldon Brown. Accelerating physics in large, continuous virtual environments. Concurrency and Computation: Practice and Experience, 24(2):125–134, February 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Turek:2018:SIP

[TBH^{+}18] Wojciech Turek, Aleksander Byrski, John Hughes, Kevin Hammond, and Marek Zaionc. Special issue on parallel and distributed computing based on the functional programming paradigm. Concurrency and Computation: Practice and Experience, 30(22):e4842:1–e4842:??, November 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thiyagalingam:2006:MLC

[TBK06] Jeyarajan Thiyagalingam, Olav Beckmann, and Paul H. J. Kelly. Is Morton layout competitive for large two-dimensional arrays yet? Concurrency and Computation: Practice and Experience, 18(11):1509–1539, September 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tudruj:2015:PFD

[TBK^{+}15] Marek Tudruj, Janusz Borkowski, Damian Kopański, Eryk Laskowski, Łukasz Maśko, and Adam Smyk. PEGASUS DA framework for distributed program execution control based on application global states monitoring. Concurrency and Computation: Practice and Experience, 27(4):1027–1053, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Touhafi:2018:CCB

[TBTZ18] Abdellah Touhafi, An Braeken, Abderrahim Tahiri, and Mostapha Zbakh. CoderLabs: a cloud-based platform for real-time online labs with user collaboration. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4377.

Tylissanakis:2012:NCM

[TC12] Giorgos Tylissanakis and Yiannis Cotronis. Nonintrusive collection and management of data provenance in scientific workflows. Concurrency and Computation: Practice and Experience, 24(18):2268–2281, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2017:VMO

[TC17] Chunming Tang and Cailing Cai. Verifiable mobile online social network privacy-preserving location sharing scheme. Concurrency and Computation: Practice and Experience, 29 (24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tolosana-Calasanz:2010:AEH

[TCBR^{+}10] Rafael Tolosana-Calasanz, José A. Bañares, Omer F. Rana, Pedro Álvarez, Joaquín Ezpeleta, and Andreas Hoheisel.

Adaptive exception handling for scientific workflows. Concurrency and Computation: Practice and Experience, 22(5): 617–642, April 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tolosana-Calasanz:2011:ASP

[TCBR11] Rafael Tolosana-Calasanz, José A. Bañares, and Omer F. Rana. Autonomic streaming pipeline for scientific workflows. Concurrency and Computation: Practice and Experience, 23 (16):1868–1892, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tolosana-Calasanz:2017:CRM

[TCDMR^{+}17] Rafael Tolosana-Calasanz, Javier Diaz-Montes, Omer F. Rana, Manish Parashar, Erotokritos Xydas, Charalampos Marmaras, Panagiotis Papadopoulos, and Liana Cipcigan. Computational resource management for data-driven applications with deadline constraints. Concurrency and Computation: Practice and Experience, 29(8):??, April 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Trinder:2013:RAP

[TCH^{+}13] P. W. Trinder, M. I. Cole, K. Hammond, H-W. Loidl, and G. J. Michaelson. Resource analyses for parallel and distributed coordination. Concurrency and Computation: Practice and Experience, 25(3):309–348, March 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Taufer:2005:SHA

[TCP^{+}05] M. Taufer, M. Crowley, D. J. Price, A. A. Chien, and C. L. Brooks III. Study of a highly accurate and fast protein– ligand docking method based on molecular dynamics. Concurrency and Computation: Practice and Experience, 17(14): 1627–1641, December 10, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tardivo:2017:HPM

[TCSBMG17] María Laura Tardivo, Paola Caymes-Scutari, Germán Bianchini, and Miguel Méndez-Garabetti. Hierarchical parallel model for improving performance on differential evolution. Concurrency and Computation: Practice and Experience, 29 (10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Teing:2018:CFC

[TDC18] Yee-Yang Teing, Ali Dehghantanha, and Kim-Kwang Raymond Choo. CloudMe forensics: a case of big data forensic investigation. Concurrency and Computation: Practice and Experience, 30(5):??, March 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4277.

Tian:2018:RAB

[TDL^{+}18] Yinhua Tian, Yuyue Du, Maozhen Li, Dong Han, and Qiang Hu. Reduced alignment based on Petri nets. Concurrency and Computation: Practice and Experience, 30(23):e4411:1– e4411:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thomas:2002:APP

[TDM^{+}02] Mary Thomas, Maytal Dahan, Kurt Mueller, Steve Mock, Cathie Mills, and Ray Regno. Application portals: practice and experience. Concurrency and Computation: Practice and Experience, 14(13–15):1427–1443, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tao:2015:GAS

[TDM^{+}15] Yuan Tao, Yangdong Deng, Shuai Mu, Zhenzhong Zhang, Mingfa Zhu, Limin Xiao, and Li Ruan. GPU accelerated sparse matrix–vector multiplication and sparse matrix– transpose vector multiplication. Concurrency and Computation: Practice and Experience, 27(14):3771–3789, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Torquati:2019:PAP

[TDM^{+}19] Massimo Torquati, Daniele De Sensi, Gabriele Mencagli, Marco Aldinucci, and Marco Danelutto. Power-aware pipelining with automatic concurrency control. Concurrency and Computation: Practice and Experience, 31(5):e4652:1– e4652:??, March 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Truong:2003:SPA

[TF03] Hong-Linh Truong and Thomas Fahringer. SCALEA: a performance analysis tool for parallel programs. Concurrency and Computation: Practice and Experience, 15(11–12):1001– 1025, September 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tourino:2007:SIC

[TFDA07] Juan Touriño, Basilio B. Fraguela, Ramón Doallo, and Manuel Arenaz. Special issue: Current trends in compilers for parallel computers. Concurrency and Computation: Practice and Experience, 19(18):2313–2316, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tejedor:2012:HPT

[TFG^{+}12] Enric Tejedor, Montse Farreras, David Grove, Rosa M. Badia, Gheorghe Almasi, and Jesus Labarta. A high-productivity task-based programming model for clusters. Concurrency and Computation: Practice and Experience, 24(18):2421–2448, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Turek:2010:FRH

[TGB^{+}10] Stefan Turek, Dominik Göddeke, Christian Becker, Sven H. M. Buijssen, and Hilmar Wobker. FEAST — realization of hardware-oriented numerics for HPC simulations with finite elements. Concurrency and Computation: Practice and Experience, 22(16):2247–2265, November 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Titmus:2014:ADD

[TGS14] Matthew A. Titmus, James Gurtowski, and Michael C. Schatz. Answering the demands of digital genomics. Concurrency and Computation: Practice and Experience, 26(4): 917–928, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tolle:2010:SIC

[TH10] Kristin M. Tolle and Anthony J. G. Hey. Special issue: Concurrency and computation: Practice and experience from the Microsoft eScience Workshop. Concurrency and Computation: Practice and Experience, 22(17):2297–2299, December 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thota:2019:SIC

[TH19] Abhinav Thota and Yun He. Special issue of the Cray User Group (CUG 2018). Concurrency and Computation: Practice and Experience, 31(16):e5117:1–e5117:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Theodoropoulos:2001:SAH

[The01] Georgios K. Theodoropoulos. Simulating asynchronous hardware on multiprocessor platforms: the case of AMULET1. Concurrency and Computation: Practice and Experience, 13 (10):869–904, August 25, 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract/85007181/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=85007181&PLACEBO=IE.pdf.

Tang:2015:HNA

[THF15] Bing Tang, Haiwu He, and Gilles Fedak. HybridMR: a new approach for hybrid MapReduce combining desktop grid and cloud infrastructures. Concurrency and Computation: Practice and Experience, 27(16):4140–4155, November 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tilak:2011:MTM

[THM^{+}11] Omkar Tilak, Andrew Hoblitzell, Snehasis Mukhopadhyay, Qian You, Shiaofen Fang, Yuni Xia, and Joseph Bidwell. Multilevel text mining for bone biology. Concurrency and Computation: Practice and Experience, 23(17):2355–2364, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thomas:2007:SIW

[Tho07] Mary Thomas. Special issue: Workshop on Grid Computing Portals (GCE 2005). Concurrency and Computation: Practice and Experience, 19(12):1563–1570, August 25, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2019:SAO

[THQ19] Junjian Tang, Shunli Hao, and Wenqi Qu. Sentiment analysis of online Chinese comments based on statistical learning combining with pattern matching. Concurrency and Computation: Practice and Experience, 31(10):e4765:1–e4765:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2017:SSI

[TJ17a] Chao Tan and Genlin Ji. Semi-supervised incremental feature extraction algorithm for large-scale data stream. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2017:SLP

[TJ17b] Chao Tan and Genlin Ji. Semisupervised local preserving embedding algorithm based on maximum margin criterion for large-scale data streams. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2017:RDS

[TJD^{+}17] Qiu Tang, Lei Jiang, Qiong Dai, Majing Su, Hongtao Xie, and Binxing Fang. RICS-DFA: a space and time-efficient signature matching algorithm with Reduced Input Character Set. Concurrency and Computation: Practice and Experience, 29 (20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thoman:2014:CMA

[TJF14] Peter Thoman, Herbert Jordan, and Thomas Fahringer. Compiler multiversioning for automatic task granularity control. Concurrency and Computation: Practice and Experience, 26(14):2367–2385, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2019:OAD

[TJZ^{+}19] Donghai Tian, Xiaoqi Jia, Zhaolong Zhang, Li Zhan, Changzhen Hu, and Jingfeng Xue. An online approach to defeating ROP attacks. Concurrency and Computation: Practice and Experience, 31(22):e4775:1–e4775:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Totok:2010:OUR

[TK10] Alexander Totok and Vijay Karamcheti. Optimizing utilization of resource pools in web application servers. Concurrency and Computation: Practice and Experience, 22(18): 2421–2444, December 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Theobald:2002:IEC

[TKA^{+}02] Kevin B. Theobald, Rishi Kumar, Gagan Agrawal, Gerd Heber, Ruppa K. Thulasiram, and Guang R. Gao. Implementation and evaluation of a communication intensive application on the EARTH multithreaded system. Concurrency and Computation: Practice and Experience, 14(3):183–201, March 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/93513486/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
93513486{\&}PLACEBO=IE.pdf.

Talukder:2009:MDE

[TKB09] A. K. M. Khaled Ahsan Talukder, Michael Kirley, and Rajkumar Buyya. Multiobjective differential evolution for scheduling workflow applications on global Grids. Concurrency and Computation: Practice and Experience, 21(13):1742–1756, September 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Toosi:2016:SSI

[TKB16] Adel Nadjaran Toosi, Farzad Khodadadi, and Rajkumar Buyya. SipaaS: Spot instance pricing as a Service framework and its implementation in OpenStack. Concurrency and Computation: Practice and Experience, 28(13):3672–3690, September 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Taha:2013:EFS

[TKHA13] Abd-Elhamid M. Taha, Pandeli Kolomitro, Hossam Hassanein, and Najah Abu Ali. Evaluating frame structure design in WiMAX relay networks. Concurrency and Computation: Practice and Experience, 25(5):608–625, April 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tariq:2011:MSD

[TKK^{+}11] Muhammad Adnan Tariq, Boris Koldehofe, Gerald G. Koch, Imran Khan, and Kurt Rothermel. Meeting subscriberdefined QoS constraints in publish/subscribe systems. Concurrency and Computation: Practice and Experience, 23(17): 2140–2153, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Topcuoglu:2018:GPI

[TKS18] Cem Topcuoglu, Kamer Kaya, and Erkay Savas. A generic Private Information Retrieval scheme with parallel multi-exponentiations on multicore processors. Concurrency and Computation: Practice and Experience, 30(21):e4685:1– e4685:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Taneja:2017:TAT

[TKZQ17] Shubbhi Taneja, Sanjay Kulkarni, Yi Zhou, and Xiao Qin. Thermal-aware task assignments in high performance computing clusters. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Totoo:2014:PHI

[TL14] Prabhat Totoo and Hans-Wolfgang Loidl. Parallel Haskell implementations of the N-body problem. Concurrency and Computation: Practice and Experience, 26(4):987–1019, March 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2019:TCE

[TL19] Zhaojun Tang and Chuandong Li. Tracking control with event-triggered strategy for multi-agent systems with noises. Concurrency and Computation: Practice and Experience, 31 (12):e4704:1–e4704:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2017:BCS

[TLF17] Xiaoyong Tang, Xiaochun Li, and Zhuojun Fu. Budgetconstraint stochastic task scheduling on heterogeneous cloud systems. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thevenoux:2017:ASS

[TLM17] Laurent Thévenoux, Philippe Langlois, and Matthieu Martel. Automatic source-to-source error compensation of floatingpoint programs: code synthesis to optimize accuracy and time. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2018:TAN

[TLPS18] Gang Tian, Pengfei Liu, Yanjun Peng, and Chengai Sun. Tagging augmented neural topic model for semantic sparse Web service discovery. Concurrency and Computation: Practice and Experience, 30(16):??, August 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4448.

Tian:2014:SEM

[TLWZ14] Chunqi Tian, Xiaojian Liu, Lisheng Wang, and Shihong Zou. Spreading evidence models for trust propagation and aggregation in peer-to-peer networks. Concurrency and Computation: Practice and Experience, 26(2):601–614, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2017:PMC

[TLX^{+}17] Zhuo Tang, Kunkun Liu, Jinbo Xiao, Li Yang, and Zheng Xiao. A parallel k-means clustering algorithm based on redundance elimination and extreme points optimization employing MapReduce. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tari:2001:SID

[TM01] Zahir Tari and Robert Meersman. Special issue: Distributed Objects and Applications ’99. Concurrency and Computation: Practice and Experience, 13(7):503–505, June 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/83002171/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
83002171&PLACEBO=IE.pdf.

Tremblay:2003:IEP

[TMAG03] G. Tremblay, C. J. Morrone, J. N. Amaral, and G. R. Gao. Implementation of the EARTH programming model on SMP clusters: a multi-threaded language and runtime system. Concurrency and Computation: Practice and Experience, 15 (9):821–844, August 10, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2010:CUT

[TMF^{+}10] Wei Tan, Paolo Missier, Ian Foster, Ravi Madduri, David De Roure, and Carole Goble. A comparison of using Taverna and BPEL in building scientific workflows: the case of caGrid. Concurrency and Computation: Practice and Experience, 22 (9):1098–1117, June 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tolooee:2016:SFC

[TMP16] Cameron Tolooee, Matthew Malensek, and Sangmi Lee Pallickara. A scalable framework for continuous query evaluations over multidimensional, scientific datasets. Concurrency and Computation: Practice and Experience, 28(8):2546–2563, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Trebon:2007:PMC

[TMR^{+}07] Nick Trebon, Allen Morris, Jaideep Ray, Sameer Shende, and A. D. Malony. Performance modeling of component assemblies. Concurrency and Computation: Practice and Experience, 19(5):685–696, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Torres:2012:RIS

[TMS^{+}12] Erik Torres, Germán Moltó, Damià Segrelles, Ignacio Blanquer, and Vicente Hernández. A replicated information system to enable dynamic collaborations in the Grid. Concurrency and Computation: Practice and Experience, 24(14): 1668–1683, September 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tu:2007:DEH

[TMZ07] Wanqing Tu, Jogesh K. Muppala, and Hai Zhuge. Distributed end-host multicast algorithms for the Knowledge Grid. Concurrency and Computation: Practice and Experience, 19(15): 2013–2029, October 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tanaka:2015:AME

[TNH15] Kazuaki Tanaka, Hideyuki Nakanishi, and Ishiguro Hiroshi. Appearance, motion, and embodiment: unpacking avatars by fine-grained communication analysis. Concurrency and Computation: Practice and Experience, 27(11):2706–2724, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See errata [TNI16].

Tanaka:2016:EAM

[TNI16] Kazuaki Tanaka, Hideyuki Nakanishi, and Hiroshi Ishiguro. Errata: Appearance, motion, and embodiment: unpacking avatars by fine-grained communication analysis. Concurrency and Computation: Practice and Experience, 28(9):2747, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [TNH15].

Takafuji:2017:CCC

[TNIB17] Daisuke Takafuji, Koji Nakano, Yasuaki Ito, and Jacir Bordim. C2CU: a CUDA–C program generator for bulk execution of a sequential algorithm. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thulasiram:2014:ESI

[TP14] Ruppa K. Thulasiram and Manish Parashar. Editorial: Special issue on computational finance. Concurrency and Computation: Practice and Experience, 26(9):1607–1608, June 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tabik:2015:DTB

[TPGC15] Siham Tabik, Maurice Peemen, Nicolas Guil, and Henk Corporaal. Demystifying the 16 × 16 thread-block for stencils on the GPU. Concurrency and Computation: Practice and Experience, 27(18):5557–5573, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tesfatsion:2018:PPO

[TPT^{+}18] Selome Kostentions Tesfatsion, Julio Proaño, Luis Tomás, Blanca Caminero, Carmen Carrión, and Johan Tordsson. Power and performance optimization in FPGA-accelerated clouds. Concurrency and Computation: Practice and Experience, 30(18):e4526:1–e4526:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tripathi:2017:IAN

[TPV17] Atul Tripathi, Isha Pathak, and Deo Prakash Vidyarthi. Integration of analytic network process with service measurement index framework for cloud service provider selection. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2014:PAB

[TQL^{+}14] Ming Tang, Zhenlong Qiu, Weijie Li, Weijin Sun, Xiaobo Hu, and Huanguo Zhang. Power analysis based reverse engineering on the secret round function of block ciphers. Concurrency and Computation: Practice and Experience, 26(8):1531–1545, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tiampo:2002:PLS

[TRH^{+}02] K. F. Tiampo, J. B. Rundle, P. Hopper, J. Sá Martins, S. Gross, and S. McGinnis. Parallelization of a large-scale computational earthquake simulation program. Concurrency and Computation: Practice and Experience, 14(6–7):531–550, May/June 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/94515745/START;   http:

//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
94515745{\&}PLACEBO=IE.pdf.

Trunfio:2015:TLM

[Tru15] Paolo Trunfio. A two-layer model for improving the energy efficiency of file sharing peer-to-peer networks. Concurrency and Computation: Practice and Experience, 27(13): 3166–3183, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tirumalai:2007:UPH

[TRW07] Vijay Tirumalai, Kenneth G. Ricks, and Keith A. Woodbury. Using parallelization and hardware concurrency to improve the performance of a genetic algorithm. Concurrency and Computation: Practice and Experience, 19(4):443–462, March 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tajiki:2019:SDS

[TSA^{+}19] M. M. Tajiki, Mohammad Shojafar, Behzad Akbari, Stefano Salsano, and Mauro Conti. Software defined service function chaining with failure consideration for fog computing. Concurrency and Computation: Practice and Experience, 31(8): e4953:1–e4953:??, April 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Toharia:2010:GGC

[TSBR10] Pablo Toharia, Alberto Sánchez, José Luis Bosque, and Oscar D. Robles. GCViR: grid content-based video retrieval with work allocation brokering. Concurrency and Computation: Practice and Experience, 22(11):1450–1475, August 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tripathi:2018:IPC

[TSKM18] Anand Tripathi, Rahul R. Sharma, Manu Khandelwal, and Tanmay Mehta. Incremental parallel computing for continuous queries in dynamic graphs using a transactional model. Concurrency and Computation: Practice and Experience, 30(18):e4500:1–e4500:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2015:MTA

[TSL15] Bo Tang, Ravi Sandhu, and Qi Li. Multi-tenancy authorization models for collaborative cloud services. Concurrency and Computation: Practice and Experience, 27(11):2851–2868, August 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tesser:2019:PMG

[TSL^{+}19] Rafael Keller Tesser, Lucas Mello Schnorr, Arnaud Legrand, Franz Christian Heinrich, Fabrice Dupros, and Philippe O. A. Navaux. Performance modeling of a geophysics application to accelerate over-decomposition parameter tuning through simulation. Concurrency and Computation: Practice and Experience, 31(11):e5012:1–e5012:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Trunfio:2018:SIH

[TSS18] Giuseppe A. Trunfio, William Spataro, and Georgios Ch. Sirakoulis. Special issue on high performance computing in modeling and simulation. Concurrency and Computation: Practice and Experience, 30(7):??, April 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4427.

Thomas:2005:PPF

[TTD^{+}05] Shawna Thomas, Gabriel Tanase, Lucia K. Dale, Jose M. Moreira, Lawrence Rauchwerger, and Nancy M. Amato. Parallel protein folding with STAPL. Concurrency and Computation: Practice and Experience, 17(14):1643–1656, December 10, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Taboada:2011:DLC

[TTD^{+}11] Guillermo L. Taboada, Juan Touriño, Ramón Doallo, Aamir Shafi, Mark Baker, and Bryan Carpenter. Device level communication libraries for high-performance computing in Java. Concurrency and Computation: Practice and Experience, 23 (18):2382–2403, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thain:2005:DCP

[TTL05] Douglas Thain, Todd Tannenbaum, and Miron Livny. Distributed computing in practice: the Condor experience. Concurrency and Computation: Practice and Experience, 17(2– 4):323–356, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thain:2006:HML

[TTL06] Douglas Thain, Todd Tannenbaum, and Miron Livny. How to measure a large open-source distributed system. Concurrency and Computation: Practice and Experience, 18(15): 1989–2019, December 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thulasiram:2016:NIS

[TTPJ16] Ruppa K. Thulasiram, Parimala Thulasiraman, Hari Prasain, and Girish K. Jha. Nature-inspired soft computing for financial option pricing using high-performance analytics. Concurrency and Computation: Practice and Experience, 28(3): 707–728, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Toya:2010:PIA

[TTR^{+}10] Y. Toya, K. F. Tiampo, J. B. Rundle, Chien chih Chen, Hsien-Chi Li, and W. Klein. Pattern informatics approach to earthquake forecasting in 3D. Concurrency and Computation: Practice and Experience, 22(12):1569–1592, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Talia:2008:WFD

[TTV08] Domenico Talia, Paolo Trunfio, and Oreste Verta. The Weka4WS framework for distributed data mining in serviceoriented Grids. Concurrency and Computation: Practice and Experience, 20(16):1933–1951, November 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Turner:2004:SID

[Tur04] Stephen John Turner. Special issue: Distributed simulation and real-time applications. Concurrency and Computation: Practice and Experience, 16(15):1477–1481, December 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tiwari:2014:OEI

[TV14] Pawan Kumar Tiwari and Deo Prakash Vidyarthi. Observing the effect of interprocess communication in auto controlled ant colony optimization-based scheduling on computational grid. Concurrency and Computation: Practice and Experience, 26(1):241–270, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tellez-Velazquez:2018:CSI

[TVCB18] Arturo Téllez-Velázquez and Raúl Cruz-Barbosa. A CUDAstreams inference machine for non-singleton fuzzy systems. Concurrency and Computation: Practice and Experience, 30(8):??, April 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4382.

Tellez-Velazquez:2019:SIP

[TVCB19] Arturo Téllez-Velázquez and Raúl Cruz-Barbosa. A Spark image processing toolkit. Concurrency and Computation: Practice and Experience, 31(17):e5283:1–e5283:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tlais:2007:DCM

[TW07] M. Tlais and F. Weis. Distributed communication model in hierarchical Infostation systems. Concurrency and Computation: Practice and Experience, 19(8):1183–1192, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Touati:2013:SIP

[TWB13] Sid-Ahmed-Ali Touati, Julien Worms, and Sébastien Briais. Special issue papers: The Speedup-Test: a statistical methodology for programme speedup analysis and computation. Concurrency and Computation: Practice and Experience, 25(10):1410–1426, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2007:PAI

[TWN07] Ye Tian, Di Wu, and Kam-Wing Ng. Performance analysis and improvement for BitTorrent-like file sharing systems. Concurrency and Computation: Practice and Experience, 19 (13):1811–1835, September 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Taylor:2005:DCT

[TWSM05] Ian Taylor, Ian Wang, Matthew Shields, and Shalil Majithia. Distributed computing with Triana on the Grid. Concurrency and Computation: Practice and Experience, 17(9):1197–1214, August 10, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2019:LCI

[TWW^{+}19] Gang Tian, Qibo Wang, Jian Wang, Keqing He, Weidong Zhao, Panpan Gao, and Yanjun Peng. Leveraging contextual information for cold-start Web service recommendation. Concurrency and Computation: Practice and Experience, 31(17): e5195:1–e5195:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tai:2019:MLB

[TWZ^{+}19] Yonghang Tai, Lei Wei, Hailing Zhou, Qiong Li, Xiaoqiao Huang, Junsheng Shi, and Saeid Nahavandi. Machine learning-based haptic-enabled surgical navigation with security awareness. Concurrency and Computation: Practice and Experience, 31(19):e4908:1–e4908:??, October 10, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2016:CTN

[TXY^{+}16] Xiaonan Tian, Rengan Xu, Yonghong Yan, Sunita Chandrasekaran, Deepak Eachempati, and Barbara Chapman. Compiler transformation of nested loops for general purpose GPUs. Concurrency and Computation: Practice and Experience, 28(2):537–556, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tao:2017:RAG

[TXZ^{+}17] Jun Tao, Yifan Xu, Ziyi Zhang, Fuqin Feng, Fei Tong, and Fang Dong. A resource allocation game with restriction mechanism in VANET cloud. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2015:MTB

[TY15] Wenhong Tian and Chee Shin Yeo. Minimizing total busy time in offline parallel scheduling with application to energy efficiency in cloud computing. Concurrency and Computation: Practice and Experience, 27(9):2470–2488, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2012:APM

[TYHL12] Wanbin Tang, Huogen Yu, Yanfeng Han, and Shaoqian Li. An analytical performance model considering access strategy of an opportunistic spectrum sharing system. Concurrency and Computation: Practice and Experience, 24(11): 1200–1212, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2015:IBH

[TYL^{+}15] Guangping Tang, Wangdong Yang, Kenli Li, Yu Ye, Guoqing Xiao, and Keqin Li. An iteration-based hybrid parallel algorithm for tridiagonal systems of equations on multi-core architectures. Concurrency and Computation: Practice and Experience, 27(17):5076–5095, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2015:PBT

[TYTY15] Feilong Tang, Ilsun You, Can Tang, and Shui Yu. A profiling based task scheduling approach for multicore network processors. Concurrency and Computation: Practice and Experience, 27(4):855–869, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2016:QPC

[TZ16] Xiaoqing Tan and Xiaoqian Zhang. Quantum private comparison protocol with cloud quantum computing. Concurrency and Computation: Practice and Experience, 28(10): 3006–3020, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tasoulas:2019:ERR

[TZG^{+}19] Evangelos Tasoulas, Feroz Zahid, Ernst Gunnar Gran, Kyrre Begnum, Bjørn Dag Johnsen, and Tor Skeie. Efficient routing and reconfiguration in virtualized HPC environments with vSwitch-enabled lossless networks. Concurrency and Computation: Practice and Experience, 31(2):e4443:1–e4443:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Taheri:2016:GAF

[TZK16] Javid Taheri, Albert Y. Zomaya, and Samee U. Khan. Genetic algorithm in finding Pareto frontier of optimizing data transfer versus job execution in grids. Concurrency and Computation: Practice and Experience, 28(6):1715–1736, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2012:PBA

[TZKH12] Yu Tian, Zili Zhu, Fima C. Klebaner, and Kais Hamza. Pricing barrier and American options under the SABR model on the graphics processing unit. Concurrency and Computation: Practice and Experience, 24(8):867–879, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2015:CSQ

[TZLC15] Mingdong Tang, Tingting Zhang, Jianxun Liu, and Jinjun Chen. Cloud service QoS prediction via exploiting collaborative filtering and location-based data smoothing. Concurrency and Computation: Practice and Experience, 27(18): 5826–5839, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2013:FSS

[TZYL13] Wanbin Tang, Jing Zhou, Huogen Yu, and Shaoqian Li. A fair scheduling scheme based on collision statistics for cognitive radio networks. Concurrency and Computation: Practice and Experience, 25(9):1091–1100, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Unat:2018:SIH

[UA18] Didem Unat and Mehmet S. Aktas. Special issue on high performance computing conference (BASARIM-2017). Concurrency and Computation: Practice and Experience, 30(21): e4926:1–e4926:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ungrangsi:2009:SOR

[UAW09] R. Ungrangsi, C. Anutariya, and V. Wuwongse. SQORE: an ontology retrieval framework for the next generation Web. Concurrency and Computation: Practice and Experience, 21 (5):651–671, April 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Utrera:2018:AIC

[UGM18] Gladys Utrera, Marisa Gil, and Xavier Martorell. Analyzing the impact of communication imbalance in high-speed networks. Concurrency and Computation: Practice and Experience, 30(7):??, April 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4394.

Underwood:2003:API

[ULS03] Keith D. Underwood, Walter B. Ligon III, and Ron R. Sass. Analysis of a prototype intelligent network interface. Concurrency and Computation: Practice and Experience, 15(7–8): 751–777, June/July 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Urbani:2013:SRD

[UMD^{+}13] Jacopo Urbani, Jason Maassen, Niels Drost, Frank Seinstra, and Henri Bal. Scalable RDF data compression with MapReduce. Concurrency and Computation: Practice and Experience, 25(1):24–39, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ururahy:2004:PCG

[UR04] Cristina Ururahy and Noemi Rodriguez. Programming and coordinating Grid environments and applications. Concurrency and Computation: Practice and Experience, 16(5):543– 549, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ur:2007:SIP

[Ur07] Shmuel Ur. Special issue: Parallel and Distributed Systems: Testing and Debugging (PADTAD). Concurrency and Computation: Practice and Experience, 19(3):265–266, March 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Voorsluys:2007:FCB

[VAC^{+}07] William Voorsluys, Eliane Araújo, Walfredo Cirne, Carlos O. Galvão, Enio P. Souza, and Enilson P. Cavalcanti. Fostering collaboration to better manage water resources. Concurrency and Computation: Practice and Experience, 19(12): 1609–1620, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanAmesfoort:2012:PAC

[vAVS12] Alexander S. van Amesfoort, Ana Lucia Varbanescu, and Henk J. Sips. Parallel application characterization with quantitative metrics. Concurrency and Computation: Practice and Experience, 24(5):445–462, April 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Varisteas:2016:PAL

[VB16] Georgios Varisteas and Mats Brorsson. Palirria: accurate online parallelism estimation for adaptive work-stealing. Concurrency and Computation: Practice and Experience, 28(2): 472–491, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Venugopal:2006:GSB

[VBW06] Srikumar Venugopal, Rajkumar Buyya, and Lyle Winton. A Grid service broker for scheduling e-Science applications on global data Grids. Concurrency and Computation: Practice and Experience, 18(6):685–699, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Venticinque:2016:EID

[VC16] Salvatore Venticinque and David Camacho. Editorials: Intelligent distributed computing. Concurrency and Computation: Practice and Experience, 28(4):1257–1260, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Verdu:2016:DWW

[VCP16] Javier Verdú, Juan José Costa, and Alex Pajuelo. Dynamic Web worker pool management for highly parallel JavaScript Web applications. Concurrency and Computation: Practice and Experience, 28(13):3525–3539, September 10, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vu:2013:SIP

[VCW13] V. T. Vu, G. Cats, and L. Wolters. Special issue papers: Graphics processing unit optimizations for the dynamics of the HIRLAM weather forecast model. Concurrency and Computation: Practice and Experience, 25(10):1376–1393, July 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vadhiyar:2005:SAG

[VD05] Sathish S. Vadhiyar and Jack J. Dongarra. Self adaptivity in Grid computing. Concurrency and Computation: Practice and Experience, 17(2–4):235–257, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanderAalst:2010:RMG

[vdABST10] Wil van der Aalst, Carmen Bratosin, Natalia Sidorova, and Nikola Trčka. A reference model for grid architectures and its validation. Concurrency and Computation: Practice and Experience, 22(11):1365–1385, August 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vanmechelen:2009:MBG

[VDB09] Kurt Vanmechelen, Wim Depoorter, and Jan Broeckhove. Market-based grid resource co-allocation and reservation for applications with hard deadlines. Concurrency and Computation: Practice and Experience, 21(18):2270–2297, December 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vargas:2007:GTS

[VDdN^{+}07] Patrícia Kayser Vargas, Inês C. Dutra, Vinícius D. do Nascimento, Lucas A. S. Santos, Luciano C. da Silva, Cláudio F. R. Geyer, and Bruno Schulze. GRAND: toward scalability in a Grid environment. Concurrency and Computation: Practice and Experience, 19(14):1991–2009, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vdKuijl:2010:RMO

[vdKEL10] Alexander v. d. Kuijl, Michael T. M. Emmerich, and Hui Li. A robust multi-objective resource allocation scheme incorporating uncertainty and service differentiation. Concurrency and Computation: Practice and Experience, 22(3):314–328, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vapirev:2015:IRC

[VDL^{+}15] A. Vapirev, J. Deca, G. Lapenta, S. Markidis, I. Hur, and J.-L. Cambier. Initial results on computational performance of Intel many integrated core, Sandy Bridge, and graphical processing unit architectures: implementation of a 1D C++/OpenMP electrostatic particle-in-cell code. Concurrency and Computation: Practice and Experience, 27(3):581– 593, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

VanderHeyden:2003:CPJ

[VDPC03] W. B. VanderHeyden, E. D. Dendy, and N. T. Padial-Collins. CartaBlanca — a pure-Java, component-based systems simulation tool for coupled nonlinear physics on unstructured grids — an update. Concurrency and Computation: Practice and Experience, 15(3–5):431–458, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanderSteen:2006:ICF

[vdS06a] Aad J. van der Steen. Issues in computational frameworks. Concurrency and Computation: Practice and Experience, 18 (2):141–150, February 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanderSteen:2006:SIC

[vdS06b] Aad J. van der Steen. Special issue: Computational frameworks (have more fun with your computational models). Concurrency and Computation: Practice and Experience, 18(2):

137–139, February 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vetter:2005:EHP

[VdSK^{+}05] Jeffrey S. Vetter, Bronis R. de Supinski, Lynn Kissel, John May, and Sheila Vaidya. Evaluating high-performance computers. Concurrency and Computation: Practice and Experience, 17(10):1239–1270, August 25, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanEngelen:2006:PTE

[vEGW06] Robert A. van Engelen, Kyle A. Gallivan, and Burt Walsh. Parametric timing estimation with Newton–Gregory formulae. Concurrency and Computation: Practice and Experience, 18(11):1435–1463, September 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Verge:2017:HAS

[VEJD17] Adrien Vergé, Naser Ezzati-Jivan, and Michel R. Dagenais. Hardware-assisted software event tracing. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vinas:2017:FDS

[VFAD17] Moisés Viñas, Basilio B. Fraguela, Diego Andrade, and Ramón Doallo. Facilitating the development of stencil applications using the Heterogeneous Programming Library. Concurrency and Computation: Practice and Experience, 29(12): ??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vinas:2018:HDC

[VFAD18] M. Viñas, B. B. Fraguela, D. Andrade, and R. Doallo. Heterogeneous distributed computing based on high-level abstractions. Concurrency and Computation: Practice and Experience, 30(17):e4664:1–e4664:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vazquez:2011:NAS

[VFG11] F. Vázquez, J. J. Fernández, and E. M. Garzón. A new approach for sparse matrix vector product on NVIDIA GPUs.

Concurrency and Computation: Practice and Experience, 23 (8):815–826, June 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valcke:2006:PEE

[VGL06] Sophie Valcke, Eric Guilyardi, and Claes Larsson. PRISM and ENES: a European approach to Earth system modelling. Concurrency and Computation: Practice and Experience, 18 (2):247–262, February 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valadares:2016:DTD

[VGL16] Arthur Valadares, Eugenia Gabrielova, and Cristina Videira Lopes. On designing and testing distributed virtual environments. Concurrency and Computation: Practice and Experience, 28(12):3291–3312, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Verma:2016:DRD

[VGN^{+}16] Manish Verma, G. R. Gangadharan, Nanjangud C. Narendra, Ravi Vadlamani, Vidyadhar Inamdar, Lakshmi Ramachandran, Rodrigo N. Calheiros, and Rajkumar Buyya. Dynamic resource demand prediction and allocation in multi-tenant service clouds. Concurrency and Computation: Practice and Experience, 28(17):4429–4442, December 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vassev:2012:AAS

[VH12] Emil Vassev and Mike Hinchey. The ASSL approach to specifying self-managing embedded systems. Concurrency and Computation: Practice and Experience, 24(16):1860–1878, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Veldema:2003:RTO

[VHBB03] R. Veldema, R. F. H. Hofman, R. A. F. Bhoedjang, and H. E. Bal. Run-time optimizations for a Java DSM implementation. Concurrency and Computation: Practice and Experience, 15 (3–5):299–316, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanHemert:2011:GWB

[vHKT^{+}11] Jano van Hemert, Jos Koetsier, Livia Torterolo, Ivan Porro, Maurizio Melato, and Roberto Barbera. Generating webbased user interfaces for computational science. Concurrency and Computation: Practice and Experience, 23(3):256–268, March 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanHeiningen:2008:BMD

[vHMB08] Willem van Heiningen, Steve MacDonald, and Tim Brecht. Babylon: middleware for distributed, parallel, and mobile Java applications. Concurrency and Computation: Practice and Experience, 20(10):1195–1224, July 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanHees:2003:PDA

[vHvdSvL03] Fons van Hees, Aad J. van der Steen, and Peter Jan van Leeuwen. A parallel data assimilation model for oceanographic observations. Concurrency and Computation: Practice and Experience, 15(13):1191–1204, November 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vivien:2003:OFS

[Viv03] Frédéric Vivien. On the optimality of Feautrier’s scheduling algorithm. Concurrency and Computation: Practice and Experience, 15(11–12):1047–1068, September 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vijayarajan:2019:NCS

[VJ19] Subramanian Maruthathurai Vijayarajan and P. Jaganathan. A novel comparative study on breast cancer detection using different types of classification techniques. Concurrency and Computation: Practice and Experience, 31(14):e4939:1– e4939:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Veldema:2005:OCN

[VJHB05] Ronald Veldema, Ceriel J. H. Jacobs, Rutger F. H. Hofman, and Henri E. Bal. Object combining: a new aggressive optimization for object intensive programs. Concurrency and Computation: Practice and Experience, 17(5–6):439–464, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valvaag:2013:CHP

[VJK13] S. V. Valvåg, Dag Johansen, and Åge Kvalnes. Cogset: a high performance MapReduce engine. Concurrency and Computation: Practice and Experience, 25(1):2–23, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vitek:2012:ISI

[VK12] Jan Vitek and Tomas Kalibera. Introduction to the Special Issue on Java Technologies for Real-Time and Embedded Systems. Concurrency and Computation: Practice and Experience, 24(8):751–752, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vishnu:2009:TAH

[VKM^{+}09] Abhinav Vishnu, Matthew Koop, Adam Moody, Amith Mamidala, Sundeep Narravula, and Dhabaleswar K. Panda. Topology agnostic hot-spot avoidance with InfiniBand. Concurrency and Computation: Practice and Experience, 21(3): 301–319, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valero-Lara:2017:RMR

[VL17] Pedro Valero-Lara. Reducing memory requirements for large size LBM simulations on GPUs. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vonLaszewski:2007:PVG

[vLDA07] Gregor von Laszewski, Jonathan DiCarlo, and Bill Allcock. A portal for visualizing Grid usage. Concurrency and Computation: Practice and Experience, 19(12):1683–1692, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vonLaszewski:2011:EDF

[vLDW11] Gregor von Laszewski, Jai Dayal, and Lizhe Wang. eMOLST: a documentation flow for distributed health informatics. Concurrency and Computation: Practice and Experience, 23(16):

1857–1867, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vinas:2013:MGS

[VLF^{+}13] M. Viñas, J. Lobeiras, B. B. Fraguela, M. Arenaz, M. Amor, J. A. García, M. J. Castro, and R. Doallo. A multi-GPU shallow-water simulation with transport of contaminants. Concurrency and Computation: Practice and Experience, 25 (8):1153–1169, June 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vonLaszewski:2001:JCG

[vLFGL01] Gregor von Laszewski, Ian Foster, Jarek Gawor, and Peter Lane. A Java commodity grid kit. Concurrency and Computation: Practice and Experience, 13(8–9):645–662, July/ August 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/84503216/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
84503216&PLACEBO=IE.pdf.

vonLaszewski:2002:FJC

[vLGL^{+}02] Gregor von Laszewski, Jarek Gawor, Peter Lane, Nell Rehn, and Mike Russell. Features of the Java Commodity Grid Kit. Concurrency and Computation: Practice and Experience, 14(13–15):1045–1055, November/December 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valero-Lara:2017:HCA

[VLJ17] Pedro Valero-Lara and Johan Jansson. Heterogeneous CPU+GPU approaches for mesh refinement over Lattice-Boltzmann simulations. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valero-Lara:2018:CCC

[VLMPS^{+}18] Pedro Valero-Lara, Ivan Martínez-Pérez, Raül Sirvent, Xavier Martorell, and Antonio J. Peña. cuThomasBatch and cuThomasVBatch, CUDA routines to compute batch of tridiagonal systems on NVIDIA GPUs. Concurrency and Computation: Practice and Experience, 30(24):e4909:1–e4909:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vonLaszewski:2002:CSD

[vLRF^{+}02] Gregor von Laszewski, Michael Russell, Ian Foster, John Shalf, Gabrielle Allen, Greg Daues, Jason Novotny, and Edward Seidel. Community software development with the Astrophysics Simulation Collaboratory. Concurrency and Computation: Practice and Experience, 14(13–15):1289–1301, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valery:2019:CCG

[VLW19] Olivier Valery, Pangfeng Liu, and Jan-Jan Wu. A collaborative CPU–GPU approach for deep learning on mobile devices. Concurrency and Computation: Practice and Experience, 31(17):e5225:1–e5225:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanNieuwpoort:2005:IFE

[vNMW^{+}05] Rob V. van Nieuwpoort, Jason Maassen, Gosia Wrzesińska, Rutger F. H. Hofman, Ceriel J. H. Jacobs, Thilo Kielmann, and Henri E. Bal. Ibis: a flexible and efficient Java-based Grid programming environment. Concurrency and Computation: Practice and Experience, 17(7–8):1079–1107, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vonOheimb:2001:HLJ

[vO01] David von Oheimb. Hoare logic for Java in Isabelle/ HOL. Concurrency and Computation: Practice and Experience, 13(13):1173–1214, November 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/88011338/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=88011338&PLACEBO=IE.
pdf.

Vinh:2015:SDB

[VO15] Phan Van Vinh and Hoon Oh. A slot demand-based path reservation approach for the timely and reliable delivery of bursty traffic in WMSNs. Concurrency and Computation:

Practice and Experience, 27(10):2569–2587, July 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vignesh:2019:CLB

[VP19] V. Vignesh and K. Premalatha. Channel load based interference-aware scheduling in military communication environment. Concurrency and Computation: Practice and Experience, 31(12):e4738:1–e4738:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Villar-Rodriguez:2016:NML

[VRDTB^{+}16] E. Villar-Rodríguez, J. Del Ser, A. I. Torre-Bastida, M. N. Bilbao, and S. Salcedo-Sanz. A novel machine learning approach to the detection of identity theft in social networks based on emulated attack instances and support vector machines. Concurrency and Computation: Practice and Experience, 28(4):1385–1395, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vonRonne:2009:SBC

[vRGNP09] Jeffery von Ronne, Andreas Gampe, David Niedzielski, and Kleanthis Psarris. Safe bounds check annotations. Concurrency and Computation: Practice and Experience, 21(1): 41–57, January 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanReeuwijk:2003:SSE

[vRKS03] C. van Reeuwijk, F. Kuijlman, and H. J. Sips. Spar: a set of extensions to Java for scientific computation. Concurrency and Computation: Practice and Experience, 15(3–5): 277–297, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vaquero:2013:ECS

[VRMB13] Luis Miguel Vaquero, Luis Rodero-Merino, and Rajkumar Buyya. Editorials: Cloud scalability: building the Millennium Falcon. Concurrency and Computation: Practice and Experience, 25(12):1623–1627, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanReeuwijk:2005:ATJ

[vRS05] C. van Reeuwijk and H. J. Sips. Adding tuples to Java: a study in lightweight data structures. Concurrency and Computation: Practice and Experience, 17(5–6):423–438, April/ May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vega-Rodriguez:2015:EPB

[VRSJ15] Miguel A. Vega-Rodríguez and Sergio Santander-Jiménez. Editorials: Parallelism-based technologies in bioinformatics and biomedicine: a view from diverse perspectives. Concurrency and Computation: Practice and Experience, 27(18): 5473–5475, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valsalam:2002:FHP

[VS02] Vinod Valsalam and Anthony Skjellum. A framework for high-performance matrix multiplication based on hierarchical abstractions, algorithms and optimized lowlevel kernels. Concurrency and Computation: Practice and Experience, 14(10):805–839, August 25, 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/97517968/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=97517968{\&}PLACEBO=
IE.pdf.

Vasko:2011:OGF

[VŠ11] A. Vaško and M. Šrámek. Optimizing Gaussian filtering of volumetric data using SSE. Concurrency and Computation: Practice and Experience, 23(1):100–116, January 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Velayudhan:2019:CAD

[VS19] Savyan P. Velayudhan and Mary Saira Bhanu Somasundaram. Compromised account detection in online social networks: a survey. Concurrency and Computation: Practice and Experience, 31(20):e5346:1–e5346:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanStokkum:2006:PSE

[vSB06] Ivo H. M. van Stokkum and Henri E. Bal. A Problem Solving Environment for interactive modelling of multiway data. Concurrency and Computation: Practice and Experience, 18 (2):263–269, February 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vitello:2015:MAD

[VSB^{+}15] Fabio Vitello, Eva Sciacca, Ugo Becciani, Alessandro Costa, Piero Massimino, Éva Takács, and Balázs Szakál. Mobile application development exploiting science gateway technologies. Concurrency and Computation: Practice and Experience, 27(16):4361–4376, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vondra:2017:MCA

[VŠC17] T. Vondra, J. Šedivý, and J. M. Castro. Modifying CloudSim to accurately simulate interactive services for cloud autoscaling. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vashisht:2017:SRC

[VSK17] Priyanka Vashisht, Anju Sharma, and Rajesh Kumar. Strategies for replica consistency in data grid — a comprehensive survey. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vidal:2009:ASG

[VSKK09] A. C. T. Vidal, F. J. S. Silva, S. T. Kofuji, and F. Kon. Applying semantics to grid middleware. Concurrency and Computation: Practice and Experience, 21(13):1725–1741, September 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Varbanescu:2009:EAM

[VSR^{+}09] Ana Lucia Varbanescu, Henk Sips, Kenneth A. Ross, Qiang Liu, Apostol (Paul) Natsev, John R. Smith, and Lurng-Kuo Liu. Evaluating application mapping scenarios on the Cell/ B.E. Concurrency and Computation: Practice and Experience, 21(1):85–100, January 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Varadharajan:2015:SWM

[VT15] Vijay Varadharajan and Udaya Tupakula. Securing wireless mobile nodes from distributed denial-of-service attacks. Concurrency and Computation: Practice and Experience, 27(15): 3794–3815, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Voulgaris:2007:PGB

[VvSI07] Spyros Voulgaris, Maarten van Steen, and Konrad Iwanicki. Proactive gossip-based management of semantic overlay networks. Concurrency and Computation: Practice and Experience, 19(17):2299–2311, December 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanWaveren:2002:CGH

[vWAH^{+}02] Matthijs van Waveren, Cliff Addison, Peter Harrison, Dave Orange, Norman Brown, and Hidetoshi Iwashita. Code generator for the HPF Library and Fortran 95 transformational functions. Concurrency and Computation: Practice and Experience, 14(8–9):589–602, July/ August 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/95016135/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
95016135{\&}PLACEBO=IE.pdf.

VanAalsburg:2010:IED

[VYK^{+}10] Jordan Van Aalsburg, M. Burak Yikilmaz, Oliver Kreylos, Louise H. Kellogg, and John B. Rundle. Interactive editing of digital fault models. Concurrency and Computation: Practice and Experience, 22(12):1720–1731, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vadivel:2019:ECS

[VYKM19] P. Sundara Vadivel, D. Yuvaraj, S. Navaneetha Krishnan, and S. R. Mathusudhanan. An efficient CBIR system based on color histogram, edge, and texture features. Concurrency and Computation: Practice and Experience, 31(12):e4994:1– e4994:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vochin:2019:IVN

[VZB19] Marius Vochin, Sorin Zoican, and Eugen Borcoci. Intelligent vehicle navigation system with assistance and alerting capabilities. Concurrency and Computation: Practice and Experience, 31(22):e4402:1–e4402:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2012:SIT

[WAD12] Guojun Wang and Ahmed Y. Al-Dubai. Special issue of trust, security, and privacy for emerging applications in computer and information systems. Concurrency and Computation: Practice and Experience, 24(17):1975–1976, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:DDI

[Wan18a] Guang Wang. Design of damage identification algorithm for mechanical structures based on convolutional neural network. Concurrency and Computation: Practice and Experience, 30(24):e4891:1–e4891:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:TDF

[Wan18b] Xiaoying Wang. Three-dimensional flow optimization algorithm in complex scene based on differential equation. Concurrency and Computation: Practice and Experience, 30(22): e4657:1–e4657:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Walker:2007:SIS

[WAS07] D. W. Walker, M. P. Atkinson, and I. Sommerville. Special issue: Selected Papers from the 2004 U.K. e-Science All Hands Meeting (AHM 2004). Concurrency and Computation: Practice and Experience, 19(2):129–131, February 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Welch:2007:AMS

[WBB^{+}07] Von Welch, Jim Barlow, James Basney, Doru Marcusiu, and Nancy Wilkins-Diehr. A AAAA model to support science gateways with community accounts. Concurrency and Computation: Practice and Experience, 19(6):893–904, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Watson:2002:WSD

[WBC^{+}02] William A. Watson III, Ian Bird, Jie Chen, Bryan Hess, Andy Kowalski, and Ying Chen. A Web services data analysis Grid. Concurrency and Computation: Practice and Experience, 14(13–15):1303–1311, November/December 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weber:2017:WBV

[WBC^{+}17] Gunther H. Weber, Mark S. Bandstra, Daniel H. Chivers, Hamdy H. Elgammal, Valerie Hendrix, John Kua, Jonathan S. Maltz, Krishna Muriki, Yeongshnn Ong, Kai Song, Michael J. Quinlan, Lavanya Ramakrishnan, and Brian J. Quiter. Webbased visual data exploration for improved radiological source detection. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Williams:2003:MWI

[WBD^{+}03] Roy Williams, Bruce Berriman, Ewa Deelman, John Good, Joseph Jacob, Carl Kesselman, Carol Lonsdale, Seb Oliver, and Thomas A. Prince. Multi-wavelength image space: another Grid-enabled science. Concurrency and Computation: Practice and Experience, 15(6):539–549, May 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wakil:2019:FLB

[WBD^{+}19] Karzan Wakil, Arshad Badfar, Pooyan Dehghani, Seyed Mojtaba Shoja Sadati, and Nima Jafari Navimipour. A fuzzy logic-based method for solving the scheduling problem in the cloud environments using a non-dominated sorted algorithm. Concurrency and Computation: Practice and Experience, 31(17):e5185:1–e5185:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2008:SOA

[WBHW08] H. Wang, K. W. Brodlie, J. W. Handley, and J. D. Wood. Service-oriented approach to collaborative visualization. Concurrency and Computation: Practice and Experience, 20(11): 1289–1301, August 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Welch:2010:ABS

[WBM^{+}10] Peter Welch, Neil Brown, James Moores, Kevin Chalmers, and Bernhard Sputh. Alting barriers: synchronisation with choice in Java using JCSP. Concurrency and Computation: Practice and Experience, 22(8):1049–1062, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2016:FPS

[WBO16] Leyuan Wang, Sean Baxter, and John D. Owens. Fast parallel skew and prefix-doubling suffix array construction on the GPU. Concurrency and Computation: Practice and Experience, 28(12):3466–3484, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2010:TLC

[WBZ10] Guojun Wang, Md. Zakirul Alam Bhuiyan, and Li Zhang. Two-level cooperative and energy-efficient tracking algorithm in wireless sensor networks. Concurrency and Computation: Practice and Experience, 22(4):518–537, March 25, 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Watson:2008:SIS

[WC08] Paul Watson and Simon Cox. Special issue: Selected papers from the 2005 U.K. e-Science All Hands Meeting (AHM 2005). Concurrency and Computation: Practice and Experience, 20(3):191–193, March 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2008:GCS

[WCA08] Shaowen Wang, Mary Kathryn Cowles, and Marc P. Armstrong. Grid computing of spatial statistics: using the Tera-Grid for G(d) analysis. Concurrency and Computation: Practice and Experience, 20(14):1697–1720, September 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wells:2004:LIJ

[WCC04] G. C. Wells, A. G. Chalmers, and P. G. Clayton. Linda implementations in Java for concurrent systems. Concurrency and Computation: Practice and Experience, 16(10): 1005–1022, August 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2005:SOJ

[WCCL05] Pu-Chen Wei, Chung-Hsin Chen, Cheng-Wei Chen, and Jenq-Kuen Lee. Support and optimization of Java RMI over a Bluetooth environment. Concurrency and Computation: Practice and Experience, 17(7–8):967–989, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Woods:2007:PFD

[WCH^{+}07] Brent Woods, Bradley Clymer, Johannes Heverhagen, Michael Knopp, Joel Saltz, and Tahsin Kurc. Parallel fourdimensional Haralick texture analysis for disk-resident image datasets. Concurrency and Computation: Practice and Experience, 19(1):65–87, January 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2010:IGP

[WCL^{+}10] Yongwei Wu, Gang Chen, Jia Liu, Rui Fang, Guangwen Yang, and Weimin Zheng. Improving grid performance by dynamically deploying applications. Concurrency and Computation: Practice and Experience, 22(13):1945–1967, September 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wen:2013:SIP

[WCLC13] Yiping Wen, Zhigang Chen, Jianxun Liu, and Jinjun Chen. Special issue papers: Mining batch processing workflow models from event logs. Concurrency and Computation: Practice and Experience, 25(13):1928–1942, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2012:IHE

[WCLH12] Tin-Yu Wu, Wei Chen, Wei-Tsong Lee, and Yong-Ping Huang. Improving handoff efficiency by IS-FMIPv6 based on IEEE 802.21. Concurrency and Computation: Practice and Experience, 24(4):371–382, March 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2014:PAC

[WCR^{+}14] Jiechen Wang, Can Cui, Yikang Rui, Liang Cheng, Yingxia Pu, Wenzhou Wu, and Zhenyu Yuan. A parallel algorithm for constructing Voronoi diagrams based on point-set adaptive grouping. Concurrency and Computation: Practice and Experience, 26(2):434–446, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wen:2019:FSI

[WCWB19] Sheng Wen, Aniello Castiglione, Tian Wang, and Md Zakirul Alam Bhuiyan. Forward to the special issue of the 9th International Symposium on Cyberspace Safety and Security (CSS 2017). Concurrency and Computation: Practice and Experience, 31(22):e5535:1–e5535:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:STF

[WCZ^{+}18] Zijian Wang, Lei Cao, Zuo Zhang, Xiaoliang Gong, Yaoru Sun, and Haoran Wang. Short time Fourier transformation and deep neural networks for motor imagery brain computer interface recognition. Concurrency and Computation: Practice and Experience, 30(16):e4413:1–e4413:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2016:ALG

[WCZX16] Zhuowei Wang, Lianglun Cheng, Wuqing Zhao, and Naixue Xiong. An architecture-level graphics processing unit energy model. Concurrency and Computation: Practice and Experience, 28(10):2795–2810, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wilkins-Diehr:2007:SIS

[WD07] Nancy Wilkins-Diehr. Special issue: Science gateways — common community interfaces to Grid resources. Concurrency and Computation: Practice and Experience, 19(6):743– 749, April 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2014:CGH

[WDG^{+}14] Yueqing Wang, Yong Dou, Song Guo, Yuanwu Lei, and Dan Zou. CPU–GPU hybrid parallel strategy for cosmological simulations. Concurrency and Computation: Practice and Experience, 26(3):748–765, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wilkins-Diehr:2015:ESG

[WDGK15] Nancy Wilkins-Diehr, Sandra Gesing, and Tamas Kiss. Editorial: Science Gateway Workshops 2013 special issue conference publications. Concurrency and Computation: Practice and Experience, 27(2):253–257, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2010:BBA

[WDLL10] Jieping Wang, Xiaoyong Du, Jiaheng Lu, and Wei Lu. Bucket-based authentication for outsourced databases. Concurrency and Computation: Practice and Experience, 22(9): 1160–1180, June 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wilkins-Diehr:2014:EXS

[WDM14] Nancy Wilkins-Diehr and Amit Majumdar. Editorial: XSEDE13 Special Issue Conference Publications. Concurrency and Computation: Practice and Experience, 26(13): 2107–2111, September 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:PNB

[WDQ^{+}18] Lu Wang, YuYue Du, Man Qi, HongDa Qi, and ZhaoYang He. Petri net-based deviation detection between a process model with loop semantics and event logs. Concurrency and Computation: Practice and Experience, 30(23):e4419:1–e4419:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Watanabe:2018:SEE

[WDT18] Ryo Watanabe, Dilawaer Duolikun, and Makoto Takizawa. Simple estimation and energy-aware migration models of virtual machines in a server cluster. Concurrency and Computation: Practice and Experience, 30(21):e4771:1–e4771:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2015:CAM

[WDW^{+}15] Kaimin Wei, Mianxiong Dong, Jian Weng, Guangzhou Shi, Kaoru Ota, and Ke Xu. Congestion-aware message forwarding in delay tolerant networks: a community perspective. Concurrency and Computation: Practice and Experience, 27 (18):5722–5734, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:RTS

[WF18] Quan Wang and Weiping Fu. Research on traffic sign detection algorithm based on deep learning. Concurrency and Computation: Practice and Experience, 30(22):e4675:1– e4675:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:ASA

[WFHT17] Xianghui Wang, Zhiyong Feng, Keman Huang, and Wei Tan. An automatic self-adaptation framework for service-based process based on exception handling. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:ISV

[WFJ^{+}17] Jianfei Wang, Fengfeng Fan, Li Jiang, Xiaoyao Liang, and Naifeng Jing. Incorporating selective victim cache into GPGPU for high-performance computing. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:TLA

[WFKS18] Baokang Wang, Yuki Fukazawa, Toshio Kondo, and Takahiro Sasaki. Tile/line access cache memory based on a multilevel Z-order tiling data layout. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4375.

Wang:2019:VCS

[WFS^{+}19] Youquan Wang, Changjian Fang, Dongqin Shen, Zhiang Wu, and Jie Cao. Verifying the claimed sale-ranking trustworthy: a maximum marginal relevance-based ranking method. Concurrency and Computation: Practice and Experience, 31(24): e5466:1–e5466:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wroe:2007:RWS

[WGG^{+}07] Chris Wroe, Carole Goble, Antoon Goderis, Phillip Lord, Simon Miles, Juri Papay, Pinar Alper, and Luc Moreau. Recycling workflows and services through discovery and reuse. Concurrency and Computation: Practice and Experience, 19 (2):181–194, February 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2015:SSW

[WGP^{+}15] Hai H. Wang, Nick Gibbins, Terry Payne, Alina Patelli, and Yangang Wang. A survey of Semantic Web Services formalisms. Concurrency and Computation: Practice and Experience, 27(15):4053–4072, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:TSF

[WGQ^{+}18] Qicong Wang, Dingxi Gong, Man Qi, Yehu Shen, and Yunqi Lei. Temporal sparse feature auto-combination deep network for video action recognition. Concurrency and Computation: Practice and Experience, 30(16):e4487:1–e4487:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:SLB

[WGY^{+}19] Yuxiang Wang, Zhangpeng Ge, Haijiang Yan, Xiaoliang Xu, and Yixing Xia. Semantic locality-based approximate knowledge graph query. Concurrency and Computation: Practice and Experience, 31(24):e5345:1–e5345:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2006:CJS

[WGZL06] Qingjiang Wang, Xiaolin Gui, Shouqi Zheng, and Yang Liu. De-centralized job scheduling on computational Grids using distributed backfilling. Concurrency and Computation: Practice and Experience, 18(14):1829–1838, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Watson:2010:SCC

[WHW10] Paul Watson, Hugo Hiden, and Simon Woodman. e-Science Central for CARMEN: science as a service. Concurrency and Computation: Practice and Experience, 22(17):2369–2380, December 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:PPT

[WHX19] Fei Wang, Yi Hong, and Cong Xu. PPLTCAM: a parallel TCAM-based IP address lookup structure with high incremental update performance. Concurrency and Computation: Practice and Experience, 31(10):e4865:1–e4865:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:SFL

[WHXB19] Cheng Wang, Zhuo Hu, Ming Xie, and Yuxiang Bian. Sustainable facility location-allocation problem under uncertainty. Concurrency and Computation: Practice and Experience, 31(9):e4521:1–e4521:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2015:EPP

[WHXzL15] Yong Wang, Jie Hou, Yun Xia, and Hong zong Li. Efficient privacy preserving matchmaking for mobile social networking. Concurrency and Computation: Practice and Experience, 27 (12):2924–2937, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wismuller:2002:SIE

[Wis02] Roland Wismüller. Special issue: Euro-Par 2000. Concurrency and Computation: Practice and Experience, 14 (3):163–164, March 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/93513487/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=93513487{\&}PLACEBO=IE.pdf.

Wittenburg:2010:AAL

[Wit10] Peter Wittenburg. Archiving and accessing language resources. Concurrency and Computation: Practice and Experience, 22(17):2354–2368, December 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2009:IDV

[WJ09] Qin Wang and Joseph JaJa. Interactive direct volume rendering on desktop multicore processors. Concurrency and Computation: Practice and Experience, 21(17):2199–2212, December 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2012:MSI

[WJ12] Liejun Wang and Zhenhong Jia. Multi-scale image segmentation algorithm based on support vector machine approximation criteria. Concurrency and Computation: Practice and Experience, 24(11):1231–1238, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Welc:2006:RTJ

[WJH06] Adam Welc, Suresh Jagannathan, and Antony L. Hosking. Revocation techniques for Java concurrency. Concurrency and Computation: Practice and Experience, 18(12):1613– 1656, October 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:MCF

[WJJM17] Yufeng Wang, Xueyu Jia, Qun Jin, and Jianhua Ma. Mobile crowdsourcing: framework, challenges, and solutions. Concurrency and Computation: Practice and Experience, 29(3): ??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Witt:2018:SBA

[WJKS18] Michael Witt, Christoph Jansen, Dagmar Krefting, and Achim Streit. Sandboxing of biomedical applications in Linux containers based on system call evaluation. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4484.

Woodward:2009:FEC

[WJLD09] Paul R. Woodward, Jagan Jayaraj, Pei-Hung Lin, and William Dai. First experience of compressible gas dynamics simulation on the Los Alamos Roadrunner machine. Concurrency and Computation: Practice and Experience, 21(17): 2160–2175, December 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weber:2014:SOB

[WJP14] Rick Weber, David D. Jenkins, and Gregory D. Peterson. Specmaster: an OpenCL-based peptide search engine for tandem mass spectrometry. Concurrency and Computation: Practice and Experience, 26(6):1234–1248, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Waechter:2014:UGP

[WJT^{+}14] Michael Waechter, Kathrin Jaeger, Daniel Thuerck, Stephanie Weissgraeber, Sven Widmer, Michael Goesele, and Kay Hamacher. Using graphics processing units to investigate molecular coevolution. Concurrency and Computation: Practice and Experience, 26(6):1278–1296, April 25, 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:SAO

[WJXZ18] Yuxiang Wang, Jiahui Jin, Xiaoliang Xu, and Longbin Zhang. Skew-aware online aggregation over joins through guided sampling. Concurrency and Computation: Practice and Experience, 30(20):e4695:1–e4695:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2016:DLB

[WJYH16] Yuzhu Wang, Jinrong Jiang, Huang Ye, and Juanxiong He. A distributed load balancing algorithm for climate big data processing over a multi-core CPU cluster. Concurrency and Computation: Practice and Experience, 28(15):4144–4160, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wilde:2007:DHP

[WK07] Torsten Wilde and James A. Kohl. Debugging highperformance component-based applications. Concurrency and Computation: Practice and Experience, 19(5):667–684, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wellings:2012:AEH

[WK12] Andy Wellings and MinSeong Kim. Asynchronous event handling and Safety Critical Java. Concurrency and Computation: Practice and Experience, 24(8):813–832, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Walker:2019:AMI

[WKB^{+}19] David W. Walker, Stephan C. Kramer, Fabian R. A. Biebl, Paul D. Ledger, and Malcolm Brown. Accelerating magnetic induction tomography-based imaging through heterogeneous parallel computing. Concurrency and Computation: Practice and Experience, 31(17):e5265:1–e5265:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Woo:2011:CUC

[WKL^{+}11] Jung-Hun Woo, HyungSeok Kim, Sang Boem Lim, Jae-Jin Kim, Jonghyun Lee, Rina Ryoo, Hansoo Kim, and Le Dinh Minh. Constructing u-City of Seoul by future foresight analysis. Concurrency and Computation: Practice and Experience, 23(10):1114–1126, July 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weeks:2014:ICL

[WKL14] Nathan T. Weeks, Marina Kraeva, and Glenn R. Luecke. ipcmd: a command-line interface to System V semaphores and message queues. Concurrency and Computation: Practice and Experience, 26(2):396–411, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2008:PEV

[WKT08] Lizhe Wang, Marcel Kunze, and Jie Tao. Performance evaluation of virtual machine-based Grid workflow system. Concurrency and Computation: Practice and Experience, 20(15):

1759–1771, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:ACP

[WKZL19] Pengfei Wang, Jens Krinke, Xu Zhou, and Kai Lu. AVPredictor: Comprehensive prediction and detection of atomicity violations. Concurrency and Computation: Practice and Experience, 31(15):e5160:1–e5160:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weissman:2002:VSG

[WL02] Jon B. Weissman and Byoung-Dai Lee. The Virtual Service Grid: an architecture for delivering high-end network services. Concurrency and Computation: Practice and Experience, 14(4):287–319, April 10, 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/93513490/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=93513490{\&}PLACEBO=
IE.pdf.

Wang:2011:DDA

[WL11a] Yang Wang and Paul Lu. Dataflow detection and applications to workflow scheduling. Concurrency and Computation: Practice and Experience, 23(11):1261–1283, August 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wittie:2011:TSC

[WL11b] Lea Wittie and Jonathan Lockhart. Type-safe concurrent resource sharing. Concurrency and Computation: Practice and Experience, 23(8):767–795, June 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2012:MCE

[WL12] Jinwei Wang and Shiguo Lian. On multiwatermarking in cloud environment. Concurrency and Computation: Practice and Experience, 24(17):2151–2164, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2008:SKF

[WLDL08] Liping Wang, Qing Li, Guozhu Dong, and Yu Li. Semantic knowledge facilities for a web-based recipe database system supporting personalization. Concurrency and Computation: Practice and Experience, 20(7):753–782, May 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:ACA

[WLF19] Junchang Wang, Tao Li, and Xiong Fu. Accurate counting algorithm for high-speed parallel applications. Concurrency and Computation: Practice and Experience, 31(13):e5090:1– e5090:??, July 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:CAS

[WLFX17] Changji Wang, Yuan Li, Jian Fang, and Jianguo Xie. Cloudaided scalable revocable identity-based encryption scheme with ciphertext update. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wong:2003:APG

[WLL03a] Yuk-Yin Wong, Kin-Hong Lee, and Kwong-Sak Leung. An adaptive parallel genetic algorithm system for i-Computing environment. Concurrency and Computation: Practice and Experience, 15(6):581–606, May 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wong:2003:SLB

[WLL03b] Yuk-Yin Wong, Kwong-Sak Leung, and Kin-Hong Lee. A stochastic load balancing algorithm for i-Computing. Concurrency and Computation: Practice and Experience, 15(1): 55–78, January 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2014:RST

[WLL14] Chung-Ju Wu, Chia-Han Lu, and Jenq Kuen Lee. Register spilling via transformed interference equations for PAC DSP architecture. Concurrency and Computation: Practice and Experience, 26(3):779–799, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wan:2015:GIP

[WLLL15] Lanjun Wan, Kenli Li, Jing Liu, and Keqin Li. GPU implementation of a parallel two-list algorithm for the subset-sum problem. Concurrency and Computation: Practice and Experience, 27(1):119–145, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wan:2016:ECG

[WLLL16] Lanjun Wan, Kenli Li, Jing Liu, and Keqin Li. Efficient CPU–GPU cooperative computing for solving the subset-sum problem. Concurrency and Computation: Practice and Experience, 28(2):492–516, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:SDF

[WLLZ18] Pengfei Wang, Kai Lu, Gen Li, and Xu Zhou. A survey of the double-fetch vulnerabilities. Concurrency and Computation: Practice and Experience, 30(6):??, March 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4345.

Wang:2017:VNE

[WLP^{+}17] Cong Wang, Guohua Liu, Sancheng Peng, Ying Yuan, Guorui Li, and Cong Wan. Virtual network embedding with pretransformation and incentive convergence mechanism. Concurrency and Computation: Practice and Experience, 29(14): ??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2016:TCC

[WLQL16] Xinyang Wang, Jiarong Liang, Deyu Qi, and Weiwei Lin. The twisted crossed cube. Concurrency and Computation: Practice and Experience, 28(5):1507–1526, April 10, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Willcock:2005:UMC

[WLR05] Jeremiah Willcock, Andrew Lumsdaine, and Arch Robison. Using MPI with C# and the Common Language Infrastructure. Concurrency and Computation: Practice and Experience, 17(7–8):895–917, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2011:AFG

[WLW11] Guojun Wang, Qin Liu, and Jie Wu. Achieving fine-grained access control for secure data sharing on cloud servers. Concurrency and Computation: Practice and Experience, 23(12): 1443–1464, August 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2014:GMP

[WLW14] Jingbo Wei, Dingsheng Liu, and Lizhe Wang. A general metric and parallel framework for adaptive image fusion in clusters. Concurrency and Computation: Practice and Experience, 26(7):1375–1387, May 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2014:SRS

[WLWX14] Yilei Wang, Zhe Liu, Hao Wang, and Qiuliang Xu. Social rational secure multi-party computation. Concurrency and Computation: Practice and Experience, 26(5):1067–1083, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See errata [WLWX16].

Wang:2016:ESR

[WLWX16] Yilei Wang, Zhe Liu, Hao Wang, and Qiuliang Xu. Errata: Social rational secure multi-party computation. Concurrency and Computation: Practice and Experience, 28(9):2748, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [WLWX14].

Wahid:2011:SSC

[WLZ11] Alif Wahid, Christopher Leckie, and Chenfeng Zhou. Selfsimilar characteristics of network intrusion attempts and the implications for predictability. Concurrency and Computation: Practice and Experience, 23(12):1367–1385, August 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:KPD

[WLZ17] Jiaze Wang, Anfeng Liu, and Shigeng Zhang. Key parameters decision for cloud computing: Insights from a multiple game model. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:IGO

[WLZ^{+}18] Xu An Wang, Yudong Liu, Jindan Zhang, Xiaoyuan Yang, and Minqing Zhang. Improved group-oriented proofs of cloud storage in IoT setting. Concurrency and Computation: Practice and Experience, 30(21):e4781:1–e4781:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2007:MEC

[WMA07] Lan Wang, Geyong Min, and Irfan Awan. Modeling and evaluation of congestion control for different classes of network traffic. Concurrency and Computation: Practice and Experience, 19(8):1141–1156, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:PPD

[WMC17] Xiaofen Wang, Yi Mu, and Rongmao Chen. Privacypreserving data search and sharing protocol for social networks through wireless applications. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wolf:2007:AAI

[WMDM07] Felix Wolf, Bernd Mohr, Jack Dongarra, and Shirley Moore. Automatic analysis of inefficiency patterns in parallel applications. Concurrency and Computation: Practice and Experience, 19(11):1481–1496, August 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:AML

[WML^{+}19] Yu Wang, Weizhi Meng, Wenjuan Li, Zhe Liu, Yang Liu, and Hanxiao Xue. Adaptive machine learning-based alarm reduction via edge computing for distributed intrusion detection systems. Concurrency and Computation: Practice and Experience, 31(19):e5101:1–e5101:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2009:CRT

[WMvP^{+}09] Peng Wu, Maged M. Michael, Christoph von Praun, Takuya Nakaike, Rajesh Bordawekar, Harold W. Cain, Calin Cascaval, Siddhartha Chatterjee, Stefanie Chiras, Rui Hou, Mark Mergen, Xiaowei Shen, Michael F. Spear, Hua Yong Wang, and Kun Wang. Compiler and runtime techniques for software transactional memory optimization. Concurrency and Computation: Practice and Experience, 21(1):7–23, January 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2015:BMR

[WNN^{+}15] Jie Wu, Krishnaprasad Narayanan, Lars Nagel, Christoph Fiehe, Anna Litvina, Jakob Tonn, Carsten Zoth, Hans-Joachim Goltz, Steffen Unger, Fabian Pursche, Michael Scheel, Andre Brinkmann, and Wolfgang Thronicke. Building a medical research cloud in the EASI-CLOUDS project. Concurrency and Computation: Practice and Experience, 27 (16):4465–4477, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weber:2002:FFC

[WNT02] S. Weber, P. A. Nixon, and B. Tangney. A flexible framework for consistency management. Concurrency and Computation: Practice and Experience, 14(1):33–53, January 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/91014113/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
91014113&PLACEBO=IE.pdf.

Wada:2002:EAT

[WO02] Yoshitaka Wada and Hiroshi Okuda. Effective adaptation technique for hexahedral mesh. Concurrency and Computation: Practice and Experience, 14(6–7):451–463, May/ June 2002. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/94515739/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
94515739{\&}PLACEBO=IE.pdf.

Williamson:2014:PPP

[WO14] Todd Williamson and Ronald A. Olsson. PySy: a Python package for enhanced concurrent programming. Concurrency and Computation: Practice and Experience, 26(2):309–335, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wolstencroft:2013:SAE

[WOH^{+}13] Katherine Wolstencroft, Stuart Owen, Matthew Horridge, Simon Jupp, Olga Krebs, Jacky Snoep, Franco du Preez, Wolfgang Mueller, Robert Stevens, and Carole Goble. Stealthy annotation of experimental biology by spreadsheets. Concurrency and Computation: Practice and Experience, 25(4): 467–480, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Withana:2012:SUR

[WP12] Eran Chinthaka Withana and Beth Plale. Sigiri: uniform resource abstraction for grids and clouds. Concurrency and Computation: Practice and Experience, 24(18):2362–2380, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:GFB

[WQL^{+}18] Xinyang Wang, Deyu Qi, Weiwei Lin, MinCong Yu, Zhishuo Zheng, Naqin Zhou, and Pengguang Chen. A general framework for big data knowledge discovery and integration. Concurrency and Computation: Practice and Experience, 30(13):??, July 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4422.

Wang:2016:LBL

[WQS^{+}16] Ke Wang, Kan Qiao, Iman Sadooghi, Xiaobing Zhou, Tonglin Li, Michael Lang, and Ioan Raicu. Load-balanced and locality-aware scheduling for data-intensive workloads at extreme scales. Concurrency and Computation: Practice and Experience, 28(1):70–94, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:EFS

[WR17] Xiaofei Wang and Jianji Ren. Editorial: Foreword to the special issue on networked system security and efficiency. Concurrency and Computation: Practice and Experience, 29(16): ??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2009:TBR

[WRC09] Mingzhong Wang, Kotagiri Ramamohanarao, and Jinjun Chen. Trust-based robust scheduling and runtime adaptation of scientific workflow. Concurrency and Computation: Practice and Experience, 21(16):1982–1998, November 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2013:DBD

[WRDZ13] Guangen Wu, Pinyi Ren, Qinghe Du, and Chao Zhang. A DOF-based dynamic spectrum auction algorithm in cognitive femtocell. Concurrency and Computation: Practice and Experience, 25(9):1126–1143, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2012:CLB

[WRLS12] Yichen Wang, Pinyi Ren, Fan Li, and Zhou Su. Cross-layer based power allocation over cognitive wireless relay link with statistical delay QoS guarantees. Concurrency and Computation: Practice and Experience, 24(11):1239–1251, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Waliullah:2009:SAS

[WS09] M. M. Waliullah and Per Stenstrom. Schemes for avoiding starvation in transactional memory systems. Concurrency and Computation: Practice and Experience, 21(7):859–873, May 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wyrzykowski:2017:EAA

[WS17] Roman Wyrzykowski and Boleslaw K. Szymanski. Editorials: Algorithmic advances for parallel architectures. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Woodworth:2019:SSS

[WS19a] Jason W. Woodworth and Mohsen Amini Salehi. S3BD: Secure semantic search over encrypted big data in the cloud. Concurrency and Computation: Practice and Experience, 31 (11):e5050:1–e5050:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wyrzykowski:2019:AAP

[WS19b] Roman Wyrzykowski and Boleslaw K. Szymanski. Algorithmic advances in parallel architectures and energy-efficient computing. Concurrency and Computation: Practice and Experience, 31(19):e5260:1–e5260:??, October 10, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2015:BIC

[WSL15] Lijuan Wang, Jun Shen, and Junzhou Luo. Bio-inspired costaware optimization for data-intensive service provision. Concurrency and Computation: Practice and Experience, 27(18): 5662–5685, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Williams:2017:VLS

[WSP17] Nadya Williams, Aimee Stewart, and Phil Papadopoulos. Virtualizing Lifemapper software infrastructure for biodiversity expedition. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weston:2012:RCV

[WSRM12] Stephen Weston, James Spooner, Sébastien Racanière, and Oskar Mencer. Rapid computation of value and risk for derivatives portfolios. Concurrency and Computation: Practice and Experience, 24(8):880–894, June 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wong:2017:IME

[WST^{+}17] Lok Wong, Shinji Shimojo, Yuuichi Teranishi, Tomoki Yoshihisa, and Jason H. Haga. Interactive museum exhibits with embedded systems: a use-case scenario. Concurrency and Computation: Practice and Experience, 29(13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2012:QAR

[WSW^{+}12] Jan-Jan Wu, Shu-Fan Shih, Hsiangkai Wang, Pangfeng Liu, and Chien-Min Wang. QoS-aware replica placement for grid computing. Concurrency and Computation: Practice and Experience, 24(3):193–213, March 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2012:JSC

[WSWL12] Xiaoxin Wu, Zhiming Shen, Ryan Wu, and Yunfeng Lin. Jump-start cloud: efficient deployment framework for largescale cloud applications. Concurrency and Computation: Practice and Experience, 24(17):2120–2137, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:EAC

[WSZ^{+}18] Xinggang Wang, Buyun Sheng, Chenglei Zhang, Zheng Xiao, Hui Wang, and Feiyu Zhao. An effective application of 3D cloud printing service quality evaluation in BM-MOPSO. Concurrency and Computation: Practice and Experience, 30(24):e4977:1–e4977:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wheeler:2010:VMM

[WT10] Kyle B. Wheeler and Douglas Thain. Visualizing massively multithreaded applications with ThreadScope. Concurrency and Computation: Practice and Experience, 22(1):45–67, January 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wyrzykowski:2015:EIC

[WT15] Roman Wyrzykowski and Marek Tudruj. Editorials: 10th international conference on Parallel Processing and Applied Mathematics, PPAM 2013. Concurrency and Computation: Practice and Experience, 27(4):882–884, March 25, 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2018:HSS

[WT18] Hsin-Te Wu and Chun-Wei Tsai. A home security system for seniors based on the beacon technology. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4496.

Walid:2017:TSM

[WTEG17] Elgenaidi Walid, Newe Thomas, O’Connell Eoin, and Dooly Gerard. Trust security mechanism for maritime wireless sensor networks. Concurrency and Computation: Practice and Experience, 29(23):??, December 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2016:IBP

[WTL^{+}16] Song Wu, Songqiao Tao, Xiao Ling, Hao Fan, Hai Jin, and Shadi Ibrahim. iShare: Balancing I/O performance isolation and disk I/O efficiency in virtualized environments. Concurrency and Computation: Practice and Experience, 28(2): 386–399, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2007:ASO

[WTN07] Di Wu, Ye Tian, and Kam Wing Ng. An analytical study on optimizing the lookup performance of distributed hash table systems under churn. Concurrency and Computation: Practice and Experience, 19(4):543–569, March 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:EMA

[WTY^{+}19] Yi-Lei Wang, Wen-Zhe Tang, Xian-Jun Yang, Ying-Jie Wu, and Fu-Ji Chen. An efficient method for autoencoder-based collaborative filtering. Concurrency and Computation: Practice and Experience, 31(23):e4507:1–e4507:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2018:ICM

[Wu18] Shaofei Wu. The 2018 International Conference on Modern Computer Science and Applications (MCSA2018). Concurrency and Computation: Practice and Experience, 30(24): e4996:1–e4996:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wood:2008:SIL

[WW08] J. D. Wood and H. Wright. Steering via the image in local, distributed and collaborative settings. Concurrency and Computation: Practice and Experience, 20(3):265–276, March 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:NSA

[WW19] Yan Wang and Mengnan Wu. A novel systematic algorithm paradigm for the electric vehicle data anomaly detection based on association data mining. Concurrency and Computation: Practice and Experience, 31(9):e5073:1–e5073:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:NVB

[WWC^{+}19] Liangliang Wang, Mi Wen, Kefei Chen, Zhongqin Bi, and Yu Long. A new VRSA-based pairing-free certificateless signature scheme for fog computing. Concurrency and Computation: Practice and Experience, 31(22):e4919:1–e4919:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wilke:2011:ERP

[WWG^{+}11] Andreas Wilke, Jared Wilkening, Elizabeth M. Glass, Narayan L. Desai, and Folker Meyer. An experience report: porting the MG-RAST rapid metagenomics analysis pipeline to the cloud. Concurrency and Computation: Practice and Experience, 23(17):2250–2257, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2015:GWA

[WWL^{+}15] Liya Wang, Doreen Ware, Carol Lushbough, Nirav Merchant, and Lincoln Stein. A genome-wide association study platform built on iPlant cyber-infrastructure. Concurrency and Computation: Practice and Experience, 27(2):420–432, February 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:MSD

[WWL^{+}17a] Hui Wang, Xiaofei Wang, Keqiu Li, Jianji Ren, Xiaohong Zhang, and Tianpeng Jiang. A measurement study of deviceto-device sharing in mobile social networks based on Spark. Concurrency and Computation: Practice and Experience, 29 (16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2017:DSA

[WWL^{+}17b] Tingmin Wu, Sheng Wen, Shigang Liu, Jun Zhang, Yang Xiang, Majed Alrubaian, and Mohammad Mehedi Hassan. Detecting spamming activities in Twitter based on deep-learning technique. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:ECG

[WWLD18] Yueqing Wang, Fang Wang, Rongchun Li, and Yong Dou. An efficient CPU–GPU hybrid parallel implementation for DVB– RCS2 receiver. Concurrency and Computation: Practice and Experience, 30(19):e4529:1–e4529:??, October 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wagner:2012:BTA

[WWS^{+}12] Cynthia Wagner, Gerard Wagener, Radu State, Alexandre Dulaunoy, and Thomas Engel. Breaking Tor anonymity with game theory and data mining. Concurrency and Computation: Practice and Experience, 24(10):1052–1065, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:EMS

[WWX^{+}19] Yeqin Wang, Zhen Wu, Aoyun Xia, Chang Guo, Yuyan Chen, Yan Yang, and Zhongyi Tang. Energy management strategy for HEV based on KFCM and neural network. Concurrency and Computation: Practice and Experience, 31(10):e4838:1– e4838:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:EHS

[WWY19] Kun Wang, Wei Wang, and Xiong Yan. Effect of hole on sound absorption coefficient of micro-perforated panels investigated by Melling computing. Concurrency and Computation: Practice and Experience, 31(12):e4725:1–e4725:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2019:PFB

[WXLM19] Yan Wu, Wei Xu, Lu Liu, and Dejun Miao. Performance formula-based optimal deployments of multilevel indices for service retrieval. Concurrency and Computation: Practice and Experience, 31(3):e4265:1–e4265:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2019:NTB

[WXSH19] Chen Wang, Lu Xiao, Jian Shen, and Rui Huang. Neighborhood trustworthiness-based vehicle-to-vehicle authentication scheme for vehicular ad hoc networks. Concurrency and Computation: Practice and Experience, 31(21):e4643:1–e4643:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2010:AAS

[WXY10] Yu Wang, Yang Xiang, and Shun-Zheng Yu. An automatic application signature construction system for unknown traffic. Concurrency and Computation: Practice and Experience, 22(13):1927–1944, September 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2007:PPR

[WYAB07] X. D. Wang, X. Yang, R. J. Allan, and M. Baker. Plug-and-play remote portlet publishing. Concurrency and Computation: Practice and Experience, 19(12):1693–1702, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2016:POD

[WYBS16] Hua Wang, Xun Yi, Elisa Bertino, and Lili Sun. Protecting outsourced data in cloud computing through access management. Concurrency and Computation: Practice and Experience, 28(3):600–615, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2014:IDC

[WYL14] Yuechuan Wei, Xiaoyuan Yang, and Chao Li. Impossible differential cryptanalysis on cipher E2. Concurrency and Computation: Practice and Experience, 26(8):1477–1489, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2013:BDS

[WYQ^{+}13] Lei Wang, Zhuxiu Yuan, Zhenquan Qin, Yuanfang Chen, Lei Shu, and Xiang-Yang Li. A backoff differentiation scheme for contention resolution in wireless converge-cast networks. Concurrency and Computation: Practice and Experience, 25 (1):112–128, January 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:RCD

[WYW^{+}17] Zhou-Feng Wang, Yang Yang, Yu-Jun Wang, Ting-Shan Zhang, and Cheng-Wu Wang. Research in the calculation of the debris flow prevention and Geographic Information System control engineering. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2019:MBM

[WYY^{+}19] Chong Wu, Yanlei Yin, Xuemei Yang, Lijuan Feng, Haixia Tang, and Jihan Tao. A Markov-based model for predicting the development trend of soil microbial communities in salinealkali land in Wudi County. Concurrency and Computation: Practice and Experience, 31(10):e4754:1–e4754:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wan:2012:EBE

[WYZ12] Shuai Wan, Kaifang Yang, and Haiyong Zhou. Efficient bitstream extraction for scalable video based on simulated annealing. Concurrency and Computation: Practice and Experience, 24(11):1223–1230, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:DBS

[WYZ^{+}17] Qian Wang, Chao Yu, Yiming Zhang, Huiba Li, and Ping Zhong. Delay-bounded skyline computing for large-scale realtime online data analytics. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2004:PLF

[WZ04] Xiaofang Wang and Sotirios G. Ziavras. Parallel LU factorization of sparse matrices on FPGA-based configurable computing engines. Concurrency and Computation: Practice and Experience, 16(4):319–343, April 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2016:EES

[WZ16] Xiao Wei and Daniel Dajun Zeng. ExNa: an efficient search pattern for semantic search engines. Concurrency and Computation: Practice and Experience, 28(15):4107–4124, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2016:SAP

[WZC16] Minqian Wang, Zhenfeng Zhang, and Cheng Chen. Security analysis of a privacy-preserving decentralized ciphertextpolicy attribute-based encryption scheme. Concurrency and Computation: Practice and Experience, 28(4):1237–1245, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2013:FDA

[WZJD13] Liejun Wang, Taiyi Zhang, Zhenhong Jia, and Liang Ding. Face detection algorithm based on hybrid Monte Carlo method and Bayesian support vector machine. Concurrency and Computation: Practice and Experience, 25(9):1064–1072, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:NCM

[WZL^{+}17a] Jinkun Wang, Shasha Zhang, Xiao Liu, Yuanchun Jiang, and Min Zhang. A novel collective matrix factorization model for recommendation with fine-grained social trust prediction. Concurrency and Computation: Practice and Experience, 29 (19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2017:DDI

[WZL^{+}17b] Jinxia Wei, Ru Zhang, Jianyi Liu, Jing Li, Xinxin Niu, and Yuangang Yao. Dynamic data integrity auditing for secure outsourcing in the cloud. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2018:SPS

[WZLL18] Zhoukai Wang, Yinliang Zhao, Yang Liu, and Cuocuo Lv. A speculative parallel simulated annealing algorithm based on Apache Spark. Concurrency and Computation: Practice and Experience, 30(14):??, July 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4429.

Wei:2016:FFR

[WZLQ16] Wei Wei, Yuhong Zhang, Yang Liu, and Zhiguang Qin. FRP: a fast resource placement algorithm in distributed cloud computing platform. Concurrency and Computation: Practice and Experience, 28(5):1399–1416, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2015:MMS

[WZS^{+}15] Jin Wang, Liwu Zuo, Jian Shen, Bin Li, and Sungyoung Lee. Multiple mobile sink-based routing algorithm for data dissemination in wireless sensor networks. Concurrency and Computation: Practice and Experience, 27(10):2656–2667, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2011:BIS

[WZT11] Wei Wang, Guosun Zeng, and Daizhong Tang. Bayesian intelligent semantic mashup for tourism. Concurrency and Computation: Practice and Experience, 23(8):850–862, June 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wen:2012:CNL

[WZXZ12] Sheng Wen, Wei Zhou, Yang Xiang, and Wanlei Zhou. CAFS: a novel lightweight cache-based scheme for large-scale intrusion alert fusion. Concurrency and Computation: Practice and Experience, 24(10):1137–1153, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2019:IPS

[WZYG19] Qinghua Wu, Zhixin Zhu, Xuesong Yan, and Wenyin Gong. An improved particle swarm optimization algorithm for AVO elastic parameter inversion problem. Concurrency and Computation: Practice and Experience, 31(9):e4987:1–e4987:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2013:SIP

[WZZL13] Mingzhong Wang, Xuyun Zhang, Liehuang Zhu, and Lejian Liao. Special issue papers: Trust-based workflow refactoring for concurrent scheduling in service-oriented environment. Concurrency and Computation: Practice and Experience, 25 (13):1879–1893, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2015:EFT

[XADLC15] Yang Xiang, Ahmed Al-Dubi, Lei Liu, and Xiaowen Chu. Editorial: Frontier technologies of trust computing and network security. Concurrency and Computation: Practice and Experience, 27(12):2907–2909, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2016:ESP

[XAK16] Yang Xiang, Man Ho Au, and Miroslaw Kutylowsky. Editorials: Security and privacy in big data. Concurrency and Computation: Practice and Experience, 28(10):2856–2857, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xing:2013:CCS

[XBB13] Qi Xing and Estela Blaisten-Barojas. A cloud computing system in Windows Azure platform for data analysis of crystalline materials. Concurrency and Computation: Practice and Experience, 25(15):2157–2169, October 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2019:FSI

[XBCW19] Yang Xiang, Md Zakirul Alam Bhuiyan, Aniello Castiglione, and Yu Wang. Foreword to the special issue on security, privacy, and social networks. Concurrency and Computation: Practice and Experience, 31(21):e5278:1–e5278:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2017:ESP

[XBK17] Yang Xiang, Elisa Bertino, and Miroslaw Kutylowski. Editorials: Security and privacy in social networks. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2014:ESN

[XBM14] Li Xu, Elisa Bertino, and Yi Mu. Editorial: Security of new generation computing systems. Concurrency and Computation: Practice and Experience, 26(8):1475–1476, June 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2015:SLT

[XBW^{+}15] Mande Xie, Urmila Bhanja, Guiyi Wei, Yun Ling, Mohammad Mehedi Hassan, and Atif Alamri. SecNRCC: a losstolerant secure network reprogramming with confidentiality consideration for wireless sensor networks. Concurrency and Computation: Practice and Experience, 27(10):2668–2680, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2013:SIP

[XBXS13] Qingjun Xiao, Kai Bu, Bin Xiao, and Limin Sun. Special issue papers: Efficient protocol design for dynamic tag population monitoring in large-scale radio frequency identification systems. Concurrency and Computation: Practice and Experience, 25(14):2080–2097, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2010:FAD

[XBZ10] Hengheng Xie, Azzedine Boukerche, and Ming Zhang. A formalized approach for designing a P2P-based dynamic load balancing scheme. Concurrency and Computation: Practice and Experience, 22(10):1223–1239, July 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xhafa:2014:EAS

[XCHK14] Fatos Xhafa, Xiaofeng Chen, Xinyi Huang, and Vladi Kolici. Editorial: Advances in secure and intelligent data processing. Concurrency and Computation: Practice and Experience, 26 (5):1021–1022, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2013:EAP

[XCHY13] Yang Xiang, Alfredo Cuzzocrea, Michael Hobbs, and Laurence T. Yang. Editorial: Advances in parallel, distributed, embedded, and ubiquitous systems. Concurrency and Computation: Practice and Experience, 25(14):1985– 1986, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xavier:2009:DAC

[XCL09] Percival Xavier, Wentong Cai, and Bu-Sung Lee. A dynamic admission control scheme to manage contention on shared computing resources. Concurrency and Computation: Practice and Experience, 21(2):133–158, February 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2004:HSO

[XDE^{+}04] Xiang Xiao, Ernst R. Dow, Russell Eberhart, Zina Ben Miled, and Robert J. Oppelt. A hybrid self-organizing maps and particle swarm optimization approach. Concurrency and Computation: Practice and Experience, 16(9):895–915, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiong:2018:HHA

[XDJL18] Runqun Xiong, Yao Du, Jiahui Jin, and Junzhou Luo. HaDaap: a hotness-aware data placement strategy for improving storage efficiency in heterogeneous Hadoop clusters. Concurrency and Computation: Practice and Experience, 30 (20):e4830:1–e4830:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2011:MDA

[XDL^{+}11] Yunni Xia, Gang Dai, Jia Li, Tianhao Sun, and Qingsheng Zhu. A model-driven approach to predicting dependability of WS-CDL based service composition. Concurrency and Computation: Practice and Experience, 23(10):1127–1145, July 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2018:HIC

[XDP18] Pu Xu, Zhijun Ding, and MeiQin Pan. A hybrid interpretable credit card users default prediction model based on RIP-PER. Concurrency and Computation: Practice and Experience, 30(23):e4445:1–e4445:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2019:NBR

[XGC19] Junxu Xia, Deke Guo, and Geyao Cheng. In-network block repairing for erasure coding storage systems. Concurrency and Computation: Practice and Experience, 31(24):e5432:1– e5432:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2015:EEB

[XGXH15] Jinhai Xu, Songtao Guo, Bin Xiao, and Jing He. Energyefficient big data storage and retrieval for wireless sensor networks with nonuniform node distribution. Concurrency and Computation: Practice and Experience, 27(18):5765–5779, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xhafa:2018:ESI

[Xha18] Fatos Xhafa. Editorial: Special issue on advanced techniques for cloud data storage and collaborative systems. Concurrency and Computation: Practice and Experience, 30(1): ??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xing:2015:OIB

[XHCL15] Bin Xing, Zhen Han, Xiaolin Chang, and Jiqiang Liu. OB–IMA: out-of-the-box integrity measurement approach for guest virtual machines. Concurrency and Computation: Practice and Experience, 27(5):1092–1109, April 10, 2015.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2019:RTT

[XHD^{+}19] Fusheng Xu, Zhongxiang Huang, Hao Dai, Xueying Zhu, Hongwei Wu, and Jun Zhang. A real-time traffic index model for expressways. Concurrency and Computation: Practice and Experience, 31(24):e5155:1–e5155:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xi:2012:MDA

[XHH12] Kai Xi, Jiankun Hu, and Fengling Han. Mobile device access control: an improved correlation based face authentication scheme and its Java ME application. Concurrency and Computation: Practice and Experience, 24(10):1066–1085, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2019:MII

[XHW^{+}19] Chuan Xu, Zhenzhen Han, Qianyun Wang, Guofeng Zhao, and Shui Yu. Modelling the impact of interference on the energy efficiency of WLANs. Concurrency and Computation: Practice and Experience, 31(17):e5217:1–e5217:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2012:ANS

[XHZ12] Yang Xiang, Jiankun Hu, and Wanlei Zhou. Advances in network and system security. Concurrency and Computation: Practice and Experience, 24(10):1035–1036, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2018:IPT

[XJAJ18] Yang Xiao, Thireshan Jeyakumaran, Ehsan Atoofian, and Ali Jannesari. Improving performance of transactional memory through machine learning. Concurrency and Computation: Practice and Experience, 30(10):??, May 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4397.

Xu:2013:SIP

[XJZ13] Li Xu, Lei Ji, and Shu Ming Zhou. Special issue papers: An efficient self-diagnosis protocol for hierarchical wireless mesh networks. Concurrency and Computation: Practice and Experience, 25(14):2036–2051, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2017:PPL

[XL17] Thomas Canhao Xu and Ville Leppänen. PEN: a power law– enhanced network design for high efficiency multicore architecture. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2019:NIF

[XLG19] Wenhao Xie, Gongqian Liang, and Qiao Guo. A new improved FSVM algorithm based on SVDD. Concurrency and Computation: Practice and Experience, 31(9):e4893:1– e4893:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xing:2017:CMO

[XLHT17] Lining Xing, Wen Li, Minfan He, and Xu Tan. Comprehensive multi-objective model to remote sensing data processing task scheduling problem. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xue:2012:DRD

[XLL^{+}12] Weilian Xue, Yaqiong Liu, Keqiu Li, Zhongxian Chi, Geyong Min, and Wenyu Qu. DHTrust: a robust and distributed reputation system for trusted peer-to-peer networks. Concurrency and Computation: Practice and Experience, 24(10): 1037–1051, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2015:PPD

[XLL^{+}15] Thomas Canhao Xu, Ville Leppänen, Pasi Liljeberg, Juha Plosila, and Hannu Tenhunen. PDNOC: Partially diagonal network-on-chip for high efficiency multicore systems. Concurrency and Computation: Practice and Experience, 27(4):

1054–1067, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xue:2018:FTF

[XLL^{+}18] Zhengyuan Xue, Ruixuan Li, Yuhua Li, Lin Huo, Xiwu Gu, and Weijun Xiao. FRFB: Top-k Followee Recommendation by exploring the Following Behaviors in social networks. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4514.

Xu:2014:MSD

[XLMH14] Zheng Xu, Xiangfeng Luo, Lin Mei, and Chuanping Hu. Measuring the semantic discrimination capability of association relations. Concurrency and Computation: Practice and Experience, 26(2):380–395, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2018:EVC

[XLQL18] Xinping Xu, Wenxin Li, Heng Qi, and Keqiu Li. On efficient virtual cluster scaling across geo-distributed data centers. Concurrency and Computation: Practice and Experience, 30(10):??, May 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4383.

Xiao:2017:SAM

[XLT^{+}17] Zheng Xiao, Pijun Liang, Zhao Tong, Kenli Li, Samee U. Khan, and Keqin Li. Self-adaptation and mutual adaptation for distributed scheduling in benevolent clouds. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2011:SIS

[XLWZ11] Yang Xiang, Javier Lopez, Haining Wang, and Wanlei Zhou. Special issue: Securing distributed networks and systems. Concurrency and Computation: Practice and Experience, 23 (12):1311–1312, August 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2016:TIE

[XLY^{+}16] Xiao Xu, Ge Li, Mei Yang, Rusheng Ju, and Kedi Huang. Towards the integration of engagement and engineering-level simulation over real-time and heterogeneous systems. Concurrency and Computation: Practice and Experience, 28(12): 3390–3408, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2017:STD

[XLYL17] Guoqi Xie, Liangjiao Liu, Liu Yang, and Renfa Li. Scheduling trade-off of dynamic multiple parallel workflows on heterogeneous distributed computing systems. Concurrency and Computation: Practice and Experience, 29(2):??, January 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2011:MSS

[XLYX11a] Zheng Xu, Xiangfeng Luo, Jie Yu, and Weimin Xu. Measuring semantic similarity between words by removing noise and redundancy in Web snippets. Concurrency and Computation: Practice and Experience, 23(18):2496–2510, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2011:MWS

[XLYX11b] Zheng Xu, Xiangfeng Luo, Jie Yu, and Weimin Xu. Mining Web search engines for query suggestion. Concurrency and Computation: Practice and Experience, 23(10):1101–1113, July 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2013:ISI

[XLZD13] Jiajie Xu, Chengfei Liu, Xiaohui Zhao, and Zhiming Ding. Incorporating structural improvement into resource allocation for business process execution planning. Concurrency and Computation: Practice and Experience, 25(3):427–442, March 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xing:2002:FEM

[XM02] H. L. Xing and A. Makinouchi. Finite-element modeling of multibody contact and its application to active faults. Concurrency and Computation: Practice and Experience, 14(6– 7):431–450, May/June 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/94515748/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=94515748{\&}PLACEBO=IE.pdf.

Xu:2017:CCG

[XMJ17] Ying Xu, Huimin Ma, and Rongling Jiang. Collaborating CPU and GPU for the electromagnetic simulations with the FDTD algorithm. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xhafa:2011:GPM

[XPBS11] Fatos Xhafa, Sabri Pllana, Leonard Barolli, and Evjola Spaho. Grid and P2P middleware for wide-area parallel processing. Concurrency and Computation: Practice and Experience, 23(5):458–476, April 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xhafa:2015:EIG

[XPS^{+}15] Fatos Xhafa, Alina-Diana Potlog, Evjola Spaho, Florin Pop, Valentin Cristea, and Leonard Barolli. Evaluation of intragroup optimistic data replication in P2P groupware systems. Concurrency and Computation: Practice and Experience, 27 (4):870–881, March 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2015:EAR

[XPWF15] Yang Xiang, Mukaddim Pathan, Guiyi Wei, and Giancarlo Fortino. Editorial: Availability, resilience, and fault tolerance of Internet and distributed computing systems. Concurrency and Computation: Practice and Experience, 27(10): 2503–2505, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xavier:2017:MSG

[XRD^{+}17] Miguel G. Xavier, Fábio D. Rossi, César A. F. De Rose, Rodrigo N. Calheiros, and Danielo G. Gomes. Modeling and simulation of global and sleep states in ACPI-compliant energyefficient cloud environments. Concurrency and Computation:

Practice and Experience, 29(4):??, February 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2019:RAC

[XS19] Yong Xie and Huiming Su. Research on the application of cloud computing mode in athletes’ technical movement capture. Concurrency and Computation: Practice and Experience, 31(10):e4747:1–e4747:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2016:ESR

[XSMZ16] Yang Xiang, Ivan Stojmenovic, Peter Mueller, and Jun Zhang. Editorials: Security and reliability in big data. Concurrency and Computation: Practice and Experience, 28(3): 581–582, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2017:SLB

[XTB17] Minxian Xu, Wenhong Tian, and Rajkumar Buyya. A survey on load balancing algorithms for virtual machines placement in cloud computing. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2008:FGS

[XTLG08] Jie Xu, Paul Townend, Nik Looker, and Paul Groth. FT-Grid: a system for achieving fault tolerance in grids. Concurrency and Computation: Practice and Experience, 20(3):297–309, March 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2010:MCM

[XTZ10] Yang Xiang, Daxin Tian, and Wanlei Zhou. A microscopic competition model and its dynamics analysis on network attacks. Concurrency and Computation: Practice and Experience, 22(4):503–517, March 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2008:SIS

[Xu08] Jie Xu. Special issue: Selected papers from the U.K. e-Science All Hands Meeting in 2006. Concurrency and Computation:

Practice and Experience, 20(11):1257–1258, August 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2019:AMS

[Xu19] Feihong Xu. Accurate measurement of structural vibration based on digital image processing technology. Concurrency and Computation: Practice and Experience, 31(10):e4767:1– e4767:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2013:SIP

[XW13] Yongming Xie and Guojun Wang. Special issue papers: Practical distributed secret key generation for delay tolerant networks. Concurrency and Computation: Practice and Experience, 25(14):2067–2079, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2012:NMS

[XWD^{+}12] Yunni Xia, Neng Wan, Gang Dai, Xin Luo, and Tianhao Sun. A non-Markovian stochastic Petri net-based approach to performance evaluation of ontology-based service composition. Concurrency and Computation: Practice and Experience, 24(18):2255–2267, December 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2008:QMA

[XWFH08] Yunni Xia, Hanpin Wang, Wangsen Feng, and Yu Huang. QoS modeling and analysis of component-based software systems: a stochastic approach. Concurrency and Computation: Practice and Experience, 20(12):1359–1385, August 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2017:NOD

[XWH^{+}17] Rongbin Xu, Yeguo Wang, Wei Huang, Dong Yuan, Ying Xie, and Yun Yang. Near-optimal dynamic priority scheduling strategy for instance-intensive business workflows in cloud computing. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2019:LLA

[XWPM19] Fangliang Xu, Yijie Wang, Xiaoqiang Pei, and Xingkong Ma. LAR: Locality-Aware Reconstruction for erasure-coded distributed storage systems. Concurrency and Computation: Practice and Experience, 31(11):e5031:1–e5031:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2017:MRS

[XWX^{+}17] Jianxun Xia, Fei Wu, Zenggang Xiong, Meikang Qiu, and Changsheng Xie. Modeling recommender systems via weighted bipartite network. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2019:DMR

[XWX^{+}19] Liyang Xu, Miao Wang, Xinhai Xu, Chen Cui, and Xuejun Yang. Dynamic mesh re-partitioning considering iterative convergence rate for multiphase flows in OpenFOAM. Concurrency and Computation: Practice and Experience, 31(21): e5412:1–e5412:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2014:SCP

[XWXC14] Qi Xie, Guilin Wang, Fubiao Xia, and Deren Chen. Selfcertified proxy convertible authenticated encryption: formal definitions and a provably secure scheme. Concurrency and Computation: Practice and Experience, 26(5):1038–1051, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2019:LAM

[XXCY19] Zisang Xu, Cheng Xu, Haixian Chen, and Fang Yang. A lightweight anonymous mutual authentication and key agreement scheme for WBAN. Concurrency and Computation: Practice and Experience, 31(14):e5295:1–e5295:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2017:SLM

[XXLL17] Guoqi Xie, Xiongren Xiao, Renfa Li, and Keqin Li. Schedule length minimization of parallel applications with energy consumption constraints using heuristics on heterogeneous distributed systems. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiong:2015:SRE

[XXX15] Lizhi Xiong, Zhengquan Xu, and Yanyan Xu. A secure reencryption scheme for data services in a cloud computing environment. Concurrency and Computation: Practice and Experience, 27(17):4573–4585, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xue:2016:ESM

[XXY^{+}16] Yunlan Xue, Lingyu Xu, Jie Yu, Lei Wang, and Gaowei Zhang. Event space model in virtual and real society based on special field. Concurrency and Computation: Practice and Experience, 28(15):4161–4176, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2017:OIA

[XY17] Wanru Xu and Panlong Yang. Optimizing the interested area coverage with efficient mobile advertisement user selection. Concurrency and Computation: Practice and Experience, 29 (19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2016:USC

[XYER16] Chengwei Xiao, Jiaqi Ye, Rui Máximo Esteves, and Chunming Rong. Using Spearman’s correlation coefficients for exploratory data analysis on big dataset. Concurrency and Computation: Practice and Experience, 28(14):3866–3878, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2018:DAH

[XYL18] Min Xie, Liangcheng Ye, and Jiarong Liang. A t/k diagnosis algorithm on hypercube-like networks. Concurrency and Computation: Practice and Experience, 30(6):??, March 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4358.

Xu:2017:ESI

[XYS17] Zheng Xu, Neil Yen, and Vijayan Sugumaran. Editorial: Special issue on selected technologies and applications in smart city computing. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2018:ISR

[XYSW18] Mengxi Xu, Yun Yang, Quansen Sun, and Xiaobin Wu. Image super-resolution reconstruction based on adaptive sparse representation. Concurrency and Computation: Practice and Experience, 31(7):e4968:1–e4968:??, December 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2009:SIM

[XZ09] Yang Xiang and Wanlei Zhou. Special issue: Multi-core supported network and system security. Concurrency and Computation: Practice and Experience, 21(10):1251–1253, July 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2016:BKB

[XZH^{+}16] Zheng Xu, Hui Zhang, Chuanping Hu, Lin Mei, Junyu Xuan, Kim-Kwang Raymond Choo, Vijayan Sugumaran, and Yiwei Zhu. Building knowledge base of urban emergency events based on crowdsourcing of social media. Concurrency and Computation: Practice and Experience, 28(15):4038–4052, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2017:COG

[XZH^{+}17] Guangbing Xiao, Haibo Zhang, Houcine Hassan, Yawen Chen, Zhiyi Huang, and Ning Sun. A cooperative offloading game on data recovery for reliable broadcast in VANET. Concurrency and Computation: Practice and Experience, 29 (14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2009:NRA

[XZHW09] Yunni Xia, Qingsheng Zhu, Yu Huang, and Zizhen Wang. A novel reduction approach to analyzing QoS of workflow processes. Concurrency and Computation: Practice and Experience, 21(2):205–223, February 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2011:GFC

[XZJ11] Jing Xu, Wen-Tao Zhu, and Wen-Ting Jin. A generic framework for constructing cross-realm C2C-PAKA protocols based on the smart card. Concurrency and Computation: Practice and Experience, 23(12):1386–1398, August 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xuan:2011:CRS

[XZT^{+}11] Kefeng Xuan, Geng Zhao, David Taniar, Maytham Safar, and Bala Srinivasan. Constrained range search query processing on road networks. Concurrency and Computation: Practice and Experience, 23(5):491–504, April 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2016:SRB

[XZZ^{+}16a] Min Xie, Hongyan Zhao, Nana Zhu, Chengju Liu, and Tianying Jiang. Study on the relationship between network position and organizational performance based on meta-analysis. Concurrency and Computation: Practice and Experience, 28 (15):4082–4092, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiong:2016:SAL

[XZZ16b] Ping Xiong, Lefeng Zhang, and Tianqing Zhu. Semantic analysis in location privacy preserving. Concurrency and Computation: Practice and Experience, 28(6):1884–1899, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yanev:2004:PDC

[YA04] Nicola Yanev and Rumen Andonov. Parallel divide and conquer approach for the protein threading problem. Concurrency and Computation: Practice and Experience, 16(9):961– 974, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2007:DPP

[YAA07] X. Yang, A. Akram, and R. J. Allan. Developing portals/ portlets using Enterprise JavaBeans for Grid users. Concurrency and Computation: Practice and Experience, 19(12): 1633–1641, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yagci:2019:PPL

[YAG19] A. Murat Yagci, Tevfik Aytekin, and Fikret S. Gürgen. Parallel pairwise learning to rank for collaborative filtering. Concurrency and Computation: Practice and Experience, 31(15): e5141:1–e5141:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:APM

[Yan19a] Bo Yang. Analysis on profit model of multi-information products logistics using evolutionary game algorithm. Concurrency and Computation: Practice and Experience, 31(9): e4752:1–e4752:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:RRS

[Yan19b] Qin Yang. A robust recommended system based on attack detection. Concurrency and Computation: Practice and Experience, 31(12):e4660:1–e4660:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yzelman:2012:OOB

[YB12] A. N. Yzelman and Rob H. Bisseling. An object-oriented bulk synchronous parallel library for multicore programming. Concurrency and Computation: Practice and Experience, 24 (5):533–553, April 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Youn:2007:GPD

[YBB^{+}07] Choonhan Youn, Chaitan Baru, Karan Bhatia, Sandeep Chandra, Kai Lin, Ashraf Memon, Ghulam Memon, and Dogan Seber. GEONGrid portal: design and implementations. Concurrency and Computation: Practice and Experience, 19 (12):1597–1607, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yoon:2007:PPT

[YBC^{+}07] Young Yoon, James C. Browne, Mathew Crocker, Samit Jain, and Nasim Mahmood. Productivity and performance through components: the ASCI Sweep3D application. Concurrency and Computation: Practice and Experience, 19(5):721–742, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yahya:2010:EEQ

[YBO10] B. Yahya and J. Ben-Othman. Energy efficient and QoS aware medium access control for wireless sensor networks. Concurrency and Computation: Practice and Experience, 22(10): 1252–1266, July 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2017:HCD

[YBX^{+}17] Hong Yao, Changmin Bai, Muzhou Xiong, Deze Zeng, and Zhangjie Fu. Heterogeneous cloudlet deployment and usercloudlet association toward cost effective fog computing. Concurrency and Computation: Practice and Experience, 29(16): ??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2015:MEV

[YBZ^{+}15] Hong Yao, Changmin Bai, Deze Zeng, Qingzhong Liang, and Yuanyuan Fan. Migrate or not? Exploring virtual machine migration in roadside cloudlet-based vehicular cloud. Concurrency and Computation: Practice and Experience, 27(18): 5780–5792, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:SBM

[YC19a] Yubin Yang and Liming Cheng. An SDN-based MTD model. Concurrency and Computation: Practice and Experience, 31(21):e4897:1–e4897:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ye:2019:MUV

[YC19b] Xijun Ye and Bingcong Chen. Model updating and variability analysis of modal parameters for super high-rise structure. Concurrency and Computation: Practice and Experience, 31 (12):e4712:1–e4712:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2011:DIA

[YCL11] Chao-Tung Yang, Keng-Yi Chou, and Kuan-Chou Lai. Design and implementation of an adaptive job allocation strategy for heterogeneous multi-cluster computing systems. Concurrency and Computation: Practice and Experience, 23(15): 1701–1722, October 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2019:DEP

[YCQF19] Yan Yao, Jian Cao, Shiyou Qian, and Shanshan Feng. Decentralized executions of privacy awareness data analytics workflows in the cloud. Concurrency and Computation: Practice and Experience, 31(15):e5063:1–e5063:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2019:IBD

[YCSY19] Haiyang Yu, Yongquan Cai, Richard O. Sinnott, and Zhen Yang. ID-based dynamic replicated data auditing for the cloud. Concurrency and Computation: Practice and Experience, 31(11):e5051:1–e5051:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2008:EGC

[YCW08] Jin Yang, Jiannong Cao, and Weigang Wu. Efficient global checkpointing algorithms for mobile agents. Concurrency and Computation: Practice and Experience, 20(7):825–838, May 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Younas:2007:QAM

[YCWH07] Muhammad Younas, Kuo-Ming Chao, Ping Wang, and Chun-Lung Huang. QoS-aware mobile service transactions in a wireless environment. Concurrency and Computation: Practice and Experience, 19(8):1219–1236, June 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2013:BBS

[YCZ^{+}13] Yang Yan, Xiang Chen, Chunhui Zhou, Xiaofeng Zhong, Ming Zhao, and Jing Wang. Bargaining-based spectrum sharing in cognitive radio network. Concurrency and Computation: Practice and Experience, 25(9):1034–1049, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2013:SDI

[YDB^{+}13] Xiaoyu Yang, Martin T. Dove, Richard P. Bruin, Andrew Walkingshaw, Richard Sinclair, Dan J. Wilson, and Peter Murray-Rust. An e-Science data infrastructure for simulations within Grid computing environment: methods, approaches and practice. Concurrency and Computation: Practice and Experience, 25(3):385–409, March 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2009:NSC

[YDL09] Zhi Yang, Yafei Dai, and Xiaoming Li. The Neutralizer: a self-configurable failure detector for minimizing distributed storage maintenance cost. Concurrency and Computation: Practice and Experience, 21(2):185–204, February 2009. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yamazaki:2014:TDS

[YDS^{+}14] Ichitaro Yamazaki, Tingxing Dong, Raffaele Solcà, Stanimire Tomov, Jack Dongarra, and Thomas Schulthess. Tridiagonalization of a dense symmetric matrix on multiple GPUs and its application to symmetric eigenvalue problems. Concurrency and Computation: Practice and Experience, 26(16): 2652–2666, November 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yebenes:2017:PDS

[YESG^{+}17a] Pedro Yébenes, Jesus Escudero-Sahuquilllo, Pedro J. García, Francisco-J. Alfaro-Cortés, and Francisco J. Quiles. Providing differentiated services, congestion management, and deadlock freedom in dragonfly networks with adaptive routing. Concurrency and Computation: Practice and Experience, 29 (13):??, July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See erratum [YESG^{+}17b].

Yebenes:2017:EPD

[YESG^{+}17b] Pedro Yébenes, Jesus Escudero-Sahuquillo, Pedro J. García, Francisco-J. Alfaro-Cortés, and Francisco J. Quiles. Erratum: Providing differentiated services, congestion management, and deadlock freedom in dragonfly networks with adaptive routing. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [YESG^{+}17a].

Yebenes:2019:HLB

[YESG^{+}19] Pedro Yébenes, Jesus Escudero-Sahuquillo, Pedro J. García, Francisco J. Quiles, and Torsten Hoefler. Head-of-line blocking avoidance in Slim Fly networks using deadlock-free nonminimal and adaptive routing. Concurrency and Computation: Practice and Experience, 31(2):e4441:1–e4441:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yildiz:2013:TME

[YF13] Beytullah Yildiz and Geoffrey C. Fox. Toward a modular and efficient distribution for Web service handlers. Concurrency and Computation: Practice and Experience, 25(3):410–426, March 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2018:ISS

[YFL18] Fengqi Yan, Qiaoqing Fan, and Mingming Lu. Improving semantic similarity retrieval with word embeddings. Concurrency and Computation: Practice and Experience, 30(16): e4489:1–e4489:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yue:2019:RMM

[YG19] Guo Yue and Getachew Nathan Girma. Research on the mass media effects to the learning efficiency based on social agents. Concurrency and Computation: Practice and Experience, 31(9):e5076:1–e5076:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2014:IPE

[YGG14] Xian Yang, Yike Guo, and Li Guo. An iterative parameter estimation method for biological systems and its parallel implementation. Concurrency and Computation: Practice and Experience, 26(6):1249–1267, April 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2005:UOM

[YGL05] Yue Yang, Ganesh Gopalakrishnan, and Gary Lindstrom. UMM: an operational memory model specification framework with integrated model checking capability. Concurrency and Computation: Practice and Experience, 17(5–6):465–487, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2017:CSI

[YGW17] Xuesong Yan, Wenyin Gong, and Qinghua Wu. Contaminant source identification of water distribution networks using cultural algorithm. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:EED

[YHASZ19] Kang Yang, Hongye He, Kamal Al-Sabahi, and Zuping Zhang. EcForest: Extractive document summarization through enhanced sentence embedding and cascade forest. Concurrency and Computation: Practice and Experience, 31(17):e5206:1–e5206:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2013:QMB

[YHH13] Xiao Yu, Zhengyu He, and Bo Hong. A queuing model-based approach for the analysis of transactional memory systems. Concurrency and Computation: Practice and Experience, 25 (6):808–825, April 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2016:IHA

[YHHS16] Xu Yang, Xinyi Huang, Jinguang Han, and Chunhua Su. Improved handover authentication and key pre-distribution for wireless mesh networks. Concurrency and Computation: Practice and Experience, 28(10):2978–2990, July 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yi:2014:EMC

[YHJ^{+}14] Xin Yi, Yingjie Hu, Zhenhong Jia, Liejun Wang, Jie Yang, and Nikola Kasabov. An enhanced multiphase Chan–Vese model for the remote sensing image segmentation. Concurrency and Computation: Practice and Experience, 26(18): 2893–2906, December 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2009:GBG

[YHK09] Chao-Tung Yang, Tsu-Fen Han, and Heng-Chuan Kan. G-BLAST: a Grid-based solution for mpiBLAST on computational Grids. Concurrency and Computation: Practice and Experience, 21(2):225–255, February 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2019:WIS

[YHYY19] Zhigang Yan, Jiazheng Han, Jieqing Yu, and Yuanxuan Yang. Water inrush sources monitoring and identification based on mine IoT. Concurrency and Computation: Practice and Experience, 31(10):e4843:1–e4843:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yamashita:2019:BED

[YIN19] Kohei Yamashita, Yasuaki Ito, and Koji Nakano. Bulk execution of the dynamic programming for the optimal polygon triangulation problem on the GPU. Concurrency and Computation: Practice and Experience, 31(19):e4947:1–e4947:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2012:RTQ

[YJL12] Fuzheng Yang, Liangliang Jiang, and Xiao Li. Real-time quality assessment for Voice over IP. Concurrency and Computation: Practice and Experience, 24(11):1192–1199, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2019:SSI

[YJZZ19] Dongjin Yu, Yike Jin, Yuqun Zhang, and Xi Zheng. A survey on security issues in services communication of microservicesenabled fog applications. Concurrency and Computation: Practice and Experience, 31(22):e4436:1–e4436:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Youn:2010:MPS

[YK10] Choonhan Youn and Tim Kaiser. Management of a parameter sweep for scientific applications on cluster environments. Concurrency and Computation: Practice and Experience, 22 (18):2381–2400, December 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:MGL

[YKA^{+}19] Bing Yang, Kenneth B. Kent, Eric Aubanel, Stephen MacKay, and Tobi Agila. A multi-granularity locking scheme for Java PackedObjects based on a concurrent multiway tree. Concurrency and Computation: Practice and Experience, 31 (11):e5024:1–e5024:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2015:PSL

[YKD^{+}15] Lin Yao, Lin Kang, Fangyu Deng, Jing Deng, and Guowei Wu. Protecting source-location privacy based on multirings in wireless sensor networks. Concurrency and Computation: Practice and Experience, 27(15):3863–3876, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yue:2001:IDP

[YL01] Kelvin K. Yue and David J. Lilja. Implementing a dynamic processor allocation policy for multiprogrammed parallel applications in the Solaris^{T M} [operating system]. Concurrency and Computation: Practice and Experience, 13(6):449–464, May 2001. CODEN CCPEBO. ISSN 1532-0626 (print),

1532-0634 (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/79503113/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
79503113&PLACEBO=IE.pdf.

Yuan:2016:UCS

[YL16] Bianqing Yuan and Jiqiang Liu. A universally composable secure grouping-proof protocol for RFID tags. Concurrency and Computation: Practice and Experience, 28(6):1872–1883, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2019:RFD

[YL19] Wei Yu and Shijun Li. Research on financial data analysis based on data mining algorithm. Concurrency and Computation: Practice and Experience, 31(10):e4780:1–e4780:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2011:RBU

[YLC11] Chao-Tung Yang, Fang-Yie Leu, and Sung-Yi Chen. Resource brokering using a multi-site resource allocation strategy for computational grids. Concurrency and Computation: Practice and Experience, 23(6):573–594, April 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2013:SIPa

[YLD13] Jun Yang, Wenmin Lin, and Wanchun Dou. Special issue papers: An adaptive service selection method for cross-cloud service composition. Concurrency and Computation: Practice and Experience, 25(18):2435–2454, December 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Youn:2014:DWB

[YLEB14] Choonhan Youn, Jinchi Lu, Ahmed Elgamal, and Chaitan Baru. Development of web-based science portal for large-scale computing collaboration in earthquake engineering. Concurrency and Computation: Practice and Experience, 26(18): 2907–2916, December 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2019:CJR

[YLH^{+}19] Dandong Yin, Yan Liu, Hao Hu, Jeff Terstriep, Xingchen Hong, Anand Padmanabhan, and Shaowen Wang. CyberGIS– Jupyter for reproducible and scalable geospatial analytics. Concurrency and Computation: Practice and Experience, 31 (11):e5040:1–e5040:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2013:SIP

[YLJZ13] Yanli Yu, Keqiu Li, Yingwei Jin, and Yong Zhang. Special issue papers: A trust management model for service-oriented distributed networks. Concurrency and Computation: Practice and Experience, 25(14):2098–2111, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2018:CLC

[YLL^{+}18] Jie Yu, Guangming Liu, Xiaoyong Li, Wenrui Dong, and Qiong Li. Cross-layer coordination in the I/O software stack of extreme-scale systems. Concurrency and Computation: Practice and Experience, 30(10):??, May 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4396.

Yang:2018:DNN

[YLLC18] Lin Yang, Xuchen Lin, Ting-Jie Lu, and Xia Chen. The debate on network neutrality regulation: a technological innovation perspective. Concurrency and Computation: Practice and Experience, 30(24):e4930:1–e4930:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2009:TLO

[YLLZ09] Jin-Min Yang, Kin Fun Li, Wen-Wei Li, and Da-Fang Zhang. Trading off logging overhead and coordinating overhead to achieve efficient rollback recovery. Concurrency and Computation: Practice and Experience, 21(6):819–853, April 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2013:SIPb

[YLR^{+}13] Chao-Tung Yang, Jung-Chun Liu, Rajiv Ranjan, Wen-Chung Shih, and Chih-Hao Lin. Special issue papers: On construction of heuristic QoS bandwidth management in clouds. Concurrency and Computation: Practice and Experience, 25(18): 2540–2560, December 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2018:IOP

[YLWZ18] Xiaoyuan Yang, Xiaoshuang Luo, Xu An Wang, and Shuaiwei Zhang. Improved outsourced private set intersection protocol based on polynomial interpolation. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yoon:2004:SMM

[YLY04] W. Y. Yoon, D. Lee, and H. Y. Youn. On the scalability of many-to-many reliable multicast sessions. Concurrency and Computation: Practice and Experience, 16(13):1353–1363, November 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2018:NSS

[YLZ18] Haipeng Yao, Huiwen Liu, and Peiying Zhang. A novel sentence similarity model with word embedding based on convolutional neural network. Concurrency and Computation: Practice and Experience, 30(16):e4415:1–e4415:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2016:ERA

[YMLR16] Yong Yu, Yi Mu, Rongxing Lu, and Jian Ren. Editorials: Recent advances in security and privacy in large-scale networks. Concurrency and Computation: Practice and Experience, 28 (4):1080–1082, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2016:SSD

[YNX^{+}16] Yong Yu, Jianbing Ni, Qi Xia, Xiaofen Wang, Haomiao Yang, and Xiaosong Zhang. SDIVIP^{2}: shared data integrity verification with identity privacy preserving in mobile clouds. Concurrency and Computation: Practice and Experience, 28(10): 2877–2888, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2016:POS

[YOBS16] Ting Yu, Julian Oppermann, Chris Bradley, and Oliver Sinnen. Performance optimisation strategies for automatically generated FPGA accelerators for biomedical models. Concurrency and Computation: Practice and Experience, 28(5): 1480–1506, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yoshida:2006:OTA

[Yos06] Akimasa Yoshida. An overlapping task assignment scheme for hierarchical coarse-grain task parallel processing. Concurrency and Computation: Practice and Experience, 18(11): 1335–1351, September 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2010:JIP

[YP10] L. Yang and M. R. Poppleton. Java implementation platform for the integrated state- and event-based specification in PROB. Concurrency and Computation: Practice and Experience, 22(8):1007–1022, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2011:TSH

[YPLJ11] Yang Yu, Maolin Pan, Xuguang Li, and Huan Jiang. Tabu search heuristics for workflow resource allocation simulation optimization. Concurrency and Computation: Practice and Experience, 23(16):2020–2033, November 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2015:DCC

[YQL^{+}15] Hao Yin, Bo Qiao, Yan Luo, Chen Tian, and Y. Richard Yang. Demystifying commercial content delivery networks in China. Concurrency and Computation: Practice and Experience, 27(13):3523–3538, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2015:BAD

[YR15] Lawrence Yao and Fethi A. Rabhi. Building architectures for data-intensive science using the ADAGE framework. Concurrency and Computation: Practice and Experience, 27(5): 1188–1206, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yarmolenko:2007:TIE

[YS07] Viktor Yarmolenko and Rizos Sakellariou. Towards increased expressiveness in service level agreements. Concurrency and Computation: Practice and Experience, 19(14):1975–1990, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2017:QAI

[YSC^{+}17] Bowei Yang, Guanghua Song, Yining Chen, Yao Zheng, and Yue Wu. QoS-aware indiscriminate volume storage cloud. Concurrency and Computation: Practice and Experience, 29 (7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2015:DWS

[YSL^{+}15] Minzhi Yan, Hailong Sun, Xudong Liu, Ting Deng, and Xu Wang. Delivering Web service load testing as a service with a global cloud. Concurrency and Computation: Practice and Experience, 27(3):526–545, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:ISO

[YSQM19] Haining Yang, Jiameng Sun, Jing Qin, and Jixin Ma. An improved scheme for outsourced computation with attributebased encryption. Concurrency and Computation: Practice and Experience, 31(21):e4833:1–e4833:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:CPS

[YSW19] Jingya Yang, Linfu Sun, and Qishi Wu. Constraint projections for semi-supervised spectral clustering ensemble. Concurrency and Computation: Practice and Experience, 31(20): e5359:1–e5359:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2017:MLM

[YSWZ17] Lu Yang, Qing Song, Zhihui Wang, and Shihui Zhang. Maximum linear matching: Intelligent and automatic wavelength calibration method. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yangui:2015:SAD

[YT15] Sami Yangui and Samir Tata. The SPD approach to deploy service-based applications in the cloud. Concurrency and Computation: Practice and Experience, 27(15):3943–3960, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yelmewad:2019:PIH

[YT19] Pramod Yelmewad and Basavaraj Talawar. Parallel iterative hill climbing algorithm to solve TSP on GPU. Concurrency and Computation: Practice and Experience, 31(7):e4974:1– e4974:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yamazaki:2017:NGR

[YTD17] Ichitaro Yamazaki, Stanimire Tomov, and Jack Dongarra. Non-GPU-resident symmetric indefinite factorization. Concurrency and Computation: Practice and Experience, 29(5): ??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2001:UKB

[YTF^{+}01] Chao-Tung Yang, Shian-Shyong Tseng, Yun-Woei Fann, Ting-Ku Tsai, Ming-Huei Hsieh, and Cheng-Tien Wu. Using knowledge-based systems for research on parallelizing compilers. Concurrency and Computation: Practice and Experience, 13(3):181–208, March 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/78003111/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=78003111&PLACEBO=IE.pdf.

Yang:2019:REA

[YTL19] Juan Yang, Michael S. C. Thomas, and Hongtao Liu. Rule extraction from autoencoder-based connectionist computational models. Concurrency and Computation: Practice and Experience, 31(3):e4262:1–e4262:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2018:CSB

[Yu18] Lan Yu. Cloud storage-based personalized sports activity management in Internet plus O2O sports community. Concurrency and Computation: Practice and Experience, 30(24): e4932:1–e4932:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2007:IWS

[YWA07] X. Yang, X. D. Wang, and R. Allan. Investigation of WSRP support in selected open-source portal frameworks. Concurrency and Computation: Practice and Experience, 19(12): 1729–1738, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yildiz:2019:PMQ

[YWBS19] Beytullah Yildiz, Kesheng Wu, Suren Byna, and Arie Shoshani. Parallel membership queries on very large scientific data sets using bitmap indexes. Concurrency and Computation: Practice and Experience, 31(15):e5157:1–e5157:??, August 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2011:PBP

[YWC11] Chao-Tung Yang, Chao-Chin Wu, and Jen-Hsiang Chang. Performance-based parallel loop self-scheduling using hybrid OpenMP and MPI programming on multicore SMP clusters. Concurrency and Computation: Practice and Experience, 23 (8):721–744, June 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2017:PIS

[YWL^{+}17a] Jianhua Yan, Licheng Wang, Jing Li, Muzi Li, Yixan Yang, and Wenbin Yao. Pre-image sample algorithm with irregular Gaussian distribution and construction of identity-based signature. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2017:DMB

[YWL^{+}17b] Ming Yang, Shan Wang, Zhen Ling, Yaowen Liu, and Zhenyu Ni. Detection of malicious behavior in Android apps through API calls and permission uses analysis. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2018:FEW

[YWLQ18] Chun Yan, Lin Wang, Wei Liu, and Man Qi. Financial early warning of non-life insurance company based on RBF neural network optimized by genetic algorithm. Concurrency and Computation: Practice and Experience, 30(23):e4343:1– e4343:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2010:SMR

[YWM^{+}10] Hao Yin, Yang Wang, Geyong Min, Sebastien Berton, Rui Guo, and Chuang Lin. A secure multipath routing protocol in mobile ad hoc networks. Concurrency and Computation: Practice and Experience, 22(4):481–502, March 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2019:LPJ

[YWR^{+}19] Ziqi Yan, Qiong Wu, Meng Ren, Jiqiang Liu, Shaowu Liu, and Shuo Qiu. Locally private Jaccard similarity estimation. Concurrency and Computation: Practice and Experience, 31(24): e4889:1–e4889:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2012:MAS

[YWT^{+}12] Zhiwei Yu, Chaokun Wang, Clark Thomborson, Jianmin Wang, Shiguo Lian, and Athanasios V. Vasilakos. Multimedia applications and security in MapReduce: Opportunities and challenges. Concurrency and Computation: Practice and Experience, 24(17):2083–2101, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2010:DSP

[YWY^{+}10] Yanqin Yang, Meng Wang, Haijin Yan, Zili Shao, and Minyi Guo. Dynamic scratch-pad memory management with data pipelining for embedded systems. Concurrency and Computation: Practice and Experience, 22(13):1874–1892, September 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2017:PPD

[YXL17] Lei Yang, Kui Xu, and Shi Liu. PADP: a parallel data possession audit model for cloud storage. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2016:MAM

[YXLZ16] Hong Yao, Jinlai Xu, Zhongwen Luo, and Deze Zeng. MEMoMR: Accelerate MapReduce via reuse of intermediate results. Concurrency and Computation: Practice and Experience, 28(14):3814–3829, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:SFD

[YY19] Guang Yang and Shuofeng Yu. Synthesized fault diagnosis method reasoned from rough set-neural network and evidence theory. Concurrency and Computation: Practice and Experience, 31(9):e4944:1–e4944:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2010:RDI

[YYC10] Chao-Tung Yang, I-Hsien Yang, and Chun-Hsiang Chen. RACAM: design and implementation of a recursively adjusting co-allocation method with efficient replica selection in Data Grids. Concurrency and Computation: Practice and Experience, 22(15):2170–2194, October 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2019:EVM

[YYC^{+}19] Feng Yao, Yiping Yao, Huangke Chen, Tianlin Li, Menglong Lin, and Xiaoxiong Zhang. An efficient virtual machine allocation algorithm for parallel and distributed simulation applications. Concurrency and Computation: Practice and Experience, 31(17):e5237:1–e5237:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2010:ARA

[YYCH10] Chao-Tung Yang, Ming-Feng Yang, Yao-Chun Chi, and Ching-Hsien Hsu. An Anticipative Recursively Adjusting Mechanism for parallel file transfer in data grids. Concurrency and Computation: Practice and Experience, 22(15): 2144–2169, October 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:NPP

[YYK^{+}19] Xu Yang, Xun Yi, Ibrahim Khalil, Hui Cui, Xuechao Yang, Surya Nepal, Xinyi Huang, and Yali Zeng. A new privacypreserving authentication protocol for anonymous web browsing. Concurrency and Computation: Practice and Experience, 31(21):e4706:1–e4706:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yuan:2012:DDB

[YYL^{+}12] Dong Yuan, Yun Yang, Xiao Liu, Gaofeng Zhang, and Jinjun Chen. A data dependency based strategy for intermediate data storage in scientific cloud workflow systems. Concurrency and Computation: Practice and Experience, 24(9): 956–976, June 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ye:2019:NAR

[YYLC19] Ayong Ye, Xiaoliang Yang, Qing Li, and Aimin Chen. A novel adaptive radio map for RSS-based indoor positioning. Concurrency and Computation: Practice and Experience, 31(23): e4486:1–e4486:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yeh:2015:IPP

[YYS15] Tsozen Yeh, Shuwen Yang, and Yifeng Sun. Improving the program performance through prioritized memory management and disk operation. Concurrency and Computation: Practice and Experience, 27(13):3345–3361, September 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yong:2019:SCE

[YYWQ19] Xiang Yong, Cui Yan, Weiyao, and Xu Qiang. A study on the construction of evaluative system of Smart City based on entropy-weighted matter-element model method. Concurrency and Computation: Practice and Experience, 31(9): e5010:1–e5010:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:AOP

[YYXL19] Fan Yang, Limei Yan, Jianjun Xu, and Hongyu Li. Analysis of an optimal PMU configuration method based on incomplete observation. Concurrency and Computation: Practice and Experience, 31(12):e4835:1–e4835:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2017:WIS

[YYZ^{+}17] Haipeng Yao, Hao Yang, Anqi Zhang, Chao Fang, and Yiru Guo. WLAN interference self-optimization using SOM neural networks. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2019:RCR

[YYZH19] Qing Yang, Zhong Yang, Tianyi Zhang, and Guoxiong Hu. A random chemical reaction optimization algorithm based on dual containers strategy for multi-rotor UAV path planning in transmission line inspection. Concurrency and Computation: Practice and Experience, 31(12):e4658:1–e4658:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2010:PAE

[YZ10] Huai-Zhong Yu and Qing-Yong Zhu. A probabilistic approach for earthquake potential evaluation based on the load/unload response ratio method. Concurrency and Computation: Practice and Experience, 22(12):1520–1533, August 25, 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2019:DAB

[YZ19] Chensi Yu and Wei Zha. Detection and application of breaking of automobile mechanical transmission rod based on ant colony algorithm. Concurrency and Computation: Practice and Experience, 31(10):e4759:1–e4759:??, May 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2017:SKS

[YZCT17] Yang Yang, Xianghan Zheng, Victor Chang, and Chunming Tang. Semantic keyword searchable proxy re-encryption for postquantum secure cloud storage. Concurrency and Computation: Practice and Experience, 29(19):??, October 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2014:CLE

[YZR14] Chao Yu, Minjie Zhang, and Fenghui Ren. Coordinated learning by exploiting sparse interaction in multiagent systems. Concurrency and Computation: Practice and Experience, 26 (1):51–70, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2015:DBT

[YZW^{+}15] Fan Yang, Fengli Zhang, Jiahao Wang, Zhiguang Qin, and Xiaolu Yuan. Distance-bounding trust protocol in anonymous radio-frequency identification systems. Concurrency and Computation: Practice and Experience, 27(16):4211– 4229, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2017:IAM

[YZXW17] Chunyong Yin, Sun Zhang, Jinwen Xi, and Jin Wang. An improved anonymity model for big data security based on clustering algorithm. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2010:PPL

[YZZ^{+}10] Xiang-Chu Yin, Lang-Ping Zhang, Yongxian Zhang, Keyin Peng, Haitao Wang, Zhiping Song, Xiaotao Zhang, and Shuai Yuan. The peak point of LURR and its significance. Concurrency and Computation: Practice and Experience, 22(12): 1549–1558, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhai:2019:ZSB

[ZAB^{+}19] Xiaojun Zhai, Abbes Amira, Faycal Bensaali, AlMaha Al-Shibani, Asma Al-Nassr, Asmaa El-Sayed, Mohammad Eslami, Sarada Prasad Dakua, and Julien Abinahed. Zynq SoC based acceleration of the lattice Boltzmann method. Concurrency and Computation: Practice and Experience, 31(17): e5184:1–e5184:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zaourar:2018:TMF

[ZABP18] Lilia Zaourar, Massinissa Ait Aba, David Briand, and Jean-Marc Philippe. Task management on fully heterogeneous micro-server system: Modeling and resolution strategies. Concurrency and Computation: Practice and Experience, 30(23):e4798:1–e4798:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2016:IPS

[ZACG16] Xianghan Zheng, Dongyun An, Xing Chen, and Wenzhong Guo. Interest prediction in social networks based on Markov chain modeling on clustered users. Concurrency and Computation: Practice and Experience, 28(14):3895–3909, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2007:COI

[ZBC^{+}07] Shujia Zhou, V. Balaji, Carlos Cruz, Arlindo da Silva, Chris Hill, Erik Kluzek, Shep Smithline, Atanas Trayanov, and Weiyu Yang. Cross-organization interoperability experiments of weather and climate models with the Earth System Modeling Framework. Concurrency and Computation: Practice and Experience, 19(5):583–592, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zbakh:2017:ECC

[ZBE17] Mostapha Zbakh, Mohamed Bakhouya, and Mohamed Essaaidi. Editorial: Cloud computing and big data: Technologies and applications. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zbakh:2018:CCB

[ZBEM18] Mostapha Zbakh, Mohamed Bakhouya, Mohamed Essaaidi, and Pierre Manneback. Cloud computing and big data: Technologies and applications. Concurrency and Computation: Practice and Experience, 30(12):??, June 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4517.

Zaia:2006:UGP

[ZBP06] Angelo Zaia, Dario Bruneo, and Antonio Puliafito. Using the Grid paradigm for multimedia applications. Concurrency and Computation: Practice and Experience, 18(8):899–910, July 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2007:UVM

[ZBP07] Xin Zhao, Kevin Borders, and Atul Prakash. Using a virtual machine to protect sensitive Grid resources. Concurrency and Computation: Practice and Experience, 19(14): 1917–1935, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zeljkovic:2015:EOD

[ZBZ^{+}15] Vesna Zeljković, Milena Bojic, Shengwei Zhao, Claude Tameze, and Ventzeslav Valev. Exudates and optic disk detection in retinal images of diabetic patients. Concurrency and Computation: Practice and Experience, 27(1):172–192, January 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2018:SCM

[ZBZ^{+}18] Guoliang Zhang, Weidong Bao, Xiaomin Zhu, Weiwei Zhao, and Huining Yan. A server consolidation method with integrated deep learning predictor in local storage based clouds. Concurrency and Computation: Practice and Experience, 30(16):e4503:1–e4503:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2011:ASP

[ZBZH11] Zhangbing Zhou, Sami Bhiri, Hai Zhuge, and Manfred Hauswirth. Assessing service protocol adaptability based on protocol reduction and graph search. Concurrency and Computation: Practice and Experience, 23(9):880–904, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zolkiewski:2019:FSI

[ZC19] Slawomir Zolkiewski and Leszek Chybowski. Foreword to the special issue of the World Conference on Information Systems and Technologies (WorldCIST 2017). Concurrency and Computation: Practice and Experience, 31(22):e5419:1–e5419:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2006:TBA

[ZCC^{+}06] Yang Zhang, Jiannong Cao, Xiaolin Chen, Sanglu Lu, and Li Xie. Threshold-based admission control for a multimedia Grid: analysis and performance evaluation. Concurrency and Computation: Practice and Experience, 18(14): 1747–1758, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2012:ACW

[ZCD^{+}12] Shujia Zhou, Carlos Cruz, Daniel Duffy, Robert Tucker, and Mark Purcell. Accelerating climate and weather simulations through hybrid computing. Concurrency and Computation: Practice and Experience, 24(1):54–61, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2018:MPA

[ZCH^{+}18] Zhou Zhou, Jian Chang, Zhigang Hu, Junyang Yu, and Fangmin Li. A modified PSO algorithm for task scheduling optimization in cloud computing. Concurrency and Computation: Practice and Experience, 30(24):e4970:1–e4970:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2014:PBF

[ZCL14] Zhuoyao Zhang, Ludmila Cherkasova, and Boon Thau Loo. Parameterizable benchmarking framework for designing a MapReduce performance model. Concurrency and Computation: Practice and Experience, 26(12):2005–2026, August 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2018:MGP

[ZCL^{+}18] Wei Zhou, Zhanxiu Cai, Bo Lian, Jincai Wang, Jianping Ma, Bin Sun, and Qian Yu. A multi-GPU protein database search model with hybrid alignment manner on distributed GPU clusters. Concurrency and Computation: Practice and Experience, 30(18):e4522:1–e4522:??, September 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2019:RHP

[ZCL^{+}19] Taoyun Zhou, Yun Cheng, Baowang Lian, Bibi Irfana, and Hongwei Zhao. Research on high-precision passive localization based on phase difference changing rate. Concurrency and Computation: Practice and Experience, 31(23):e4492:1– e4492:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2019:ECT

[ZCLL19] Qiansheng Zhao, Zi Chen, Chang Liu, and Nianxue Luo. Extracting and classifying typhoon disaster information based on volunteered geographic information from Chinese Sina microblog. Concurrency and Computation: Practice and Experience, 31(9):e4910:1–e4910:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2006:TBP

[ZCS06] Hai Zhuge, Xue Chen, and Xiaoping Sun. Trust-based probabilistic search with the view model of peer-to-peer networks. Concurrency and Computation: Practice and Experience, 18 (14):1839–1855, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2018:CMT

[ZCW^{+}18] Meng Zhao, Liyi Chen, Honggang Wu, Wei Guo, and Changwen Ye. Computer microscopic test study on the suitability of modified polyimide grease used to improve collapsible loess railway subgrade. Concurrency and Computation: Practice and Experience, 30(24):e4900:1–e4900:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2017:EBA

[ZCXH17] Ning Zhou, Riqing Chen, Yuanqing Xia, and Jie Huang. Estimator-based adaptive neural network control of leaderfollower high-order nonlinear multiagent systems with actuator faults. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:GPB

[ZCZ^{+}19] Jinghui Zhang, Jian Chen, Jun Zhan, Jiahui Jin, and Aibo Song. Graph partition-based data and task co-scheduling of scientific workflow in geo-distributed datacenters. Concurrency and Computation: Practice and Experience, 31(24): e5245:1–e5245:??, December 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2007:APC

[ZDA^{+}07] Chongjie Zhang, Chirag Dekate, Gabrielle Allen, Ian Kelley, and Jon MacLaren. An application portal for collaborative coastal modeling. Concurrency and Computation: Practice and Experience, 19(12):1571–1581, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2014:XPF

[ZDB^{+}14] Fan Zhang, Ciprian Docan, Hoang Bui, Manish Parashar, and Scott Klasky. XpressSpace: a programming framework for coupling partitioned global address space simulation codes. Concurrency and Computation: Practice and Experience, 26 (3):644–661, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2009:IIC

[ZDC^{+}09] Shujia Zhou, Daniel Duffy, Thomas Clune, Max Suarez, Samuel Williams, and Milton Halem. The impact of IBM Cell technology on the programming paradigm in the context of computer systems for climate and weather models. Concurrency and Computation: Practice and Experience, 21(17): 2176–2186, December 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2015:SEM

[ZDC15] Shaoqian Zhang, Wanchun Dou, and Jinjun Chen. A service evaluation method for cross-cloud service choreography. Concurrency and Computation: Practice and Experience, 27(9): 2373–2389, June 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zou:2014:SSC

[ZDG^{+}14] Dan Zou, Yong Dou, Song Guo, Rongchun Li, and Lin Deng. Supernodal sparse Cholesky factorization on graphics processing units. Concurrency and Computation: Practice and Experience, 26(16):2713–2726, November 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2018:EIE

[ZDHJ18] Bo Zhang, Yuanyuan Dou, Quan Hong, and Honghu Ji. Experimental investigation of effect of tooth geometrical parameters to flow characteristics in slant labyrinth seals. Concurrency and Computation: Practice and Experience, 30(24): e4951:1–e4951:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2007:NSR

[ZDL07] Hai Zhuge, Lianhong Ding, and Xiang Li. Networking scientific resources in the Knowledge Grid environment. Concurrency and Computation: Practice and Experience, 19(7): 1087–1113, May 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:PHM

[ZDL19] Xinwei Zhang, Shaohua Dong, and Yanhui Liu. p-hub median location optimization of hub-and-spoke air transport networks in express enterprise. Concurrency and Computation: Practice and Experience, 31(9):e4981:1–e4981:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2018:ACA

[ZDR^{+}18] Naweiluo Zhou, Gwenaël Delaval, Bogdan Robu, Éric Rutten, and Jean-François Méhaut. An autonomic-computing approach on mapping threads to multi-cores for software transactional memory. Concurrency and Computation: Practice and Experience, 30(18):e4506:1–e4506:??, September 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zou:2012:OSP

[ZDX12] Dan Zou, Yong Dou, and Fei Xia. Optimization schemes and performance evaluation of Smith–Waterman algorithm on CPU, GPU and FPGA. Concurrency and Computation: Practice and Experience, 24(14):1625–1644, September 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2010:SLB

[ZEB10] Ming Zhang, Elie El Ajaltouni, and Azzedine Boukerche. A scheduling and load balancing scheme for dynamic P2P-based system. Concurrency and Computation: Practice and Experience, 22(10):1325–1334, July 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zeng:2019:ORA

[Zen19a] Chenxi Zeng. Optimal resource allocation in mobile satellite networks: a noncooperative dynamic game-based approach. Concurrency and Computation: Practice and Experience, 31 (10):e4857:1–e4857:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zeng:2019:MCB

[Zen19b] Wen Zeng. A methodology for cost-benefit analysis of information security technologies. Concurrency and Computation: Practice and Experience, 31(7):e5004:1–e5004:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zakay:2014:WRP

[ZF14] Netanel Zakay and Dror G. Feitelson. Workload resampling for performance evaluation of parallel job schedulers. Concurrency and Computation: Practice and Experience, 26(12): 2079–2105, August 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zdunek:2018:DGN

[ZF18] Rafal Zdunek and Krzysztof Fonal. Distributed geometric nonnegative matrix factorization and hierarchical alternating least squares-based nonnegative tensor factorization with the MapReduce paradigm. Concurrency and Computation: Practice and Experience, 30(17):e4473:1–e4473:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2016:CER

[ZFJ16] Jingya Zhou, Jianxi Fan, and Juncheng Jia. A cost-efficient resource provisioning algorithm for DHT-based cloud storage systems. Concurrency and Computation: Practice and Experience, 28(18):4485–4506, December 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2008:TWS

[ZFT08] Yanchong Zheng, Yushun Fan, and Wei Tan. Towards workflow simulation in service-oriented architecture: an eventbased approach. Concurrency and Computation: Practice and Experience, 20(4):315–330, March 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2017:TTM

[ZFW^{+}17] Jingya Zhou, Jianxi Fan, Jin Wang, Baolei Cheng, and Juncheng Jia. Towards traffic minimization for data placement in online social networks. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2019:DSD

[ZFWJ19] Jingya Zhou, Jianxi Fan, Jin Wang, and Juncheng Jia. Dynamic service deployment for budget-constrained mobile edge computing. Concurrency and Computation: Practice and Experience, 31(24):e5436:1–e5436:??, December 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2019:TTA

[ZFXJ19] Xianqian Zhou, Liyuan Fang, Haiming Xie, and Wanchun Jiang. TAP: Timeliness-aware predication-based replica selection algorithm for key–value stores. Concurrency and Computation: Practice and Experience, 31(17):e5171:1–e5171:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2004:HRT

[ZG04] Hui Zhao and Nicolas D. Georganas. HLA real-time extension. Concurrency and Computation: Practice and Experience, 16(15):1503–1525, December 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2007:PEK

[ZGL07] Hai Zhuge, Weiyu Guo, and Xiang Li. The potential energy of knowledge flow. Concurrency and Computation: Practice and Experience, 19(15):2067–2090, October 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:STI

[ZGL19] Yuanchao Zhang, Jingwei Gao, and Qiao Li. Study on tire– ice traction using a combined neural network and secondary development finite element modeling method. Concurrency and Computation: Practice and Experience, 31(9):e5045:1– e5045:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhi-Gang:2010:STP

[ZGRSC10] Shao Zhi-Gang, Fu Rong-Shan, and Jiang Changsheng. Short-time postseismic deformation of 2001 Ms8.1 Kunlun (China) earthquake. Concurrency and Computation: Practice and Experience, 22(12):1803–1812, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2017:DSD

[ZGS17] Xu Zhang, Naijie Gu, and Junjie Su. DCUDP: scalable data transfer for high-speed long-distance networks. Concurrency and Computation: Practice and Experience, 29(4):??, February 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2008:MTS

[ZGST08] Jun Zhao, Carole Goble, Robert Stevens, and Daniele Turi. Mining Taverna’s Semantic Web of provenance. Concurrency and Computation: Practice and Experience, 20(5):463–472, April 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2009:HCE

[zGWXT09] Chong zhi Gao, Baodian Wei, Dongqing Xie, and Chunming Tang. How to construct efficient on-line/off-line threshold signature schemes through the simulation approach. Concurrency and Computation: Practice and Experience, 21(10): 1351–1372, July 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2011:SED

[ZGX11] Wen Tao Zhu, Fei Gao, and Yang Xiang. A secure and efficient data aggregation scheme for wireless sensor networks. Concurrency and Computation: Practice and Experience, 23 (12):1414–1430, August 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:FPI

[ZGZ19] Shunli Zhang, Guohua Geng, and Jian Zhao. Fast parallel image reconstruction for cone-beam FDK algorithm. Concurrency and Computation: Practice and Experience, 31(10): e4697:1–e4697:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2008:RFS

[ZH08] Jiaying Zhang and Peter Honeyman. A replicated file system for Grid computing. Concurrency and Computation: Practice and Experience, 20(9):1113–1130, June 25, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2015:ADC

[ZH15] Wei Zheng and Shouhui Huang. An adaptive deadline constrained energy-efficient scheduling heuristic for workflows in clouds. Concurrency and Computation: Practice and Experience, 27(18):5590–5605, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2016:DAN

[ZH16] Jinson Zhang and Mao Lin Huang. Density approach: a new model for BigData analysis and visualization. Concurrency and Computation: Practice and Experience, 28(3):661–673, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2008:SIR

[Zha08] Ben Y. Zhao. Special issue: Recent advances in peer-to-peer systems and security. Concurrency and Computation: Practice and Experience, 20(2):125–126, February 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhai:2019:DVP

[Zha19] Sheping Zhai. Distance variety preserving hashing for largescale retrieval. Concurrency and Computation: Practice and Experience, 31(12):e4829:1–e4829:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2018:SRS

[ZHC^{+}18] Lu Zhang, Gaofeng He, Jie Cao, Haiting Zhu, and Bingfeng Xu. Spotting review spammer groups: a cosine pattern and network based method. Concurrency and Computation: Practice and Experience, 30(20):e4686:1–e4686:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2016:AAI

[Zhe16] Jie Zheng. An association analysis and identification for unknown protocol of bitstream oriented. Concurrency and Computation: Practice and Experience, 28(15):4067–4081, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2016:EGD

[ZHGX16] Wei Zhou, Jizhong Han, Yun Gao, and Zhiyong Xu. An efficient graph data processing system for large-scale social network service applications. Concurrency and Computation: Practice and Experience, 28(3):729–747, March 10, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:TDS

[ZHJ19] Maosheng Zhang, Ruimin Hu, and Lin Jiang. Threedimensional sound reproduction in vehicle based on data mining technique. Concurrency and Computation: Practice and Experience, 31(4):e4936:1–e4936:??, February 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2017:MAF

[ZHM^{+}17] Quanxin Zhao, Shuai Han, Yuming Mao, Supeng Leng, Geyong Min, Jia Hu, and Noushin Najjari. A Markovian analytical framework for public-safety video sharing by deviceto-device communications. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2006:CCM

[Zho06] S. J. Zhou. Coupling climate models with the Earth System Modeling Framework and the Common Component Architecture. Concurrency and Computation: Practice and Experience, 18(2):203–213, February 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2008:ATP

[ZHT08] Ying-Nan Zhang, Qing-Ni Hu, and Hong-Fei Teng. Active target particle swarm optimization. Concurrency and Computation: Practice and Experience, 20(1):29–40, January 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2007:ASL

[Zhu07] Hai Zhuge. Autonomous semantic link networking model for the Knowledge Grid. Concurrency and Computation: Practice and Experience, 19(7):1065–1085, May 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2015:ESK

[Zhu15] Hai Zhuge. Editorials: Semantics, knowledge and grids. Concurrency and Computation: Practice and Experience, 27(15): 3912–3914, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2018:CPS

[Zhu18] Dingju Zhu. Cloud parallel spatial-temporal data model with intelligent parameter adaptation for spatial-temporal big data. Concurrency and Computation: Practice and Experience, 30(22):e4497:1–e4497:??, November 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2019:DGN

[Zhu19] Dapeng Zhu. Digital generation of non-Gaussian random vibration signals in railway transportation and package response analysis. Concurrency and Computation: Practice and Experience, 31(10):e4795:1–e4795:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2016:PFI

[ZHW^{+}16] Wei Zhou, Jiankun Hu, Song Wang, Ian Petersen, and Mohammed Bennamoun. Partial fingerprint indexing: a combination of local and reconstructed global features. Concurrency and Computation: Practice and Experience, 28(10): 2940–2957, July 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2019:LST

[ZHX^{+}19] Yu-Long Zhou, Ren-Jie Han, Qian Xu, Qi-Jie Jiang, and Wei-Ke Zhang. Long short-term memory networks for CSI300 volatility prediction with Baidu search volume. Concurrency and Computation: Practice and Experience, 31(10):e4721:1– e4721:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2013:SIP

[ZHZ^{+}13] Wei Zhou, Jizhong Han, Zhang Zhang, Jiao Dai, and Zhiyong Xu. Special issue papers: HDKV: supporting efficient high-dimensional similarity search in key-value stores. Concurrency and Computation: Practice and Experience, 25(12): 1675–1698, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zicari:2012:MWC

[Zic12] Roberto V. Zicari. Managing Web 2.0 content. Concurrency and Computation: Practice and Experience, 24(17): 2165–2166, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zahavi:2010:OIF

[ZJKL10] Eitan Zahavi, Gregory Johnson, Darren J. Kerbyson, and Michael Lang. Optimized InfiniBand^{T M} fat-tree routing for shift all-to-all communication patterns. Concurrency and Computation: Practice and Experience, 22(2):217–231, February 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2013:EER

[ZJL13] Jianhua Zhang, Yun Jiang, and Xiaofan Li. Energy-efficient resource allocation in multiuser relay-based OFDMA networks. Concurrency and Computation: Practice and Experience, 25(9):1113–1125, June 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2015:SSS

[ZJL15] Liang Zhu, Hai Jin, and Xiaofei Liao. SymS: a symmetrical scheduler to improve multi-threaded program performance on NUMA systems. Concurrency and Computation: Practice and Experience, 27(18):5810–5825, December 25, 2015.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zima:2011:FTB

[ZJS11] Hans P. Zima, Mark L. James, and Paul L. Springer. Faulttolerant on-board computing for robotic space missions. Concurrency and Computation: Practice and Experience, 23(17): 2192–2204, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2017:MSD

[ZJS^{+}17] Fan Zhang, Tong Jin, Qian Sun, Melissa Romanus, Hoang Bui, Scott Klasky, and Manish Parashar. In-memory staging and data-centric task placement for coupled scientific simulation workflows. Concurrency and Computation: Practice and Experience, 29(12):??, June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:HCM

[ZJT^{+}19] Zhewei Zhang, Tao Jing, Chunhua Tian, Meilin Gao, and Xuejing Li. Hybrid context modeling and wavelet transform for region of interest detection and rate control based on VP9 video codec. Concurrency and Computation: Practice and Experience, 31(10):e4855:1–e4855:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zaharia:2008:GBS

[ZK08] M. Zaharia and S. Keshav. Gossip-based search selection in hybrid peer-to-peer networks. Concurrency and Computation: Practice and Experience, 20(2):139–153, February 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2007:GPS

[ZKA07] Chongjie Zhang, Ian Kelley, and Gabrielle Allen. Grid portal solutions: a comparison of GridPortlets and OGCE. Concurrency and Computation: Practice and Experience, 19(12): 1739–1748, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2007:HSN

[ZKJ^{+}07] S. Zhou, W. Kuang, W. Jiang, P. Gary, J. Palencia, and G. Gardner. High-speed network and Grid computing for high-end computation: application in geodynamics ensemble simulations. Concurrency and Computation: Practice and Experience, 19(5):573–582, April 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhizhin:2007:IMD

[ZKR^{+}07] M. Zhizhin, E. Kihn, R. Redmon, A. Poyda, D. Mishin, D. Medvedev, and V. Lyutsarev. Integrating and mining distributed environmental archives on Grids. Concurrency and Computation: Practice and Experience, 19(16):2157–2170, November 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zin:2017:ACI

[ZKWK17] HyeonCheol Zin, ChongGun Kim, Mary Wu, and Shinheon Kim. Avoidance of channel interference in polygonal IoT networks. Concurrency and Computation: Practice and Experience, 29(11):??, June 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2006:LAC

[ZL06] Hai Zhuge and Yanyan Li. Learning with an active e-course in the Knowledge Grid environment. Concurrency and Computation: Practice and Experience, 18(3):333–356, March 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2009:SIW

[ZL09] Hai Zhuge and Qing Li. Special issue: Web 2.0, semantics, knowledge and Grid. Concurrency and Computation: Practice and Experience, 21(5):557–559, April 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2012:PEF

[ZL12] X. Zhao and H. Liang. A performance evaluation framework for packet scheduling algorithms in wireless system. Concurrency and Computation: Practice and Experience, 24(11): 1180–1191, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:IAQ

[ZL19] Bi Zhang and Wei Li. Intelligent air quality detection based on genetic algorithm and neural network: an urban China case study. Concurrency and Computation: Practice and Experience, 31(10):e4673:1–e4673:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2015:ACM

[ZLA^{+}15] Jun Zhu, Alexey Lastovetsky, Shoukat Ali, Rolf Riesen, and Khalid Hasanov. Asymmetric communication models for resource-constrained hierarchical Ethernet networks. Concurrency and Computation: Practice and Experience, 27(6): 1575–1590, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2017:KDC

[ZLAa^{+}17] Yichao Zhang, Xing Li, M. A. Aziz-alaoui, Cyrille Bertelle, Jihong Guan, and Shuigeng Zhou. Knowledge diffusion in complex networks. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2015:CSI

[ZlC15] Yang Zhang and Jun liang Chen. Constructing scalable Internet of Things services based on their event-driven models. Concurrency and Computation: Practice and Experience, 27(17):4819–4851, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhong:2017:MMG

[ZLC17a] Hong Zhong, Weiya Luo, and Jie Cui. Multiple multicast group key management for the Internet of People. Concurrency and Computation: Practice and Experience, 29(3): ??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

ZhuanSun:2017:MRA

[ZLC17b] Zongliang ZhuanSun, Keqiu Li, and Guolong Chen. Multipath routing algorithm for application-specific wormhole NoCs. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zahn:2019:LWS

[ZLF19] Felix Zahn, Steffen Lammel, and Holger Fröning. On link width scaling for energy-proportional direct interconnection networks. Concurrency and Computation: Practice and Experience, 31(2):e4439:1–e4439:??, January 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2019:IGB

[ZLG^{+}19] Fan Zhao, Xiangyang Luo, Yong Gan, Shuodi Zu, Qingfeng Cheng, and Fenlin Liu. IP geolocation based on identification routers and local delay distribution similarity. Concurrency and Computation: Practice and Experience, 31(22):e4722:1– e4722:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2015:SSP

[ZLH^{+}15] Yonglong Zhang, Bin Li, Zhiqiu Huang, Jin Wang, and Junwu Zhu. SGAM: strategy-proof group buying-based auction mechanism for virtual machine allocation in clouds. Concurrency and Computation: Practice and Experience, 27(18): 5577–5589, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2018:TLB

[ZLH^{+}18] Baobing Zhang, Siguang Li, Zhengwen Huang, Babak H. Rahi, Qicong Wang, and Maozhen Li. Transfer learningbased online multiperson tracking with Gaussian process regression. Concurrency and Computation: Practice and Experience, 30(23):e4917:1–e4917:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zefreh:2017:DDP

[ZLKK17] Ebrahim Zarei Zefreh, Shahriar Lotfi, Leyli Mohammad Khanli, and Jaber Karimpour. 3-D data partitioning for 3level perfectly nested loops on heterogeneous distributed systems. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:VFS

[ZLL19] Guijuan Zhang, Dianjie Lu, and Hong Liu. Visualization of fluid simulation: an SPH-based multi-resolution method. Concurrency and Computation: Practice and Experience, 31(23):e4509:1–e4509:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2011:ERQ

[ZLLL11] Yiming Zhang, Ling Liu, Xicheng Lu, and Dongsheng Li. Efficient range query processing over DHTs based on the balanced Kautz tree. Concurrency and Computation: Practice and Experience, 23(8):796–814, June 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2013:SIPc

[ZLN^{+}13] Xuyun Zhang, Chang Liu, Surya Nepal, Chi Yang, Wanchun Dou, and Jinjun Chen. Special issue papers: SaC-FRAPP: a scalable and cost-effective framework for privacy preservation over big data on cloud. Concurrency and Computation: Practice and Experience, 25(18):2561–2576, December 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2018:TDM

[ZLQ^{+}18] Taotao Zhao, Xiangfeng Luo, Wei Qin, Subin Huang, and Shaorong Xie. Topic detection model in a single-domain corpus inspired by the human memory cognitive process. Concurrency and Computation: Practice and Experience, 30(19): e4642:1–e4642:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2016:DCA

[ZLT^{+}16] Shuoyi Zhao, Ruixuan Li, Wenlong Tian, Weijun Xiao, Xinhua Dong, Dongjie Liao, Samee U. Khan, and Keqin Li. Divide-and-conquer approach for solving singular value decomposition based on MapReduce. Concurrency and Computation: Practice and Experience, 28(2):331–350, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2018:DCN

[ZLW^{+}18] Gang Zhao, Shan Liu, Qi Wang, Tao Hu, Yawen Chen, Luyu Lin, and Dasheng Zhao. Deep convolutional neural network for drowsy student state detection. Concurrency and Computation: Practice and Experience, 30(23):e4457:1–e4457:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2019:RHF

[ZLW^{+}19] Yuefeng Zheng, Ying Li, Gang Wang, Yupeng Chen, Qian Xu, Jiahao Fan, and Xueting Cui. Retracted: A hybrid feature selection algorithm for microarray data. Concurrency and Computation: Practice and Experience, 31(12):e4716:1– e4716:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2013:SIPb

[ZLY^{+}13] Mingwei Zhang, Chengfei Liu, Jian Yu, Zhiliang Zhu, and Bin Zhang. Special issue papers: A correlation context-aware approach for composite service selection. Concurrency and Computation: Practice and Experience, 25(13):1909–1927, September 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2015:MCV

[ZLZ15] Xinyan Zhang, Keqiu Li, and Yong Zhang. Minimum-cost virtual machine migration strategy in datacenter. Concurrency and Computation: Practice and Experience, 27(17): 5177–5187, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2017:PAA

[ZLZ^{+}17] Jieyu Zhao, Jianxin Li, Baojian Zhou, Feng Chen, Paul Tomchik, and Wuyang Ju. Parallel algorithms for anomalous subgraph detection. Concurrency and Computation: Practice and Experience, 29(3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2019:HPM

[ZLZ^{+}19] Jing Zhao, Jianqi Liu, Zhenting Zhao, Miao Xin, and Yu Chen. A high-performance maintenance strategy for stochastic selective maintenance. Concurrency and Computation: Practice and Experience, 31(12):e4840:1–e4840:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2013:SLB

[ZM13] Anping Zhao and Yan Ma. Semantic link-based constructing of multi-dimensional service resource space for collaboration. Concurrency and Computation: Practice and Experience, 25 (11):1586–1603, August 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2010:VSH

[ZMJZ10] Long-Quan Zhou, Hong-Sheng Ma, Chang-Sheng Jiang, and Jun-Jie Zhou. The velocity structure of the hypocenter area of 2003 M 6.2 and M 6.1 Dayao earthquake. Concurrency and Computation: Practice and Experience, 22(12):1796–1802, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2018:PLF

[ZMYA18] Wei Zhang, Gang Mu, Gangui Yan, and Jun An. A power load forecast approach based on spatial-temporal clustering of load data. Concurrency and Computation: Practice and Experience, 30(16):e4386:1–e4386:??, December 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2011:SMA

[ZMZD11] Qi Zhang, Yi Mu, Minjie Zhang, and Robert H. Deng. Secure mobile agents with controlled resources. Concurrency and Computation: Practice and Experience, 23(12):1348–1366, August 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2016:PMR

[ZNT^{+}16] Yafang Zhang, Jianbing Ni, Xiaoling Tao, Yong Wang, and Yong Yu. Provable multiple replication data possession with full dynamics for secure cloud storage. Concurrency and Computation: Practice and Experience, 28(4):1161–1173, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2014:AOP

[ZO14] Bowen Zhang and Cornelis W. Oosterlee. Acceleration of option pricing technique on graphics processing units. Concurrency and Computation: Practice and Experience, 26(9): 1626–1639, June 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2006:SDG

[ZP06] L. Zhang and M. Parashar. Seine: a dynamic geometry-based shared-space interaction framework for parallel scientific applications. Concurrency and Computation: Practice and Experience, 18(15):1951–1973, December 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zarrelli:2007:PCM

[ZP07] R. Zarrelli and M. Petrone. Parallel computation on multilayer cluster grids. Concurrency and Computation: Practice and Experience, 19(13):1865–1878, September 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:EIS

[ZP19] Yan Zhang and Haitao Pu. Environmental indicators of sustainable computing applications for Smart City. Concurrency and Computation: Practice and Experience, 31(9):e4751:1– e4751:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2010:PAE

[ZPG10] Gu-Chun Zhang, Xiao-Hong Peng, and Xuan-Ye Gu. Performance analysis of an experimental wireless relay sensor network. Concurrency and Computation: Practice and Experience, 22(4):462–480, March 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2016:SBA

[ZQD16] Yue Zhang, Jing Qin, and Lihua Du. A secure biometric authentication based on PEKS. Concurrency and Computation: Practice and Experience, 28(4):1111–1123, March 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2017:RSH

[ZQD^{+}17] Haihui Zhao, Yaoguang Qi, Hongwei Du, Ningning Wang, Guofu Zhang, Wenbao Liu, and Hailong Lu. Running state of the high energy consuming equipment and energy saving countermeasure for Chinese petroleum industry in cloud computing. Concurrency and Computation: Practice and Experience, 29(14):??, July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2012:GEI

[ZQH12] Shujia Zhou, Judy Qiu, and Kenneth Hawick. Guest Editor’s introduction: Special section on challenges and solutions in multicore and many-core computing. Concurrency and Computation: Practice and Experience, 24(1):1–2, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2015:HSG

[ZQK15] F. Zhang, K. Qiu, and J. S. Kim. Hyper-star graphs: Some topological properties and an optimal neighbourhood broadcasting algorithm. Concurrency and Computation: Practice and Experience, 27(16):4186–4193, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2012:ERM

[ZQLZ12] Sihai Zhang, Hongyang Qiu, Yuan Liu, and Wuyang Zhou. Evolutionary reputation model for node selfishness resistance in opportunistic networks. Concurrency and Computation: Practice and Experience, 24(11):1261–1269, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhong:2019:MBF

[ZQMC19] Lunlong Zhong, Zhongjun Qu, and Félix Mora-Camino. A model-based flight qualities evaluation approach for civil aircraft. Concurrency and Computation: Practice and Experience, 31(10):e4723:1–e4723:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2019:UAB

[ZQR^{+}19] Yun Zhou, Lianyong Qi, Alexander Raake, Tao Xu, Marta Piekarska, and Xuyun Zhang. User attitudes and behaviors toward personalized control of privacy settings on smartphones. Concurrency and Computation: Practice and Experience, 31(22):e4884:1–e4884:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2017:LSA

[ZQW^{+}17] Naqin Zhou, Deyu Qi, Xinyang Wang, Zhishuo Zheng, and Weiwei Lin. A list scheduling algorithm for heterogeneous systems based on a critical node cost table and pessimistic cost table. Concurrency and Computation: Practice and Experience, 29(5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2016:EMC

[ZQZ^{+}16] Dongfang Zhao, Kan Qiao, Zhou Zhou, Tonglin Li, Xiaobing Zhou, and Ioan Raicu. Exploiting multi-cores for efficient interchange of large messages in distributed systems. Concurrency and Computation: Practice and Experience, 28(13): 3568–3585, September 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2019:FSI

[ZRB19] Xianghan Zheng, Chunming Rong, and Tuyatsetseg Badarch. Foreword to the special issue of green cloud computing: Methodology and practice. Concurrency and Computation: Practice and Experience, 31(23):e5425:1–e5425:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2001:HJA

[ZS01] Xiaolan Zhang and Margo Seltzer. HBench:Java: an application-specific benchmarking framework for Java Virtual Machines. Concurrency and Computation: Practice and Experience, 13(8–9):775–792, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/84503219/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=84503219&PLACEBO=IE.
pdf.

Zhang:2017:NEP

[ZS17] Yi Zhang and Jin Sun. Novel efficient particle swarm optimization algorithms for solving QoS-demanded bag-of-tasks scheduling problems with profit maximization on hybrid clouds. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2019:SKG

[ZS19] Hai Zhuge and Xiaoping Sun. Semantics, knowledge, and grids at the age of big data and AI. Concurrency and Computation: Practice and Experience, 31(3):e5066:1–e5066:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zechar:2010:CSE

[ZSL^{+}10] J. Douglas Zechar, Danijel Schorlemmer, Maria Liukis, John Yu, Fabian Euchner, Philip J. Maechling, and Thomas H. Jordan. The Collaboratory for the Study of Earthquake Predictability perspective on computational earthquake science. Concurrency and Computation: Practice and Experience, 22 (12):1836–1847, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2015:EMS

[ZSL^{+}15] Jinlin Zhang, Jun Shao, Yun Ling, Min Ji, Guiyi Wei, and Bishan Ying. Efficient multiple sources network coding signature in the standard model. Concurrency and Computation: Practice and Experience, 27(10):2616–2636, July 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2018:RAD

[ZSS18] Ying Zheng, Wangan Song, and Longfeng Shen. Research on application of data mining in fast character recognition based on big data. Concurrency and Computation: Practice and Experience, 30(24):e4925:1–e4925:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhan:2018:UTL

[ZSWS18] Jianzhou Zhan, Mei Sun, Huidan Wu, and Haojun Sun. Using triangles and latent factor cosine similarity prior to improve community detection in multi-relational social networks. Concurrency and Computation: Practice and Experience, 30(16):??, August 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4453.

Zhang:2014:EFP

[ZSZ^{+}14] Yu Zhang, Yue Sun, Jianzhong Zhang, Jingdong Xu, and Ying Wu. An efficient framework for parallel and continuous frequent item monitoring. Concurrency and Computation: Practice and Experience, 26(18):2856–2879, December 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2015:TDC

[ZSZ15] Li Zhang, Xiaoping Sun, and Hai Zhuge. Topic discovery of clusters from documents with geographical location. Concurrency and Computation: Practice and Experience, 27(15): 4015–4038, October 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2009:DPS

[ZT09] Ying-Nan Zhang and Hong-Fei Teng. Detecting particle swarm optimization. Concurrency and Computation: Practice and Experience, 21(4):449–473, March 25, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zou:2017:ASM

[ZTGW17] Yuxue Zou, Yueyang Tian, Songtao Guo, and Yan Wu. Active synchronization of multi-domain controllers in softwaredefined networks. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2012:RTH

[ZTM12] Richong Zhang, Thomas Tran, and Yongyi Mao. Real-time helpfulness prediction based on voter opinions. Concurrency and Computation: Practice and Experience, 24(17):2167– 2178, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2018:RTB

[ZTZ^{+}18] Jun Zhang, Lai Tu, Fan Zhang, Benxiong Huang, and Jin Liu. A real-time bus-subway transfer scheme recommendation systems. Concurrency and Computation: Practice and Experience, 30(20):e4640:1–e4640:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zeng:2009:EBI

[ZW09] Guosun Zeng and Wei Wang. An evidence-based iterative content trust algorithm for the credibility of online news. Concurrency and Computation: Practice and Experience, 21 (15):1857–1881, October 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2017:SSA

[ZW17] Zhipeng Zhao and Bin Wu. Scalable SDN architecture with distributed placement of controllers for WAN. Concurrency and Computation: Practice and Experience, 29(16):??, August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2006:VDG

[ZWF^{+}06] Yong Zhao, Michael Wilde, Ian Foster, Jens Voeckler, James Dobson, Eric Gilbert, Thomas Jordan, and Elizabeth Quigg. Virtual data Grid middleware services for data-intensive science. Concurrency and Computation: Practice and Experience, 18(6):595–608, May 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2017:NFA

[ZWH^{+}17] Ao Zhou, Shangguang Wang, Ching-Hsien Hsu, Myung Ho Kim, and Kok seng Wong. Network failure-aware redundant virtual machine placement in a cloud data center. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2013:SIPa

[ZWL^{+}13] Wanfeng Zhang, Lizhe Wang, Dingsheng Liu, Weijing Song, Yan Ma, Peng Liu, and Dan Chen. Special issue papers: Towards building a multi-datacenter infrastructure for massive remote sensing image processing. Concurrency and Computation: Practice and Experience, 25(12):1798–1812, August 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2015:TOS

[ZWL^{+}15] Jinghui Zhang, Mingjun Wang, Junzhou Luo, Fang Dong, and Junxue Zhang. Towards optimized scheduling for dataintensive scientific workflow in multiple datacenter environment. Concurrency and Computation: Practice and Experience, 27(18):5606–5622, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2017:OAP

[ZWL^{+}17] Huming Zhu, Yanfei Wu, Pei Li, Peng Zhang, Zhe Ji, and Maoguo Gong. An OpenCL-accelerated parallel immunodominance clone selection algorithm for feature selection. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuang:2019:OCM

[ZWL^{+}19] Yuan Zhuang, Xiaohui Wei, Hongliang Li, Yongfang Wang, and Xubin He. An optimal checkpointing model with online OCI adjustment for stream processing applications. Concurrency and Computation: Practice and Experience, 31(20): e5347:1–e5347:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2016:MED

[ZWLY16] Shunxiang Zhang, Yin Wang, Weidong Liu, and Xiaobo Yin. A model for estimating the out-degree of nodes in associated semantic network from semantic feature view. Concurrency and Computation: Practice and Experience, 28(15): 4177–4193, October 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ziermann:2012:DSO

[ZWMT12] Tobias Ziermann, Stefan Wildermann, Nina Mühleis, and Jürgen Teich. Distributed self-organizing bandwidth allocation for priority-based bus communication. Concurrency and Computation: Practice and Experience, 24(16):1903–1917, November 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2018:EET

[ZWT^{+}18] Yujian Zhang, Yun Wang, Xueyan Tang, Xin Yuan, and Yifan Xu. Energy-efficient task scheduling on heterogeneous computing systems by linear programming. Concurrency and Computation: Practice and Experience, 30(19):e4731:1– e4731:??, October 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2014:IMR

[ZWW14] Leqing Zhu, Dadong Wang, and Huiyan Wang. An improved method for the removal of ring artifacts in synchrotron radiation images by using GPGPU computing with compute unified device architecture. Concurrency and Computation: Practice and Experience, 26(18):2880–2892, December 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2018:LBT

[ZWW^{+}18] Hai-Tao Zheng, Zhe Wang, Wei Wang, Arun Kumar Sangaiah, Xi Xiao, and Congzhi Zhao. Learning-based topic detection using multiple features. Concurrency and Computation: Practice and Experience, 30(15):??, August 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4444.

Zhao:2019:EMD

[ZWXS19] Hongzhi Zhao, Yongchang Wang, Ke Xiong, and Lihua Song. An efficient Manhattan-distance-constrained disjoint paths algorithm for incomplete mesh network. Concurrency and Computation: Practice and Experience, 31(1):e4799:1– e4799:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:MAA

[ZWY^{+}19] Xiao Zhang, Faguo Wu, Wang Yao, Zhao Wang, and Wenhua Wang. Multi-authority attribute-based encryption scheme with constant-size ciphertexts and user revocation. Concurrency and Computation: Practice and Experience, 31(21): e4678:1–e4678:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2018:BWM

[ZWZ^{+}18] Fu Zhang, Yafei Wang, Limin Zheng, Tianle Ma, and Jiajia Wang. Biomimetic walking mechanisms: Kinematic parameters of goats walking on different slopes. Concurrency and Computation: Practice and Experience, 31(7):e4913:1– e4913:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2009:WCM

[ZX09] Yanchun Zhang and Guandong Xu. On Web communities mining and recommendation. Concurrency and Computation: Practice and Experience, 21(5):561–582, April 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2011:BOC

[ZX11] Hai Zhuge and Bei Xu. Basic operations, completeness and dynamicity of cyber physical socio semantic link network CPSocio-SLN. Concurrency and Computation: Practice and Experience, 23(9):924–939, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:PPB

[ZXC^{+}19] Jiangjiang Zhang, Fei Xue, Xingjuan Cai, Zhihua Cui, Yu Chang, Wensheng Zhang, and Wuzhao Li. Privacy protection based on many-objective optimization algorithm. Concurrency and Computation: Practice and Experience, 31(20): e5342:1–e5342:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:DPM

[ZXRZ19] Lefeng Zhang, Ping Xiong, Wei Ren, and Tianqing Zhu. A differentially private method for crowdsourcing data submission. Concurrency and Computation: Practice and Experience, 31(19):e5100:1–e5100:??, October 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zeng:2016:VEF

[ZXW16a] Lingfang Zeng, Shijie Xu, and Yang Wang. VMBackup: an efficient framework for online virtual machine image backup and recovery. Concurrency and Computation: Practice and Experience, 28(9):2630–2643, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2016:DGK

[ZXW16b] Wei Zhou, Yang Xu, and Guojun Wang. Decentralized group key management for hierarchical access control using multilinear forms. Concurrency and Computation: Practice and Experience, 28(3):631–645, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:CSP

[ZXW19] Weipeng Zhang, Ying Xu, and Chanti Wu. Comparative study of PLS path algorithm in water utilization rate of sponge city. Concurrency and Computation: Practice and Experience, 31(10):e4768:1–e4768:??, May 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2017:TDP

[ZXX17] Gaowei Zhang, Lingyu Xu, and Yunlan Xue. The time dependency predictive model on the basis of community detection and long-short term memory. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2006:IBS

[ZXXN06] Yanmin Zhu, Lijuan Xiao, Zhiwei Xu, and Lionel M. Ni. Incentive-based scheduling in Grid computing. Concurrency and Computation: Practice and Experience, 18(14):1729– 1746, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2012:EEL

[ZY12] Lingwen Zhang and Gang Yang. An energy-efficient localization algorithm for mobile vehicles in vehicle to vehicle network. Concurrency and Computation: Practice and Experience, 24(11):1213–1222, August 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhan:2019:SPS

[ZYFZ19] Dongyang Zhan, Lin Ye, Binxing Fang, and Hongli Zhang. SAVM: a practical secure external approach for automated in-VM management. Concurrency and Computation: Practice and Experience, 31(23):e4482:1–e4482:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2009:SIE

[ZYH09] Shujia Zhou, Yelena Yesha, and Milton Halem. Special issue: Exploring the frontiers of computing science and technology: Adapting emerging multi- and many-core processors. Concurrency and Computation: Practice and Experience, 21(17):

2141–2142, December 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2012:SIE

[ZYH12] Shujia Zhou, Yelena Yesha, and Milton Halem. Special issue: Exploring the frontiers of computing science and technology: efficiently utilizing multicore and many-core processors. Concurrency and Computation: Practice and Experience, 24(2): 109–110, February 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2008:MLT

[ZYL^{+}08] Hai Zhuge, Kehua Yuan, Jin Liu, Junsheng Zhang, and Xiaofeng Wang. Modeling language and tools for the semantic link network. Concurrency and Computation: Practice and Experience, 20(7):885–902, May 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2010:RBL

[ZYL10] Lang-Ping Zhang, Xiang-Chu Yin, and Nai-Gang Liang. Relationship between load/unload response ratio and damage variable and its application. Concurrency and Computation: Practice and Experience, 22(12):1534–1548, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2006:GGE

[ZYLT06] Haifang Zhou, Xujun Yang, Hengzhu Liu, and Yu Tang. GPGC: a Grid-enabled parallel algorithm of geometric correction for remote-sensing applications. Concurrency and Computation: Practice and Experience, 18(14):1775–1785, December 15, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2018:VMM

[ZYLY18] Zhou Zhou, Junyang Yu, Fangmin Li, and Fei Yang. Virtual machine migration algorithm for energy efficiency optimization in cloud computing. Concurrency and Computation: Practice and Experience, 31(7):e4942:1–e4942:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2007:AFG

[ZYN^{+}07] N. Zhang, L. Yao, A. Nenadic, J. Chin, C. Goble, A. Rector, D. Chadwick, S. Otenko, and Q. Shi. Achieving finegrained access control in virtual organizations. Concurrency and Computation: Practice and Experience, 19(9):1333–1352, June 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2016:EPI

[ZYW^{+}16] Ji-Lin Zhang, Jun-Feng Yuan, Jian Wan, Jie Mao, Li-Ting Zhu, Li Zhou, Cong-Feng Jiang, Peng Di, and Jue Wang. Efficient parallel implementation of incompressible pipe flow algorithm based on SIMPLE. Concurrency and Computation: Practice and Experience, 28(6):1751–1766, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:VDS

[ZYY^{+}19] Hong Zhang, Yifan Yang, Ding Yuan, Daniel Sun, Jun Zhang, Guoqiang Li, and Mingui Sun. Video denoising for security and privacy in fog computing. Concurrency and Computation: Practice and Experience, 31(22):e4763:1–e4763:??, November 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2006:JEJ

[ZYZ06] Bao-Yin Zhang, Guang-Wen Yang, and Wei-Min Zheng. Jcluster: an efficient Java parallel environment on a largescale heterogeneous cluster. Concurrency and Computation: Practice and Experience, 18(12):1541–1557, October 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2012:OQE

[ZYZ^{+}12] Yansong Zhang, Lisheng Yu, Xiao Zhang, Shan Wang, and Hui Li. Optimizing queries with expensive video predicates in cloud environment. Concurrency and Computation: Practice and Experience, 24(17):2102–2119, December 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2016:PPU

[ZYZ16] Yang Zhao, Liang Yao, and Yin Zhang. Purchase prediction using Tmall-specific features. Concurrency and Computation: Practice and Experience, 28(14):3879–3894, September 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2017:STS

[ZYZC17] Xinkui Zhao, Jianwei Yin, Chen Zhi, and Zuoning Chen. Sim-Mon: a toolkit for simulation of monitoring mechanisms in cloud computing environment. Concurrency and Computation: Practice and Experience, 29(1):??, January 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhuge:2011:ACS

[ZZ11] Hai Zhuge and Junsheng Zhang. Automatically constructing semantic link network on documents. Concurrency and Computation: Practice and Experience, 23(9):956–971, June 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2014:SSC

[ZZ14] Bo Zhang and Fangguo Zhang. Secure similarity coefficients computation for binary data and its extensions. Concurrency and Computation: Practice and Experience, 26(5):1023–1037, April 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2015:SED

[ZZ15] Jiang Zhang and Zhenfeng Zhang. Secure and efficient datasharing in clouds. Concurrency and Computation: Practice and Experience, 27(8):2125–2143, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2016:EWH

[ZZ16] Yunquan Zhang and Ji-Lin Zhang. Editorials: Workshop on high performance data intensive computing. Concurrency and Computation: Practice and Experience, 28(6):1695–1696, April 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2017:EFS

[ZZ17] Xiaobo Zhou and Laiping Zhao. Editorial: Foreword to the special issue on parallel and distributed computing with its applications. Concurrency and Computation: Practice and Experience, 29(17):??, September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2018:ICS

[ZZ18] Lei Zhang and Kaibo Zhang. An interactive control system for mobile robot based on cloud services. Concurrency and Computation: Practice and Experience, 31(7):e4983:1–e4983:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2015:EPT

[ZZC15] Minghui Zheng, Huihua Zhou, and Jing Chen. An efficient protocol for two-party explicit authenticated key agreement. Concurrency and Computation: Practice and Experience, 27 (12):2954–2963, August 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2017:SFS

[ZZC^{+}17] Yizhi Zhang, Zhengdong Zhu, Honglin Cui, Xiaoshe Dong, and Heng Chen. Small files storing and computing optimization in Hadoop parallel rendering. Concurrency and Computation: Practice and Experience, 29(20):??, October 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2017:SSP

[ZZD^{+}17] Hao Zheng, Zhengdong Zhu, Xiaoshe Dong, Baoke Chen, and Chengzhe Liu. Studying shadow page cache to improve isolated drivers’ performance. Concurrency and Computation: Practice and Experience, 29(10):??, May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2017:PGB

[ZZL^{+}17a] Haoyang Zhu, Peidong Zhu, Xiaoyong Li, Qiang Liu, and Peng Xun. Parallelization of group-based skyline computation for multi-core processors. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2017:PSP

[ZZL^{+}17b] Haoyang Zhu, Peidong Zhu, Xiaoyong Li, Qiang Liu, and Peng Xun. Parallelization of skyline probability computation over uncertain preferences. Concurrency and Computation: Practice and Experience, 29(18):??, September 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2018:CAC

[ZZL^{+}18] Biaokai Zhu, Jumin Zhao, Dengao Li, Hong Wang, Ruiqin Bai, Yanxia Li, and Hao Wu. Cloud access control authentication system using dynamic accelerometers data. Concurrency and Computation: Practice and Experience, 30(20):e4474:1– e4474:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2019:ICR

[ZZL^{+}19] Youwen Zhu, Yue Zhang, Xingxin Li, Hongyang Yan, and Jing Li. Improved collusion-resisting secure nearest neighbor query over encrypted data in cloud. Concurrency and Computation: Practice and Experience, 31(21):e4681:1–e4681:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2015:MRM

[ZZY^{+}15] Shuai Zhao, Yang Zhang, Le Yu, Bo Cheng, Yang Ji, and Junliang Chen. A multidimensional resource model for dynamic resource matching in Internet of Things. Concurrency and Computation: Practice and Experience, 27(8):1819–1843, June 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhanjun:2019:FDR

[ZZY^{+}19] Liu Zhanjun, Li Zhaoyi, Cheng Yuxia, Huang Xiaoge, and Chen Qianbin. A full-duplex relay selection strategy based on potential game in cognitive cooperative networks. Concurrency and Computation: Practice and Experience, 31(9): e4635:1–e4635:??, May 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2010:SFE

[ZZYW10] Yongxian Zhang, Xiaotao Zhang, Xiangchu Yin, and Yongjia Wu. Study on the forecast effects of PI method to the North and Southwest China. Concurrency and Computation: Practice and Experience, 22(12):1559–1568, August 25, 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2015:PML

[ZZZ^{+}15] Leqing Zhu, Yadong Zhou, Daxing Zhang, Dadong Wang, Huiyan Wang, and Xun Wang. Parallel multi-level 2D-DWT on CUDA GPUs and its application in ring artifact removal. Concurrency and Computation: Practice and Experience, 27 (17):5188–5202, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2019:RNN

[ZZZX19] Xiangwei Zheng, Yuang Zhang, Hui Zhang, and Qingshui Xue. An RBF neural network-based dynamic virtual network embedding algorithm. Concurrency and Computation: Practice and Experience, 31(23):e4516:1–e4516:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).